A Complete Bibliography of Publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences

halieutiques et aquatiques (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.11

Title word cross-reference

3 [2037]. ^{+} [490, 938]. ^{−} [2633]. ^{106} [2488]. ^{109} [1976]. ^{12} [2874]. ^{13} [2874, 1873, 1093, 2876, 14, 1122]. ^{134} [1731, 2487]. ^{137} [2488, 2284, 66, 1858, 2285, 2883, 2743]. ^{14} [2874]. ^{15} [2874, 2842, 1873, 1093, 2876, 14, 1122, 2203]. ^{18} [2842]. ^{18}O [1818]. ^{210} [17, 207, 66, 1398]. ^{226} [2787, 17, 1398]. ^{3} [2719]. ^{31} [1374]. ^{34} [1654, 1093]. ^{60} [1731, 2487]. _{1}^{3} [1098]. _{2} [1726, 397, 2395, 1164, 1165, 2635, 1391, 1390]. _{3} [1726, 560, 44]. _{4} [1726, 560, 261]. _{inf} [2185]. _{n} [1814]. a [1560, 1370, 2719, 1523, 1965, 826]. b [317, 281, 2894, 2641]. : [2874, 1390, 1192, 422, 1398]. δ [2842, 1654, 1873, 1093, 2876, 14, 1122, 2203, 1818]. K [1319]. L_{∞} [1319]. O [2586, 2418, 1610]. p [474, 1996, 652, 651, 1303]. × [1452, 2290, 1364, 1417, 1081, 406].

-ATPase [938]. -D [2037]. -deethylase [2586, 2418, 1610]. -dioxin [474, 1996, 1303]. -dioxins [652, 651]. -Fixation [1391, 1390].

0-Group [245, 1491]. 0-year [1802].

10-day [2865]. 100-kHz [2252]. 12- [819, 2719]. 13C [2213]. 18-month [819]. 1960s [762]. 1988 [663]. 1990s [2247, 2902]. 1992 [2013]. 1996 [2868, 2869]. 1997 [2258]. 1r* [1293].

2 [1754, 1475]. 2-Methylisoborneol [836, 61]. 20-yr [614, 1245]. 219-lake [1943]. 221 [1441]. 223 [1871]. 227 [1391, 1396, 1389, 1390, 2720, 1397, 1393, 1395]. 24S [2446]. 25th [1389].

3-trifluoromethyl-4-nitrophenol [2598].

4 [1532]. 4’-tetrachlorobiphenyl [1532]. 4WX [1162].

5 [2719]. 5’- [2719]. 5’-Hexachlorobiphenyl [2719]. 50-yr [1270, 1253]. 5992 [2452].

6- [819]. 6-Yr [1407].

7-ethoxyresorufin- [1610].

8-TCDD [1610]. 8-Tetrachlorodibenzo- [474, 1303, 1996]. 8-Tetrachlorodibenzo-p-dioxin [378].

âge [1234]. Îles [1026].

Aanalysis [1451]. Abalone [1087, 2822]. Abalones [1088]. Abdomen [2432]. Aberdeen [2861]. Aberdeenshire [2059]. Abilities [129, 1325, 342, 2782]. ability [1751, 1581, 2156, 2326]. Abiotic [1365, 449, 682, 1446, 11, 936, 2264, 2320, 1746]. Abitibi [2797]. Ablation [1358, 1012, 2298, 1618, 2854]. Abnormal [1605]. abnormalities [2025, 2828]. Aboriginal [1170]. above [1566]. Abramis [276, 918, 917, 2071, 945]. Absence [212]. absolute [1786]. Absorption [365, 2147]. Abundance [1491, 810, 556, 425, 342, 1258, 380, 1448, 101, 1145, 198, 613, 1097, 1391, 349, 592, 83, 1120, 1072, 1232, 287, 723, 1372, 1073, 848, 2237, 268, 1241, 1246, 443, 1264, 1172, 1956, 545, 114, 1053, 312, 936, 902, 103, 973, 41, 475, 201, 1670, 2413, 2701, 2306, 2307, 2242, 2046, 2011, 1501, 2305, 2896, 1487, 2013, 2221, 2528, 2174, 2069, 1583, 2696, 2025, 1622, 2200, 2345, 2829, 2178, 1862, 1800, 2139, 2205, 1790, 2197, 1519, 2131, 2843, 1626].

Abundances [112, 1573, 1487]. Acadian [527]. acanthias [525]. Acanthocephala [799, 1107]. Acanthocheilonema [2166]. Acanthodiaptomus [1021]. accelerated [2351]. Access [1251]. acclimated [2791, 1705]. Acclimation [885, 1139, 510, 1761, 420, 513, 514, 808, 512, 1217, 1218, 1648, 2020, 1865]. accordance [2754]. account [2603, 2657, 2651]. Accounting [1612, 1953]. Accumulation [2342, 2173, 995, 1144, 561, 2720, 1742, 875, 1154, 2082, 2704, 2165, 2487, 2194, 2193, 2621, 2876, 1863, 2496, 1595, 2009]. Accuracy [1242, 2202, 869, 2351, 2298]. Acetone [826]. Acetylcholinesterase [188]. achieve [2650]. Acid [937, 885, 1290, 411, 1004, 271, 785, 1316, 170, 580, 515, 1023, 1296, 1350, 834, 306, 1056, 186, 1282, 184, 605, 36, 615, 1067, 377, 511, 327, 893, 1925, 1509, 1590, 2677, 2173, 2797, 2879, 2525, 2800, 2864, 2802, 2329, 1855, 2402, 2375, 1806, 2226, 2880, 1955, 1730, 1865, 1916]. acid-base [1855]. Acid-Exposed [515]. Acid-Sensitive [1290, 1316, 834]. Acid-Tolerant [834]. Acidic [55, 395, 1324, 48, 1155, 1017, 942, 125, 7, 350, 72, 272, 6, 95, 45, 843, 185, 270, 269, 1166, 2722, 1650, 2678, 2225]. Acidification [179, 510, 989, 1036, 585, 394, 576, 847, 566, 567, 1229, 1003, 131, 120, 11, 577, 89, 967, 1276, 2138, 1970, 2874, 2141, 1780, 1844, 1963, 1945, 2602, 2292, 1727, 1728]. Acidified [157, 47, 26, 973, 1414, 2468, 1885, 1729, 2246]. Acidify [1316]. Acidifying [1316]. Acidity [664, 492, 880, 151, 181, 182, 2636]. Acids [157, 1131, 978, 324, 1301, 2466, 2801, 1564, 2453]. Acipenser [1830, 611, 2377, 2797, 1900, 787, 719, 939, 2832, 2244, 1131, 1532, 2158, 2224, 100, 1288, 2854]. Acoustic [810, 1174, 573, 380, 320, 665, 147, 284, 965, 902, 483, 93, 216, 1983, 1934, 2040]. acoustics [2660, 1808]. Acquired [2448, 1655]. acquisition [1516]. acrolepis [2788]. across [2859, 2727, 1180, 635, 1252, 2014, 2320, 1550, 1551]. act [1678]. ACTH [2430]. Actiononaias [2129]. Activation [1139, 181]. Active [1163]. Actively [853]. Activities [407, 328, 553, 1221, 2381, 2020]. Activity [905, 853, 170, 697, 861, 382, 938, 188, 1363, 438, 743, 399, 913, 1414, 1298, 2380, 2095, 2766, 1997, 1642, 2168, 2706, 1743, 2230, 2284, 2598, 2103, 2856, 1733, 2586, 2418, 1858, 2285, 2035, 1608, 1636, 1960, 2379, 1610]. Activity-Specific [697]. actual [1582]. Acuarioidea [630]. Acuity [948]. aculeatus [1607, 1922, 2365, 139]. acuminata [2446]. Acute [2129, 442, 1361, 2741, 1191, 1200, 337, 817, 1860, 2801, 1739, 1912]. acutely [1704]. Adams [97, 341]. Adaptability [396, 691, 582, 622]. adaptation [1778, 2094]. adaptations [2855]. Adapted [1409]. Adaptive [427, 1375, 1589, 1822, 2681]. added [1731, 2487, 1976]. Addition [395, 2576, 2225, 1974]. Additional [220, 1006]. Additions [29, 1726, 2876, 1975, 1897, 1717, 1917]. Additive [722, 2648, 2362]. Address [1450]. Adductor [407, 1001]. adducts [2469]. Adhesion [167]. Adipocytes [1812]. Adirondack [639, 2225, 585, 1316, 207, 880, 1155, 1837, 782, 2875, 967, 151, 1239, 2270, 1166]. Adjacent [1384, 133, 1469, 2906]. Adjusting [2161, 1126, 1234].

Adjustment [276, 917, 1948]. Adjustments [1139]. Administration [88]. adolescent [2102]. Adrenergic [327]. Adriatic [2322]. Adsorption [490]. adspersus [1678]. Adult [1444, 389, 932, 393, 370, 869, 221, 269, 1303, 1278, 1178, 2678, 2421, 2186, 2871, 2085, 2442, 1481, 2700, 1630, 1583, 2099, 2646, 2829, 2223, 1894, 2102, 2280, 2212, 2480]. adulthood [2865]. Adults [609, 2555]. Advance [1355]. Advantage [204]. Advection [1026, 1291, 2748]. Advections [814]. Advective [687]. Advice [526, 1433, 2476]. aeglefinus [573, 598, 40, 2011, 2662, 1579, 627, 2895, 2383, 849]. Aerial [1528]. Aerobic [1218, 2134, 1633]. Aeromonas [2592]. aeruginosa [2850]. aetiopicus [1288]. Affect [879, 559, 2737, 2269, 2581, 2441, 1952]. Affected [967, 2847, 1780, 2030, 1938]. Affecting [639, 391, 809, 758, 2597, 1807, 1957, 2388, 2113, 2609, 2378, 1825, 2287, 2009, 2695]. Affects [755]. affinis [1615, 51]. Affinities [766, 2085]. African [117]. After [44, 1387, 1388, 1380, 935, 28, 343, 407, 1536, 1753, 48, 674, 1597, 2791, 47, 2442, 2026, 254, 253, 2274, 2700, 2022, 794, 765, 88, 1989, 1239, 1721, 893]. Ag [2093, 2833]. against [1655, 2677, 2473, 2699, 1800]. agardhii [467]. Age [2788, 958, 670, 732, 473, 12, 279, 32, 547, 813, 1399, 465, 2707, 903, 742, 1134, 896, 2382, 191, 263, 393, 175, 928, 1251, 1583, 642, 1047, 1335, 1129, 30, 351, 258, 64, 551, 1227, 244, 650, 755, 1521, 150, 1005, 1427, 446, 961, 1206, 1002, 1398, 784, 2787, 1627, 2257, 2366, 2373, 2687, 2830, 1611, 1741, 2615, 1546, 2601, 1810, 2814, 2630, 1234, 2519, 1740, 1777, 1695, 1786, 1518, 2367, 1837, 1615, 1490, 2849, 2524, 2107, 1603, 2400, 1962, 2070, 2169, 2759, 2255, 2483, 2749, 1894, 1872, 1712, 2895, 2513, 2512, 2197, 1828, 2091, 1703, 1737]. age [1503]. Age- [1227, 1005]. Age-0 [784, 2366, 2687, 2630, 2849, 2400, 2169, 2749, 1872, 1703]. age-1 [1490]. age-at-recruitment [1627, 2257]. age-classes [1810]. age-O [1741, 1490]. Age-Selective [958]. Age-Specific [813, 1583]. Age-Structured [465, 263, 1251, 2830, 1611, 2601, 1894]. Ageing [344, 770, 2788, 1521]. agent [2194]. Agents [831]. Ages [392, 1733, 1964]. agestructured [2779]. Aggregating [949]. Aggregation [573, 680, 871, 2299, 2056, 2772]. Aggregations [58, 320, 755, 2318, 2325, 1554]. Aggregative [1289]. Aggressiveness [105]. Agonistic [941, 65, 242, 1987]. Agreement [2125]. agreements [2541]. Agricultural [39, 2618, 2227]. aguabonita [2812]. Air [699, 1343, 1344]. airborne [1587]. al [2916, 2913, 2825, 2911, 879]. alalunga [2601]. alanine [2137]. alarm [2737]. alascanus [1845]. Alaska [193, 1465, 392, 1983, 2021, 685, 899, 957, 1250, 2199, 2539, 2088, 287, 228, 793, 1861, 860, 14, 1122, 2536, 1457, 345, 2274, 2700, 2022, 2222, 614, 606, 2025, 2245, 2726, 2759, 2829, 2767, 949, 1463, 1459, 2214, 2537, 2815, 1600, 2201]. Alaskan [241, 2177, 2614, 465, 2715, 2889, 1124, 882, 2638]. albacares [1234]. albacore [2601]. Alberni [2074]. Alberta [2309, 2308, 1577, 2306, 2307, 2305, 2302, 2304]. albida [586, 587]. albula [969, 2081, 1756]. Alevin [1017]. Alevins [937, 182]. Alewife [766, 794, 538, 1517, 1946, 1572]. Algae [1230, 159, 592, 125, 46, 584, 230, 554, 1898, 1520, 2665, 2068, 2287, 1730, 1729]. Algal [501, 90, 1892, 753, 1443, 847, 904, 116, 993, 781, 858, 2106, 1582, 1523, 2505, 2665, 2408, 1728]. Algonquin [1036]. Algorithm [1429, 2107]. algorithms [1694]. alien [1764]. Aliphatic [1774]. Alkaline [1296, 1761]. Alkalinity [1349, 1440, 1018, 2296, 1742, 1945]. All-Female [1362]. all-male [2115]. Allantoic [36]. Allele [111, 2228]. Allocation [901, 1112, 2461, 2460, 1792, 2467, 1904, 1719]. Allochthonous [634, 1950]. Allometric [365, 1499, 2323, 2608]. allométrique [575]. Allometry [2279, 1624, 63, 1011, 1267, 2432, 1773, 2217]. Allopatric [692, 740, 1498]. Allopodocotyle [1667]. allow [2534]. Allozyme [586, 587, 126, 1466, 766, 1310, 1460, 1671, 1827, 1497, 2559]. Allozymes [1465]. Allozymic [1248]. alone [2620, 1568]. along [783, 1167, 342, 1562, 2469, 645, 35, 1875, 1553, 1513, 531, 2299, 2778, 290, 2378, 2574, 823, 563, 2731, 2007]. Alosa [1641, 772, 766, 1517, 2770, 1802, 601, 794, 1572, 478, 538, 178, 2326, 2097]. Alpine [1742, 1863, 1403, 2636, 1765, 2823, 2794]. alpinus [1170, 2034, 2380, 597, 392, 1498, 2514, 2761, 1680, 2882, 1356, 445, 1144, 1476, 2385, 2791, 1742, 1863, 1285, 622, 2846, 1966, 621, 1151, 1301, 406]. Alps [2623]. Altantic [1955]. Alter [541, 2719, 1939, 2028]. Alteration [1237, 481, 811, 1881, 2297, 1880, 1879, 1878]. Alterations [1456, 1166]. altered [2410]. Alternate [751]. Alternative [1431, 68, 370, 1126, 267, 725, 2721, 2267]. Alters [817]. altivelis [1845]. aluminium [1969, 2192, 2678, 1636]. Aluminum [510, 785, 23, 24, 1296, 834, 186, 180, 183, 584, 704, 513, 184, 514, 185, 176, 512, 1067, 511, 163, 1217, 1218, 327, 181, 182, 2722, 2138, 1561, 1747, 1516, 1865]. amago [2121]. ambient [1772]. ambiguity [2785]. Amblema [2129]. Ameiurus [2236]. America [1115, 656, 470, 165, 655, 1070, 533, 2282, 1513, 2133, 2557, 1563, 749, 440, 1068, 1311, 1462, 1943, 2062, 2489]. American [2714, 2443, 1854, 379, 911, 1670, 2014, 1773, 1212, 302, 1098, 1341, 1656, 2281, 1248, 1641, 868, 2265, 1495, 772, 1211, 1292, 2621, 2632, 1192, 2770, 1802, 1691, 601, 1512, 2251, 1406, 478, 2733, 765, 178, 588, 734, 505, 1464, 1779, 1257, 850, 1790, 2197, 2642, 1735, 267, 693, 2892, 1949, 1599, 812, 2326, 2097]. americana [2621, 816, 358, 630, 1083, 351, 202]. americanus [1384, 909, 2142, 1773, 58, 137, 520, 2404, 302, 1098, 1341, 24, 868, 2265, 1292, 2632, 1192, 1254, 1774, 734, 505, 494, 811, 1960, 1298, 267, 693, 2892]. Amia [352]. Amino [580]. Amisk [2309, 2308, 2303, 2302, 2301, 2304]. Ammocoetes [201, 1508]. Ammodytes [1076]. ammonia [1713, 2444, 2540]. Ammonium [1004, 2579, 422, 2837]. Amnicola [194]. Among [1466, 113, 1457, 2033, 316, 60, 1472, 1632, 1826, 126, 1655, 635, 797, 2484, 2177, 2706, 1448, 124, 1032, 256, 1641, 1810, 2393, 791, 2518, 1127, 2265, 450, 2085, 1720, 2185, 2420, 1280, 2821, 2824, 2490, 1099, 2069, 1563, 1015, 1047, 478, 1083, 881, 869, 2646, 2829, 1025, 2178, 2239, 2503, 2846, 1966, 780, 2191, 1222, 2574, 277, 1284, 971]. Among-lake [2033]. amont [2884]. amount [1911]. Amphibian [1034]. Amphipod [1365, 1074, 576, 1713, 2723]. Amphipoda [829, 1991, 1145, 2008, 1107, 2794]. amphipods [1674]. amplification [1869]. Amplified [1099, 2265]. amplitudes [2095]. Amundsen [2013]. anachoronism [1973]. anadromes [534]. Anadromous [425, 60, 347, 1356, 491, 393, 33, 34, 1124, 178, 1171, 1451, 1473, 621, 1823, 2514, 2873, 534, 2429, 2229, 2640, 2195]. Anadromy [1285]. Anaerobic [2508, 1509]. Anal [356]. analogue [2286, 2802]. Analyses [379, 584, 178, 594, 441, 2140, 2777, 1786, 1722, 2603, 2657, 2559]. Analysing [1518]. Analysis [317, 60, 1121, 344, 892, 929, 986, 315, 524, 771, 944, 1145, 291, 37, 999, 995, 1334, 349, 672, 435, 1022, 1339, 1295, 855, 2715, 1292, 596, 1423, 1255, 350, 1169, 1197, 263, 727, 795, 284, 59, 552, 532, 716, 928, 642, 606, 568, 687, 1078, 1, 258, 965, 109, 770, 1425, 1137, 3, 620, 1077, 923, 70, 993, 21, 722, 1305, 1455, 1308, 1368, 1883, 2000, 1256, 111, 234, 1267, 181, 2201, 1531, 2564, 2499, 1958, 2514, 1776, 1671, 1827, 1813, 2264, 1758, 2842, 1578, 2150, 1561, 2276, 1978, 1993, 2873, 1829, 1629, 2450, 1641, 2615]. analysis [1803, 1545, 1979, 2641, 2691, 1843, 1596, 2243, 2382, 1763, 1547, 1397, 2288, 1802, 2724, 1662, 1934, 2650, 1796, 2250, 1586, 2051, 2431, 1693, 1603, 2424, 2016, 2681, 2804, 2558, 2679, 2255, 2483, 2907, 2398, 1914, 1483, 2751, 2408, 2492, 1891, 1949, 2343, 2885, 1575, 2372, 1494]. analytic [2334]. Analytical [2437, 483]. Analyzing [1304, 2003]. Anarhichas [2683]. Anatomical [1807]. ancestral [2085]. ancestry [1751, 1987]. Anchoveta [765]. anchovy [2506]. Andrews [2908]. Androgenetic [1569]. anemones [2411]. anglers [2751]. angling [1509, 1990]. Anguilla [1212, 2281, 1611, 1211, 2367, 1581, 2733, 2625, 1215]. anguillicaudatus [1557]. angularity [2741]. animal [2592, 1622, 2017]. animals [2874, 2320, 2534]. Anisakidae [486, 951, 1673, 609, 2069]. Anisakinae [1674]. anisakine [1670]. Anisakis [862, 1673]. Annelida [1321]. Anniversary [1389]. Annual [692, 249, 907, 622, 2648, 2494, 1846, 2104, 633]. annually [1818]. annuelle [633]. annulus [1521]. Anodonta [2082]. anomalies [1626]. Anomalous [2429, 1932]. Anoplopoma [175, 653]. Anotopterus [555]. Anoxia [1111, 2445]. Antarctic [2472, 355]. antarcticus [1962, 2355]. Antennae [413]. anthonyi [146]. anthracene [2719]. Anthropogenic [120, 11, 1745]. Antifreeze [172, 631, 1444, 2855, 1848]. antillarum [1627]. Antioxidant [881]. antipodarum [2162]. antipredator [2737, 1939, 2028]. antipredatory [2054]. Antisapstain [1342]. Aphanizomenon [1392]. API [992]. Aplodinotus [591, 502]. Appalachian [1875]. apparent [1997, 2401, 2592]. appetite [2444]. applicability [2631]. Application [1701, 2787, 318, 2689, 1110, 229, 1030, 1546, 901, 27, 772, 557, 1423, 1615, 727, 1062, 833, 2793, 1312, 869, 1223, 178, 1198, 2398, 164, 413, 982, 495, 883, 1749, 2352, 2150, 2746, 2601, 1234, 2170, 1934, 2108, 1693, 2627, 2652, 2748, 2611, 7]. Applications [1090, 1138, 1137, 1930, 1640, 1869]. applied [2488, 2440]. Appraisal [1105, 1981]. appraising [2631]. Approach [156, 1304, 465, 952, 857, 827, 875, 1429, 1375, 997, 444, 3, 972, 1136, 2516, 2807, 2391, 2660, 1887, 2314, 1877, 2152, 2724, 1520, 2108, 2649, 2650, 2387, 2151, 2334, 1964, 2541, 2004, 1914, 1505]. Approaches [701, 1346, 263, 517, 956]. Appropriate [1266]. April [2905]. Aquaculture [299, 1374, 2542, 2067]. aquarium [1663]. Aquatic [553, 739, 807, 697, 450, 1159, 203, 879, 95, 997, 62, 867, 2, 1011, 644, 1166, 1775, 2342, 2612, 2213, 2872, 2594, 2517, 2376, 2850, 1985, 1928, 1752, 2793, 1702, 2560, 2241, 2457, 2775, 2776]. Aqueous [644]. Archipelago [643]. Architecture [542]. archival [2786]. Arctic [1206, 2034, 2380, 597, 2499, 392, 2333, 174, 1498, 2514, 2761, 685, 1680, 779, 2106, 241, 2882, 1356, 445, 1144, 1476, 1596, 2647, 2385, 1535, 2266, 196, 228, 792, 793, 1378, 2889, 1185, 2791, 829, 282, 1742, 1863, 1170, 1285, 2561, 338, 2866, 2670, 2538, 2697, 622, 2846, 1966, 154, 1328, 621, 1762, 1206, 1835, 2187, 1151, 1301, 406]. arcticus [685, 228, 792, 793, 1378, 2889, 282]. Arctocephalus [2472]. Area [1732, 1437, 123, 790, 1321, 1350, 468, 1394, 1397, 1179, 864, 8, 1976, 1580, 2649, 2131, 1798, 1789, 410, 90, 1441, 1389, 1434, 1392, 2703, 1973, 1931, 1725]. area-swept [2649]. Area-Under-the-Curve [790]. Areas [360, 789, 1345, 1143, 891, 1724]. arenaria [189, 428]. argentatus [2691]. Argentina [1683, 107]. Arginine [306, 1056]. argus [2122, 2231]. Argyle [137]. arid [135]. Arkansas [2604]. Armorhead [833]. aromatic [2865, 2469]. Arrival [1899]. ARROW [1245]. Arsenate [247, 298]. Arsenite [1200]. artedi [2309, 2308, 2502, 780]. artedii [60, 346, 34, 177]. Artemia [500]. Arterial [668]. Artificial [104, 91, 128, 1246, 902, 2631, 1638, 1824, 2122, 1752, 2826]. artificially [2720]. Ascaridoid [676]. Ascaridoidea [1673, 862, 1671, 1827, 1813, 77, 799, 630, 1675, 466, 870, 255]. Ascendency [387]. Ascorbate [170, 500]. Ascorbic [170, 2375, 2677]. Asia [1462, 1460]. Asian [2543, 1589]. Asiatic [1524]. Aspect [483, 2637]. Aspects [500, 168, 100]. Assay [712, 2446]. assays [2298]. Assemblage [375, 80, 1405, 2549, 2778, 2682]. Assemblages [701, 343, 156, 735, 908, 239, 90, 1070, 196, 238, 2557, 1345, 577, 426, 863, 1276, 218, 1326, 1028, 1395, 2859, 2727, 2564, 2636, 2264, 2497, 1907, 2604, 1483, 1825, 2419, 2246, 1832, 1943]. Assess [672, 1229, 538]. Assessing [2564, 1244, 2803, 908, 2701, 2673, 703, 7, 350, 584, 6, 2602, 95, 526, 1943, 1561, 2796, 2568]. Assessment [1226, 237, 1110, 712, 757, 1242, 697, 1030, 259, 387, 1996, 963, 1232, 1421, 778, 1429, 1187, 1367, 1191, 577, 1433, 1077, 935, 388, 2648, 1823, 2364, 2048, 2612, 2830, 1814, 2423, 2232, 2601, 1617, 1618, 1780, 2295, 2476, 2234, 1822, 1683, 2107, 2914, 2663, 1895, 2852, 1851, 1906, 1820, 2438, 2717, 2786]. Assessments [958, 752, 1138, 1775, 2023, 2762]. assigning [1964]. assignment [2459]. Assimilation [313, 1112]. Associated [488, 404, 1159, 358, 546, 805, 1480, 497, 2753, 1489, 1933, 1815, 1821, 2047, 2050, 2378, 2205, 1554, 2211]. Association [1837, 417, 869, 1293, 1679, 2121]. Associations [814, 403, 1223, 11, 1206, 2204, 2030, 2642, 1728]. Assumption [1125]. assumptions [1749]. assurance [2498]. Astaxanthin [330]. Asterionella [877]. asymptotes [2425]. Atherinidae [2503]. Atlantic [2215, 2682, 1969, 900, 486, 1106, 1020, 2027, 1834, 2760, 2057, 1504, 1352, 396, 1770, 379, 2192, 2268, 132, 347, 1140, 1997, 1263, 1670, 1509, 797, 598, 862, 1813, 1664, 1644, 2373, 1990, 2678, 271, 2912, 2784, 1358, 1748, 2077, 2844, 281, 732, 315, 814, 172, 1986, 2412, 1840, 1788, 1918, 2421, 2861, 499, 2546, 473, 773, 94, 470, 245, 1006, 1058, 1741, 1075, 2031, 1993, 2873, 296, 819, 2680, 698, 2466, 2839, 560, 2615, 1038, 2332, 922, 1810, 1173, 2504, 1452, 1981, 1234, 329, 1667, 631, 1444, 2855, 1194, 2219, 1787, 2518, 2519, 2230, 1839, 1614, 373]. Atlantic [1747, 1529, 1777, 2054, 2785, 1292, 1134, 1133, 1376, 2816, 2669, 1333, 273, 2294, 79, 798, 2422, 191, 776, 1471, 2740, 1499, 2279, 2207, 1681, 2525, 138, 548, 1351, 1846, 1547, 2834, 1229, 2864, 2041, 1844, 1317, 2327, 1073, 2662, 2913, 1249, 2572, 1606, 430, 330, 2676, 531, 2478, 954, 2170, 272, 1692, 2602, 1821, 2683, 1691, 2820, 2509, 786, 2553, 654, 1172, 2530, 1586, 2058, 1696, 1187, 188, 759, 1363, 2137, 1044, 1690, 1548, 927, 1047, 2400, 478, 2750, 1497, 2016, 968, 2866, 2100, 2099, 2670, 1059, 2350, 1078, 2144, 1223, 535, 2863, 1221, 1293]. Atlantic [2402, 2891, 275, 374, 594, 1101, 326, 1848, 2358, 270, 668, 2503, 2565, 815, 2076, 1141, 529, 2617, 2436, 891, 1067, 1890, 2858, 523, 1130, 121, 1162, 2593, 86, 1005, 946, 2642, 1735, 2591, 2529, 805, 2883, 2743, 377, 1688, 269, 2899, 856, 2271, 372, 2059, 1379, 2668, 2667, 2809, 581, 2128, 163, 333, 50, 1278, 2372, 1178, 1084, 1812, 849, 1161]. atlanticus [672, 87, 1521, 1048]. atlantique [575]. atmosphere [2635]. Atmospheric [1290, 72, 323, 1114, 2578]. atmospherically [2441]. ATPase [938]. atratulus [967, 985]. Atresia [548]. attachment [1492]. Attacks [555]. attempts [1967]. Attenuation [2789, 1652]. Attraction [384]. Atyid [1306]. Atyidae [1306]. atypical [2807]. audax [1339]. Aufwuchs [808]. Auke [2088, 949]. Auklet [1076]. aurata [1739]. auratus [1482, 970, 1496, 2781]. Aurelia [1922, 2365]. aurita [1922, 2365, 369]. auritus [2247]. Australia [1815, 2551, 2060]. Australian [1271, 401, 1049]. australis [1683]. Author [2909]. Autochthonous [634]. Autocorrelated [574, 57, 2084]. Autocorrelation [1232, 2825, 2603, 2657]. Automated [1974]. Autotrophic [117, 2033]. autres [926]. Autumn [86, 1485]. autumnalis [241]. aux [1234]. auxiliary [1786]. Availability [1438, 648, 1007, 985, 190, 2620, 2210]. Available [1140, 91]. aval [2884]. Average [1415]. avoid [1751]. Avoidance [1969, 1238, 861, 890, 2465, 1556, 2690, 1541, 2782, 2156, 1721, 1706]. Axenic [580]. axial [2168]. Axis [1139]. ayresi [1544]. azteca [1713, 398, 1567, 1566, 2194, 2193, 2803, 576, 2008]. aztecus [1238].

B [1650, 2148, 633]. B2 [1475]. B2-2 [1475]. BACI [935]. Bacillus [1193, 1265, 913, 1307]. Back [246, 684, 729, 1695, 2651]. Back-Calculation [684, 729, 1695, 2651]. Back-Calculations [246]. background [1567, 1566]. Backscattering [320]. Backwaters [591]. Bacteria [831, 992]. Bacterial [905, 223, 206, 837, 2160, 1265, 117, 1386, 1141, 2408]. Bacterioplankton [382, 209]. bacterivory [2160]. Baculovirus [1163]. baie [2869, 1026, 2868]. bairdi [1035, 1289]. Bait [128]. Baited [952, 363]. Balaena [2647]. Balance [895, 2079, 2625, 1982]. balances [1647, 2369]. baleen [2647]. Ballast [1095]. Baltic [691, 416, 2469, 578, 1487, 1999, 2740, 2834, 2718, 1785, 643]. bands [2355, 1818]. Bank [2861, 2237, 1526, 581, 40, 626, 2662, 432, 928, 627, 1227, 2383, 121, 748]. bankfull [2871]. Banks [1277, 2321]. barbatus [123, 2174]. Barents [531]. Barkley [2018, 1014]. barrage [2884]. barren [2378]. Barrier [2060]. Barrow [159, 27, 8]. Base [395, 1169, 377, 1509, 2225, 1855, 1917]. Based [246, 362, 12, 32, 1399, 596, 451, 875, 526, 433, 878, 684, 729, 1422, 190, 1531, 2736, 690, 1748, 2077, 1590, 2412, 1840, 2150, 2853, 1629, 2450, 790, 2601, 1617, 855, 2888, 1764, 2420, 1220, 2645, 2314, 1864, 848, 2613, 330, 2649, 1902, 2554, 1586, 2424, 2627, 2811, 1649, 852, 2446, 426, 956, 1894, 1585, 2870, 2606, 2076, 2651, 2408, 2354, 2892, 2167, 2786, 1600]. Baseline [1460]. Basin [710, 1228, 539, 1588, 443, 1507, 1849, 2835, 1829, 2477, 2749, 2105, 325, 1297, 617, 453, 1702, 1401]. Basins [320, 939, 1168, 2303, 2881, 2838, 2627]. Basis [1399, 1249, 535, 1310, 2075, 2268, 1773, 1999, 1966]. Bass [1012, 476, 84, 704, 262, 178, 383, 729, 767, 873, 1096, 943, 1314, 1313, 2687, 2328, 2630, 1640, 2656, 2500, 1957, 2646, 2790, 2440, 2291, 1853, 1575, 2372]. bassin [325]. Batch [1004, 2286]. batch-spawning [2286]. Bathymetric [1845, 1005]. bathymetry [2702]. Bathypelagic [555]. baudiniana [1669]. Bay [2039, 1384, 2264, 1228, 2413, 2763, 162, 161, 249, 1250, 2460, 1597, 1408, 1936, 1815, 295, 1469, 2277, 2587, 1122, 2536, 2079, 1420, 1622, 1585, 290, 1112, 211, 949, 1178, 1600, 2201, 1291, 2199, 2256, 939, 1640, 2283, 816, 2024, 33, 2174, 1420, 884, 2868, 2869, 64, 312, 1915, 2671, 898, 1575, 240]. Bayes [1233]. Bayesian [2807, 2321, 1346, 1822, 2718, 2288, 1429, 2108, 2107, 2650, 2762, 2663, 1964, 2681, 2398, 2123]. Bays [1420]. Be [1376, 1088, 2620, 1568, 2852, 2806]. beach [2264, 2539]. beach-spawning [2539]. Beaked [356, 1968]. beaks [2089]. bear [1762]. Bearded [123, 2174]. Bearing [444]. Beaufort [241, 645, 2013, 974]. Beaver [1111, 2511, 1608]. become [1859]. Bed [93, 1893, 2350]. Bedford [443]. Beds [1448, 610, 681, 844, 2219, 1815, 2378]. Beef [1940]. Before [44, 1380, 935, 1536, 1753, 1597, 430, 794]. Before-After-Control-Impact [935]. Begin [1316]. Behavior [499, 1329, 390, 254, 199, 284, 472, 623, 720, 589, 949, 65, 2443, 1987, 2737, 1637, 1680, 2349, 1522, 2028, 2071, 2400, 2506, 1534, 2102, 1736, 2454]. Behavioral [219, 2688, 1156, 2339]. Behaviour [220, 396, 454, 662, 172, 105, 890, 941, 358, 733, 622, 744, 1141, 242, 1014, 1969, 2248, 2698, 2705, 2054, 2669, 1939, 2359, 1696, 1584]. Behavioural [2722, 2169, 1935]. behaviours [2761]. Belgian [2457]. bellum [987]. Beluga [625, 2013]. Belugas [1120, 1328]. Benedenia [1663]. Bénéfices [534]. Benefits [2521, 534]. Benthic [1152, 757, 889, 1350, 382, 1150, 1193, 2493, 532, 230, 679, 2045, 519, 503, 80, 1265, 151, 1276, 913, 581, 1751, 2269, 2377, 2087, 1700, 2392, 2413, 1501, 2176, 633, 1693, 2629, 2261, 1616, 1571, 1808, 2492, 1832, 2187, 2379, 2537, 2220, 1703]. benthiques [633]. Benthivorous [91]. Benthos [158, 224, 222, 541, 777, 2175, 2345]. Benzo [96, 2791]. Bering [1086, 606, 150, 2108, 1534, 722, 1519]. Bermuda [360, 1177]. Bernard [2439]. Berried [58]. Bertalanffy [160, 524, 708, 1331, 2757, 2293, 1612, 2425, 1319]. beta [96, 420]. Better [1236]. Between [1391, 1454, 476, 1471, 1478, 1380, 235, 544, 701, 1648, 1870, 1167, 2118, 1445, 2027, 1834, 379, 1776, 656, 853, 398, 692, 1360, 2494, 2395, 1370, 2392, 2281, 2502, 1145, 707, 872, 130, 2614, 2190, 527, 753, 861, 279, 582, 2539, 1810, 2714, 87, 480, 1714, 2623, 1045, 2468, 1023, 2545, 2705, 1873, 2067, 2186, 445, 1487, 280, 2080, 402, 1836, 1518, 2860, 2616, 2294, 2185, 1837, 1453, 282, 1001, 1733, 286, 1547, 1542, 393, 2303, 2782, 1175, 1630, 2240, 786, 882, 1420, 679, 2127, 1595, 1373, 202, 2453, 2329, 2111, 2112, 1293, 760, 643, 1508, 684]. between [1046, 1872, 1848, 2358, 2897, 2348, 2126, 815, 151, 1806, 1699, 1762, 65, 484, 1790, 1735, 1718, 242, 777, 376, 1417, 2836, 886, 1206, 978, 2270, 858, 1151, 1503, 292, 1319, 2901, 1515, 2457, 895, 945]. Beverton [1905, 1867]. Bias [810, 404, 1065, 164, 57, 74, 1488, 2154, 1722, 2649, 1502, 2390, 2153]. Biased [131]. Biases [497]. Bicarbonate [752]. bichir [1288]. Big [1940]. Bigger [1378]. Bight [2682]. biliary [2719]. bilinearis [2688, 2476]. Bimodal [691, 529]. Binding [1153, 1154, 399, 2612, 1297, 1761]. Binomial [755, 982, 2679]. Bioaccumulation [801, 318, 398, 2008, 1402, 2499, 2029, 2803, 1866, 2793]. Bioassay [474, 285, 1342, 1610]. Bioassay-Derived [474, 1610]. Bioassays [1110, 1033, 967]. Bioavailability [879, 688, 1282, 2801, 652, 2612, 2093]. Biochemical [486, 2546, 922, 1001, 1143, 1461, 2776]. Biochemistry [1336, 2032]. biochimique [2833]. Biocontrol [831]. Biodiversity [1329, 2249]. bioeconomic [1698]. Bioeconomics [1430]. Bioenergetic [63, 1417, 1494, 2284, 2416, 1858, 2285, 2743]. Bioenergetics [1121, 690, 331, 2266, 857, 2217, 2625, 726, 190, 2736, 2516, 2391, 1640, 1639, 2420, 1520, 1630, 1585, 2354]. Bioenergetics-Based [190, 2736, 2420, 2354]. Biofilm [1793, 2537, 2730]. Biogeochemical [2878, 1114]. Biogeochemieal [1654]. Biogeochemistry [897, 2853]. Biogeographic [1621]. biogeography [2641, 2557]. bioindicator [2005]. Bioindicators [1384, 807]. Biological [328, 989, 237, 1269, 1408, 561, 1229, 1175, 568, 1029, 1554, 2075, 2736, 1551, 1681, 1682, 1683, 1801, 2560, 2908]. Biologically [102, 674]. Biology [927, 1148, 2822, 2356, 2526, 2506, 2402]. Biomagnification [2499, 1768]. Biomanipulation [1441, 1279]. Biomarker [1056]. biomarkers [2017]. Biomass [739, 706, 610, 501, 1886, 1225, 37, 613, 1419, 567, 542, 731, 705, 107, 374, 594, 574, 289, 506, 507, 1011, 858, 2309, 2410, 1840, 2462, 2713, 2307, 2599, 2262, 1582, 1892, 1898, 2561, 2649, 1693, 2033, 2261, 2772, 2368, 2555]. Biomass-size [1886]. Biomonitoring [1373, 2488]. biomonitors [1508]. biophysical [2901]. Biopsy [938]. Biota [95, 2487, 2342, 2878, 2793, 2869, 2378]. biote [2869]. Biotic [449, 682, 1446, 1345, 563, 2264, 1746, 2491, 1717]. biotransformation [2381]. Biotreated [960]. Biovolume [93, 1523, 2123]. biphenyl [2664, 2496, 2793]. biphenyls [1903, 2501, 2578, 2441, 1937, 2836]. Birch [332]. birds [1669]. Birth [961, 1816, 1777]. Bivalve [149, 1550, 1551, 1508, 1486]. Bivalves [865, 1387, 1388, 1620, 1927, 2731]. Bivalvia [1277, 1763, 1118, 1572, 916, 1528, 1309]. Bivariate [1126]. bjoerkna [2071, 945]. Black [229, 1323, 502, 1307, 2884]. Blackberry [1249]. blackfish [1757]. Blackflies [664, 1036]. Blacknose [967, 985]. Blackwater [9, 1524]. bladder [1807]. Bleached [1196, 674, 438, 743, 1383, 960, 1525, 2346, 2344, 1610]. Blicca [2071, 945]. Bloater [59]. Blood [667, 1165, 377, 1855]. bloom [2313, 2338, 1729, 2695, 2289]. bloom-forming [1729]. blooms [2263, 2505, 2287, 1730]. Blubber [142]. Blue [556, 605, 1764, 2505, 1533, 2287]. blue-green [2505, 2287]. Bluefin [32, 436, 2676]. Bluefish [1338, 2687, 1695, 1640, 1992, 2556]. Bluegill [1297, 953, 366, 190, 1637, 1816, 1814, 1738, 2424, 1971, 2339]. Bodied [518, 56]. bodies [1804]. Body [1020, 1244, 553, 1164, 1213, 2461, 490, 941, 1372, 491, 600, 548, 661, 1216, 100, 1011, 181, 182, 2859, 1749, 1816, 2190, 1511, 1900, 2630, 2442, 2770, 1583, 2829]. Body-Size [1011]. Boeckella [401]. bog [2686, 1539]. bog-rich [1539]. bomb [1615]. Bone [1023, 272]. Bongo [316, 2157]. Bonne [2413, 2763, 249]. bonnethead [1856]. Bonneville [1281]. Bony [392]. Bootstrap [2139]. Boreal [638, 803, 2618, 448, 1169, 1268, 1028, 1950, 2692, 1951, 2130]. borealis [349, 2464, 231, 612, 1176, 171, 92, 574]. Boreogadus [1535, 2266]. borne [2893]. Bosmina [55]. Bosminidae [1248]. Bottlenecked [994]. Bottom [669, 223, 807, 498, 432, 2039, 1810, 1907]. Bottom-Trawl [498, 1907]. Bottomed [67]. bound [2093]. Boundary [845, 581]. bouvieri [1479]. Bovine [148]. Bowfin [352]. bowhead [2647]. Bowland [46, 47, 44]. Brachyura [311, 2432, 1571, 2102, 2498]. Brain [1056, 188]. Brains [1288]. Brama [1776, 1148, 276, 918, 917, 2071, 945]. Branchial [885, 1139, 515, 276, 945]. branchialis [2422]. Break [645]. Breakdown [920]. breaks [2774]. Breakup [1408]. Bream [276, 918, 917, 945, 1496, 2071, 1739]. Breeding [249, 1034, 19, 2247, 2498]. brevis [2382]. Brief [699, 569, 2725]. Brine [500]. Bristol [1122, 1600, 2201]. British [425, 1258, 2248, 1055, 35, 2315, 260, 13, 2074, 1657, 82, 403, 774, 2235, 284, 671, 778, 2345, 1302, 608, 1337, 221, 2018, 537, 1475, 2189, 129, 754, 783, 2118, 149, 1322, 1672, 647, 999, 386, 1210, 1457, 199, 303, 716, 715, 1375, 1583, 114, 2906, 2638, 1400, 56, 2759, 2829, 1744, 2399, 1463, 2717, 2289]. Broad [1093, 2245]. broadcast [2669]. Broadscale [748]. brochets [2884]. bromide [1982]. Brood [345, 146]. Broodstock [1480]. Brook [667, 1010, 639, 803, 179, 853, 479, 1324, 2434, 105, 445, 1056, 186, 180, 183, 384, 158, 682, 254, 253, 225, 528, 513, 1190, 122, 184, 514, 185, 967, 1063, 376, 1239, 511, 1307, 181, 182, 2138, 1854, 1498, 1967, 2155, 2225, 1646, 1645, 1561, 534, 2620, 1511, 1920, 2243, 2063, 1837, 2467, 1564, 2473, 1532, 1699, 1718, 2743, 2270]. Brooks [41]. Brown [487, 1238, 991, 1355, 692, 740, 741, 929, 995, 718, 1452, 491, 393, 786, 2491, 277, 1307, 1706, 1084, 1970, 2394, 2366, 2373, 1758, 2620, 1707, 2010, 2183, 1720, 2294, 1939, 2028, 1490, 2130, 2796, 1705, 1704, 1687, 1635, 2236, 1989, 1699, 1703, 2195]. Browns [581]. Browsing [332]. Brunswick [2063, 2344, 2434]. Brush [496]. Btk [1265]. Bubble [216]. Buccinum [1447, 363]. Buckley [2913]. Budget [1391, 1951, 2363]. budgeting [2899]. Budgets [1437, 1057, 1555, 1931, 2743]. Bufo [24]. bugensis [2810, 2216, 1956, 2331, 1309, 1310, 1716, 1948, 2885]. BUGS [2762]. Bull [533, 2118, 2747]. Bullhead [487, 2236]. Bunker [1245]. Buoys [556]. burbot [1585]. Burdens [1052, 962, 1753]. Burial [1354, 2259, 1893]. Burrowing [280, 1941]. butyltin [2411]. By-Catch [196, 2671]. Bycatch [1086, 1430]. bypass [2006]. Bythotrephes [1556, 83, 794, 2843, 618, 694, 1494].

c [1415, 1648, 2874, 1873, 1093, 2876, 14, 1122, 1245]. C. [34, 780]. CaCO [44]. caddisflies [2860]. Caddisfly [586, 587, 600]. Cadmium [398, 1297, 118, 280, 887, 1973, 881, 1191, 1153, 1154, 644, 1359, 1146, 1637, 1605, 2501, 1977, 1975, 2699, 2768, 2479, 1976, 2731]. caerulea [477]. caged [1610]. Calamoecia [401]. Calanoid [1021, 1413, 1765]. Calanoida [322, 401]. Calanus [316, 2215, 2905]. Calcium [1219, 180, 183, 184, 194, 1146, 181, 182, 2504, 1761, 2296, 1912]. calculated [2203]. Calculation [684, 729, 1233, 1695, 2651]. Calculations [246]. Calibration [424, 2354]. California [1033, 1565, 2566, 2636, 1601, 1638, 477, 1489, 2714, 288, 2812, 2493, 2915, 146, 1088]. Caligidae [2074]. Caligus [954, 955, 1678]. Calliopius [1145]. Callorhinus [559, 961]. calva [352]. camtschatica [219, 345, 150]. camtschaticus [1086, 465, 949, 1600, 2201]. Can [1376, 518, 642, 1088, 2852, 2506, 1]. Canada [2148, 2315, 1753, 787, 2283, 321, 443, 747, 1752, 2111, 2112, 2164, 1744, 2333, 2514, 1748, 2077, 58, 1500, 2413, 159, 208, 2262, 592, 1196, 2124, 1321, 939, 1356, 868, 1764, 539, 116, 13, 561, 7, 350, 2063, 2359, 72, 2207, 1681, 2344, 2703, 894, 1678, 1692, 6, 95, 1968, 2906, 1851, 2558, 2076, 1976, 850, 2908, 2455, 1891, 1539, 1574, 554, 1461, 618, 2189, 2289]. canadense [1364]. canadensis [899]. Canadian [1206, 486, 1106, 1650, 1731, 2487, 976, 1813, 223, 501, 1370, 2635, 2232, 527, 241, 2470, 196, 2234, 829, 1299, 2166, 927, 2670, 1413, 975, 2559, 2405, 461, 1328, 891, 1762, 974, 2246, 1359, 1835, 2187, 1151, 873, 406]. Canadohta [305]. Cancer [906, 1097, 311, 723, 910, 796, 146, 1995]. Cancridae [311]. Cannibalism [1596, 518]. cannot [2620, 1568]. canonical [2564]. Cap [1932, 2200]. capability [2796]. Capable [464]. Capacity [597, 233, 1857, 2508]. Cape [498, 1157]. Capelin [1145, 1038, 838, 536, 517, 1227, 1291, 1686, 2159, 1922, 2365, 1932, 2451, 2170, 2840]. capensis [1286, 498, 522]. capital [2336]. Capsalidae [1663]. Captive [773, 548, 2355, 720, 1480, 1986, 1836, 1846, 2478]. Capture [741, 628, 852, 115, 998, 985, 2157, 2556, 2311, 2712, 1506]. capture-recapture [2311, 2712]. Captured [1384, 1211]. captures [1234]. Carassius [1482, 970]. Carbon [761, 173, 1979, 877, 1295, 539, 1093, 1936, 1169, 1951, 1031, 1062, 572, 1153, 1112, 1018, 1797, 1485, 1978, 1545, 2877, 1836, 2647, 1768, 2878, 2577, 1945, 2801, 1928, 2369, 2875, 2120, 2906, 2686, 2035, 2565, 1769, 2605, 1912]. Carbon-13 [1979]. carbunculus [660]. carcass [2730]. Carcasses [300, 2576, 2537]. Carcharhinus [732]. Carcinonemertes [146]. carcinophagus [1740]. cardiac [2678]. Caribbean [2698]. Carlo [2681]. Carolina [2338, 767]. Carotenoid [2195]. carp [2507]. carpio [2507]. carpionis [1679]. carps [2083]. carterae [2313]. Cascade [299, 2068, 1755]. Cascading [10]. Case [639, 672, 465, 1385, 574, 2764, 2127, 2255, 2126, 2438]. Cases [1427]. Caspian [2659]. Castle [1489]. Catamaran [2434]. cataractae [1656]. Catastrophic [301]. Catch [316, 573, 2171, 196, 128, 1129, 258, 717, 902, 947, 721, 21, 1256, 2708, 1509, 1990, 2614, 2011, 1803, 2055, 2764, 2688, 2172, 2777, 1786, 2107, 2231, 2530, 1568, 1603, 2679, 2003, 2255, 2483, 1712, 2513, 2399, 2671, 1883, 2000, 2510, 699]. Catch-Age [258, 2107, 2483, 1712]. Catch-at-Age [1129, 1786, 1603]. Catch-at-Length [21]. Catch-effort [2171]. Catch-per-Unit-Effort [721]. Catchability [740, 669, 40, 442, 192, 498, 374, 923, 1266, 2622, 2698, 1580]. Catches [771, 742, 594, 2548, 1234, 1487, 1796, 2681]. Catchment [1031, 893, 2312, 1647, 1885, 2881, 2838, 1789]. catchments [1726]. Catfish [836, 1282, 61, 88, 1926]. Cations [490, 2867]. Catostomus [4, 5, 1642, 2494, 1371, 1196, 2098, 2295, 2188, 2133, 2884, 679, 438, 1889, 376, 446, 1610]. Catskill [2138]. caudal [2814]. caudatum [130]. Caught [405, 1192, 1198, 863, 1657]. Cause [1036, 1354]. Caused [906, 1970, 1544, 2257, 1702]. Causes [40, 896, 1410, 2125]. cautions [1979]. CBE [1905]. Cd [1365, 1550, 1551, 2093, 800, 1742, 2884, 2833, 1976, 2704]. cederstroemi [83, 794, 618, 1494]. cederstroemii [694]. celery [2621]. Cell [1318, 870, 1903, 1821]. Cell-mediated [870]. Cells [365, 421, 977, 430, 1163, 1409, 2192, 2342, 2677, 2173, 1592]. Cellular [481]. cellularity [1548]. central [2566, 2306, 2307, 1780, 2905, 1523, 864, 2369, 1534, 2445, 1616, 2060, 2135, 1891, 2189, 405, 145, 241, 1262, 1072, 566, 293, 538]. Centrarchid [1039, 2461]. Centrarchid-Dominated [1039]. Centrarchidae [2773]. Centrarchids [760]. Centroberyx [1615]. Centropomus [1884]. cepedianum [761, 1872, 1364]. Cephalopod [866, 926]. Cephalopoda [1123, 2382, 1243]. Céphalopode [926]. Cephalopods [596]. cephalus [2821]. Ceratium [648]. Ceratomyxa [778]. Cercopagis [2659]. cernuus [342, 2104, 749]. Cerorhinca [1076]. ces [2833]. Chaetoceros [783, 1410]. chaetognath [1835]. Chain [317, 1201, 281, 709, 944, 833, 227, 2136, 2681, 2164, 1906, 1820, 2446]. chalcogramma [1983, 287, 606, 1519]. Chalinura [1667]. Challenge [482, 511, 2719, 2430]. challenges [1817]. chalumnae [1288]. Chamber [1238]. Chance [2805]. Change [454, 1030, 387, 1093, 1320, 1214, 388, 1366, 1709, 1710, 1767, 2563, 1708, 1929, 2341, 2455, 1795, 2562]. Change-in-Ratio [1030]. Changes [1357, 2309, 667, 407, 1244, 342, 989, 375, 1036, 2635, 2204, 2175, 560, 249, 762, 1119, 498, 1390, 887, 896, 457, 47, 1009, 1179, 2527, 283, 759, 1363, 743, 658, 2198, 1195, 1100, 1327, 1381, 744, 1141, 523, 41, 28, 1027, 961, 2211, 213, 948, 2731, 111, 898, 240, 2132, 2486, 2320, 2684, 2421, 1550, 2281, 2673, 2644, 1875, 1614, 1780, 2064, 2130, 2020, 1622, 1689, 2378, 2409, 1798, 1832, 1910, 2306, 2307]. changing [2643, 2314]. Channel [127, 836, 1282, 61, 2696, 2350, 88, 1168, 235, 1783, 1926, 2840]. Channels [674, 695, 221, 1525, 2176]. Chaoboridae [1299]. Chaoborus [909, 2141, 119, 1299, 2582]. chaos [1808]. Char [392, 1356, 445, 1144, 1476, 1170, 1285, 622, 621, 1151, 2034, 2380, 597, 1498, 2514, 2761, 1680, 2882, 2385, 2791, 1742, 1863, 2813, 2804, 2846, 1966, 1301]. Character [728, 1773]. characteristic [1706]. Characteristics [565, 94, 1213, 296, 1055, 485, 448, 1031, 438, 2100, 1651, 193, 974, 1063, 1135, 1081, 201, 1562, 1986, 2485, 2615, 1714, 2707, 2283, 2567, 2188, 2104, 2820, 2527, 2733, 2099, 1718, 1729, 1717, 2775]. Characterization [486, 2352, 1070, 1333, 1464, 177, 2408, 1813, 1869, 1779]. characterize [1768]. Characterizing [1050]. Characters [601, 1665]. Charr [105, 406]. Chatham [672]. Check [87]. Checkley [2916]. Chehalis [988]. Chela [1272]. Chelonia [1666]. Chemical [2737, 683, 989, 237, 1211, 942, 2396, 2736, 2735, 2448, 2142, 1562, 2281, 2862, 1555, 2263, 2493, 2429, 2111, 2112, 2776]. Chemicals [1270, 1113]. Chemistry [929, 806, 162, 957, 1072, 169, 1179, 44, 218, 10, 1536, 1562, 1500, 1540, 2312, 1988, 2435, 1885, 1963, 2602, 2053, 1539]. Chemometry [1023]. Chemosensory [384]. Chena [685]. cherry [2571]. Chesapeake [1640]. Chesterfield [554]. chez [1660, 719, 2833, 575]. Chile [1683]. Chilko [389]. China [2083, 2101]. Chinese [972]. Chinook [2911, 1465, 2465, 474, 937, 1565, 1819, 1244, 2737, 1258, 2248, 756, 822, 2352, 944, 464, 1619, 2714, 890, 889, 1454, 903, 2715, 1541, 912, 169, 809, 860, 1453, 1362, 2580, 2442, 1542, 1342, 778, 1838, 2156, 2547, 521, 1025, 2749, 950, 2758, 1275, 960, 1327, 2626, 1040, 1300, 242, 996, 1135, 2211, 2212, 2771, 1475, 1410, 2454, 1515]. Chionoecetes [2432, 1272, 841, 2413, 2763, 750, 540, 1030, 249, 414, 714, 1103, 1571, 2102, 1035, 1289, 2498, 1136]. chironomid [1732]. Chironomidae [1809, 388, 2445, 1727, 1574, 413, 821]. Chironomids [973]. Chironomini [1809]. Chironomus [2307, 2457, 821]. Chlamydomonas [1516]. Chlamys [1874, 1001, 303]. Chloramine [1336, 1855]. Chloramine-T [1336]. Chloride [421, 977, 430, 62, 1409, 1482, 1591, 2032]. chlorinated [1937]. chloromonad [2313]. Chlorophyll [1370, 509, 13, 195, 1202, 1203, 293, 29, 972, 826, 1560, 1653, 2206, 2296, 1523, 2672, 2552, 1965, 1759, 2674, 1789]. Chlorophyta [125]. Chlorophyte [1332]. Chlorpyrifos [135]. Choice [1238, 1242, 228, 190, 1638, 2896, 2669]. Chondrus [610]. Choosing [1510, 2108]. Chorion [970]. Chromatography [1138]. Chromosomal [464, 177]. Chromosome [297, 2804]. Chronic [398, 800, 1361, 2001, 2444, 247, 1200, 2722, 1566, 2543, 2592, 1774, 2292, 2540, 1948, 1941]. Chronically [488, 811]. chronologies [2832]. Chronology [22, 549, 727, 1100]. chronometer [1615]. Chrysophyrys [1496]. Chrysophyte [1395]. Chrysophytes [153]. Chub [769]. Chum [944, 896, 1453, 1457, 1214, 1458, 357, 1222, 1305, 823, 1459, 1460, 2559, 1658]. chydorid [2497]. Ciliates [1021]. Circannual [582]. Circulation [210, 196, 295, 1073, 2734, 2901]. Circulus [649]. circumlitus [2333]. Circumneutral [55, 843]. Circumventing [1965]. Cisco [60, 241, 1250, 34, 177, 2309, 2308]. Ciscoes [780]. CL [2601, 2633]. Cladocera [1396, 834, 2497, 2282]. Cladoceran [2427, 752, 936, 1495, 2659, 2479, 2843]. Cladoceran- [2427]. Cladocerans [578, 83, 1267]. Clair [1760, 442, 1387, 66, 1486]. Clam [854, 189, 428, 2543, 1589, 2082, 1524]. clapper [1669]. Clarity [669, 2012]. Clark [1707, 1705, 1704, 1702, 1706, 1703]. clarki [635, 1479, 2575, 1933, 2600, 1923, 1577, 2516, 2575, 2871, 1933, 2600, 437, 364, 139, 2348]. Class [416, 435, 969, 627, 1746, 1606, 2676, 2158]. Classes [653, 1810, 2080]. Classification [426, 1204, 2721, 2879, 2322, 2210, 2226, 2880]. Clay [844]. cleaner [1678]. Clear [565, 634, 1650, 2622, 2692, 2686]. Clear-Cut [565, 634]. clear-cutting [2692]. clear-water [1650, 2686]. Clearance [2103]. Clearwater [395]. cleavage [2329]. clés [1234]. Climate [1115, 979, 454, 1070, 1247, 1420, 686, 2132, 2708, 2614, 1767, 2739, 2444, 2368, 1795, 2904]. Climatic [991, 653, 122, 2158, 388, 2277, 2587, 2341, 2455]. Clinch [757]. Clinostomus [423]. clonal [1901]. clonal-sexual [1901]. clones [2582]. Cloning [1474]. Closely [987]. Closure [1262]. Clupea [1445, 797, 2021, 2619, 172, 2861, 2031, 578, 214, 168, 167, 1487, 266, 1471, 2024, 1073, 2478, 1934, 677, 882, 2073, 2726, 56, 643, 2767, 551, 1037, 121, 1162, 974, 2368, 216, 2529, 264, 50, 1278]. clupeaformis [656, 2502, 527, 655, 166, 34, 743]. clupeafotmis [886]. Clupeidae [601]. clustering [2702]. Clutch [322]. Clutches [401]. Co [1400, 2395, 1165, 2635, 1731, 2487, 2093]. Co-management [1400]. Coal [1298, 1960]. Coarse [608, 193, 2707]. Coast [1115, 473, 268, 1464, 525, 823, 1475, 1748, 2077, 1764, 1513, 1907, 2235, 2557, 1779, 2076, 2378, 2368, 2372, 783, 35, 1210, 942, 478]. Coastal [9, 371, 334, 872, 1090, 647, 1250, 386, 1262, 437, 364, 1254, 838, 659, 658, 1114, 370, 193, 1383, 461, 891, 541, 2738, 2264, 2519, 2575, 2871, 1933, 2600, 2160, 1542, 2530, 2627, 2866, 1909, 1483, 1825]. coastline [2469]. Coasts [863]. cobalamin [1938]. Cobble [807, 2350]. cobble-bed [2350]. Cobble-Bottom [807]. Cod [1140, 1263, 598, 862, 1358, 281, 698, 1038, 922, 192, 631, 1444, 1134, 1133, 1376, 138, 548, 1351, 1317, 1249, 128, 1157, 1172, 1044, 968, 1078, 374, 594, 1101, 1005, 946, 805, 537, 856, 372, 1178, 1161, 1491, 1701, 2760, 2057, 1997, 1670, 2912, 1748, 2077, 2844, 1986, 1840, 1788, 2546, 1743, 2466, 2839, 2011, 2397, 1810, 2855, 2219, 2518, 2519, 1839, 1535, 2266, 2785, 1882, 2816, 2669, 2422, 1681, 2525, 1846, 2718, 2327, 2662, 2913, 2572, 1606, 2237, 2170, 2561, 2524, 1526, 2530, 1586, 2058, 1690, 1604, 2866, 2100, 2099, 2144, 2329, 2863, 2891, 1848, 2056, 2358]. cod [2565, 2806, 2076, 2617, 2096, 1756, 2642, 1688]. Coded [737, 771, 2019, 2584]. coded-wire [2019, 2584]. Codends [804]. Coding [994, 431]. Coefficient [923]. Coefficients [1359]. Coelacanth [1288]. Coho [2708, 1797, 2576, 835, 1382, 756, 2177, 2597, 1824, 256, 899, 200, 999, 2896, 1454, 1541, 243, 809, 941, 789, 2119, 774, 2235, 2741, 1850, 482, 733, 1799, 472, 659, 658, 2627, 623, 1940, 2774, 2897, 2348, 950, 337, 724, 98, 1327, 1381, 1380, 308, 1300, 65, 484, 624, 1051, 1410]. Cohort [1097, 1796, 2250, 1942, 2169, 2060, 2056]. Cohorts [1097, 171, 2481]. coincide [2674]. Cold [1360, 1444, 1207, 2404, 2238, 2218, 2286, 2817, 2020, 2002, 2100, 2819, 2851]. Cold-water [1360, 2238, 2286, 2002]. Collaboration [410]. Collapse [1376, 765, 1882, 2327, 1604, 2726]. collected [2216, 1968]. collecting [2542]. collection [2842, 2157]. Collections [552]. collective [1483]. colloid [2322]. colloids [2322]. Cololabis [2196]. Colonization [2631, 920, 641, 1118]. colonne [2868]. Color [412]. coloradensis [1156]. Colorado [635, 1920, 1651, 1484, 1923]. Coloration [861]. Colour [951, 2369]. Colpoys [1112]. Columbia [783, 1055, 2315, 1210, 13, 2906, 1281, 1744, 2189, 2289, 2911, 129, 754, 2118, 2075, 425, 149, 1322, 1258, 2248, 1672, 647, 999, 386, 35, 260, 2074, 1657, 82, 403, 1457, 774, 2235, 199, 284, 303, 671, 778, 1838, 716, 715, 1375, 1583, 114, 2345, 1302, 852, 1400, 56, 2759, 2829, 2223, 608, 1337, 221, 2758, 2006, 2535, 2399, 2018, 537, 1463, 564, 2717, 1475]. Columbian [2638]. Column [753, 1190, 1975, 1752, 2868, 2304, 1847]. columnaris [2353]. columns [2313]. Colville [1250, 2245]. combination [1514]. Combined [2804, 2699]. combining [2386]. Comment [2912, 1709, 2360, 2232, 1978, 2181, 1710, 2879, 2284, 2233, 2182, 2916, 2913, 2051, 2151, 2180, 1931, 2126, 1878, 2076, 2153, 2562, 2654, 2825]. Commentaire [1852]. commentaires [25, 252, 274, 340, 409, 434, 516, 570, 593, 689, 713, 738, 788, 1260, 1283, 1377, 543, 914, 964, 990, 1016, 1041]. Commentary [1881, 1880, 1852]. Comments [469, 935, 2911, 25, 252, 274, 340, 409, 434, 516, 570, 593, 689, 713, 738, 788, 1260, 1283, 1377, 543, 914, 964, 990, 1016, 1041]. comments/Letters [1041]. comments/Lettres [25, 252, 274, 340, 409, 434, 516, 570, 593, 689, 713, 738, 788, 1260, 1283, 1377, 543, 914, 964, 990, 1016]. Commercial [771, 210, 745, 603, 863, 663, 558, 1552, 2055, 2764, 1861, 2699, 2513, 2550, 1883, 2000, 2510]. commersoni [4, 1642, 2494, 1371, 1196, 2098, 2295, 2188, 2133, 2884, 679, 438, 1889, 376, 446, 1610]. commetsoni [5]. Common [1175, 945, 2507, 1687, 2071, 1884, 2212, 1828]. Communities [1174, 905, 1152, 757, 703, 157, 1350, 116, 871, 457, 1169, 1404, 567, 1179, 532, 646, 1274, 1109, 506, 1811, 1573, 2269, 2427, 2106, 2148, 1875, 2415, 2160, 2371, 2552, 2086, 2261, 2347, 2004, 2408, 2135, 2114, 1916]. Community [156, 371, 55, 1228, 757, 730, 762, 394, 387, 855, 1396, 702, 825, 1150, 1372, 47, 51, 1003, 1203, 1253, 808, 687, 139, 1265, 956, 1383, 781, 858, 1166, 1950, 2225, 2320, 2495, 2713, 1700, 2392, 2799, 1871, 1977, 2315, 1714, 1961, 2064, 1944, 2452, 2403, 2515, 2569, 1628, 1938, 2017, 2492, 1685, 2110, 2304, 2605, 2187, 2794, 2220]. Como [886]. Comparative [1263, 1609, 360, 1856, 1639, 1282, 254, 1692, 687, 109, 1081, 42, 2639, 2217, 1568, 2866]. compare [2743].

Compared [356]. Comparing [1331, 1147, 2203, 883, 1916]. Comparison [701, 316, 801, 2766, 1226, 2298, 1060, 708, 1318, 259, 922, 2055, 1804, 2098, 1482, 1786, 125, 335, 53, 1471, 238, 1763, 1394, 1630, 677, 248, 1420, 110, 261, 843, 526, 869, 1105, 178, 2603, 1508, 729, 1464, 815, 891, 1473, 974, 2591, 446, 1417, 1935, 826, 1459, 945, 1776, 2521, 2765, 1487, 2744, 1936, 2382, 2725, 2509, 1965, 2329, 2390, 1779, 1521, 1608, 1599, 2091, 2657]. Comparisons [160, 23, 747, 1059, 1418, 1040, 2494]. compartimentation [2833]. Compartmental [1267]. compartments [1985]. compensation [2117]. compensatory [2694]. Competing [277]. Competition [2502, 711, 476, 1421, 253, 2086, 2348, 971, 2395, 2669, 2891, 1872, 2498]. Competitive [129, 1325, 342, 2034, 2666, 2574]. competitively [2102]. competitiveness [2468]. Competitors [56]. complanata [1801, 1818, 1146]. Complex [2257, 707, 867, 2737, 1694, 1144, 2317, 1901]. Complexation [24, 163]. complexity [2242, 1940, 2483]. Component [353, 1749, 2779]. Components [923, 2570, 1864, 1739]. composites [1804]. Composition [937, 1174, 1061, 892, 806, 1448, 768, 90, 249, 613, 32, 547, 1321, 1159, 1372, 548, 681, 1007, 248, 30, 936, 755, 940, 781, 1239, 348, 1205, 2859, 1811, 2689, 2410, 1590, 1991, 2713, 1617, 1618, 2862, 2064, 1547, 2703, 1606, 2577, 1496, 2875, 2833, 2329, 1938, 2056, 1684, 2089, 1685]. Compound [61]. Compounds [1024, 2818]. compromise [2316]. Computer [1208, 2764]. concavicornis [783, 1410]. Concentration [750, 170, 951, 1334, 678, 323, 2840, 882, 657, 1126, 1029, 1300, 1969, 2511, 1985]. Concentrations [976, 835, 1370, 173, 1219, 1194, 509, 984, 1240, 677, 1564, 679, 1373, 820, 2538, 973, 507, 2854, 213, 1206, 511, 1146, 324, 292, 1567, 1605, 2842, 1550, 1977, 1653, 1804, 1597, 2475, 2664, 1975, 2426, 1801, 2793, 1532, 2441, 2351, 2789, 1610]. concept [2709, 1920, 2888]. Conception [1622, 1291]. Concepts [361]. Conceptual [1089, 1709, 1710, 1708]. Concern [1345, 826]. concerning [1881, 1880]. Concerns [935]. Concordance [2859, 2727, 1150]. concordant [2559]. Condition [1020, 1355, 2327, 2509, 882, 612, 1336, 1198, 931, 1132, 2546, 2466, 2839, 2450, 1900, 1529, 1606, 2237, 2693, 2025, 2891]. Conditioned [943]. Conditioning [997, 1096, 1314, 1313, 2448]. Conditions [889, 723, 942, 1292, 358, 1317, 171, 2607, 2370, 2861, 1932, 2896, 1839, 2703, 2431, 2002, 2902, 2897, 1658]. Conducting [789]. confidence [2139, 2042]. confinement [1959]. confluentus [2118, 533, 2747]. confound [2494]. Confounding [752, 1427]. Congener [1008, 1937, 2793]. Congener-specific [1937]. Congeners [1318, 323, 493]. Connection [530]. Consequence [1011, 1092]. Consequences [2066, 813, 2515, 175, 1343, 1217, 1218, 2218, 2125, 2532, 2050, 2169, 2585, 2482]. Conservation [2063, 1170, 1087, 461, 2360, 2356, 2272, 2083, 1631, 2361]. Consideration [2607]. Considerations [947, 1552]. Conspecifics [384]. Constance [390]. constant [2314]. Constants [1154, 1760]. Constrained [1205]. constraints [1841]. Consumed [1061, 1104]. Consumers [230, 867, 1866]. Consumption [1182, 397, 756, 1164, 718, 331, 1417, 564, 1648, 2391, 2116, 1585, 1890, 2858, 2379, 1494]. contact [2570]. Containing [474]. Contaminant [454, 237, 310, 2384, 413, 2005, 2043, 2047]. Contaminant-Stressed [413]. Contaminants [487, 459, 2029, 2330, 1508, 1507, 1725]. Contaminated [1297, 752, 85, 323, 679, 1298, 978, 821, 2194, 2364, 2494, 2430, 1960, 1703]. Contamination [807, 549, 1211, 1775, 1550, 1551, 2281, 2488, 2411, 2164, 1916]. Contemporaneous [742]. Content [1357, 1155, 836, 1031, 1113, 728, 142, 1775]. Contents [2910, 549, 972, 2150, 2096]. Contested [2438]. contexts [2034]. Continent [1460]. continental [2046, 1907, 2299, 2131, 756, 1539]. continued [2396]. Continuous [597, 1482, 2120]. Continuously [1051, 181]. continuum [1875, 2007]. Contracaecum [1671, 1827, 609, 255]. contractile [1821]. Contradictory [1077]. Contrasting [920, 321, 1202, 1203, 1687, 2758, 242, 2911, 2766, 2552, 2214, 2220]. contributing [1984]. Contribution [4, 613, 13, 1187, 2164, 308, 2823, 1485, 1555, 1622, 2056, 2591]. Contributions [634, 1446, 1052, 433]. Control [761, 229, 822, 762, 561, 1419, 338, 935, 966, 28, 1051, 2075, 1644, 2579, 2460, 2208, 2505, 1895, 1698, 2681, 1598]. Controlled [394, 798]. Controlling [1095, 1440, 2767, 2289]. Controls [411, 173, 2338, 2007]. Contwoyto [2112]. conundrums [1979]. Convergence [902]. convergent [2237]. Conversion [1317, 62, 1231, 1733]. convolutus [783]. Conwy [1696]. Cooperative [1375]. Copepod [696, 1021, 653, 640, 2427, 2521, 2522, 1677, 2819, 2851, 1765]. copepod-dominated [2427, 2819, 2851]. Copepoda [322, 2074, 401, 1679]. Copepodid [955]. copepodids [1549]. Copepodite [316]. Copepods [215, 578, 1413, 2903, 1724]. coping [1795]. Copper [104, 118, 130, 679, 1052, 1153, 1154, 1018, 1567, 1566, 2193, 2801, 926, 2699, 2479, 1912]. Coracinidae [1286]. Coracinus [1286]. coral [2060, 2480]. Corbicula [104, 2543, 1589, 1524]. Cordillera [655]. Core [549, 1393]. Coregonid [545]. Coregoninae [1099, 1723]. Coregonines [33]. Coregonus [2309, 2308, 60, 1472, 1632, 656, 346, 2502, 390, 527, 241, 1250, 655, 969, 2081, 1093, 166, 34, 743, 2245, 2239, 177, 780, 1756, 1028, 1354, 886]. cormorants [2238, 2247]. Cornwall [2884]. Coronary [1817, 668]. Correction [1065, 57, 826, 1722]. Corrections [49, 187, 319, 463, 619, 763, 1353, 115, 915, 1066, 1189, 319, 463, 763, 915]. Correlate [1246]. Correlated [217, 871, 475, 1603]. Correlates [951, 431, 922, 2546, 1511, 1499]. Correlation [2017, 1319, 1595, 2603, 2657]. Correlations [2281]. correspondence [2564]. corridor [1760]. Cortisol [488, 1381, 624, 2364, 2385]. Corynosoma [799]. Coryphaena [1959]. Coryphaenoides [2788, 1667]. Cost [33, 2735, 2284, 2038, 1858, 2285]. costatum [2833]. Costs [106, 1217, 1218, 2521, 2816]. cottage [1780]. Cottus [2539]. Council [1851, 569]. Countergradient [2328]. countries [2310]. Counts [435, 1220, 356, 2027, 1834, 2744, 1962, 1521]. coupled [2298, 2882, 1894, 2854]. Coupling [598, 1386, 1267]. Courage [2745]. courante [633]. Course [513, 759, 514]. court [534]. courtship [2142]. Covariance [1022]. covariates [1800]. Covariation [1126, 2638, 2759]. Cover [1347, 98, 2896]. Covered [675, 2370]. CPUE [2393, 717]. Crab [1086, 712, 1272, 841, 750, 540, 1030, 219, 906, 249, 1097, 465, 910, 345, 796, 143, 831, 1103, 146, 150, 949, 1136, 2432, 2413, 723, 2127, 1571, 2126, 2102, 1995, 2498, 1600, 2201, 2275]. crabeater [1740]. crabes [414]. Crabs [236, 714, 1035, 1289, 2763, 2143, 311, 414, 1512, 2458]. Crappie [502]. Crassadoma [303]. Crassostrea [141, 325, 407, 1322, 1469, 2490]. Crayfish [547, 934, 1008, 874, 493, 760, 2072, 2090, 2492]. Crayfishes [1032, 2343]. Create [659]. Creek [14, 286, 1940, 1100, 1275]. crested [2247]. crispus [610]. cristata [851]. Criteria [1125, 1204]. Critical [73, 1098, 540, 1545, 1933, 1749, 2150, 2886, 1978]. Critique [165]. croissance [325, 719]. Cross [1065, 1807, 1869]. cross-section [1807]. cross-species [1869]. Crossing [81]. Crude [811, 978, 2865, 2619]. Crue [2868, 2869]. Crustacea [1204, 55, 2282, 311, 2008, 630, 1311, 2794, 1917]. Crustacean [2544, 492, 468, 1404, 607, 686, 1149, 444, 2675, 1556, 2305, 2105, 1916, 2901]. Crustaceans [1128, 553, 2163]. Cryptic [861]. Cryptobia [1655]. crysoleucas [662]. crystal [2847]. Cs [1731, 2487, 2488, 2284, 66, 1858, 2285, 2883, 2743]. Cu [1365, 1550, 879, 1742, 2884, 2833, 2704]. Cucumaria [1805]. cucumber [1805]. cue [1723]. cues [2448]. cuivre [926]. Cultivation [609]. Cultural [1276]. Culturally [1420]. Culture [411, 1004, 1071, 580, 1069, 972]. Cultured [937, 385, 420, 1096, 2173, 1663]. Cumulative [2817, 808, 1840]. cunner [1678]. Current [1323, 456, 963, 461, 39, 1787, 2621, 477, 2915]. currently [2462]. Currents [1208, 2738]. Curvature [69, 530]. Curve [790]. Curves [751, 50, 1568, 2781]. Cury [1852]. Cut [565, 634]. Cuticle [1341]. Cutthroat [635, 1479, 437, 364, 139, 1577, 2516, 2575, 2871, 1933, 2600, 2348, 1923]. cutting [2692]. cuttings [1774]. cuttlefish [926]. cyaneum [1447]. cyanobacteria [2703, 1847]. Cyanobacterial [1332, 2263, 2082, 2123]. Cyanobacterium [467, 2395]. cyanoprokaryote [1392]. Cycle [1145, 582, 1022, 907, 671, 924, 1103, 821, 2087, 2104, 2627, 2375, 2035, 1818]. Cycles [1348, 819, 2178, 2257, 2855, 2518, 2624]. Cyclic [68, 2808, 602, 97]. Cycling [207, 447, 857, 1627, 2253, 2079, 1759, 2324]. Cyclotella [13]. Cyfluthrin [1207]. cygnus [152, 2481, 1049]. Cynocion [336]. cynoglossus [598]. Cyprinid [583, 972, 2249, 2474]. Cyprinidae [276, 917, 2623, 583]. Cyprinids [310, 2821]. Cyprinus [2507]. Cystophora [851]. Cytochrome [317, 281, 1903, 1194, 1254, 1525, 2894, 2641]. cytogenetic [2023, 2274, 2025]. Cytological [926]. cytologique [926]. cytotoxicity [1551]. Cyymothoidae D [1098, 2037]. dab [2586, 2418, 2375]. Dace [423, 967, 1284, 1417, 985, 2290]. Daily [415, 1082, 404, 1142, 1246, 916, 494, 775, 525, 564, 2362, 2596, 2754, 1739, 2892]. d’ajustement [1234]. Dakota [1527]. Dälalven [2294]. d’albacores [1234]. Dam [1281, 1326, 1838, 2884, 2006, 495]. Damage [7, 350, 6, 95, 2677]. damaged [2402]. dams [2114]. damselfish [1503]. Danish [2497, 2010, 2183, 1332]. dans [325, 414, 926, 2868, 2869]. Daphnia [1781, 1495, 112, 589, 2907, 2479, 41, 1102, 2582, 42, 1917]. Daphniidae [1917]. dark [1517]. Darter [987]. Darters [359]. Data [1235, 690, 892, 708, 1304, 908, 1466, 12, 32, 349, 547, 1331, 1065, 1426, 1220, 1394, 848, 552, 747, 1129, 1312, 1223, 965, 717, 1198, 1425, 164, 3, 1077, 1037, 574, 103, 70, 722, 1256, 234, 1002, 1205, 1671, 1827, 2807, 2259, 1803, 2795, 2746, 2055, 2688, 2321, 2611, 2534, 2533, 2785, 2422, 2297, 2645, 2170, 1934, 2107, 2231, 1568, 1603, 1649, 2804, 2679, 2603, 2657, 2003, 2390, 2748, 2513, 2857, 2203, 1883, 2000, 2510, 1505, 2042, 2628, 1943]. Date [996, 1816, 2841, 2646]. Dating [207, 1393]. Day [810, 505, 1104, 2865, 2622, 2295, 2899]. Day-to-Day [1104]. Daylength [86]. daytime [2856]. DC [1343]. DDAC [2032]. DDE [690, 1768]. Dead [1281, 2503]. Dean [221]. Death [704]. d’eau [633, 2868]. deauratum [1277]. Debris [565, 127, 647, 301, 1168, 193, 2566, 2707, 2575, 2871, 2747, 2176, 2527, 1651]. Decadal [2684, 1033, 2132]. decadal-scale [2132]. Decapoda [1238, 1306]. Decapods [1026]. Decay [102, 1158]. decentralized [1953]. decipiens [486, 1813, 77, 951, 1673, 799, 630, 1674, 1675, 2065, 466, 870, 255]. Decision [716, 2402, 1822, 2650, 2398, 2751]. Decision-making [2402, 2751]. Decisions [963, 1429]. Decline [1294, 1387, 1388, 961, 2222, 2752, 1634, 2826]. Declines [198, 2629, 2708, 2640]. Declining [1212]. Decomposition [332, 300, 782, 2697, 2730]. decrease [2664, 2147]. Decreased [10]. deductive [1877, 2152, 2151]. Dee [2059]. Deep [675, 1159, 13, 236, 625, 29, 660, 1112, 2309, 1811, 1664, 2373, 2305, 1667, 2206, 2616, 1897, 2302, 2301, 2035, 2357, 2091]. Deep-Sea [236, 1664, 1667]. deep-water [2616]. Deepwater [1123, 2331, 1035, 823, 2175, 1909]. deethylase [2586, 2418, 1610]. defaunated [2604]. Defenses [881]. Deficiency [1290, 124, 259, 1896]. Defining [2241]. Definition [863, 1412, 2360, 1631, 2361]. Definitions [1431]. deforestation [2218]. Deformities [1732, 413, 821, 1809, 2457]. Deformity [590]. Deglaciated [710]. Degree [1285, 2295]. degree-day [2295]. Dehydroascorbic [170]. Dehydrogenase [356, 1902]. Delaware [1853]. delay [2663]. delayed [2334]. delivered [1655]. Delphinapterus [1120, 625, 1328]. Delta [1565, 848]. Delta-Lognornial [848]. demand [1640, 2710]. Demersal [2262, 791, 1042, 1043, 1195, 374, 2557]. Demethylation [572, 1633]. Demic [220, 921]. Demographic [813, 41, 1917, 2117]. demography [1856, 1845, 2646]. Demonstrate [755]. Denitrification [397].

Denitrifïcations [133]. Denmark [2464]. Dennys [2027, 1834]. Densities [1193, 2796, 1379, 2714, 1835, 2815]. Density [1199, 692, 740, 346, 2373, 2784, 610, 768, 1071, 40, 200, 331, 791, 197, 27, 382, 436, 934, 1493, 1579, 338, 1042, 152, 859, 1053, 923, 1364, 1005, 946, 784, 821, 2117, 2687, 2106, 2190, 2476, 1882, 1944, 2800, 2467, 2694, 1502, 2334, 2169, 1620, 1734]. Density-Dependence [200]. Density-Dependent [346, 40, 1042, 923, 1005, 946, 2784, 1493, 1579, 2687, 2476, 2694, 1502, 2334, 2169, 1734]. dentatus [775]. denticornis [1021]. Department [2232, 2234]. Dependence [925, 200, 436, 117, 2915]. Dependent [346, 40, 522, 791, 20, 1042, 869, 1105, 878, 923, 150, 475, 422, 1005, 946, 564, 1252, 2687, 2784, 1816, 1794, 2365, 2080, 2476, 2656, 2500, 2916, 2116, 2782, 1493, 1734, 1579, 2694, 2147, 1502, 2334, 2841, 2866, 2811, 2200, 2169, 2774, 2503, 2323, 2608]. Depends [1355]. Depensation [2288, 2806]. Depensatory [1348]. déplacements [414]. depleted [2476]. depletion [2568]. Deposition [1290, 7, 350, 72, 323, 6, 95, 1114, 176, 845, 893, 1697, 2501, 2015]. Depositional [1306]. Depression [480, 1739]. Depressions [1239]. Deprivation [1369, 2680]. Deprived [817, 931, 1051]. Depth [1125, 360, 1071, 108, 203, 542, 681, 467, 1015, 845, 475, 2305, 2218, 2621, 1576, 1956]. Depth-Dependent [475]. depths [2259, 1893]. Depuration [836, 605]. Deriso [1308]. derivation [1903]. Derivatives [972]. Derived [474, 1318, 747, 594, 1587, 2441, 1610]. describe [1562]. describing [2387, 2454]. Description [551, 2269, 1664, 2038]. Desert [1247, 1919]. Design [147, 1375, 1552, 2108, 2224, 2555]. Designed [716, 715]. Designing [2423]. Designs [789, 1000, 2049, 1488, 2154, 2153, 1913, 2275]. destination [2581]. Destructive [2890]. Detect [649, 2386, 1559]. Detected [1012, 944, 2490]. Detecting [1385, 2883]. Detection [281, 1241, 314, 1027, 2252, 2529, 2440]. Determinant [1201, 33]. Determinants [1098, 1341, 1900, 2088, 682, 820, 2860, 2158, 2034]. Determination [1174, 1010, 17, 71, 195, 363, 1398, 2787, 1794, 1518, 2729, 1962, 1649, 1521, 1737]. Determinations [868, 1750, 1582]. Determine [1169, 138, 729]. Determined [123, 623, 202, 1113, 2876, 1568, 1964]. Determines [227, 2835]. Determining [464, 1197, 1320, 1335, 1204, 1460, 1737, 1488, 2154, 2586, 2418, 2563, 2153, 2196, 2562]. detoxification [2047]. Detritivory [5]. Detritus [4, 67, 367, 608, 1401]. Detroit [2330]. Deux [787, 2833]. Developing [500, 1219, 1330, 2549, 2310, 2650, 2745, 2653]. Development [343, 597, 1504, 1672, 167, 942, 1186, 2359, 1673, 1229, 1581, 1447, 26, 339, 1380, 529, 494, 378, 1303, 554, 250, 943, 1313, 694, 2432, 1560, 2732, 1987, 1859, 1660, 1780, 1805, 1639, 2600, 2595, 2313, 1535, 2580, 2041, 2065, 1548, 2002, 2453, 2626, 1995, 6]. Developmental [712, 471, 1910, 1921, 2098, 2295, 2526, 2573]. Developments [1425]. développement [1660]. deviations [1892]. Device [720, 721]. Devils [1527]. d’hydrologie [410]. diadromous [1627]. diagnostic [1672]. Diagnostics [131]. diagrams [2718]. diameter [1888]. Diamond [804]. Diaphanosoma [55]. Diaptomus [909, 55, 322]. Diatom [701, 411, 580, 1070, 877, 604, 577, 1395, 2351, 1832, 1943]. diatom-inferred [2351]. Diatoms [783, 1349, 2673, 2531, 1847]. Dibenzo [652, 651]. Dibenzo- [652, 651]. Dicentratchus [1496]. Did [885, 1316, 2726]. didecyldimethylammonium [2032]. Didinium [130]. Didymocystis [1668]. Didymozoidae [1668]. Didymozoinae [1668]. Diel [22, 2095, 667, 1121, 2622, 1209, 1416, 696, 2518, 1549, 636, 727, 199, 204, 284, 80, 1531, 2690, 2290, 2230, 1689]. Diet [4, 22, 407, 1019, 1076, 135, 1758, 1181, 106, 1093, 682, 272, 1190, 2222, 2666, 1195, 270, 1401, 525, 784, 2691, 1836, 2279, 1526, 2247, 1862, 2280, 2089]. Dietary [900, 842, 170, 1194, 818, 2525, 1446, 1052, 644, 1301, 2067, 1836, 2203]. diets [2592, 2416]. Differ [1301]. Difference [621, 804, 2496, 2663]. Differences [754, 1655, 408, 582, 87, 1720, 1542, 236, 1705, 1478, 1259, 1380, 660, 484, 2270, 250, 1515, 1986, 2213, 2855, 2661, 2691, 1518, 2616, 1772, 2251, 2503, 2089]. Different [416, 1448, 1477, 1416, 1023, 583, 195, 1285, 330, 1149, 891, 523, 1314, 2034, 1970, 1504, 2761, 1990, 2713, 2328, 2711, 2896, 2705, 1996, 2098, 2581, 2002, 1904, 2503, 2605, 1911]. Differential [586, 587, 909, 1281]. Differentials [1237, 1881, 1880, 1879, 1878]. Differentiation [754, 656, 787, 766, 282, 2228, 357, 1284, 1463, 886, 1470, 849, 1778, 2623, 2834, 2041, 2821, 1691, 1635, 1848, 2358, 2559]. Differing [681, 802, 1158, 2466, 2012]. Differs [1360]. Diffusion [1291, 2748]. Digenea [1664, 1667, 1669]. digenean [1665, 1668]. digestibility [2592]. digestion [2089]. digestive [1581, 926]. Digitized [1065]. Diked [550]. Dilute [490]. Dilution [1110, 144]. dimension [1816]. dimensional [1701]. dimensions [2660, 2442, 2778]. Dimethyl [826]. dimethylbenz [2719]. Dimethylsulfoxide [2818]. Dimictic [1443, 981]. dinoflagellate [2446, 1924]. Dinophysis [2446]. dinucleotide [2016]. d’invertébrés [633]. dioxide [2686]. Dioxin [1318, 474, 1996, 2793, 378, 1303]. dioxins [652, 651]. Diploid [1454, 1453, 1455, 2270]. diploids [1569]. Diporeia [1991]. Diptera [1727, 1574, 664, 1036, 1809, 1299, 2445, 413, 821]. Direct [317, 1099, 1008, 230, 2078, 1765, 50, 1307, 2729]. direction [2581]. directions [2541]. directly [2513]. Discarding [1430, 2550, 2386]. Discharge [1445, 768, 251, 2189, 2007]. discontinuous [1573]. discounted [1861]. Discriminant [178, 1671, 1827]. Discriminate [1001, 1171]. Discriminates [1465, 2352]. Discriminating [601, 1735]. Discrimination [379, 1773, 986, 315, 2216, 478, 517, 2670]. Discriminations [1173]. discriminators [2637]. Disease [911, 2074, 988, 950, 1657, 2073, 2890, 2353, 2032]. disease-induced [2073]. Diseased [891]. Disentangling [2811, 1864]. Disequilibrium [1398]. dismissed [2806]. Dispersal [1286, 2472, 1250, 813, 439, 1095, 864, 2895, 2887, 2248, 2860, 2477, 1528, 2436, 1724, 2489, 2885]. dispersed [2284, 1858, 2285]. Dispersion [214, 1783, 1291]. display [1803]. Disposition [61]. Disruption [1728]. Dissolved [591, 675, 173, 877, 1320, 572, 519, 1153, 657, 940, 1018, 898, 2803, 2893, 2817, 2802, 1945, 2801, 2563, 2369, 1752, 2120, 1938, 2699, 1769, 2405, 2789, 1912, 2562]. distance [2770, 2581, 2390]. distances [2783]. Distantly [376]. Distinct [347, 2236]. distinguish [2195]. Distinguishes [470]. Distinguishing [1023, 1547, 1806]. Distribution [129, 783, 598, 360, 123, 473, 665, 1477, 613, 1038, 431, 386, 592, 901, 522, 439, 1294, 260, 373, 1476, 2013, 2905, 210, 287, 723, 1347, 196, 910, 1907, 2422, 688, 1280, 1073, 1588, 2174, 802, 323, 2069, 1264, 338, 2025, 538, 1214, 305, 231, 1176, 1312, 2144, 8, 88, 962, 80, 312, 171, 529, 650, 1328, 440, 946, 1266, 663, 913, 748, 996, 1239, 1832, 1359, 2317, 982, 2555, 1161, 2308, 1491, 1841, 2373, 2847, 2377, 2413, 1741, 2615, 2305, 2746, 1810, 2896, 2393, 1805, 2283, 2313, 2889, 1898, 1576, 2696, 2100, 2350, 2558, 1926, 2679, 2405]. distribution [1689, 2895, 1862, 2378, 2205, 1790, 2197, 2131, 1883, 2510, 1530, 1911, 2668, 1812]. Distributional [239, 236]. Distributions [1160, 626, 533, 450, 1302, 2035, 764, 92, 650, 1233, 1701, 1573, 1694, 2242, 2493, 2755, 2099, 2291, 2512, 2018, 2249]. Disturbance [189, 224, 222, 1405, 2564, 2485, 1745]. Disturbances [251, 1350, 327, 1855]. Diurnal [2380, 861, 1320, 2563, 2562]. Diver [552]. Divergence [1442, 113, 1641, 2882, 1868, 2526]. Diversification [812]. Diversity [735, 611, 1248, 81, 1468, 1271, 1478, 1458, 780, 1823, 2514, 2306, 2792, 2282, 2083, 2524, 2222]. Diving [625, 1584]. division [1821]. DMA [1575]. DMS [1249]. DMS-Related [1249]. DMSO [2818]. DNA [1465, 754, 1819, 2689, 379, 60, 1472, 1823, 126, 347, 1322, 994, 1813, 1644, 611, 2177, 281, 2216, 2352, 1814, 944, 1477, 464, 2031, 2276, 1829, 1641, 2469, 1468, 2641, 113, 1127, 1271, 939, 2265, 1476, 1843, 1192, 1170, 2133, 1474, 1099, 2083, 677, 1902, 1635, 1799, 606, 2236, 1478, 478, 1497, 1113, 2245, 828, 1467, 1848, 2239, 2559, 177, 780, 1473, 1040, 1735, 1305, 2591, 1557, 1891, 1949, 873, 2372, 1475, 849]. Dnieper [1310]. Do [1861, 2581, 956, 728, 2320, 1859, 1378, 1860, 2676]. Does [353, 1139, 1967, 2269, 2494, 2143, 2119, 2310, 2554, 2441, 559, 1510, 1952]. Dogfish [525]. Dolly [2118, 256, 899, 2199, 533, 1124, 2804]. dolomieu [2291]. dolomieui [383]. dolphin [1959]. d’omble [534]. Domestic [307, 254, 253]. Domesticated [919]. Domestication [1939, 2028]. Dominance [346, 941, 792, 793, 1332, 602, 97, 2774, 1847]. Dominated [761, 711, 1039, 2427, 2819, 2851, 2492]. Dominica [1627]. Domoic [411, 1004, 580, 605, 615, 2173]. Doppler [2257]. Doppler-like [2257]. Dorosoma [761, 1872, 1364]. Dorsal [483, 2070]. Dose [134, 2719]. double [2795, 2247, 1972]. double-crested [2247]. double-tagging [2795]. doubled [2635]. Dover [2131]. Down [803, 762, 2701]. Downstream [2005, 396, 132, 1324, 1320, 304, 2374, 1014, 2884, 2344, 2563, 2456, 2562]. Dr. [2658, 2827]. Dracunculoidea [1672]. Dragging [734]. Drainage [939]. Draining [638, 565]. Dramatic [1388]. dramatically [1613]. dredge [2887]. Dredging [906, 2711]. Dreissena [2810, 1648, 1182, 2728, 1849, 2129, 980, 2216, 1117, 2175, 2501, 2488, 2723, 439, 865, 2296, 1597, 2103, 2543, 2015, 2253, 1280, 1782, 518, 1520, 2426, 864, 1246, 1264, 1759, 1956, 1928, 1760, 2071, 1387, 616, 1589, 588, 859, 1620, 1927, 2347, 2331, 1309, 2870, 1388, 726, 1310, 1716, 2492, 2363, 440, 1899, 1948, 2280, 2885]. Dreissena-dominated [2492]. Dreissenid [1116, 736, 2810, 1572, 2674, 1771, 1948, 2409, 1793]. Dreissenidae [1118, 1572, 1528, 1309]. dreissenids [2216]. Drift [371, 1416, 626, 228, 792, 793, 1193, 503, 985, 1778, 1746, 1842, 2144, 2210, 1919]. Drift-Feeding [228, 792, 793]. Drifting [864, 2330]. drill [1774]. driven [2718, 1486]. droebachiensis [966]. Drum [591, 502]. Drunella [1156]. Dry [984]. dual [2729]. due [810, 2548, 7, 350, 6, 95, 1195, 2351]. d’un [1852]. Dungeness [906, 1097, 2143, 723, 910, 796, 1995]. Duplicated [234]. durant [575]. Duration [835, 1602, 1382, 2087, 2763, 1561, 2751]. During [159, 1454, 251, 1381, 2309, 1726, 1925, 1352, 656, 1382, 785, 1437, 675, 161, 560, 1438, 2305, 655, 645, 1932, 2877, 1294, 2871, 2013, 969, 169, 887, 1408, 1333, 2327, 2134, 984, 430, 443, 796, 733, 2553, 1108, 1363, 2137, 1595, 2020, 2002, 1931, 2355, 2453, 1221, 1937, 1855, 2375, 1757, 2018, 663, 961, 2498, 378, 1935, 2304, 1745, 2695, 1865]. Dwarf [177, 886]. Dwarfed [1144]. Dwelling [256, 400, 359, 2348]. dybowskii [2841]. Dynamic [1060, 615, 1533, 883, 1852]. Dynamics [761, 2039, 9, 1180, 690, 806, 1428, 1111, 599, 901, 280, 460, 2072, 825, 16, 1419, 278, 52, 731, 1003, 907, 518, 1108, 67, 1243, 640, 30, 1904, 869, 603, 2416, 746, 608, 1381, 86, 366, 1308, 781, 42, 1049, 2810, 2136, 2257, 1841, 2566, 2780, 2320, 2462, 1540, 2391, 1629, 2397, 2141, 2685, 1999, 2038, 1625, 2471, 1782, 2613, 2702, 2515, 1760, 2624, 1532, 2506, 2338, 2686, 1894, 2808, 2890, 1524, 1899, 2319]. dynamique [1852]. Dystrophic [843].

E. [987, 1081]. each [2574]. Early [1465, 1833, 179, 1263, 598, 473, 1921, 1805, 2600, 243, 46, 1338, 1083, 339, 2329, 1259, 950, 2121, 378, 1303, 1307, 250, 1014, 2021, 1695, 2041, 1957, 2553, 2841, 2247, 2866, 2897, 2895, 1899, 1995]. Early-Stage [598]. east [2235, 2557, 1234, 1459]. Eastern [58, 665, 196, 636, 1469, 2333, 1748, 2077, 2175, 2282, 2647, 2063, 1860, 1678, 2108, 1772, 2058, 2670, 2076, 2062, 1086, 116, 7, 350, 6, 283, 1172, 95, 606, 150, 1206, 369]. Eau [648]. écaille [575]. Ecdysone [750]. Echinodermata [1805]. Echinogammarus [2105]. Echinorhynchus [1107]. Echo [1065, 426]. Echosounders [1101]. éclair [2868, 2869]. Ecological [334, 1694, 1117, 2539, 1186, 1342, 1138, 2050, 1730, 2415, 2532, 1649, 2343, 2826]. Ecologists [2]. Ecology [1329, 219, 1372, 254, 1118, 1577, 2879, 2272, 1851, 2226, 2880]. economics [2402]. Ecosystem [452, 299, 2213, 251, 301, 2915, 300, 765, 18, 39, 1601, 2644, 1553, 2079, 2387, 2902, 2007]. Ecosystems [739, 872, 1434, 997, 1268, 355, 1011, 1978, 1545, 2837, 2340, 2241]. ecotonal [1979]. Ecotoxicology [2612]. Ecotypes [357]. ecozones [2635]. ectothermic [2725]. Eddy [155]. Edge [531]. Edible [683, 1060]. Editorial [410, 2209, 2184, 2209, 2184]. Editorial/Éditorial [2209, 2184]. EDTA [1567]. edulis [556, 700, 2093, 1692, 605, 615, 1533]. Eel [1212, 1215, 1611, 2367, 1581, 2733, 2625]. eelgrass [2219]. Eels [1211, 2443, 2281, 1677]. Effect [2810, 2138, 2760, 1997, 740, 835, 1382, 1638, 1323, 170, 1743, 299, 130, 819, 2680, 241, 1250, 485, 2792, 1119, 890, 1262, 509, 1695, 389, 1888, 1134, 1193, 2791, 1490, 1946, 716, 715, 175, 247, 2065, 122, 1447, 2813, 2592, 997, 339, 1628, 666, 767, 337, 1327, 29, 2383, 1890, 1971, 1636, 2899, 1736, 2339, 2628, 1828, 1018, 2668, 2667, 163, 1922, 1871, 1617, 1618, 2206, 2507, 1992, 2693, 1738, 1796, 2100, 2891]. Effective [1129]. Effectiveness [745, 143, 659, 495]. Effects [2465, 1182, 488, 1199, 1950, 1004, 638, 803, 1841, 1987, 396, 1770, 991, 1509, 2874, 2738, 344, 135, 1637, 395, 2225, 2395, 237, 2159, 1164, 1165, 524, 2830, 768, 1540, 2485, 148, 2106, 2061, 2873, 296, 2242, 2046, 2011, 840, 1438, 699, 1441, 922, 2831, 331, 23, 1196, 1988, 96, 1194, 2644, 2172, 428, 423, 207, 1439, 490, 834, 847, 106, 912, 2176, 2296, 385, 2850, 180, 332, 46, 273, 16, 1361, 2543, 1372, 1516, 2740, 278, 195, 72, 52, 2475, 2876, 1426, 2015, 2253, 932, 1229, 607, 1317, 51, 2134, 731, 2770, 1994, 827, 1003, 705, 2604, 2478]. Effects [84, 2577, 2426, 1007, 1008, 2683, 230, 1344, 2447, 304, 521, 1367, 550, 545, 1532, 808, 536, 30, 720, 1897, 1302, 1336, 2302, 1306, 2292, 2540, 18, 577, 367, 1265, 2734, 505, 374, 950, 270, 846, 960, 194, 89, 2626, 721, 2293, 657, 2353, 1067, 506, 1276, 150, 86, 1326, 1027, 1688, 781, 985, 996, 1658, 269, 2341, 2783, 511, 1303, 57, 971, 2114, 2605, 1002, 2794, 333, 1941, 2730, 418, 419, 181, 182, 2032, 10, 1189, 2189, 1096, 1314, 2775, 2776, 2005, 2859, 2722, 1573, 2095, 2129, 1970, 2146, 2257, 1751, 2735, 1990, 2021, 2912, 1551, 2599]. effects [2397, 1977, 2141, 2010, 2183, 2639, 1767, 1885, 1961, 2023, 2523, 2621, 2656, 2500, 1860, 1944, 2525, 2741, 1643, 2913, 2572, 1963, 2817, 1593, 2702, 2649, 2509, 2633, 1583, 1956, 2881, 2838, 2071, 2491, 2750, 2811, 1765, 2350, 2068, 1771, 2699, 2145, 1727, 2348, 2712, 2845, 1769, 2492, 1982, 1995, 2280, 1730, 2135, 2843, 1795, 1703]. Effets [2833]. Efficiency [1231, 496, 1488, 934, 2071, 115, 985, 2711, 2088, 2515, 2157, 2417, 2006, 2154, 2153]. efficient [1737]. Effluent [1196, 674, 278, 304, 438, 743, 1383, 960, 1525, 2346, 2817, 1610]. Effluents [590, 2370]. Effort [901, 1698, 1129, 717, 721, 982, 1136, 1573, 1841, 2171, 2172, 2777, 2186, 2533, 2596, 2231, 1904, 2003, 1719, 1883, 2000, 2510]. Egg [797, 2215, 302, 167, 2580, 548, 345, 609, 552, 1447, 489, 444, 146, 1354, 2648, 2732, 2362, 1986, 2763, 2259, 2188, 1846, 2572, 1675, 1893, 2813, 2099, 2329, 2767, 1784, 2555, 2022]. Egg-Bearing [444]. Egg-Laying [609]. egg-lipid [2329]. Eggers [404]. Eggs [474, 937, 598, 1145, 167, 970, 180, 1024, 856, 1701, 2760, 1590, 1666, 2746, 2449, 1673, 2286, 2065, 1586, 2100, 1908, 2144, 2555]. Eight [149]. either [597]. ELA [1441]. elasmobranch [2150]. Electricity [1306]. Electrofishing [1343]. Electrolytes [271]. electromyograms [2168]. electron [2298, 2322, 1608]. electrophoresis [1671, 1827]. Electrophoretic [1187, 517, 1464, 441, 369, 1779]. Electroporation [1557]. Electroshock [1344]. elegans [1835]. Element [807, 450, 1407, 973, 2670]. Elemental [2847, 1358, 1177, 2298, 1617, 1618]. elementary [1803, 2051]. Elements [983, 450, 1340, 14, 1122, 324, 2571, 2538]. Elevated [488, 396, 510, 1056, 704, 1995, 1969, 2061, 2886]. Elevation [283, 2678, 1700, 1920]. Eleven [681]. Elicited [941]. Elimination [1760, 2082]. elle [719]. Elliptio [1801, 1818]. Ellptio [1146]. elongatus [954, 955, 1678, 109]. Embayment [882, 1946]. Embedded [620]. Embrace [1272]. embryo [2595, 2024, 1893]. Embryonic [535, 694, 2098, 2295, 2329, 1721]. Embryos [500, 24, 827, 1833, 2023, 26]. Emerged [1190, 65, 2248, 1564]. Emergence [415, 1036, 536, 1646, 1727]. Emergent [54, 1230]. Emerging [1114]. Emigration [2749, 1837, 2270]. emission [2298]. Emphasis [100, 2018]. Empirical [994, 1722, 997, 869, 1105, 1137, 2836, 1629, 2685, 1909, 1558, 633]. empiriques [633]. Enclosures [968, 652, 651, 1063, 135, 2452, 1897, 2280]. encounter [2570, 2811]. End [830, 1756]. End-of-Summer [830]. Endangered [527]. Endocrine [885]. Endurance [583]. energetic [2290, 2639, 2038, 2426]. Energetics [2394, 358, 303, 2735, 1639, 1535, 2771]. énergétique [325]. Energy [1128, 872, 818, 1273, 746, 100, 2343, 1314, 325, 1792, 2567, 2327, 2770, 2467, 2568, 2870, 2121, 2363]. Engineered [148]. Engineering [1851]. England [1773, 1349, 2685, 384, 1698]. English [481]. Engraulis [2506, 765]. Enhance [842, 906, 1091, 1101]. Enhanced [2401]. Enhancement [1091, 745, 1214, 2122, 1484]. Enough [176, 1641]. Enrichment [1033, 705, 455, 1950, 1650, 2087, 2213]. Enrichments [165]. Entering [932, 1014]. entrained [2892]. Entrainment [2887]. entropy [2510]. Entry [1259]. Envelope [1333]. Environment [1428, 1321, 280, 1150, 769, 1107, 601, 1306, 950, 1778, 2095, 2320, 2855, 2595, 2567, 2757]. Environmental [1139, 1010, 700, 995, 1334, 2460, 576, 933, 912, 561, 1302, 2338, 651, 966, 484, 57, 1002, 2009, 333, 292, 1607, 1775, 2861, 2094, 2839, 1511, 2792, 1614, 2740, 1499, 1588, 1593, 1621, 1693, 2086, 2050, 2569, 2292, 2378, 1800, 2084, 1519, 1832, 2731]. environmentally [2718]. Environments [94, 335, 687, 308, 1035, 413, 1987, 2366, 2314, 2145]. Enzymatic [583]. Enzyme [407, 553, 944, 431, 191, 225, 1221, 1833, 2381, 2020]. Enzyme-2 [191]. Enzymes [420, 2466]. eos [2290, 1284, 1417]. Ephemeral [825, 230]. Ephemeridae [1941]. Ephemeroptera [280, 1156, 1941]. epialti [146]. Epibenthic [156]. Epidermal [183, 1889]. epilimnetic [2287]. epilimnion [1975]. Epilinnnion [429]. Epilithic [174, 753, 206, 2799]. Epilithon [1440]. epinepheli [1663]. Epinephelinae [2060]. Epinephelus [1676, 1177]. Epiphytes [1340]. epiphytic [2320]. epiphytism [2890]. Epiphyton [542]. Episodes [1324]. Episodic [1970, 1561, 1350, 967, 1239, 1844]. epithelial [1592]. Epizoology [1075]. equal [2293]. Equation [160, 524]. Equations [465]. Equilibria [1165]. Equilibrium [526]. equipment [2517]. equivalency [1903]. Equivalent [1318, 1610]. Equivalents [474]. Erie [1849, 980, 1117, 2175, 2501, 162, 161, 1538, 83, 2644, 865, 557, 1261, 1792, 1597, 2160, 2710, 2015, 2253, 1009, 1520, 864, 224, 1760, 202, 2331, 1388, 222, 2492, 2363, 1929, 2409, 1793, 784]. eriensis [13]. Erignathus [123, 2174]. Erken [1911]. EROD [1642]. Erosion [1226]. Erratum [2389, 2027, 1632, 1827, 2183, 2656, 1711, 2028, 2587, 2586, 1734, 2881, 1906, 2657, 2311, 1779, 2388, 2609, 2608, 2585]. erroneous [2830]. Error [526, 770, 2462, 2172, 2553, 1712, 2123]. Errors [344, 164, 2250, 1603]. Errors-In-Variables [164]. erythrocytes [2094]. Erythron [887]. Escape [368, 1910]. Escaped [1084]. Escapement [790, 1220, 998, 97, 2689, 2297, 2161, 2311]. Escapements [789, 2744]. Escapes [329, 2421]. escaping [1756]. Esox [205, 2242, 2469, 2876, 2884, 1994, 590, 1373, 1081, 292]. Essential [1340]. est [719, 1234]. est-elle [719]. Establishment [617, 794, 2674]. Estimate [632, 1082, 547, 27, 91, 1273, 446, 1002, 2788, 2584, 1501, 2795, 2140, 1487, 2243, 2300, 2535, 2323, 2608, 2555]. Estimated [392, 344, 795, 1215]. Estimates [1183, 810, 994, 732, 509, 857, 1431, 305, 1154, 63, 902, 1436, 755, 1027, 1417, 2544, 2168, 1839, 1786, 2856, 2528, 1945, 2649, 2250, 2078, 2200, 1712, 2559, 2772, 1800, 2139, 2293, 1762, 2089, 2042, 2786]. Estimating [1573, 1565, 285, 2019, 2735, 309, 790, 718, 32, 1740, 404, 772, 336, 2520, 557, 2744, 789, 2886, 1426, 1220, 934, 2528, 248, 2334, 1942, 152, 717, 845, 142, 1415, 998, 1037, 574, 2513, 1000, 2463, 1912, 495, 2117, 2548, 1887, 2038, 2170, 2261, 2414, 1505, 2458, 2091, 1737, 1913, 1972, 2311]. Estimation [1235, 556, 1514, 1094, 2590, 1242, 2449, 1399, 2779, 566, 1192, 1881, 952, 75, 131, 1880, 1237, 1129, 1053, 2712, 103, 2512, 993, 475, 483, 93, 57, 1626, 74, 50, 2742, 1136, 1205, 2689, 2462, 2814, 2171, 2172, 2777, 2407, 2610, 2611, 2724, 2596, 1838, 1502, 1558, 2748, 2339, 2628, 2425, 1503, 1879, 1878]. estimator [2414]. Estimators [71, 2645, 526, 2462, 2423, 1488, 2154, 2153]. Estuaries [872]. Estuarine [1012, 94, 1107, 472, 1191, 232, 308, 2443, 1639]. Estuary [156, 556, 1181, 1211, 1073, 2344, 2338, 517, 1298, 1686, 2687, 1741, 2013, 2104, 1802, 2840, 1696, 2558, 1899, 1960, 2682, 2289, 1075, 214, 1481, 1083, 966, 2901]. esturgeon [719]. Esva [1187]. Etelis [660]. eternal [1852]. éternel [1852]. Ethanol [826]. Etheostoma [987]. ethoxyresorufin [2586, 2418, 1610]. ethoxyresorufin- [2586, 2418]. eukaryotic [2395]. Eumetopias [2222]. Euphausia [2856]. Euphausiid [902]. euphausiids [2299]. euphotic [1951]. Europe [470, 1563, 1311]. European [379, 1611, 2367, 2821, 2754, 1496, 1691, 2625, 1215, 2848, 891, 1735, 618]. Euryhaline [1191]. Eutrophic [1443, 1394, 110, 1216, 18, 1354, 2309, 2306, 2307, 2305, 2876, 2578, 2302, 2301, 2287, 1847]. Eutrophication [1396, 1114, 1195, 2135, 1395, 1744, 2351, 2531]. Eutrophication-Induced [1195]. Eutrophied [1420, 2720]. Evacuation [383]. Evaluate [1269, 972]. Evaluated [745]. Evaluating [2736, 1237, 2681, 1831, 2631, 2549, 1881, 1880, 1879, 1878, 1913]. Evaluation [2075, 1020, 632, 368, 2637, 2410, 1348, 1665, 757, 309, 2888, 108, 72, 1229, 6, 715, 294, 370, 2070, 1872, 1433, 356, 1162, 1104, 1432, 2653, 1531, 2149, 2024, 1844, 2263, 2274, 2228, 2456, 2006, 2096]. Even [480]. Even- [480]. event [2617]. Events [1408, 16, 1162, 1921, 2739]. Everglades [637]. Evidence [220, 810, 2118, 598, 1485, 102, 1794, 655, 35, 445, 903, 1123, 1142, 14, 1122, 2169, 455, 277, 555, 1797, 2687, 1551, 1680, 2230, 2067, 1580, 2905, 2691, 1868, 2523, 2669, 2396, 2571, 2236, 2358, 2808, 1929, 2249, 1917, 1861]. Evidenced [1444]. Evokes [1139]. Evolution [710, 656, 675, 655, 1450, 2773, 1851]. evolving [2583]. exact [1972, 2742]. Examination [1749, 1978, 1545, 2104, 2473, 2191, 1943]. examine [2896]. Example [198, 336, 131, 374, 74, 1918, 2512]. Examples [1089, 441]. exceeding [2886]. Exceptional [1115]. Exceptions [2771]. excess [1566]. Exchange [280, 402, 2303, 1095, 432, 140, 235, 2787, 1714, 2435, 2134]. Exclusion [1306, 2507]. Excretion [38, 428, 88, 1648, 2719, 2134, 2120]. excursions [1932]. Exercise [1362, 1344, 377, 1096, 943, 1314, 1313]. Exercise-Conditioned [943]. Exercised [699]. Exhaustive [377]. Exhaustively [699]. Exhibiting [782]. Exogenous [744]. exophytic [2828]. exotic [1495, 2253, 2090, 2162]. Expansion [970, 1784]. Expected [524]. Expected-Value [524]. Expenditure [1273]. experience [1751, 2207]. experienced [1541]. Experiences [1400]. Experiment [2043, 716, 715, 1411, 1973, 1754, 1972]. Experimental [174, 989, 664, 15, 291, 165, 999, 394, 1617, 1618, 674, 1396, 702, 1434, 1961, 307, 43, 51, 1394, 1003, 1674, 68, 503, 1889, 1306, 867, 1727, 89, 255, 1276, 1976, 1728, 1368, 1766, 971, 1395, 1525, 2784, 1605, 2877, 2594, 1888, 1975, 2120, 2198, 2492, 1913, 410, 1732, 1437, 1441, 1321, 1350, 1389, 1434, 1392, 1397, 2703, 1973, 1931, 1725]. Experimentally [157, 1393, 893, 2604, 2547, 1976, 1729]. expérimentaux [410]. Experiments [751, 1185, 144, 351, 620, 222, 1432, 1969, 2450, 2795, 1397, 2078, 2742]. Explain [1370, 176]. explained [2620]. Explanation [1348, 2806]. explicit [1953, 1630]. Exploitation [362, 1049, 2235, 2417, 1795]. Exploited [1237, 226, 1881, 1880, 1879, 2073, 1878]. Exploration [569]. Explore [874, 725]. exploring [1755]. export [2618]. Exposed [911, 1297, 515, 1324, 800, 1207, 704, 262, 188, 743, 1406, 881, 1383, 481, 811, 856, 327, 181, 1525, 2193, 2678, 2619, 1605, 1707, 1705, 2274, 2002, 1774]. Exposure [885, 835, 1382, 699, 674, 310, 887, 183, 1361, 1456, 1343, 1344, 184, 1336, 728, 1380, 657, 1239, 1307, 2722, 2192, 2736, 2735, 2865, 1551, 1561, 1482, 2621, 2791, 2043, 2024, 2444, 2047, 1959, 1760, 1937, 1855, 1528, 2699, 2353, 1955, 2582, 1721, 2776]. exposure-introduction [2043]. exposures [1704, 2156, 2050]. Expression [2619, 1833]. Extant [153]. Extended [86, 1796, 2304]. Extension [892, 2752]. extensive [1644]. Extent [7, 350, 6, 95]. External [700, 38, 175, 743, 2278]. Extinction [1255]. Extracellular [328]. Extract [870, 481]. Extractions [826]. Extracts [474, 870]. extrapolation [2777]. extravariation [2042]. extreme [2739]. extremes [1745]. Extrinsic [356]. Exxon [2021, 2023, 2026, 2274, 2700, 2022, 2025]. Eye [2041]. Eyes [955, 2387].

f. [411, 580]. Fabens [2425]. Fabricius [714]. Face [68]. Facilities [1131]. Factor [1355, 752, 33, 57, 1132, 2586, 2418, 2891]. Factors [1365, 639, 700, 2366, 322, 1318, 391, 479, 1489, 995, 1334, 449, 809, 1332, 238, 567, 286, 2403, 1446, 1957, 1015, 1264, 758, 11, 832, 1402, 1039, 2113, 1825, 2287, 292, 2289, 2264, 2410, 1903, 2597, 1661, 2415, 1746, 2740, 1588, 2493, 1807, 1621, 1801, 2086, 2767, 2845, 2378, 1519, 2009, 2388, 2609]. factory [2406]. Facts [1470]. Facultative [5]. FADs [1426]. failed [2641]. Failure [842, 297, 704]. failures [1634]. Falkland [1683]. Fall [521, 1100, 2871, 2559, 2270]. fall-run [2559]. fallout [2578]. Family [236, 1665, 2356, 1860]. Far [1459]. Farm [1475, 2421]. Farm-reared [1475]. Farmed [1442, 329, 330, 1084, 1678]. fasciatus [1365, 1074, 1968, 356]. Fast [176]. Fat [836, 643, 326]. Fate [1167, 1731, 2487, 651]. Fathead [1199, 135, 1207, 1017, 1113, 1018, 2391, 2227, 1482, 1643, 1912]. fatness [2626]. Fatty [937, 1023, 1131, 2226, 324, 1301, 1925, 1590, 2525, 2329, 1806, 2879, 2880]. Fauna [156, 2202]. Faunal [224, 2636]. favors [2846]. Fe [879]. Features [127, 2085]. Fecundity [797, 2763, 1715, 311, 548, 595, 225, 134, 1103, 2580, 2478, 1947]. Fed [383, 978, 2173, 2705, 2435, 2625]. Feed [215, 1249, 2592]. Feedback [1308, 2336]. Feeding [22, 1352, 342, 853, 499, 1209, 773, 1021, 276, 918, 917, 228, 792, 793, 816, 1421, 1069, 1372, 1338, 1351, 91, 727, 2030, 531, 443, 968, 202, 359, 589, 1100, 89, 744, 972, 525, 372, 1092, 1217, 945, 2810, 2380, 2761, 2150, 2655, 2716, 2545, 2705, 2654, 1517, 1815, 2740, 1581, 2170, 1738, 2071, 1757, 2474]. Fellodistomidae [1664]. Female [346, 1477, 560, 1362, 1447, 1046, 1103, 767, 949, 1303, 74, 579, 2432, 1925, 2472, 2763, 719, 1777, 2143, 2498]. Females [909, 1124, 264, 1084, 2102]. femelle [719]. Fen [638, 1727, 1539]. Fence [1220]. Fenitrothion [188]. ferrugineus [2732, 2286]. fertilisation [2591]. Fertility [970, 805]. Fertilization [2822, 174, 779, 168, 2206, 99, 829, 1456, 181, 2106, 1921, 2756, 2891]. Fertilized [1393]. fertilizers [2110]. Fibre [1243]. fibropapillomas [1666]. Fidelity [1356]. Field [73, 801, 1550, 1551, 1056, 1192, 110, 1137, 222, 1417, 2822, 1556, 1967, 2887, 2216, 2238, 1839, 2611, 1526, 1906, 1820, 1620, 2436, 2743]. Field-Caught [1192]. field-collected [2216]. Filamentous [46, 1730, 1729]. Filter [276, 918, 917, 945, 2810, 2003]. Filter-Feeding [276, 918, 917, 945]. Filtering [55, 980, 2103]. Filters [2]. Filtration [1782, 1648, 2729, 1486]. fimbria [175, 653]. Fin [1226, 535, 1336, 356, 2854]. Final [442]. finding [1549]. Fine [2406, 154, 2510, 2738]. fine-resolution [2738]. Fine-scale [2406, 2510]. Fingerling [960]. Fingerlings [900]. Fingerprinting [1358, 1474, 1473, 1040, 1814]. Finland [2081, 969, 2135]. finmarchicus [316, 2215, 2905]. Finned [208, 2092, 2070]. Finnish [1031]. Fires [638, 1435]. First [1352, 1355, 499, 1097, 1454, 1069, 1351, 704, 2828, 1221, 1411, 961, 618, 1777, 2239, 2908, 2498, 2304, 1375]. First-Feeding [499, 1069]. First-Time [1351]. First-Year [704, 1411]. Fish [343, 810, 571, 353, 1174, 318, 1094, 1115, 735, 634, 683, 976, 632, 1082, 1231, 38, 1252, 737, 360, 669, 77, 708, 1318, 442, 1145, 707, 779, 872, 1032, 291, 217, 876, 198, 861, 1416, 813, 1219, 1399, 1449, 1468, 791, 465, 449, 404, 539, 71, 402, 1072, 1232, 385, 825, 306, 196, 561, 1150, 436, 871, 1421, 1185, 448, 147, 546, 238, 72, 567, 1009, 91, 673, 857, 1273, 1299, 731, 848, 1446, 1008, 1237, 1345, 300, 1216, 1042, 1043, 1335, 646, 1191, 339, 489, 869, 1105, 115, 1274, 1195, 433, 176, 1153, 1154, 178, 107, 466]. Fish [1543, 227, 965, 18, 426, 684, 1337, 1143, 936, 374, 1402, 992, 1415, 1383, 3, 1077, 337, 352, 1000, 744, 1126, 891, 2419, 1266, 972, 649, 483, 1104, 1163, 781, 213, 544, 555, 1398, 581, 2005, 2034, 2386, 1811, 1573, 1531, 1870, 1525, 1514, 1556, 2109, 2736, 2735, 2636, 2264, 2448, 2637, 2298, 1816, 2159, 2462, 2898, 1859, 2423, 2150, 2655, 1886, 1694, 2106, 2238, 2232, 2599, 2262, 1922, 2365, 2716, 2792, 1953, 1667, 2179, 1804, 1817, 2723, 2644, 2098, 1487, 2654, 1695, 1488, 2154, 2284, 2583, 2081, 1553, 2598, 2182, 2567, 1999, 2037, 2549, 2889, 2234, 2371]. fish [1907, 2420, 2064, 1746, 2886, 1881, 2359, 1733, 2536, 2725, 2314, 1864, 2288, 2604, 2613, 2676, 1678, 1593, 1621, 2702, 2467, 2515, 2240, 1880, 2649, 1674, 2509, 2554, 2553, 1947, 2433, 1879, 2335, 2431, 2229, 2900, 2825, 2020, 2002, 1964, 1942, 1908, 2811, 1649, 2441, 1622, 2200, 1897, 2086, 2666, 2569, 2778, 1906, 1820, 2603, 2657, 1940, 2068, 1628, 2417, 2477, 1858, 2285, 2004, 2060, 1483, 1878, 2606, 2748, 1825, 1506, 2084, 2191, 2574, 2211, 1685, 2203, 1883, 2510, 1736, 1901, 1910, 2220, 1623, 1737, 1913]. fish-amphipod [2723]. Fish-Size [1126]. Fished [213]. Fisher [2124, 221, 2698, 2127, 2126]. Fisheries [1086, 179, 510, 739, 462, 1709, 1315, 1304, 908, 1079, 1242, 2232, 1090, 699, 361, 1710, 2234, 607, 1175, 212, 833, 1375, 2852, 1, 1400, 1425, 374, 1433, 1430, 1330, 1236, 1708, 28, 1308, 441, 558, 883, 2101, 2439, 2590, 1578, 2807, 2360, 2397, 2542, 2601, 2181, 2260, 1510, 1930, 2660, 2124, 2779, 2685, 1861, 2359, 2207, 2310, 2650, 2745, 1631, 2726, 2652, 2541, 2414, 2361, 2606, 1798, 2653]. Fishermen [129, 1522]. Fishery [137, 15, 672, 901, 498, 628, 809, 72, 584, 1249, 1251, 248, 1367, 603, 1101, 1137, 923, 721, 1162, 111, 234, 2780, 1552, 2406, 2055, 2698, 2688, 1822, 2108, 1698, 2398, 2551, 2437, 1915, 2671, 1883, 2000, 2717, 2319]. Fishes [1325, 754, 801, 1360, 375, 1089, 1448, 288, 1347, 727, 1108, 661, 545, 1223, 381, 446, 1092, 812, 2787, 1659, 1671, 1827, 1813, 1664, 2687, 1656, 2150, 1501, 2067, 2186, 2888, 2272, 2523, 2475, 1527, 1876, 2116, 2557, 1496, 1624, 1663, 2474, 2205, 1685, 2153, 2317, 2695, 2091, 2682, 2220, 1952]. Fishing [1348, 210, 716, 715, 1059, 863, 363, 1430, 244, 663, 1841, 2590, 2406, 2393, 2143, 2407, 2533, 1888, 1887, 1861, 1904, 2551, 2845, 2513, 2857, 2742]. Fishless [1209, 1765, 2823]. fishway [2224]. Fitness [865, 528]. Fitting [285, 1331, 1346, 1147]. Five [392, 1416, 2177, 2664, 1910]. Fixation [174, 1391, 1390, 1498, 1485]. Fixed [594, 1000, 1795]. Fixed-Location [1000]. Fjord [1471, 546, 2818, 2763, 2868, 2869, 1447, 2869]. flagellate [2313]. Flatfish [626]. flatfishes [2447]. flavescens [220, 921, 2766, 2364, 2392, 101, 2430, 48, 106, 2384, 358, 843, 202, 351, 63, 1411, 2169, 1872, 2870, 2808, 502, 2280, 1068, 948, 1935, 784, 2775, 2776]. flavidus [907]. Flavor [61, 836]. Fleet [129, 233, 2780, 2685]. flemingeri [614]. Flemish [1932, 2200]. Flesh [474]. fleuve [719, 2884]. Flexibility [1054]. Floating [2761, 2753]. Flood [761, 705, 2888, 2871, 1723, 2868, 2869].

Flood-Control [761]. flood-pulse [2888]. flooded [1754]. Flooding [52, 550, 1753]. Floods [1170]. Florida [1870, 412, 897, 2716, 1045, 1873, 2122, 637, 238]. Flounder [1384, 598, 1228, 1142, 1254, 494, 775, 811, 1298, 2732, 2404, 1984, 2286, 1774, 1894, 1960]. Flow [1125, 1238, 806, 189, 301, 1050, 424, 844, 1222, 235, 1561, 2473, 2120, 2223, 2734]. flow-through [2120]. flowing [2037]. Flows [685, 2315]. Fluctuating [1428, 769, 184]. fluctuation [2189]. Fluctuations [1070, 969, 588, 355, 2376, 2805]. Fluid [951, 2572]. fluminea [104, 2543, 1589, 1524]. Fluorescence [195]. Fluoride [163, 1747, 1516]. flushing [2351]. Flutter [535]. Fluvial [1365, 282, 2615]. fluviatilis [342, 416, 2242, 2469, 278, 2135]. Flux [1183, 1167, 615, 2853, 1622, 2053]. Fluxes [429, 1267]. Fly [229, 1323, 1307]. Foci [481]. focus [2367]. Focusing [486, 981]. Folding [1341]. Following [158, 301, 1435, 2865, 2410, 1921, 2715, 2598, 1361, 468, 2025, 2666, 870, 746, 1739, 1166, 1307, 1917]. fonctionnel [1852]. fontaine [534]. fontinalis [667, 2138, 639, 803, 179, 1498, 1967, 2155, 853, 2225, 1646, 1645, 1561, 534, 479, 2620, 1511, 1324, 105, 445, 1056, 186, 180, 183, 2063, 1837, 384, 158, 682, 254, 253, 225, 528, 2467, 513, 1190, 1564, 122, 184, 514, 1532, 185, 967, 1063, 1718, 376, 2743, 1239, 511, 2270, 1307, 181, 182]. Food [1128, 1140, 756, 1201, 1074, 718, 331, 1295, 460, 492, 880, 1155, 711, 1369, 1142, 1317, 747, 794, 1216, 2400, 924, 643, 227, 289, 2858, 1755, 541, 366, 1104, 1368, 817, 1051, 978, 931, 190, 2499, 1650, 2136, 2029, 2680, 2141, 2213, 2896, 2643, 2064, 1951, 2876, 1768, 1985, 2426, 1576, 1928, 2793, 2198, 2906, 2164, 1906, 1820, 2417, 2768, 1890]. food-chain [1906, 1820]. Food-Deprived [817, 1051]. Food-Limited [1142]. Food-web [1755]. Foods [1206]. Foraging [853, 1252, 662, 291, 105, 889, 116, 636, 307, 1054, 1190, 953, 1967, 1758, 1637, 1680, 1743, 2879, 2518, 2782, 2515, 2400, 2506, 2416, 2226, 2880, 2354, 2695]. Forced [814, 1344]. forces [2086]. Forcing [723, 2264, 2662, 2145]. forecast [1840]. Forecasting [1053, 264, 2898]. Forecasts [1101, 1844]. foreign [1557]. Forest [638, 1435, 803, 1169, 300, 924, 1268, 1028, 2692, 2452]. Forested [634, 133, 837, 2618]. Forests [565, 2214]. Fork [1707, 1705, 1704, 1702, 1706, 1703, 22]. Form [683, 952, 1568, 2355]. forma [604]. Formation [700, 569, 775, 2404, 2765, 1769, 2668]. Former [1427]. forming [1729]. Forms [1465, 291, 1184, 886]. formulation [2433]. Fossil [1394, 1403, 2455, 2445]. Fouling [865]. Founder [2010, 2183]. Four [740, 265, 42, 497, 2083, 1534, 2135, 2091, 1847]. Fourier [315, 1197]. Fox [883]. fraction [2865]. fractionation [2877, 1836]. Fractions [444, 2893]. Fragment [754, 2641, 1843]. fragmentation [2135]. Framework [1330, 1913]. franciscanus [2512]. Frank [2439]. Fraser [2248, 1275, 425, 1445, 1544, 2132, 1348, 1315, 15, 2204, 1767, 354, 1857, 1481, 778, 370, 2178, 2414, 1208, 2252, 2167, 2189]. Free [901, 1273, 625, 1363, 1707, 2896, 2393, 144, 2724, 2065]. free-living [2065]. Free-Ranging [625, 1707]. Free-swimming [1273]. Freeze [631, 1393].

French [325, 1852, 534, 1660, 1234, 719, 2884, 414, 926, 633, 2868, 2869]. Frequencies [487, 2014]. Frequency [320, 32, 349, 547, 705, 1436, 92, 529, 650, 111, 1205, 821, 1776, 2382, 2170, 2040, 2512, 2529, 2628]. fresh [1806]. fresh- [1806]. Freshwater [1325, 73, 1365, 425, 591, 669, 1230, 1012, 959, 1225, 680, 296, 560, 1452, 576, 1285, 33, 34, 1095, 1446, 758, 424, 746, 502, 308, 1146, 298, 1409, 1525, 1659, 2269, 1990, 1656, 1550, 1551, 2539, 2521, 2522, 2415, 2837, 2103, 2116, 1801, 2729, 2229, 1679, 2086, 2082, 1938, 2704, 2323, 2608, 1636, 2203, 1818, 2731, 2605]. Freshwater-Adapted [1409]. Freshwaters [165, 2825, 1543]. Frog [23]. Frond [610]. frondosa [1805]. front [2237]. frontinalis [2243]. fronts [2840]. FRS [1905]. Fry [415, 912, 180, 183, 53, 1453, 184, 846, 277, 2212, 1239, 1303, 1379, 1969, 1751, 1987, 2865, 2516, 1996, 1591]. Fry-to-adult [2212]. Fuel [1245, 2134]. fullness [2401]. fulvescens [1830, 2377, 2797, 1900, 787, 719, 939, 2832, 1131, 1532, 2158, 2224, 100]. fumigant [1601]. Function [2566, 565, 1272, 750, 678, 798, 1362, 759, 743, 686, 444, 950, 1137, 668, 1427, 1298, 333, 1671, 1827, 2598, 1493, 1734, 2137, 1651]. Functional [952, 661, 852, 290, 755, 1852, 1568, 2210]. Functions [708, 1331]. funduloides [423]. Fundulus [2344, 1721]. Fundy [1228, 2063, 64, 1915, 2671, 898, 1575, 240]. fungal [2073]. Fungicide [1342]. Furan [1318, 2793]. Further [568]. fused [1569]. Future [1423, 1746, 2560]. fuzzy [2724].

gadoids [2550]. Gadus [1491, 2760, 2057, 1140, 1997, 1263, 1670, 598, 862, 1358, 1748, 2077, 2844, 281, 1986, 1840, 1788, 2546, 1743, 698, 2466, 2839, 2011, 1038, 2397, 922, 631, 1444, 2855, 2219, 2518, 2519, 1839, 2785, 1134, 1133, 1376, 2816, 2669, 2422, 1681, 2525, 138, 548, 1351, 1846, 2718, 1317, 2327, 2662, 1249, 2572, 1606, 2237, 128, 2170, 2524, 1172, 2530, 1586, 1044, 968, 2866, 2100, 2099, 1078, 2144, 2329, 2863, 2891, 374, 594, 1101, 1848, 2056, 2358, 2565, 2806, 2076, 2617, 2096, 1005, 946, 2642, 805, 1688, 537, 856, 372, 1178, 1161]. gagei [1010]. Galápagos [579]. galaxias [1687]. galeata [41, 1102, 1917]. Galeorhinus [1962]. galeus [1962]. Galjoen [1286]. Galle [648]. Gambusia [1860, 51, 283, 1108, 1772]. gamete [1509, 2891]. Gametogenesis [141]. Gaming [1432]. Gammaridae [1145]. Gammarus [1365, 1074, 1107, 2823, 2794, 2848]. GAP [2145]. Gape [1252, 502]. Gape-dependent [1252]. Gas [216, 377, 2048, 2046, 2134, 2156]. gas-producing [2046]. gas-supersaturated [2156]. Gasbladder [356]. Gasterosteiformes [2841]. Gasterosteus [1607, 473, 1922, 2365]. Gastric [383]. gastrically [2465]. gastropod [2824]. Gastropoda [194, 2162]. Gastrosteus [139]. Gauge [1076]. Gauntlet [59]. gazella [2472]. Gear [2530, 78, 594, 1427, 2078]. Gear-independent [2530]. gemfish [1684]. Gender [1206]. Gene [281, 1222, 234, 1470, 1833, 2514, 2894, 1763, 2734]. Genera [450]. General [1186, 134, 1330, 2606, 1972]. Generality [635, 1755]. Generalized [1198, 722, 2648, 2362, 2412, 1864]. Generating [162]. generation [1828]. Generations [480]. Genes [317, 470, 1058]. Genetic [656, 994, 1813, 737, 822, 2421, 1466, 464, 1511, 582, 361, 2623, 69, 1339, 2265, 912, 1843, 2272, 1868, 385, 766, 191, 776, 1471, 1001, 1457, 2821, 2490, 528, 2115, 1563, 833, 736, 2244, 716, 715, 1772, 530, 1237, 1478, 2245, 1078, 1458, 1848, 2358, 1989, 1966, 1451, 70, 850, 484, 1068, 1284, 1311, 1463, 1462, 886, 602, 1480, 441, 1459, 1460, 250, 1461, 1778, 1607, 1823, 2268, 1694, 2542, 2332, 2792, 2882, 2010, 2183, 2595, 2464, 1881, 1547, 2133, 1880, 1691, 1879, 1791, 2016, 2734, 2056, 2390, 2503, 1869, 1878, 1884, 1917]. Genetically [347, 148, 2271]. Genetics [854, 1449, 1981, 2063, 2083]. Geneva [2798]. genome [1498, 1557]. genomic [1813]. Genotype [586, 587, 2595, 1860]. Genotypes [586, 587, 1764]. Genotypic [81]. Genus [1476, 1467, 2894, 2276, 2539, 2282, 1677]. Geochemical [2044]. Geochemistry [785, 894]. Geographic [1061, 1133, 932, 345, 1241, 2378, 946, 324, 1590, 1580, 1790]. Geographical [1670]. Geology [453, 686, 1588]. geometry [2038]. Geophysical [720, 721]. geoposition [2786]. George [1384, 747, 1815, 321]. Georges [40, 2662, 928, 627, 2383, 121, 748]. Georgia [260, 2189, 1445, 2708, 2751, 1029, 2289]. Georgian [1112, 2256, 2460]. geostatistical [1934, 2772]. Geostatistics [1312, 2746]. germ [2677]. germline [1613]. Geryonidae [236]. Get [1378, 642]. Giant [916, 1749, 1921, 1554]. gibberum [55, 42]. gibbosus [1213, 485, 1054]. gigantea [303]. gigas [141, 325, 1322, 2490]. Gill [407, 271, 490, 557, 977, 430, 938, 988, 1154, 512, 1218, 1409, 2570, 1761, 1888, 1887, 2422, 1593, 1592, 1589, 1915, 2032, 514]. gill-net [1915]. Gillis [1881]. Gillnet [740, 741, 2337, 1487, 336, 557, 2785, 1101, 50, 2520, 1568, 2671]. Gills [273, 930, 176, 1153, 1154, 399, 1409, 2731]. gilt [1496]. gilt-head [1496]. gilthead [1739]. GIS [1280]. Gizzard [761, 1364, 1738, 1872]. Glacial [655, 1068, 2236, 2374]. Glaciations [611]. Glacier [1599]. gladius [2070, 1644]. gland [926]. glande [926]. global [2292, 2540, 2551, 2585, 2482]. globally [2284, 1858, 2285]. Globicephala [1157]. Glycogen [1221, 817, 1051]. Glycoprotein [1444]. Glyptocephalus [598]. Gmelin [802]. GnRH [2286]. GnRH-A [2286]. Go [1378]. gobies [1627]. Gobiidae [617]. goby [1829]. Going [2473]. Golden [662, 2812]. Goldfish [970, 1482]. golfe [414]. Gonad [797, 841]. Gonadotropin [2286]. Gonadotropin-releasing [2286]. gonads [1676, 2375]. GOOMEX [2043]. gorbuscha [2865, 2449, 2356, 480, 2595, 1453, 716, 715, 19, 2414, 1464, 1779, 978, 441, 1470]. Gorge [286]. governed [2892]. Governing [322]. government [2208]. Gradient [1167, 342, 290, 563, 821, 1550, 1551, 2469, 1539, 2731]. Gradients [782, 1405, 2005, 2564, 1553, 1625, 2574]. grading [1522]. Grand [626, 1277, 1227, 58, 2321, 432, 2531]. Grande [1754]. Grande-2 [1754]. grandis [2082, 375, 1638, 1550, 1551, 2704, 2731]. grands [2884]. Granite [495]. Graph [229]. Graphical [952]. grassland [2007]. Gray [2162].

Grayling [415, 685, 228, 792, 793, 1378, 282, 2889, 2754, 1689]. graze [2427]. Grazed [116]. Grazer [781, 2484, 2552]. Grazers [629]. Grazing [905, 55, 1209, 730, 136, 2665, 1924, 2890]. Great [487, 1116, 431, 1070, 2207, 1118, 461, 211, 2105, 2162, 2835, 1829, 2125, 439, 1996, 2691, 1408, 1936, 766, 617, 2403, 2507, 736, 1345, 2236, 2247, 2674, 1771, 2060, 726, 1310, 1401, 10, 201]. Green [1137, 966, 1666, 2505, 1926, 2287, 1730, 1729, 2264, 2079, 1420, 1585, 290]. Greenback [1142]. Greenland [1106, 1120, 1888, 1887, 2271]. grenadier [2788]. Grey [1061, 1777, 2058, 2428, 2069, 1937]. grilse [1614]. groenlandica [1020, 1019, 676, 1023, 531, 2069, 1130]. Ground [1262]. Groundfish [362, 427, 799, 1241, 1302, 863, 103, 722, 1552, 2831, 1698, 2139, 2717]. Grounds [56, 2378]. Groundwater [1646, 1645, 1555, 140, 2435]. groundwater-fed [2435]. Group [245, 985, 1311, 1491, 2095]. grouped [2742]. grouper [1676]. Groups [95, 815, 2590]. Growing [1093, 2404]. Growth [2728, 556, 1199, 1004, 1830, 104, 700, 1776, 565, 634, 1263, 346, 135, 670, 246, 732, 160, 708, 2413, 1297, 1071, 842, 245, 148, 1145, 1074, 1058, 291, 2839, 165, 189, 753, 214, 922, 32, 428, 48, 1529, 1331, 912, 243, 180, 596, 1123, 436, 1142, 919, 53, 99, 491, 1338, 1192, 1362, 1285, 1317, 144, 1802, 303, 451, 2754, 1850, 874, 175, 882, 1243, 110, 1216, 521, 1047, 1335, 968, 1215, 30, 202, 916, 1411, 1336, 36, 878, 1259, 1046, 1177, 1571, 1039, 668, 726, 3, 729, 171, 815, 494, 1364, 523, 150, 86, 21, 649, 277]. Growth [559, 1163, 1440, 1729, 624, 978, 74, 1379, 2091, 1217, 1865, 1278, 418, 250, 1319, 1189, 1301, 943, 1313, 406, 240, 2309, 2432, 2788, 2766, 1650, 325, 1770, 1602, 2066, 1637, 2912, 2494, 1748, 2077, 2684, 2328, 2546, 2092, 1613, 2061, 2680, 2466, 1629, 1489, 2615, 2199, 1984, 1900, 2504, 2140, 2814, 2630, 2877, 2545, 2705, 1839, 719, 2888, 1805, 1695, 2575, 2476, 1767, 2296, 2367, 2243, 2816, 2850, 2382, 2632, 1815, 1939, 2028, 2119, 2710, 1733, 1490, 2832, 2030, 2662, 2913, 2116, 2444, 2426, 1692, 2316, 2693, 2694, 2240, 2683, 1902, 2554, 2781, 1896, 2447, 2431, 2813]. growth [2841, 2424, 1962, 1964, 2866, 2453, 2070, 2646, 2757, 1940, 2774, 2697, 2846, 2848, 2626, 2076, 2651, 1521, 2293, 2512, 2353, 1890, 2858, 2593, 1524, 2368, 2280, 1971, 1688, 1658, 2341, 1766, 1853, 1612, 1505, 2425, 1835, 2667, 1941, 2214, 2730, 1952]. Growth/Mortality [874]. grunniens [591, 502]. Grunow [604]. grypus [1061, 1777, 2428, 2069, 1937]. GSI [441]. GT [1814]. Guest [410]. Guide [1101]. Gulf [556, 814, 172, 610, 1225, 1181, 287, 236, 606, 231, 714, 1103, 171, 1005, 946, 369, 1686, 2432, 156, 1874, 1019, 1670, 1983, 2844, 2048, 1212, 1918, 245, 90, 2839, 2046, 349, 1468, 851, 2428, 2013, 2616, 2382, 420, 2043, 2327, 414, 2676, 2069, 614, 2045, 1968, 2127, 1176, 2100, 2099, 2050, 517, 594, 356, 2126, 666, 2035, 92, 2205, 850, 1790, 2642]. gulls [2691]. gummy [1962, 2355]. gustatory [2661]. Gut [972, 1775, 2401, 2343]. Guts [1082]. guttatus [1177]. Gymnocephalus [342, 2104, 749]. Gymnophallidae [1665]. gynogenesis [1994]. Gypsum [424]. Gyrodactylus [2273, 2436].

H [1726, 2719]. Habitat [735, 375, 1228, 2377, 291, 906, 647, 1097, 2080, 310, 437, 1232, 1292, 1123, 1054, 727, 2702, 733, 545, 659, 658, 924, 400, 1223, 433, 1053, 1100, 2767, 98, 190, 1874, 1854, 1638, 2784, 2799, 2242, 2615, 2620, 2088, 2714, 2896, 2434, 2085, 1720, 2221, 2812, 2297, 2030, 2604, 2130, 2796, 1579, 1630, 2491, 2627, 2350, 1940, 1484, 1808, 1483, 2642, 1688, 1736, 2809, 1923]. habitat-based [2627]. Habitat-dependent [2080]. Habitats [710, 1092, 2737, 2264, 2784, 1694, 2530, 2335, 2668]. Habits [531, 366]. Haddock [573, 598, 40, 627, 849, 2912, 2011, 2662, 2913, 1579, 2693, 2694, 2895, 2383]. Haematocrit [1257]. haemoflagellate [1655]. Haemoglobin [951]. Haemohormidium [1257]. Haemorrhagic [417]. Hake [498, 522, 244, 650, 755, 525, 74, 2406, 2688, 1580, 2476, 2616, 2299]. hakes [1683]. Halibut [1106, 1352, 1333, 654, 927, 1312, 1198, 1141, 2684, 2041, 2399, 2271]. Halichoerus [1061, 1777, 2428, 2069, 1937]. Halifax [549]. Haliotis [2822]. Hamilton [2036]. Handling [559, 817, 1959]. Haplotype [1205, 2014]. Haplotypes [1477]. Harbour [1181, 114, 64, 2036, 558, 240, 1598, 1806, 1915, 2671, 1949, 1584, 549]. Hard [854, 1002, 1761, 2347]. Hard-part [1002]. Hardness [1296, 1713, 2296]. hardwater [1847]. harengus [172, 168, 167, 266, 56, 643, 551, 121, 216, 264, 50, 797, 172, 2861, 2031, 578, 214, 1487, 1471, 1073, 2478, 1934, 2073, 121, 1162, 2529, 50, 1278]. Harmful [783, 260, 1410, 2018]. Harp [1020, 1019, 676, 1023, 531, 1130, 1529, 2069]. Harvest [1089, 15, 1085, 1224, 963, 673, 1429, 1431, 1088, 341, 2019, 2584, 1510, 2681, 2201]. Harvesting [610, 1255, 769, 1059, 68, 2124, 2084]. Hasle [580, 604]. hastata [303]. hatch [2732]. hatched [1586]. Hatchery [1226, 1382, 81, 673, 552, 1478, 164, 767, 1381, 1380, 308, 65, 484, 996, 1135, 406, 1751, 1987, 2737, 1958, 2448, 2459, 2715, 2509, 2440, 1989, 2211, 2212, 1515, 1597]. Hatchery-Origin [552]. hatchery-produced [2212]. Hatchery-Reared [1381, 1380, 1382, 2448, 2440, 2211, 1515]. Hatching [805, 1833, 1527, 1723, 2329, 1766]. Haul [751]. Hawaiian [2091]. Hawks [1275]. Haynes [2658, 321]. head [2840, 1496]. headwater [1745]. Health [2109, 2005, 2447]. Heaphy [1905]. heart [2678]. Heartworm [2166]. Heat [1207, 1997, 1994, 1593]. Heavy [757, 1334, 887, 1206, 2364, 1700, 1714, 2457]. Hecate [155, 1302, 537]. Held [560, 270]. Helminth [265, 2163]. help [2650]. helvomaculatus [2734]. Hemolymph [750]. hemorrhagic [2619]. Henry [1954, 22]. Hepatic [1075, 2598, 743, 481, 817, 1051, 333, 1610, 2017, 1960]. Hepatocytes [118, 96, 420]. Hepatopancreas [407]. Hepatotoxic [467, 2263]. hepatotoxins [2082]. herbicide [2110]. Herbivore [136, 2468, 1553]. Herbivory [1202, 506, 2346, 2012]. Hermaphroditic [1191]. herpetofaunal [2725]. Herring [346, 797, 334, 172, 578, 214, 168, 167, 266, 1471, 1073, 677, 56, 643, 551, 1037, 121, 1162, 804, 974, 216, 264, 50, 1278, 1445, 1544, 2021, 2619, 2861, 2502, 2031, 1487, 2691, 2023, 2024, 2026, 2478, 1934, 2274, 2700, 882, 2022, 2025, 2073, 2726, 2767, 1862, 2368, 2529]. herrings [2429]. Hesperodiaptomus [1413, 1765]. heteroclitus [2344, 1721]. Heterocyst [1391]. Heterogeneity [735, 1477, 2799, 2490, 2798]. heterogeneous [2795]. Heterosigma [2313]. Heterotrophic [117, 1386]. Heterozygosity [225, 1048]. Hexachlorobiphenyl [2719, 313]. Hexagenia [280, 1941]. hexapterus [1076]. Hg [2884]. hierarchic [2288]. Hierarchies [792, 793]. High [675, 148, 2620, 1338, 1138, 820, 1328, 1410, 1522, 1920, 1596, 2286, 1934, 2881, 2838, 2040, 2625, 2302, 1808, 2529, 1835, 2187, 154]. high-elevation [1920]. high-frequency [2529]. high-grading [1522]. High-Mountain [675]. High-Performance [1138]. high-quality [2286]. Higher [257, 1551]. highland [1832]. Highly [1240, 2643]. Hind [1177]. Hippoglossoides [911, 1670, 1406, 1257, 850, 1790, 2197, 2642, 1106, 2271]. Hippoglossus [1352, 2684, 1333, 2041, 654, 927, 1312, 1198, 1141]. hippurus [1959]. Hirudinea [1321]. hirundinella [648]. hispida [2333, 27, 562, 2174, 1762]. Histological [168, 2104]. Histology [271, 514]. histopathological [2469, 2023]. Histopathology [674, 2274, 2700, 2026]. Historic [762, 2318, 1702]. Historical [1299, 1507, 455, 1621, 1626]. Histories [729, 242, 2716, 1744]. History [362, 822, 959, 485, 1479, 1389, 860, 283, 339, 1025, 974, 1184, 812, 2021, 2792, 1695, 2272, 2085, 2451, 1867, 2185, 2420, 1738, 2217, 2813, 2127, 2491, 2733, 1649, 1293, 2903, 2907, 2126, 2606, 1966, 2211, 2899, 2582]. hoferi [2073]. hoki [1883, 2000, 2555]. holbrooki [1860, 283, 1108, 1772]. holding [1787]. Holopedium [55, 2282, 42]. Holothuroidea [1805]. Holt [1867]. homari [302]. Homarus [2142, 1773, 137, 520, 302, 1098, 1341, 868, 2265, 1292, 2632, 1192, 734, 505, 2848, 267, 693, 2892]. Homatus [58]. Home [384, 1624]. Homeostasis [834]. Homing [205, 1534, 1222, 996, 2443]. homogeneous [2271]. homology [2244]. Hooded [851]. Hook [78, 923, 721, 1568]. Hook-and-Line [721]. Hooks [952]. Hoplostethus [672, 87, 1521, 1048]. Horizontal [1073, 221, 2755]. Horizontally [1029]. Hormonal [1051]. Hormone [148, 1058, 306, 270, 523, 1163, 624, 1939, 2028, 2286, 1902, 2593, 1610]. Hormones [271, 1221, 624]. Host [954, 630, 1107, 1549, 2422, 1679]. host-finding [1549]. Hoxie [286]. hoyi [780]. HPLC [993, 2408]. HPLC-PDA [2408]. huîtres [325]. hubbsi [1683]. Hudson [1019, 665, 636, 1408, 1936, 312, 54, 2687, 1741, 1075, 2283, 2656, 2500, 1802, 2174, 687, 1083, 1899]. hudsonius [459]. Human [611, 2356]. Humic [1296, 1155, 1414, 1950, 2850, 2643]. humoral [1739]. huntsmani [527]. Huron [613, 2460, 2277, 2587]. Hyalella [1713, 398, 1567, 1566, 2194, 2193, 2803, 576, 2008]. Hybrid [254, 253, 1081]. Hybridization [445, 2294, 919, 786, 1102, 1151, 1084, 2118]. Hybrids [1452, 480, 1454, 1453, 250, 406, 1933, 2600]. hydraulics [2434]. Hydroacoustic [810, 390, 1000, 1531, 1514, 2637, 1487, 2772]. Hydroacoustics [314]. Hydrobiidae [2162]. Hydrocarbon [549, 2865, 2469, 2050]. hydrocarbons [2024, 2274, 1774]. Hydrochemical [1726]. hydrodynamic [2038]. Hydroelectric [894, 1131, 1720, 1754, 2165]. Hydrogen [102, 1769, 2353]. hydrogeology [1647]. Hydrographic [723, 231, 171, 1932]. hydrologic [1745]. Hydrological [1714, 2579, 2158]. Hydrologically [503]. Hydrolyzable [447]. Hydropsychid [600, 2860]. Hydropsychidae [600]. hydroxylation [2797]. hyperbenthic [2163]. hypereutrophic [1582, 2886, 2403]. hypermineralised [2355]. hyperosmoregulatory [2326]. Hyperoxia [408]. Hypertrophic [117, 1386]. Hypervariable [1474]. hypochlorite [1855]. Hypolimnetic [397, 2309, 2308, 2306, 2307, 2305, 1877, 2152, 2151, 1595, 2302, 2445, 2304]. hypolimnion [1731, 2487, 1594]. Hypoosmoregulatory [597]. Hypoptychus [2841]. Hyporheic [133, 469, 2435, 913, 235, 2706, 2202]. hyposaline [1582, 1959]. hypothèse [1852]. Hypotheses [1370, 68, 2521, 1882, 2726, 1559]. Hypothesis [1348, 1105, 1852, 1823, 2753, 2490, 1755, 299]. Hypoxanthine [36]. Hypoxia [1195, 2543, 2592, 1941].

IBI [1345]. Ice [223, 675, 159, 90, 192, 1408, 295, 531, 8, 154, 856, 554, 2370, 2668]. Ice-Covered [675, 2370]. Iceland [1874, 1614, 1001]. Icelandic [2882, 2464, 2524, 2096]. ICES [569]. Ichthyobodo [1658]. Ichthyomyzon [1010]. Ichthyophonus [2073]. Ichthyoplankton [155, 2610, 2611, 2157]. ICPMS [1358, 1617, 1618]. Ictalurus [487, 836, 1282, 61, 88, 1926]. Idaho [22, 1100]. Idea [540]. Ideal [2896, 901, 2393]. Identification [317, 2194, 1671, 737, 1358, 69, 736, 530, 178, 1309, 992, 1310, 1305, 1135, 1818, 441, 1460, 1475, 1819, 1958, 1590, 2862, 1799, 1891, 1827]. Identify [1412, 1764]. Identifying [1223]. identity [2560]. Idiosepius [866]. II [1650, 587, 1341, 1551, 2307, 2611, 7, 1122, 514, 2776]. III [95]. Iliamna [2539]. illecebrosus [2092, 2070]. Illex [2092, 2070]. Illustrated [32]. Image [965]. imaging [2040]. Immature [664, 579, 1775]. immigration [2575]. Immune [870, 950, 1739]. Impact [2308, 1601, 2687, 610, 757, 159, 2306, 2307, 703, 2692, 2376, 1332, 345, 253, 778, 443, 1759, 640, 438, 808, 493, 209, 2090, 2569, 1927, 935, 2370, 1640, 2058, 2073, 526, 1620, 2560, 2551, 734, 1808, 1989]. Impacts [1228, 980, 2807, 2711, 1538, 1045, 2476, 95, 2110, 2602, 2868, 2869]. impair [2143]. Impaired [2364]. impairment [1707]. impairs [1955]. implanted [2465]. implementation [2663]. implication [1784]. Implications [474, 1650, 1167, 126, 1252, 1117, 762, 699, 1522, 2393, 2049, 1421, 282, 144, 432, 2694, 736, 305, 1176, 1373, 1400, 36, 363, 1087, 1288, 777, 111, 2810, 1748, 2077, 2342, 2853, 2259, 2839, 2290, 2356, 2723, 1695, 2313, 1836, 2103, 2384, 2064, 2834, 2426, 1962, 1528, 2035, 1533, 2076, 2089, 1891, 812, 1503]. Importance [412, 285, 38, 1230, 456, 1111, 37, 251, 243, 918, 1429, 529, 1268, 1158, 1132, 2893, 1822, 2107, 2768, 1924]. important [2171]. Impoundment [1039, 2410]. imprinting [1824]. Improve [1]. Improved [697, 943, 2788, 2127, 2126, 2425]. Improvement [659]. Improvements [468, 1320, 2563, 2562]. improves [1750]. Improving [2362, 1945, 2524, 2042]. in-river [1514]. In-season [1918, 2528]. In-stream [1050]. inbreeding [1860]. incidence [1809]. Incidental [1915, 558]. include [892]. Including [963, 1456, 11, 2524, 2681]. incompatible [2208]. Incomplete [1383, 70, 2534]. inconnu [2847]. Incorporates [21]. incorporating [2533, 1796]. Incorporation [1797, 1945]. Increase [571, 2678, 558, 1604]. Increased [2501, 2305, 871, 99, 417, 1650, 2410, 1613, 2401, 2351]. Increases [296, 307, 519, 1561, 1594]. Increment [2404, 1423, 494, 775, 1002, 1997]. Incremental [1125, 1871]. Increments [1142, 1335, 684, 1962, 1964]. Incubation [2526, 536, 2763, 1646]. independent [2530]. Index [2909, 501, 757, 1030, 217, 1269, 1120, 789, 203, 1345, 504, 471, 2231, 1895, 1896]. Index-Removal [1030]. Indexing [413]. India [268]. Indicate [297, 1644]. Indicated [1334, 916, 28, 1684]. Indicating [405]. Indicator [306, 887, 2564, 237, 1963, 2854]. Indicators [1349, 539, 1143, 388, 1870, 1681, 2536, 1682, 2246, 2455, 1832, 1574, 2776]. Indices [848, 594, 1749, 2549, 2796, 1564]. Indies [1627]. Indigenous [282, 1989]. Indirect [2141, 1008, 230, 367, 50, 1307, 1859, 1765]. Individual [2248, 628, 2545, 402, 824, 802, 451, 875, 677, 852, 878, 2870, 729, 1750, 2853, 2716, 2124, 2705, 1846, 2026, 2314, 1864, 1606, 2613, 1586, 2424, 2506, 2811, 1894, 1506, 2191, 2399, 2167, 1612, 2628, 2425]. Individual-Based [451, 875, 878, 852, 2870, 2853, 2314, 1864, 2613, 1586, 2424, 2811, 1894, 2167]. individually [1940]. Individuals [1329, 1240, 1758, 2774, 2846]. Induced [1199, 104, 1026, 283, 1254, 1363, 1195, 1141, 1650, 1970, 2298, 1996, 1513, 2074, 2422, 2137, 2073, 2907, 2870]. Induces [1456, 2286, 2750]. inducible [2586, 2418]. Induction [1333, 481, 1903, 2598, 1994]. inductively [2298, 2854]. Industrial [855, 945]. industries [2857]. industry [2127, 2126]. infaunal [2045]. infected [2547, 2592, 1658]. infecting [1672]. Infection [862, 954, 2750, 1257, 255, 2422, 1674]. Infections [1141, 2333, 2470, 2166]. Inference [153, 2807, 1943]. Inferences [878, 2688, 2422, 2613, 2445]. Inferred [1466, 113, 1299, 1099, 828, 1467, 2894, 1843, 2351, 2239, 2895]. Inferring [650, 2903]. Infestation [345, 1387, 1388, 2129, 1620]. infested [2066]. infinite [2228]. Inflate [795]. Influence [129, 1325, 2867, 1713, 415, 958, 611, 669, 1775, 806, 685, 2087, 101, 1071, 2312, 959, 1225, 1074, 162, 2365, 189, 995, 1334, 592, 2698, 265, 1296, 1747, 933, 2575, 2871, 266, 723, 1369, 1347, 2816, 2632, 2643, 818, 836, 600, 1054, 1362, 2511, 2800, 1285, 1287, 34, 1073, 1247, 2703, 430, 518, 2467, 653, 1947, 2250, 1801, 1576, 953, 572, 1548, 646, 924, 627, 2790, 140, 1402, 1364, 2368, 1063, 218, 1208, 235, 1685, 1717, 1146, 2304, 982, 2537, 2815, 292, 1132, 1565, 2136, 2732, 1700, 1647, 2272, 2610, 2611, 1746, 2718, 1850, 1821, 1893, 1908]. influence [1940, 2210, 2556, 1712, 1911]. influenced [2346, 2349]. Influences [2636, 877, 449, 2665, 2128, 2701, 2218, 2872, 2764, 1761, 1844, 2878]. Influencing [1365, 897, 1039, 2366, 238]. Information [551, 93, 2208].

Infrared [1062]. ingens [1123]. Ingestion [1323, 518, 1967, 2912, 1643, 2913]. Inhabited [300]. Inhabiting [1469, 601]. Inheritance [903]. Inhibition [970]. Inhibitors [970]. Initiatives [461]. Injury [1702]. Inland [7, 350, 1280, 6, 95, 780, 618, 2145, 2606]. Inlet [2074, 554, 2856]. innate [1655]. Inner [320]. Inorganic [159, 490, 519, 185, 728, 940, 324, 2138, 2756, 2426]. inquiry [2232, 2234, 2208]. Insect [1163, 2768, 2165]. Insecticide [229, 158, 1265, 2452]. Insects [807, 768, 450, 519, 1775, 2594]. Insemination [267]. Inshore [1444, 1172, 734, 1101, 1491, 1840, 1848, 2358, 2617]. Insights [956, 2499, 2873, 2641, 1596, 2343]. Institutional [695]. Instream [1125, 98]. Instrumental [181]. Instruments [904]. intake [2592]. integer [2877]. integerrimum [1662]. Integrated [972, 2535]. integrating [2735]. Integration [426, 1607]. Integrity [1269, 1345, 2549]. intense [2143]. Intensely [116]. intensity [2577, 1620]. Intensively [213]. Interaction [2732, 449, 912, 476, 51, 290, 2342, 130, 1580, 2595, 2901]. Interactions [2706, 707, 1810, 2705, 702, 816, 321, 1446, 1176, 881, 202, 603, 2646, 512, 867, 374, 28, 541, 1011, 376, 1770, 2392, 2723, 2518, 2080, 2782, 1512, 2491, 2808, 2890, 2241, 1699, 2574, 1972]. Interactive [1007, 953, 506, 1944]. Interannual [2675, 1445, 1033, 243, 354, 1247, 1298, 2007, 2614, 1489, 2434, 2693, 2229]. Interbasin [140]. Interbreeding [307]. Intercalibration [2652]. Interdecadal [2573, 2614]. Interdependence [908]. Interface [847, 2320]. Interference [119, 2393]. Interindividual [2725]. Interior [228, 793, 2889]. Interlake [2424]. Intermediate [1317, 630, 1107]. Intermittent [1336, 2706, 1482, 2353]. Intermolt [267]. Intermontane [304]. Internal [305, 2278, 2578, 2302, 2351, 2239]. International [2589, 1400, 2541, 1085, 569]. interoceanic [1598]. Interplay [2860, 2233, 2234, 2402]. Interpretable [160]. Interpretation [37, 599, 1241, 494, 2393, 2049, 2321]. interpretations [2125]. Interpreting [2399]. Interrelation [649]. interrelationships [2545]. interrenal [1139, 2430]. Interspecific [408, 711, 253, 1102]. interspersed [2571]. interstrain [2381]. Intertidal [1228, 906, 1097]. intervals [2293, 2042]. Intervention [611]. interviewing [2852]. Intestinal [1352, 1532]. Intra [408, 253]. Intra- [408, 253]. intraannual [1489]. Intrabiome [239]. Intrapopulation [19, 2293]. Intraspecific [281, 2133, 2824, 406, 2808]. Intrastrain [2381]. Intravascular [88]. introduce [1557]. Introduced [1322, 865, 749, 2636, 2216, 1998, 1868, 2580, 1687, 493, 1771, 1736]. Introduction [2021, 452, 1085, 839, 1449, 1434, 2301, 2145, 1725, 1418, 269, 2225, 2043, 2442]. Introgression [1498, 2623]. invaded [1928]. Invader [766, 618, 2105]. invaders [2560]. Invading [1008]. invariants [1867]. Invasion [1117, 1495, 2659, 466, 2175, 2723, 1927]. Invasive [736, 2885]. Inventory [461]. Inverse [1368, 1627, 2424]. Invertebrate [371, 683, 1231, 920, 1145, 1169, 1842, 1253, 644, 2106, 1714, 1944, 2493, 2210].

Invertebrates [1230, 1150, 679, 550, 503, 151, 633, 1693, 2090, 1616, 1533, 2379, 2214, 2220, 1703]. Investigate [995]. investigating [2879, 2226, 2880]. investigation [2324]. Involve [1139]. Involved [421]. Involving [1375, 1836]. Ion [835, 957, 521, 1239, 2787, 2053]. ion-exchange [2787]. Ionic [1070, 1982]. ionization [2787]. Ions [181, 182, 1988]. Ireland [1981, 1635]. Irish [610, 2421]. Iron [706, 145, 1194, 207, 1057, 1650, 2850, 1595]. Irradïance [411]. irradiation [1994]. ischnus [2105]. ishikawai [2121]. Island [665, 1055, 636, 1657, 82, 403, 774, 2235, 1337, 1475, 210, 2299, 2902, 244, 525, 2368, 663, 1014]. islandica [1874, 1001]. Islands [90, 1426, 777, 1683, 579]. Isoelectric [486]. Isolated [118, 1476, 288, 323, 230, 891, 2094, 1825]. isolating [1963]. Isolation [464, 1333, 1869, 2604]. Isomers [330]. isopod [1992, 1669]. Isopoda [1992]. Isotope [1169, 2499, 2874, 1978, 1993, 2873, 1545, 1979, 2877, 1873, 2067, 2691, 1596, 2647, 1733, 2536, 2878, 2577, 2670, 2565, 1806, 2479, 2203]. Isotopes [539, 1797, 2872, 1768, 1928, 2906, 2343]. Isotopic [2862, 1295, 1654, 1836, 2875]. israelensis [1193, 913, 1307]. Issues [1086, 1089, 2588]. Isurus [1843]. ITS1 [2804, 2239]. IV [350].

Jackass [1271]. Jacobson [2916]. James [939, 33]. Japan [2221, 1668, 769, 1679, 1663, 823]. Japanese [994, 1581, 2752, 1214, 2634, 1634, 1784]. japonica [1776, 1581, 1148]. japonicus [769]. jaune [719]. Jaw [1023]. jellyfish [1922, 2365]. Jersey [473]. jet [1492]. Joaquin [1565, 2493]. John [2827, 1905]. jordani [933, 1715, 1580]. jubatus [2222]. juillet [2868, 2869]. July [2013, 2868, 2869, 663]. Junction [515]. Juvenile [4, 5, 1199, 1244, 853, 1382, 756, 980, 822, 1071, 473, 94, 245, 899, 819, 983, 999, 215, 922, 800, 631, 197, 890, 889, 287, 809, 941, 1857, 2221, 1374, 1142, 1361, 860, 307, 99, 1342, 204, 778, 930, 796, 513, 1172, 1447, 521, 514, 1042, 1043, 659, 658, 1078, 852, 1281, 152, 1221, 1293, 63, 1336, 512, 1100, 1046, 383, 950, 1275, 98, 1327, 232, 1381, 1380, 815, 657, 494, 150, 86, 775, 1063, 242, 564, 1092, 978, 581, 163, 495, 1217, 1218, 406, 2465, 2722, 2039, 1830, 2576, 1983, 2448, 1637, 2687, 2784, 2377, 2546]. juvenile [2061, 256, 2349, 1984, 2896, 2855, 1817, 2219, 1787, 2518, 2545, 2705, 2230, 2080, 1933, 1541, 1517, 1939, 2028, 1499, 2279, 1490, 1542, 2741, 2134, 2130, 2796, 2467, 1493, 1734, 1579, 2694, 1850, 2683, 2508, 2509, 2156, 2547, 2530, 1576, 2447, 1959, 2374, 2866, 2456, 1940, 2292, 2540, 2774, 2556, 2535, 2353, 1995, 1718, 2743, 1658, 2899, 1853, 2196, 2695, 1865, 2326]. Juveniles [65, 2737, 2216, 1859, 2116, 2846, 2895].

Kalman [2003]. Kamchatka [1463]. Kapuscinski [1880]. Kasilof [1465]. Katsuwonus [402, 1668, 1426, 2748]. Kattegat [1195]. kelp [2890, 2378]. Kelt [773]. Kelts [560, 79, 798]. Kenai [1465, 957]. Kepone(R) [1593]. keta [944, 896, 1453, 1214, 2559, 357, 1222, 1305, 823, 1658, 1460]. Ketone [100]. Key [1233, 2586, 2418]. Keynote [1450]. keys [1234]. kHz [2252]. Kidney [272, 1657]. killer [1862]. kindti [83]. Kinetics [1566, 2008, 1533]. King [1086, 219, 465, 345, 150, 949, 1600, 2201, 2275, 1815]. kiss [1307]. kisutch [2708, 2576, 835, 1382, 756, 2177, 2597, 1824, 256, 899, 200, 999, 2896, 243, 809, 941, 789, 2119, 774, 2235, 2741, 1850, 482, 733, 1799, 472, 659, 658, 2627, 623, 1940, 2774, 2897, 2348, 950, 337, 724, 98, 1327, 1381, 1380, 308, 1300, 65, 484, 624, 1051, 1410]. knowable [2125]. Knowledge [456]. Known [1214, 1656]. Kokanee [1121, 582, 53, 284, 2354, 250]. Krabbe [255]. Kraft [1196, 674, 438, 743, 1383, 960, 1525, 2346, 2344, 1610]. Krill [2589, 2772]. Krøyer [2066]. Kuroshio [2752]. kurstaki [1265]. Kvichak [1122].

L [2185, 2137]. L-alanine [2137]. L. [1969, 1970, 1611, 189, 2754, 926, 1821, 1635]. Labelled [62]. labile [2120]. Laboratory [801, 1140, 889, 110, 351, 314, 2102, 1969, 2822, 1556, 1839, 2024, 2274, 1512, 2273, 2436]. Labrador [2215, 2514, 676, 862, 445, 1376, 275]. labrx [1496]. Lac [787]. l’Achigan [1562]. Lack [328, 2806]. lacking [2512]. lacs [410]. Lactation [562, 1937]. Lacustrine [692, 282, 956, 2615]. lacustris [2823, 2794]. laden [2447]. laevigata [2822]. laeviusculus [1145]. Laflamme [1697]. Laguna [1469]. Lake [1183, 2308, 1128, 397, 285, 179, 510, 656, 690, 2194, 346, 2264, 1601, 989, 395, 929, 980, 1201, 675, 709, 145, 259, 162, 161, 527, 1489, 198, 840, 839, 1438, 2148, 599, 1334, 655, 2539, 2088, 2460, 83, 1045, 1127, 1295, 207, 787, 939, 1439, 48, 1356, 865, 1350, 1279, 1443, 476, 1597, 335, 816, 13, 818, 1255, 871, 829, 47, 43, 1404, 52, 2277, 2587, 584, 2079, 82, 403, 1009, 1527, 1285, 166, 1555, 321, 542, 34, 2832, 984, 1247, 284, 875, 1240, 1007, 552, 131, 224, 1131, 747, 338, 1420, 572, 1253, 26, 45, 44, 110, 1760]. Lake [743, 843, 538, 884, 305, 924, 686, 2112, 120, 1013, 2302, 2301, 962, 227, 18, 577, 1386, 140, 936, 1872, 728, 2036, 1585, 2870, 1412, 290, 1388, 652, 651, 29, 89, 100, 780, 993, 222, 1276, 76, 211, 422, 429, 1728, 1239, 1354, 213, 886, 1359, 378, 1303, 2304, 971, 1166, 1018, 1151, 1393, 10, 2309, 2859, 1950, 1830, 2389, 2254, 2029, 1731, 2487, 2155, 1781, 2225, 1605, 2377, 2087, 1680, 2502, 2318, 2306, 2307, 2797, 1629, 1582, 1892, 2795, 2305, 2141, 1900, 2218, 2290, 2206, 719, 1996, 2707, 2337, 1652, 2639, 1596, 1974, 2371, 2643, 2756]. lake [1951, 2886, 2876, 1863, 2471, 2452, 2849, 2578, 1975, 1397, 1973, 2147, 2515, 1965, 1624, 1595, 1594, 1532, 2158, 2755, 2441, 1765, 2224, 1897, 2090, 2033, 2665, 2646, 1628, 2351, 1719, 2808, 1980, 2405, 2606, 2789, 2287, 2408, 1730, 1729, 1891, 2783, 2110, 2270, 2823, 2794, 1943, 2062, 2489, 474, 1849, 1950, 318, 1536, 690, 1231, 346, 454, 380, 462, 706, 1485, 908, 1079, 1697, 456, 442, 2318, 452, 1117, 2175, 2501, 1998, 390, 198, 762, 1441, 1391, 1871, 1538, 592, 2716, 1033, 2644, 1479, 1127, 1396, 2337, 460, 2893, 1389, 969, 2081, 389, 557, 1261, 1792, 1868]. Lake [2085, 2526, 1390, 382, 1885, 2160, 46, 2720, 457, 1625, 1951, 47, 2710, 2739, 1768, 2015, 2253, 166, 321, 2664, 2012, 2505, 2303, 1394, 2878, 112, 2104, 59, 2659, 1520, 552, 864, 453, 1946, 2515, 747, 1895, 1896, 794, 1572, 44, 1216, 1760, 2793, 438, 458, 1387, 2629, 19, 2228, 2828, 202, 686, 209, 2798, 832, 2416, 2417, 2291, 66, 2331, 455, 1980, 780, 2278, 289, 1976, 381, 2492, 2363, 1929, 28, 1300, 2123, 459, 2409, 1486, 1793, 1911, 1393, 784, 1494, 1395]. Lake-Level [1045]. lake-outlet [2090]. lake-resident [1680]. lake-spawning [2155]. lake-wide [1628]. Lakes [1365, 410, 1290, 353, 639, 412, 735, 174, 1732, 55, 976, 395, 1201, 1437, 496, 501, 1370, 897, 585, 1316, 124, 1349, 1032, 876, 37, 479, 957, 753, 157, 1441, 1070, 2125, 1321, 1294, 1350, 509, 450, 492, 880, 1155, 1389, 1434, 1072, 1056, 942, 125, 1159, 561, 1332, 1421, 1419, 566, 1185, 7, 350, 617, 238, 468, 1404, 567, 1392, 1031, 1280, 682, 934, 1299, 1394, 1403, 1397, 2703, 199, 1179, 323, 1062, 1973, 6, 1015, 1264, 95, 1253, 45, 293, 830, 843, 1057, 924, 1149, 1413, 1274, 120, 11, 1407, 107, 962, 588, 227, 1771]. Lakes [820, 1725, 845, 973, 1976, 1401, 507, 1414, 1028, 1158, 1440, 1463, 1368, 42, 292, 618, 895, 1811, 2727, 2675, 1650, 1870, 2636, 2497, 2364, 2484, 2495, 1991, 2853, 2701, 1886, 2427, 2485, 2312, 2673, 2721, 1653, 2468, 1873, 1780, 2282, 2296, 2103, 2523, 1857, 1837, 2475, 1742, 2800, 1985, 1523, 2824, 2396, 1621, 2552, 1759, 2369, 2881, 2838, 1931, 2875, 603, 2111, 2112, 2033, 1559, 2445, 1543, 1744, 1616, 2686, 2538, 2697, 1769, 2405, 2388, 2113, 2609, 2419, 2323, 2608, 2357, 1798, 2135, 2246, 2455, 1832, 1574, 1599, 2249, 2220, 1917, 1916, 1847, 487, 2835, 1116]. Lakes [1829, 613, 431, 439, 1996, 2691, 766, 2207, 2403, 2507, 1630, 736, 1118, 1345, 2236, 2247, 2674, 1628, 726, 461, 1310, 1418, 1422, 2105, 618, 201, 2162]. Lakewater [153]. Laminar [1238]. Laminar-Flow [1238]. Laminated [1396]. Lampetra [425, 1544, 2894, 931]. Lamprey [425, 1010, 1006, 373, 1369, 1188, 1255, 28, 931, 201, 1132, 1544, 1854, 1602, 2894, 1592, 1508, 1507, 1980]. Lampreys [1401]. lampricide [2598]. Lampsilis [1801, 1486]. Lance [1076]. Land [2218, 2859, 2835, 1588, 1965]. land-use [1965]. Landed [1415]. Landings [1292, 745, 1840]. Landlocked [1124, 622, 621, 201, 1602]. Landscape [1448]. Lanthanide [983]. Large [316, 156, 565, 920, 391, 259, 198, 647, 329, 2747, 1292, 16, 1007, 716, 715, 2798, 2575, 2871, 1652, 2176, 2821, 2490, 2115, 1838, 2527, 2120, 1651, 2789, 2408]. Large-Lake [198]. Large-Scale [391, 329, 16, 716, 715, 1292, 2798, 2115]. Largemouth [476, 704, 262, 2630, 2646, 2790]. larger [2889]. larger-fish-upstream [2889]. Larkin [1255]. Larus [2691]. Larvae [1352, 662, 477, 229, 1323, 951, 119, 868, 723, 1372, 600, 609, 955, 518, 799, 677, 882, 831, 968, 231, 612, 1176, 536, 744, 1141, 748, 856, 413, 821, 2760, 1732, 2320, 2404, 2401, 2746, 2023, 1809, 1673, 1581, 1643, 2116, 2274, 2065, 1586, 1908, 2811, 1622, 2144, 1507, 2651, 2165, 1793, 1721, 2555, 2457]. Larval [486, 1854, 862, 135, 1252, 77, 473, 2450, 214, 626, 449, 1017, 287, 1188, 1026, 436, 1746, 1073, 2662, 2913, 796, 953, 489, 869, 1105, 627, 63, 56, 466, 870, 878, 502, 729, 767, 1364, 1401, 775, 2196, 1018, 2682, 1291, 2732, 1670, 1671, 1827, 2502, 1922, 2365, 2860, 1864, 2490, 1606, 2237, 1592, 1946, 2240, 1674, 1526, 2025, 2866, 2200, 2558, 2440, 2056, 1784, 1721, 1912, 1910, 2912, 2022]. Larval/Early [473]. Larvicidal [913]. Larvicide [1307]. Laser [1358, 1012, 2298, 1618, 2854]. Late [1465, 473, 168, 2283, 1509, 2230, 2013]. late-season [1509]. Late-summer [2283]. Lateral [473, 1517, 2350]. Latimeria [1288]. latipinna [1108]. Latitude [412, 932]. latitudes [2328, 2503]. Latitudinal [2268, 650]. l’Atlantique [1234]. Laurent [719, 2884, 414]. Laurentian [431, 487, 1152, 1116, 479, 1064, 2207, 682, 2840, 2105, 2162]. lavaretus [390]. Lavinia [1638]. law [2681]. Lawrence [1365, 1180, 1270, 1212, 2281, 90, 2797, 214, 787, 939, 1211, 2376, 2621, 2384, 2884, 1801, 1968, 2158, 2558, 1507, 1771, 1927, 2772, 966, 850, 2901, 2432, 156, 556, 1874, 1686, 1019, 1670, 2844, 814, 172, 610, 1918, 245, 1225, 2839, 1181, 719, 851, 2428, 2616, 2327, 414, 2840, 2069, 2127, 231, 1176, 2100, 2099, 517, 594, 356, 714, 1103, 2126, 666, 2035, 171, 92, 1005, 946, 1790, 2642]. lawrencianus [1107]. Layer [1438, 119, 76, 581, 2035]. Layers [382]. Laying [609]. Lead [280, 2803, 1753, 2008, 2670, 1636]. leading [1993]. Leaf [920, 332, 782, 997, 1158]. Leaky [2]. Learned [1376, 2852]. Learning [695, 1091]. Least [1250, 1205]. Leech [959, 1655]. Leeches [408, 1321]. Leidy [2166]. Leiostomus [215, 1576]. Length [754, 1226, 32, 349, 742, 932, 261, 642, 134, 684, 529, 244, 650, 21, 649, 446, 1278, 2136, 1627, 1776, 2159, 1991, 2601, 2641, 1488, 2154, 1843, 2382, 2759, 2829, 2895, 2513, 2153, 2628, 1600]. length- [1627]. Length-at-Age [642, 244, 2895]. length-based [2601, 1600]. length-frequency [2382, 2628]. Length-specific [1278]. Lengths [684, 650, 755]. Lengths-at-Age [650, 755]. Lentic [408, 2631, 2849]. Leo [2258]. Leopard [23, 2117]. leopardus [2764, 2060, 2480]. Lepeophteirus [2074]. Lepeophtheirus [2066, 1549, 2750]. Lepidosiren [1288]. Lepomis [1637, 1816, 1814, 1297, 1213, 485, 1054, 953, 2424, 366, 1366, 2339, 190]. Leptodora [83]. Lernaeocera [2422]. Lernaeopodidae [1679]. Lesions [743, 668, 481, 2469, 2828]. Lessons [1091]. Lethal [1771, 1704]. Lethality [337]. lethostigma [1984]. Letters [25, 252, 274, 340, 409, 434, 516, 570, 593, 689, 713, 738, 788, 1260, 1283, 1377, 543, 914, 964, 990, 1016, 1041]. Lettres [25, 252, 274, 340, 409, 434, 516, 570, 593, 689, 713, 738, 788, 1260, 1283, 1377, 543, 914, 964, 990, 1016]. leucas [1120, 625, 1328]. leucichthys [2847]. Leuciscus [2623, 2821]. leucomaenis [2813]. Leucopenia [1410]. Level [1079, 703, 1045, 539, 321, 747, 1418, 1028, 1997, 2376, 2169, 1848, 2617]. Levels [1365, 667, 1201, 172, 560, 1416, 1119, 1444, 933, 257, 169, 1056, 448, 1362, 1007, 654, 438, 743, 227, 29, 811, 1551, 2385, 1535, 2384, 2856, 2015, 2012, 1821, 2729, 2625, 1938, 1507]. liberty [1972, 2742]. Lice [954, 1678]. LIDAR [1930]. Life [179, 1263, 362, 822, 959, 485, 23, 1479, 583, 860, 283, 2733, 1083, 924, 339, 1293, 1025, 974, 242, 1184, 378, 1303, 2582, 812, 821, 1970, 1602, 2021, 2087, 2088, 2792, 1695, 2272, 2085, 2451, 1867, 2185, 2420, 2217, 1957, 2553, 2841, 2491, 2627, 1649, 2907, 2895, 2606, 1966, 2211, 2899, 1828]. life-cycle [2087]. Life-History [1479, 812, 1293, 1649]. life-span [1828]. ligamentina [2129]. Ligands [1153]. Light [365, 597, 415, 2690, 257, 955, 1007, 1420, 953, 820, 902, 76, 971, 1531, 1652, 2577, 2147, 1896, 2405, 2783, 2786, 2868, 2869]. light-based [2786]. light-mediated [1531]. Ligia [1669]. Ligulae [413]. Like [82, 2257, 2828]. Likelihood [32, 71, 1220, 530, 1094, 2171, 1505]. Likelihood-Based [32]. Limanda [2586, 2418, 2375]. Limb [712]. limbata [1941]. Lime [343]. Limed [271, 942, 273, 1414, 2864]. Liming [1885, 272, 577, 270, 1166, 1536]. Limit [854, 197, 629]. Limitation [1110, 706, 165, 711, 1247, 209, 502, 1092, 1485, 2346, 1489, 2141, 2122, 2850, 2643, 1896, 2345, 1924]. Limitations [508, 1979]. Limited [1142, 293]. Limiting [832]. Limits [451, 1716, 2139, 1530, 2585, 2482]. Limnetic [905, 18, 1267]. Limnocorrals [298]. Limnological [2765, 1274]. limnology [2111, 2112]. Limnoplankton [508]. limosa [194]. Lindegren [2916]. Line [78, 923, 721, 1501, 1517]. lineages [2834]. Linear [2679, 74]. Lines [1318, 1903]. Lingcod [109]. Link [544, 2127, 2126]. Linkage [753, 760, 777]. linkages [2218]. Linking [653]. links [1551]. lions [2222]. lip [1889]. Lipid [1357, 855, 818, 1606, 907, 1496, 1108, 612, 100, 1301, 2029, 1991, 2475, 2329, 2903, 2191, 1812]. Lipid-based [855]. Lipids [1128, 937, 1288]. lipofuscin [2481]. Lipoproteins [84]. Liquid [1138]. Lironeca [1992]. list [1656]. literature [2725, 2203]. litre [2729]. Litter [1230, 920, 332]. Little [989]. Littoral [735, 135, 1448, 1072, 1232, 1347, 1345, 493, 115, 1274, 1109, 1143, 973, 1276, 2320, 2392, 2485, 2707, 2471, 2452, 1669, 1616, 2004, 1728, 1730, 1729, 2317, 2114]. Littoral-zone [1347, 1616, 2004]. Live [605, 1281, 870]. Livebearing [1108]. lived [2787]. Liver [1318, 420, 272, 438, 881, 1221, 1903, 1804, 2586, 2418, 2453, 1610]. living [2342, 2190, 2344, 2065, 2731]. loach [1557]. Loading [38, 709, 1390, 305, 2599, 2505, 1965, 2302, 2351, 2278]. loads [2278]. Lobodon [1740]. Lobster [137, 520, 302, 1098, 1341, 868, 1292, 143, 734, 505, 267, 693, 1049, 883, 2142, 1773, 1590, 2265, 2122, 2632, 2231, 1512, 2251, 2481, 2848, 2892]. Lobsters [58, 1192, 152]. Local [391, 1400, 1778, 2011, 2567, 1934]. Local-Scale [391]. Localisation [926]. localization [926]. localized [2519]. Locally [433]. located [2111, 2112]. Location [1000]. Locations [87]. Loci [431, 69, 530, 234, 2244, 1691, 1869]. locomotor [2725]. Locus [586, 191, 1293, 1040, 1981]. log [1717]. Logged [127, 1168, 2747]. logic [2724]. Logistic [131, 178]. Logit [1198]. Lognornial [848]. Loliginid [1243]. Loliginidae [2382, 1243]. Loligo [208]. Lolliguncula [2382]. Long [885, 2486, 208, 162, 161, 1977, 1119, 674, 1369, 1185, 2064, 747, 1253, 1083, 2440, 511, 298, 544, 2787, 1780, 2523, 2444, 1895, 2045, 2050, 1798, 2585, 2482, 1917]. Long-Finned [208]. long-lived [2787]. Long-Term [885, 674, 747, 1083, 511, 298, 544, 2486, 1977, 1119, 1369, 1185, 2064, 1253, 2440, 1780, 2523, 1895, 2045, 2050, 1798, 2585, 2482, 1917]. Longevity [17, 491, 2848]. longiceps [268]. longimanus [2843]. Longitudinal [1560, 1875, 39, 1842, 2574]. Longline [1312]. longlines [1506]. Longlining [128]. longueur [575]. longueur/écaille [575]. Look [540]. Loss [906, 296, 913, 1027, 2687, 2820, 2335, 2767, 2326]. Losses [818, 2249]. Lost [1424, 2449]. Lota [1585]. Lotic [239, 629, 1169, 641, 1385, 506, 2366, 1866, 2340]. Lough [1635]. Louis [2104]. louse [2066, 1549, 2074, 2750]. Low [1139, 510, 421, 911, 698, 509, 273, 1317, 2596, 704, 184, 1406, 2040, 151, 811, 1440, 856, 372, 1018, 1307, 1969, 2192, 1605, 2842, 2427, 1977, 2266, 2817, 2900, 2750, 2625, 2617, 2789]. Low-effort [2596]. Low-Order [151]. low-productivity [2427]. Lower [2714, 29, 1741, 2247, 1915, 2671, 2728, 137, 431, 539, 1118, 525, 495, 2901]. lowland [2457]. LR [1985]. lucioperca [670]. lucius [2242, 2469, 2876, 2884, 590, 1373, 1081, 292]. lumholtzi [1495]. Lunenburg [2039]. lungfish [1288]. Lungworm [2333, 2470]. lupus [2683]. Lutjanidae [78]. Lutra [899]. Lymnaea [1636]. Lymphocystis [911]. Lysine [900].

M. [1683]. mâles [414]. MacCall [2916]. maccoyii [32, 436]. Mackenzie [539, 2013, 2477]. Mackerel [315, 814, 245, 769, 1059, 815, 1788]. Macro [1450]. Macrobenthic [703]. Macrobenthos [458, 956]. macrocephalus [537, 579]. macrochirus [1637, 1816, 1814, 1297, 953, 2424, 366, 2339, 190]. macroepifauna [2046]. Macrofauna [8, 581]. Macrograzers [592]. Macroinvertebrate [239, 532, 1109, 1276, 1158, 1745, 2269, 2087, 2306, 1875, 1944, 2629, 2347, 2492]. Macroinvertebrates [127, 1193, 641, 924, 2631, 2306, 2307, 2176, 2537]. Macrophyte [291, 37, 1072, 542, 681, 844, 2315]. Macrophytes [54, 1448, 456, 1279, 46, 879, 493, 2507, 2330, 2836, 2317]. macropterus [1399, 1271]. Macrotidal [898]. macrozooplankton [2157, 2798]. Macruronus [2555]. maculatum [987]. made [2473]. Madeleine [1026]. Madre [1469]. Magdalen [90]. magellanicus [2039, 1749, 1071, 1921, 2173, 1783, 1374, 1001, 1197, 802, 928, 916, 734, 1554, 748]. magister [906, 1097, 723, 910, 796, 1995]. magna [2479]. magnitude [2087]. mahimahi [1959]. Main [1944]. Maine [527, 2027, 1834, 334, 349, 1170, 1073, 2676]. maintenance [1901]. Majidae [2432, 1571, 2102, 2498]. Major [957, 198, 1333, 1659, 1601, 2081]. making [1822, 2402, 2751]. mako [1843]. malachite [1926]. Malacostraca [1357]. Malate [356]. Male [396, 1355, 1272, 841, 750, 540, 560, 19, 326, 811, 1184, 1136, 1925, 1770, 2472, 2677, 414, 2115, 1571, 2591, 2771]. Males [691, 909, 2267, 1777, 2143, 2102]. Malic [191]. Mallotus [1686, 2159, 1145, 1922, 2365, 1038, 1932, 2451, 2170, 2840, 838, 536, 517, 1227, 1291]. malma [2118, 256, 899, 2199, 533, 1124, 2804]. Mammalian [1318]. Mammals [1360, 1836]. Management [1709, 427, 1090, 672, 361, 2583, 607, 1175, 212, 1375, 526, 1, 226, 1236, 1162, 1708, 777, 1432, 725, 2607, 1918, 2360, 2839, 2260, 1710, 1953, 2834, 2718, 1965, 2745, 2127, 1631, 1400, 2560, 2541, 2398, 2126, 2336, 1914, 2414, 2361, 2606, 2653]. managers [2181, 2310]. Managing [2316, 2826, 2650]. Manan [58]. Mandibles [1299]. maneuvers [2038]. Manganese [1183]. Manifestation [911]. Manipulated [476, 1916]. Manipulation [79, 798, 867, 693, 2115]. Manipulations [367, 971, 2484, 2849]. Manitoba [561, 1746]. Manitouwadge [2194]. Mantle [407]. many [1562]. Mapping [616, 574, 1934, 2510]. Maps [828].

Marennes [325]. Marginal [192]. Margins [122]. Margolis [2258]. margolisi [1664, 1667, 1668, 1676]. marina [2219]. Marine [1286, 149, 1263, 683, 1089, 1145, 872, 580, 1090, 1085, 813, 696, 1468, 373, 243, 436, 601, 1042, 1043, 1047, 339, 489, 1223, 1433, 481, 2772, 744, 308, 569, 1088, 355, 2187, 418, 419, 1014, 2480, 2499, 2874, 1671, 1827, 1813, 2356, 1767, 1836, 2837, 2163, 2322, 1947, 2229, 1663, 2819, 2851, 2145, 2897, 1533, 1955, 575, 2908, 2854, 2640, 1910]. marinus [1602, 1006, 373, 1369, 1188, 1255, 1592, 1980, 201, 1132]. Maritime [527]. Maritimes [868]. maritimus [1762]. Mark [1235, 309, 983, 1220, 620, 2300, 1800, 2293]. mark-recapture [2300]. Marked [1094, 414, 1940]. marker [2352, 1672, 1735]. Markers [1001, 2542, 1764, 2245, 2559, 2372]. market [1841]. Marking [892, 1591, 657]. Markov [2681]. marks [2440]. Marlin [1339]. marmoratus [1829, 1191]. marqués [414]. Marsh [1230, 1654]. Marshes [550, 2835]. masou [1833, 2221, 2571, 1184]. masquinongy [205, 1994, 1081]. Mass [1012, 983, 2787, 2298, 1647, 2190, 2283, 2079, 2369, 2854]. Massachusetts [1157]. Massena [2884]. Masu [1184, 1833, 2221, 2121]. Match [1145, 214]. Match/Mismatch [1145, 214]. matching [2517]. Mate [1124, 2669]. mated [2102]. Material [1045]. Maternal [1986, 1024, 2732, 2813, 2845]. Mathematical [532]. Mating [1272, 401, 714, 1289, 267, 1814, 2267, 2143]. Matriarchal [1998]. Matter [1167, 353, 634, 675, 984, 154, 1950, 2342, 2706, 2501, 1510, 2802, 2755, 2851, 2164, 2699, 2789]. Maturation [1355, 346, 1371, 819, 1134, 1333, 643, 326, 1770, 2494, 2092, 2504, 2476, 2382, 2478, 1571, 2667]. Mature [396, 691]. maturing [2858]. Maturity [1272, 841, 750, 208, 903, 191, 393, 175, 743, 1047, 64, 109, 1046, 3, 1035, 446, 1204, 2204, 2680, 1740, 1777, 2759, 2121, 1828, 2270, 2458]. Maurice [1196]. Maxima [467]. Maximal [553]. Maximizing [934]. Maximum [71, 13, 530, 2900, 29, 483, 564, 2171, 2206, 2510, 1505]. maximus [2459, 1141]. May [1076, 119]. mayflies [2817, 1941]. Mayfly [421, 1209, 1416, 280, 2581]. McAllister [1881]. McClatchie [2916]. McLeod [211]. meal [2170]. Mean [203, 650, 755, 961, 1740, 2139]. Means [1095, 1978, 1545, 2003, 2531]. Measure [496, 1266, 1278, 1492, 1606, 2561, 1630, 2743]. Measurement [2495, 387, 1062, 938, 261, 2172, 2553]. Measurements [1174, 922, 424, 1380, 972, 2887, 1906, 1820, 1912]. Measures [380, 1419, 1053, 1844, 2390, 1935]. Measuring [1219, 2284, 904, 206, 677, 1147, 1858, 2285, 233, 74, 2149, 2837, 2241]. mécanisme [1852]. Mechanics [348]. Mechanism [155, 204, 1852, 2094, 1549]. Mechanisms [1984, 2723, 59, 1203, 493, 2227, 2755, 2169]. mechanistic [1906, 1820]. median [2367]. Mediated [102, 1531, 870]. mediation [2574]. Mediterranean [2579, 2623, 1496]. mediterraneus [1667]. Medusae [443]. meeting [2753]. Megalopal [910, 1995]. Megalophallus [1669]. Megaloptera [2768]. MeHg [2164]. Meiobenthic [215, 640]. meiofauna [2819, 2851]. melaena [1157]. Melanogrammus [573, 598, 40, 2011, 2662, 1579, 627, 2895, 2383, 849]. melanostictus [2752, 2634, 1634, 1784]. melanostomus [1829]. Melvin [1635]. member [1823]. member-vagrant [1823]. Members [113, 1099]. Membership [214]. Membership/Vagrancy [214]. membras [643]. Mendota [2123, 2012, 2505, 2278]. mendotae [41, 1102, 1917]. Menidia [2268, 2503]. menta [1732]. mentella [1968, 356]. Menten [285]. Merced [2714]. Mercenaria [854]. mercuric [1482]. Mercury [586, 587, 683, 976, 690, 1201, 118, 995, 1064, 1753, 1866, 561, 448, 894, 1633, 572, 2729, 1754, 924, 1373, 962, 975, 62, 1126, 1268, 213, 1599, 292, 2499, 1750, 2692, 1804, 2594, 2471, 1752, 1906, 1820, 1938, 2165, 2009, 2220]. Meristic [527, 601, 1078, 2616]. Merluccius [2406, 2688, 498, 522, 1580, 2476, 1683, 2299, 244, 650, 755, 525, 74]. Meromictic [1015]. Meserve [1663]. Mesh [804]. Mesocosm [51, 519, 867, 2819, 2851]. Mesodesma [1277]. Mesodesmatidae [1277]. Mesoscale [1983]. Mesotrophic [843, 429]. meta [2288, 2334, 2907]. meta-analysis [2288, 2907]. meta-analytic [2334]. Metabolic [488, 148, 1876, 2002, 843, 63, 352, 1217, 1218, 2722, 1950, 2568]. Metabolism [1140, 500, 697, 96, 800, 1374, 688, 1320, 521, 270, 100, 817, 1051, 1301, 1509, 2149, 2168, 2912, 2061, 2913, 2444, 2563, 1926, 2593, 2007, 2562]. Metabolizing [420]. Metal [1365, 1297, 1334, 549, 877, 1017, 887, 2621, 679, 1154, 512, 399, 1409, 2194, 2193, 2803, 1775, 1550, 1551, 1988, 2253, 1863, 2731, 1917, 1916]. Metal-binding [1297]. Metal-Contaminated [679, 2194]. metal-spiked [2193]. Metalimnetic [16, 1421]. metallothionein [1550, 1551]. Metals [1270, 757, 1157, 1113, 1206, 1703, 2364, 2612, 1700, 1707, 1714, 1944, 1742, 2884, 1705, 1704, 926, 1801, 1702, 2419, 1706, 2457]. metals-acclimated [1705]. Metals-contaminated [1703]. metamorphic [2892]. Metamorphosis [1369, 1132, 2041]. Metaphytic [592]. Metaphyton [125]. metapopulation [2332, 2826]. Metastrongyloidea [2470]. métaux [2884, 926, 2833]. metazoan [2261]. methane [2389, 2254, 2686, 2036]. Methanogenesis [397]. Method [632, 2548, 309, 790, 718, 32, 349, 1219, 336, 557, 138, 1220, 330, 264, 495, 2117, 2648, 2362, 2495, 2842, 1561, 2721, 2814, 1234, 2837, 2261, 2651, 2425, 2458, 1737]. méthode [1234]. methodologies [2725]. Methodology [1125, 81]. Methods [73, 2655, 1030, 259, 1399, 69, 2654, 1241, 938, 677, 248, 530, 1385, 1147, 1053, 356, 363, 998, 729, 767, 446, 1398, 2150, 1501, 2765, 2744, 1518, 1558, 2603, 2657, 2311, 2091]. Methyl [2594, 2471, 975, 1938, 1268, 2692, 2729, 2165, 2009]. Methylation [572, 1633]. Methylisoborneol [836, 61]. methylmercury [1750, 2072, 1593, 1721]. metrics [2017]. meuniers [2884]. Mexico [2043, 2048, 2046, 1468, 2382, 236, 2045, 1676, 2050, 2205, 369]. MFO [438, 1298]. Michaelis [285]. Michigan [2079, 1630, 1628, 1418, 474, 380, 2264, 1991, 2318, 198, 762, 613, 469, 382, 1421, 1419, 2664, 112, 59, 1420, 2629, 2828, 209, 1585, 290, 289, 1422, 381, 31, 2053, 1300, 1494]. Michx.

[2621]. Micro [1450]. Microalgae [36, 418, 419, 2146]. Microalgal [1138]. microalgues [2833]. Microbes [1158]. Microbial [855, 572, 782, 1414, 1650, 2706, 1938, 2110]. Microchemistry [1171]. microcosm [1961]. Microcosms [1080, 1938]. microcrustacean [1572]. Microcystin [1985]. Microcystin-LR [1985]. Microcystis [2850]. Microdistributions [1125]. microevolution [1661]. Microflora [1352]. Microfossil [211]. Microgadus [1741, 1075]. Micrograzer [209]. Microgrowth [700]. Microhabitat [635, 375, 256, 999, 423, 987, 80, 1327, 2400, 1623]. Microhomogenization [1750]. microincrement [2091]. microlepidotus [1757]. Micronucleus [237]. Microphallidae [1669]. microplankton [2018, 1486]. microprobe [2298]. Micropterus [2630, 476, 704, 262, 2790, 2291, 383]. Microsatellite [2057, 2514, 2773, 2689, 2244, 1691, 2228, 2245, 1848, 2390, 1869, 2885]. Microsatellites [1360, 1791, 1814, 2459, 2332, 2016]. microscopy [2322]. microsomal [1960]. Microstomus [2131]. microstructural [1802]. Microstructure [1012, 596, 1123, 494, 1515]. Microtox(R) [1113]. mid [2676]. mid-Atlantic [2676]. Middle [2682]. Midge [119, 2320, 1809]. midges [2445]. midnorthern [2164]. Midwater [360, 2406]. migrant [2535]. Migrating [810, 132, 265, 53, 778, 1565, 2567, 2385, 1547, 2300, 2568, 2252, 2326]. Migration [1121, 888, 991, 872, 309, 249, 696, 539, 16, 648, 33, 34, 199, 204, 284, 733, 838, 370, 80, 1037, 41, 1531, 1925, 2066, 2738, 1990, 2844, 2873, 2690, 2290, 2521, 2522, 1549, 2740, 2536, 2770, 1802, 2649, 2456, 2223, 2883, 2167, 2319, 2901]. Migrations [425, 832, 92, 823, 2429, 2374, 2854, 2585, 2482]. Migratory [396, 172, 1055, 472, 258, 374, 650, 277, 1014, 2443, 1696, 1904, 2319, 2454]. Mill [1384, 1196, 674, 278, 590, 438, 743, 1383, 960, 1525, 2494, 2370, 2346, 2344, 1610]. Miller [1880]. Minas [1228]. Mine [304, 2030]. mineralization [2851]. Minisateliite [1305]. Minisatellite [1474, 1819, 1981, 1799, 1735, 2591]. Minnesota [1290, 2620, 1555, 603, 177, 1699]. Minnow [1199, 1207, 1017, 1018, 2227, 1482, 1643, 1912, 2249]. Minnows [135, 800, 1113, 2391]. minutus [55]. Miramichi [2344]. mirex [1768]. Mirror [941]. Mirror-Elicited [941]. Misgurnus [1557]. Mismatch [1145, 214]. misread [2712]. missed [2161]. Mississippi [2728, 1297, 2888, 761, 591, 2477]. Mitchill [1823]. Mitigate [906]. Mitigation [2671, 269, 2826]. mitigative [1844]. Mitochondrial [1465, 317, 379, 60, 1472, 1823, 126, 347, 1322, 994, 1644, 611, 2177, 281, 944, 1477, 1829, 1641, 1468, 113, 1127, 1271, 939, 1170, 1099, 2083, 1635, 606, 2236, 1478, 478, 1497, 177, 780, 1891, 1949, 873, 849, 1498, 2216, 2894, 2276, 2641, 1843, 2133, 1902, 2245, 1598, 1575, 2372]. Mixed [15, 1438, 1262, 673, 248, 743, 1473, 972, 1298, 111, 234, 333, 1575, 2372, 1205, 2450, 1641, 2598]. Mixed-Function [743]. mixed-prey [2450]. Mixed-Species [1262]. Mixed-Stock [15, 111, 234, 1575, 2372, 1641]. Mixing [2844, 108, 1015, 1644, 2534, 2304]. mixta [2877]. Mixture [800, 70, 1704]. Mixtures [1017, 1113]. Mn [879, 2884]. Mobile [1920, 374]. Mobility [356, 1893]. Mobilization [785]. Mode [866, 2695]. Model [690, 741, 334, 1315, 1164, 12, 1250, 331, 404, 276, 354, 228, 793, 1255, 7, 350, 1229, 451, 875, 6, 1251, 294, 526, 852, 1411, 588, 62, 504, 878, 109, 134, 1198, 2483, 383, 1430, 726, 164, 615, 755, 935, 1029, 21, 483, 1417, 1308, 1368, 1049, 725, 190, 1701, 325, 2736, 2738, 1748, 2077, 2412, 1840, 2898, 2853, 1611, 1892, 2601, 2488, 1522, 2295, 2685, 1999, 2517, 2785, 2038, 2420, 2662, 1864, 2613, 2724, 2170, 2602, 2231, 2663, 2793, 2424, 2627, 2811, 1649, 1909, 1906, 1820, 2679, 2757, 2416, 1585, 2870, 2336, 2606, 2748, 2076]. model [2354, 2123, 1971, 2341, 1612, 2042, 2425, 1901, 1972, 2454, 2901, 1600]. Model-Based [526]. model-free [2724]. Modèle [325]. Modèles [633]. Modeling [2780, 1428, 2795, 465, 402, 2710, 1426, 1280, 2718, 2649, 2781, 2506, 444, 1137, 537, 1136, 2238, 2639, 2916, 2800]. Modelling [571, 318, 708, 607, 1373, 2073, 1154, 1337, 1291, 2648, 2887, 2462, 2314, 2754, 2058, 1533, 2096, 2167, 2904]. Models [353, 635, 1060, 925, 1224, 108, 1346, 2534, 824, 848, 370, 869, 1147, 63, 859, 770, 717, 1330, 103, 28, 722, 602, 1319, 1189, 883, 1470, 2648, 2362, 2912, 2888, 2098, 1640, 1639, 2407, 1711, 1537, 2913, 1945, 1965, 1587, 633, 1502, 2228, 1831, 2769, 1894, 1712, 2437, 2323, 2608, 2717, 1943]. Moderate [703, 509]. Modernized [1383]. Modification [1153]. Modifications [368, 2344, 512]. modified [2781]. Mohn [2052]. Molecular [1450, 2571, 1672]. mollusc [2162]. Molluscs [149, 1508, 2409]. Mollusk [700, 926]. Mollusque [926]. Molt [540, 1204, 2892]. moluccensis [1503]. monitor [2872]. Monitored [1374]. Monitoring [2043, 1083, 2531, 1000, 1930, 2049, 2384, 2297]. monocerata [1076]. monoceros [1120, 312]. Monodon [1120, 312]. Monogenea [1659, 1660, 1662, 1663, 2273, 2436]. monogenean [1661]. Monomeric [185]. monophyletic [2571]. Montagnes [787]. montagui [665, 636]. Montana [282, 1707, 2747, 1705, 1704, 1702, 1599, 1706, 1703]. Montane [367, 608, 2756, 1923]. Monte [2681]. month [819]. Montréal [1642]. Moored [556]. Moose [332]. Morality [139]. moray [1677]. mordax [1531, 1823, 1572, 2506, 2558, 2651]. morhua [1491, 2760, 2057, 1140, 1997, 1263, 1670, 598, 862, 1358, 1748, 2077, 2844, 281, 1986, 1840, 1788, 2546, 1743, 698, 2466, 2839, 2011, 1038, 2397, 922, 631, 1444, 2855, 2219, 2518, 2519, 1839, 2785, 1134, 1133, 1376, 2816, 2669, 2422, 1681, 2525, 138, 548, 1351, 1846, 2718, 1317, 2327, 2662, 1249, 2572, 1606, 2237, 128, 2170, 2524, 1172, 2530, 1586, 1044, 968, 2866, 2100, 2099, 1078, 2144, 2329, 2863, 2891, 374, 594, 1101, 1848, 2056, 2358, 2565, 2806, 2076, 2617, 2096, 1005, 946, 2642, 805, 1688, 856, 372, 1178, 1161]. morio [1676]. Morone [2687, 2328, 1012, 816, 2656, 2500, 358, 84, 1083, 202, 351, 178, 2440, 729, 767, 1853, 873, 2372, 1096, 943, 1314, 1313]. Moroteuthis [1123]. Morozov [1665]. Morphoedaphic [501, 217, 203, 504]. Morphological [885, 922, 1399, 787, 1078, 535, 484, 886, 1048, 2023, 1662, 2025, 1935, 2457]. Morphology [1244, 710, 1173, 941, 1285, 604, 1190, 686, 1168, 471, 660, 1218, 1623, 1607, 2216, 2349, 1933, 2085, 2071, 2350, 2771, 2032]. Morphometric [1272, 527, 1022, 2616, 273, 601, 517, 1671, 1827, 2458]. Morphometrics [868]. Morphometry [841, 1404, 120, 140, 76, 1035]. morphs [1966]. Mortalities [1410, 1544, 1942, 2742]. Mortality [670, 477, 309, 485, 1324, 106, 358, 1069, 546, 345, 795, 874, 175, 882, 1042, 1043, 1059, 30, 339, 1105, 152, 146, 1257, 1427, 1027, 1354, 558, 2728, 2138, 1970, 2066, 1781, 2590, 2830, 1996, 2054, 1580, 1783, 2407, 2610, 2611, 2656, 2500, 2422, 2022, 2554, 2553, 2250, 1502, 1604, 2334, 2073, 2646, 2503, 2513, 2512, 1915, 1658, 2695]. Morwong [1271]. Mosaicism [1456]. Moses [2884]. Mosquitofish [51, 283, 1860]. mosqyitofish [1772]. Moss [610]. mosses [2697]. most [1737]. Motility [418, 2572]. Moulting [714]. moults [1675]. Mountain [675, 880, 1155, 228, 793, 235, 2881, 2838, 1717, 1700, 1944, 1299, 1651, 1599]. Mountains [151, 2138, 1413]. mouth [2844]. mouthpart [1809]. mouvements [534]. Movement [1235, 1324, 901, 1424, 71, 600, 1426, 1367, 538, 363, 31, 1920, 2871, 1684, 2748, 1913, 2480]. Movements [814, 1190, 733, 221, 1327, 1328, 1531, 2109, 534, 2037, 414, 2581, 2473, 1534, 2000, 1923]. mRNA [1821]. mtDNA [1632, 2014, 2558, 1205]. Much [585, 1641]. mucous [2192]. Mucus [918]. Mud [845]. Muitimesh [557]. Müller [1874]. Multi [1975, 1963]. multi-taxonomic [1963]. Multi-year [1975]. Multiagency [441]. multiannual [2316]. Multibasin [140, 2739]. multibeam [2660]. Multicollinearity [131]. multicomponent [2227, 2779]. MULTIFAN [1776, 32, 349, 2601]. MULTIFAN-CL [2601]. Multifilament [557]. Multilocus [1040, 587]. Multiparous [1103, 2763]. Multiple [638, 320, 32, 2613, 401, 2068, 867, 1028, 2885, 1393, 2859, 2727, 2544, 2461, 2286, 2817, 2078, 2454, 1418]. Multiple-Frequency [320]. multiple-reach [2454]. multiple-year [2544]. multiplexing [2016]. multiseries [604, 411, 580, 2173]. Multispecies [1119, 1105, 2717, 1999]. Multistage [742]. Multivariate [156, 1022, 588, 109, 2897, 551, 1126, 2269, 2264]. Multiyear [2533]. mummichog [2344, 1721]. municipal [2370]. Muscle [407, 272, 1243, 356, 2168, 1804, 1817, 1821, 1548, 2020, 2002, 1812]. Muscles [1001]. Muscular [943, 1313]. Musculature [862]. Muskellunge [205, 1081, 1994]. Muskoka [618]. Mussel [1182, 556, 680, 439, 865, 1280, 864, 1246, 736, 1264, 1118, 1387, 616, 859, 1309, 1388, 726, 1310, 440, 1146, 2728, 1849, 2129, 1538, 2093, 2723, 2415, 1764, 2296, 2103, 2543, 1763, 2015, 2253, 1782, 1520, 2426, 1563, 1759, 2071, 1589, 2347, 2870, 1533, 1716, 2704, 1899, 2280, 1818, 2885]. Mussels [980, 1116, 518, 605, 1492, 2810, 2129, 1605, 2501, 2488, 1801, 1760, 2674, 2090, 1528, 1771, 2331, 2363, 1948]. Mustelus [1962, 2355]. mutation [2228]. my [1307]. Mya [189, 428]. Myctophiformes [555]. mydas [1666]. Myelin [1288]. mykiss [2812, 2380, 2722, 2867, 488, 22, 1569, 885, 1139, 1125, 1751, 1987, 1958, 2576, 2168, 1638, 2448, 2677, 1540, 842, 170, 405, 148, 2061, 1998, 2459, 999, 431, 699, 2719, 515, 81, 2227, 2545, 2705, 423, 1296, 1761, 674, 490, 1482, 2661, 1933, 2600, 1017, 476, 1361, 307, 919, 2119, 2756, 977, 417, 688, 1282, 1490, 2381, 2849, 2001, 2134, 2444, 2801, 247, 2325, 1052, 300, 1343, 1344, 521, 1791, 2592, 2228, 881, 2453, 1336, 1855, 870, 1200, 512, 399, 2292, 2540, 2699, 1100, 221, 98, 255, 2353, 1473, 1982, 817, 348, 1217, 1218, 1865, 327, 2032, 1301, 1409]. Myriophyllum [1340]. Mysidacea [1357, 630, 1311]. Mysis [1357, 1811, 2690, 2877, 1421, 1625, 313, 112, 284, 2479, 1311]. mysticetus [2647]. Mytilus [556, 700, 2093, 1764, 1763, 1692, 605, 615, 1533]. Myxobolus [1656]. Myxosporea [1656]. Myxosporean [778].

N [2874, 2842, 1111, 1391, 1873, 1390, 1093, 2876, 14, 1122, 2345, 36, 422, 2203]. n.sp [1664, 1667, 1668, 1676]. N.W.T [1356, 554]. NADH [1902]. Nahanni [655]. naive [1541, 1705]. Namakan [1079]. namaycush [2029, 690, 1201, 2318, 2795, 1127, 1996, 48, 2337, 818, 1255, 47, 875, 552, 2515, 338, 26, 227, 1412, 1980, 2405, 2606, 1891, 2783, 378, 1303, 1151, 2062, 2489]. Namibia [791]. Namibian [522]. Nanophyetus [405]. Nanoplankton [971, 2033]. Nanticoke [162, 161]. Narrow [2846]. Narwhals [1120, 312]. Nass [2689]. nasus [1093, 2245]. nasutum [130]. National [2063, 2340, 1599]. Nations [1375]. Native [865, 384, 1008, 122, 1387, 1475, 1868, 1720, 2523, 2083, 1687, 2090, 1620, 1927, 2162]. Natural [597, 329, 877, 445, 1434, 385, 776, 2536, 673, 795, 827, 802, 786, 1043, 120, 11, 1558, 2699, 1327, 1427, 1092, 1018, 1151, 2095, 2118, 2709, 2636, 2590, 2342, 2830, 1824, 2148, 1977, 1580, 2407, 2429, 2801, 2250, 1752, 2431, 1604, 2336, 2212, 2742, 1851]. Naturalized [225, 528]. Naturally [940, 2211]. Nature [1329, 513, 514]. Nauyuk [1356]. Navigation [556, 1326, 2114]. ND3 [2894]. Near [48, 1062, 244, 554, 1642, 2844, 665, 2044, 2884, 2396, 898]. Near-Infrared [1062]. Nearctic [1472, 1632, 1659]. Nearshore [2519, 457, 864, 44, 1178, 2518, 2674]. nebulosus [487, 2236]. necator [1658]. necessary [1562]. Nechako [1258, 2315, 2345]. Nécrologie [2658, 2827]. Nectopsyche [586, 587]. Need [441]. Needed [456, 2505]. Needs [462, 2310]. Negative [982, 2679]. nehulosus [336]. neiges [414]. Nektonic [550]. Nemadactylus [1399, 1271]. Nematoda [754, 486, 862, 1671, 1827, 1813, 77, 1672, 951, 2470, 1673, 609, 799, 630, 1674, 1675, 1676, 466, 870, 255]. Nematodes [676, 630, 1670, 2069]. Nemertean [302, 146]. Neocalanus [2573, 614]. neogaeus [2290, 1417]. Neogastropod [1447]. Neogobius [1829]. Neomysis [2877, 630]. neopilchardus [279, 1546]. Neoplasia [1075]. Nepheloid [382]. Nephelopsis [959, 746]. nerka [1925, 2689, 2132, 2486, 1121, 2516, 1348, 1315, 15, 2204, 1794, 582, 2267, 2539, 2140, 2862, 265, 389, 1868, 2085, 2526, 695, 2567, 354, 2074, 53, 932, 1682, 1722, 204, 284, 1481, 2528, 2554, 2334, 2624, 370, 2638, 2759, 2829, 2223, 2568, 1259, 2178, 1171, 2398, 846, 2354, 1208, 1463, 1462, 2167, 602, 97, 341, 2585, 2482, 250, 1014, 1461]. Nestling [1076]. nests [2291]. Net [902, 898, 2335, 2679, 1915]. Net-Catch [902]. Nets [557, 1374, 2570, 1888, 1887]. network [2898, 2769]. neural [2898]. Neurotoxic [580]. Neurotoxin [615]. Neuse [2338]. Neutralization [48, 46, 47, 43, 45]. Neutron [181]. Nevada [1033, 984, 2566]. Newfoundland [2413, 1491, 1701, 1384, 676, 862, 2410, 2763, 2546, 208, 2092, 1006, 249, 2518, 2519, 1277, 1376, 1681, 786, 1172, 1044, 1078, 1622, 2144, 1848, 2358, 1178, 1161, 1291]. Newly [749, 65, 2248, 2534, 1586]. newmani [696]. Newnans [1045]. NH [1726]. Ni [1365, 2704]. Niche [1188, 545, 1412, 1845, 2778]. Night [810, 2622]. Nighttime [20]. nigromaculatus [502]. Nile [299]. niloticus [299, 2115, 471]. Nine [311]. Nitrate [1004, 1029, 235, 2842, 2094, 1982]. Nitric [157]. Nitrification [2036, 2819]. nitrite [2094, 1982]. Nitrogen [1849, 1290, 638, 174, 145, 165, 1391, 1295, 539, 1093, 818, 857, 1247, 1062, 293, 2838, 1114, 36, 940, 2499, 1797, 1485, 2618, 2877, 1836, 2647, 1768, 2703, 2577, 1738, 1928, 2625, 2875, 2324, 2203, 2214, 2881]. Nitrogen/Phosphorus [293]. nitrogenous [1648, 2134]. nitrophenol [2598]. Nitzschia [411, 1004, 580, 386, 604, 2173]. Njupfatet [1536]. nm [2760]. NMR [1374]. NO [1726, 2335]. Nocturnal [2230, 2809, 2380]. noirs [2884]. Noise [1287]. NOLSS [839]. Nomenclatural [1309]. Non [74, 1720, 2474, 1989]. non-indigenous [1989]. Non-linear [74]. non-native [1720]. non-suspension-feeding [2474]. Nonanadromous [347, 300, 2195]. noncompetitively [2102]. Nonessential [1340]. Nonesterified [1131, 1925]. Nongrowing [649]. Nonhomologous [870]. Nonlethal [938]. Nonlinear [1308, 2520, 2383, 2084]. Nonlinearity [524]. Nonmigrating [53]. Nonmonotone [1427]. Nonnatal [1275]. Nonparametric [1053, 1840]. nonrandom [1741]. nonswitching [2761]. Nontarget [1193]. Normal [1144, 755, 886]. North [1115, 785, 732, 2861, 405, 1070, 16, 1527, 1696, 2338, 849, 1653, 1523, 2824, 2665, 379, 1826, 656, 1656, 470, 1248, 2614, 165, 655, 1804, 533, 1495, 2282, 2785, 1513, 896, 2133, 2676, 2573, 2557, 1934, 1563, 833, 1691, 2325, 1690, 2073, 1148, 1407, 749, 588, 1046, 2375, 1464, 1779, 891, 569, 2550, 440, 1735, 1305, 1068, 1311, 1462, 2271, 1949, 1599, 1943, 2062, 2489, 812, 2904]. north-central [1523]. North-Temperate [16, 1653]. northcentral [2118]. northeast [2866, 2144, 354, 531, 2561, 544, 2249]. northeastern [1701, 2727, 1664, 2484, 2673, 1667, 2133, 1724, 2419, 127, 236]. Northern [1164, 1165, 854, 349, 1193, 1457, 590, 1187, 1176, 1373, 765, 852, 1281, 594, 171, 92, 574, 1284, 559, 961, 823, 564, 1081, 292, 1874, 2075, 1830, 2377, 2839, 2721, 2469, 2397, 2199, 1882, 2221, 2464, 2876, 2884, 2327, 2672, 2130, 1754, 1604, 2506, 2100, 2099, 2245, 2646, 1651, 2686, 1484, 1570, 2551, 2806, 1862, 2617, 2205, 2053, 1290, 814, 1349, 1181, 655, 1070, 631, 1444, 260, 445, 469, 561, 1285, 705, 1752, 1047, 231, 312, 1227, 1311, 856]. Northwest [1106, 405, 27, 2111, 2112, 211, 581, 2057, 2747, 1020, 797, 2215, 1788, 123, 2421, 2597, 1006, 2839, 839, 373, 1529, 1777, 1134, 1133, 2422, 1059, 1223, 1458, 8, 1103, 815, 1130]. Northwestern [1437, 1321, 2432, 2046, 2382, 1392, 1931, 1727, 2196, 556, 26, 714, 973]. Norway [1970, 995, 1334, 1885, 1285, 1862]. Norwegian [929, 329, 1471, 546, 2866, 2538, 1862, 2378, 406]. nostocacean [1392].

Notemigonus [662]. Notice [804]. Notropis [459]. Nouveau [534]. Nova [2039, 137, 443, 208, 2905, 1292, 1374, 272, 2251, 666, 2890]. novaezelandiae [2555]. Novel [169, 972, 2814, 2452, 1869]. Nuclear [1475, 2276, 1764, 1763, 2372]. Nucleic [2466, 978, 1564, 2453]. Nuisance [592]. Number [430, 1415, 1573, 2268, 1675, 2813, 2414]. numbers [1655, 2300, 2750]. Numerical [354, 640, 863, 2887]. nurseries [1576]. Nursery [152, 1490, 2131]. Nutrient [1290, 638, 1110, 501, 872, 259, 2346, 1438, 1439, 460, 1419, 984, 705, 2672, 1007, 2552, 1407, 455, 846, 29, 506, 2484, 2713, 1993, 2599, 1489, 2256, 1597, 1936, 2643, 1897]. Nutrient-chlorophyll-Secchi [2552]. Nutrients [411, 1270, 159, 753, 1441, 629, 1332, 820, 154, 1440, 898, 2012, 2592, 1765, 1847]. Nutritional [5, 922, 931, 2873, 1606, 1564]. NWT [159]. Nymphs [280, 1156].

O [1741, 1490, 397, 1164, 1165, 2842]. O. [756, 944, 999, 809, 919, 1453, 714, 950, 98, 1327, 1300, 1410, 1470]. Obituary [2658, 2827, 1954, 1905, 2258]. Obituary/Nécrologie [2658, 2827]. objects [2753]. obscura [959, 746]. obscure [1860]. observation [2462, 2123]. observation-error [2462]. Observations [573, 167, 552, 110, 1411, 216, 1178, 2766, 2504, 2828]. Observed [663, 2902]. observer [2688]. observers [2779]. Observing [1434]. Obstinate [1329]. Obtained [751, 2168]. Occasion [569]. Occur [885]. Occurred [585]. Occurrence [208, 1006, 1333, 546, 1992, 859, 1862, 201, 1666, 1825]. occurs [1673]. Ocean [360, 933, 243, 860, 2235, 1214, 1208, 2132, 2607, 1715, 1580, 2829, 2897, 2368, 2167, 2585, 2482, 2865, 1644, 405, 1339, 354, 896, 2573, 833, 2325, 1148, 891, 974, 1305, 544, 2904]. Ocean-type [860]. Oceanic [210, 796, 2902, 1530]. Oceanographic [1249, 1101, 2734]. Oceanography [244, 663, 748]. Oceans [2232, 2234]. Octopus [1766]. Odd [480, 1464, 1779]. Odd-Year [480, 1464, 1779]. Odds [1256]. Odhner [1664]. Odobenus [2670, 1206]. odors [1824]. Off [836, 61, 863, 1726, 1115, 756, 360, 1773, 1748, 2077, 58, 137, 473, 791, 2186, 210, 1120, 636, 1867, 1681, 671, 2299, 1496, 1172, 1604, 2906, 1528, 1337, 1848, 2358, 2076, 525, 663, 1634, 579, 2555, 1161]. Off-Flavor [61, 836]. officinalis [926]. offs [2468, 874]. Offshore [473, 1117, 457, 2043, 1254, 1388, 1784, 1491, 2048, 2290, 2905, 2160, 2044, 2047, 2045, 2050, 1848, 2358]. Offspring [330]. Ohio [777, 1326, 2409, 2114]. Oil [268, 811, 1245, 978, 2865, 2021, 2619, 2048, 2023, 2026, 2274, 2700, 2022, 2447, 2025, 1774, 2775, 2776]. oil-laden [2447]. Oiled [536, 2024]. Okanagan [284]. Okeechobee [2716]. Oklahoma [449]. Old [701, 565, 634, 2214]. Old-Growth [565, 634, 2214]. Oléron [325, 325]. olfaction [2443]. Olfactory [759, 1363, 2137]. Oligochaetes [801, 920]. Oligotrophic [1128, 259, 567, 82, 403, 1062, 1007, 29, 422, 1440, 42, 971, 2495, 2290, 2206, 2756, 2876, 1742, 1863, 2578, 2808]. Olympic [2527]. Ommastrephidae [2070]. omnivorous [2599]. onboard [2386]. oncogene [1821]. oncorhynchi [1657]. Oncorhynchus [1465, 2465, 2380, 2722, 2867, 488, 22, 474, 1569, 1833, 937, 885, 1565, 1925, 1139, 1819, 2689, 1445, 2132, 2708, 1125, 1244, 2486, 1751, 1987, 2737, 1958, 1121, 1826, 2576, 2865, 1655, 635, 1258, 2248, 835, 1382, 2168, 756, 1638, 1577, 2448, 2177, 2516, 1348, 1315, 2677, 822, 2352, 1540, 15, 2597, 842, 2204, 1794, 944, 170, 405, 148, 464, 1824, 2061, 256, 899, 2276, 1998, 200, 1619, 2459, 999, 431, 699, 582, 2267, 2539, 1224, 2719, 515, 2449, 2140, 2714, 2862, 265, 1988, 2356, 480, 2896, 81, 2227, 890, 889, 2545, 2705, 1479, 423, 1296, 1761, 674]. Oncorhynchus [490, 903, 2661, 2715, 2575, 2871, 1933, 2600, 912, 2595, 437, 364, 389, 1868, 2085, 2526, 1017, 695, 2567, 476, 169, 243, 941, 354, 2221, 789, 896, 1361, 860, 2074, 307, 919, 2119, 53, 2756, 1453, 977, 417, 1362, 2580, 2442, 688, 1282, 2812, 1490, 932, 1682, 2381, 1722, 1342, 774, 2235, 2645, 2741, 2849, 2001, 2134, 204, 284, 1481, 2528, 2444, 778, 2801, 1850, 930, 1838, 716, 715, 247, 482, 2325, 733, 2554, 2156, 2547, 1052, 1799, 300, 1343, 1344, 521, 1791, 472, 2571, 2374, 2334, 2624, 2592, 659, 658, 2627, 1478, 19, 2228, 623, 1214, 1534, 881]. Oncorhynchus [370, 2453, 2638, 1336, 1855, 2759, 2829, 1025, 1940, 2223, 870, 2568, 1200, 512, 399, 2292, 2540, 139, 2774, 1259, 2699, 2178, 1100, 1171, 2398, 2749, 221, 2897, 2348, 950, 2758, 846, 2414, 2559, 1275, 337, 724, 960, 1464, 1779, 98, 1327, 1381, 1380, 2121, 2626, 255, 308, 2353, 1473, 1040, 2354, 1982, 1300, 65, 484, 357, 1222, 242, 1305, 1208, 1184, 823, 996, 1135, 2211, 2212, 2771, 1658, 1463, 1462, 624, 817, 1051, 2167, 602, 978, 348, 97, 555, 341, 1530, 2585, 2482, 441, 1217, 1218, 1865, 1460, 1307, 1475, 327, 250, 1014, 1461, 2032, 1301, 1410, 1923]. Oncorhynchus [2454, 2911, 1515, 1470, 1409]. Oncothynchus [1482, 809]. One [994, 1701, 2779, 2104, 1595]. one-dimensional [1701]. Oneida [1183, 747, 1872, 2870]. Only [282, 804, 300]. Onset [1100, 1046, 744, 1542]. Onshore [814, 2386, 2290]. onshore-offshore [2290]. Ontario [839, 1321, 321, 2884, 453, 747, 2036, 1422, 1811, 318, 1560, 1732, 55, 690, 1231, 2194, 454, 462, 1201, 1437, 1991, 1371, 2377, 456, 452, 145, 1998, 592, 1127, 48, 1350, 1780, 460, 1072, 2296, 837, 816, 457, 1421, 1625, 1419, 566, 43, 2739, 1768, 1392, 1394, 2878, 2659, 323, 2396, 1621, 552, 1630, 1946, 1420, 1895, 1896, 794, 1572, 26, 45, 293, 2369, 2793, 458, 616, 1931, 2228, 538, 884, 2445, 832, 1628, 2416, 2417, 227, 1616, 2291, 1727, 2531, 455, 1980, 2606, 1418, 973, 459, 507, 2357, 1284, 886, 292]. Ontogenetic [494, 948, 784]. Ontogenic [1874]. Ontogeny [499, 106, 2639, 661, 1695, 2849, 1876, 2532, 1935, 2097]. Onychoteuthidae [1123]. Oocyte [168, 1351]. Oocytes [1303]. opaque [1818]. Opecoelidae [1667]. Open [360, 1437, 707, 161, 837, 1251, 623, 1931]. Open-Access [1251]. Open-Water [1437, 707, 161, 1931]. Opeongo [2739, 2291]. operating [2780]. Operation [162]. Operations [2043, 1803, 2051]. Ophiodon [109]. opilio [2432, 1272, 841, 2413, 2763, 750, 540, 1030, 249, 414, 714, 1103, 1571, 2102, 2498, 1136]. Opossum [1357, 2479]. opportunities [2774]. Opsanus [96, 420]. Optical [395, 330, 902]. Optimal [888, 1653, 673, 769, 1719, 2084, 97, 341, 2423, 1867, 2336]. Optimization [229]. Optimizing [742, 574]. Optimum [1431, 1313]. option [2124]. options [2263]. oquassa [1170]. Oral [88, 2791]. Orange [672, 87, 1048, 1521]. Orangefin [987]. Orconectes [1032, 2072, 493, 2492]. Order [151]. Oregon [127, 2749, 756, 1552, 1542, 659, 658, 2627, 863]. oregonensis [2075, 1164, 1165, 2521, 2522, 852, 1281, 564]. Oreochromis [299, 2115, 471]. Organic [761, 318, 412, 634, 785, 1270, 173, 24, 877, 1031, 2330, 572, 2755, 1113, 1153, 728, 1919, 1401, 1018, 2029, 2342, 2706, 2853, 2501, 2800, 2322, 2802, 1945, 2801, 2369, 2851, 2120, 2699, 2531, 1769, 2789, 2605, 1912]. Organisation [1660]. Organismal [1138]. Organisms [697, 1095, 2517]. Organization [1360, 1479, 887, 352, 1551, 1660, 569]. Organochlorine [1024, 459, 507, 2366, 2475, 1508, 1507]. organs [1581]. Orientation [265, 1287, 414, 802, 414]. oriented [1914]. Origin [379, 405, 2185, 552, 1171, 484, 1460, 1590, 1654, 2348, 2212]. Origins [1659, 1068, 2062]. Ormetoprim [88]. Orthodon [1757]. Oscillatoria [467]. osculatum [1671, 1827, 609, 255]. Osmerus [1531, 1823, 1572, 2558, 2651]. osmorégulateur [1660]. Osmoregulation [302, 1955]. Osmoregulatory [306, 704, 1660]. Osmotic [2054]. Ossification [649]. Osteichthyes [1288]. Ostreichthyes [601]. Other [360, 443, 1264, 227, 481, 1936, 1588, 1963, 926, 1669, 2157, 1533, 2574, 2280, 1599]. Otolith [986, 1358, 315, 279, 1173, 87, 1399, 1142, 2553, 1171, 729, 660, 494, 1776, 1546, 1695, 1733, 1802, 2554, 2651, 2091, 1503, 1952, 1515]. Otolith-Based [729, 2651]. Otoliths [246, 1219, 1215, 1177, 277, 2787, 2847, 2298, 2404, 1617, 1618, 2504, 2862, 1988, 2565, 2089]. Otostrongylus [2333]. Otters [899]. ouest [414]. outbreak [2073]. Outbreeding [480]. Outcome [2617]. outdoor [2120]. outer [2299, 2131]. Outflow [1350, 666]. outlet [2090]. Outlook [453]. outmigrating [2463]. Outperform [1454]. Output [311, 866, 2678, 1936]. ova [2591]. ovalis [1992]. ovarian [2572]. ovaries [2104]. Ovary [167, 2432, 1535]. over-winter [1940]. overexploitation [2792]. overfishing [2316]. overgrowth [1771]. overland [1528]. Overlap [545, 734, 2067]. Overview [43, 2043, 1397, 1702]. Overwinter [485, 2630, 2646]. Overwintering [959, 631, 1444, 1374, 1178, 2844, 2855]. Ovigerous [949]. oviparous [2186]. Ovulation [168]. ovulations [2286]. Ovulatory [138]. owing [1634]. Oxidase [1298]. Oxidases [333]. Oxidation [2094, 136, 2389, 2254, 2036]. Oxidative [1143]. Oxolinic [1282]. Oxygen [1648, 1182, 149, 591, 20, 1320, 830, 898, 2370, 2303, 1877, 2152, 2817, 2563, 2151, 1938, 2035, 2405, 2379, 2562]. Oxygenase [743, 2598, 1960]. oxygenation [2309, 2308, 2306, 2307, 2305, 2302, 2301, 2304]. oxyrinchus [1843]. oxytetracycline [2440]. Oyster [407, 1322, 906]. Oysters [141, 1469, 325, 2490]. Ozark [705, 2672, 2343].

p [378, 1111, 1390, 1374, 2573, 633, 2345, 506, 422]. P. [2290, 1417, 1503]. P450 [1194]. P4501A [1525, 2791, 1254]. P4501A1 [1903]. Pääjärvi [1950, 1951]. Pacific [405, 354, 896, 833, 2325, 1148, 569, 974, 1305, 544, 2904, 1778, 141, 2788, 979, 1445, 1544, 2709, 1076, 1776, 1826, 2865, 1322, 2738, 1655, 1609, 2021, 477, 2780, 2619, 2684, 2597, 1058, 1613, 2031, 2276, 2406, 2614, 1224, 2601, 168, 167, 2124, 1339, 1580, 1210, 266, 1711, 1537, 2916, 1471, 14, 1122, 2024, 2026, 1220, 2645, 2490, 2915, 2299, 2573, 930, 2274, 2700, 2914, 677, 882, 2022, 1902, 2374, 2040, 1534, 1148, 2726, 1312, 1458, 1198, 2767, 551, 1037, 1464, 1779, 244, 650, 755, 2399, 525, 974, 2368, 216, 537, 823, 111, 1480, 264, 2196, 74, 555, 1530]. Pacific [931, 2826, 1470]. pacifica [2856]. pacificus [2131]. Pagrus [2781]. PAHs [2720]. Paired [1035]. palaeosalinity [1574]. Palearctic [1472, 1632]. paleoecology [2853]. Paleolimnologic [1929]. Paleolimnological [1780, 584, 1397, 577, 455, 1798, 2765]. Pallas [1928]. pallasi [1445, 2021, 2619, 2031, 168, 167, 266, 1471, 2024, 677, 882, 2726, 56, 2767, 551, 1037, 974, 2368, 216, 264]. palp [1589]. Pamunkey [1957]. pan [2883]. pan-Atlantic [2883]. Pancreatic [1221]. Pandalid [636]. Pandalus [368, 665, 349, 933, 1715, 1580, 636, 2464, 231, 612, 1176, 171, 92, 574, 324]. panel [2857]. Panulirus [2122, 2231, 152, 2481, 1049]. Paper [1384]. papers [2145]. papillomas [1889]. Paradigm [1424, 2179]. paradigms [2182, 2180]. paradoxa [1288]. paradoxus [522]. Paralichthys [1984, 775]. Paralithodes [1086, 219, 465, 345, 150, 949, 1600, 2201]. Paramecium [130]. Parameter [994, 2462, 57, 74, 2293]. Parameterizations [524]. Parameters [667, 285, 160, 362, 32, 1196, 1426, 363, 1233, 2831, 2171, 2517, 2185, 2420, 2748, 2512, 2211, 2628, 2425, 2196]. Parasagitta [1835]. Parasite [405, 778, 2422, 1679, 2073, 1685, 2135]. parasite-induced [2422]. Parasites [1106, 1686, 265, 1681, 1682, 1285, 607, 1683, 1968, 2486, 1666, 2163, 1684, 1685]. Parasitic [425, 1602, 1677, 1992]. parasitizing [1656]. Parasitological [1680]. paratoluenesulphonamide [1855]. Parent [928]. Parent-Age [928]. Parentage [2459]. Park [2063, 2340, 1599, 1036, 2772]. Parr [1125, 396, 132, 1355, 835, 1285, 326, 1067, 523, 2027, 1834, 1770, 2615, 2626, 2668, 2667, 2809]. Part [765, 2359, 1002]. Partial [394, 1333, 2446, 947, 57]. Participating [249]. Particle [1231, 276, 144, 1757, 2474, 2147]. Particle-free [144]. Particle-size-conversion [1231]. Particles [1294, 82, 144, 930, 1911]. Particulate [761, 1167, 675, 1652, 984, 608, 154, 2893, 1517, 2164, 1757]. Partitioning [2570, 375, 256, 310, 2610, 2611, 62, 2230, 2778, 2731]. parvum [546]. Passage [77, 772, 1000, 1514, 1838]. Passages [356]. Passing [495]. passive [2535]. Past [1070, 388, 2445]. Paternal [417]. paternity [2498]. Path [995, 1334, 1483]. Pathogenic [992]. pathological [2281]. paths [2319]. Pathways [454, 1978, 1993, 1545, 1661]. Patinopecten [994]. Patophrys [481]. Pattern [691, 822, 855, 1186, 231, 31, 1005, 1958, 2522, 2739, 2834, 2467, 2329, 1971]. Patterns [508, 371, 573, 2264, 1448, 239, 1213, 1225, 37, 861, 1416, 297, 791, 1321, 1443, 469, 71, 932, 774, 34, 641, 1202, 1264, 1385, 838, 1113, 2638, 1458, 2759, 1025, 1109, 356, 1177, 92, 461, 537, 996, 624, 812, 821, 2859, 2727, 2095, 1781, 1983, 2199, 2647, 2889, 2656, 2500, 1939, 2028, 2044, 1863, 2572, 1692, 1621, 2530, 2665, 2758, 1862, 2512, 2167, 2911, 2480]. Pb [17, 1365, 207, 1742, 66, 2704, 1398]. Pb/ [17]. PCB [474, 690, 1768, 323, 1240, 227, 1300, 2191]. PCB-Containing [474]. PCBs [801, 318, 1903, 2720, 85, 875, 1760]. PCR [944, 1764, 2490, 1099, 2016]. PCR-Amplified [1099]. PCR-based [1764]. PCR-SSCP [2490]. PDA [2408]. peak [2159]. pealei [208]. Pearl [1374]. Peat [637]. Pectinidae [916]. Pectoral [535, 2854]. Pelagia [1422]. Pelagic [380, 1201, 390, 880, 1155, 636, 796, 747, 339, 227, 1337, 2405, 289, 541, 804, 1491, 2320, 2305, 2792, 1951, 1768, 2702, 2906, 2068, 1628, 2417, 2902]. pelamis [402, 1668, 1426, 2748]. pelicans [2238]. Penaeidae [1238]. Penaeus [1238]. pengoi [2659]. Peninsula [957, 2527, 2111]. Pennsylvania [1193, 305]. Penny [123]. Pens [898]. Pentachlorophenol [1238]. Perca [220, 921, 2766, 342, 416, 2364, 2392, 101, 2242, 2469, 2430, 48, 106, 2384, 358, 278, 843, 202, 351, 63, 1411, 2169, 1872, 2870, 2808, 502, 2280, 1068, 2135, 948, 1935, 784, 2775, 2776]. Percentage [2835]. Perch [220, 921, 416, 101, 48, 106, 816, 358, 278, 843, 202, 351, 63, 1411, 502, 1068, 948, 784, 2766, 2364, 2392, 2242, 2469, 2430, 2384, 1946, 2169, 1872, 2870, 2808, 2280, 2135, 1935, 2775, 2776]. Percidae [126]. Percids [342]. Performance [141, 835, 1382, 1452, 253, 1138, 1344, 1218, 943, 1313, 2275, 2465, 1958, 1613, 1787, 2545, 2705, 2725, 2508, 2509, 2387, 2224, 2456, 2593, 1721, 1910, 2032]. Period [675, 984, 1653, 2344, 2498]. Periodic [355, 2744]. périodicité [719]. periodicities [2243]. Periodicity [1209, 525, 719]. Periods [931]. Periphyton [73, 174, 496, 2713, 855, 257, 705, 1007, 758, 519, 506, 2246, 298, 2149, 2468, 2577]. Perlid [861]. permeability [1761]. permit [1861]. Peroxide [102, 1769, 2353]. Persistence [1245, 1868]. Persistent [1144, 1785]. Perspective [452, 451, 1698, 2050, 1702]. Perspectives [1733, 568, 1114, 2848]. perturbation [2569]. perturbations [2050]. Peru [765]. Peter [2827]. Peters [1954]. Petersen [2414]. petroleum [2024, 2274, 2205]. Petromyzon [1602, 1006, 373, 1369, 1188, 1255, 1592, 1980, 201, 1132]. PGM [1293]. PGM-1r* [1293]. pH [1969, 1713, 1139, 510, 2192, 421, 989, 2484, 1349, 2061, 1871, 592, 877, 1296, 1761, 251, 2296, 180, 183, 879, 273, 2886, 2800, 1179, 1963, 1945, 704, 572, 782, 185, 1156, 194, 153, 1158, 1239, 1018, 1912, 333, 327]. PHABSIM [2714]. PHABSIM-predicted [2714]. Phaeopigment [826]. Phalacrocorax [2247]. Phantom [119]. pharao [555]. Pharmacokinetics [688, 1282, 61]. Phase [87, 2043, 796, 2443, 575]. phenetic [1662]. Phenotypic [1607, 431, 1442, 30, 357, 1102, 2882]. Phillip [1815]. Philometra [1676]. Philometridae [754, 1676]. Philonema [754, 1672]. Phoca [1020, 1019, 2333, 676, 1023, 27, 562, 2174, 531, 2069, 114, 1806, 1130, 1762]. Phocascaris [1671, 1827]. Phocid [142, 355, 2166]. phocids [2470]. Phocoena [1181, 64, 1598, 1915, 2671, 1949, 1584, 558, 240]. Phosphatase [447]. Phosphate [285, 706, 637, 422, 940, 2579]. Phosphate-Poor [940]. Phosphate-Rich [940]. phosphores [1653]. Phosphorus [638, 174, 38, 1437, 1370, 124, 145, 165, 678, 1294, 509, 1374, 16, 648, 1419, 637, 52, 447, 286, 857, 51, 2505, 1062, 1202, 1203, 1420, 293, 261, 519, 884, 305, 429, 1789, 858, 971, 1267, 1849, 1950, 2618, 1892, 2893, 1961, 1516, 2511, 2079, 2703, 1738, 1965, 1759, 1895, 1896, 2881, 2838, 1595, 2674, 2302, 2351, 2278, 2123, 2823, 1931]. Phosphorus-Limited [293]. phosphorus-loading [1965]. phosphorus-rich [1892]. Photo [136]. Photo-Oxidation [136]. photoadaptive [2256]. photoautotrophic [2408]. Photochemical [1769]. Photololigo [1243]. Photoperiod [597, 819, 560, 1369, 79, 798, 430, 529, 693, 1132]. photoperiodically [2892]. Photoperiods [523, 1504]. photoreduction [1650]. photosynthate [2460]. Photosynthesis [840, 1420, 1436, 972, 2146, 2256, 1895]. Photosynthetic [1180, 1112]. Phototrophic [613, 36, 2455]. Phoxinus [2290, 800, 1284, 1417]. phycotoxin [1533]. Phylogenetic [1472, 126, 2276, 113, 1099, 2239, 1288, 2804]. Phylogenic [1632]. Phylogeny [2894, 1902, 828]. Phylogeographic [2834, 2062]. Phylogeography [2489, 2734]. Physeter [579]. Physical [454, 942, 1332, 448, 2111, 2112, 728, 1029, 663, 748, 1694, 2610, 2611, 2493, 2145, 1554, 1717]. Physicochemical [469, 1264, 1402, 2703]. Physics [10]. Physiological [1990, 1116, 698, 699, 2148, 168, 855, 482, 262, 1959, 1406, 1259, 1739, 1217, 1218, 327, 2129, 2735, 1707, 1839, 1807, 2633, 2547, 2699, 2775]. Physiology [1217, 419, 2128, 513]. phytobenthos [1588]. Phytofaun [681]. Phytoflagellate [546]. Phytoplankter [1410]. Phytoplankton [365, 1290, 412, 38, 2484, 1060, 1110, 706, 259, 165, 840, 394, 1441, 1439, 260, 1394, 1403, 1895, 45, 120, 820, 1386, 1570, 290, 846, 1436, 1112, 298, 895, 2395, 2599, 1489, 157, 1871, 1977, 2256, 2460, 2283, 2893, 2643, 2703, 2147, 1896, 1587, 2338, 2304, 2114]. pico [2033]. pico- [2033]. picocyanobacterial [1650]. picogram [2729]. Picoplankton [1180, 613, 403, 971, 1485, 1486]. Pigment [335, 136, 1403, 1138, 2853, 2408, 2195]. Pigments [54, 1159, 1394, 993, 2147, 2755, 2665, 2455]. Pike [590, 1373, 1081, 292, 2242, 2469, 2876, 2884]. Pikeperch [670]. Pilchard [279, 1546]. pilchardus [1807]. Pilot [1157]. Pimephales [1199, 135, 2391, 2227, 1482, 1207, 1017, 1643, 1113, 1018, 1912]. Pink [480, 1453, 716, 715, 19, 1464, 1451, 978, 441, 1470, 2865, 2088, 2449, 2356, 2595, 2414, 1779]. pinniped [2089]. Pinus [2707]. pipiens [23, 24]. Piroplasm [1257]. Pisces [215, 1992, 1676]. Piscicola [1655]. Piscine [237]. Piscivore [702]. piscivores [1639, 2782]. piscivorous [2556]. Piscivory [670, 479, 1338, 1364, 2532, 2790]. Pituitary [1139]. Pituitary-interrenal [1139]. placement [2176]. Placing [1256]. Placopecten [2039, 1749, 1071, 1921, 2173, 1783, 1374, 1001, 1197, 802, 928, 916, 734, 1554, 748]. Plaice [911, 1406, 1046, 1257, 850, 1670, 1790, 2197, 2642]. Plain [9, 371, 2618, 1483, 1825]. Plains [1070]. Plaisance [1026]. Plan [435]. Planktivore [707, 198, 702, 731, 41, 1755]. Planktivores [380, 731, 199]. Planktivorous [578, 1811, 2067, 2081, 1897, 2068]. Planktivory [1572, 2843, 1628]. Plankton [285, 1536, 598, 162, 1295, 1279, 942, 1159, 51, 731, 686, 1149, 1337, 507, 10, 895, 1811, 1950, 2136, 2225, 2701, 2148, 1873, 1961, 2605, 2794]. Planktonic [1167, 696, 877, 889, 748, 1368, 1601, 2198, 1622, 1766, 1847]. planktonio [1560]. Planning [2584, 229, 824, 433]. Plant [54, 1230, 93, 1166, 898]. Plasma [1925, 835, 271, 172, 560, 1444, 169, 2385, 1333, 1131, 654, 743, 624, 2298, 1535, 1496, 2854, 1610]. Plasmid [891]. Plastic [968]. Plate [473, 1693]. platessa [1046]. platessoides [911, 1670, 1406, 1257, 850, 1790, 2197, 2642]. platform [2205]. platforms [2048, 2046, 2044, 2047, 2045]. platyceros [368, 324]. Platyhelminthes [2273, 2436]. plea [1709, 1710, 1877, 2152, 2151, 1708]. Plectropomus [2764, 2060, 2480]. Pleistocene [656, 611, 655, 1724]. pleuriticus [635, 1923]. Pleuronectes [1384, 2732, 2404, 2286, 1254, 1774, 1046, 811, 1960, 1298]. plicata [2129]. Ploidy [1456, 1455]. plotters [2551]. plumbeus [732]. plumchrus [2573, 614]. plume [1544, 1956]. plumes [2887]. Podding [219]. Poecilia [1108]. Poeciliidae [1772, 1901]. Poeciliopsis [1901]. Poecilostomatoida [1677]. Point [162, 161, 1059, 2005, 2753, 2764]. Points [1175, 568, 2437]. polar [1762]. Policies [15, 226, 341, 1510, 2653]. Policy [769, 2233, 2234, 2319]. politics [2233, 2234]. Pollock [287, 606, 1983, 1519]. pollutant [2876]. Pollutants [318, 1144, 1785]. Polluted [1143]. Pollution [586, 587, 416, 1270, 2531, 2135]. Polychlorinated [652, 651, 1903, 2501, 2664, 2578, 2496, 2793, 2441, 2836]. polycyclic [2469]. Polymerase [317, 281, 944, 833, 2446]. polymorpha [2810, 1648, 1182, 2728, 1849, 2129, 980, 2216, 2501, 2488, 2723, 439, 865, 2296, 1597, 2103, 2543, 2015, 2253, 1280, 1782, 518, 1520, 2426, 864, 1246, 1264, 1759, 1956, 1928, 1760, 2071, 1387, 616, 1589, 588, 859, 2347, 2870, 1388, 726, 1716, 2363, 440, 1899, 1948, 2280]. Polymorphic [1691, 2847, 2265]. Polymorphida [799]. Polymorphism [470, 297, 2057, 2641, 1843, 2133, 1662]. Polynomial [708, 1331]. Polypterus [1288]. Polystoma [1662]. Polyunsaturated [1301]. Pomacentrus [1503]. Pomatomidae [1992]. Pomatomus [2687, 1695, 1338, 1992, 2556]. Pomfret [1148, 1776]. Pomoxis [502]. Pond [1263, 1111, 1170, 75, 20, 972, 1608, 2280, 1636]. Ponds [299, 1034, 1069, 2067, 2511, 1772, 2120]. Ponto [2659]. Pool [429, 2014]. pooled [1972]. Pools [874, 230, 1190, 2373, 2575, 2604]. Poor [940, 1678, 1727]. pop [2676]. pop-up [2676]. Population [220, 921, 1778, 371, 692, 740, 741, 344, 1348, 1030, 2092, 1145, 854, 2391, 12, 1619, 331, 1981, 1953, 2855, 1468, 197, 1339, 702, 2889, 1277, 2464, 278, 282, 1197, 263, 1469, 2442, 158, 795, 2314, 225, 528, 1003, 236, 518, 833, 928, 2325, 1118, 640, 438, 2627, 2752, 30, 765, 588, 1137, 177, 615, 1451, 70, 121, 1524, 850, 366, 1305, 1081, 42, 1049, 1233, 1256, 1634, 1835, 579, 812, 558, 1205, 2319, 2309, 2117, 1607, 2057, 1498, 534, 1998, 2190, 1629, 1803, 2397, 2141, 2862, 2521, 2171, 2428, 2010, 2183, 1783, 1843, 2816, 2420, 1625, 2188]. population [2063, 2849, 1785, 2083, 1520, 1662, 2231, 2222, 2051, 2300, 2900, 2506, 2073, 2681, 1904, 2169, 1484, 2331, 1894, 1585, 2060, 2390, 2348, 2808, 2017, 1712, 1869, 2559, 1899, 2211, 2892, 2042, 2628, 1737, 1600, 2775]. population-based [2892]. population-level [2169]. population-phenetic [1662]. Population-specific [2442, 2325]. Populations [60, 1472, 634, 347, 656, 416, 1322, 740, 994, 797, 611, 1371, 427, 1213, 309, 680, 876, 485, 547, 813, 1196, 329, 877, 1023, 465, 385, 1185, 776, 282, 1009, 1457, 1285, 34, 1299, 802, 601, 1237, 300, 843, 1478, 478, 226, 1025, 517, 1412, 1464, 153, 1473, 621, 65, 484, 1222, 242, 1284, 446, 1462, 324, 849, 2766, 1970, 2027, 1834, 1632, 1644, 2177, 2421, 1856, 2459, 2792, 2519, 2085, 2122, 2616, 1837, 2886, 1881, 2580, 2297, 2821, 2613, 2573, 1563, 1880, 1691, 1635, 2629, 1497, 2228, 2558, 2402, 1598, 2870, 1924, 2758, 2559, 1779]. populations [1989, 1806, 2084, 1762, 1891, 2911, 1879, 1878]. popylations [1641]. Pore [1064, 752]. porewater [2048]. porphyrin [1605]. Porpoise [1181, 558, 1598, 1915, 2671, 1949]. Porpoises [64, 240, 1584]. Port [1815]. Portunus [831]. Position [228, 646, 2716, 2475, 2203, 1853]. Positioning [1209, 2551]. Positions [792, 793]. Possibility [421]. Possible [474, 155, 1277, 896, 2094, 2218, 1549, 2812, 2491, 2755, 2100, 2806, 1832]. post [2407, 2668]. post-season [2407]. post-young-of-the-year [2668]. Postembryonic [694]. Posterior [1233]. postexercise [2128]. Postfire [2340]. Postglacial [2477, 1724, 2489]. posthatch [1548]. Postlarvae [505]. Postlarval [1238, 1069, 1192, 916]. postmaturation [2771]. Postmetamorphic [864]. Postmolt [1098, 1341]. Postsettlement [1744, 2632, 1815, 1688]. Postsmolt [275, 2750]. pot [2275]. Potamopyrgus [2162]. Potential [1231, 897, 435, 1767, 752, 306, 1056, 879, 1423, 7, 350, 112, 6, 443, 1203, 95, 640, 616, 859, 2845, 2436, 277, 581, 558, 2637, 1590, 1814, 2315, 2639, 2122, 1683, 1678, 2561, 2554, 2222, 2169]. Potentially [260]. pour [534]. Power [680, 758, 1, 1888, 1887, 1831, 2568, 2551]. practices [2509, 2550]. Prairie [1279, 820, 1386, 39, 1654, 2391, 1582, 1570, 2110]. Prairies [1642]. Prawn [368, 1262, 324, 2551]. Pre [2407, 171]. Pre- [2407]. Pre-Recruit [171]. Precambrian [1726, 1605, 1647, 1404]. precautionary [2650, 2541]. Precipitation [176, 2053, 269, 1500, 1595, 2402]. precise [2837]. Precision [392, 680, 934, 2362, 2298, 2202, 2200]. Predaceous [861]. predacious [2843]. Predation [1325, 909, 375, 2238, 1032, 899, 1416, 1021, 702, 1255, 871, 307, 1009, 1342, 1008, 640, 852, 1281, 589, 2774, 878, 2556, 1980, 146, 381, 1092, 1638, 2687, 2516, 1922, 2365, 2450, 2141, 2522, 1553, 1944, 2782, 1850, 1946, 1687, 2547, 1576, 2058, 1908, 2811, 1765, 2090, 2086, 2646, 2383, 2084, 1736]. Predator [302, 1428, 130, 1629, 1038, 890, 1541, 2156, 1385, 746, 2907, 1011, 2695, 2465, 2075, 2136, 1751, 2448, 2782, 1908, 2556, 1721]. Predator-induced [2907]. Predators [1145, 696, 119, 443, 56, 644, 2039, 2581]. Predatory [83, 1640, 825, 59, 1413, 400, 2710, 2659, 2768, 1685]. Predict [501, 859, 383, 1967, 2098, 2295, 2119, 2812, 1594, 2506, 1909, 2123]. Predicted [1582, 2714, 95]. Predicting [1125, 435, 108, 2523, 448, 1280, 830, 588, 1620, 2560, 1694, 1653, 2206]. Prediction [293, 1886, 1877, 2152, 2151, 2793, 2416, 2769]. Predictions [331, 1119, 2507, 1202, 1411, 1556, 1892, 1965, 2341]. Predictive [294, 2796, 1831]. Predictor [197]. Predictors [442, 941, 185, 107]. Predominant [369]. Preexposure [511]. Preference [1125, 1124, 2130, 2796, 1793]. Preferences [437]. Preferenda [442]. Pregnancy [64]. prehatch [1548]. Preindustrial [585, 2538]. Preliminary [1455]. premature [2066]. Premolt [1341]. Preovigerous [693]. Prerelease [925]. près [2884]. preschooling [2196]. Preseason [1315]. Presence [639, 212, 743, 57, 2572, 1934, 2801, 2704]. Presence-Absence [212]. Present [459]. presented [1687]. Preservation [970, 335]. preservatives [2874]. presmolt [2119]. prespawning [2619]. Pressure [871, 1553, 2845]. Pressures [149]. pretreatments [2719]. Prevalence [405]. prevention [2353]. Previously [691, 1185]. Prey [1061, 662, 1428, 101, 130, 2349, 1038, 578, 917, 871, 1054, 1009, 91, 112, 882, 1526, 1385, 351, 1274, 760, 359, 589, 2417, 1585, 502, 1011, 985, 948, 581, 2136, 2377, 2392, 2391, 1629, 2242, 2450, 2756, 2662, 2782, 2515, 2240, 1908, 1628, 2556, 1685, 2783, 2339, 2214, 1494]. Prey-Size [359]. preying [2479]. Pribilof [961]. prices [1861]. Primary [1167, 412, 217, 629, 832, 1029, 2495, 2206, 1873, 1866, 1723, 1897, 1570, 2145, 2189]. primer [2244]. Primiparous [1103, 2102]. Primitive [359, 352]. Primor’e [1184]. Prince [2021, 2024, 2026, 2274, 2700, 2022, 2025, 2726, 2767]. principles [2360, 1631, 2361]. Prior [1204]. priority [2560]. pristine [2324]. Prix [410]. Prize [410]. probabilistic [2751]. Probabilities [741, 1964]. Probability [1355, 1233, 2886, 2769]. Probe [464, 1040, 1475]. probes [1981]. Problem [1249]. problems [1803, 1930, 833, 2101]. Procedure [790]. Procedures [1432]. Process [1186, 1158, 2549, 2819]. Processes [102, 1434, 641, 1950, 2631, 2346, 2610, 2611, 2889, 1682, 1842, 2578, 2426, 2145]. Processing [997]. Procladius [413]. produced [1569, 1287, 2212, 1818]. producers [1866]. Producing [604, 2046]. Production [556, 1167, 411, 1004, 1115, 979, 1445, 2132, 1231, 573, 38, 598, 223, 779, 1613, 580, 217, 876, 613, 631, 851, 1091, 1346, 206, 648, 567, 802, 653, 1447, 640, 2120, 869, 1013, 1147, 975, 117, 820, 1386, 767, 1422, 1130, 381, 235, 544, 883, 895, 2648, 2709, 2362, 2215, 2495, 2048, 2412, 2106, 2391, 1582, 2855, 2428, 1999, 2044, 1846, 2115, 2047, 2433, 2045, 633, 2825, 1897, 2050, 2417, 1543, 1938, 1570, 2145, 2388, 2113, 2609, 1524, 2323, 2608, 2196, 1745, 2189, 2555]. Productive [1110, 1543]. Productivity [412, 342, 217, 904, 629, 837, 832, 1029, 2117, 2427, 1892, 2206, 1873, 1553, 2160, 1625, 2251, 2335, 2758, 2537, 2815, 2911]. Products [328, 944, 1594]. productus [2406, 1580, 2299, 244, 650, 755, 525, 74]. profile [1605]. Profiles [1023, 904, 830, 1925, 1753, 1877, 2152, 2151, 1806, 2195]. Profiling [891, 2591]. profitability [2556]. Profundal [728, 2306, 2307]. prognosis [2207]. Program [229, 1091, 43, 745, 2075, 2399]. Programs [824, 574, 1480, 441, 2584]. Progress [462, 1214, 2101]. Progression [668]. Project [179, 510, 893, 2301]. Projected [980, 440]. Prolactin [1363, 2385]. Proliferative [1657]. Prolonged [887, 550, 2725]. promelas [1199, 135, 2391, 2227, 1207, 1017, 1643, 1113, 1018, 1912]. prone [2774].

Propagation [767]. Properties [360, 395, 1269, 210, 2564, 2389, 2254, 1483]. Proportionality [684]. proposal [1852]. proposition [1852]. propulsive [2473]. prospects [2101]. Protease [970]. protectant [2227]. Protection [433]. Protective [2677, 2633]. protects [2699]. Protein [818, 776, 1113, 1821, 2683, 1564, 2592, 2453, 2292, 2540, 1865]. Proteins [486, 1297, 1333, 1288]. Proterorhinus [1829]. protistan [2160]. proto [1821]. proto-oncogene [1821]. Protocol [509]. Proton [977, 2298]. proton-induced [2298]. Protopterus [1288]. Protozoan [905]. provide [2559]. provides [2774]. Provinces [527]. Provisioning [168]. Proximate [548]. Prudhoe [1250, 2199, 2536, 2024]. Prymnesium [546]. Pseudo [2173]. Pseudo-nitzschia [2173]. Pseudocalanus [696]. Pseudocarcinonemertes [302]. Pseudocoelomic [951]. pseudoharengus [766, 1517, 794, 1572, 538]. Pseudopentaceros [833]. Pseudopleuronectes [494]. Pseudotaeniacanthus [1677]. Pseudoterranova [486, 1813, 77, 951, 799, 630, 1675, 2065, 466, 870, 255]. Pseudotertanova [1674]. Pseudotetranova [1673]. ptomelas [1482]. Ptychocheilus [2075, 375, 1638, 1164, 1165, 852, 1281, 564]. publications [49, 187, 319, 463, 619, 763, 1353, 915, 1066]. publications/Corrections [49, 187, 619, 1353, 1066, 319, 463, 763, 915]. published [2259]. Publishing [2588]. Puerto [1177]. Puget [703]. pulex [589, 2582]. Pulp [1384, 278, 590, 2494, 2370, 2346, 2344, 1610]. Pulse [785, 2637, 2888]. Pulsed [1343, 1704]. Pulses [664]. Pumping [149, 980]. Pumpkinseed [1213, 485, 1054]. Pumps [1340]. punctata [1815]. punctatum [1627]. punctatus [836, 1282, 61, 88, 1926]. pungens [411, 580, 1004, 386, 604]. Puntledge [1657]. Pup [851, 2428, 1130]. Pups [559, 1937]. Pure [406]. Purine [36]. Purine-N [36]. purse [2780, 2698]. purse-seine [2780]. purse-seiner [2698]. putative [2623]. Putrescine [842]. Pyganodon [1550, 1551, 2704, 2731]. pygmaeus [866]. Pyhäjärvi [969, 2081]. Pyramid [984]. pyrene [96, 2791]. Pyrethroid [1207].

Quadrats [934]. Quagga [1118, 1528, 2331, 1716, 2885, 1309, 1310]. Qualitative [2371, 1255]. Quality [2125, 468, 924, 433, 805, 2835, 2673, 1780, 2272, 2785, 2864, 2286, 2507, 1630, 2302, 2891, 2769, 2775]. Quantification [2318]. Quantify [1342, 2735]. Quantifying [2516, 707, 2335, 1808, 2479, 74, 1593]. Quantitative [900, 1045, 2595, 1169, 1592, 1306, 2445, 1425, 388, 945, 1958, 2352, 2371, 1574]. Quantity [1104, 2527]. Quasi [1094]. Quasi-likelihood [1094]. Quebec [1196, 787, 1752, 2868, 2869, 1026, 2868, 2869, 785, 1697, 534, 2797, 1753, 939, 1936, 2818, 894, 1754, 1274, 2558, 120, 11, 2033, 2164, 2312, 534, 173, 1507]. Queen [1901]. Queensland [1262]. quelques [926]. Quinte [816, 1420, 884]. Quota [628, 824, 2399, 2653]. Quotas [947]. quotients [1594].

Ra [2787, 17, 1398]. Race [655]. radiata [1801, 1801, 1486]. radiated [2401]. radiation [2146, 2760, 2148, 2256, 2147, 1769, 2789, 2640]. Radio [62, 2465].

Radio-Labelled [62]. radiocarbon [1615]. Radiocesium [1402, 1540]. Radioisotope [718]. radiolabelled [2072]. radiometric [2788, 2787, 1521]. Radiophosphate [599]. Radiotracer [261, 2284, 1858, 2285]. rail [1669]. RAIN [893]. Rainbow [488, 22, 885, 1139, 842, 170, 148, 1477, 1058, 431, 699, 515, 81, 423, 1296, 674, 490, 1017, 476, 1361, 919, 977, 417, 688, 1282, 247, 1052, 300, 1343, 1344, 521, 881, 1336, 870, 1200, 512, 399, 1100, 255, 1473, 817, 348, 1217, 1218, 327, 1301, 1409, 2380, 2722, 2867, 1531, 1569, 1823, 2168, 1638, 2448, 2677, 1540, 2061, 1998, 2459, 2719, 2227, 2545, 2705, 1761, 1482, 2661, 2523, 2756, 2381, 2849, 2001, 2134, 2444, 1705, 1704, 2801, 2633, 1572, 1791, 2592, 2228, 2453, 2558, 1855, 2292, 2540, 2699, 2651, 2353, 1982, 2341, 1865, 2032, 1703]. Rainy [908, 1079]. rajids [2814]. Rakaia [1135]. Ram [1577]. Raman [1219]. Rana [23, 24]. ranched [1442, 2591]. Randall [2825]. Random [353, 789, 147, 1000, 2265, 2757]. Range [285, 989, 854, 122, 1367, 314, 374, 93, 1624]. Ranging [625, 1707]. Ranking [69, 792]. Rapid [2882, 1338, 2016, 2261, 2846, 1929]. rapport [633]. rat [1903]. Rate [1164, 922, 1317, 75, 20, 1216, 589, 878, 1177, 668, 1427, 341, 1967, 2678, 2830, 2466, 2620, 2688, 2545, 2098, 2295, 1695, 2407, 2595, 2816, 2119, 2580, 1643, 1850, 2431, 1760, 2900, 2646, 2651, 1718, 1853, 1795, 1952]. Rates [316, 1235, 149, 853, 980, 1323, 362, 309, 332, 1338, 673, 2664, 795, 432, 1246, 572, 997, 916, 339, 869, 64, 383, 1037, 89, 961, 558, 1648, 2728, 2766, 2732, 55, 1609, 2215, 2427, 1629, 1922, 2011, 2795, 2831, 2764, 1839, 1777, 2407, 2779, 2610, 2611, 2103, 2243, 2116, 2496, 2453, 2638, 1906, 1820, 2302, 2758, 2712, 2463, 2383, 2084, 2640, 2187, 2091, 1913, 1972, 2911]. Rather [728]. Ratio [1030, 1390, 1614, 2115, 633, 2556]. Ration [1082, 404, 1362, 525, 2061, 2478, 2596, 2625, 1890]. rations [2705]. Ratios [203, 1192, 330, 293, 422, 2874, 2504, 1873, 1733, 1768, 2703, 2577, 2674, 1806, 2203]. Ray [356, 2298]. Raymond [1905]. rays [2854]. Re [2914, 2456]. Re-assessment [2914]. re-evaluation [2456]. reach [2454]. reaches [2566, 2434]. Reacidification [46]. react [2181]. Reaction [317, 281, 944, 833, 2446, 2748, 2783]. Reactive [286]. reanalysis [2180]. Reappearance [121, 2861]. Reappraisal [598, 1102, 2167]. rear [2640]. Reared [597, 148, 968, 1381, 1380, 523, 1504, 1382, 2448, 1679, 2440, 2211, 1475, 1515]. Rearing [1199, 659, 831, 950, 2732, 1987, 2737, 1857, 2509, 2350, 1658, 1766]. reassessment [1558]. rebuild [2476]. Rebuilding [341, 2383]. Rebuilt [1088]. Recapture [1094, 309, 1331, 772, 1220, 363, 998, 2645, 2300, 2311, 2712, 1800, 2293, 1505]. Receiving [855, 832]. récentes [319, 463, 763, 915]. Recently [710, 1190, 1564, 1928]. récents [49, 187, 619, 1353, 1066]. Reciprocal [199, 2600]. reclamation [2775, 2776]. Recognition [855, 2448]. recognize [2448]. Recolonization [121, 2826]. Recombinant [1163]. Reconditioning [773, 560, 79]. Reconstructed [1253]. Reconstructing [1070]. Reconstruction [2497]. reconstructions [1836, 1798]. Record [1270, 1407].

Recorded [1396]. recorder [2379]. recording [2379]. Records [1006, 1403, 1253, 618]. Recoveries [71, 1913]. Recovery [1235, 925, 394, 468, 301, 1179, 188, 1383, 620, 1037, 563, 182, 1916, 1096, 1871, 2831, 1844, 2604, 2396, 2806, 2246, 1745, 1917]. recreational [2708]. Recruit [1262, 568, 526, 171, 2117, 2181, 2916, 2914]. Recruitment [1076, 692, 344, 1252, 477, 334, 1212, 685, 391, 241, 1160, 1224, 435, 1119, 933, 1443, 1351, 595, 59, 864, 796, 928, 1044, 68, 489, 627, 764, 947, 150, 537, 57, 1092, 1161, 1627, 2257, 1816, 1578, 2898, 2684, 2179, 2644, 2081, 2122, 1786, 2656, 2500, 1711, 1537, 1815, 1722, 2915, 2724, 2561, 2693, 2694, 2524, 1957, 2231, 1947, 1690, 1502, 2180, 2229, 2100, 2863, 2646, 1924, 1914, 2845, 2634, 1626, 1634, 2667, 1600, 2182]. Recruitments [795]. Recurrent [467]. Recycling [14, 1122, 1407, 1849, 2720, 2823]. Red [1086, 219, 465, 345, 603, 1177, 660, 150, 949, 1676, 2512, 1600, 2201, 2275, 1804, 1677, 1901]. Redbelly [1284]. Redberry [2148]. redd [2027, 1834, 2155, 2620, 2714]. redds [2449]. Redeposition [981]. Redescription [1656, 1663]. Redfish [17, 1615]. Redfishes [356, 1968]. Rediscovery [1679]. Redox [879]. Reduce [119, 902]. Reduced [709, 865, 128, 2674, 893, 1697, 2061, 2351, 2704, 1941]. Reduces [1097, 919]. Reducing [2315, 1594]. Reduction [1183, 460, 1390, 158, 506, 1961, 2378]. Reductions [104, 1156, 2505]. Reef [1089, 2060]. Reefs [152]. reevaluating [1920]. reevaluation [1901]. Reference [416, 454, 1175, 568, 1059, 1470, 1642, 2494, 2098, 2661, 1682, 2437, 1916]. Refinement [2354]. Reflect [172, 728, 1177, 2855, 1861]. Reflectance [1062]. Reflection [483]. refractory [2120]. Refuge [1086, 655, 2477, 1724]. Refuges [1089]. Refugia [1090, 1050, 1087, 1088, 2236, 1085]. Regeneration [712, 1738, 846]. Regime [922, 1351, 2709, 2846, 2904]. Regimes [560, 2634]. Region [585, 162, 161, 655, 1034, 439, 1044, 748, 1644, 2797, 2199, 2623, 2868, 2869, 1598, 410, 2868, 2869]. Regional [1500, 35, 7, 350, 1229, 1247, 224, 660, 2586, 2418, 1945, 1598, 2606, 2249]. Regional-Scale [1229, 1945, 2249]. Regions [994, 1996, 2905, 2906, 1599]. Regression [1304, 131, 178, 1800, 1126, 1578, 1892, 2520, 2596, 1558, 1831, 2897]. regulate [2403]. Regulated [563, 1919]. Regulating [479]. Regulation [136, 521, 1413, 1924, 377, 812, 1847, 2694, 1755]. Regulations [1079, 1522]. Regulatory [1293, 2688]. Rehabilitation [460]. reinhardtii [1516]. Reinhardtius [1106, 2271]. Reintroduction [47]. rejection [2681]. Related [1290, 976, 1346, 567, 987, 1009, 542, 1249, 590, 1179, 1015, 734, 171, 376, 784, 2708, 2410, 719, 2367, 2785, 1863, 2818, 2820, 2750, 1689, 1366, 2775, 2776]. Relating [2224, 1519]. Relation [1128, 979, 735, 1140, 739, 696, 1321, 1614, 868, 210, 196, 648, 548, 1351, 286, 1523, 907, 1190, 1264, 110, 1216, 1214, 936, 1257, 663, 748, 575, 292, 201, 2132, 2027, 1834, 2799, 1787, 2283, 2085, 1535, 2185, 1742, 2578, 1588, 2493, 2572, 1621, 2431, 2400, 2752, 2197]. Relations [1870, 342, 2457, 2185, 2420]. Relationship [701, 1445, 853, 398, 692, 1370, 730, 1666, 1058, 861, 1391, 279, 2714, 83, 1873, 1294, 1031, 595, 679, 2453, 1013, 2829, 643, 1046, 244, 1718, 507, 978, 39, 1566, 2861, 1546, 2539, 1714, 1695, 1733, 2524, 2240, 2222, 2881, 2838, 2191, 2357, 1789, 2059, 1503]. Relationships [1167, 1472, 520, 1466, 2614, 2190, 113, 1072, 1232, 1150, 1457, 112, 1099, 1175, 1202, 1203, 882, 747, 1385, 1047, 1083, 1373, 1274, 359, 134, 151, 494, 1762, 1790, 858, 1632, 2194, 2839, 2181, 2265, 1488, 2154, 2639, 2221, 1746, 1722, 1898, 2672, 2724, 2552, 2914, 2554, 2553, 2369, 1558, 1559, 2897, 2239, 2845, 1483, 2836, 2634, 2153, 1600]. Relative [1325, 1226, 38, 1230, 1145, 2397, 349, 251, 2893, 1120, 848, 1446, 1704, 1052, 107, 2768, 1158, 201, 1573, 1751, 2522, 1786, 1705, 1693, 1908, 2456, 2348, 1924, 1862, 1626]. Release [699, 216, 996, 2865, 1509, 1990, 2612, 2831, 2645, 2666, 1989, 2479]. release-recapture [2645]. Released [308]. Releases [261, 2211]. releasing [2286]. Relevance [1391]. Relevant [1089]. Reliability [958, 751, 1786, 1943]. Reliable [246, 772, 2462]. Relict [1277]. relicta [1357, 1811, 2690, 1421, 1625, 313, 112, 284, 2479, 1311]. reliée [719]. Remaining [282]. Rematuration [396, 773, 560]. remediation [2612]. remobilization [1754]. Remote [976, 975, 1236, 1587, 1599]. Removal [1030, 27, 637, 1520, 18, 1971, 2042]. removing [1678]. remuneration [1841]. Renal [88]. Renewable [1376, 1236, 2336]. Renibacterium [328, 2547]. Rent [628]. Repeat [805, 2394]. Repeat-Spawning [805]. Repeats [1474]. replace [2574]. Replacement [1553, 1093, 1008, 874, 2262, 1988]. Replacements [1032]. Replicate [968]. Replicated [427, 441]. Replicates [620]. Reply [2077, 2052, 2154, 2234, 2913, 2152, 2563, 1879, 2127, 2285, 2361, 2880, 2655, 2234]. reporting [2407, 2779]. representative [1705]. Reproducing [578]. Reproduction [474, 48, 1792, 33, 1108, 654, 270, 240, 1538, 719, 2816, 1867, 2217, 2341]. Reproductive [2155, 205, 1329, 1213, 485, 1196, 1022, 106, 1261, 311, 79, 798, 2026, 34, 2104, 866, 303, 907, 927, 1103, 194, 1699, 961, 811, 2461, 2186, 2415, 1777, 2296, 2314, 2344, 2490, 2561, 2900, 2331, 2375]. Requirement [900]. requirements [2412, 2864]. Resampling [1429, 2746, 2107]. Research [462, 1, 2055, 1979, 1887, 2523, 2359, 2652, 1851, 1725, 2513, 2908]. Reserve [1128, 1286]. Reserves [1314, 2327, 2480]. Reservoir [761, 16, 648, 981, 117, 366, 2410, 2594, 2765, 1720, 1754, 2198, 1079]. Reservoirs [561, 448, 894, 107, 2765, 1495, 2164, 2165]. Residence [295, 232, 2645]. Resident [1424, 1171, 1680, 2871, 2385, 1547]. Residual [1073]. Resistance [586, 587, 631, 950, 471, 1655, 2468, 2032]. Resolute [2187]. Resolution [1366, 2738, 2269, 1808, 2339, 1935, 665, 636]. resolve [2641]. Resolving [1029, 2481]. Resource [452, 1376, 72, 1274, 2778, 1236, 1841, 2852, 2560, 1924]. Resources [1128, 1269, 584, 91, 2108, 1702, 2319]. Respect [1292, 2405]. Respiration [1164, 959, 428, 75, 20, 572, 895, 2146, 2495, 1757, 1718, 1835, 2187]. Respiratory [1594, 1855, 2593, 1948, 2134, 2035]. respond [2320]. Response [1199, 174, 2576, 375, 709, 1697, 408, 698, 394, 855, 865, 1396, 912, 1390, 1093, 183, 1625, 1419, 829, 158, 954, 955, 224, 2411, 661, 1253, 45, 188, 1406, 808, 852, 466, 870, 134, 950, 222, 693, 511, 327, 784, 1301, 1726, 2109, 2736, 2087, 1551, 2401, 2315, 2520, 2791, 2277, 2587, 2297, 1895, 2045, 1484, 2902, 2278, 1948, 2455, 2582, 1910, 2537]. Responses [885, 1525, 2192, 271, 664, 773, 198, 576, 1350, 1279, 702, 2081, 257, 583, 79, 2756, 1860, 1342, 272, 1241, 482, 262, 184, 519, 1156, 503, 724, 1405, 41, 2409, 563, 190, 182, 1096, 1314, 1395, 2005, 2484, 1700, 1629, 2148, 1817, 2723, 2043, 2452, 2012, 2047, 2633, 2547, 2340, 1959, 2137, 1739, 2775]. responsible [2889, 1604, 2640]. Responsiveness [2430]. Resting [667, 1217]. restoration [2402]. Restored [232, 2294]. Restricted [1424, 1367, 1092, 1920]. Restricted-Range [1367]. Restriction [754, 379, 60, 944, 828, 2177, 2641, 1843]. result [1867]. resulting [2421, 1975]. Results [2676, 893, 2738, 2049, 1397, 1754]. Resurgence [121]. Resuspension [145, 981, 1754]. Retained [645]. Retardative [1189]. Retention [155, 145, 276, 917, 844, 940, 945, 2570, 2579, 2575, 2440, 1757, 2474]. retina [2041]. Retinoic [2797]. retinoid [1532]. retour [1852]. Retropinna [1828]. Return [1210, 308, 1208, 1852, 2066]. Returns [771]. Revealed [1472, 1305, 60, 1632, 2265, 2133, 2429, 2885, 1575]. Reveals [347]. Review [73, 165, 641, 453, 1105, 1400, 1309, 569, 1826, 1609, 2150, 2259, 2260, 1882, 2725, 2127, 1908, 2652, 2126, 2726]. Reviews [360]. Revised [732, 1656, 2915]. Revisited [762, 1160, 2915, 2086]. revisiting [2630]. Rexea [1684]. RFLP [1813]. RH [2452]. RH-5992 [2452]. Rheotactic [282, 622]. rhinichthidis [1656]. Rhinichthys [1656, 967]. Rhinicthys [985]. Rhinoceros [1076]. Rhizosolenia [13]. Rhombosolea [1142]. Rhythm [415, 138, 2095, 2892]. Ribosomal [470, 828, 1467, 2031, 2804, 2239]. Rich [940, 1892, 2772, 1539]. Richardson [1828]. Richness [238, 468, 95, 2727, 2544, 1621]. Rico [1177]. Rideau [1560]. Ridge [700]. Ridges [916]. Riffles [874]. rights [1953, 2124]. rigida [280]. Rigler [285]. Ring [87]. Ringed [27, 562, 2333, 2174, 1762]. ringens [765]. Riparian [127, 304, 2218, 2214]. Ripeness [264]. Rise [672]. Risk [1325, 1242, 672, 963, 1431, 589, 1433, 1412, 1092, 1480, 2612, 2522, 1850, 1649, 2568, 2774, 2260]. risk-prone [2774]. risks [1687]. River [22, 392, 1565, 1560, 425, 2027, 1834, 104, 710, 371, 2109, 1258, 2248, 271, 1577, 2412, 920, 685, 757, 1297, 779, 1006, 1075, 899, 1707, 2315, 1196, 297, 1804, 469, 539, 1408, 1936, 1193, 860, 99, 1746, 1657, 1122, 85, 1588, 2493, 1481, 2083, 778, 1705, 1704, 1957, 1187, 1083, 2345, 1281, 1702, 2338, 1507, 2749, 2531, 270, 2006, 2463, 1310, 1135, 1379, 39, 1706, 1703, 1514, 1544, 1830, 2373, 2106, 2238, 2010, 2183, 1542, 2821, 2030, 2130, 1687, 1624, 2374, 1723, 2350, 2324, 1919, 1524, 2007, 2728, 1365, 2689, 1445, 1544, 2132, 1180, 54, 591, 635, 2687, 1270]. River [1348, 1212, 1315, 15, 2281, 2204, 1741, 2797, 1619, 1250, 1981, 787, 719, 939, 2888, 1767, 1211, 354, 2376, 2621, 1857, 2656, 2500, 2384, 2294, 2884, 2104, 1802, 2330, 1838, 1801, 1696, 988, 2158, 687, 370, 852, 2558, 2223, 2477, 1771, 1927, 2178, 2414, 2559, 1275, 1899, 1208, 777, 1326, 2409, 2252, 564, 1853, 2167, 2114, 97, 341, 495, 2189, 1923, 2454]. River-specific [2412]. riverbank [1714]. Riverbed [806]. Riverine [94, 2259, 1074, 472, 777, 2784]. Rivers [1465, 807, 785, 229, 329, 393, 272, 1264, 873, 2637, 2494, 2370, 2713, 2377, 1918, 2421, 2714, 1936, 2889, 2864, 1844, 2426, 2796, 1543, 2758, 2535, 1575, 2911, 2457, 2075, 2245]. Rivière [1642, 1562]. Rivulus [1191]. RNA [470, 288, 1192, 677, 1113]. Roach [110, 1638, 2071, 2135, 1935]. Robert [2658, 1954]. Robust [1304, 790, 1578, 2807]. Rock [989, 152, 146, 1049, 2481]. Rockfish [1065, 907, 720, 134, 721, 1845, 2734]. Rocky [1299, 1413, 1700, 1944, 1651, 1599]. Role [2029, 591, 897, 707, 1032, 2542, 24, 2415, 460, 825, 2163, 1114, 2819, 326, 1087, 98, 1102, 1409, 2443, 2677, 2087, 2141, 2630, 2567, 2851, 1938, 2347, 2808, 2343, 1917]. Roles [1440]. Rooted [879, 2836]. Rootwads [98]. rosea [42]. rosethorn [2734]. rosmarus [2670, 1206, 2670, 1206]. rostrata [1212, 2281, 1211, 2733]. Rosyside [423]. Rotifers [1413]. Roughy [672, 87, 1048, 1521]. round [1829]. Routes [1055]. Routine [348]. Røynelandsvatn [1885]. rRNA [1763, 2446]. Ru [2488]. rubida [303]. Rudolphi [255]. Ruffe [749, 2104]. rule [2190]. Run [991, 1355, 191, 2559]. running [633, 2436]. Runoff [1225, 1111, 2635, 2889]. Runs [14, 1122]. Russia [1464, 1779, 1184, 1463]. Russian [2538, 1459]. rusticus [493, 2492]. ruthenus [1288]. Rutilus [2821, 110, 2071, 2135, 1935].

S [1654, 1093]. S. [445, 2294, 786, 356, 1364, 1151]. s.l [1671, 1827]. Saanich [2856]. Sabiefish [175]. Sablefish [653]. Sac [1141, 1303, 182, 1996, 2041]. Sacramento [1565, 375, 1638]. Sagavanirktok [2245]. sagax [477, 1546, 2746, 1210, 1711, 1537, 2916, 2915, 2914, 765]. Saguenay [2818, 1447, 2868, 2869, 2772]. saiar [377]. saida [1535, 2266]. Saint [714, 1103, 719, 2884, 414]. Saint-Laurent [719, 2884, 414]. Saint-Lawrence [719]. saira [2196]. saisonniers [534]. salar [1969, 900, 2027, 1834, 1504, 396, 691, 1770, 379, 2192, 132, 347, 1509, 2373, 1990, 2678, 271, 2784, 2412, 1918, 2421, 499, 773, 94, 470, 2873, 296, 819, 2680, 560, 2332, 1173, 2504, 1452, 1981, 297, 329, 1194, 2230, 1614, 1747, 2054, 273, 2294, 79, 798, 191, 776, 2740, 1499, 2279, 1547, 2834, 1229, 2864, 1785, 430, 330, 954, 272, 2602, 1821, 1691, 2820, 2509, 786, 2553, 1696, 1187, 188, 759, 1363, 2137, 1548, 1047, 2750, 1497, 2016, 535, 1221, 1293, 275, 326, 270, 668, 529, 2436, 1067, 1955, 1890, 2858, 523, 86, 1735, 2591, 575]. salar [2883, 2743, 269, 2899, 2059, 1379, 2128, 163, 333, 1084, 1812]. salaris [2273, 2436]. salina [500]. Saline [1128, 820, 1892, 1527, 2110]. Salinities [911, 195]. Salinity [407, 302, 698, 1070, 1317, 262, 521, 1406, 1044, 2273, 242, 10, 1504, 1617, 1618, 1860, 1542, 2572, 2065, 1716, 2455, 1943, 2097]. Salmincola [1679, 405]. Salmo [1969, 900, 1970, 2027, 1834, 2394, 1504, 396, 691, 1770, 2366, 379, 2192, 2066, 132, 347, 888, 991, 1355, 1509, 692, 740, 741, 2373, 1758, 1990, 2678, 271, 929, 2784, 2412, 1918, 2421, 499, 773, 94, 470, 2873, 296, 819, 2680, 560, 2620, 995, 2332, 718, 1173, 2504, 1452, 1981, 297, 329, 1194, 2230, 1614, 1747, 2054, 2010, 2183, 1720, 273, 2294, 1939, 2028, 79, 798, 491, 191, 776, 2740, 1499, 2279, 1490, 1547, 2834, 1229, 2864, 393, 1785, 430, 330, 954, 272, 2130, 2796, 2602, 1821, 1691, 2820, 2509, 786, 1635, 2553, 1696, 1187]. Salmo [188, 759, 1363, 2137, 2491, 1548, 1047, 2750, 1497, 2016, 535, 1221, 1293, 275, 326, 270, 668, 1989, 529, 2436, 1067, 1955, 1890, 2858, 523, 86, 1735, 2591, 277, 575, 2883, 2743, 377, 269, 2899, 2059, 1379, 2128, 163, 333, 1307, 1084, 2195, 1812]. salmoides [2630, 476, 704, 262, 2790]. Salmon [129, 1465, 900, 474, 937, 979, 1445, 1244, 396, 691, 379, 1121, 132, 347, 1258, 835, 1382, 756, 271, 785, 1348, 1315, 822, 15, 499, 773, 944, 94, 470, 464, 1058, 256, 899, 296, 819, 560, 200, 790, 999, 1442, 582, 1224, 1173, 1452, 265, 329, 480, 1194, 890, 889, 1454, 903, 912, 389, 695, 1091, 169, 243, 809, 941, 354, 789, 896, 273, 860, 79, 798, 191, 776, 1453, 1362, 14, 1122, 932, 1457, 1229, 1342, 774, 1220, 204, 430, 330, 954, 272, 778, 930, 716, 715, 482, 786, 733, 1187, 188, 521, 759, 1363, 472, 1047]. Salmon [659, 658, 19, 623, 1214, 370, 535, 1458, 1400, 1221, 1293, 1025, 832, 275, 1259, 383, 326, 950, 270, 668, 846, 164, 998, 1275, 337, 724, 960, 1464, 923, 98, 1327, 232, 1381, 1380, 529, 1451, 308, 1067, 523, 1040, 86, 1300, 65, 484, 357, 1222, 242, 1305, 1208, 1184, 377, 823, 996, 1135, 269, 1463, 1462, 624, 1051, 602, 978, 111, 97, 555, 341, 982, 1379, 441, 163, 333, 1459, 495, 1460, 1475, 250, 1014, 1461, 1410, 1084, 1470, 1969, 2465, 1833, 1565, 1925, 1819, 2689, 1544, 2132, 2709, 2708, 2027, 1834, 2486, 2737, 1504, 1770]. salmon [2019, 2584, 2192, 1826, 1797, 2576, 2066, 2865, 2607, 1509, 2738, 1655, 1609, 2248, 2373, 1990, 2678, 2784, 2177, 2516, 2412, 1918, 2421, 2352, 2597, 2204, 1794, 1824, 2276, 1993, 2873, 2614, 2680, 1619, 2615, 2332, 2267, 2539, 2504, 2088, 2449, 2140, 2714, 1981, 2862, 2356, 2896, 2124, 1787, 2230, 1614, 1747, 2054, 2715, 1541, 2595, 1868, 2085, 2526, 1549, 1767, 2567, 1861, 2221, 2294, 2074, 2119, 2740, 1499, 2279, 2580, 2442, 1547, 2834, 1682, 1722, 2297, 1542, 2864, 2235, 1844, 2645, 2741, 1785, 1481, 2528, 1678, 2602, 1850, 1838, 1821, 1691, 2820, 2509, 1902, 2554, 2553]. salmon [2156, 2547, 1799, 2387, 1696, 2137, 2571, 2374, 2334, 2624, 1548, 2040, 2627, 2400, 2750, 1497, 2016, 1534, 2350, 2638, 2402, 2759, 2829, 1940, 2223, 2161, 2568, 2774, 2178, 2398, 2749, 2805, 2897, 2348, 2758, 1591, 2311, 2414, 1869, 2559, 1779, 2121, 2626, 2463, 2436, 1955, 1890, 2858, 2593, 1735, 2591, 2252, 2883, 2743, 2211, 2212, 2771, 1658, 2899, 2167, 2059, 1530, 2585, 2482, 2668, 2667, 2809, 2128, 2537, 2815, 2730, 2826, 2454, 2911, 1515, 1812]. salmonicida [2592, 891]. Salmonid [754, 710, 647, 1055, 288, 1091, 99, 827, 1456, 828, 1467, 898, 2259, 1817, 1893, 2696, 2783, 2214]. Salmonids [318, 983, 1424, 197, 228, 793, 1474, 988, 852, 1281, 657, 1455, 564, 1480, 1672, 1613, 2190, 2661, 1657, 2812, 2664, 1493, 1734, 2508, 1791, 1679, 2456, 2210, 2535, 2640]. salmoninarum [328, 2547]. Salmonine [1240, 1013, 381, 2416, 2417]. Salmonines [1024, 1630]. salmonis [2066, 1549, 2074, 1107, 2750]. salmons [1778]. salmositica [1655]. Salt [10, 1970, 1806]. salt-water [1806]. saltatrix [2687, 1695, 1338, 1992, 2556]. Saltwater [1180, 809, 1456, 1380]. Salvelinus [2034, 2380, 597, 667, 392, 2138, 2118, 639, 803, 2029, 179, 1498, 2514, 1967, 2155, 853, 690, 2761, 2225, 1201, 1466, 1646, 1645, 1680, 2318, 1561, 256, 899, 534, 479, 2620, 2795, 2199, 1511, 1324, 2882, 105, 113, 1127, 1996, 48, 1356, 533, 445, 1144, 2337, 1476, 2747, 2385, 1056, 2243, 186, 180, 183, 818, 1255, 2063, 2791, 1837, 384, 47, 1742, 1863, 1170, 158, 1285, 682, 254, 253, 225, 528, 875, 1124, 2467, 552, 2515, 513, 1190, 1564, 122, 338, 26, 2813, 184, 514, 1532, 185, 1467, 2804, 227, 1412, 622, 1980, 2405, 2846, 2606]. Salvelinus [967, 1966, 621, 1063, 1718, 376, 1028, 2743, 1239, 1891, 2783, 511, 378, 1303, 2270, 1151, 2062, 2489, 1307, 181, 182, 1301, 406]. Same [1240]. Sample [614, 1415]. Sampler [316, 496]. Samplers [497, 2157]. Samples [349, 294, 1473, 1562, 2490]. Sampling [1552, 1038, 27, 509, 742, 789, 745, 212, 1429, 758, 526, 115, 1436, 574, 1000, 982, 1136, 2386, 1573, 1514, 2548, 1653, 2202, 1488, 2154, 2371, 1822, 2107, 2681, 1719, 2390, 2854, 2153]. Sampling/Importance [1429]. San [1565, 2493]. Sand [1076, 67]. Sand-Bottomed [67]. Sandbar [732]. sands [2775, 2776]. sapidissima [1641, 772, 2770, 1802, 601, 478, 178, 2326, 2097]. Sardina [1807]. Sardine [477, 1210, 268, 765, 369, 2746, 2641, 1711, 1537, 2916, 2915, 1807, 2914, 2752, 2634, 1634, 1784]. Sardinella [268, 369, 1250]. Sardinops [477, 279, 1546, 2746, 2641, 1210, 1711, 1537, 2916, 2915, 2914, 2752, 765, 2634, 1634, 1784]. Sashin [14]. Saskatchewan [2148]. Satellite [1476, 2676]. Saugeye [1364]. saumon [575]. Saunders [2884, 2884]. saury [2196]. saxatilis [2687, 2328, 1012, 2656, 2500, 84, 1083, 178, 2440, 729, 767, 1853, 873, 2372, 1096, 943, 1314, 1313]. Scale [391, 37, 2140, 329, 16, 1229, 716, 715, 2200, 684, 471, 657, 649, 1135, 2859, 2132, 1958, 2406, 2262, 2615, 1292, 1643, 2115, 1945, 1385, 2798, 2510, 2249]. Scale-dependent [2200]. Scaled [153]. Scales [156, 1012, 1237, 92, 2614, 1881, 1880, 1879, 1690, 2229, 1598, 1878]. Scaling [365, 553, 1142]. Scallop [994, 1374, 1001, 1197, 303, 928, 916, 734, 748, 1874, 2039, 2887, 1749, 2711, 1921, 2173, 1783, 1554]. Scallops [1071]. Scaphirhynchus [2244]. Scattering [365]. scenario [2292, 2540]. Schaefer [883]. Schaller [2911]. scheme [2322]. schemes [1904]. schindleri [1392]. Schnute [1308]. Schoedler [618]. Scholarship [2439]. School [1065, 965, 2506, 1962]. Schools [402, 2698, 2529]. Schwinghamer [2827]. Science [1709, 569, 2439, 1710, 2233, 2234, 2359, 2310, 2127, 2541, 2126]. Sciences [203, 1851, 1708]. scientific [2208, 2908, 2232, 2234]. scientists [2181]. Scomber [315, 814, 1788, 769, 1059, 815]. Scombrus [245, 315, 814, 1788, 1059, 815]. Scope [387, 1218]. Scophthalmus [2459, 1141]. Scotia [443, 2039, 137, 208, 2905, 1292, 272, 2251, 666, 2890]. Scotian [320, 2237, 1526, 1061, 598, 2262, 1932, 2688, 1374, 799, 1579, 1586, 2058]. Scotland [776, 2059]. Scottish [297, 1187]. scour [2259, 1893]. sculpins [2539]. scums [2287]. Sea [888, 991, 1355, 223, 2215, 2861, 159, 1071, 1006, 200, 1442, 35, 373, 1369, 1188, 1374, 1255, 191, 1001, 1197, 932, 393, 236, 954, 928, 1047, 623, 734, 154, 966, 569, 28, 748, 982, 201, 1132, 2039, 1970, 1602, 2066, 1664, 2173, 1667, 1805, 1783, 2074, 1678, 1496, 2411, 2222, 2750, 2752, 2890, 1980, 2512, 2436, 1739, 1086, 416, 2469, 241, 645, 2698, 1487, 2013, 2785, 1677, 2740, 2834, 1785, 2322, 531, 1934, 2108, 606, 1534, 2073, 643, 275, 1046, 2375, 150, 2550, 722, 1519, 974]. Sea-ranched [1442]. Sea-Run [991, 1355, 191]. sea-running [2436]. sea-salt-induced [1970]. seabass [2095, 1496]. seagrass [1815]. Seal [1020, 12, 1023, 562, 851, 559, 961, 725, 1740, 1529, 1777, 2428, 2058, 1937, 1762]. Seals [1019, 1061, 676, 123, 27, 531, 114, 142, 1130, 355, 2333, 2472, 1813, 2879, 1836, 2174, 2069, 2166, 1806, 2226, 2880]. Sealworm [486, 630, 870, 1674, 1675, 2065]. Search [662]. seas [2040]. Season [1437, 161, 600, 444, 337, 1509, 1918, 2764, 2407, 2528, 1931, 2463]. season-wide [2463]. Seasonal [1357, 1365, 667, 1019, 1627, 1602, 1781, 1061, 1991, 94, 1213, 408, 1145, 698, 1196, 83, 1033, 1787, 1262, 251, 437, 1419, 919, 1863, 671, 2130, 1131, 654, 1896, 658, 2686, 1436, 29, 2205, 2642, 357, 1326, 1417, 1539, 693, 554, 898, 1560, 1986, 534, 1535, 1625, 1796, 2503, 1862, 2363, 1960, 1818]. Seasonality [1134]. Seasonally [1222]. Seasons [1477]. Seatrout [336]. Seaward [265, 778, 733]. Seawater [396, 691, 835, 1382, 159, 582, 1453, 1362, 144, 1221, 1259, 622, 621, 1067, 624, 898, 2680, 2119, 1955, 1890]. Sebastes [907, 1968, 720, 2200, 2734, 356, 721]. Sebastolobus [1845]. Sebates [1968]. Secchi [2552]. Second [2589, 565]. Second-Growth [565]. secondary [1773, 2591]. secretion [2364]. Secretory [1163]. Section [1065, 2294, 1807]. Sediment [487, 1435, 2048, 806, 189, 645, 549, 1064, 1045, 428, 1443, 879, 981, 85, 827, 930, 679, 536, 728, 481, 1080, 1359, 821, 2109, 2887, 2835, 2485, 2673, 2093, 1714, 1873, 2015, 2741, 2426, 2875, 2441, 2869, 1719, 2036, 2280]. sediment-bound [2093]. Sedimentary [335, 2079, 884]. Sedimentation [1294, 1159, 1898, 136, 1407, 728, 1976]. Sedimented [1349]. sedimenting [2665]. Sediments [801, 1297, 1117, 133, 1334, 1064, 207, 1396, 752, 335, 1340, 286, 313, 1253, 305, 962, 66, 337, 724, 211, 811, 1393, 1395, 2194, 2193, 2803, 2706, 2612, 2044, 2471, 2447, 2755, 1774, 2538, 1924, 1608, 2836, 2457, 2869]. See [642]. seed [2227]. seeded [1783]. seeding [2039]. Seepage [962, 2471]. Segregation [987]. seiche [926]. Seine [115, 2780]. seiner [2698]. seismic [2011]. select [2554, 1136]. Selected [1318, 789, 659, 1464, 1659, 2888, 1523, 1583, 1779]. Selecting [530]. Selection [4, 22, 639, 751, 101, 999, 437, 793, 2802, 1237, 400, 78, 502, 277, 948, 1092, 50, 2784, 1645, 2349, 1881, 2702, 1579, 1880, 1568, 1879, 1878, 1736, 2668]. Selective [958, 670, 204, 1481, 716, 715, 30, 1741, 2553]. Selectively [578]. Selectivity [1235, 741, 336, 557, 952, 351, 89, 804, 1427, 1136, 2570, 2401, 2337, 2520, 1887, 2785, 2279, 2240, 2078]. Selenate [418, 419]. Selenite [418, 419]. Selenium [1206]. self [2190]. self-thinning [2190]. semelparity [2771]. semifasciata [2117]. senegalus [1288]. sensibility [2661]. Sensing [1236, 1587]. Sensitive [1290, 1316, 834, 2561]. Sensitivity [421, 757, 1431, 1436, 2298, 2690, 2024, 1705, 1704]. Sensitization [870]. Sensory [204]. Separable [263]. Separate [51]. Separated [1222]. Separating [1106, 330]. Separation [1116, 910, 356, 2787, 1845]. Sepia [926]. Sepioids [1423]. Septicaemia [417]. septicemia [2619]. Sequence [317, 281, 1058, 2031, 1763, 2244, 2446, 2239]. Sequences [1476, 1467, 1644, 2894, 2276, 2641, 1902, 2804]. Sequencing [1099]. Sequential [344, 645, 263, 1330, 1803, 2051]. Serial [77]. Series [344, 839, 614, 594, 28, 1722, 2297, 1502, 1559, 1626]. Serranidae [1676, 2060, 2480]. Serum [84, 1363, 438, 1336]. serve [2310]. Serving [36]. Seston [1062, 2015, 2729, 2875]. sestonic [2672]. set [1943]. Sets [32, 747]. Setting [947]. Settlement [1246, 2131, 2887, 2519, 1805, 2632, 1688, 1766]. settling [1911]. Seven [95, 418]. Several [193, 1614, 1752, 2390]. Sewage [709]. Sex [1010, 464, 2367, 2451, 654, 743, 1227, 755, 2197, 1773, 2352, 1794, 1535, 2115, 1901]. Sex-related [2367]. Sex-Specific [1227, 2451, 2197]. Sexual [396, 1355, 750, 208, 819, 1333, 2496, 643, 64, 1046, 446, 1770, 2494, 2092, 2680, 1740, 2626, 2270, 1901]. Shad [761, 772, 601, 478, 178, 1364, 2014, 1641, 2770, 1802, 1738, 1872, 2326, 2097]. Shaded [257]. Shadowing [810]. Shallow [58, 1279, 1332, 295, 110, 1216, 1386, 714, 1035, 1781, 2519, 2616, 1897, 2408, 2529]. Shallow- [1035, 2616]. Shannon [1981]. shanny [2401]. Shape [986, 315, 764, 945]. sharing [2142]. Shark [732, 2117, 1856, 1962]. sharks [2355]. sharp [1968]. sharp-beaked [1968]. Shasta [1601, 778]. shedding [2795, 1972]. Shelf [1701, 2262, 645, 2905, 1078, 2144, 666, 2046, 1907, 2299, 2906, 2131, 1061, 598, 756, 320, 1932, 2688, 2237, 799, 1579, 1526, 1586, 2058]. shelfwide [2262]. Shelikof [606]. Shell [700, 906, 189, 1097, 428, 2216, 2824, 1818]. Shellfish [1449]. Shells [1197]. Shelter [2142]. Sheridan [1880]. Shield [1726, 2675, 1650, 1731, 2487, 976, 1605, 2312, 1647, 479, 1404, 682, 975, 2405, 2246, 1359]. Shift [1352, 1123, 1332, 784, 2401, 2615, 2666, 2904]. shifts [1874, 2257, 2907, 2695]. Shiner [662, 459]. Ships [1095]. Shoal [1277, 1743, 432]. shock [1994, 1593]. shooting [2011]. Shore [785]. Short [835, 1382, 387, 251, 1156, 1955, 298, 2092, 2243, 2571, 2070, 534]. Short-Duration [835, 1382]. short-finned [2092, 2070]. Short-Term [387, 251, 1156, 298, 1955, 2243, 534]. Shortcomings [203]. Shortening [515]. shortfin [1843]. Should [2806, 2181]. Shown [1191]. Shredders [1158]. Shrimp [1357, 1238, 665, 500, 349, 933, 196, 636, 231, 612, 1176, 1306, 171, 92, 574, 1715, 1580, 2464, 2479]. shunt [2870]. Sialis [2768]. Sicydium [1627]. side [2637]. side-aspect [2637]. sided [325]. sidescan [2252, 2529]. Sierra [2566]. Sieve [276, 917]. Sieves [945]. Sigmoid [383, 858]. Sigmoidal [1370]. Signals [2739, 2737, 2142]. signature [2525]. Signatures [1335, 1654, 2879, 2878, 1964, 2670, 2226, 2880]. Significance [5, 573, 920, 299, 1342, 864, 489, 1401, 2736, 2803, 2772]. significant [2020]. Siliceous [584, 211]. Silicon [897]. siliquoidea [1486]. Sillaginodes [1815]. silver [2688, 2476, 2367, 2633, 1100]. silverside [2268, 2503]. Similar [956, 2427]. Similarities [236]. Similarity [92, 1563]. Simple [1224, 6, 2624, 1411, 2830, 1234, 2837, 2602, 1909, 1971]. simplex [862, 1673]. simplified [2433]. simpsoniana [2082]. Simulated [358, 2902, 2611, 2444, 2540]. Simulating [1894, 2801, 2568]. Simulation [1315, 1250, 2420, 827, 725, 2753, 2777, 2613, 2554, 2341]. Simulations [1262, 1208, 2764, 2870]. Simuliidae [664, 1036]. Simultaneously [629, 2379]. since [585, 2175, 2012]. SINEs [2571]. Single [1040, 369, 2544, 1981, 325, 586]. single- [2544]. Single-Locus [1040, 586]. Single-sided [325]. Sinking [1167]. sire [2891]. Site [1374, 6, 1172, 1642, 2177, 2518, 2800, 1587]. site-specific [1587]. Sites [635, 679, 1254, 975, 2155, 1645, 2430, 2024, 2378]. Situ [63, 1174, 2168, 237, 1788, 2098, 2295, 2284, 904, 2024, 2263, 1858, 2285, 967, 1104]. Situk [860]. Six [669, 942, 482, 2033, 1744, 1484]. Size [508, 1811, 1235, 1556, 285, 553, 905, 976, 888, 1061, 380, 1252, 360, 670, 1152, 1164, 1098, 1341, 322, 685, 101, 473, 1213, 208, 840, 839, 1438, 1224, 547, 522, 791, 197, 1439, 1294, 389, 243, 896, 1372, 491, 600, 548, 393, 2116, 2782, 716, 715, 338, 1172, 661, 614, 1237, 1447, 110, 1216, 1047, 30, 1149, 765, 339, 489, 869, 1105, 359, 78, 1109, 139, 444, 878, 956, 936, 551, 767, 1689, 337, 92, 1451, 1126, 289, 1422, 150, 422, 804, 1063, 218, 1011, 985, 886, 1366, 784, 201, 2859, 2570, 2732, 1770, 1826, 1231, 2373, 1816]. size [2204, 1886, 1611, 1741, 2401, 2262, 2365, 2046, 2461, 2630, 1234, 2356, 1695, 2080, 2656, 2500, 2632, 1881, 1499, 2580, 2849, 1898, 2662, 2676, 2170, 1493, 1734, 1880, 1583, 2554, 2553, 2781, 1879, 1624, 2881, 2838, 2813, 2424, 2866, 1908, 2811, 2441, 1589, 2778, 1940, 2891, 2261, 2774, 2348, 2556, 2503, 2846, 1878, 2512, 2205, 2357, 2089, 2783, 2843, 2339, 2042, 2458, 1952]. Size- [2116]. Size-at-Age [150, 2373]. Size-Based [1422, 956]. size-classes [2080]. Size-Dependent [522, 791, 869, 1105, 878, 422, 2782, 1816, 2365, 2656, 2500, 2866, 2811, 2774, 2503]. Size-Related [784, 1689, 1366]. Size-Selective [670, 716, 715, 30, 1741, 2553]. Size-specific [1556]. Size-Structured [1224, 139, 2778]. size/age [1234]. Sized [128]. Sizes [429, 1986, 2900, 1942, 1626]. Skeletal [590]. Skeletonema [2833]. skill [2857]. skin [2192]. Skipjack [402, 1426, 1668, 2748]. Skipper [2857]. Skistodiaptomus [2521, 2522]. Slave [211]. Slope [660, 2485, 1907, 2906, 2091]. Slow [2697, 2404]. slow-growing [2404]. sluicing [2109]. Small [316, 1435, 174, 133, 647, 288, 437, 364, 981, 567, 518, 1320, 532, 1385, 550, 608, 1275, 804, 235, 1797, 2485, 2290, 2080, 2523, 1763, 1643, 2310, 2396, 2563, 2805, 2405, 1699, 2640, 2379, 1923, 2562]. Small-Bodied [518]. Small-scale [1385, 1643]. small-subunit [1763]. Smallmouth [383, 2291]. smelt [1531, 1823, 2523, 1572, 2558, 2651, 1828]. Smolt [888, 991, 94, 296, 389, 243, 932, 1047, 1380, 1067, 523, 2412, 2820, 2463, 2667]. Smolt-to-Adult [389, 932]. Smolte [733]. Smoltification [1244, 691, 1382, 169, 430, 759, 1363, 1381, 1770, 2137]. Smolting [132, 819, 1787, 2864]. Smolts [265, 472, 1014, 1565, 2486, 1504, 2192, 2199, 2054, 1838, 1696, 2749, 1955, 1890]. smooth [1817]. Snail [711, 338, 1636]. Snail-Dominated [711]. Snails [493]. Snake [2075, 2911, 22, 1619, 2758, 2463, 2535, 495, 2454]. Snapper [660, 2781]. Snappers [78]. snook [1884]. Snow [1272, 841, 750, 540, 1030, 249, 714, 1103, 31, 1136, 2432, 2413, 2763, 414, 2127, 1571, 2126, 2102, 2498]. Snowmelt [785, 1726]. snowpack [2053]. social [2034]. Socioeconomic [462]. Sockeye [1348, 1315, 15, 1058, 582, 265, 389, 695, 354, 932, 204, 370, 1259, 846, 1208, 1463, 1462, 602, 97, 341, 250, 1014, 1461, 1925, 2689, 2132, 2486, 2516, 2204, 1794, 2267, 2539, 2140, 2862, 1868, 2085, 2526, 1767, 2567, 1857, 2074, 1682, 1722, 1481, 2528, 1583, 2554, 2334, 2624, 2638, 2759, 2829, 2223, 2568, 2178, 2398, 2167, 2585, 2482]. Sodium [834]. Soft [189, 428, 1017, 704, 56, 2722, 1550, 2711, 1761, 1912]. Soft-Bodied [56]. Soft-Shell [189, 428]. Softwater [753, 1595]. soil [1601, 1754]. Soils [102, 637, 1753, 1754]. solandri [1684]. Solar [2147, 2760, 2640]. Sole [1287, 481, 2131]. Solea [1287]. Soluble [486, 195, 286, 2865]. Solute [31]. solutes [2342]. solution [1617]. solution-based [1617]. Solutions [490, 1591]. Somatic [33, 494, 2140, 2431, 1952]. Some [2254, 55, 501, 500, 479, 1661, 974, 1206, 1580, 926, 2389, 2125]. Sonar [2517, 2856, 2660]. sonars [2252, 2529]. soufia [2623]. Sound [2021, 2274, 2700, 2022, 2025, 2726, 2767, 2018, 1014, 2024, 2026]. Sounder [1065]. Sounds [703, 720, 721]. Source [85, 424, 36, 70, 2005, 2905, 2442]. Sources [1962, 1215, 1077, 1268, 644, 1960, 2710, 2768, 2605, 2343, 2214]. south [1780, 2445, 1616, 2601, 2277, 2587, 538, 765, 2749, 767, 1184]. south-central [1780, 2445, 1616, 538]. Southeast [1457, 2013, 899, 1277, 774, 432, 193, 2815]. Southeastern [14, 1500, 2312, 2214, 2537]. Southern [1010, 32, 436, 122, 1773, 1918, 1984, 1885, 2616, 1785, 1683, 1572, 2629, 2111, 1980, 1790, 2642, 172, 610, 1225, 173, 1334, 1055, 626, 837, 384, 1274, 146, 244, 1005, 946, 507, 1088]. Southwest [969, 2902, 2368, 210, 268]. Southwestern [137, 1122, 2576, 2081, 414, 1579, 2127, 2126, 565, 671, 608, 1337]. sp [1472, 1632, 55, 1671, 1827, 1144, 1243, 2200, 1354, 1470]. sp.nov [1392]. Space [686, 2663, 1904]. spacer [2239]. Spacing [649]. Spain [1187, 1047]. span [1828]. Spanish [297, 369]. Sparus [1739]. Spatial [129, 1183, 783, 735, 1748, 806, 2597, 1225, 680, 2799, 2746, 901, 522, 2691, 364, 1232, 1211, 1347, 1809, 1882, 910, 1142, 1815, 776, 1280, 2849, 984, 1446, 1934, 532, 928, 2251, 1385, 1690, 2229, 2825, 2247, 646, 305, 1149, 1312, 8, 962, 1616, 2291, 734, 2004, 2056, 92, 722, 2018, 996, 1853, 1049, 1573, 1841, 2780, 2413, 2242, 2522, 1805, 2384, 1625, 2314, 1587, 2798, 2895, 2205, 1883, 2717, 2077, 2076]. Spatially [427, 574, 475, 1630]. Spatiotemporal [156, 2648, 2558]. Spawn [2841, 2676]. spawner [2669, 2561, 2815]. spawners [2524, 2414, 1989]. Spawning [598, 1577, 1477, 1145, 390, 2267, 2764, 933, 266, 695, 1535, 1134, 1133, 2669, 789, 138, 1351, 866, 552, 2099, 56, 643, 998, 815, 966, 121, 805, 693, 1084, 1925, 2394, 1797, 2155, 2159, 1986, 2412, 2861, 1645, 2318, 2873, 2539, 2461, 2088, 2295, 2747, 2744, 2812, 2286, 2613, 2676, 1893, 2696, 1723, 2350, 2646, 2311, 2845, 1862, 2617, 2211, 1784, 2059]. Special [416, 100, 2588, 2098, 2661, 1682]. spéciale [410]. specialization [2002]. specialized [1680]. Speciation [1359, 1866, 2322]. Species [1325, 156, 317, 735, 149, 126, 909, 1061, 669, 322, 768, 101, 1466, 408, 1248, 1032, 1416, 265, 1468, 791, 1321, 1262, 260, 311, 583, 238, 468, 303, 1446, 874, 736, 95, 305, 1191, 869, 603, 11, 936, 1430, 28, 972, 644, 2089, 376, 1311, 369, 418, 419, 1811, 1573, 2544, 2675, 1750, 2146, 1781, 2874, 2637, 2410, 1656, 2262, 2831, 1553, 1677, 2664, 2116, 2703, 2577, 1621, 2702, 2002, 1534, 1669, 1649, 2239, 1869, 2241, 2574, 2018, 1910]. species-dependent [2116]. Specific [697, 435, 813, 1227, 1556, 2412, 2451, 2740, 2442, 2770, 1583, 2325, 1587, 1937, 2197, 1278, 1972]. Specificity [1217]. Spectra [508, 739, 1152, 1886, 2262, 2261]. Spectral [1055, 1219, 2690, 2243]. Spectrometry [1012, 2787, 2298, 2854]. Spectrophotometric [826]. Spectroscopy [1062]. Spectrum [289, 1011, 1693]. Speed [1140, 632]. speeds [2567, 2568]. Sperm [2572, 2891, 805, 2498, 579, 1569, 1994, 1557]. Sphaerospora [1657]. Sphyrna [1856]. spicatum [1340]. Spicules [897]. spiked [2193]. Spill [1245, 495, 1601, 2021, 2023, 2026, 2030, 2274, 2700, 2022, 2025]. spined [1556]. Spiny [525, 2122, 2231]. spirocauda [2166]. Spirorchidae [1666]. Spirurid [630]. Spleens [930]. splines [1693]. splitting [2169]. Sponge [897, 1771]. Sport [462, 809, 2751]. Spot [324, 1576]. Spottail [459]. Spotted [336, 987, 2813]. Spotter [848]. spp [754, 1445, 1826, 55, 1655, 1991, 1672, 2307, 1224, 2905, 1764, 13, 2645, 930, 2374, 1534, 720, 721, 1366, 555, 1530]. sprat [1487]. Spread [616, 1899]. Spring [1244, 159, 443, 1723, 538, 1770, 1862, 2123, 1530, 2695, 2289]. spring-spawning [1862]. Sproat [13]. Sptattus [1487]. Squalus [525]. Square [804]. Squares [1205]. Squawfish [375, 1164, 1165, 852, 1281, 564, 2075, 1638]. Squid [208, 1123, 1243, 2092, 2070]. Squids [1423]. SSCP [2490]. St [2432, 156, 556, 1874, 1019, 1670, 2844, 814, 172, 610, 1918, 245, 442, 1225, 90, 2839, 1181, 851, 2428, 2616, 321, 2327, 414, 2069, 747, 1968, 2127, 1760, 1387, 231, 1176, 2100, 2099, 1771, 66, 517, 594, 356, 2126, 666, 2035, 171, 92, 2772, 850, 1005, 946, 1790, 2642, 1486, 2901, 1270]. St. [1365, 1686, 1384, 1180, 1212, 2281, 2797, 214, 1196, 787, 939, 1211, 2376, 2621, 2384, 2884, 2104, 2840, 1801, 2158, 2558, 1507, 1927, 966, 2908]. Stability [343, 768, 456, 364, 402, 1185, 1154, 471, 2014, 2350, 2056, 2358, 2682, 1847]. Stabilization [1045]. Stable [1993, 1295, 2067, 539, 1836, 2647, 1169, 1928, 261, 2875, 2906, 503, 2565, 2499, 1797, 2874, 1978, 2873, 1545, 1979, 2872, 2877, 2691, 1596, 1768, 2577, 1806, 2479, 2203, 2343]. stable-isotope [2873, 1596, 1806]. Stackhouse [610]. Stag [23]. Stage [486, 598, 1098, 745, 138, 609, 378, 1303, 1602, 1673, 2300, 2895]. stage-structured [2300]. Stages [316, 179, 1263, 1083, 339, 1307, 182, 1970, 1990, 2021, 2413, 1695, 2041, 1957, 2841, 1571, 2436]. stagnalis [1636]. stamp [2542]. standing [2106]. Standings [1157]. Starvation [84]. State [710, 452, 1870, 2497, 2663, 2331, 527]. States [2727, 2673, 1523, 1724]. Station [162, 2573, 2908]. stationary [1514]. stations [1895, 1896]. Statistical [929, 160, 1269, 203, 715, 2431, 1385, 1603, 1, 770, 717, 620, 2006, 103, 2096, 649, 982, 1552, 2888, 2004]. Statistically [504]. statistics [2399].

Statolith [596, 1123, 1423, 2382]. Status [132, 929, 501, 1439, 509, 941, 166, 542, 681, 1299, 458, 1013, 728, 459, 1204, 1460, 1509, 2713, 2873, 2256, 1839, 1985, 2396, 1564, 1929, 2357, 1798]. Steelhead [1125, 405, 999, 307, 919, 521, 221, 98, 1307, 1751, 1987, 1958, 2576, 2080, 1933, 2600, 2119, 1490, 2325]. Steller [2222]. Stenodus [2847]. stenolepis [2684, 1312, 1198]. stenotherms [2218]. Step [1186, 1237, 1887, 1881, 1880, 1879, 1878]. stepwise [2228]. Stereo [1273]. Stereo-video [1273]. Stereocinematographic [632]. Steringophorus [1664]. Steroid [438, 743, 270, 1535, 1610]. Steroids [654, 811]. stewardship [1953]. Stickleback [139, 1607]. Sticklebacks [473]. Still [1316, 1190, 2037]. Still-Water [1190]. Stimulates [779, 1650]. Stimulation [169]. Stizostedion [2570, 487, 126, 670, 1538, 331, 557, 1261, 1069, 1372, 2188, 2710, 1009, 2030, 451, 2496, 2596, 1373, 2870, 1039, 1364, 292]. Stochastic [21]. Stock [1819, 958, 737, 892, 986, 1358, 315, 685, 15, 1173, 297, 361, 69, 933, 868, 963, 1537, 910, 1471, 745, 595, 774, 1429, 928, 248, 530, 1799, 1367, 838, 1214, 2670, 1059, 869, 1078, 178, 594, 1433, 1077, 1684, 1087, 475, 1135, 57, 111, 234, 441, 1459, 1460, 1205, 2648, 2689, 2462, 2898, 2830, 2423, 2232, 1641, 2397, 2601, 2181, 1580, 1999, 1882, 2234, 2916, 2740, 1822, 1547, 1722, 2327, 2724, 2524, 2107, 2914, 2762, 2663, 1502, 1914, 2845, 2806, 2617, 1506, 2383, 2438, 1884, 1891, 2634, 1626, 1575, 2372, 1600, 1711]. stock-production [1999]. stock-recruit [2181]. Stock-recruitment [1537, 1722, 2724, 2524, 1502, 2634, 1711]. stock-specific [2740]. Stocked [1127, 1039, 1379, 1720, 1403]. Stockholm [410]. Stocking [254, 253]. Stocks [1106, 147, 673, 482, 68, 765, 1259, 109, 594, 551, 1088, 1459, 2014, 1773, 2844, 1590, 2106, 2179, 1614, 2583, 2476, 2182, 1999, 2074, 1681, 1547, 2235, 2288, 2310, 1583, 2638, 2329, 2759, 2805, 2336, 2606]. Stoichiometry [1437, 1931]. Stomach [676, 2150, 2170, 2069, 2096]. Stomachs [1038]. Stoneflies [861, 400]. Storage [67, 746, 66, 2903, 2121]. Stores [843, 326, 1642]. Storms [162]. Strain [81, 307, 2661]. Strains [1263, 1127, 186, 254, 253, 891, 1239, 406]. Strait [665, 260, 27, 2283, 636, 606, 1302, 537, 1445, 2708, 1019, 123, 159, 155, 2464, 8, 2751, 1029, 2189, 2289]. Strategies [342, 672, 116, 607, 1175, 1059, 863, 725, 1758, 2406, 2839, 2415, 2451, 2314, 2316, 2217, 2733, 2000, 1795, 2885, 2201]. Strategy [106, 1375, 1431, 267, 2386, 1830, 2792, 2899]. Stratification [925, 1438, 1294, 1202, 2277, 2587, 2739, 1911]. Stratified [397, 309, 789, 147, 830, 1000, 2389, 2254, 1594, 2755, 1897, 2414]. stratified-Petersen [2414]. Straying [774, 1222, 996, 2715]. Strays [996]. Stream [1435, 803, 375, 685, 768, 2618, 1209, 256, 133, 647, 999, 2434, 1424, 855, 251, 674, 1350, 449, 437, 364, 257, 825, 332, 384, 286, 158, 727, 301, 827, 1320, 874, 532, 230, 67, 300, 1893, 519, 659, 400, 503, 1306, 760, 80, 359, 1168, 367, 1265, 1327, 31, 235, 985, 563, 39, 201, 497, 1573, 2138, 1525, 1607, 2566, 2579, 1562, 2706, 1978, 2190, 1501, 2620, 1545, 1979, 2088, 1875, 2080, 2575, 2893, 2272, 2176, 2435, 1885, 2221, 1944, 2511, 1490, 2645, 1050, 1481, 2604, 2602, 2563, 2581, 2090, 2348, 2758, 1608, 1699, 2053, 2574, 1789]. stream [1717, 1745, 2343, 2537, 2815, 1623, 2562, 2911]. Stream-bed [1893]. stream-borne [2893]. Stream-Drift [985]. Stream-Dwelling [256, 400, 359, 2348]. stream-fish [1573]. stream-living [2190]. stream-subsurface [2435]. stream-type [2758, 2911]. stream-water [2053]. streambed [2799]. Streamflow [503, 1327]. Streamflows [98]. Streams [638, 104, 9, 565, 634, 635, 807, 127, 496, 1152, 920, 1036, 173, 1324, 197, 711, 206, 837, 228, 793, 384, 447, 705, 1124, 1320, 261, 659, 658, 782, 608, 193, 967, 151, 913, 940, 1797, 2576, 1638, 1700, 1920, 2747, 2435, 1842, 2672, 1493, 1734, 2563, 2527, 1687, 2473, 2369, 1693, 2491, 1651, 1484, 1483, 1789, 2640, 2214, 1923, 2562]. Strength [1174, 416, 483, 1492, 2637, 2701, 2321, 2824, 2158]. strengths [1788, 1746]. Stress [1199, 1384, 387, 855, 912, 306, 1056, 887, 1143, 950, 471, 817, 1096, 1314, 2735, 2054, 1860, 2547, 1959, 2592, 2750, 1739]. stress-related [2750]. Stressed [366, 413]. Stresses [959, 746]. stressor [2736]. stressor-response [2736]. stressors [2817, 1593]. Striped [1012, 1339, 84, 178, 729, 767, 873, 1096, 943, 1314, 1313, 2687, 2328, 1640, 2656, 2500, 1957, 2440, 1853, 1575, 2372]. strobus [2707]. Strong [653, 2060]. Strongylocentrotus [2512, 966]. Strontium [1012, 2504, 1219, 657, 1591, 2854]. structural [2681]. structurally [2317]. Structure [921, 156, 1174, 735, 1180, 905, 1201, 730, 90, 12, 762, 1468, 1339, 855, 492, 880, 1155, 1197, 1054, 1469, 51, 1446, 833, 1203, 1118, 838, 646, 1078, 227, 1109, 965, 140, 956, 936, 1402, 154, 1422, 506, 850, 541, 1305, 218, 1011, 858, 1166, 579, 39, 1459, 2309, 1811, 1950, 2057, 2029, 2269, 2847, 2092, 1743, 1998, 2046, 2332, 1714, 2010, 2183, 1843, 1596, 1961, 2643, 1837, 2876, 2133, 2849, 2821, 2403, 2083, 2424, 2070, 2734, 2331, 2390, 1808, 2772, 1825, 1884, 1883, 2843, 2794, 2682, 1737]. Structured [1224, 465, 263, 1251, 139, 2830, 1611, 2601, 2300, 2778, 1559, 1894]. Structures [392, 742, 2220]. Structuring [1028, 2415, 2063, 2086, 2347]. Studies [801, 456, 1022, 43, 758, 775, 269, 2822, 2590, 2021, 2048, 2259, 2533, 2464, 2180, 2070, 2907, 1719, 2712, 2273, 2436, 2438]. Study [156, 639, 1140, 665, 390, 672, 702, 354, 307, 51, 1385, 869, 1105, 965, 574, 441, 369, 181, 2753, 2238, 2877, 2049, 2639, 1961, 2444, 1675, 1791, 2866, 2099, 2120, 1774, 1622, 2200, 2255, 2904]. studying [2150, 2655, 2654]. Stunted [741]. Stunting [624]. Sturgeon [611, 787, 939, 1131, 100, 1830, 2377, 2797, 1900, 719, 2832, 2244, 1532, 2158, 2224, 1288, 2854]. Styles [583]. Sub [8, 1206]. Sub-Arctic [1206]. Sub-Ice [8]. subalpine [1651]. Subarctic [223, 709, 1449, 1295, 228, 793, 332, 882, 2475, 2573, 2796, 2695]. subbifurcata [2401]. subcellular [2731]. Subfossil [1299, 2497]. Subgenus [1413, 1765]. subgroups [2804]. Subice [904]. Subject [30]. Subjected [558, 2156]. Sublethal [885, 2023, 1361, 2043, 1342, 2047, 1113, 512, 724, 511, 2722, 2444, 2156, 2540, 1771]. Submerged [456, 37, 844, 2836]. Submergent [54, 93, 2507]. submersed [2330]. Subpopulation [166]. Subpopulations [1463, 1758, 2265, 1949]. Subsequent [861, 1097, 846, 150, 1957, 1955]. subspecies [2623]. substances [2850]. substrata [2347]. Substrate [285, 768, 552, 1246, 1264, 209, 497, 2485, 1793]. Substrates [91, 913, 2631, 2711, 1810]. Subsurface [67, 2435]. Subtropical [238, 1652, 2552, 2408]. subunit [1763, 1902]. Success [210, 194, 805, 1278, 2034, 1814, 2267, 2143, 2296, 2026, 2490, 2329, 2891, 2556]. successful [2034]. Succession [993, 211, 2765]. successional [2125]. Successive [1435]. Sucker [4, 5, 1371, 1196, 679, 438, 376, 446, 1642, 2494, 2098, 2295, 2188, 2133, 1889, 1610]. suckers [2884]. sud [414]. sud-ouest [414]. Sudbury [1034, 48, 468, 1179, 2396, 1798, 1917]. suecica [2833]. Sufficient [703]. Suggests [369]. Suitability [635, 727]. Suitable [659]. Sulfadimethoxine [688]. Sulfate [170, 1527]. sulfate-saline [1527]. sulfide [1595]. Sulfohydrolase [170]. Sulfoxide [826]. Sulfoxide/Acetone [826]. Sulfur [539, 1093, 1080, 2818]. Sulfuric [157]. Sulphate [1647, 820, 1697]. Sumari [2086]. Summer [1250, 2199, 1438, 1181, 1294, 531, 830, 659, 1221, 244, 650, 775, 1923, 2061, 2230, 2283, 2791, 1595, 2292, 2540, 1995]. summer-acclimated [2791]. summers [2018]. Summertime [432]. Sunfish [1366, 1816, 1814, 2773]. superimposed [2449]. superimposition [2620]. Superior [346, 2337, 382, 2104, 2515, 438, 19, 780, 28]. supersaturated [2156]. Supplementation [1480]. supplements [2588]. Supply [761, 38, 422, 971, 2703]. Support [634, 1556]. supported [1851]. Supporting [220]. supports [2490, 2804]. Surface [223, 814, 35, 889, 1294, 490, 469, 512, 1112, 76, 1500, 2698, 1761, 2520, 1633, 1893, 2752, 2755, 2070, 2036, 2287]. surfaces [2342]. Surficial [305, 962]. surgically [2465]. Surplus [1346, 1147]. Surveillance [690]. Survey [745, 720, 1312, 1223, 965, 721, 103, 722, 1256, 1461, 2548, 1803, 2746, 1786, 1934, 2108, 2652, 2513, 2139, 1506, 2555, 929]. surveying [2108]. Surveys [810, 958, 680, 789, 147, 224, 426, 2161, 1800]. Survival [2865, 1263, 591, 214, 48, 865, 449, 834, 106, 772, 389, 116, 243, 180, 1453, 1527, 932, 827, 778, 26, 968, 612, 339, 878, 1039, 1227, 308, 1756, 856, 1239, 1303, 1379, 1565, 2737, 2394, 1509, 1609, 1816, 2597, 2630, 2356, 2186, 2600, 1767, 2296, 2816, 1867, 1490, 2235, 1864, 1692, 1493, 1734, 1946, 1838, 2240, 1957, 2065, 1893, 2841, 2400, 2638, 1940, 2805, 2897, 2758, 2463, 2535, 1955, 1688, 2212, 2771, 2341, 2640, 1784, 1941, 2911]. Survivorship [3, 2543]. Susceptibility [417, 2811]. Suspended [1167, 102, 1071, 428, 930, 337, 724, 2741, 2164, 2665]. Suspension [1021, 1757, 2474]. Suspension-Feeding [1021, 1757]. Susquehanna [1193]. Sustainable [1256]. Sustained [1362]. Sweden [2294, 1195]. Swedish [2469, 942, 2800, 2491, 924, 577, 1383]. swept [2649]. Swim [2567, 1344, 2568, 181, 182, 1807, 2508]. swim-bladder [1807]. Swim-Up [181, 182]. Swimming [1140, 632, 835, 1382, 623, 505, 1063, 1051, 348, 1721, 1218, 943, 1314, 1313, 2465, 2095, 1817, 2545, 2705, 1933, 2038, 2037, 1273, 2134, 2224, 2456, 2593, 1718, 2032]. Switch [545]. switching [2761]. swordfish [1644]. Symbionts [345]. symmetricus [1638]. symmetry [1518]. Sympatric [347, 656, 582, 1327, 250, 1639, 1490, 1635, 2558, 1966, 1699]. Symposium [1085]. Synergistic [2599]. Synopsis [1138]. Synthesis [958, 1163, 2185, 2683, 2292, 2540, 2907]. Synthetic [877, 1207, 1153]. Sysposium [2589]. System [392, 697, 1194, 465, 1169, 546, 1273, 85, 304, 992, 28, 39, 1841, 1797, 1660, 2088, 1981, 2888, 2010, 2183, 1857, 1657, 1593, 1771, 1608, 1739]. System-of-Equations [465]. Systematic [1000]. Systematics [533, 1467]. système [1660]. Systemic [583]. Systems [104, 1252, 1186, 1469, 62, 2517, 1542, 1752, 2551].

T [1855, 560, 1336]. Tactics [1261, 359, 1430, 2267]. Tadpoles [24]. Taeniacanthidae [1677]. Tag [771, 405, 925, 1331, 772, 71, 175, 363, 1137, 1037, 1027, 2795, 2779, 2712, 1505, 1913, 1972, 2742]. tag-misread [2712]. Tag-Recovery [925, 1037]. tag-specific [1972]. Tagged [737, 2831, 2534]. Tagging [1426, 559, 1189, 2590, 2795, 2407, 2534, 2533, 2422, 2676, 2078, 2748, 1972]. tags [2019, 2584, 1683, 2161, 2535, 2786]. Tahoe [1033, 706, 1485, 2893]. tailings [2030]. taille [1234]. taille/âge [1234]. Tailrace [1281]. Taking [1086, 361, 1092]. tanks [2761, 1512]. Tanner [1035, 1289]. Tannins [997]. tapirina [1142]. Tar [1298, 1960]. Target [1174, 483, 2637, 2412, 1788, 2321]. Targets [1175, 314]. Tasmanian [1832]. Tautogolabrus [1678]. taxa [1659, 1763]. Taxon [2282, 956]. Taxon-based [956]. Taxonomic [1116, 703, 506, 2859, 2727, 2269, 1656, 1809, 1963]. TCDD [378, 1610]. TCMTB [1342]. tebufenozide [2452]. Technique [285, 772, 1320, 2788, 2787, 2631, 2563, 1506, 2562]. Techniques [131, 659, 965, 1460, 2149, 2037]. teeth [2670]. Telemetry [623, 2168]. Teleost [1360, 2286, 1876]. Teleostei [1271, 1286, 2623, 583, 1048]. Temperate [1089, 16, 238, 1057, 489, 1407, 1653, 2471, 2824, 2552, 2411, 2665, 2323, 2608, 1789, 1599]. Temperature [1182, 1128, 415, 396, 591, 416, 520, 1164, 1165, 148, 442, 161, 296, 96, 1369, 1292, 79, 798, 420, 2418, 2915, 20, 704, 2820, 247, 110, 1216, 2866, 339, 2863, 1013, 2223, 820, 505, 1177, 337, 150, 2323, 2574, 856, 564, 693, 190, 1132, 1565, 2136, 2732, 2912, 2579, 1748, 2077, 2159, 1840, 1540, 2204, 1794, 1922, 2365, 1617, 1618, 1988, 1787, 2526, 2916, 2543, 2791, 1863, 2662, 2913, 2754, 1821, 2683, 2914, 2065, 2781, 1956, 2431, 1548, 2752, 2100, 2829, 2405, 2076, 1716, 1995, 2197, 1718, 2409, 2341, 1486, 2582, 2059, 2128, 1941, 1847, 2586, 2608]. Temperature- [2866]. temperature-based [1748, 2077, 2076]. Temperature-Dependent [150, 564, 2323, 1794, 2916, 2608]. temperature-driven [1486]. temperature-modified [2781]. Temperature-related [2820]. Temperatures [35, 432, 372, 1648, 2061, 2266, 2817]. templet [2799]. Temporal [1183, 783, 2014, 1562, 806, 678, 1211, 1419, 566, 776, 286, 345, 532, 1172, 1693, 1335, 8, 734, 2060, 459, 355, 111, 1816, 2597, 2799, 2242, 2522, 2691, 1882, 1587, 2247, 1964, 2665, 1616, 2291, 2004, 2056, 2358, 2018, 2682, 1550]. temporal-spatial [2522]. Temporary [232, 1570]. tenuis [2616]. teratogenesis [1642, 2797]. Term [885, 387, 251, 674, 747, 1083, 1156, 511, 298, 544, 2486, 534, 1977, 1119, 1780, 1369, 2243, 2523, 1185, 2064, 1895, 1253, 2045, 2050, 2440, 1955, 1798, 2585, 2482, 1917]. terme [534]. Terminal [540]. terranovae [1257]. Terrestrial [2214]. territoriality [1850]. Territories [2111, 2112, 277, 123, 27, 8, 211]. Territory [197, 1499, 2475]. Test [1125, 237, 1370, 279, 480, 1056, 228, 793, 528, 716, 715, 2332, 2430, 2122, 2812, 1559, 2416, 2358]. Testing [1587, 190, 2521, 2802]. Tests [737, 160, 482, 1191, 1518]. tetrachlorobiphenyl [1532]. Tetrachlorodibenzo [474, 378, 1303, 1996]. Tetrachloroguaiacol [1361]. tetranucleotide [2016]. Tetrapturus [1339]. Tetraselmis [2833]. Teuthoidea [2070]. Texas [1469]. TFM [2598]. Their [573, 1349, 1030, 1454, 345, 1241, 1024, 1400, 1221, 760, 1415, 973, 507, 1135, 1081, 1301, 883, 2039, 317, 1566, 2392, 1666, 1810, 1933, 2038, 2756, 2442, 112, 2444, 1683, 443, 2222, 1586, 1791, 2355, 2755, 2035, 1832, 250, 2457]. theoretical [1722]. Theory [229, 1186, 845, 1137, 39, 2896, 1808, 2826, 2610]. Theragra [1983, 287, 606, 1519]. Thermal [83, 1188, 1202, 283, 140, 2482, 2787, 2277, 2587, 2739, 1772, 2002, 1530, 2585]. Thermally [397, 366, 2389, 2254]. thermocline [2218]. Thin [119, 1693]. thin-plate [1693]. thinning [2190]. thiocyanate [2001]. Third [486, 609, 1673]. Third-Stage [486, 609, 1673]. thorns [2814]. thornyhead [1845]. Thoughts [1329]. Thousand [929]. Threat [307]. Threaten [329]. Threatened [987]. Threats [456]. Three [126, 395, 677, 1407, 1259, 109, 461, 1473, 419, 1970, 2660, 2434, 1857, 2116, 303, 2804, 2778, 1616, 1570, 2574, 2582]. threespine [1607]. Threshold [226, 529]. Thresholds [1175]. Throughout [959]. Thunnus [317, 32, 2601, 1234, 436]. thuringiensis [1193, 1265, 913, 1307]. Thymallus [415, 685, 228, 792, 793, 1378, 2889, 282, 2754, 1689]. Thyroid [271, 1362, 2137, 624]. Thyroxine [169, 1363, 2385]. tiburo [1856]. Tidal [266, 2738, 1783, 1481, 2840, 1576]. Tight [515]. tilapia [2115, 299, 471]. Timber [127]. Time [888, 344, 925, 266, 295, 1351, 513, 514, 1502, 686, 643, 505, 594, 657, 28, 2614, 2632, 1722, 2297, 2645, 1904, 1909, 1559, 2651, 1506, 2899, 1626, 1828, 2059]. time-series [1722, 1559]. time-varying [2651]. Times [585, 1208, 1972, 2742]. Timing [991, 1824, 2615, 2295, 1844, 852, 2892, 264, 2155, 2159, 2087, 2573, 2287, 2131, 2289]. Tissue [683, 170, 1093, 61, 1052, 1243, 88, 466, 2430, 2639, 2770, 1496, 1926, 2419, 1812]. tissue-specific [2770]. Tissues [679, 1206, 1550, 1836, 1801]. Tjeukemeer [1216]. Toadfish [96, 420]. Tolerance [421, 302, 877, 1207, 1453, 1052, 808, 621, 1067, 242, 1504, 1542, 1772, 1528, 2273, 1716, 1955, 2097]. Tolerances [83]. Tolerant [834]. Tomcod [1075, 1741, 1741, 1075]. tomography [2510]. Tool [1070, 1236, 2896]. Tools [1038, 2542, 2879, 2226, 2880]. Top [803, 762, 2701]. Top-Down [803, 762, 2701]. topography [2312, 2702]. Total [1357, 1370, 678, 1294, 637, 52, 1031, 1752, 594, 1415, 978, 2570, 2351, 2123, 2165]. Toxaphene [1253]. Toxic [1318, 546, 1157, 808, 2735, 2194, 2342, 1903, 1561, 2173, 2446, 1610]. Toxicant [1199, 1974]. Toxicant-Induced [1199]. Toxicity [398, 2193, 712, 118, 24, 1034, 877, 800, 1296, 1017, 752, 186, 247, 1191, 185, 1113, 1200, 378, 298, 1018, 163, 1307, 419, 2867, 1713, 1567, 1566, 2194, 2803, 2048, 1747, 2001, 2008, 2802, 2801, 1774, 1912, 1917]. Toxicological [2227, 2125]. Toxin [604]. Toxin-Producing [604]. Trace [1365, 807, 1270, 450, 1017, 1340, 1742, 2538]. Traced [1093]. tracer [2862, 2324]. Tracers [54, 215, 2429, 2906, 2479]. Trachichthyidae [1048]. tracing [1978, 1545]. tracked [2095]. tracking [2037, 2060]. Trade [2468, 2186, 874, 1867]. Trade-off [2186, 1867]. Trade-offs [2468, 874]. Training [583, 2134]. traits [1649, 2210]. trajectories [2466]. TRANSAQUA [2488]. transboundary [2374]. transcribed [2239]. Transect [147, 1514, 1501]. Transfer [1045, 1024, 2220, 1890]. Transferable [628, 824]. Transformation [1067, 523, 940]. Transformations [1080]. transgenic [1613, 2593]. Transition [1180]. Translocation [1303]. transmission [2163, 2322, 1889, 1685]. transmitted [1674]. transmitters [2465]. transmontanus [611, 2854]. Transparent [119]. transplantation [1550, 1551, 2715]. Transplanted [297, 1187]. transponder [2535]. Transport [1435, 872, 287, 1193, 977, 1073, 827, 796, 1176, 627, 66, 940, 377, 1354, 1741, 1842, 1481, 1622, 1608]. Transported [14, 1122]. Trap [594, 363, 1136, 2397]. Trapping [1512]. Traps [368, 645, 143]. Trawl [1086, 751, 498, 742, 1198, 103, 2622, 2548, 2406, 2397, 2685, 1907, 2108, 2157, 2139, 1756, 2717]. Trawl-Caught [1198]. trawlers [2406]. Trawling [1262, 1808]. Trawls [1228, 669, 804]. Treated [674]. Treatment [343, 158, 1253, 44, 234, 2598, 2263]. Treatments [229]. trees [2879, 2226, 2880]. Treespine [139]. Trematoda [1666]. Trend [35]. Trends [979, 1258, 963, 1597, 1419, 584, 532, 458, 114, 1300, 475, 459, 2709, 1826, 1500, 2428, 2739, 2396, 2513, 1519]. Triakis [2117]. trials [2039]. Tributaries [487, 832, 2114]. Tributary [425, 19, 646, 1275, 1326]. Trichoptera [829, 600]. tridentata [931]. tridentate [425]. trifluoromethyl [2598]. Trinity [1178]. trip [1522, 2751]. Tripfoid [141]. Triploid [1452, 1454, 1453, 1455, 2270]. trituberculatus [831]. trivittatus [119]. Troll [129, 210, 663]. Trophic [701, 761, 1167, 634, 2392, 2716, 509, 539, 257, 476, 816, 1169, 321, 542, 681, 747, 1176, 1114, 1013, 760, 728, 1402, 290, 29, 1418, 1028, 2810, 1870, 1797, 2497, 2087, 2213, 1580, 1596, 2475, 1985, 2068, 2903, 1755, 1929, 2357, 2203, 1853, 299]. Trophodynamics [1337, 2872]. Tropical [2753, 596, 866, 1243, 1306, 940, 2569, 2778, 1503]. trossulus [1692]. Trough [1064]. Trouser [751]. Trout [488, 22, 667, 885, 1139, 639, 803, 179, 888, 991, 1355, 853, 690, 692, 740, 741, 1226, 635, 929, 1201, 1209, 842, 170, 405, 148, 1477, 118, 1058, 479, 431, 699, 995, 718, 515, 1324, 1452, 81, 1127, 423, 1296, 48, 533, 674, 445, 490, 437, 364, 1017, 476, 1056, 186, 180, 183, 818, 1255, 1361, 307, 919, 491, 977, 417, 384, 47, 688, 1282, 158, 393, 682, 254, 253, 1403, 225, 528, 875, 552, 513, 247, 786, 1190, 122, 338, 1052, 300, 1343, 1344, 26, 521, 184, 514, 881, 185, 1336, 870, 1200, 227].

Trout [512, 399, 139, 1100, 1484, 221, 1412, 98, 967, 255, 1473, 1063, 277, 376, 1239, 817, 511, 378, 1303, 348, 1151, 1217, 1218, 1307, 181, 182, 1301, 1084, 1409, 2380, 2722, 2867, 1569, 2138, 1970, 2118, 2029, 1751, 1987, 2394, 2366, 1498, 2066, 1967, 2155, 2373, 1758, 2168, 1638, 1577, 2448, 2225, 2516, 2677, 1903, 1540, 1646, 1645, 2318, 1561, 2061, 534, 1998, 1629, 2620, 2459, 2795, 1707, 1511, 2719, 1920, 2227, 2545, 2705, 1479, 1996, 1761, 1482, 2010, 2183, 2337, 2661, 2575, 2871, 2747, 1933, 2600, 2639, 1720, 2243, 2294, 1939, 2028, 2119, 2756, 2063, 1837, 2812, 1490]. trout [2381, 2849, 2001, 2134, 2444, 2130, 2796, 2467, 1705, 1704, 2801, 2515, 2633, 1687, 1902, 2325, 1564, 2473, 1635, 1791, 2491, 1532, 2592, 2228, 2453, 1855, 2292, 2540, 2699, 2060, 2348, 1980, 2405, 2606, 1989, 1699, 2353, 1982, 1718, 2743, 2341, 1891, 2783, 2270, 1865, 2062, 2489, 2032, 1706, 1703, 1923, 2195, 2480, 335]. Trout-stocked [1403]. Troutf [327]. Trouts [510]. trutta [1970, 2394, 2366, 2066, 888, 991, 1355, 692, 740, 741, 2373, 1758, 929, 2620, 995, 718, 1452, 2010, 2183, 1720, 2294, 1939, 2028, 491, 1490, 393, 2130, 2796, 786, 1635, 2491, 1989, 277, 1307, 1084, 2195]. tshawytscha [1465, 2465, 474, 937, 1565, 1819, 1244, 2737, 1258, 2248, 756, 822, 2352, 944, 464, 1619, 2714, 1988, 890, 889, 903, 2715, 912, 169, 809, 860, 1453, 1362, 2580, 2442, 1342, 778, 1838, 2156, 2547, 521, 1025, 2749, 950, 2758, 1275, 960, 1327, 2626, 1040, 1300, 242, 996, 1135, 2211, 2212, 2771, 1475, 1410, 2454, 2911, 1515]. tubenose [1829]. Tucker [2157]. Tuktoyaktuk [2111]. Tumor [487, 2828]. tumor-like [2828]. Tuna [317, 32, 402, 436, 1426, 2780, 2753, 2698, 1234, 1668, 2676, 2748]. Tundra [779, 2324]. Turbid [889, 2403, 2426]. Turbidity [1182, 890, 1279, 953, 2790, 1948, 2409, 2783]. Turbine [772, 1838, 2006]. Turbot [1141, 2459]. turbulence [2401, 1643, 2662]. Turnover [2578, 2675, 2639, 2453, 1909, 1865]. turtle [1666]. Turton [1277]. twine [1888]. Two [1325, 1365, 1435, 342, 909, 1263, 994, 1437, 920, 322, 408, 1116, 560, 480, 583, 125, 186, 1185, 1677, 1362, 745, 33, 236, 894, 802, 1124, 1108, 1420, 1253, 820, 1430, 815, 242, 644, 376, 1239, 931, 2129, 2146, 2576, 2874, 2761, 2590, 2216, 1856, 2660, 1920, 2067, 2765, 2080, 2435, 1518, 1887, 2523, 384, 2580, 1542, 2821, 2303, 2796, 1496, 1895, 1896, 1931, 2875, 2329, 2558, 2686, 1989, 2438, 1575, 1503, 2319, 2275]. Two-Species [1430]. Two-Stage [745]. two-step [1887]. Type [1074, 1058, 1054, 731, 924, 860, 1902, 2696, 2758, 2556, 2911]. type-2 [1902]. Types [1293]. typical [1704].

U.S. [2780, 405, 2476, 1907, 1433, 2559, 2419]. U.S.A [1870]. U.S.A. [2576, 127, 2527, 2340, 2338, 2686, 2749, 1825, 2249, 2214, 2537]. Ukraine [1310]. Ultra [82, 403]. Ultra-Oligotrophic [82, 403]. Ultraphytoplankton [144]. Ultrasonic [623]. Ultrastructural [167, 1593, 1675]. Ultrastructure [1139, 430, 955, 419, 1592]. Ultraviolet [2256, 2146, 2760, 2148, 2147, 1769, 2789]. ultraviolet-B [2148]. Ulvaria [2401]. Umpqua [2749]. uncaught [1953]. uncertainties [2718, 1906, 1820].

Uncertainty [963, 1147, 1433, 74, 2181, 1510, 2681, 350]. Uncoupled [700]. undatum [363]. undecimalis [1884]. Under-Ice [295]. Under-Sized [128]. underlie [2624]. underlying [2723]. Underwater [123, 2766]. Underyearling [1259, 1504]. Undisturbed [127]. Unequal [1136]. Unexpected [961]. Unexploited [1185]. uniboite [325]. Unidirectional [786]. Unified [3]. Unionid [1387, 1388, 2129, 1620, 1927, 1486]. Unique [988]. Unit [721]. United [2727, 2673, 1523, 1724]. unlimited [2061]. Unlogged [1168]. Unmanipulated [1437, 2701, 1931]. unoiled [2024]. Unregulated [371]. Unusual [627]. upland [1726]. upon [611, 1649, 2479, 2128]. Upper [197, 1530, 1907, 1927, 2749, 1825, 2248, 1297, 2888, 1275, 31]. upriver [2567, 2568]. upriver-migrating [2567, 2568]. Upstream [1378, 14, 1122, 1320, 2889, 2884, 2563, 2562]. Uptake [23, 836, 261, 1926, 1820, 506, 1982, 422, 644, 1146, 1409, 2867, 2366, 2479, 1906]. Upwelling [268, 765, 544, 1513, 2840]. Upwellings [162]. Urban [481]. Urchin [116, 966, 2890, 2512]. Urophycis [2616]. ursinus [559, 961]. Ursus [1762]. USA [747, 710, 585, 1316, 527, 703, 1033, 207, 880, 288, 984, 193]. usage [2134]. Use [1106, 149, 501, 1349, 906, 983, 1501, 995, 547, 672, 1173, 2219, 2777, 423, 2243, 1517, 1169, 1001, 91, 1273, 1190, 733, 658, 538, 1306, 134, 1100, 1171, 1101, 1412, 363, 902, 2414, 2481, 1327, 153, 2535, 103, 755, 657, 649, 1135, 1092, 1432, 2859, 2155, 2835, 1979, 2356, 2660, 2567, 2812, 1588, 2770, 2802, 2130, 1965, 1791, 2491, 2400, 1559, 2568, 70, 1832, 2809, 1623, 1923]. Used [1269, 1175, 446, 2462, 2488, 1888, 1887]. Useful [1191, 2542]. Usefulness [315, 601]. user [2590]. users [2852]. uses [1694]. Using [754, 1174, 1244, 737, 986, 1358, 1318, 1803, 2872, 404, 354, 789, 1768, 857, 934, 128, 874, 1429, 2650, 530, 1138, 2300, 478, 1191, 1223, 1053, 863, 992, 998, 1037, 1000, 935, 993, 1473, 475, 722, 972, 1088, 446, 1163, 2743, 1256, 1002, 1398, 1460, 1475, 1136, 2788, 486, 2766, 2648, 1514, 1819, 2631, 379, 2736, 2362, 2194, 2497, 2298, 2495, 2516, 160, 2216, 1918, 1814, 2673, 2276, 2459, 2332, 2267, 32, 1617, 1618, 672, 2814, 1981, 2284, 2520, 1786, 71, 1822, 2812, 2724, 2107, 2231, 2663, 2781, 1799, 2228, 2670, 2120, 314, 2681, 2679, 2445, 2907]. using [1894, 1858, 2285, 2311, 1808, 1806, 2479, 1506, 2408, 93, 2591, 2252, 2529, 2883, 2203, 1818, 1883, 2000, 2510, 2372]. Utah [10, 1226]. Utility [2297, 2903]. Utilization [36, 1275, 1236, 233, 1720]. UV [1650]. UV-B [1650].

Vagrancy [214]. vagrant [1823]. Valdez [2021, 2023, 2026, 2274, 2700, 2022, 2025]. Valid [504]. Validation [104, 494, 775, 2788, 1615]. validity [1587]. Valley [1746]. Vallisneria [2621]. Value [524, 1953, 1861, 2528, 1683]. values [2886]. Valve [916]. Vancouver [1055, 210, 1657, 774, 1337, 244, 525, 1475, 1014, 82, 403, 2235, 2299, 2902, 2368, 663]. var [1193, 1265, 913, 1307]. Varden [2118, 256, 899, 2199, 533, 1124, 2804]. Variability [797, 696, 2356, 1033, 243, 354, 1247, 1864, 796, 758, 30, 489, 2157, 1411, 1598, 92, 242, 355, 1068, 1048, 544, 1319, 2675, 1758, 1562, 2548, 2792, 2434, 2661, 2188, 2277, 2587, 2381, 2586, 2418, 2116, 1934, 1947, 1586, 1690, 2228, 1558, 2895, 2772, 2293, 2084, 1960, 2582, 1612, 2628]. Variable [795, 1240, 1043, 115, 1984, 2081]. Variables [701, 1072, 1274, 936, 164, 1963, 1554]. Variance [571, 678, 795, 294, 526, 2490, 2649, 2390]. Variation [129, 1465, 1286, 1019, 60, 1472, 126, 347, 991, 1322, 994, 1061, 281, 473, 944, 124, 405, 1248, 951, 161, 698, 527, 431, 435, 1127, 939, 251, 933, 1580, 2085, 1133, 766, 1142, 919, 191, 776, 1846, 345, 393, 528, 1446, 1241, 1015, 1131, 1172, 614, 19, 305, 1215, 686, 1149, 120, 11, 503, 1025, 878, 244, 461, 650, 1310, 475, 65, 1102, 1305, 1104, 1462, 693, 2059, 873, 1503, 1819, 2689, 1632, 2268, 2248, 1813, 2494, 2177, 2622, 1991, 2328, 2031, 1489, 1641, 1511, 2688, 2545, 2715, 2595, 2272, 1809, 1882, 2384, 1815, 2710, 2442, 2725]. variation [2824, 2573, 2693, 2244, 2530, 1799, 1693, 2247, 2424, 1497, 2016, 1962, 2245, 1589, 1616, 2004, 1848, 2056, 2503, 2205, 2642, 1789, 1539, 1853, 2007]. Variations [415, 416, 94, 192, 450, 1211, 984, 1622, 622, 525, 1670, 2877, 2229, 2200, 2033, 2565, 1966]. Various [667, 1590]. Varved [1393]. Varying [1213, 492, 880, 1155, 98, 1379, 2484, 2256, 1985, 2845, 2651]. vascular [1817]. Vasotocin [306, 1056]. vaterite [1988]. Vector [1163, 1655, 2649]. vector-based [2649]. vegetal [2592]. vegetation [2376, 1971]. velata [2768]. Veliger [1246]. veligers [1572]. Velocities [1314]. Velocity [1125, 1323, 505, 985, 1313, 2810, 1933, 2577]. Vendace [969, 2081, 1756]. Venezuelan [2698]. Ventilatory [511]. Verification [435]. Vermont [1379]. Versions [883]. versus [1491, 634, 38, 2579, 391, 1582, 2397, 2125, 2382, 2552, 1896, 1897, 1507, 484, 1158]. vertebrae [1962, 2355]. Vertebral [435]. vertebrate [2268]. vertebrates [2725]. Vertical [1174, 1121, 360, 696, 2313, 648, 1073, 199, 204, 284, 1176, 221, 154, 41, 2308, 1701, 1531, 2690, 2521, 2522, 1487, 1549, 1534, 2100, 2901]. Vertically [810, 53]. Vertically-Migrating [53]. Vessel [663, 2000]. Vessels [210, 2393, 2652]. vetulus [481]. VHS [417]. VI [2145]. Viability [1452, 1509, 1619, 2314, 2627]. Vibrio [891]. Vicinity [1131, 2048]. video [2766, 2037, 1273]. Viewing [1448]. villosus [1686, 2159, 1145, 1922, 2365, 1038, 1932, 2451, 2170, 2840, 838, 536, 517, 1227, 1291]. Viral [417, 2619]. Virgin [805, 2394]. Virginia [757, 1957]. virginica [407, 1469]. virilis [2072]. Virtual [795, 1256, 2681]. Virus [288, 82, 2619]. Virus-Like [82]. Visibility [119]. visiting [2761]. Visual [789, 948, 1366, 2339, 1935]. Visualizing [2255]. Visually [91]. Vital [452, 869]. Vitamin [1642, 1532]. Vitelline [1333]. Vitellogenin [654, 2598]. vitreum [331, 557, 451, 292, 2570, 487, 1538, 331, 557, 1261, 1069, 1372, 2188, 2710, 1009, 2030, 451, 2496, 2596, 1373, 2870, 1039, 1364, 292]. vitro [2364, 609]. vitulina [114, 1806]. viviparous [2186]. vivo [2364, 2430]. VNTR [1981]. Vocalizations [123]. Volume [2909, 2910]. volumes [2903]. VPA [795]. vulgaris [1766]. Vulnerability [809, 2547, 1908, 2450, 1736]. Vulnerable [567].

Wafer [1095]. Wales [1696, 1497]. wall [1492]. Walleye [487, 331, 557, 1261, 287, 1069, 1372, 1009, 451, 606, 1373, 1039, 292, 2570, 1983, 2349, 1538, 1792, 2188, 2710, 2030, 2496, 2596, 2870, 1519]. Walrus [1206, 2670]. wandering [2350]. wardi [1503]. Warmed [814]. Warming [35, 122, 2444, 2292, 2540, 2585, 2482]. Warmwater [727, 646, 1956, 2002]. Warning [2]. warp [1693]. Was [1604]. Washington [703, 1868, 2527, 710, 565, 2576, 756, 1868, 2085, 2526, 1588, 1940, 1168, 863, 2006, 2463]. waste [2134]. Water [410, 285, 149, 669, 1437, 1079, 707, 162, 161, 189, 753, 1269, 2125, 1064, 1045, 1443, 210, 469, 1017, 1072, 169, 752, 1369, 196, 273, 384, 468, 584, 286, 2864, 2012, 1179, 704, 1190, 679, 572, 44, 185, 844, 433, 2769, 820, 505, 993, 1067, 218, 985, 1359, 181, 182, 1132, 2722, 1650, 2865, 1536, 2678, 1360, 2404, 1840, 1500, 1540, 2835, 2312, 2238, 2673, 2218, 1761, 1780, 2283, 2272, 1885, 2376, 2621, 2037, 2616, 2511, 1555, 2286, 1975, 2303, 1963, 2507, 2577, 2754, 2683, 2633, 2156, 1752, 633, 2002, 1931, 2100, 2441, 2868, 2302, 2686, 2768]. water [1806, 1955, 2053, 1718, 1539, 2304, 1912, 1706, 1847, 2775]. water-chemistry [1963]. Water-Column [1190]. water-quality [1780]. water-soluble [2865]. Waterborne [1052, 1200, 2867, 2001, 2633, 1906, 1820]. Waterborne-Copper [1052]. Waters [486, 271, 58, 334, 1110, 814, 102, 599, 386, 1444, 522, 1271, 1296, 1294, 927, 714, 1388, 541, 1018, 1178, 1969, 1491, 2842, 2519, 1888, 1887, 2464, 1527, 1683, 1633, 2674, 1909, 2529, 1949, 2776]. Watershed [1122, 2118, 2835, 1998, 2872, 2053]. Watersheds [638, 2618, 1697, 2747]. Wave [845, 2485]. weakfish [1640]. Weakly [1296, 1761]. weather [2278]. Weathered [1245, 2619]. Web [1295, 460, 492, 880, 1155, 289, 541, 2499, 1650, 2029, 2643, 1951, 2876, 1768, 1985, 1928, 2793, 2198, 2906, 2417, 1755]. Webs [747, 1368, 2213, 2064]. wegeneri [799]. Weight [571, 797, 279, 1415, 1278, 1991, 1511, 1546, 1488, 2154, 2153, 1503]. Weighted [131]. well [1774]. Wells [2006]. West [1627, 942, 525, 2013, 1764, 1513, 1907, 1891, 1115, 1120, 864, 1475]. west-central [1891, 864]. Western [152, 1112, 1539, 1459, 1849, 1970, 1983, 2780, 2501, 1538, 2644, 2905, 1597, 2015, 2104, 2174, 1520, 2481, 2492, 2363, 2455, 1574, 1943, 1019, 980, 732, 83, 865, 287, 1009, 2237, 224, 1526, 1168, 1388, 171, 222, 1049, 849]. Westland [2555]. Wetland [232, 2264, 2403, 2507]. Wetlands [461, 1268, 1647, 2391, 1885, 1523, 1570, 1825]. Whale [1408, 1936]. Whales [625, 1157, 579, 2013, 2647, 2772, 1862]. wheeleri [833]. Whelk [363]. Where [215, 2779, 1427]. Which [885]. White [4, 5, 611, 1371, 1196, 816, 358, 679, 438, 202, 351, 376, 446, 945, 1642, 2494, 2098, 2295, 2616, 2188, 2133, 2813, 2071, 1889, 2854, 1610, 1379]. white-spotted [2813]. Whitefish [1472, 656, 390, 527, 655, 1093, 166, 34, 743, 1354, 886, 1632, 2502, 2245]. whiting [1815]. Whole [553, 395, 779, 99, 47, 420, 75, 1320, 45, 44, 962, 181, 182, 893, 2087, 1804, 1974, 2525, 2849, 1397, 2770, 1973, 2563, 2017, 2562]. Whole-Body [553, 2770]. Whole-Lake [395, 47, 45, 44, 962, 2087, 2849, 1397, 1973]. Whole-River [779, 99]. Whole-Stream [1320, 2563, 2562]. wide [1628, 2463]. width [2637]. Widths [1065, 2091]. Wild [937, 888, 1355, 1226, 994, 1382, 560, 1442, 254, 253, 330, 625, 1478, 1381, 1380, 65, 484, 377, 1135, 624, 1096, 1751, 1987, 1958, 1509, 1590, 1839, 1614, 2385, 2621, 2074, 1657, 2820, 2509, 1696, 2813, 1679, 1989, 2591, 1515, 1610]. wild-caught [1657]. wilderness [2747]. wildfire [2692]. Will [804]. William [2021, 2024, 2026, 2274, 2700, 2022, 2025, 2726, 2767]. Williams [1555]. Wind [814, 241, 723, 1513, 1026, 1292, 1287, 2109]. Wind-Forced [814]. Wind-Induced [1026, 1513]. Wind-produced [1287]. Winds [921, 1250]. Winkler [2495]. Winnipeg [166, 686]. Winter [1384, 396, 591, 2370, 123, 1111, 645, 2500, 358, 1254, 659, 746, 1100, 494, 811, 1298, 2404, 2680, 2871, 2791, 2355, 1774, 1940, 1894, 2503, 1960, 2668, 2809, 2656]. winter- [2791]. Wintering [1120]. Wire [737, 771, 2019, 2584]. Wisconsin [989, 1032, 2721, 335, 648, 1280, 962, 2686, 2123]. Witch [598]. Within [2188, 1463, 1970, 1813, 375, 2847, 2177, 791, 469, 1868, 2526, 1123, 393, 284, 2083, 1563, 2033, 1559, 242]. within-lake [2033]. within-lakes [1559]. Within-population [2188]. without [1811, 2678, 2592, 1204]. wolffish [2683]. Wood [9, 2730, 392]. Woodland [151]. Woody [565, 127, 647, 1168, 193, 2566, 2707, 2575, 2871, 2747, 2176, 2527, 1651]. workshop [2145]. Wyoming [2109, 635].

X [2298]. X-ray [2298]. xanthurus [215, 1576]. Xenobiotic [420, 2586, 2418]. xenobiotic-inducible [2586, 2418]. Xiphias [1644].

Y-Chromosomal [464]. Yakima [1588]. Yamaguti [1677, 1663]. Yamasu [1677]. Yangtze [2083]. Year [416, 331, 435, 480, 105, 1454, 772, 969, 358, 1338, 451, 704, 653, 627, 1411, 1464, 1364, 459, 1096, 943, 1314, 1313, 2544, 1967, 1097, 2644, 2656, 2500, 1746, 1975, 2030, 1802, 2444, 1992, 2158, 1894, 1779, 2899, 2668]. Year-Class [416, 969, 1746, 2158]. Year-Classes [653]. year-long [2444]. yearling [2454]. Years [635, 1036, 1434, 1394, 931, 893, 2014, 2442, 2026, 2908, 2304]. Yellow [220, 921, 101, 48, 106, 358, 843, 202, 351, 63, 1411, 502, 146, 1068, 948, 784, 2766, 2443, 2364, 2392, 2430, 2384, 1946, 2169, 1872, 2870, 2808, 2280, 1935, 2775, 2776]. yellow-phase [2443]. yellowfin [1234]. Yellowknife [2112]. Yellowstone [1479, 2340]. Yellowtail [907, 2732, 2286]. yessoensis [994]. Yield [353, 739, 520, 1262, 568, 526, 956, 2117]. Yield-per-Recruit [1262]. Yields [638]. Yolk [1141, 182, 2041]. Yolk-Sac [1141, 182]. York [585, 1316, 2884, 747, 1183, 2138, 473, 1155, 286, 1872, 2870, 151]. Young [635, 101, 1250, 1097, 331, 105, 772, 358, 1338, 451, 1364, 459, 1967, 2373, 2328, 2644, 2656, 2500, 2030, 2596, 1992, 2473, 2554, 1894, 1689, 2668, 2214]. young-growth [2214]. Young-of-the-Year [331, 105, 772, 358, 1338, 451, 459, 1096, 943, 1314, 1313, 1097, 1967, 2644, 2656, 2500, 1992, 1894]. Young-of-Year [1364, 2030]. Yukon [2475, 2111, 2559].

Zealand [672, 1271, 2580, 2442, 1687, 1025, 996, 1135, 1883, 2000, 1828, 2162, 2555]. Zebra [1182, 980, 439, 865, 1280, 518, 864, 1246, 1264, 1387, 616, 859, 1388, 726, 440, 2280, 1492, 2728, 1849, 2129, 2501, 1538, 2488, 2723, 2296, 2103, 2543, 2015, 2253, 1782, 1520, 2426, 1563, 1759, 1760, 2071, 1589, 2090, 1528, 2347, 2870, 1716, 2363, 1899]. Zinc [280, 679, 808, 644, 2867, 1567, 1566, 2193]. Zn [1365, 1550, 1551, 800, 879, 1742, 2884, 2704]. Zone [1180, 1448, 133, 192, 469, 1232, 1345, 1109, 973, 913, 235, 2706, 2392, 2485, 2707, 2435, 1347, 1951, 1616, 2004, 1730, 1729]. Zones [1123, 2114]. Zoobenthos [47, 2280]. Zoogeographical [126]. Zoogeography [1311]. zooplanktivorous [2064]. Zooplankton [55, 1252, 320, 730, 1225, 707, 762, 645, 492, 847, 457, 1372, 468, 1404, 2452, 199, 1003, 1179, 671, 518, 443, 1203, 545, 538, 294, 687, 11, 56, 936, 290, 666, 846, 89, 1405, 2357, 218, 1326, 784, 945, 2544, 2675, 2564, 1750, 2410, 2427, 2502, 1582, 1892, 2305, 2692, 1553, 2403, 2507, 1572, 2828, 2120, 2441, 2164, 2388, 2113, 2609, 2419, 2323, 2608, 2843, 2009, 1916, 2901]. Zooplanktonic [1128, 2187]. Zostera [2219]. Zygnematacean [125]. ZYM [992].

References

Peterman:1990:SPA

[1] Randall M. Peterman. Statistical power analysis can improve fisheries research and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (1):2–15, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

      1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
      10.1139/f90-001.

Stockner:1990:LFW

[2] John G. Stockner, M. Emilia Klut, and William P. Cochlan. Leaky filters: a warning to aquatic ecologists. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):16–23, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

      com/doi/full/10.1139/f90-002.

Schnute:1990:UAA

[3] Jon T. Schnute and Laura J. Richards. A unified approach to the analysis of fish growth, maturity, and survivorship data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):24–40, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-003.

Ahlgren:1990:DSC

[4] Molly O. Ahlgren. Diet selection and the contribution of detritus to the diet of the juvenile white sucker (Catostomus commersoni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):41–48, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-004.

Ahlgren:1990:NSF

[5] Molly O. Ahlgren. Nutritional significance of facultative detritivory to the juvenile white sucker (Catostomus commetsoni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):49–54, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-005.

Marmorek:1990:APE

[6] David R. Marmorek, Michael L. Jones, Charles K. Minns, and Floyd C. Elder. Assessing the potential extent of damage to inland lakes in Eastern Canada due to acidic deposition. I. Development and evaluation of a simple “Site” model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (1):55–66, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-006.

Jones:1990:APE

[7] Michael L. Jones, Charles K. Minns, David R. Marmorek, and Floyd C. Elder. Assessing the potential extent of damage to inland lakes in Eastern Canada due to acidic deposition: II. Application of the regional model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):67–80, January 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-007.

Pike:1990:STD

[8] Daniel G. Pike and Harold E. Welch. Spatial and temporal distribution of sub-ice macrofauna in the Barrow Strait area, Northwest Territories.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):81–91, January 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-008.

Benke:1990:WDC

[9] Arthur C. Benke and J. Bruce Wallace. Wood dynamics in coastal plain blackwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (1):92–99, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-009.

Wurtsbaugh:1990:CED

[10] Wayne A. Wurtsbaugh and Therese Smith Berry. Cascading effects of decreased salinity on the plankton chemistry, and physics of the Great Salt Lake (Utah). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):100–109, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-010.

Pinel-Alloul:1990:ZSA

[11] Bernadette Pinel-Alloul, Ginette Méthot, Ghislain Verreault, and Yvan Vigneault. Zooplankton species associations in Quebec lakes: Variation with abiotic factors, including natural and anthropogenic acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):110–121, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-011.

Eberhardt:1990:SPM

[12] L. L. Eberhardt. A fur seal population model based on age structure data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):122–127, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-012.

Jackson:1990:CRE

[13] Leland J. Jackson, John G. Stockner, and Paul J. Harrison. Contribution of Rhizosolenia eriensis and Cyclotella spp. to the deep chlorophyll maximum of Sproat Lake, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):128–135, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-013.

Kline:1990:RET

[14] Thomas C. Kline, Jr., John J. Goering, Ole A. Mathisen, Patrick H. Poe, and Patrick L. Parker. Recycling of elements transported upstream by runs of Pacific salmon: I, δ ^{15}N and δ ^{13}C evidence in Sashin Creek, Southeastern Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1): 136–144, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-014.

Collie:1990:EHP

[15] Jeremy S. Collie, Randall M. Peterman, and Carl J. Walters. Experimental harvest policies for a mixed-stock fishery: Fraser River sockeye salmon, Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):145–155, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-015.

James:1990:ELS

[16] William F. James, Robert H. Kennedy, and Robert F. Gaugush. Effects of large-scale metalimnetic migration events on phosphorus dynamics in a north-temperate reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (1):156–162, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-016.

Campana:1990:DLR

[17] S. E. Campana, K. C. T. Zwanenburg, and J. N. Smith. ^{210}Pb/^{226}Ra determination of longevity in redfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):163–165, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-017.

Reinertsen:1990:EFR

[18] Helge Reinertsen, Arne Jensen, Jan Ivar Koksvik, Arnfinn Langeland, and Yngvar Olsen. Effects of fish removal on the limnetic ecosystem of a eutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1): 166–173, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-018.

Noltie:1990:IVB

[19] Douglas B. Noltie. Intrapopulation variation in the breeding of male pink salmon (Oncorhynchus gorbuscha) from a Lake Superior tributary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):174–179, January 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-019.

Madenjian:1990:NPR

[20] Charles P. Madenjian. Nighttime pond respiration rate: Oxygen or temperature dependent? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1): 180–183, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-020.

Sullivan:1990:CLA

[21] Patrick J. Sullivan, Han-Lin Lai, and Vincent F. Gallucci. A catch-at-length analysis that incorporates a stochastic model of growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):184–198, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-021.

Angradi:1990:DFC

[22] T. R. Angradi and J. S. Griffith. Diel feeding chronology and diet selection of rainbow trout (Oncorhynchus mykiss) in the Henry’s Fork of the Snake River, Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1): 199–209, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-022.

Freda:1990:EAL

[23] Joseph Freda and D. Gordon McDonald. Effects of aluminum on the leopard frog, Rana pipiens: Life stag comparisons and aluminum uptake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):210–216, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-023.

Freda:1990:ROC

[24] Joseph Freda, Vandy Cavdek, and D. Gordon McDonald. Role of organic complexation in the toxicity of aluminum to Rana pipiens embryos and Bufo americanus tadpoles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (1):217–224, January 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-024.

Anonymous:1990:LCLa

[25] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(1):225–233, January 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-025.

Mohr:1990:SDL

[26] L. C. Mohr, K. H. Mills, and J. F. Klaverkamp. Survival and development of lake trout (Salvelinus namaycush) embryos in an acidified lake in Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2): 236–243, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-026.

Hammill:1990:ARS

[27] M. O. Hammill and T. G. Smith. Application of removal sampling to estimate the density of ringed seals (Phoca hispida) in Barrow Strait, Northwest Territories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2): 244–250, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-027.

Stone:1990:CSI

[28] Jeff N. Stone and Yosef Cohen. Changes in species interactions of the Lake Superior fisheries system after the control of sea lamprey as indicated by time series models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (2):251–261, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-028.

Shortreed:1990:ENA

[29] Ken S. Shortreed and John G. Stockner. Effect of nutrient additions on lower trophic levels of an oligotrophic lake with a seasonal deep chlorophyll maximum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):262–273, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-029.

Parma:1990:DAS

[30] Ana M. Parma and Richard B. Deriso. Dynamics of age and size composition in a population subject to size-selective mortality: Effects of phenotypic variability in growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (2):274–289, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-030.

Stottlemyer:1990:PSM

[31] R. Stottlemyer and D. Toczydlowski. Pattern of solute movement from snow into an Upper Michigan stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):290–300, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-031.

Fournier:1990:MLB

[32] D. A. Fournier, John R. Sibert, Jacek Majkowski, and John Hampton. MULTIFAN a likelihood-based method for estimating growth parameters and age composition from multiple length frequency data sets illustrated using data for southern bluefin tuna (Thunnus maccoyii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):301–317, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-032.

Lambert:1990:FMD

[33] Yvan Lambert and Julian J. Dodson. Freshwater migration as a determinant factor in the somatic cost of reproduction of two anadromous coregonines of James Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (2):318–334, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-033.

Lambert:1990:IFM

[34] Yvan Lambert and Julian J. Dodson. Influence of freshwater migration on the reproductive patterns of anadromous populations of cisco (Coregonus artedii) and lake whitefish (C. clupeaformis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):335–345, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-034.

Freeland:1990:SST

[35] Howard J. Freeland. Sea surface temperatures along the Coast of British Columbia: Regional evidence for a warming trend. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):346–350, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-035.

Qliveira:1990:PGM

[36] Luis Qliveira and Hanh Huynh. Phototrophic growth of microalgae with allantoic acid or hypoxanthine serving as nitrogen source, implications for purine-N utilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2): 351–356, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-036.

Duarte:1990:PSM

[37] Carlos M. Duarte and Jacob Kalff. Patterns in the submerged macrophyte biomass of lakes and the importance of the scale of analysis in the interpretation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):357–363, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-037.

Brabrand:1990:RIP

[38] Åge Brabrand, Bjørn Andreas Faafeng, and Jens Petter Moritz Nilssen. Relative importance of phosphorus supply to phytoplankton production: Fish excretion versus external loading. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):364–372, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-038.

Wiley:1990:LSA

[39] M. J. Wiley, L. L. Osborne, and R. W. Larimore. Longitudinal structure of an agricultural prairie river system and its relationship to current stream ecosystem theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (2):373–384, February 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-039.

Crecco:1990:CDD

[40] Victor Crecco and William J. Overholtz. Causes of density-dependent catchability for Georges Bank haddock Melanogrammus aeglefinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):385–394, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-040.

Stirling:1990:VMD

[41] D. G. Stirling, D. J. McQueen, and M. R. S. Johannes. Vertical migration in Daphnia galeata mendotae (Brooks): Demographic responses to changes in planktivore abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):395–400, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-041.

Walters:1990:CPD

[42] Carl J. Walters, Donald C. E. Robinson, and Thomas G. Northcote. Comparative population dynamics of Daphnia rosea and Holopedium gibberum in four oligotrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):401–409, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-042.

Keller:1990:ELN

[43] W. Keller, D. P. Dodge, and G. M. Booth. Experimental lake neutralization program: Overview of neutralization studies in Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):410–411, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-043.

Molot:1990:WLN

[44] Lewis A. Molot, Peter J. Dillon, and Gillian M. Booth. Whole-lake and nearshore water chemistry in Bowland Lake before and after treatment with CaCO_{3}. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):412–421, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-044.

Molot:1990:RPA

[45] Lewis A. Molot, L. Heintsch, and K. H. Nicholls. Response of phytoplankton in acidic lakes in Ontario to whole-lake neutralization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):422–431, February 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-045.

Jackson:1990:ENE

[46] Michael B. Jackson, Elizabeth M. Vandermeer, Nigel Lester, Jacqueline A. Booth, Lewis Molot, and Isobelle M. Gray. Effects of neutralization and early reacidification on filamentous algae and macrophytes in Bowland Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):432–439, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-046.

Keller:1990:CZC

[47] W. Keller, Lewis A. Molot, Ronald W. Griffiths, and N. D. Yan. Changes in the zoobenthos community of acidified Bowland Lake after wholelake neutralization and lake trout (Salvelinus namaycush) reintroduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):440–445, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-047.

Gunn:1990:SGR

[48] J. M. Gunn, J. G. Hamilton, G. M. Booth, C. D. Wren, G. L. Beggs, H. J. Rietveld, and J. R. Munro. Survival, growth, and reproduction of lake trout (Salvelinus namaycush) and yellow perch (Perca flavescens) after neutralization of an acidic lake near Sudbury, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):446–453, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-048.

Anonymous:1990:CRPa

[49] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(2):455–456, February 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-049.

Winters:1990:DIE

[50] G. H. Winters and J. P. Wheeler. Direct and indirect estimation of gillnet selection curves of Atlantic herring (Clupea harengus harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):460–470, March 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-050.

Lancaster:1990:EMS

[51] Holli F. Lancaster and Ray W. Drenner. Experimental mesocosm study of the separate and interaction effects of phosphorus and mosquitofish (Gambusia affinis) on plankton community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):471–479, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-051.

Kenney:1990:LDE

[52] Bernard C. Kenney. Lake dynamics and the effects of flooding on total phosphorus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):480–485, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-052.

Johnston:1990:CGV

[53] N. T. Johnston. A comparison of the growth of vertically-migrating and nonmigrating Kokanee (Oncorhynchus nerka) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):486–491, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-053.

Bianchi:1990:PPT

[54] Thomas S. Bianchi and Stuart Findlay. Plant pigments as tracers of emergent and submergent macrophytes from the Hudson River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):492–494, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-054.

Bleiwas:1990:FRD

[55] Agnes H. Bleiwas and Pamela M. Stokes. Filtering rates of Diaptomus minutus, Bosmina spp., Diaphanosoma sp., Holopedium gibberum (Crustacea), and zooplankton community grazing rates in some acidic and circumneutral Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (3):495–504, March 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-055.

Purcell:1990:SBZ

[56] Jennifer E. Purcell. Soft-bodied zooplankton predators and competitors of larval herring (Clupea harengus pallasi) at herring spawning grounds in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (3):505–515, March 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-056.

Walters:1990:PBC

[57] Carl J. Walters. A partial bias correction factor for stock–recruitment parameter estimation in the presence of autocorrelated environmental effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):516–519, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-057.

Campbell:1990:ABL

[58] Alan Campbell. Aggregations of berried lobsters (Homatus americanus) in shallow waters off Grand Manan, Eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):520–523, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-058.

Luecke:1990:RMB

[59] Chris Luecke, James A. Rice, Larry B. Crowder, Steven E. Yeo, and Fred P. Binkowski. Recruitment mechanisms of bloater in Lake Michigan: an analysis of the predatory gauntlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):524–532, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-059.

Bernatchez:1990:MDV

[60] Louis Bernatchez and Julian J. Dodson. Mitochondrial DNA variation among anadromous populations of cisco (Coregonus artedii) as revealed by restriction analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (3):533–543, March 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-060.

Martin:1990:PTD

[61] John F. Martin, Steven M. Plakas, Janis H. Holley, Joseph V. Kitzman, and Anthony M. Guarino. Pharmacokinetics and tissue disposition of the off-flavor compound 2-methylisoborneol in the channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (3):544–547, March 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-061.

Regnell:1990:CPR

[62] Olof Regnell. Conversion and partitioning of radio-labelled mercury chloride in aquatic model systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):548–553, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-062.

Post:1990:MAL

[63] John R. Post. Metabolic allometry of larval and juvenile yellow perch (Perca flavescens): In situ estimates and bioenergetic models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):554–560, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-063.

Read:1990:ASM

[64] Andrew J. Read. Age at sexual maturity and pregnancy rates of harbour porpoises Phocoena phocoena from the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):561–565, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-064.

Swain:1990:VAB

[65] D. P. Swain and B. E. Riddell. Variation in agonistic behavior between newly emerged juveniles from hatchery and wild populations of Coho salmon, Oncorhynchus kisutch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):566–571, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-065.

Robbins:1990:TSS

[66] John A. Robbins, Alena Mudroch, and Barry G. Oliver. Transport and storage of ^{137}Cs and ^{210}Pb in sediments of Lake St. Clair. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):572–587, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-066.

Metzler:1990:SDS

[67] Glenn M. Metzler and Leonard A. Smock. Storage and dynamics of subsurface detritus in a sand-bottomed stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):588–594, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-067.

Parma:1990:EHC

[68] Ana M. Parma and Richard B. Deriso. Experimental harvesting of cyclic stocks in the face of alternative recruitment hypotheses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):595–610, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-068.

Gomulkiewicz:1990:RLG

[69] Richard Gomulkiewicz, Jon K. T. Brodziak, and Marc Mangel. Ranking loci for genetic stock identification by curvature methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):611–619, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-069.

Smouse:1990:GMA

[70] Peter E. Smouse, Robin S. Waples, and Joseph A. Tworek. A genetic mixture analysis for use with incomplete source population data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):620–634, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-070.

Hilborn:1990:DFM

[71] Ray Hilborn. Determination of fish movement patterns from tag recoveries using maximum likelihood estimators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):635–643, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-071.

Kelso:1990:EEA

[72] J. R. M. Kelso, M. A. Shaw, C. K. Minns, and K. H. Mills. An evaluation of the effects of atmospheric acidic deposition on fish and the fishery resource of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (3):644–655, March 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-072.

Aloi:1990:CRR

[73] Jane E. Aloi. A critical review of recent freshwater periphyton field methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(3):656–670, March 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-073.

Welch:1990:QGF

[74] David W. Welch and Gordon A. McFarlane. Quantifying the growth of female Pacific hake (Merluccius productus): an example of measuring uncertainty and bias in non-linear parameter estimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):672–681, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-074.

Madenjian:1990:EWP

[75] Charles P. Madenjian, Gary L. Rogers, and Arlo W. Fast. Estimation of whole pond respiration rate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (4):682–686, April 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-075.

Sterner:1990:LML

[76] Robert W. Sterner. Lake morphometry and light in the surface layer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):687–692, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-076.

Burt:1990:SPL

[77] M. D. B. Burt, J. D. Campbell, C. G. Likely, and J. W. Smith. Serial passage of larval Pseudoterranova decipiens (Nematoda: Ascaridoidea) in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):693–695, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-077.

Ralston:1990:SSS

[78] Stephen Ralston. Size selection of snappers (Lutjanidae) by hook and line gear. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):696–700, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-078.

Johnston:1990:RRR

[79] C. E. Johnston, S. R. Farmer, R. W. Gray, and M. Hambrook. Reconditioning and reproductive responses of Atlantic salmon kelts (Salmo salar) to photoperiod and temperature manipulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):701–710, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-079.

Rader:1990:DMM

[80] Russell B. Rader and James V. Ward. Diel migration and microhabitat distribution of a benthic stream assemblage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):711–718, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-080.

Gile:1990:CMG

[81] Stephen R. Gile and Moira M. Ferguson. Crossing methodology and genotypic diversity in a hatchery strain of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):719–724, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-081.

Klut:1990:VLP

[82] M. Emilia Klut and John G. Stockner. Virus-like particles in an ultraoligotrophic lake on Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):725–730, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-082.

Garton:1990:TTP

[83] David W. Garton, David J. Berg, and Robert J. Fletcher. Thermal tolerances of the predatory cladocerans Bythotrephes cederstroemi and Leptodora kindti: Relationship to seasonal abundance in Western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):731–738, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-083.

MacFarlane:1990:SLS

[84] R. Bruce MacFarlane, H. Rodger Harvey, Michael J. Bowers, and John S. Patton. Serum lipoproteins in striped bass (Morone saxatilis): Effects of starvation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):739–745, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-084.

Larsson:1990:CSS

[85] Per Larsson, Lennart Okla, Sven-Olof Ryding, and Bengt Westöö. Contaminated sediment as a source of PCBs in a river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):746–754, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-085.

Stewart:1990:EED

[86] Mark W. Stewart, Richard L. Saunders, and A. J. Wiggs. Effects of extended daylength on autumn growth dynamics of juvenile Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (4):755–759, April 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-086.

Gauldie:1990:PDB

[87] R. W. Gauldie. Phase differences between check ring locations in the orange roughy otolith (Hoplostethus atlanticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):760–765, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-087.

Plakas:1990:TDR

[88] S. M. Plakas, R. W. Dickey, M. G. Barron, and A. M. Guarino. Tissue distribution and renal excretion of ormetoprim after intravascular and oral administration in the channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):766–771, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-088.

Sierszen:1990:EEL

[89] Michael E. Sierszen and Thomas M. Frost. Effects of an experimental lake acidification on zooplankton feeding rates and selectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):772–779, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-089.

DeSeve:1990:SCI

[90] Michèle A. De Sève and Maxwell J. Dunbar. Structure and composition of ice algal assemblages from the Gulf of St. Lawrence, Magdalen Islands Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):780–788, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-090.

Konkle:1990:UAS

[91] Blake R. Konkle, Nicholas C. Collins, and Robert L. Baker. Use of artificial substrates to estimate prey resources available to a visually feeding benthivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (4):789–793, April 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-091.

Simard:1990:VSP

[92] Yvan Simard and Louise Savard. Variability, spatial patterns and scales of similarity in size–frequency distributions of the northern shrimp (Pandalus borealis) and its migrations in the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):794–804, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-092.

Thomas:1990:ESP

[93] G. L. Thomas, S. L. Thiesfeld, S. A. Bonar, R. N. Crittenden, and G. B. Pauley. Estimation of submergent plant bed biovolume using acoustic range information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (4):805–812, April 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-093.

Cunjak:1990:SVS

[94] R. A. Cunjak, R. L. Saunders, and E. M. P. Chadwick. Seasonal variations in the smolt characteristics of juvenile Atlantic salmon (Salmo salar) from estuarine and riverine environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):813–820, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-094.

Minns:1990:APE

[95] C. K. Minns, J. E. Moore, D. W. Schindler, and M. L. Jones. Assessing the potential extent of damage to inland lakes in Eastern Canada due to acidic deposition. III. Predicted impacts on species richness in seven groups of aquatic biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (4):821–830, April 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-095.

Gill:1990:ETM

[96] Kenneth A. Gill and Patrick J. Walsh. Effects of temperature on metabolism of benzo[a]pyrene by toadfish (Opsanus beta) hepatocytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):831–837, April 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-096.

Welch:1990:CDO

[97] D. W. Welch and D. J. Noakes. Cyclic dominance and optimal escapement of Adams River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(4):838–849, April 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-097.

Shirvell:1990:RIR

[98] C. S. Shirvell. Role of instream rootwads as juvenile Coho salmon (Oncorhynchus kisutch) and steelhead trout (O. mykiss) cover habitat under varying streamflows. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (5):852–861, May 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-098.

Johnston:1990:IJS

[99] N. T. Johnston, C. J. Perrin, P. A. Slaney, and B. R. Ward. Increased juvenile salmonid growth by whole-river fertilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):862–872, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-099.

Singer:1990:AEM

[100] Thomas D. Singer, Vhundi G. Mahadevappa, and James S. Ballantyne. Aspects of the energy metabolism of lake sturgeon, Acipenser fulvescens, with special emphasis on lipid and ketone body metabolism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):873–881, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-100.

Confer:1990:IPA

[101] John L. Confer, Edward L. Mills, and Linda O’Bryan. Influence of prey abundance on species and size selection by young yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):882–887, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-101.

Cooper:1990:HPD

[102] William J. Cooper and Richard G. Zepp. Hydrogen peroxide decay in waters with suspended soils: Evidence for biologically mediated processes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):888–893, May 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-102.

Smith:1990:USM

[103] Stephen J. Smith. Use of statistical models for the estimation of abundance from groundfish trawl survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):894–903, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-103.

Belanger:1990:VCF

[104] Scott E. Belanger, Jerry L. Farris, Donald S. Cherry, and John Cairns, Jr. Validation of Corbicula fluminea growth reductions induced by copper in artificial streams and river systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):904–914, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-104.

Grant:1990:AFB

[105] James W. A. Grant. Aggressiveness and the foraging behaviour of young-of-the-year brook charr (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):915–920, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-105.

Hayes:1990:RSY

[106] Daniel B. Hayes and William W. Taylor. Reproductive strategy in yellow perch (Perca flavescens): Effects of diet ontogeny, mortality, and survival costs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):921–927, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
106.

Quiros:1990:PRF

[107] Rolando Quiros. Predictors of relative fish biomass in lakes and reservoirs of Argentina. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):928–939, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
107.

Hanna:1990:EMP

[108] Micheline Hanna. Evaluation of models predicting mixing depth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):940–947, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-108.

Richards:1990:MMM

[109] Laura J. Richards, Jon T. Schnute, and Claudia M. Hand. A multivariate maturity model with a comparative analysis of three lingcod (Ophiodon elongatus) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):948–959, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-

Mooij:1990:GRR

[110] W. M. Mooij and O. F. R. van Tongeren. Growth of 0+ roach (Rutilus rutilus) in relation to temperature and size in a shallow eutrophic lake: Comparison of field and laboratory observations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):960–967, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-110.

Waples:1990:TCA

[111] Robin S. Waples. Temporal changes of allele frequency in Pacific salmon: Implications for mixed-stock fishery analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):968–976, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-111.

Lehman:1990:MRL

[112] John T. Lehman, James A. Bowers, Robert W. Gensemer, Glenn J. Warren, and Donn K. Branstrator. Mysis relicta in Lake Michigan: Abundances and relationships with their potential prey, Daphnia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):977–983, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-112.

Grewe:1990:PRA

[113] Peter M. Grewe, Neil Billington, and Paul D. N. Hebert. Phylogenetic relationships among members of Salvelinus inferred from mitochondrial DNA divergence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):984–991, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
113.

Olesiuk:1990:RTA

[114] Peter F. Olesiuk, Michael A. Bigg, and Graeme M. Ellis. Recent trends in the abundance of harbour seals, Phoca vitulina, in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):992–1003, May 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-114.

Pierce:1990:SLF

[115] Clay L. Pierce, Joseph B. Rasmussen, and William C. Leggett. Sampling littoral fish with a seine: Corrections for variable capture efficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):1004–1010, May 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-115.

Himmelman:1990:UFA

[116] John H. Himmelman and Henri Nédélec. Urchin foraging and algal survival strategies in intensely grazed communities in Eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):1011–1026, May 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-116.

Robarts:1990:HBP

[117] Richard D. Robarts and Richard J. Wicks. Heterotrophic bacterial production and its dependence on autotrophic production in a hypertrophic African reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):1027– 1037, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-117.

Denizeau:1990:TCC

[118] F. Denizeau and M. Marion. Toxicity of cadmium, copper, and mercury to isolated trout hepatocytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (5):1038–1042, May 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-118.

Giguere:1990:TLI

[119] Louis A. Giguère and Robert L. Dunbrack. Thin layer interference may reduce the visibility of transparent phantom midge larvae (Chaoborus trivittatus) to predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (5):1043–1046, May 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-119.

Pinel-Alloul:1990:PQL

[120] Bernadette Pinel-Alloul, Ginette Méthot, Ghislain Verrault, and Yvan Vigneault. Phytoplankton in Quebec lakes: Variation with lake morphometry, and with natural and anthropogenic acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(5):1047–1057, May 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-120.

Stephenson:1990:RSA

[121] Robert L. Stephenson and Irv Kornfield. Reappearance of spawning Atlantic herring (Clupea harengus harengus) on Georges Bank: Population resurgence not recolonization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1060–1064, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-121.

Meisner:1990:ECW

[122] J. Donald Meisner. Effect of climatic warming on the southern margins of the native range of brook trout, Salvelinus fontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1065–1070, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-122.

Cleator:1990:WDB

[123] Holly J. Cleator and Ian Stirling. Winter distribution of bearded seals (Erignathus barbatus) in the Penny Strait area, Northwest Territories, as determined by underwater vocalizations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1071–1076, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-123.

Currie:1990:PDV

[124] David J. Currie. Phosphorus deficiency and its variation among lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1077–1084, June 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-124.

Howell:1990:CZC

[125] E. T. Howell, M. A. Turner, R. L. France, M. B. Jackson, and P. M. Stokes. Comparison of zygnematacean (Chlorophyta) algae in the metaphyton of two acidic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (6):1085–1092, June 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-125.

Billington:1990:AMD

[126] Neil Billington, Paul D. N. Hebert, and Robert D. Ward. Allozyme and mitochondrial DNA variation among three species of Stizostedion (Percidae): Phylogenetic and zoogeographical implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1093–1102, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-126.

Carlson:1990:WDC

[127] J. Y. Carlson, C. W. Andrus, and H. A. Froehlich. Woody debris, channel features, and macroinvertebrates of streams with logged and undisturbed riparian timber in Northeastern Oregon, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1103–1111, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-127.

Lokkeborg:1990:RCU

[128] Svein Løkkeborg. Reduced catch of under-sized cod (Gadus morhua) in longlining by using artificial bait. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1112–1115, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-128.

Abrahams:1990:VCA

[129] M. V. Abrahams and M. C. Healey. Variation in the competitive abilities of fishermen and its influence on the spatial distribution of the British Columbia salmon troll fleet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (6):1116–1121, June 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-129.

Doucet:1990:ECI

[130] Christine M. Doucet and Edward J. Maly. Effect of copper on the interaction between the predator Didinium nasutum and its prey Paramecium caudatum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1122–1127, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
130.

Marx:1990:WML

[131] Brian D. Marx and Eric P. Smith. Weighted multicollinearity in logistic regression: Diagnostics and biased estimation techniques with an example from lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (6):1128–1135, June 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-131.

Birt:1990:SSD

[132] Tim P. Birt, John M. Green, and William S. Davidson. Smolting status of downstream migrating Atlantic salmon (Salmo salar) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1136–1139, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-132.

Duff:1990:DSH

[133] John H. Duff and Frank J. Triska. Denitrifïcations in sediments from the hyporheic zone adjacent to a small forested stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1140–1147, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-133.

Richards:1990:UGD

[134] Laura J. Richards and Jon T. Schnute. Use of a general dose– response model for rockfish fecundity–length relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1148–1156, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-134.

Brazner:1990:ECD

[135] John C. Brazner and Edward R. Kline. Effects of chlorpyrifos on the diet arid growth of larval fathead minnows, Pimephales promelas, in littoral enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1157–1165, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
135.

Leavitt:1990:RPS

[136] P. R. Leavitt and S. R. Carpenter. Regulation of pigment sedimentation by photo-oxidation and herbivore grazing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1166–1176, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-136.

Campbell:1990:LHA

[137] Alan Campbell. The lobster (Homarus americanus) fishery off Lower Argyle, Southwestern Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1177–1184, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-137.

Kjesbu:1990:ORM

[138] O. S. Kjesbu, P. R. Witthames, P. Solemdal, and M. Greer WalKer. Ovulatory rhythm and a method to determine the stage of spawning in Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (6):1185–1193, June 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-138.

Reimchen:1990:SSM

[139] T. E. Reimchen. Size-structured morality in a treespine stickleback (Gastrosteus aculeatus)–cutthroat trout (Oncorhynchus clarki) community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1194–1205, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-139.

Robertson:1990:TSM

[140] Dale M. Robertson and Robert A. Ragotzkie. Thermal structure of a multibasin lake: Influence of morphometry, interbasin exchange, and groundwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1206– 1212, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-140.

Allen:1990:PTP

[141] Standish K. Allen, Jr. and Sandra L. Downing. Performance of tripfoid Pacific oysters, Crassostrea gigas: Gametogenesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1213–1222, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-141.

Ryg:1990:EBC

[142] Morten Ryg, Christian Lydersen, Nina H. Markussen, Thomas G. Smith, and Nils Are Øritsland. Estimating the blubber content of phocid seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1223–1227, June 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-142.

Miller:1990:ECL

[143] Robert J. Miller. Effectiveness of crab and lobster traps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(6):1228–1251, June 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-143.

Li:1990:PPF

[144] William K. W. Li. Particles in “Particle-free” seawater: Growth of ultraphytoplankton and implications for dilution experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1258–1268, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-144.

Dillon:1990:RRP

[145] P. J. Dillon, R. D. Evans, and Lewis A. Molot. Retention and resuspension of phosphorus, nitrogen, and iron in a Central Ontario lake.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1269–1274, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-145.

Shields:1990:BME

[146] Jeffrey D. Shields, Robert K. Okazaki, and Armand M. Kuris. Brood mortality and egg predation by the Nemertean, Carcinonemertes epialti, on the yellow rock crab, Cancer anthonyi, in Southern California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1275–1281, July 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-146.

Jolly:1990:SRT

[147] G. M. Jolly and I. Hampton. A stratified random transect design for acoustic surveys of fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1282–1291, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-147.

Danzmann:1990:MEB

[148] Roy G. Danzmann, Glen J. Van Der Kraak, Thomas T. Chen, and Dennis A. Powers. Metabolic effects of bovine growth hormone and genetically engineered rainbow trout growth hormone in rainbow trout (Oncorhynchus mykiss) reared at a high temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1292–1301, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-148.

Bernard:1990:PRW

[149] F. R. Bernard and D. J. Noakes. Pumping rates, water pressures, and oxygen use in eight species of marine bivalve molluscs from British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1302–1306, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
149.

Stevens:1990:TDG

[150] Bradley G. Stevens. Temperature-dependent growth of juvenile red king crab (Paralithodes camtschatica) and its effects on size-at-age and subsequent recruitment in the Eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1307–1317, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-150.

Smith:1990:RBA

[151] Michael E. Smith, Barbara J. Wyskowski, Carol M. Brooks, Charles T. Driscoll, and Christina C. Cosentini. Relationships between acidity and benthic invertebrates of low-order woodland streams in the Adirondack Mountains, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (7):1318–1329, July 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-151.

Phillips:1990:EDM

[152] Bruce F. Phillips. Estimating the density and mortality of juvenile western rock lobsters (Panulirus cygnus) in nursery reefs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1330–1338, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-152.

Siver:1990:UEP

[153] Peter A. Siver and Josephine S. Hamer. Use of extant populations of scaled chrysophytes for the inference of lakewater pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1339–1347, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-153.

Smith:1990:VFS

[154] Ralph E. H. Smith, W. Glen Harrison, Leslie R. Harris, and Alex W. Herman. Vertical fine structure of particulate matter and nutrients in sea ice of the High Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (7):1348–1355, July 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-154.

Crawford:1990:PER

[155] William R. Crawford, Albert V. Tyler, and Richard E. Thomson. A possible eddy retention mechanism for ichthyoplankton in Hecate Strait.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1356–1363, July 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-155.

Ardisson:1990:MAS

[156] P.-L. Ardisson, E. Bourget, and P. Legendre. Multivariate approach to study species assemblages at large spatiotemporal scales: The community structure of the epibenthic fauna of the estuary and Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1364–1377, July 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-156.

Findlay:1990:PCL

[157] D. L. Findlay and S. E. M. Kasian. phytoplankton communities of lakes experimentally acidified with sulfuric and nitric acids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1378–1386, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-157.

Kreutzweiser:1990:RBT

[158] D. P. Kreutzweiser. Response of a brook trout (Salvelinus fontinalis) population to a reduction in stream benthos following an insecticide treatment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1387–1401, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-158.

Cota:1990:IIA

[159] Glenn F. Cota, Jeffrey L. Anning, Leslie R. Harris, W. Glen Harrison, and Ralph E. H. Smith. Impact of ice algae on inorganic nutrients in seawater and sea ice in Barrow Strait, NWT, Canada, during spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1402–1415, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-159.

Cerrato:1990:IST

[160] Robert M. Cerrato. Interpretable statistical tests for growth comparisons using parameters in the von Bertalanffy equation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1416–1426, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-160.

Dunstall:1990:VLW

[161] T. G. Dunstall, D. W. Lawler, R. Farooqui, and G. T. Haymes. Variation in lake water temperature in the Nanticoke Region of Long Point Bay, Lake Erie, during the open-water season. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1427–1433, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-161.

Dunstall:1990:IUS

[162] T. G. Dunstall, J. C. H. Carter, B. P. Monroe, G. T. Haymes, R. R. Weiler, and G. J. Hopkins. Influence of upwellings, storms, and generating station operation on water chemistry and plankton in the Nanticoke Region of Long Point Bay, Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1434–1445, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-162.

Wilkinson:1990:EFC

[163] Kevin J. Wilkinson, Peter G. C. Campbell, and Pierre Couture. Effect of fluoride complexation on aluminum toxicity towards juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (7):1446–1452, July 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-163.

Schnute:1990:EVB

[164] J. T. Schnute, T. J. Mulligan, and B. R. Kuhn. An errors-in-variables bias model with an application to salmon hatchery data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1453–1467, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-164.

Elser:1990:PNL

[165] James J. Elser, Erich R. Marzolf, and Charles R. Goldman. Phosphorus and nitrogen limitation of phytoplankton growth in the freshwaters of North America: a review and critique of experimental enrichments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(7):1468–1477, July 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-165.

Kristofferson:1990:SSL

[166] A. H. Kristofferson and J. W. Clayton. Subpopulation status of lake whitefish (Coregonus clupeaformis) in Lake Winnipeg. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1484–1494, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-166.

Gillis:1990:UOO

[167] D. J. Gillis, B. A. McKeown, and D. E. Hay. Ultrastructural observations on the ovary and eggs, and the development of egg adhesion in Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1495–1504, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-167.

Gillis:1990:PHA

[168] D. J. Gillis, B. A. McKeown, and D. E. Hay. Physiological and histological aspects of late oocyte provisioning, ovulation, and fertilization in Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1505–1512, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-168.

Hoffnagle:1990:SPT

[169] Timothy L. Hoffnagle and Albert J. Fivizzani. Stimulation of plasma thyroxine levels by novel water chemistry during smoltification in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1513–1517, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-169.

Dabrowski:1990:EDA

[170] Konrad Dabrowski, Reinhard Lackner, and Cristine Doblander. Effect of dietary ascorbate on the concentration of tissue ascorbic acid, dehydroascorbic acid, ascorbic sulfate, and activity of ascorbic sulfate sulfohydrolase in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1518–1525, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-170.

Simard:1990:DGP

[171] Yvan Simard, Pierre Brunel, and Johanne Lacelle. Distribution and growth of pre-recruit cohorts of the northern shrimp (Pandalus borealis) in the Western Gulf of St. Lawrence as related to hydrographic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1526–1533, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
171.

Chadwick:1990:PAL

[172] E. M. P. Chadwick, D. K. Cairns, H. M. C. Dupuis, K. V. Ewart, M. H. Kao, and G. L. Fletcher. Plasma antifreeze levels reflect the migratory behaviour of Atlantic herring (Clupea harengus harengus) in the Southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (8):1534–1536, August 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-172.

Eckhardt:1990:CDO

[173] B. W. Eckhardt and T. R. Moore. Controls on dissolved organic carbon concentrations in streams, Southern Québec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1537–1544, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-173.

Bergmann:1990:NFE

[174] Martin A. Bergmann and Harold E. Welch. Nitrogen fixation by epilithic periphyton in small Arctic lakes in response to experimental nitrogen and phosphorus fertilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (8):1545–1550, August 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-174.

McFarlane:1990:EET

[175] Gordon A. McFarlane and Richard J. Beamish. Effect of an external tag on growth of sabiefish (Anoplopoma fimbria), and consequences to mortality and age at maturity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (8):1551–1557, August 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-175.

Playle:1990:PAF

[176] Richard C. Playle and Chris M. Wood. Is precipitation of aluminum fast enough to explain aluminum deposition on fish gills? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1558–1561, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-176.

Shields:1990:CMD

[177] Barbara A. Shields, Kevin S. Guise, and James C. Underhill. Chromosomal and mitochondrial DNA characterization of a population of dwarf cisco (Coregonus artedii) in Minnesota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1562–1569, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-177.

Prager:1990:CLR

[178] Michael H. Prager and Mary C. Fabrizio. Comparison of logistic regression and discriminant analyses for stock identification of anadromous fish, with application to striped bass (Morone saxatilis) and American shad (Alosa sapidissima). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8): 1570–1577, August 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-178.

Bergman:1990:LAF

[179] H. L. Bergman and J. S. Mattice. Lake acidification and fisheries project: Brook trout (Salvelinus fontinalis) early life stages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1578–1579, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-179.

Ingersoll:1990:EPA

[180] C. G. Ingersoll, D. R. Mount, D. D. Gulley, T. W. La Point, and H. L. Bergman. Effects of pH, aluminum, and calcium on survival and growth of eggs and fry of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1580–1592, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-180.

Wood:1990:EWA

[181] C. M. Wood, D. G. McDonald, C. G. Ingersoll, D. R. Mount, O. E. Johannsson, S. Landsberger, and H. L. Bergman. Effects of water acidity, calcium, and aluminum on whole body ions of brook trout (Salvelinus fontinalis) continuously exposed from fertilization to swimup: a study by instrumental neutron activation analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1593–1603, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-181.

Wood:1990:WBI

[182] C. M. Wood, D. G. McDonald, C. G. Ingersoll, D. R. Mount, O. E. Johannsson, S. Landsberger, and H. L. Bergman. Whole body ions of brook trout (Salvelinus fontinalis) alevins: Responses of yolk-sac and swim-up stages to water acidity, calcium, and aluminum, and recovery effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1604–1615, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-182.

Ingersoll:1990:ERP

[183] C. G. Ingersoll, D. A. Sanchez, J. S. Meyer, D. D. Gulley, and J. E. Tietge. Epidermal response to pH, aluminum, and calcium exposure in brook trout (Salvelinus fontinalis) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1616–1622, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-183.

Mount:1990:RBT

[184] D. R. Mount, M. J. Swanson, J. E. Breck, A. M. Farag, and H. L. Bergman. Responses of brook trout (Salvelinus fontinalis) fry to fluctuating acid, aluminum, and low calcium exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1623–1630, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-184.

Parkhurst:1990:IMA

[185] Benjamin R. Parkhurst, Harold L. Bergman, Joseph Fernandez, David D. Gulley, J. Russell Hockett, and Deborah A. Sanchez. Inorganic monomeric aluminum and pH as predictors of acidic water toxicity to brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1631–1640, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-185.

Ingersoll:1990:AAT

[186] C. G. Ingersoll, D. D. Gulley, D. R. Mount, M. E. Mueller, J. D. Fernandez, J. R. Hockett, and H. L. Bergman. Aluminum and acid toxicity to two strains of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1641–1648, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-186.

Anonymous:1990:CRPb

[187] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(8):1649– 1650, August 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f90-187.

Morgan:1990:RRB

[188] M. J. Morgan, L. L. Fancey, and J. W. Kiceniuk. Response and recovery of brain acetylcholinesterase activity in Atlantic salmon (Salmo salar) exposed to fenitrothion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9): 1652–1654, September 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-188.

Emerson:1990:ISD

[189] Craig W. Emerson. Influence of sediment disturbance and water flow on the growth of the soft-shell clam, Mya arenaria L. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1655–1663, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-189.

Wildhaber:1990:TBB

[190] Mark L. Wildhaber and Larry B. Crowder. Testing a bioenergeticsbased habitat choice model: Bluegill (Lepomis macrochirus) responses to food availability and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1664–1671, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-190.

Jordan:1990:GVM

[191] W. C. Jordan, A. F. Youngson, and J. H. Webb. Genetic variation at the malic enzyme-2 locus and age at maturity in sea-run Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9): 1672–1677, September 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-191.

Frechet:1990:CVC

[192] Alain Fréchet. Catchability variations of cod in the marginal ice zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1678–1683, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-192.

Robison:1990:CCW

[193] E. George Robison and Robert L. Beschta. Characteristics of coarse woody debris for several coastal streams of Southeast Alaska, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1684–1693, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-193.

Shaw:1990:ECP

[194] Margo A. Shaw and Gerry L. Mackie. Effects of calcium and pH on the reproductive success of Amnicola limosa (Gastropoda). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1694–1699, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-194.

Keller:1990:SFE

[195] Aimee A. Keller, Lynn L. Beatty, Laura E. Weber, and Cynthia A. Heil. Soluble fluorescence: Effects on chlorophyll determination at different salinities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1700– 1709, September 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-195.

Hudon:1990:DSF

[196] Christiane Hudon. Distribution of shrimp and fish by-catch assemblages in the Canadian Eastern Arctic in relation to water circulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1710–1723, September 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-196.

Grant:1990:TSP

[197] James W. A. Grant and Donald L. Kramer. Territory size as a predictor of the upper limit to population density of juvenile salmonids in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1724–1737, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-197.

Evans:1990:LLR

[198] Marlene S. Evans. Large-lake responses to declines in the abundance of a major fish planktivore — the Lake Michigan example. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1738–1754, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-198.

Levy:1990:RDV

[199] David A. Levy. Reciprocal diel vertical migration behavior in planktivores and zooplankton in British Columbia lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1755–1764, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-199.

Emlen:1990:DDS

[200] J. M. Emlen, R. R. Reisenbichler, A. M. McGie, and T. E. Nickelson. Density-dependence at sea for Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1765–1772, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-200.

Young:1990:ORA

[201] R. J. Young, J. R. M. Kelso, and J. G. Weise. Occurrence, relative abundance, and size of landlocked sea lamprey (Petromyzon marinus) ammocoetes in relation to stream characteristics in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1773–1778, September 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-201.

Parrish:1990:IBW

[202] Donna L. Parrish and F. Joseph Margraf. Interactions between white perch (Morone amerićana) and yellow perch (Perca flavescens) in Lake Erie as determined from feeding and growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1779–1787, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-202.

Jackson:1990:RAS

[203] Donald A. Jackson, Harold H. Harvey, and Keith M. Somers. Ratios in aquatic sciences: Statistical shortcomings with mean depth and the morphoedaphic index. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9): 1788–1795, September 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-203.

Levy:1990:SMS

[204] David A. Levy. Sensory mechanism and selective advantage for diel vertical migration in juvenile sockeye salmon, Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1796–1802, September 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-204.

Crossman:1990:RHM

[205] E. J. Crossman. Reproductive homing in muskellunge, Esox masquinongy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1803–1812, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-205.

Hudson:1990:MEB

[206] J. J. Hudson, J. C. Roff, and B. K. Burnison. Measuring epilithic bacterial production in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (9):1813–1820, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f90-206.

Gubala:1990:EIC

[207] Chad P. Gubala, Daniel R. Engstrom, and Jeffrey R. White. Effects of iron cycling on ^{210}Pb dating of sediments in an Adirondack lake, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1821–1829, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-207.

Dawe:1990:OSS

[208] E. G. Dawe, J. C. Shears, N. E. Balch, and R. K. O’Dor. Occurrence, size, and sexual maturity of long-finned squid (Loligo pealei) at Nova Scotia and Newfoundland, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1830–1835, September 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-208.

Pernie:1990:MIS

[209] Gwenyth Laird Pernie, Donald Scavia, Michael L. Pace, and Hunter J. Carrick. Micrograzer impact and substrate limitation of bacterioplankton in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(9):1836– 1841, September 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-209.

Healey:1990:DCT

[210] Michael C. Healey, Richard E. Thomson, and John F. T. Morris. Distribution of commercial troll fishing vessels off Southwest Vancouver Island in relation to fishing success and oceanic water properties and circulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1846–1864, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-210.

Stoermer:1990:SMS

[211] Eugene F. Stoermer, Claire L. Schelske, and Julie A. Wolin. Siliceous microfossil succession in the sediments of McLeod Bay, Great Slave Lake, Northwest Territories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10): 1865–1874, October 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-211.

Mangel:1990:PAS

[212] Marc Mangel and Paul E. Smith. Presence-absence sampling for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10): 1875–1887, October 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-212.

Verta:1990:CFM

[213] Matti Verta. Changes in fish mercury concentrations in an intensively fished lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1888–1897, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-213.

Fortier:1990:LHC

[214] Louis Fortier and Jacques A. Gagné. Larval herring (Clupea harengus) dispersion, growth, and survival in the St. Lawrence Estuary: Match/ mismatch or membership/vagrancy? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1898–1912, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-214.

Feller:1990:MCT

[215] Robert J. Feller, Bruce C. Coull, and Brian T. Hentschel. Meiobenthic copepods: Tracers of where juvenile Leiostomus xanthurus (Pisces) feed? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1913–1919, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-215.

Thorne:1990:AOG

[216] Richard E. Thorne and G. L. Thomas. Acoustic observations of gas bubble release by Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1920–1928, October 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-216.

Downing:1990:FPC

[217] John A. Downing, Céline Plante, and Sophie Lalonde. Fish production correlated with primary productivity, not the morphoedaphic index. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1929–1936, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-217.

Tessier:1990:IWC

[218] Alan J. Tessier and Richard J. Horwitz. Influence of water chemistry on size structure of zooplankton assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1937–1943, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-218.

Dew:1990:BEP

[219] C. Braxton Dew. Behavioral ecology of podding red king crab, Paralithodes camtschatica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10): 1944–1958, October 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-219.

Aalto:1990:AES

[220] S. K. Aalto and G. E. (Buck) Newsome. Additional evidence supporting demic behaviour of a yellow perch (Perca flavescens) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1959–1962, October 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-220.

Ruggerone:1990:HVM

[221] Gregory T. Ruggerone, Thomas P. Quinn, Ian A. McGregor, and Thomas D. Wilkinson. Horizontal and vertical movements of adult steelhead trout, Oncorhynchus mykiss, in the Dean and Fisher Channels, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10): 1963–1969, October 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-221.

Soster:1990:BRD

[222] Frederick M. Soster and Peter L. McCall. Benthos response to disturbance in Western Lake Erie: Field experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1970–1985, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-222.

Bunch:1990:BPB

[223] J. N. Bunch and R. C. Harland. Bacterial production in the bottom surface of sea ice in the Canadian subarctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1986–1995, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-223.

McCall:1990:BRD

[224] Peter L. McCall and Frederick M. Soster. Benthos response to disturbance in Western Lake Erie: Regional faunal surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):1996–2009, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-224.

Liskauskas:1990:EHF

[225] Arunas P. Liskauskas and Moira M. Ferguson. Enzyme heterozygosity and fecundity in a naturalized population of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):2010–2015, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f90-
225.

Quinn:1990:TMP

[226] Terrance J. Quinn II, Robert Fagen, and Jie Zheng. Threshold management policies for exploited populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):2016–2029, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-226.

Rasmussen:1990:FCS

[227] J. B. Rasmussen, D. J. Rowan, D. R. S. Lean, and J. H. Carey. Food chain structure in Ontario lakes determines PCB levels in lake trout (Salvelinus namaycush) and other pelagic fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):2030–2038, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-227.

Hughes:1990:PCD

[228] Nicholas F. Hughes and Lawrence M. Dill. Position choice by driftfeeding salmonids: Model and test for Arctic grayling (Thymallus arcticus) in subarctic mountain streams, interior Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):2039–2048, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-228.

Chalifour:1990:OIT

[229] Alain Chalifour, Jacques Boisvert, and Christian Back. Optimization of insecticide treatments in rivers: an application of graph theory for planning a black fly larvae control program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):2049–2056, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-229.

McCormick:1990:DIE

[230] Paul V. McCormick. Direct and indirect effects of consumers on benthic algae in isolated pools of an ephemeral stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(10):2057–2065, October 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-230.

Ouellet:1990:DSP

[231] Patrick Ouellet, Denis Lefaivre, and Vladimir Koutitonsky. Distribution of shrimp (Pandalus borealis) larvae and hydrographic pattern in the Northern Gulf of St Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2068–2078, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-231.

Shreffler:1990:TRJ

[232] David K. Shreffler, Charles A. Simenstad, and Ronald M. Thom. Temporary residence by juvenile salmon in a restored estuarine wetland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2079–2084, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-232.

Smith:1990:MFC

[233] Courtland L. Smith and Susan S. Hanna. Measuring fleet capacity and capacity utilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11): 2085–2091, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-233.

Waples:1990:TDD

[234] Robin S. Waples and Paul B. Aebersold. Treatment of data for duplicated gene loci in mixed-stock fishery analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2092–2098, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-234.

Triska:1990:IEF

[235] Frank J. Triska, John H. Duff, and Ronald J. Avanzino. Influence of exchange flow between the channel and hyporheic zone on nitrate production in a small mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2099–2111, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-235.

Lockhart:1990:DDP

[236] F. D. Lockhart, W. J. Lindberg, N. J. Blake, R. B. Erdman, H. M. Perry, and R. S. Waller. Distributional differences and population similarities for two deep-sea crabs (family Geryonidae) in the Northeastern Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2112– 2122, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-236.

Carrasco:1990:APM

[237] Kenneth R. Carrasco, Karen L. Tilbury, and Mark S. Myers. Assessment of the piscine micronucleus test as an in situ biological indicator of chemical contaminant effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11): 2123–2136, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-237.

Keller:1990:FIF

[238] A. E. Keller and T. L. Crisman. Factors influencing fish assemblages and species richness in subtropical Florida lakes and a comparison with temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2137– 2146, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-238.

Corkum:1990:IDP

[239] Lynda D. Corkum. Intrabiome distributional patterns of lotic macroinvertebrate assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):

2147–2157, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-239.

jRead:1990:CGR

[240] Andrew j. Read and David E. Gaskin. Changes in growth and reproduction of harbour porpoises, Phocoena phocoena, from the Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2158–2163, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-240.

Fechhelm:1990:EWR

[241] Robert G. Fechhelm and William B. Griffiths. Effect of wind on the recruitment of Canadian Arctic cisco (Coregonus autumnalis) into the Central Alaskan Beaufort Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47 (11):2164–2171, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f90-241.

Taylor:1990:VAB

[242] Eric B. Taylor. Variability in agonistic behaviour and salinity tolerance between and within two populations of juvenile Chinook salmon, Oncorhynchus tshawytscha, with contrasting life histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2172–2180, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-242.

Holtby:1990:ISS

[243] L. Blair Holtby, Bruce C. Andersen, and Ronald K. Kadowaki. Importance of smolt size and early ocean growth to interannual variability in marine survival of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2181–2194, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-243.

Smith:1990:VPH

[244] Barry D. Smith, Gordon A. McFarlane, and Mark W. Saunders. Variation in Pacific hake (Merluccius productus) summer length-at-age near Southern Vancouver Island and its relationship to fishing and oceanography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2195–2211, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-244.

Damours:1990:GJG

[245] Denis D’amours, Jean G. Landry, and Timothy C. Lambert. Growth of juvenile (0-group) Atlantic mackerel (Scombrus) in the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2212–2218, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-245.

Campana:1990:HRG

[246] Steven E. Campana. How reliable are growth back-calculations based on otoliths? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2219– 2227, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-246.

McGeachy:1990:ETC

[247] S. M. McGeachy and D. G. Dixon. Effect of temperature on the chronic toxicity of arsenate to rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2228–2234, November 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-247.

Millar:1990:CME

[248] R. B. Millar. Comparison of methods for estimating mixed stock fishery composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2235– 2241, November 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-248.

Ennis:1990:CCS

[249] G. P. Ennis, R. G. Hooper, and D. M. Taylor. Changes in the composition of snow crab (Chionoecetes opilio) participating in the annual breeding migration in Bonne Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2242–2249, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-249.

Wood:1990:GDE

[250] Chris C. Wood and Chris J. Foote. Genetic differences in the early development and growth of sympatric sockeye salmon and Kokanee (Oncorhynchus nerka), and their hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2250–2260, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-250.

Hall:1990:RIS

[251] Ronald J. Hall. Relative importance of seasonal, short-term pH disturbances during discharge variation on a stream ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2261–2274, November 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-251.

Anonymous:1990:LCLb

[252] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(11):2275, November 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-252.

Lachance:1990:PDH

[253] Stephanie Lachance and Pierre Magnan. Performance of domestic, hybrid, and wild strains of brook trout, Salvelinus fontinalis, after stocking: the impact of intra- and interspecific competition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2278–2284, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-253.

Lachance:1990:CEB

[254] Stephanie Lachance and Pierre Magnan. Comparative ecology and behavior of domestic, hybrid, and wild strains of brook trout, Salvelinus fontinalis, after stocking. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):

2285–2292, December 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-254.

Smith:1990:EIR

[255] J. W. Smith, A. E. Elarifi, R. Wootten, A. W. Pike, and M. D. B. Burt. Experimental infection of rainbow trout, Oncorhynchus mykiss, with Contracaecum osculatum (Rudolphi, 1802) and Pseudoterranova decipiens (Krabbe, 1878) (Nematoda; Ascaridoidea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2293–2296, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-255.

Dolloff:1990:MPA

[256] C. Andrew Dolloff and Cordon H. Reeves. Microhabitat partitioning among stream-dwelling juvenile Coho salmon, Oncorhynchus kisutch, and Dolly Varden, Salvelinus malma. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2297–2306, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-256.

Hill:1990:PRH

[257] Walter R. Hill and Bret C. Harvey. Periphyton responses to higher trophic levels and light in a shaded stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2307–2314, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-257.

Quinn:1990:MCA

[258] Terrance J. Quinn II, Richard B. Deriso, and Philip R. Neal. Migratory catch-age analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12): 2315–2327, December 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-258.

Dodds:1990:CMA

[259] Walter K. Dodds and John C. Priscu. A comparison of methods for assessment of nutrient deficiency of phytoplankton in a large oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2328–2338, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-259.

Haigh:1990:DPH

[260] R. Haigh and F. J. R. Taylor. Distribution of potentially harmful phytoplankton species in the Northern Strait of Georgia, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2339–2350, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-260.

Mulholland:1990:MPU

[261] Patrick J. Mulholland, Alan D. Steinman, and Jerry W. Elwood. Measurement of phosphorus uptake length in streams: Comparison of radiotracer and stable PO_{4} releases. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2351–2357, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f90-261.

Meador:1990:PRL

[262] M. R. Meador and W. E. Kelso. Physiological responses of largemouth bass, Micropterus salmoides, exposed to salinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2358–2363, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-262.

Kimura:1990:AAS

[263] Daniel K. Kimura. Approaches to age-structured separable sequential population analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12): 2364–2374, December 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-263.

Ware:1990:MFT

[264] D. M. Ware and R. W. Tanasichuk. A method for forecasting the timing of ripeness in Pacific herring (Clupea harengus pallasi) females. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2375–2379, December 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-264.

Garnick:1990:IFS

[265] E. Garnick and L. Margolis. Influence of four species of helminth parasites on orientation of seaward migrating sockeye salmon (Oncorhynchus nerka) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2380– 2389, December 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-265.

Hay:1990:TIS

[266] D. E. Hay. Tidal influence on spawning time of Pacific herring (Clupea harengus pallasi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12): 2390–2401, December 1990. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f90-266.

Waddy:1990:IIA

[267] S. L. Waddy and D. E. Aiken. Intermolt insemination, an alternative mating strategy for the American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2402–2406, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-267.

Longhurst:1990:AOS

[268] Alan R. Longhurst and Warren S. Wooster. Abundance of oil sardine (Sardinella longiceps) and upwelling on the Southwest Coast of India. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2407–2419, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-268.

Uthe:1990:SME

[269] John F. Uthe. Studies on mitigation of the effects of acidic precipitation on adult Atlantic salmon (Salmo salar) — introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2420–2421, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-269.

Sangalang:1990:EDL

[270] G. B. Sangalang, H. C. Freeman, J. F. Uthe, and L. S. Sperry. Effects of diet or liming on steroid hormone metabolism and reproduction in Atlantic salmon (Salmo salar) held in an acidic river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2422–2430, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f90-270.

Brown:1990:RPE

[271] S. B. Brown, R. E. Evans, H. S. Majewski, G. B. Sangalang, and J. F. Klaverkamp. Responses of plasma electrolytes, thyroid hormones, and gill histology in Atlantic salmon (Salmo salar) to acid and limed river waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2431–2440, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f90-271.

Majewski:1990:RKL

[272] H. S. Majewski, S. B. Brown, R. E. Evans, H. C. Freeman, and J. F. Klaverkamp. Responses of kidney, liver, muscle, and bone in Atlantic salmon (Salmo salar) to diet and liming in acidic Nova Scotia rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2441–2450, December 1990. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f90-272.

Jagoe:1990:MEL

[273] Charles H. Jagoe and Terry A. Haines. Morphometric effects of low pH and limed water on the gills of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2451–2460, December 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-273.

Anonymous:1990:LCLc

[274] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 47(12):2461–2467, December 1990. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f90-274.

Reddin:1991:PAS

[275] D. G. Reddin and P. B. Short. Postsmolt Atlantic salmon (Salmo salar) in the Labrador Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (1):2–6, January 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-001.

Hoogenboezem:1991:NMP

[276] Wim Hoogenboezem, Jos G. M. van den Boogaart, Ferdinand A. Sibbing, Eddy H. R. R. Lammens, Arie Terlouw, and Jan W. M. Osse. A new model of particle retention and branchial sieve adjustment in filterfeeding bream (Abramis brama, cyprinidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):7–18, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-002.

Titus:1991:SGP

[277] Robert G. Titus and Henrik Mosegaard. Selection for growth potential among migratory brown trout (Salmo trutta) fry competing for territories: Evidence from otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (1):19–27, January 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-003.

Karaas:1991:EPM

[278] Peter Karås, Erik Neuman, and Olof Sandström. Effects of a pulp mill effluent on the population dynamics of perch, Perca fluviatilis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):28–34, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-004.

Fletcher:1991:TRB

[279] W. J. Fletcher. A test of the relationship between otolith weight and age for the pilchard Sardinops neopilchardus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):35–38, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-005.

Hare:1991:DCL

[280] Landis Hare, Erwan Saouter, Peter G. C. Campbell, André Tessier, Francis Ribeyre, and Alain Boudou. Dynamics of cadmium, lead, and zinc exchange between nymphs of the burrowing mayfly Hexagenia rigida (Ephemeroptera) and the environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):39–47, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-006.

Carr:1991:DID

[281] Steven M. Carr and H. Dawn Marshall. Detection of intraspecific DNA sequence variation in the mitochondrial cytochrome b gene of Atlantic cod (Gadus morhua) by the polymerase chain reaction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):48–52, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-007.

Kaya:1991:RDB

[282] Calvin M. Kaya. Rheotactic differentiation between fluvial and lacustrine populations of Arctic grayling (Thymallus arcticus), and implications for the only remaining indigenous population of fluvial “Montana Grayling”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):53–59, January 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-008.

Meffe:1991:LHC

[283] Gary K. Meffe. Life history changes in Eastern mosquitofish (Gambusia holbrooki) induced by thermal elevation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):60–66, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-009.

Levy:1991:AAD

[284] David A. Levy. Acoustic analysis of diel vertical migration behavior of Mysis relicta and Kokanee (Oncorhynchus nerka) within Okanagan Lake, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (1):67–72, January 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-010.

Bentzen:1991:EMM

[285] E. Bentzen and W. D. Taylor. Estimating Michaelis–Menten parameters and lake water phosphate by the rigler bioassay: Importance of fitting technique, plankton size, and substrate range. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):73–83, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-011.

Klotz:1991:TRB

[286] R. L. Klotz. Temporal relation between soluble reactive phosphorus and factors in stream water and sediments in Hoxie Gorge Creek, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):84–90, January 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-012.

Hinckley:1991:TDA

[287] Sarah Hinckley, Kevin M. Bailey, Susan J. Picquelle, James D. Schumacher, and Phyllis J. Stabeno. Transport, distribution, and abundance of larval and juvenile walleye pollock (Theragra chalcogramma) in the Western Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (1):91–98, January 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-013.

Hedrick:1991:SRV

[288] R. P. Hedrick, S. Yun, and W. H. Wingfield. A small RNA virus isolated from salmonid fishes in California, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):99–104, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-014.

Sprules:1991:BSS

[289] W. Gary Sprules, S. B. Brandt, D. J. Stewart, M. Munawar, E. H. Jin, and J. Love. Biomass size spectrum of the Lake Michigan pelagic food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):105–115, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-015.

Sager:1991:FIP

[290] Paul E. Sager and Sumner Richman. Functional interaction of phytoplankton and zooplankton along the trophic gradient in Green Bay, Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):116–122, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
016.

Dionne:1991:EAM

[291] Michele Dionne and Carol L. Folt. An experimental analysis of macrophyte growth forms as fish foraging habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):123–131, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-017.

Wren:1991:RBM

[292] C. D. Wren, W. A. Scheider, D. L. Wales, B. W. Muncaster, and I. M. Gray. Relation between mercury concentrations in walleye (Stizostedion vitreum vitreum) and northern pike (Esox lucius) in Ontario lakes and influence of environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):132–139, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-018.

Molot:1991:NPR

[293] Lewis A. Molot and P. J. Dillon. Nitrogen/phosphorus ratios and the prediction of chlorophyll in phosphorus-limited lakes in Central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):140–145, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-019.

Pace:1991:VZS

[294] Michael L. Pace, Stuart E. G. Findlay, and David Lints. Variance in zooplankton samples: Evaluation of a predictive model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):146–151, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-020.

Kenney:1991:UIC

[295] Bernard C. Kenney. Under-ice circulation and the residence time of a shallow bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(1):152–162, January 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
021.

Duston:1991:EIF

[296] J. Duston, R. L. Saunders, and D. E. Knox. Effects of increases in freshwater temperature on loss of smolt characteristics in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):164–169, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-022.

Garcia-Vazquez:1991:CPP

[297] E. García-Vázquez, P. Morán, and A. M. Pendás. Chromosome polymorphism patterns indicate failure of a Scottish stock of Salmo salar transplanted into a Spanish river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (2):170–172, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-023.

Wangberg:1991:LTS

[298] Sten-Åke Wängberg, Ulf Heyman, and Hans Blanck. Long-term and short-term arsenate toxicity to freshwater phytoplankton and periphyton in limnocorrals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2): 173–182, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-024.

Diana:1991:ENT

[299] James S. Diana, Daniel J. Dettweiler, and C. Kwei Lin. Effect of Nile Tilapia (Oreochromis niloticus) on the ecosystem of aquaculture ponds, and its significance to the Trophic Cascade Hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):183–190, February 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-025.

Minshall:1991:DRT

[300] G. Wayne Minshall, Evelyn Hitchcock, and James R. Barnes. Decomposition of rainbow trout (Oncorhynchus mykiss) carcasses in a forest stream ecosystem inhabited only by nonanadromous fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):191–195, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-026.

Lamberti:1991:SER

[301] Gary A. Lamberti, Stanley V. Gregory, Linda R. Ashkenas, Randall C. Wildman, and Kelly M. S. Moore. Stream ecosystem recovery following a catastrophic debris flow. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (2):196–208, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-027.

Charmantier:1991:STO

[302] G. Charmantier, M. Charmantier-Daures, S. L. Waddy, and D. E. Aiken. Salinity tolerance and osmoregulation in the Nemertean Pseudocarcinonemertes homari, an egg predator of American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):209–214, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-028.

MacDonald:1991:GRE

[303] B. A. MacDonald, R. J. Thompson, and N. F. Bourne. Growth and reproductive energetics of three scallop species from British Columbia (Chlamys hastata, Chlamys rubida, and Crassadoma gigantea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):215–221, February 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-029.

Moore:1991:DEM

[304] Johnnie N. Moore, Samuel N. Luoma, and Donald Peters. Downstream effects of mine effluent on an intermontane riparian system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):222–232, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-030.

Ostrofsky:1991:SVD

[305] M. L. Ostrofsky and G. G. McGee. Spatial variation in the distribution of phosphorus species in the surficial sediments of Canadohta Lake, Pennsylvania; implications for internal phosphorus loading estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):233–237, February 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-031.

Hontela:1991:AVO

[306] A. Hontela, J. B. Rasmussen, D. Ko, K. Lederis, and G. Chevalier. Arginine vasotocin, an osmoregulatory hormone, as a potential indicator of acid stress in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2): 238–242, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-032.

Johnsson:1991:IDS

[307] Jörgen I. Johnsson and Mark V. Abrahams. Interbreeding with domestic strain increases foraging under threat of predation in juvenile steelhead trout (Oncorhynchus mykiss): an experimental study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):243–247, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-033.

Solazzi:1991:SCR

[308] Mario F. Solazzi, Thomas E. Nickelson, and Steven L. Johnson. Survival, contribution, and return of hatchery Coho salmon (Oncorhynchus kisutch) released into freshwater, estuarine, and marine environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):248–253, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-034.

Dorazio:1991:EMR

[309] Robert M. Dorazio and Paul J. Rago. Evaluation of a mark–recapture method for estimating mortality and migration rates of stratified populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):254–260, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-035.

Hebert:1991:HPC

[310] Craig E. Hebert and G. D. Haffner. Habitat partitioning and contaminant exposure in cyprinids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (2):261–266, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-036.

Hines:1991:FRO

[311] Anson H. Hines. Fecundity and reproductive output in nine species of Cancer crabs (Crustacea, Brachyura, Cancridae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):267–275, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-037.

Richard:1991:ADN

[312] Pierre R. Richard. Abundance and distribution of narwhals (Monodon monoceros) in Northern Hudson Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):276–283, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-038.

Klump:1991:MRA

[313] J. Val Klump, Jerry L. Kaster, and Michael E. Sierszen. Mysis relicta assimilation of hexachlorobiphenyl from sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):284–289, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-039.

Patrick:1991:RDT

[314] P. H. Patrick, B. Sim, and G. Hunt. Range detection of targets using hydroacoustics in the laboratory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):290–295, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-040.

Castonguay:1991:UFA

[315] Martin Castonguay, Patrick Simard, and Pierre Gagnon. Usefulness of Fourier analysis of otolith shape for Atlantic mackerel (Scomber scombrus) stock discrimination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (2):296–302, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-041.

Anderson:1991:CCR

[316] John T. Anderson and William G. Warren. Comparison of catch rates among small and large bongo sampler for Calanus finmarchicus Copepodite stages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):303–308, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-042.

Bartlett:1991:ITT

[317] Sylvia E. Bartlett and William S. Davidson. Identification of Thunnus tuna species by the polymerase chain reaction and direct sequence analysis of their mitochondrial cytochrome b genes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):309–317, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-043.

Barber:1991:MBO

[318] M. Craig Barber, Luis A. Suárez, and Ray R. Lassiter. Modelling bioaccumulation of organic pollutants in fish with an application to PCBs in Lake Ontario salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2): 318–337, February 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-044.

Anonymous:1991:CRPa

[319] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(2):338, February 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-045.

Cochrane:1991:MFA

[320] N. A. Cochrane, D. Sameoto, A. W. Herman, and J. Neilson. Multiplefrequency acoustic backscattering and zooplankton aggregations in the Inner Scotian Shelf basins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (3):340–355, March 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-046.

Lafontaine:1991:CTL

[321] Nathalie Lafontaine and Donald J. McQueen. Contrasting trophic level interactions in Lake St. George and Haynes Lake (Ontario, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):356–363, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-047.

Chow-Fraser:1991:FGC

[322] Patricia Chow-Fraser and Edward J. Maly. Factors governing clutch size in two species of Diaptomus (Copepoda: Calanoida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):364–370, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-048.

Macdonald:1991:CDP

[323] C. R. Macdonald and C. D. Metcalfe. Concentration and distribution of PCB congeners in isolated Ontario lakes contaminated by atmospheric deposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (3):371–381, March 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-049.

Whyte:1991:CIE

[324] J. N. C. Whyte and J. A. Boutillier. Concentrations of inorganic elements and fatty acids in geographic populations of the spot prawn Pandalus platyceros. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):382–390, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-050.

Bacher:1991:MEU

[325] C. Bacher, M. Héral, J. M. Deslous-Paoli, and D. Razet. Modèle énergétique uniboite de la croissance des huîtres (Crassostrea gigas) dans le bassin de Marennes–Oléron. (French) [Single-sided energy model of the growth of oysters (Crassostrea gigas) in the Marennes–Oléron Basin]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):391–404, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-051.

Rowe:1991:RFS

[326] D. K. Rowe, J. E. Thorpe, and A. M. Shanks. Role of fat stores in the maturation of male Atlantic salmon (Salmo salar) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):405–413, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-052.

Witters:1991:ARP

[327] H. E. Witters, S. Van Puymbroeck, and O. L. J. Vanderborght. Adrenergic response to physiological disturbances in rainbow troutf Oncorhynchus mykiss, exposed to aluminum at acid pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):414–420, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-053.

Bandin:1991:LBA

[328] I. Bandin, Y. Santos, D. W. Bruno, R. S. Raynard, A. E. Toranzo, and J. L. Barja. Lack of biological activities in the extracellular products of Renibacterium salmoninarum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):421–425, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-054.

Gausen:1991:LSE

[329] D. Gausen and V. Moen. Large-scale escapes of farmed Atlantic salmon (Salmo salar) into Norwegian rivers threaten natural populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):426–428, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-055.

Lura:1991:MSO

[330] Harald Lura and Harald Saegrov. A method of separating offspring from farmed and wild Atlantic salmon (Salmo salar) based on different ratios of optical isomers of astaxanthin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):429–433, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-056.

Fox:1991:FCB

[331] Michael G. Fox. Food consumption and bioenergetics of young-of-the-year walleye (Stizostedion vitreum vitreum): Model predictions and population density effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (3):434–441, March 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-057.

Irons:1991:EMB

[332] John G. Irons III, John P. Bryant, and Mark W. Oswood. Effects of moose browsing on decomposition rates of birch leaf litter in a subarctic stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):442–444, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-058.

Willis:1991:EEP

[333] Douglas E. Willis, Allison J. Edwards, and Richard F. Addison. Effects of environmental pH on the hepatic mixed function oxidases in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (3):445–447, March 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-059.

Campbell:1991:HRM

[334] Daniel E. Campbell and Joseph J. Graham. Herring recruitment in Maine coastal waters: an ecological model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):448–471, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-060.

Hurley:1991:PPL

[335] James P. Hurley and David E. Armstrong. Pigment preservation in lake sediments: a comparison of sedimentary environments in Trout Lake, Wisconsin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):472–486, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-061.

Helser:1991:NME

[336] Thomas E. Helser, Richard E. Condrey, and James P. Geaghan. A new method of estimating gillnet selectivity, with an example for spotted seatrout, Cynocion nehulosus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):487–492, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-062.

Servizi:1991:ETS

[337] James A. Servizi and Dennis W. Martens. Effect of temperature, season, and fish size on acute lethality of suspended sediments to Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (3):493–497, March 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-063.

Merrick:1991:LTS

[338] Glenn W. Merrick, Anne E. Hershey, and Michael E. McDonald. Lake trout (Salvelinus namaycush) control of snail density and size distribution in an Arctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (3):498–502, March 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-064.

Pepin:1991:ETS

[339] Pierre Pepin. Effect of temperature and size on development, mortality, and survival rates of the pelagic early life history stages of marine fish.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):503–518, March 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-065.

Anonymous:1991:LCLa

[340] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(3):519–522, March 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-066.

Welch:1991:OHR

[341] David W. Welch and Donald J. Noakes. Optimal harvest rate policies for rebuilding the Adams River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):526–535, April 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-067.

Bergman:1991:CAT

[342] Eva Bergman. Changes in abundance of two percids, Perca fluviatilis and Gymnocephalus cernuus, along a productivity gradient: Relations to feeding strategies and competitive abilities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):536–545, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-068.

Appelberg:1991:DSF

[343] Magnus Appelberg and Erik Degerman. Development and stability of fish assemblages after lime treatment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):546–554, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-069.

Bradford:1991:EAE

[344] Michael J. Bradford. Effects of ageing errors on recruitment time series estimated from sequential population analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):555–558, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-070.

Kuris:1991:IBS

[345] Armand M. Kuris, S. Forrest Blau, A. J. Paul, Jeffrey D. Shields, and Daniel E. Wickham. Infestation by brood symbionts and their impact on egg mortality of the red king crab, Paralithodes camtschatica, in Alaska: Geographic and temporal variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):559–568, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-071.

Bowen:1991:DDM

[346] Stephen H. Bowen, Donna J. D’angelo, Stephen H. Arnold, Michael J. Keniry, and Ronald J. Albrecht. Density-dependent maturation, growth, and female dominance in Lake Superior lake herring (Coregonus artedii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):569–576, April 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-072.

Birt:1991:MDV

[347] Tim P. Birt, John M. Green, and William S. Davidson. Mitochondrial DNA variation reveals genetically distinct sympatric populations of anadromous and nonanadromous Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):577–582, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-073.

Webb:1991:CMR

[348] Paul W. Webb. Composition and mechanics of routine swimming of rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):583–590, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-074.

Fournier:1991:ALF

[349] David A. Fournier, John R. Sibert, and Mark Terceiro. Analysis of length frequency samples with relative abundance data for the Gulf of Maine northern shrimp (Pandalus borealis) by the MULTIFAN method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):591–598, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-075.

Jones:1991:APE

[350] Michael L. Jones, Charles K. Minns, David R. Marmorek, and Kim J. Heltcher. Assessing the potential extent of damage to inland lakes in Eastern Canada due to acidic deposition. IV. Uncertainty analysis of a regional model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):599–606, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
076.

Parrish:1991:PSA

[351] Donna L. Parrish and F. Joseph Margraf. Prey selectivity by age 0 white perch (Morone americana) and yellow perch (Perca flavescens) in laboratory experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (4):607–610, April 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-077.

Singer:1991:MOP

[352] Thomas D. Singer and James S. Ballantyne. Metabolic organization of a primitive fish, the bowfin (Amia calva). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):611–618, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-078.

Bajdik:1991:MFY

[353] Chris D. Bajdik and David C. Schneider. Models of the fish yield from lakes: Does the random component matter? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):619–622, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-079.

Hsieh:1991:UNM

[354] William W. Hsieh, Warren G. Lee, and Lawrence A. Mysak. Using a numerical model of the Northeast Pacific Ocean to study the interannual variability of the Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):623–630, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-080.

Testa:1991:TVA

[355] J. Ward Testa, Gary Oehlert, David G. Ainley, John L. Bengtson, Donald B. Siniff, Richard M. Laws, and David Rounsevell. Temporal variability in Antarctic marine ecosystems: Periodic fluctuations in the phocid seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):631–639, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-081.

Rubec:1991:EMS

[356] Peter J. Rubec, Jacquie M. McGlade, Bertin L. Trottier, and André Ferron. Evaluation of methods for separation of Gulf of St. Lawrence beaked redfishes, Sebastes fasciatus and S. mentella: Malate dehydrogenase mobility patterns compared with extrinsic gasbladder muscle passages and anal fin ray counts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (4):640–660, April 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-082.

Tallman:1991:PDS

[357] Ross F. Tallman and Michael C. Healey. Phenotypic differentiation in seasonal ecotypes of chum salmon, Oncorhynchus keta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):661–671, April 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-083.

Johnson:1991:BEA

[358] Timothy B. Johnson and David O. Evans. Behaviour, energetics, and associated mortality of young-of-the-year white perch (Morone americana) and yellow perch (Perca flavescens) under simulated winter conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):672–680, April 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-084.

Rakocinski:1991:PSR

[359] Chet Rakocinski. Prey-size relationships and feeding tactics of primitive stream-dwelling darters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (4):681–693, April 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-085.

Brooks:1991:VSD

[360] Albert L. Brooks and R. Alfred Saenger. Vertical size–depth distribution properties of midwater fish off Bermuda, with comparative reviews for other open ocean areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (4):694–721, April 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-086.

Gauldie:1991:TSG

[361] R. W. Gauldie. Taking stock of genetic concepts in fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(4):722–731, April 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-087.

Clark:1991:GER

[362] William G. Clark. Groundfish exploitation rates based on life history parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):734–750, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-088.

Sainte-Marie:1991:WBU

[363] Bernard Sainte-Marie. Whelk (Buccinum undatum) movement and its implications for the use of tag–recapture methods for the determination of baited trap fishing parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):751–756, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-089.

Heggenes:1991:SSC

[364] Jan Heggenes, Thomas G. Northcote, and Armin Peter. Spatial stability of cutthroat trout (Oncorhynchus clarki) in a small, coastal stream.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):757–762, May 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-090.

Agusti:1991:ASL

[365] Susana Agustí. Allometric scaling of light absorption and scattering by phytoplankton cells. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (5):763–767, May 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-091.

Taylor:1991:PDF

[366] Barbara E. Taylor, John M. Aho, Diane L. Mahoney, and Ruth A. Estes. Population dynamics and food habits of bluegill (Lepomis macrochirus) in a thermally stressed reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):768–775, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f91-092.

Richardson:1991:IED

[367] John S. Richardson and William E. Neill. Indirect effects of detritus manipulations in a montane stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):776–783, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-093.

Boutillier:1991:EEM

[368] J. A. Boutillier and N. A. Sloan. Evaluation of escape modifications in traps for prawn, Pandalus platyceros. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):784–791, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-094.

Wilson:1991:ESS

[369] Raymond R. Wilson, Jr. and Peter D. Alberdi, Jr. An electrophoretic study of Spanish sardine suggests a single predominant species in the Eastern Gulf of Mexico, Sardinella aurita. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):792–798, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-095.

Pascual:1991:EAM

[370] Miguel A. Pascual and Thomas P. Quinn. Evaluation of alternative models of the coastal migration of adult Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (5):799–810, May 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-096.

Benke:1991:PCP

[371] Arthur C. Benke, Keith A. Parsons, and Sunita M. Dhar. Population and community patterns of invertebrate drift in an unregulated coastal plain river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):811–823, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-097.

Waiwood:1991:FAC

[372] K. G. Waiwood, S. J. Smith, and M. R. Petersen. Feeding of Atlantic cod (Gadus morhua) at low temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):824–831, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-098.

Halliday:1991:MDS

[373] R. C. Halliday. Marine distribution of the sea lamprey (Petromyzon marinus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):832–842, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-099.

Rose:1991:EBR

[374] G. A. Rose and W. C. Leggett. Effects of biomass–range interactions on catchability of migratory demersal fish by mobile fisheries: an example of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):843–848, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-100.

Brown:1991:CHM

[375] Larry R. Brown and Peter B. Moyle. Changes in habitat and microhabitat partitioning within an assemblage of stream fishes in response to predation by Sacramento squawfish (Ptychocheilus grandis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):849–856, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-101.

Tremblay:1991:IBT

[376] Serge Tremblay and Pierre Magnan. Interactions between two distantly related species, brook trout (Salvelinus fontinalis) and white sucker (Catostomus commersoni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (5):857–867, May 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-102.

Tufts:1991:EEW

[377] B. L. Tufts, Y. Tang, K. Tufts, and R. G. Boutilier. Exhaustive exercise in “Wild” Atlantic salmon (Salmo saiar): Acid–base regulation and blood gas transport. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (5):868–874, May 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-103.

Walker:1991:TPD

[378] Mary K. Walker, Jan M. Spitsbergen, James R. Olson, and Richard E. Peterson. 2,3,7,8-tetrachlorodibenzo-p-dioxin (TCDD) toxicity during early life stage development of lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):875–883, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-104.

Bermingham:1991:DBA

[379] Eldredge Bermingham, Stephen H. Forbes, Kevin Friedland, and Carles Pla. Discrimination between Atlantic salmon (Salmo salar) of North American and European origin using restriction analyses of mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):884–893, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-105.

Brandt:1991:AMA

[380] Stephen B. Brandt, Doran M. Mason, E. Vincent Patrick, Ray L. Argyle, L. Wells, Philip A. Unger, and Donald J. Stewart. Acoustic measures of the abundance and size of pelagic planktivores in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):894–908, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-106.

Stewart:1991:PPS

[381] Donald J. Stewart and Myriam Ibarra. Predation and production by salmonine fishes in Lake Michigan, 1978–88. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):909–922, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-107.

Hicks:1991:BDA

[382] Randall E. Hicks and Christopher J. Owen. Bacterioplankton density and activity in benthic nepheloid layers of Lake Michigan and Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):923–932, May 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-108.

Rogers:1991:SMP

[383] Jean Beyer Rogers and Craig C. Burley. A sigmoid model to predict gastric evacuation rates of smallmouth bass (Micropterus dolomieui) fed juvenile salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):933–937, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
109.

Keefe:1991:CAH

[384] MaryLouise Keefe and Howard E. Winn. Chemosensory attraction to home stream water and conspecifics by native brook trout, Salvelinus fontinalis, from two Southern New England streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):938–944, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-110.

Hindar:1991:GEC

[385] Kjetil Hindar, Nils Ryman, and Fred Utter. Genetic effects of cultured fish on natural fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(5):945–957, May 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-111.

Forbes:1991:DNP

[386] J. R. Forbes and K. L. Denman. Distribution of Nitzschia pungens in coastal waters of British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):960–967, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-112.

Genoni:1991:MSC

[387] Giulio P. Genoni and Claudia Pahl-Wostl. Measurement of scope for change in ascendency for short-term assessment of community stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):968–974, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-113.

Walker:1991:ACQ

[388] Ian R. Walker, John P. Smol, Daniel R. Engstrom, and H. J. B. Birks. An assessment of chironomidae as quantitative indicators of past climatic change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):975–987, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-114.

Henderson:1991:ESS

[389] M. A. Henderson and A. J. Cass. Effect of smolt size on smolt-to-adult survival for Chilko Lake sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):988–994, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-115.

Eckmann:1991:HSP

[390] Reiner Eckmann. A hydroacoustic study of the pelagic spawning behavior of whitefish (Coregonus lavaretus) in Lake Constance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):995–1002, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-116.

Cohen:1991:LSV

[391] Edward B. Cohen, David G. Mountain, and Robert O’Boyle. Localscale versus large-scale factors affecting recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1003–1006, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-117.

Baker:1991:PAE

[392] Timothy T. Baker and L. Saree Timmons. Precision of ages estimated from five bony structures of Arctic char (Salvelinus alpinus) from the Wood River System, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (6):1007–1014, June 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-118.

LAbee-Lund:1991:VWB

[393] Jan Henning L’Abée-Lund. Variation within and between rivers in adult size and sea age at maturity of anadromous brown trout, Salmo trutta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1015–1021, June 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-119.

Findlay:1991:RPC

[394] D. L. Findlay and S. E. M. Kasian. Response of a phytoplankton community to controlled partial recovery from experimental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1022–1029, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-120.

Bukaveckas:1991:EWL

[395] Paul A. Bukaveckas and Charles T. Driscoll. Effects of whole-lake base addition on the optical properties of three clearwater acidic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1030–1040, June 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-121.

Berglund:1991:EEW

[396] I. Berglund, L. P. Hansen, H. Lundqvist, B. Jonsson, T. Eriksson, J. E. Thorpe, and L.-O. Eriksson. Effects of elevated winter temperature on seawater adaptability, sexual rematuration, and downstream migratory behaviour in mature male Atlantic salmon parr (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1041–1047, June 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-122.

Bedard:1991:HCD

[397] Charles Bédard and Roger Knowles. Hypolimnetic O_{2} consumption, denitrification, and methanogenesis in a thermally stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1048–1054, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-123.

Borgmann:1991:RBC

[398] U. Borgmann, W. P. Norwood, and I. M. Babirad. Relationship between chronic toxicity and bioaccumulation of cadmium in Hyalella azteca. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1055–1060, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-124.

Reid:1991:MBA

[399] Scott D. Reid and D. G. McDonald. Metal binding activity of the gills of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1061–1068, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-125.

Peckarsky:1991:HSS

[400] Barbara L. Peckarsky. Habitat selection by stream-dwelling predatory stoneflies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1069–1076, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-126.

Maly:1991:MMC

[401] Edward J. Maly. Mating and multiple clutches in Australian Boeckella and Calamoecia (Copepoda: Calanoida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1077–1080, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-127.

Hilborn:1991:MSF

[402] Ray Hilborn. Modeling the stability of fish schools: Exchange of individual fish between schools of skipjack tuna (Katsuwonus pelamis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1081–1091, June 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-128.

Klut:1991:PAU

[403] M. E. Klut and J. G. Stockner. Picoplankton associations in an ultraoligotrophic lake on Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1092–1099, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-129.

Hayward:1991:BAU

[404] Robert S. Hayward. Bias associated with using the eggers model for estimating fish daily ration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (6):1100–1103, June 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-130.

Dalton:1991:VPN

[405] Timothy J. Dalton. Variation in prevalence of Nanophyetus salmincola, a parasite tag indicating U.S. northwest origin, in steelhead trout (Oncorhynchus mykiss) caught in the Central North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1104–1108, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-131.

deMarch:1991:HGP

[406] B. G. E. de March. Hatchery growth of pure strains and intraspecific hybrids of juvenile Arctic charr, Salvelinus alpinus (Canadian × Norwegian charr). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1109–1116, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-132.

Ballantyne:1991:EAG

[407] J. S. Ballantyne and J. A. Berges. Enzyme activities of gill, hepatopancreas, mantle, and adductor muscle of the oyster (Crassostrea virginica) after changes in diet and salinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1117–1123, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-133.

Davies:1991:IID

[408] Ronald W. Davies and T. E. Gates. Intra- and interspecific differences in the response of two lentic species of leeches to seasonal hyperoxia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1124–1127, June 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-134.

Anonymous:1991:LCLb

[409] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(6):1128–1129, June 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-135.

Anonymous:1991:SWP

[410] Anonymous. The Stockholm Water Prize and the Experimental Lakes Area guest editorial: Le Prix d’hydrologie de Stockholm et la région des lacs expérimentaux collaboration spéciale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1132–1135, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-136.

Bates:1991:CDA

[411] S. S. Bates, A. S. W. de Freitas, J. E. Milley, R. Pocklington, M. A. Quilliam, J. C. Smith, and J. Worms. Controls on domoic acid production by the diatom Nitzschia pungens f. multiseries in culture: Nutrients and irradïance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1136–1144, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-137.

Beaver:1991:ILO

[412] John R. Beaver and Thomas L. Crisman. Importance of latitude and organic color on phytoplankton primary productivity in Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1145–1150, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-138.

Warwick:1991:IDL

[413] W. F. Warwick. Indexing deformities in ligulae and antennae of Procladius larvae (Diptera: Chironomidae): Application to contaminantstressed environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (7):1151–1166, July 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-139.

Lefebvre:1991:ODC

[414] Louise Lefebvre and Jean-Claude F. Brêthes. Orientation des déplacements de crabes des neiges mâles (Chionoecetes opilio) marqués dans le sud-ouest du golfe du Saint-Laurent. (French) [Orientation of movements of male snow crabs (Chionoecetes opilio) marked in the southwestern Gulf of St. Lawrence]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1167–1175, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-140.

Bardonnet:1991:IDV

[415] Agnès Bardonnet and Philippe Gaudin. Influence of daily variations of light and temperature on the emergence rhythm of grayling fry (Thymallus thymallus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1176– 1180, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-141.

Bohling:1991:VYC

[416] Paula Böhling, Richard Hudd, Hannu Lehtonen, Peter Karås, Erik Neuman, and Gunnar Thoresson. Variations in year-class strength of different perch (Perca fluviatilis) populations in the Baltic Sea with special reference to temperature and pollution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1181–1187, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-142.

Kaastrup:1991:PAI

[417] P. Kaastrup, V. Hørlyck, N. J. Olesen, N. Lorenzen, P. E. Vestergaard Jørgensen, and P. Berg. Paternal association of increased susceptibility to viral haemorrhagic septicaemia (VHS) in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1188–1191, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-143.

Wong:1991:ESSa

[418] Donald Wong and Luis Olivesra. Effects of selenite and selenate on the growth and motility of seven species of marine microalgae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1193–1200, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-144.

Wong:1991:ESSb

[419] Donald Wong and Luis Oliveira. Effects of selenite and selenate toxicity on the ultrastructure and physiology of three species of marine microalgae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1201–1211, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-145.

Kennedy:1991:TAX

[420] Christopher J. Kennedy, Kenneth A. Gill, and Patrick J. Walsh. Temperature acclimation of xenobiotic metabolizing enzymes in cultured hepatocytes and whole liver of the Gulf toadfish, Opsanus beta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1212–1219, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-146.

Berrill:1991:CCI

[421] Michael Berrill, Graeme Taylor, and Helene Savard. Are chloride cells involved in low pH tolerance and sensitivity? The mayfly possibility. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1220–1225, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-147.

Suttle:1991:SDA

[422] Curtis A. Suttle, William P. Cochlan, and John G. Stockner. Sizedependent ammonium and phosphate uptake, and N:P supply ratios in an oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (7):1226–1234, July 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-148.

Grossman:1991:ERD

[423] Gary D. Grossman and Vital Boulé. Effects of rosyside dace (Clinostomus funduloides) on microhabitat use of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1235–1243, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-149.

Petticrew:1991:CGS

[424] Ellen L. Petticrew and Jacob Kalff. Calibration of a gypsum source for freshwater flow measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1244–1249, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-150.

Beamish:1991:AFM

[425] Richard J. Beamish and Colin D. Levings. Abundance and freshwater migrations of the anadromous parasitic lamprey, Lampetra tridentate, in a tributary of the Fraser River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1250–1263, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-151.

Richards:1991:CFA

[426] Laura J. Richards, Robert Kieser, Timothy J. Mulligan, and John R. Candy. Classification of fish assemblages based on echo integration surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1264–1272, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-152.

Collie:1991:AMS

[427] Jeremy S. Collie and Carl J. Walters. Adaptive management of spatially replicated groundfish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1273–1284, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-153.

Grant:1991:ESS

[428] Jonathan Grant and Bruce Thorpe. Effects of suspended sediment on growth, respiration, and excretion of the soft-shell clam (Mya arenaria). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1285–1292, July 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-154.

Taylor:1991:PPS

[429] William D. Taylor and David R. S. Lean. Phosphorus pool sizes and fluxes in the epilinnnion of a mesotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1293–1301, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-155.

Lubin:1991:IPN

[430] Robert T. Lubin, A. W. Rourke, and Richard L. Saunders. Influence of photoperiod on the number and ultrastructure of gill chloride cells of the Atlantic salmon (Salmo salar) before and during smoltification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1302–1307, July 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-156.

Ferguson:1991:DPC

[431] Moira M. Ferguson and Peter E. Ihssen. Distribution and phenotypic correlates of variation at enzyme coding loci in rainbow trout (Oncorhynchus mykiss) from the Lower Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1308–1315, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-157.

Loder:1991:SBT

[432] John W. Loder. Summertime bottom temperatures on the Southeast Shoal of the Grand Bank, and implications for exchange rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1316–1325, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-158.

Pinkerton:1991:LBW

[433] Evelyn Pinkerton. Locally based water quality planning: Contributions to fish habitat protection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (7):1326–1333, July 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-159.

Anonymous:1991:LCLc

[434] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(7):1334–1344, July 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-160.

Frank:1991:PRV

[435] Kenneth T. Frank. Predicting recruitment variation from year class specific vertebral counts: an analysis of the potential and a plan for verification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1350–1357, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
161.

Jenkins:1991:DDL

[436] Greg P. Jenkins, Jock W. Young, and Tim L. O. Davis. Density dependence of larval growth of a marine fish, the southern bluefin tuna, Thunnus maccoyii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8): 1358–1363, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-162.

Heggenes:1991:SHS

[437] Jan Heggenes, Thomas G. Northcote, and Armin Peter. Seasonal habitat selection and preferences by cutthroat trout (Oncorhynchus clarki) in a small coastal stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8): 1364–1370, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-163.

Munkittrick:1991:IBK

[438] K. R. Munkittrick, C. B. Portt, G. J. Van Der Kraak, I. R. Smith, and D. A. Rokosh. Impact of bleached kraft mill effluent on population characteristics, liver MFO activity, and serum steroid levels of a Lake Superior white sucker (Catostomus commersoni) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1371–1380, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-164.

Griffiths:1991:DDZ

[439] Ronald W. Griffiths, Donald W. Schloesser, Joseph H. Leach, and William P. Kovalak. Distribution and dispersal of the zebra mussel (Dreissena polymorpha) in the Great Lakes region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1381–1388, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-165.

Strayer:1991:PDZ

[440] David L. Strayer. Projected distribution of the zebra mussel, Dreissena polymorpha, in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1389–1395, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-166.

White:1991:NRE

[441] Bruce A. White and James B. Shaklee. Need for replicated electrophoretic analyses in multiagency genetic stock identification (GSI) programs: Examples from a pink salmon (Oncorhynchus gorbuscha) GSI fisheries study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8): 1396–1407, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-167.

Danzmann:1991:AFT

[442] Roy G. Danzmann, Don S. MacLennan, Don G. Hector, Paul D. N. Hebert, and Jurek Kolasa. Acute and final temperature preferenda as predictors of Lake St. Clair fish catchability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1408–1418, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-168.

Matsakis:1991:AFM

[443] Sophie Matsakis and Robert J. Conover. Abundance and feeding of medusae and their potential impact as predators on other zooplankton in Bedford Basin (Nova Scotia, Canada) during spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1419–1430, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-169.

Restrepo:1991:AMC

[444] Victor R. Restrepo and Reg A. Watson. An approach to modeling crustacean egg-bearing fractions as a function of size and season. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1431–1436, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-170.

Hammar:1991:NHB

[445] Johan Hammar, J. Brian Dempson, and Eric Verspoor. Natural hybridization between Arctic char (Salvelinus alpinus) and brook trout (S. fontinalis): Evidence from Northern Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1437–1445, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-171.

Trippel:1991:CMU

[446] Edward A. Trippel and Harold H. Harvey. Comparison of methods used to estimate age and length of fishes at sexual maturity using populations of white sucker (Catostomus commersoni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1446–1459, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-172.

Klotz:1991:CPH

[447] R. L. Klotz. Cycling of phosphatase hydrolyzable phosphorus in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1460–1467, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-173.

Johnston:1991:PFM

[448] Thomas A. Johnston, R. A. Bodaly, and J. A. Mathias. Predicting fish mercury levels from physical characteristics of boreal reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1468–1475, August 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-174.

Harvey:1991:IAB

[449] Bret C. Harvey. Interaction of abiotic and biotic factors influences larval fish survival in an Oklahoma stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1476–1480, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-175.

Hare:1991:TED

[450] Landis Hare, André Tessier, and Peter G. C. Campbell. Trace element distributions in aquatic insects: Variations among genera, elements, and lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1481–1491, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
176.

Madenjian:1991:LGY

[451] Charles P. Madenjian. Limits to growth of young-of-the-year walleye (Stizostedion vitreum vitreum): an individual-based model perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1492–1499, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-177.

DePinto:1991:SLO

[452] Joseph V. DePinto. State of the Lake Ontario ecosystem: Introduction to an ecosystem perspective on a vital resource. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1500–1502, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-178.

Martini:1991:GLO

[453] I. Peter Martini and John R. Bowlby. Geology of the Lake Ontario Basin: a review and outlook. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (8):1503–1516, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-179.

Boyce:1991:PBL

[454] F. M. Boyce, W. M. Schertzer, P. F. Hamblin, and C. R. Murthy. Physical behaviour of Lake Ontario with reference to contaminant pathways and climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8): 1517–1528, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-180.

Schelske:1991:HNE

[455] Claire L. Schelske. Historical nutrient enrichment of Lake Ontario: Paleolimnological evidence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8): 1529–1538, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-181.

Crowder:1991:SML

[456] A. Crowder and D. S. Painter. Submerged macrophytes in Lake Ontario: Current knowledge, importance, threats to stability, and needed studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1539–1545, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-182.

Johannsson:1991:CNO

[457] Ora E. Johannsson, Edward L. Mills, and Robert O’Gorman. Changes in the nearshore and offshore zooplankton communities in Lake Ontario: 1981–88. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1546–1557, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-183.

Nalepa:1991:STL

[458] Thomas F. Nalepa. Status and trends of the Lake Ontario macrobenthos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1558–1567, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-184.

Suns:1991:PST

[459] K. Suns, G. Hitchin, and E. Adamek. Present status and temporal trends of organochlorine contaminants in young-of-the-year spottail shiner (Notropis hudsonius) from Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1568–1573, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-185.

Hartig:1991:RLO

[460] John H. Hartig, James F. Kitchell, Donald Scavia, and Stephen B. Brandt. Rehabilitation of Lake Ontario: the role of nutrient reduction and food web dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8): 1574–1580, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-186.

Smith:1991:CWT

[461] P. G. R. Smith, V. Glooschenko, and D. A. Hagen. Coastal wetlands of three Canadian Great Lakes: Inventory, current conservation initiatives, and patterns of variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (8):1581–1594, August 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-187.

Brown:1991:LOS

[462] Tommy L. Brown, Barbara A. Knuth, and Fredric C. Menz. Lake Ontario’s sport fisheries: Socioeconomic research progress and needs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1595–1601, August 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-188.

Anonymous:1991:CRPb

[463] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(8):1602, August 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
189.

Devlin:1991:ICD

[464] Robert H. Devlin, B. Kelly McNeil, T. David D. Groves, and Edward M. Donaldson. Isolation of a Y-chromosomal DNA probe capable of determining genetic sex in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1606–1612, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-190.

Greenberg:1991:SEA

[465] Joshua A. Greenberg, Scott C. Matulich, and Ron C. Mittelhammer. A system-of-equations approach to modeling age-structured fish populations: The case of Alaskan red king crab, Paralithodes camtschaticus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1613–1622, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-191.

Ramakrishna:1991:TRF

[466] N. R. Ramakrishna and M. D. B. Burt. Tissue response of fish to invasion by larval Pseudoterranova decipiens (Nematoda; Ascaridoidea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1623–1628, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-192.

Lindholm:1991:RDM

[467] Tore Lindholm and Jussi A. O. Meriluoto. Recurrent depth maxima of the hepatotoxic cyanobacterium Oscillatoria agardhii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1629–1634, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-193.

Keller:1991:RCZ

[468] W. Keller and N. D. Yan. Recovery of crustacean zooplankton species richness in Sudbury area lakes following water quality improvements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1635–1644, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-194.

Hendricks:1991:PPW

[469] Susan P. Hendricks and David S. White. Physicochemical patterns within a hyporheic zone of a Northern Michigan river, with comments on surface water patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1645– 1654, September 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-195.

Cutler:1991:PRR

[470] Michael G. Cutler, Sylvia E. Bartlett, Sheila E. Hartley, and William S. Davidson. A polymorphism in the ribosomal RNA genes distinguishes Atlantic salmon (Salmo salar) from North America and Europe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1655–1661, September 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-196.

Shackell:1991:SMI

[471] Nancy L. Shackell and Roger W. Doyle. Scale morphology as an index of developmental stability and stress resistance of Tilapia (Oreochromis niloticus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1662–1669, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-197.

Moser:1991:REM

[472] Mary L. Moser, Alan F. Olson, and Thomas P. Quinn. Riverine and estuarine migratory behavior of Coho salmon (Oncorhynchus kisutch) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1670–1678, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-198.

Cowen:1991:ODS

[473] Robert K. Cowen, Louis A. Chiarella, Candice J. Gomez, and Michael A. Bell. Offshore distribution, size, age, and lateral plate variation of late larval/early juvenile sticklebacks (Gasterosteus) off the Atlantic Coast of New Jersey and New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (9):1679–1684, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f91-199.

Ankley:1991:BDT

[474] Gerald T. Ankley, Donald E. Tillitt, John P. Giesy, Paul D. Jones, and David A. Verbrugge. Bioassay-derived 2,3,7,8-tetrachlorodibenzo-p-dioxin equivalents in PCB-containing extracts from the flesh and eggs of Lake Michigan Chinook salmon (Oncorhynchus tshawytscha) and possible implications for reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1685–1690, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-200.

Sullivan:1991:SAE

[475] Patrick J. Sullivan. Stock abundance estimation using depth-dependent trends and spatially correlated variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1691–1703, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-201.

Hodgson:1991:TIC

[476] James R. Hodgson, Carol J. Hodgson, and Scott M. Brooks. Trophic interaction and competition between largemouth bass (Micropterus salmoides) and rainbow trout (Oncorhynchus mykiss) in a manipulated lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1704–1712, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-202.

Butler:1991:MRP

[477] John L. Butler. Mortality and recruitment of Pacific sardine, Sardinops sagax caerulea, larvae in the California Current. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1713–1723, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-203.

Nolan:1991:DAA

[478] Kathleen Nolan, Joseph Grossfield, and Isaac Wirgin. Discrimination among Atlantic Coast populations of American shad (Alosa sapidissima) using mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1724–1734, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-204.

East:1991:SFR

[479] Pierre East and Pierre Magnan. Some factors regulating piscivory of brook trout, Salvelinus fontinalis, in lakes of the Laurentian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1735–1743, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-205.

Gharrett:1991:TGH

[480] A. J. Gharrett and W. W. Smoker. Two generations of hybrids between even- and odd-year pink salmon (Oncorhynchus gorbuscha): a test for outbreeding depression? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):

1744–1749, September 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-206.

Schiewe:1991:IFC

[481] Michael H. Schiewe, Douglas D. Weber, Mark S. Myers, Frank J. Jacques, William L. Reichert, Cheryl A. Krone, Donald C. Malins, Bruce B. Mc-Cain, Sin-Lam Chan, and Usha Varanasi. Induction of foci of cellular alteration and other hepatic lesions in English sole (Patophrys vetulus) exposed to an extract of an urban marine sediment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1750–1760, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-207.

McGeer:1991:PRC

[482] James C. McGeer, Leanne Baranyi, and George K. Iwama. Physiological responses to challenge tests in six stocks of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1761–1771, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-208.

Thiebaux:1991:AMA

[483] M. L. Thiebaux, P. R. Boudreau, and L. M. Dickie. An analytical model of acoustic fish reflection for estimation of maximum dorsal aspect target strength. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1772– 1782, September 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-209.

Swain:1991:MDB

[484] D. P. Swain, B. E. Riddell, and C. B. Murray. Morphological differences between hatchery and wild populations of Coho salmon (Oncorhynchus kisutch): Environmental versus genetic origin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1783–1791, September 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-210.

Fox:1991:EOM

[485] Michael G. Fox and Allen Keast. Effect of overwinter mortality on reproductive life history characteristics of pumpkinseed (Lepomis gibbosus) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1792– 1799, September 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-211.

Appleton:1991:BCT

[486] Tracey E. Appleton and M. D. B. Burt. Biochemical characterization of third-stage larval sealworm, Pseudoterranova decipiens (Nematoda: Anisakidae), in Canadian Atlantic waters using isoelectric focusing of soluble proteins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1800– 1803, September 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-212.

Baumann:1991:TFW

[487] Paul C. Baumann, Michael J. Mac, Stephen B. Smith, and John C. Harshbarger. Tumor frequencies in walleye (Stizostedion vitreum) and brown bullhead (Ictalurus nebulosus) and sediment contaminants in tributaries of the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1804–1810, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-213.

Andersen:1991:MEA

[488] Donald E. Andersen, Scott D. Reid, Thomas W. Moon, and Steve F. Perry. Metabolic effects associated with chronically elevated cortisol in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(9):1811–1817, September 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-214.

Pepin:1991:SEL

[489] Pierre Pepin and Ransom A. Myers. Significance of egg and larval size to recruitment variability of temperate marine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1820–1828, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-215.

Handy:1991:EIC

[490] R. D. Handy and F. B. Eddy. Effects of inorganic cations on Na^{+} adsorption to the gill and body surface of rainbow trout, Oncorhynchus mykiss, in dilute solutions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10): 1829–1837, October 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-216.

Jonsson:1991:LBS

[491] Bror Jonsson, Jan Henning L’Abée-Lund, Tor G. Heggberget, Arne J. Jensen, Bjørn O. Johnsen, Tor F. Næsje, and Leif M. Sættem. Longevity, body size, and growth in anadromous brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1838–1845, October 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-217.

Havens:1991:CZF

[492] Karl E. Havens. Crustacean zooplankton food web structure in lakes of varying acidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10): 1846–1852, October 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-218.

Olsen:1991:MII

[493] T. Mark Olsen, David M. Lodge, Gregory M. Capelli, and Robert J. Houlihan. Mechanisms of impact of an introduced crayfish (Orconectes rusticus) on littoral congeners, snails, and macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1853–1861, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-219.

Sogard:1991:IOM

[494] Susan M. Sogard. Interpretation of otolith microstructure in juvenile winter flounder (Pseudopleuronectes americanus): Ontogenetic development, daily increment validation, and somatic growth relationships.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1862–1871, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-220.

Wilson:1991:MES

[495] John W. Wilson, Albert E. Giorgi, and Lowell C. Stuehrenberg. A method for estimating spill effectiveness for passing juvenile salmon and its application at Lower Granite Dam on the Snake River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1872–1876, October 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-221.

Cattaneo:1991:EBS

[496] Antonella Cattaneo and Guy Roberge. Efficiency of a brush sampler to measure periphyton in streams and lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1877–1881, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-222.

Young:1991:BAF

[497] Michael K. Young, Wayne A. Hubert, and Thomas A. Wesche. Biases associated with four stream substrate samplers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1882–1886, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-223.

Gordoa:1991:CCB

[498] Ana Gordoa and Joseph E. Hightower. Changes in catchability in a bottom-trawl fishery for Cape hake (Merluccius capensis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1887–1895, October 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-224.

Coughlin:1991:OFB

[499] David J. Coughlin. Ontogeny of feeding behavior of first-feeding Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10): 1896–1904, October 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-225.

Dabrowski:1991:SAA

[500] Konrad Dabrowski. Some aspects of ascorbate metabolism in developing embryos of the brine shrimp (Artemia salina). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1905–1908, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-226.

Chow-Fraser:1991:UMI

[501] Patricia Chow-Fraser. Use of the morphoedaphic index to predict nutrient status and algal biomass in some Canadian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1909–1918, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-227.

Schael:1991:GLP

[502] Denise M. Schael, Lars G. Rudstam, and John R. Post. Gape limitation and prey selection in larval yellow perch (Perca flavescens), freshwater drum (Aplodinotus grunniens), and black crappie (Pomoxis nigromaculatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1919–1925, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f91-
228.

Poff:1991:DRB

[503] N. LeRoy Poff and J. V. Ward. Drift responses of benthic invertebrates to experimental streamflow variation in a hydrologically stable stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1926–1936, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-229.

Rempel:1991:SVM

[504] Robert S. Rempel and Peter J. Colby. A statistically valid model of the morphoedaphic index. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10): 1937–1943, October 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-230.

Rooney:1991:ETD

[505] Patricia Rooney and J. Stanley Cobb. Effects of time of day, water temperature, and water velocity on swimming by postlarvae of the American lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1944–1950, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-231.

Steinman:1991:IEN

[506] Alan D. Steinman, Patrick J. Mulholland, and David B. Kirschtel. Interactive effects of nutrient reduction and herbivory on biomass, taxonomic structure, and P uptake in lotic periphyton communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1951–1959, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-232.

Taylor:1991:OCP

[507] W. D. Taylor, J. H. Carey, D. R. S. Lean, and D. J. McQueen. Organochlorine concentrations in the plankton of lakes in Southern Ontario and their relationship to plankton biomass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1960–1966, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-233.

Ahrens:1991:PLL

[508] Martin A. Ahrens and Robert H. Peters. Patterns and limitations in limnoplankton size spectra. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (10):1967–1978, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f91-234.

Hanna:1991:ESP

[509] Micheline Hanna and Robert Henry Peters. Effect of sampling protocol on estimates of phosphorus and chlorophyll concentrations in lakes of low to moderate trophic status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1979–1986, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-235.

Bergman:1991:LAF

[510] H. L. Bergman and J. S. Mattice. Lake acidification and fisheries project: Acclimation to low pH and elevated aluminum by trouts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1987–1988, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-236.

Walker:1991:ELT

[511] R. L. Walker, C. M. Wood, and H. L. Bergman. Effects of long-term preexposure to sublethal concentrations of acid and aluminum on the ventilatory response to aluminum challenge in brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1989–1995, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-237.

Reid:1991:ASA

[512] Scott D. Reid, D. G. McDonald, and R. R. Rhem. Acclimation to sublethal aluminum: Modifications of metal–gill surface interactions of juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):1996–2005, October 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-238.

McDonald:1991:NTC

[513] D. G. McDonald, C. M. Wood, R. G. Rhem, M. E. Mueller, D. R. Mount, and H. L. Bergman. Nature and time course of acclimation to aluminum in juvenile brook trout (Salvelinus fontinalis). I. Physiology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):2006–2015, October 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-239.

Mueller:1991:NTC

[514] M. E. Mueller, D. A. Sanchez, H. L. Bergman, D. G. McDonald, R. G. Rhem, and C. M. Wood. Nature and time course of acclimation to aluminum in juvenile brook trout (Salvelinus fontinalis). II. Gill histology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):2016–2027, October 1991.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-240.

Freda:1991:SBT

[515] J. Freda, D. A. Sanchez, and H. L. Bergman. Shortening of branchial tight junction acid-exposed rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):2028–2033, October 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-241.

Anonymous:1991:LCLd

[516] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(10):2034–2035, October 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-242.

Roby:1991:MEA

[517] D. Roby, J. D. Lambert, and J. M. Sévigny. Morphometric and electrophoretic approaches to discrimination of capelin (Mallotus villosus) populations in the estuary and Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2040–2050, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-243.

Maclsaac:1991:ISB

[518] Hugh J. Maclsaac, W. Gary Sprules, and J. H. Leach. Ingestion of smallbodied zooplankton by zebra mussels (Dreissena polymorpha): Can cannibalism on larvae influence population dynamics? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2051–2060, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-244.

Mundie:1991:RSP

[519] J. H. Mundie, K. S. Simpson, and C. J. Perrin. Responses of stream periphyton and benthic insects to increases in dissolved inorganic phosphorus in a mesocosm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11): 2061–2072, November 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-245.

Campbell:1991:TLH

[520] A. Campbell, D. J. Noakes, and R. W. Elner. Temperature and lobster, Homarus americanus, yield relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2073–2082, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-246.

Morgan:1991:ESG

[521] John D. Morgan and George K. Iwama. Effects of salinity on growth, metabolism, and ion regulation in juvenile rainbow and steelhead trout (Oncorhynchus mykiss) and fall Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2083– 2094, November 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-247.

Gordoa:1991:SDS

[522] Ana Gordoa and Carlos M. Duarte. Size-dependent spatial distribution of hake (Merluccius capensis and Merluccius paradoxus) in Namibian waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2095–2099, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-248.

Stefansson:1991:GPS

[523] Sigurd O. Stefansson, Björn Th. Bjömsson, Tom Hansen, Carl Haux, G. Lasse Taranger, and Richard L. Saunders. Growth, parr–smolt transformation, and changes in growth hormone of Atlantic salmon (Salmo salar) reared under different photoperiods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2100–2108, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-249.

Cerrato:1991:ANE

[524] Robert M. Cerrato. Analysis of nonlinearity effects in expected-value parameterizations of the von Bertalanffy equation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2109–2117, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-250.

Tanasichuk:1991:VDD

[525] R. W. Tanasichuk, D. M. Ware, W. Shaw, and G. A. McFarlane. Variations in diet, daily ration, and feeding periodicity of Pacific hake (Merluccius productus) and spiny dogfish (Squalus acanthias) off the Lower West Coast of Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2118–2128, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-251.

Pelletier:1991:AIS

[526] Dominique Pelletier and Philippe Gros. Assessing the impact of sampling error on model-based management advice: Comparison of equilibrium yield per recruit variance estimators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2129–2139, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-252.

Edge:1991:MMV

[527] Thomas A. Edge, Don E. McAllister, and Sami U. Qadri. Meristic and morphometric variation between the endangered Acadian whitefish, Coregonus huntsmani, and the lake whitefish, Coregonus clupeaformis, in the Canadian Maritime provinces and the State of Maine, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2140–2151, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-253.

Liskauskas:1991:GVF

[528] Arunas P. Liskauskas and Moira M. Ferguson. Genetic variation and fitness: a test in a naturalized population of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2152–2162, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-254.

Skilbrei:1991:ITL

[529] Ove T. Skilbrei. Importance of threshold length and photoperiod for the development of bimodal length–frequency distribution in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11): 2163–2172, November 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-255.

Millar:1991:SLG

[530] R. B. Millar. Selecting loci for genetic stock identification using maximum likelihood, and the connection with curvature methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2173–2179, November 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-256.

Lydersen:1991:FHN

[531] Christian Lydersen, Lars Anker Angantyr, Øystein Wiig, and Torger Øritsland. Feeding habits of Northeast Atlantic harp seals (Phoca groenlandica) along the summer ice edge of the Barents Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2180–2183, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-257.

Matthews:1991:MAT

[532] G. B. Matthews, R. A. Matthews, and B. Hachmöller. Mathematical analysis of temporal and spatial trends in the benthic macroinvertebrate communities of a small stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2184–2190, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-258.

Haas:1991:SDD

[533] Gordon R. Haas and J. D. McPhail. Systematics and distributions of Dolly Varden (Salvelinus malma) and bull trout (Salvelinus confluentus) in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11): 2191–2211, November 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-259.

Doyon:1991:BCT

[534] Jean-François Doyon, Christiane Hudon, Roderick Morin, and F. G. Whoriskey, Jr. Bénéfices à court terme des mouvements anadromes saisonniers pour une population d’omble de fontaine (Salvelinus fontinalis) du Nouveau Québec. (French) [Short-term benefits of seasonal anadromous movements for a population of brook trout (Salvelinus fontinalis) in New Quebec]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11): 2212–2222, November 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-260.

Peterson:1991:MBP

[535] R. H. Peterson, J. Power, and D. J. Martin-Robichaud. Morphological basis of the pectoral fin flutter of embryonic Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2223–2227, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-261.

Paine:1991:EIO

[536] Michael D. Paine, William C. Leggett, Jeff K. McRuer, and Kenneth T. Frank. Effects of incubation in oiled sediment on emergence of capelin (Mallotus villosus) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (11):2228–2239, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f91-262.

Tyler:1991:MRP

[537] A. V. Tyler and W. R. Crawford. Modeling of recruitment patterns in Pacific cod (Gadus macrocephalus) in Hecate Strait, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2240–2249, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-263.

OGorman:1991:UZA

[538] Robert O’Gorman, Edward L. Mills, and Joe DeGisi. Use of zooplankton to assess the movement and distribution of alewife (Alosa pseudoharengus) in South-Central Lake Ontario in spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2250–2257, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-264.

Hesslein:1991:SIS

[539] R. H. Hesslein, M. J. Capel, D. E. Fox, and K. A. Hallard. Stable isotopes of sulfur, carbon, and nitrogen as indicators of trophic level and fish migration in the Lower Mackenzie River Basin, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2258–2265, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-265.

Dawe:1991:CLI

[540] Earl G. Dawe, David M. Taylor, John M. Hoenig, William G. Warren, Gerald P. Ennis, Robert G. Hooper, William E. Donaldson, A. J. Paul, and Judy M. Paul. A critical look at the idea of terminal molt in male snow crab (Chionoecetes opilio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2266–2275, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-266.

Sullivan:1991:IBA

[541] Barbara K. Sullivan, Peter H. Doering, Candace A. Oviatt, Aimee A. Keller, and Jeffrey B. Frithsen. Interactions with the benthos alter pelagic food web structure in coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2276–2284, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-267.

Lalonde:1991:EBR

[542] Sophie Lalonde and John A. Downing. Epiphyton biomass is related to lake trophic status, depth, and macrophyte architecture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2285–2291, November 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-268.

Schindler:1991:LCL

[543] D. W. Schindler. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(11):2292, November 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-269.

Ware:1991:LBL

[544] Daniel M. Ware and Richard E. Thomson. Link between long-term variability in upwelling and fish production in the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2296–2306, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-270.

Naesje:1991:HSN

[545] T. F. Næsje, B. Jonsson, O. T. Sandlund, and G. Kjellberg. Habitat switch and niche overlap in coregonid fishes: Effects of zooplankton abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2307–2315, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f91-271.

Kaartvedt:1991:OTP

[546] S. Kaartvedt, T. M. Johnsen, D. L. Aksnes, U. Lie, and H. Svendsen. Occurrence of the toxic phytoflagellate Prymnesium parvum and associated fish mortality in a Norwegian fjord system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2316–2323, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-272.

France:1991:USF

[547] Robert France, John Holmes, and Alejandro Lynch. Use of size–frequency data to estimate the age composition of crayfish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2324–2332, December 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-273.

Kjesbu:1991:FAE

[548] O. S. Kjesbu, J. Klungsøyr, H. Kryvi, P. R. Witthames, and M. Greer Walker. Fecundity, atresia, and egg size of captive Atlantic cod (Gadus morhua) in relation to proximate body composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2333–2343, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-274.

Gearing:1991:HMC

[549] Juanita N. Gearing, Dale E. Buckley, and John N. Smith. Hydrocarbon and metal contents in a sediment core from Halifax Harbour: a chronology of contamination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12): 2344–2354, December 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-275.

Murkin:1991:EPF

[550] Henry R. Murkin, John A. Kadlec, and Elaine J. Murkin. Effects of prolonged flooding on nektonic invertebrates in small diked marshes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2355–2364, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-276.

Schweigert:1991:MDP

[551] J. F. Schweigert. Multivariate description of Pacific herring (Clupea harengus pallasi) stocks from size and age information. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2365–2376, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-277.

Marsden:1991:SHO

[552] J. Ellen Marsden and Charles C. Krueger. Spawning by hatchery-origin lake trout (Salvelinus namaycush) in Lake Ontario: Data from egg collections, substrate analysis, and diver observations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2377–2384, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-278.

Berges:1991:SSW

[553] Jonn A. Berges and James S. Ballantyne. Size scaling of whole-body maximal enzyme activities in aquatic crustaceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2385–2394, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-279.

Welch:1991:SDI

[554] H. E. Welch, M. A. Bergmann, T. D. Siferd, and P. S. Amarualik. Seasonal development of ice algae near Chesterfield Inlet, N.W.T., Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2395–2402, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-280.

Welch:1991:EAB

[555] D. W. Welch, L. Margolis, M. A. Henderson, and S. McKinnell. Evidence for attacks by the bathypelagic fish Anotopterus pharao (Myctophiformes) on Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2403–2407, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-281.

Ardisson:1991:AGP

[556] P.-L. Ardisson and E. Bourget. Abundance, growth, and production estimation of the blue mussel Mytilus edulis on moored navigation buoys in the estuary and Northwestern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2408–2419, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-282.

Henderson:1991:MEG

[557] Bryan A. Henderson and Jonathan L. Wong. A method for estimating gillnet selectivity of walleye (Stizostedion vitreum vitreum) in muitimesh multifilament gill nets in Lake Erie, and its application. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2420–2428, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-283.

Woodley:1991:PRI

[558] Thomas H. Woodley and Andrew J. Read. Potential rates of increase of a harbour porpoise (Phocoena phocoena) population subjected to incidental mortality in commercial fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2429–2435, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-284.

Trites:1991:DTH

[559] Andrew W. Trites. Does tagging and handling affect the growth of northern fur seal pups (Callorhinus ursinus)? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2436–2442, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-285.

Eales:1991:CPL

[560] J. G. Eales, D. G. Cyr, K. Finnson, and C. E. Johnston. Changes in plasma T_{4} and T_{3} levels during reconditioning and rematuration in male and female wild Atlantic salmon (Salmo salar) kelts held in freshwater under two photoperiod regimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2443–2448, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-286.

Jackson:1991:BEC

[561] Togwell A. Jackson. Biological and environmental control of mercury accumulation by fish in lakes and reservoirs of Northern Manitoba, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2449–2470, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f91-287.

Hammill:1991:LRS

[562] M. O. Hammill, C. Lydersen, M. Ryg, and T. G. Smith. Lactation in the ringed seal (Phoca hispida). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2471–2476, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-288.

Voelz:1991:BRA

[563] Neal J. Voelz and J. V. Ward. Biotic responses along the recovery gradient of a regulated stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (12):2477–2490, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f91-289.

Vigg:1991:TDM

[564] Steven Vigg and Craig C. Burley. Temperature-dependent maximum daily consumption of juvenile salmonids by northern squawfish (Ptychocheilus oregonensis) from the Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2491–2498, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-290.

Bilby:1991:CFL

[565] Robert E. Bilby and James W. Ward. Characteristics and function of large woody debris in streams draining old-growth, clear-cut, and second-growth forests in Southwestern Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2499–2508, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f91-291.

Johnson:1991:ETA

[566] M. G. Johnson, S. E. George, and J. R. M. Kelso. Estimation of temporal acidification of Central Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2509–2522, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-292.

Kelso:1991:FRB

[567] J. R. M. Kelso and M. G. Johnson. Factors related to the biomass and production of fish communities in small, oligotrophic lakes vulnerable to acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12): 2523–2532, December 1991. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f91-293.

Norris:1991:FPY

[568] James G. Norris. Further perspectives on yield per recruit analysis and biological reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48 (12):2533–2542, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f91-294.

Stewart:1991:BRI

[569] James E. Stewart. A brief review of the International Council for the Exploration of the Sea (ICES) on the occasion of the formation of the North Pacific Marine Science Organization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2543–2550, December 1991. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f91-295.

Anonymous:1991:LCLe

[570] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 48(12):2551–2553, December 1991. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f91-296.

Arnason:1992:MIV

[571] A. N. Arnason, M. H. Papst, and G. E. Hopky. Modelling the increase in variance of fish weight. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (1):2–16, January 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-001.

Miskimmin:1992:IDO

[572] Brenda M. Miskimmin, John W. M. Rudd, and Carol A. Kelly. Influence of dissolved organic carbon, pH, and microbial respiration rates on mercury methylation and demethylation in lake water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):17–22, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-002.

Boudreau:1992:AOP

[573] Paul R. Boudreau. Acoustic observations of patterns of aggregation in haddock (Melanogrammus aeglefinus) and their significance to production and catch. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):23–31, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-003.

Simard:1992:MEB

[574] Yvan Simard, Pierre Legendre, Gilles Lavoie, and Denis Marcotte. Mapping, estimating biomass, and optimizing sampling programs for spatially autocorrelated data: Case study of the northern shrimp (Pandalus borealis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):32–45, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-004.

Tremblay:1992:RLE

[575] Gilles Tremblay and Louis A. Giguère. Relation longueur/écaille allométrique chez le saumon atlantique (Salmo salar) durant la phase marine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):46–51, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-005.

Grapentine:1992:RFA

[576] Lee C. Grapentine and David M. Rosenberg. Responses of the freshwater amphipod Hyalella azteca to environmental acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):52–64, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-006.

Renberg:1992:PAA

[577] Ingemar Renberg and Hans Hultberg. A paleolimnological assessment of acidification and liming effects on diatom assemblages in a Swedish lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):65–72, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-007.

Flinkman:1992:PBH

[578] Juha Flinkman, Ilppo Vuorinen, and Eero Aro. Planktivorous Baltic herring (Clupea harengus) prey selectively on reproducing copepods and cladocerans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):73–77, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-008.

Whitehead:1992:PSF

[579] Hal Whitehead, Susan Waters, and Thomas Lyrholm. Population structure of female and immature sperm whales (Physeter macrocephalus) off the Galápagos Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (1):78–84, January 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-009.

Douglas:1992:PDA

[580] Donald J. Douglas and Stephen S. Bates. Production of domoic acid, a neurotoxic amino acid, by an axenic culture of the marine diatom Nitzschia pungens f. multiseries Hasle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):85–90, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-010.

Wildish:1992:BBL

[581] D. J. Wildish, A. J. Wilson, and B. Frost. Benthic boundary layer macrofauna of Browns Bank, Northwest Atlantic, as potential prey of juvenile benthic fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (1):91–98, January 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-011.

Foote:1992:CCS

[582] Chris J. Foote, Chris C. Wood, W. Craig Clarke, and John Blackburn. Circannual cycle of seawater adaptability in Oncorhynchus nerka: Genetic differences between sympatric sockeye salmon and Kokanee. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):99–109, January 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-012.

Hinterleitner:1992:SER

[583] S. Hinterleitner, M. Huber, R. Lackner, and W. Wieser. Systemic and enzymatic responses to endurance training in two cyprinid species with different life styles (Teleostei: Cyprinidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):110–115, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-013.

Kingston:1992:ATF

[584] J. C. Kingston, H. J. B. Birks, A. J. Uutala, B. F. Cumming, and J. P. Smol. Assessing trends in fishery resources and lake water aluminum from paleolimnological analyses of siliceous algae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):116–127, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-014.

Cumming:1992:HMA

[585] B. F. Cumming, J. P. Smol, J. C. Kingston, D. F. Charles, H. J. B. Birks, K. E. Camburn, S. S. Dixit, A. J. Uutala, and A. R. Selle. How much acidification has occurred in Adirondack region lakes (New York, USA) since preindustrial times? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (1):128–141, January 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-015.

Benton:1992:AGDa

[586] Michael J. Benton and Sheldon I. Guttman. Allozyme genotype and differential resistance to mercury pollution in the caddisfly, Nectopsyche albida. I. Single-locus genotypes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):142–146, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-016.

Benton:1992:AGDb

[587] Michael J. Benton and Sheldon I. Guttman. Allozyme genotype and differential resistance to mercury pollution in the caddisfly, Nectopsyche albida. II. Multilocus genotypes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):147–149, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-017.

Ramcharan:1992:MMP

[588] Charles W. Ramcharan, Dianna K. Padilla, and Stanley I. Dodson. A multivariate model for predicting population fluctuations of Dreissena polymorpha in North American lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):150–158, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-018.

Ramcharan:1992:PRP

[589] Charles W. Ramcharan, Stanley I. Dodson, and Jason Lee. Predation risk, prey behavior, and feeding rate in Daphnia pulex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):159–165, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-019.

Lindesjoo:1992:SDN

[590] Eric Lindesjöö and Jan Thulin. A skeletal deformity of northern pike (Esox lucius) related to pulp mill effluents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):166–172, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-020.

Bodensteiner:1992:RTD

[591] Leo R. Bodensteiner and William M. Lewis. Role of temperature, dissolved oxygen, and backwaters in the winter survival of freshwater drum (Aplodinotus grunniens) in the Mississippi River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):173–184, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-021.

France:1992:ILP

[592] Robert L. France and Pamela M. Welbourn. Influence of Lake pH and macrograzers on the distribution and abundance of nuisance metaphytic algae in Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (1):185–195, January 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-022.

Anonymous:1992:LCLa

[593] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(1):196–199, January 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-023.

Rose:1992:ITS

[594] G. A. Rose. Indices of total stock biomass in the “Northern” and Gulf of St. Lawrence Atlantic cod (Gadus morhua) stocks derived from time series analyses of fixed gear (trap) catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):202–209, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-024.

Koslow:1992:FSR

[595] J. Anthony Koslow. Fecundity and the stock–recruitment relationship. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):210–217, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-025.

Jackson:1992:GTC

[596] George D. Jackson and J. Howard Choat. Growth in tropical cephalopods: an analysis based on statolith microstructure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):218–228, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-026.

Arnesen:1992:DHC

[597] Arne M. Arnesen, Morten Halvorsen, and Kjell J. Nilssen. Development of hypoosmoregulatory capacity in Arctic char (Salvelinus alpinus) reared under either continuous light or natural photoperiod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):229–237, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-027.

Brander:1992:DES

[598] Keith Brander and Peter C. F. Hurley. Distribution of early-stage Atlantic cod (Gadus morhua), haddock (Melanogrammus aeglefinus), and witch flounder (Glyptocephalus cynoglossus) eggs on the Scotian Shelf: a reappraisal of evidence on the coupling of cod spawning and plankton production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):238–251, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-028.

Fisher:1992:IRD

[599] T. R. Fisher and D. R. S. Lean. Interpretation of radiophosphate dynamics in lake waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2): 252–258, February 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-029.

Kerans:1992:SBS

[600] B. L. Kerans. Season and body size influence movement of hydropsychid caddisfly larvae (Trichoptera: Hydropsychidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):259–265, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-030.

Melvin:1992:UMM

[601] Gary D. Melvin, Michael J. Dadswell, and Joseph A. McKenzie. Usefulness of meristic and morphometric characters in discriminating populations of American shad (Alosa sapidissima) (Ostreichthyes: Clupeidae) inhabiting a marine environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):266–280, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-031.

Walters:1992:GMC

[602] Carl Walters and J. C. Woodey. Genetic models for cyclic dominance in sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):281–292, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-032.

Pereira:1992:DSI

[603] Donald L. Pereira, Yosef Cohen, and George R. Spangler. Dynamics and species interactions in the commercial fishery of the red lakes, Minnesota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):293–302, February 1992.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-033.

MacPhee:1992:MTP

[604] Daniel J. MacPhee, Louis A. Hanic, Dianne L. Friesen, and David E. Sims. Morphology of the toxin-producing diatom Nitzschia pungens Grunow forma multiseries Hasle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):303–311, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-034.

Novaczek:1992:DDA

[605] I. Novaczek, M. S. Madhyastha, R. F. Ablett, A. Donald, G. Johnson, M. S. Nijjar, and D. E. Sims. Depuration of domoic acid from live blue mussels (Mytilus edulis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (2):312–318, February 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-035.

Mulligan:1992:MDA

[606] Timothy J. Mulligan, Robert W. Chapman, and Bonnie L. Brown. Mitochondrial DNA analysis of walleye pollock, Theragra chalcogramma, from the Eastern Bering Sea and Shelikof Strait, Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):319–326, February 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-036.

Kuris:1992:MCF

[607] Armand M. Kuris and Kevin D. Lafferty. Modelling crustacean fisheries: Effects of parasites on management strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):327–336, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-037.

Richardson:1992:CPD

[608] John S. Richardson. Coarse particulate detritus dynamics in small, montane streams Southwestern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):337–346, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-038.

Likely:1992:VCC

[609] C. G. Likely and M. D. B. Burt. In vitro cultivation of Contracaecum osculatum (Nematoda: Anisakidae) from third-stage larvae to egg-laying adults. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):347–348, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-039.

Chopin:1992:IHF

[610] T. Chopin, J. D. Pringle, and R. E. Semple. Impact of harvesting on frond density and biomass of Irish moss (Chondrus crispus Stackhouse) beds in the Southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):349–357, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-040.

Brown:1992:IPG

[611] James R. Brown, Andrew T. Beckenbach, and Michael J. Smith. Influence of Pleistocene glaciations and human intervention upon mitochondrial DNA diversity in white sturgeon (Acipenser transmontanus) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):358–367, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-041.

Ouellet:1992:LCS

[612] Patrick Ouellet, Christopher T. Taggart, and Kenneth T. Frank. Lipid condition and survival in shrimp (Pandalus borealis) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):368–378, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-042.

Fahnenstiel:1992:PPL

[613] Gary L. Fahnenstiel and Hunter J. Carrick. Phototrophic picoplankton in Lakes Huron and Michigan: Abundance, distribution, composition, and contribution to biomass and production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):379–388, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-043.

Miller:1992:SVN

[614] Charles B. Miller, John Fulton, and Bruce W. Frost. Size variation of Neocalanus plumchrus and Neocalanus flemingeri in a 20-yr sample series from the Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (2):389–399, February 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-044.

Silvert:1992:DMF

[615] W. Silvert and D. V. Subba Rao. Dynamic model of the flux of domoic acid, a neurotoxin, through a Mytilus edulis population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):400–405, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-045.

Neary:1992:MPS

[616] B. P. Neary and J. H. Leach. Mapping the potential spread of the zebra mussel (Dreissena polymorpha) in Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):406–415, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-046.

Jude:1992:EGG

[617] David J. Jude, Robert H. Reider, and Gerald R. Smith. Establishment of Gobiidae in the Great Lakes Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):416–421, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-047.

Yan:1992:BCS

[618] Norman D. Yan, Warren I. Dunlop, Trevor W. Pawson, and Lori E. MacKay. Bythotrephes cederstroemi (Schoedler) in Muskoka Lakes: First records of the European invader in inland lakes in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):422–426, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-048.

Anonymous:1992:CRPa

[619] Anonymous. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(2):427, February 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-049.

Schnute:1992:SAE

[620] Jon T. Schnute. Statistical analysis of embedded replicates in mark– recovery experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (3):432–442, March 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-050.

Staurnes:1992:DST

[621] Magne Staurnes, Trygve Sigholt, Grete Lysfjord, and Odd A. Gulseth. Difference in the seawater tolerance of anadromous and landlocked populations of Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):443–447, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-051.

Schmitz:1992:AVR

[622] Monika Schmitz. Annual variations in rheotactic behaviour and seawater adaptability in landlocked Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):448–452, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-052.

Ogura:1992:SBC

[623] Miki Ogura and Yukimasa Ishida. Swimming behavior of Coho salmon, Oncorhynchus kisutch, in the open sea as determined by ultrasonic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):453–457, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-053.

Varnavsky:1992:SWC

[624] Vladimir S. Varnavsky, Tatsuya Sakamoto, and Tetsuya Hirano. Stunting of wild Coho salmon (Oncorhynchus kisutch) in seawater: Patterns of plasma thyroid hormones, cortisol, and growth hormone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):458–461, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-054.

Martin:1992:DDW

[625] A. R. Martin and T. G. Smith. Deep diving in wild, free-ranging beluga whales, Delphinapterus leucas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):462–466, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-055.

Frank:1992:LFD

[626] Kenneth T. Frank, John W. Loder, James E. Carscadden, William C. Leggett, and Christopher T. Taggart. Larval flatfish distributions and drift on the Southern Grand Bank. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):467–483, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-056.

Polacheck:1992:RYC

[627] T. Polacheck, D. Mountain, D. McMillan, W. Smith, and P. Berrien. Recruitment of the 1987 year class of Georges Bank haddock (Melanogrammus aeglefinus): The influence of unusual larval transport. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):484–496, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-057.

Grafton:1992:RCI

[628] R. Quentin Grafton. Rent capture in an individual transferable quota fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):497–503, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-058.

Hill:1992:GNS

[629] Walter R. Hill, Harry L. Boston, and Alan D. Steinman. Grazers and nutrients simultaneously limit lotic primary productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):504–512, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-059.

Marcogliese:1992:NAC

[630] David J. Marcogliese. Neomysis americana (Crustacea: Mysidacea) as an intermediate host for sealworm, Pseudoterranova decipiens (Nematoda: Ascaridoidea), and spirurid nematodes (Acuarioidea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):513–515, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-060.

Goddard:1992:APF

[631] Sally V. Goddard, Ming H. Kao, and Garth L. Fletcher. Antifreeze production, freeze resistance, and overwintering of juvenile Northern Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (3):516–522, March 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-061.

Boisclair:1992:ESM

[632] Daniel Boisclair. An evaluation of the stereocinematographic method to estimate fish swimming speed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):523–531, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-062.

Morin:1992:MEP

[633] Antoine Morin and Nathalie Bourassa. Modèles empiriques de la production annuelle et du rapport P / B d’invertébrés benthiques d’eau courante. (French) [Empirical models of the annual production and the P / B ratio of running water benthic invertebrates]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):532–539, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-063.

Bilby:1992:AVA

[634] Robert E. Bilby and Peter A. Bisson. Allochthonous versus autochthonous organic matter contributions to the trophic support of fish populations in clear-cut and old-growth forested streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):540–551, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-064.

Bozek:1992:GMS

[635] Michael A. Bozek and Frank J. Rahel. Generality of microhabitat suitability models for young Colorado River cutthroat trout (Oncorhynchus clarki pleuriticus) across sites and among years in Wyoming streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):552–564, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-065.

Hudon:1992:DPF

[636] C. Hudon, D. G. Parsons, and R. Crawford. Diel pelagic foraging by a pandalid shrimp (Pandalus montagui) off Resolution Island (Eastern Hudson Strait). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):565–576, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-066.

Jones:1992:RTP

[637] Ronald D. Jones and José A. Amador. Removal of total phosphorus and phosphate by peat soils of the Florida Everglades. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):577–583, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-067.

Bayley:1992:EMF

[638] S. E. Bayley, D. W. Schindler, K. G. Beaty, B. R. Parker, and M. P. Stainton. Effects of multiple fires on nutrient yields from streams draining boreal forest and fen watersheds: Nitrogen and phosphorus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):584–596, March 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-068.

Beauchamp:1992:SFA

[639] J. J. Beauchamp, S. W. Christensen, and E. P. Smith. Selection of factors affecting the presence of brook trout (Salvelinus fontinalis) in Adirondack lakes: a case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (3):597–608, March 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-069.

Morris:1992:PDN

[640] James T. Morris and Bruce C. Coull. Population dynamics, numerical production, and potential predation impact on a meiobenthic copepod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(3):609–616, March 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-070.

Mackay:1992:CLM

[641] Rosemary J. Mackay. Colonization by lotic macroinvertebrates: a review of processes and patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (3):617–628, March 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-071.

Mulligan:1992:LAA

[642] T. J. Mulligan and B. M. Leaman. Length-at-age analysis: Can you get what you see? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (4):632–643, April 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-072.

Rajasilta:1992:RBF

[643] Marjut Rajasilta. Relationship between food, fat, sexual maturation, and spawning time of Baltic herring (Clupea harengus membras) in the Archipelago Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):644–654, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-

Timmermans:1992:CZU

[644] Klaas R. Timmermans, Elly Spijkerman, Marcel Tonkes, and Harrie Govers. Cadmium and zinc uptake by two species of aquatic invertebrate predators from dietary and aqueous sources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):655–662, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-074.

Forbes:1992:ZRS

[645] J. R. Forbes, R. W. Macdonald, E. C. Carmack, K. Iseki, and M. C. O’Brien. Zooplankton retained in sequential sediment traps along the Beaufort Sea shelf break during winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):663–670, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-075.

Osborne:1992:ITS

[646] Lewis L. Osborne and Michael J. Wiley. Influence of tributary spatial position on the structure of warmwater fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):671–681, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-076.

Fausch:1992:LWD

[647] Kurt D. Fausch and Thomas G. Northcote. Large woody debris and salmonid habitat in a small coastal British Columbia stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):682–693, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-077.

James:1992:PVM

[648] William F. James, William D. Taylor, and John W. Barko. Production and vertical migration of Ceratium hirundinella in relation to phosphorus availability in Eau Galle Reservoir, Wisconsin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):694–700, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-078.

Talbot:1992:SIL

[649] André J. Talbot and Roger W. Doyle. Statistical interrelation of length, growth, and scale circulus spacing: Use of ossification to detect nongrowing fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):701–707, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-079.

Smith:1992:ISD

[650] Barry D. Smith, Gordon A. McFarlane, and Mark W. Saunders. Inferring the summer distribution of migratory Pacific hake (Merluccius productus) from latitudinal variation in mean lengths-at-age and length frequency distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (4):708–721, April 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-080.

Servos:1992:EFP

[651] Mark R. Servos, Derek C. G. Muir, and G. R. Barrie Webster. Environmental fate of polychlorinated dibenzo-p-dioxins in lake enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):722–734, April 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-081.

Servos:1992:BPD

[652] Mark R. Servos, Derek C. G. Muir, and G. R. Barrie Webster. Bioavailability of polychlorinated dibenzo-p-dioxins in lake enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):735–742, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-082.

McFarlane:1992:CIL

[653] G. A. McFarlane and R. J. Beamish. Climatic influence linking copepod production with strong year-classes in sablefish, Anoplopoma fimbria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):743–753, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-083.

Methven:1992:SRP

[654] David A. Methven, Laurence W. Crim, Birgitta Norberg, Joseph A. Brown, Gregory P. Goff, and Ingvar Huse. Seasonal reproduction and plasma levels of sex steroids and vitellogenin in Atlantic halibut (Hippoglossus hippoglossus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (4):754–759, April 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-084.

Foote:1992:ELW

[655] C. J. Foote, J. W. Clayton, C. C. Lindsey, and R. A. Bodaly. Evolution of lake whitefish (Coregonus clupeaformis) in North America during the Pleistocene: Evidence for a Nahanni glacial refuge race in the Northern Cordillera region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (4):760–768, April 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-085.

Bodaly:1992:ELW

[656] R. A. Bodaly, J. W. Clayton, C. C. Lindsey, and J. Vuorinen. Evolution of lake whitefish (Coregonus clupeaformis) in North America during the Pleistocene: Genetic differentiation between sympatric populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):769–779, April 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-086.

Snyder:1992:UDS

[657] Randal J. Snyder, Brian A. McKeown, Konrad Colbow, and Robert Brown. Use of dissolved strontium in scale marking of juvenile salmonids: Effects of concentration and exposure time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):780–782, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-087.

Nickelson:1992:SCH

[658] Thomas E. Nickelson, Jeffrey D. Rodgers, Steven L. Johnson, and Mario F. Solazzi. Seasonal changes in habitat use by juvenile Coho salmon (Oncorhynchus kisutch) in Oregon coastal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):783–789, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-088.

Nickelson:1992:ESS

[659] Thomas E. Nickelson, Mario F. Solazzi, Steven L. Johnson, and Jeffrey D. Rodgers. Effectiveness of selected stream improvement techniques to create suitable summer and winter rearing habitat for juvenile Coho salmon (Oncorhynchus kisutch) in Oregon coastal streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):790–794, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-089.

Smith:1992:RDO

[660] M. Kimberly Smith. Regional differences in otolith morphology of the deep slope red snapper Etelis carbunculus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):795–804, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-090.

Miller:1992:BSO

[661] Thomas J. Miller, Larry B. Crowder, James A. Rice, and Fred P. Binkowski. Body size and the ontogeny of the functional response in fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):805–812, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-091.

Browman:1992:FPS

[662] Howard I. Browman and W. John O’Brien. Foraging and prey search behaviour of golden shiner (Notemigonus crysoleucas) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):813–819, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-092.

Thomson:1992:CTF

[663] Richard E. Thomson, Michael C. Healey, John F. T. Morris, and Gary A. Borstad. Commercial troll fishing vessel distribution off Vancouver Island during July 1988: Relation to observed physical oceanography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):820–832, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-093.

Chmielewski:1992:RIB

[664] C. M. Chmielewski and R. J. Hall. Responses of immature blackflies (Diptera: Simuliidae) to experimental pulses of acidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(4):833–840, April 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-094.

Crawford:1992:ASS

[665] R. E. Crawford, C. Hudon, and D. G. Parsons. An acoustic study of shrimp (Pandalus montagui) distribution near Resolution Island (Eastern Hudson Strait). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (5):842–856, May 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-095.

Sameoto:1992:EOG

[666] D. D. Sameoto and A. W. Herman. Effect of the outflow from the Gulf of St. Lawrence on Nova Scotia shelf zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):857–869, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-096.

Audet:1992:DSC

[667] Céline Audet and Guy Claireaux. Diel and seasonal changes in resting levels of various blood parameters in brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):870–877, May 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-097.

Saunders:1992:PCA

[668] R. L. Saunders, A. P. Farrell, and D. E. Knox. Progression of coronary arterial lesions in Atlantic salmon (Salmo salar) as a function of growth rate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):878–884, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
098.

Buijse:1992:IWC

[669] Anthone D. Buijse, Leendert A. Schaap, and Tammo P. Bust. Influence of water clarity on the catchability of six freshwater fish species in bottom trawls. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):885–893, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-099.

Buijse:1992:PGS

[670] Anthonie D. Buijse and Rob P. Houthuijzen. Piscivory, growth, and sizeselective mortality of age 0 pikeperch (Stizostedion lucioperca). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):894–902, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-100.

Mackas:1992:SCZ

[671] David L. Mackas. Seasonal cycle of zooplankton off Southwestern British Columbia: 1979–89. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (5):903–921, May 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-101.

Francis:1992:URA

[672] R. I. C. C. Francis. Use of risk analysis to assess fishery management strategies: a case study using orange roughy (Hoplostethus atlanticus) on the Chatham Rise, New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):922–930, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-102.

Kope:1992:OHR

[673] Robert G. Kope. Optimal harvest rates for mixed stocks of natural and hatchery fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):931–938, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
103.

Hall:1992:HRT

[674] Timothy J. Hall, Richard K. Haley, Dennis L. Borton, Alexander H. Walsh, and Richard E. Wolke. Histopathology of rainbow trout (Oncorhynchus mykiss) after long-term exposure to biologically treated bleached kraft mill effluent in experimental stream channels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):939–944, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-104.

Catalan:1992:EDP

[675] Jordi Catalan. Evolution of dissolved and particulate matter during the ice-covered period in a deep, high-mountain lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):945–955, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-105.

Brattey:1992:ANS

[676] John Brattey and I-Hsun Ni. Ascaridoid nematodes from the stomach of harp seals, Phoca groenlandica, from Newfoundland and Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):956–966, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-106.

McGurk:1992:CTM

[677] Michael D. McGurk and Wolfgang C. Kusser. Comparison of three methods of measuring RNA and DNA concentrations of individual Pacific herring, Clupea pallasi, larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (5):967–974, May 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-107.

France:1992:TVF

[678] Robert L. France and Robert H. Peters. Temporal variance function for total phosphorus concentration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):975–977, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f92-108.

Miller:1992:RBC

[679] P. A. Miller, K. R. Munkittrick, and D. G. Dixon. Relationship between concentrations of copper and zinc in water, sediment, benthic invertebrates, and tissues of white sucker (Catostomus commersoni) at metal-contaminated sites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (5):978–984, May 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-109.

Downing:1992:SAP

[680] John A. Downing and William L. Downing. Spatial aggregation, precision, and power in surveys of freshwater mussel populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):985–991, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-110.

Lalonde:1992:PEM

[681] Sophie Lalonde and John A. Downing. Phytofaun of eleven macrophyte beds of differing trophic status, depth, and composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):992–1000, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-111.

Lacasse:1992:BAD

[682] Sylvain Lacasse and Pierre Magnan. Biotic and abiotic determinants of the diet of brook trout, Salvelinus fontinalis, in lakes of the Laurentian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1001–1009, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-112.

Bloom:1992:CFM

[683] Nicolas S. Bloom. On the chemical form of mercury in edible fish and marine invertebrate tissue. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (5):1010–1017, May 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-113.

Ricker:1992:BCF

[684] William E. Ricker. Back-calculation of fish lengths based on proportionality between scale and length increments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1018–1026, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-114.

Clark:1992:ISF

[685] Robert A. Clark. Influence of stream flows and stock size on recruitment of Arctic grayling (Thymallus arcticus) in the Chena River, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1027–1034, May 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-115.

Patalas:1992:CPL

[686] K. Patalas and A. Salki. Crustacean plankton in Lake Winnipeg: Variation in space and time as a function of lake morphology, geology, and climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1035–1059, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-116.

Pace:1992:ZAE

[687] Michael L. Pace, Stuart E. G. Findlay, and David Lints. Zooplankton in advective environments: The Hudson River community and a comparative analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1060– 1069, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-117.

Kleinow:1992:PBD

[688] Kevin M. Kleinow, Wendy L. Beilfuss, Herman H. Jarboe, Brad F. Droy, and John J. Lech. Pharmacokinetics, bioavailability, distribution, and metabolism of sulfadimethoxine in the rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1070–1077, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-118.

Anonymous:1992:LCLb

[689] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(5):1078–1079, May 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-119.

Borgmann:1992:BPD

[690] U. Borgmann and D. M. Whittle. Bioenergetics and PCB, DDE, and mercury dynamics in Lake Ontario lake trout (Salvelinus namaycush): a model based on surveillance data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1086–1096, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-120.

Berglund:1992:SAB

[691] Ingemar Berglund, Monika Schmitz, and Hans Lundqvist. Seawater adaptability in Baltic salmon (Salmo salar): a bimodal smoltification pattern in previously mature males. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1097–1106, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-121.

BorgstrOm:1992:RBA

[692] Reidar BorgstrØm. Relationship between annual recruitment and density in a lacustrine population of allopatric brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1107–1113, June 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-122.

Waddy:1992:SVS

[693] S. L. Waddy and D. E. Aiken. Seasonal variation in spawning by preovigerous American lobster (Homarus americanus) in response to temperature and photoperiod manipulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1114–1117, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-123.

Yurista:1992:EPD

[694] Peder M. Yurista. Embryonic and postembryonic development in Bythotrephes cederstroemii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (6):1118–1125, June 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-124.

Hilborn:1992:ILS

[695] Ray Hilborn. Institutional learning and spawning channels for sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1126–1136, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-125.

Frost:1992:VDV

[696] Bruce W. Frost and Stephen M. Bollens. Variability of diel vertical migration in the marine planktonic copepod Pseudocalanus newmani in relation to its predators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (6):1137–1141, June 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-126.

Davies:1992:AAS

[697] Ronald W. Davies, Frederick J. Wrona, and V. Kalarani. Assessment of activity-specific metabolism of aquatic organisms: an improved system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1142–1148, June 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-127.

Dutil:1992:SVP

[698] J.-D. Dutil, J. Munro, C. Audet, and M. Besner. Seasonal variation in the physiological response of Atlantic cod (Gadus morhua) to low salinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1149–1156, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-128.

Ferguson:1992:PEB

[699] R. A. Ferguson and B. L. Tufts. Physiological effects of brief air exposure in exhaustively exercised rainbow trout (Oncorhynchus mykiss): Implications for “Catch and release” fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1157–1162, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-129.

Berard:1992:MSG

[700] Hélène Bérard, Edwin Bourget, and Marcel Fréchette. Mollusk shell growth: External microgrowth ridge formation is uncoupled to environmental factors in Mytilus edulis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1163–1170, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-130.

Agbeti:1992:RBD

[701] Michael D. Agbeti. Relationship between diatom assemblages and trophic variables: a comparison of old and new approaches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1171–1175, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-131.

He:1992:ESP

[702] Xi He and Russell A. Wright. An experimental study of piscivore– planktivore interactions: Population and community responses to predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1176–1183, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-132.

Ferraro:1992:TLS

[703] Steven P. Ferraro and Faith A. Cole. Taxonomic level sufficient for assessing a moderate impact on macrobenthic communities in Puget Sounds, Washington, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1184– 1188, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-133.

McCormick:1992:OFD

[704] J. Howard McCormick and Kathleen M. Jensen. Osmoregulatory failure and death of first-year largemouth bass (Micropterus salmoides) exposed to low pH and elevated aluminum, at low temperature in soft water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1189–1197, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-134.

Lohman:1992:ENE

[705] Kirk Lohman, John R. Jones, and Bruce D. Perkins. Effects of nutrient enrichment and flood frequency on periphyton biomass in Northern Ozark streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6): 1198–1205, June 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-135.

Chang:1992:PIL

[706] Cecily C. Y. Chang, James S. Kuwabara, and Sofie P. Pasilis. Phosphate and iron limitation of phytoplankton biomass in Lake Tahoe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1206–1215, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-136.

DeVries:1992:CIB

[707] Dennis R. DeVries and Roy A. Stein. Complex interactions between fish and zooplankton: Quantifying the role of an open-water planktivore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1216–1227, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-137.

Chen:1992:CBP

[708] Y. Chen, D. A. Jackson, and H. H. Harvey. A comparison of von Bertalanffy and polynomial functions in modelling fish growth data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1228–1235, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-138.

Choulik:1992:RSL

[709] O. Choulik and T. R. Moore. Response of a subarctic lake chain to reduced sewage loading. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (6):1236–1245, June 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-139.

Benda:1992:MES

[710] L. Benda, T. J. Beechie, R. C. Wissmar, and A. Johnson. Morphology and evolution of salmonid habitats in a recently deglaciated river basin, Washington State, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (6):1246–1256, June 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-140.

Hill:1992:FLI

[711] Walter R. Hill. Food limitation and interspecific competition in snaildominated streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6): 1257–1267, June 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-141.

Clare:1992:ACL

[712] Anthony S. Clare, John D. Costlow, Hassan M. Bedair, and George Lumb. Assessment of crab limb regeneration as an assay for developmental toxicity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1268– 1273, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-142.

Anonymous:1992:LCLc

[713] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(6):1274–1280, June 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-143.

Sainte-Marie:1992:MMS

[714] Bernard Sainte-Marie and François Hazel. Moulting and mating of snow crabs, Chionoecetes opilio (O. Fabricius), in shallow waters of the Northwestern Gulf of Saint Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1282–1293, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-144.

McAllister:1992:SEL

[715] Murdoch K. McAllister, Randall M. Peterman, and Darren M. Gillis. Statistical evaluation of a large-scale fishing experiment designed to test for a genetic effect of size-selective fishing on British Columbia pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (7):1294–1304, July 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-145.

McAllister:1992:DAL

[716] Murdoch K. McAllister and Randall M. Peterman. Decision analysis of a large-scale fishing experiment designed to test for a genetic effect of sizeselective fishing on British Columbia pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1305–1314, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-146.

Richards:1992:SME

[717] Laura J. Richards and Jon T. Schnute. Statistical models for estimating CPUE from catch and effort data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1315–1327, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-147.

Forseth:1992:RME

[718] Torbjørn Forseth, Bror Jonsson, Roger Næumann, and Ola Ugedal. Radioisotope method for estimating food consumption by brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1328– 1335, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-148.

Guenette:1992:PCC

[719] Sylvie Guénette, Daniel Goyette, Réjean Fortin, Jean Leclerc, Nelson Fournier, Gilles Roy, and Pierre Dumont. La périodicité de la croissance chez la femelle de l’esturgeon jaune (Acipenser fulvescens) du fleuve Saint-Laurent est-elle reliée à la périodicité de la reproduction?. (French) [Is the periodicity of growth in female lake sturgeon (Acipenser fulvescens) from the Saint-Lawrence River related to the periodicity of reproduction?]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1336– 1342, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-149.

Pearson:1992:ESG

[720] Walter H. Pearson, John R. Skalski, and Charles I. Malme. Effects of sounds from a geophysical survey device on behavior of captive rockfish (Sebastes spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (7):1343–1356, July 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-150.

Skalski:1992:ESG

[721] John R. Skalski, Walter H. Pearson, and Charles I. Malme. Effects of sounds from a geophysical survey device on catch-per-unit-effort in a hook-and-line fishery for rockfish (Sebastes spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1357–1365, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-151.

Swartzman:1992:SAB

[722] Gordon Swartzman, Chisheng Huang, and Stephen Kaluzny. Spatial analysis of Bering Sea groundfish survey data using generalized additive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1366–1378, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-152.

Hobbs:1992:IHC

[723] Roderick C. Hobbs, Louis W. Botsford, and Andrew Thomas. Influence of hydrographic conditions and wind forcing on the distribution and abundance of Dungeness crab, Cancer magister, larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1379–1388, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-153.

Servizi:1992:SRC

[724] James A. Servizi and Dennis W. Martens. Sublethal responses of Coho salmon (Oncorhynchus kisutch) to suspended sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1389–1395, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-154.

Wickens:1992:SSM

[725] P. A. Wickens, P. A. Shelton, J. H. M. David, J. G. Field, W. H. Oosthuizen, J-P. Roux, and A. M. Starfield. A fur seal simulation model to explore alternative management strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1396–1405, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-155.

Schneider:1992:BMZ

[726] Daniel W. Schneider. A bioenergetics model of zebra mussel, Dreissena polymorpha, growth in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1406–1416, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-156.

Kwak:1992:ADF

[727] Thomas J. Kwak, Michael J. Wiley, Lewis L. Osborne, and R. Weldon Larimore. Application of diel feeding chronology to habitat suitability analysis of warmwater stream fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1417–1430, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-157.

Rowan:1992:PSO

[728] D. J. Rowan, J. Kalff, and J. B. Rasmussen. Profundal sediment organic content and physical character do not reflect lake trophic status, but rather reflect inorganic sedimentation and exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1431–1438, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-158.

Secor:1992:COB

[729] David H. Secor and John Mark Dean. Comparison of otolith-based back-calculation methods to determine individual growth histories of larval striped bass, Morone saxatilis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1439–1454, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-159.

Cyr:1992:GZR

[730] Hélène Cyr and Michael L. Pace. Grazing by zooplankton and its relationship to community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1455–1465, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-160.

Lazzaro:1992:PPD

[731] Xavier Lazzaro, Ray W. Drenner, Roy A. Stein, and J. Durward Smith. Planktivores and plankton dynamics: Effects of fish biomass and planktivore type. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1466–1473, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-161.

Casey:1992:REA

[732] John G. Casey and Lisa J. Natanson. Revised estimates of age and growth of the sandbar shark (Carcharhinus plumbeus) from the Western North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1474–1477, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-162.

McMahon:1992:BHU

[733] Thomas E. McMahon and L. Blair Holtby. Behaviour, habitat use, and movements of Coho salmon (Oncorhynchus kisutch) smolte during seaward migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1478– 1485, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-163.

Roddick:1992:STO

[734] D. L. Roddick and R. J. Miller. Spatial and temporal overlap of the American lobster (Homarus americanus) and sea scallop (Placopecten magellanicus) as related to the impact of inshore scallop dragging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1486–1492, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-164.

Benson:1992:SHL

[735] Barbara J. Benson and John J. Magnuson. Spatial heterogeneity of littoral fish assemblages in lakes: Relation to species diversity and habitat structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1493–1500, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-165.

May:1992:GII

[736] Bernie May and J. Ellen Marsden. Genetic identification and implications of another invasive species of dreissenid mussel in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1501–1506, July 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-166.

Brodziak:1992:TGS

[737] Jon Brodziak, Boyd Bentley, Devin Bartley, Graham A. E. Gall, Richard Gomulkiewicz, and Marc Mangel. Tests of genetic stock identification using coded wire tagged fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (7):1507–1517, July 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-167.

Anonymous:1992:LCLd

[738] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(7):1518–1525, July 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-168.

Boudreau:1992:BSA

[739] P. R. Boudreau and L. M. Dickie. Biomass spectra of aquatic ecosystems in relation to fisheries yield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (8):1528–1538, August 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-169.

Borgstrom:1992:EPD

[740] Reidar Borgstrøm. Effect of population density on gillnet catchability in four allopatric populations of brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1539–1545, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-170.

Borgstrom:1992:GSM

[741] Reidar Borgstrøm and Erik Plahte. Gillnet selectivity and a model for capture probabilities for a stunted brown trout (Salmo trutta) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1546–1554, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
171.

Horppila:1992:OST

[742] Jukka Horppila and Heikki Peltonen. Optimizing sampling from trawl catches: Contemporaneous multistage sampling for age and length structures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1555–1559, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-

Munkittrick:1992:CMP

[743] K. R. Munkittrick, M. E. McMaster, C. B. Portt, G. J. Van Der Kraak, I. R. Smith, and D. G. Dixon. Changes in maturity, plasma sex steroid levels, hepatic mixed-function oxygenase activity, and the presence of external lesions in lake whitefish (Coregonus clupeaformis) exposed to bleached kraft mill effluent. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (8):1560–1569, August 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-173.

Skiftesvik:1992:CBO

[744] Anne Berit Skiftesvik. Changes in behaviour at onset of exogenous feeding in marine fish larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (8):1570–1572, August 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-174.

Kitada:1992:ESE

[745] Syuiti Kitada, Yasushi Taga, and Hirohisa Kishino. Effectiveness of a stock enhancement program evaluated by a two-stage sampling survey of commercial landings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8): 1573–1582, August 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-175.

Reddy:1992:DES

[746] D. C. Reddy, Emil Dratnal, and Ronald W. Davies. Dynamics of energy storage in a freshwater predator (Nephelopsis obscura) following winter stresses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1583–1587, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
176.

McQueen:1992:TLR

[747] Donald J. McQueen, Edward L. Mills, John L. Forney, Mark R. S. Johannes, and John R. Post. Trophic level relationships in pelagic food webs: Comparisons derived from long-term data sets for Oneida Lake, New York (USA), and Lake St. George, Ontario (Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1588–1596, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-177.

Tremblay:1992:PSS

[748] M. John Tremblay and Michael Sinclair. Planktonic sea scallop larvae (Placopecten magellanicus) in the Georges Bank region: Broadscale distribution in relation to physical oceanography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1597–1615, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-178.

Pratt:1992:RGC

[749] Dennis M. Pratt, William H. Blust, and James H. Selgeby. Ruffe, Gymnocephalus cernuus: Newly introduced in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1616–1618, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-179.

Cormier:1992:HEC

[750] Roland J. Cormier, Alan R. Fraser, Richard F. J. Bailey, and Nicole Raymond. Hemolymph ecdysone concentration as a function of sexual maturity in the male snow crab (Chionoecetes opilio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1619–1623, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-180.

Cadigan:1992:RSC

[751] N. G. Cadigan and R. B. Millar. Reliability of selection curves obtained from trouser trawl or alternate haul experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1624–1632, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-181.

Hoke:1992:BPC

[752] Robert A. Hoke, William R. Gala, James B. Drake, John P. Giesy, and Stan Flegler. Bicarbonate as a potential confounding factor in cladoceran toxicity assessments of pore water from contaminated sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1633–1640, August 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-182.

Fairchild:1992:LBE

[753] G. Winfield Fairchild and John W. Sherman. Linkage between epilithic algal growth and water column nutrients in softwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1641–1649, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-183.

Adamson:1992:DPS

[754] M. L. Adamson, D. F. Clease, and L. Margolis. Differentiation of Philonema spp. (Nematoda; Philometridae) in British Columbia salmonid fishes using DNA restriction fragment length differences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1650–1656, August 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-184.

Smith:1992:UBN

[755] Barry D. Smith, Mark W. Saunders, and Gordon A. McFarlane. Use of a binomial–normal functional model to demonstrate that the sex and age composition of Pacific hake (Merluccius productus) aggregations affects estimates of mean lengths-at-age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1657–1669, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-185.

Brodeur:1992:FCJ

[756] Richard D. Brodeur, Robert C. Francis, and William G. Pearcy. Food consumption of juvenile Coho (Oncorhynchus kisutch) and Chinook salmon (O. tshawytscha) on the Continental Shelf off Washington and Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1670–1685, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-

Clements:1992:AIH

[757] William H. Clements, Donald S. Cherry, and John H. Van Hassel. Assessment of the impact of heavy metals on benthic communities at the Clinch River (Virginia): Evaluation of an index of community sensitivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1686–1694, August 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-187.

Morin:1992:FAS

[758] Antoine Morin and Antonella Cattaneo. Factors affecting sampling variability of freshwater periphyton and the power of periphyton studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1695–1703, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-188.

Morin:1992:COF

[759] Pierre-Philippe Morin and Kjell B. Døving. Changes in the olfactory function of Atlantic salmon, Salmo salar, in the course of smoltification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1704–1713, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-189.

Rabeni:1992:TLB

[760] Charles F. Rabeni. Trophic linkage between stream centrarchids and their crayfish prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8): 1714–1721, August 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-190.

Aumen:1992:POC

[761] Nicholas G. Aumen, Cindy L. Crist, Dawn E. Miller, and Keith O. Meals. Particulate organic carbon supply and trophic dynamics in a Mississippi flood-control reservoir dominated by gizzard shad (Dorosoma cepedianum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1722–1733, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-

Evans:1992:HCL

[762] Marlene S. Evans. Historic changes in Lake Michigan zooplankton community structure: The 1960s revisited with implications for top-down control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1734–1749, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
192.

Anonymous:1992:CRPb

[763] Anonymous. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(8):1751, August 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
193.

Shelton:1992:SRD

[764] Peter A. Shelton. The shape of recruitment distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1754–1761, September 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-194.

Patterson:1992:SSA

[765] K. R. Patterson, J. Zuzunaga, and G. Cárdenas. Size of the South American sardine (Sardinops sagax) population in the northern part of the Peru upwelling ecosystem after collapse of anchoveta (Engraulis ringens) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1762– 1769, September 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-195.

Ihssen:1992:AVG

[766] Peter E. Ihssen, G. William Martin, and David W. Rodgers. Allozyme variation of Great Lakes alewife, Alosa pseudoharengus: Genetic differentiation and affinities of a recent invader. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1770–1777, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-196.

Secor:1992:EFS

[767] David H. Secor, John Mark Dean, Thomas A. Curtis, and Forrest W. Sessions. Effect of female size and propagation methods on larval production at a South Carolina striped bass (Morone saxatilis) hatchery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1778–1787, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-197.

Cobb:1992:EDS

[768] D. G. Cobb, T. D. Galloway, and J. F. Flannagan. Effects of discharge and substrate stability on density and species composition of stream insects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1788–1795, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-198.

Matsuda:1992:OHP

[769] Hiroyuki Matsuda, Tatsu Kishida, and Takashi Kidachi. Optimal harvesting policy for chub mackerel (Scomber japonicus) in Japan under a fluctuating environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9): 1796–1800, September 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-199.

Richards:1992:SMA

[770] Laura J. Richards, Jon T. Schnute, A. R. Kronlund, and Richard J. Beamish. Statistical models for the analysis of ageing error. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1801–1815, September 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-200.

Cormack:1992:ACW

[771] R. M. Cormack and J. R. Skalski. Analysis of coded wire tag returns from commercial catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1816– 1825, September 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-201.

Heisey:1992:RTR

[772] Paul G. Heisey, Dilip Mathur, and Ted Rineer. A reliable tag–recapture technique for estimating turbine passage survival: Application to youngof-the-year American shad (Alosa sapidissima). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1826–1834, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-202.

Crim:1992:FRR

[773] L. W. Crim, C. E. Wilson, Y. P. So, D. R. Idler, and C. E. Johnston. Feeding, reconditioning, and rematuration responses of captive Atlantic salmon (Salmo salar) kelt. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1835–1842, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-203.

Labelle:1992:SPC

[774] M. Labelle. Straying patterns of Coho salmon (Oncorhynchus kisutch) stock from Southeast Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1843–1855, September 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-204.

Szedlmayer:1992:VSD

[775] Stephen T. Szedlmayer and Kenneth W. Able. Validation studies of daily increment formation for larval and juvenile summer flounder, Paralichthys dentatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1856– 1862, September 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-205.

Jordan:1992:GPV

[776] W. C. Jordan, A. F. Youngson, D. W. Hay, and A. Ferguson. Genetic protein variation in natural populations of Atlantic salmon (Salmo salar) in Scotland: Temporal and spatial variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1863–1872, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-206.

Thorp:1992:LBI

[777] James H. Thorp. Linkage between islands and benthos in the Ohio River, with implications for riverine management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1873–1882, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-207.

Margolis:1992:AIM

[778] L. Margolis, T. E. McDonald, and D. J. Whitaker. Assessment of the impact of the myxosporean parasite Ceratomyxa shasta on survival of seaward migrating juvenile Chinook salmon, Oncorhynchus tshawytscha, from the Fraser River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1883–1889, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-208.

Deegan:1992:WRF

[779] Linda A. Deegan and Bruce J. Peterson. Whole-river fertilization stimulates fish production in an Arctic tundra river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1890–1901, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-209.

Snyder:1992:MDD

[780] Thomas P. Snyder, Robert D. Larsen, and Stephen H. Bowen. Mitochondrial DNA diversity among Lake Superior and inland lake ciscoes (Coregonus artedi and C. hoyi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1902–1907, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-210.

Turner:1992:EGC

[781] Andrew M. Turner and Gary G. Mittelbach. Effects of grazer community composition and fish on algal dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1908–1915, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-211.

Osgood:1992:MLD

[782] Marcy P. Osgood and Charles W. Boylen. Microbial leaf decomposition in Adirondack streams exhibiting pH gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1916–1923, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-212.

Albright:1992:TSD

[783] L. J. Albright, S. Johnson, and A. Yousif. Temporal and spatial distribution of the harmful diatoms Chaetoceros concavicornis and Chaetoceros convolutus along the British Columbia Coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1924–1931, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-213.

Wu:1992:ODS

[784] Lin Wu and David A. Culver. Ontogenetic diet shift in Lake Erie age-0 yellow perch (Perca flavescens): a size-related response to zooplankton density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1932–1937, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-214.

Campbell:1992:GQN

[785] Peter G. C. Campbell, Hans J. Hansen, Bertrand Dubreuil, and William O. Nelson. Geochemistry of Quebec North Shore salmon rivers during snowmelt: Organic acid pulse and aluminum mobilization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1938–1952, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-215.

McGowan:1992:UNH

[786] Colin McGowan and William S. Davidson. Unidirectional natural hybridization between brown trout (Salmo trutta) and Atlantic salmon (S. salar) in Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9): 1953–1958, September 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-216.

Guenette:1992:MDL

[787] Sylvie Guénette, Eric Rassart, and Réjean Fortin. Morphological differentiation of lake sturgeon (Acipenser fulvescens) from the St. Lawrence River and Lac des Deux Montagnes (Quebec, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1959–1965, September 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-217.

Anonymous:1992:LCLe

[788] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(9):1966–1969, September 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-218.

Irvine:1992:ECS

[789] J. R. Irvine, R. C. Bocking, K. K. English, and M. Labelle. Estimating Coho salmon (Oncorhynchus kisutch) spawning escapements by conducting visual surveys in areas selected using stratified random and stratified index sampling designs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10): 1972–1981, October 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-219.

English:1992:RPE

[790] K. K. English, R. C. Bocking, and J. R. Irvine. A robust procedure for estimating salmon escapement based on the area-under-the-curve method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):1982–1989, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-220.

Gordoa:1992:SDD

[791] Ana Gordoa and Carlos M. Duarte. Size-dependent density of the demersal fish off Namibia: Patterns within and among species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):1990–1993, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-221.

Hughes:1992:RFP

[792] Nicholas F. Hughes. Ranking of feeding positions by drift-feeding Arctic grayling (Thymallus arcticus) in dominance hierarchies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):1994–1998, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-222.

Hughes:1992:SPD

[793] Nicholas F. Hughes. Selection of positions by drift-feeding salmonids in dominance hierarchies: Model and test for Arctic grayling (Thymallus arcticus) in subarctic mountain streams, interior Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):1999–2008, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-223.

Mills:1992:FAA

[794] Edward L. Mills, Robert O’Gorman, Joe DeGisi, Roy F. Heberger, and Robert A. House. Food of the alewife (Alosa pseudoharengus) in Lake Ontario before and after the establishment of Bythotrephes cederstroemi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2009–2019, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-224.

Lapointe:1992:VNM

[795] Michael F. Lapointe, Randall M. Peterman, and Brian J. Rothschild. Variable natural mortality rates inflate variance of recruitments estimated from virtual population analysis (VPA). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2020–2027, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-225.

McConnaughey:1992:JDC

[796] Robert A. McConnaughey, David A. Armstrong, Barbara M. Hickey, and Donald R. Gunderson. Juvenile Dungeness crab (Cancer magister) recruitment variability and oceanic transport during the pelagic larval phase. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2028–2044, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
226.

Bradford:1992:EWF

[797] Rodney G. Bradford and Robert L. Stephenson. Egg weight, fecundity, and gonad weight variability among Northwest Atlantic herring (Clupea harengus) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10): 2045–2054, October 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-227.

Johnston:1992:MRF

[798] C. E. Johnston, M. J. Hambrook, R. W. Gray, and K. G. Davidson. Manipulation of reproductive function in Atlantic salmon (Salmo salar) kelts with controlled photoperiod and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2055–2061, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-228.

Marcogliese:1992:CWA

[799] David J. Marcogliese and Gary McClelland. Corynosoma wegeneri (Acanthocephala: Polymorphida) and Pseudoterranova decipiens (Nematoda: Ascaridoidea) larvae in Scotian Shelf groundfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2062–2069, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-229.

Glynn:1992:CTM

[800] Anders Wicklund Glynn, Carl Haux, and Christer Hogstrand. Chronic toxicity and metabolism of Cd and Zn in juvenile minnows (Phoxinus phoxinus) exposed to a Cd and Zn mixture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2070–2079, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-230.

Ankley:1992:BPS

[801] Gerald T. Ankley, Philip M. Cook, Anthony R. Carlson, Daniel J. Call, Julie A. Swenson, Henry F. Corcoran, and Robert A. Hoke. Bioaccumulation of PCBs from sediments by oligochaetes and fishes: Comparison of laboratory and field studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2080–2085, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-231.

MacDonald:1992:ODI

[802] B. A. MacDonald and C. D. Bajdik. Orientation and distribution of individual Placopecten magellanicus (Gmelin) in two natural populations with differing production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2086–2092, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-232.

Bechara:1992:TEB

[803] José A. Bechara, Guy Moreau, and Dolors Planas. Top-down effects of brook trout (Salvelinus fontinalis) in a boreal forest stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2093–2103, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-233.

Suuronen:1992:SSD

[804] Petri Suuronen and Russell B. Millar. Size selectivity of diamond and square mesh codends in pelagic herring trawls: Only small herring will notice the difference. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10): 2104–2117, October 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-234.

Trippel:1992:FSQ

[805] Edward A. Trippel and John D. Neilson. Fertility and sperm quality of virgin and repeat-spawning Atlantic cod (Gadus morhua) and associated hatching success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10): 2118–2127, October 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-235.

Chambers:1992:TSD

[806] P. A. Chambers, E. E. Prepas, and K. Gibson. Temporal and spatial dynamics in riverbed chemistry: The influence of flow and sediment composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2128–2140, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f92-
236.

Cain:1992:AIB

[807] Daniel J. Cain, Samuel N. Luoma, James L. Carter, and Steven V. Fend. Aquatic insects as bioindicators of trace element contamination in cobble-bottom rivers and streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2141–2154, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-237.

Niederlehner:1992:CRC

[808] B. R. Niederlehner and John Cairns, Jr. Community response to cumulative toxic impact: Effects of acclimation on zinc tolerance of Aufwuchs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2155–2163, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-238.

Holtby:1992:FAV

[809] L. Blair Holtby, Ronald K. Kadowaki, and Kent S. Simpson. Factors affecting the vulnerability of juvenile Coho (Oncothynchus kisutch) and Chinook salmon (O. tshawytscha) in a saltwater sport fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2164–2178, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-239.

Appenzeller:1992:BHE

[810] A. R. Appenzeller and W. C. Leggett. Bias in hydroacoustic estimates of fish abundance due to acoustic shadowing: Evidence from day–night surveys of vertically migrating fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2179–2189, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-240.

Truscott:1992:ARS

[811] Beryl Truscott, David R. Idler, and Garth L. Fletcher. Alteration of reproductive steroids of male winter flounder (Pleuronectes americanus) chronically exposed to low levels of crude oil in sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2190–2195, October 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-241.

Winemiller:1992:PLH

[812] Kirk O. Winemiller and Kenneth A. Rose. Patterns of life-history diversification in North American fishes: implications for population regulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(10):2196–2218, October 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-242.

Frank:1992:DCA

[813] Kenneth T. Frank. Demographic consequences of age-specific dispersal in marine fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11): 2222–2231, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-243.

Castonguay:1992:OMA

[814] Martin Castonguay, George A. Rose, and William C. Leggett. Onshore movements of Atlantic mackerel (Scomber scombrus) in the Northern Gulf of St. Lawrence: Associations with wind-forced advections of warmed surface waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11): 2232–2241, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-244.

Simard:1992:GCB

[815] Patrick Simard, Martin Castonguay, Denis D’Amours, and Pierre Magnan. Growth comparison between juvenile Atlantic mackerel (Scomber scombrus) from the two spawning groups of the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2242–2248, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-245.

Hurley:1992:FTI

[816] D. A. Hurley. Feeding and trophic interactions of white perch (Morone americana) in the Bay of Quinte, Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2249–2259, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-246.

Vijayan:1992:AHS

[817] M. M. Vijayan and T. W. Moon. Acute handling stress alters hepatic glycogen metabolism in food-deprived rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2260–2266, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-247.

Jayaram:1992:IDP

[818] Muthanna G. Jayaram and F. W. H. Beamish. Influence of dietary protein and lipid on nitrogen and energy losses in lake trout, Salvelinus namaycush. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2267– 2272, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-248.

Duston:1992:EMP

[819] J. Duston and R. L. Saunders. Effect of 6-, 12-, and 18-month photoperiod cycles on smolting and sexual maturation in juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11): 2273–2280, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-249.

Robarts:1992:LNW

[820] Richard D. Robarts, Marlene S. Evans, and Michael T. Arts. Light, nutrients, and water temperature as determinants of phytoplankton production in two saline, prairie lakes with high sulphate concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2281–2290, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-250.

vanUrk:1992:LCP

[821] Geert van Urk, Frans C. M. Kerkurm, and Henk Smit. Life cycle patterns, density, and frequency of deformities in Chironomus larvae (Diptera:

Chironomidae) over a contaminated sediment gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2291–2299, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-251.

Clarke:1992:GCJ

[822] W. Craig Clarke, Ruth E. Withler, and John E. Shelbourn. Genetic control of juvenile life history pattern in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2300– 2306, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-252.

Ueno:1992:DMC

[823] Yasuhiro Ueno. Deepwater migrations of chum salmon (Oncorhynchus keta) along the Pacific Coast of Northern Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2307–2312, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-253.

Lanfersieck:1992:PMI

[824] John Lanfersieck and Dale Squires. Planning models for individual transferable quota programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11): 2313–2321, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-254.

Holomuzki:1992:RPF

[825] Joseph R. Holomuzki and R. Jan Stevenson. Role of predatory fish in community dynamics of an ephemeral stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2322–2330, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-255.

Webb:1992:CCE

[826] D. J. Webb, B. K. Burnison, A. M. Trimbee, and E. E. Prepas. Comparison of chlorophyll a extractions with ethanol and dimethyl sulfoxide/acetone, and a concern about spectrophotometric phaeopigment correction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2331–2336, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-256.

Lisle:1992:EST

[827] Thomas E. Lisle and Jack Lewis. Effects of sediment transport on survival of salmonid embryos in a natural stream: a simulation approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2337–2344, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-257.

Phillips:1992:SPI

[828] Ruth B. Phillips, Kay A. Pleyte, and Michael R. Brown. Salmonid phylogeny inferred from ribosomal DNA restriction maps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2345–2353, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-258.

Jorgenson:1992:RAT

[829] John K. Jorgenson, Harold E. Welch, and Martin F. Curtis. Response of amphipoda and Trichoptera to lake fertilization in the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2354–2362, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-259.

Molot:1992:PES

[830] L. A. Molot, P. J. Dillon, B. J. Clark, and B. P. Neary. Predicting end-of-summer oxygen profiles in stratified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2363–2372, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-260.

Nogami:1992:BBA

[831] Kinya Nogami and Masachika Maeda. Bacteria as biocontrol agents for rearing larvae of the crab Portunus trituberculatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2373–2376, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-261.

Rand:1992:FLP

[832] Peter S. Rand, Charles A. S. Hall, William H. McDowell, Neil H. Ringler, and Jonathan G. Kennen. Factors limiting primary productivity in Lake Ontario tributaries receiving salmon migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2377–2385, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-262.

Martin:1992:PGS

[833] Andrew P. Martin, Robert Humphreys, and Stephen R. Palumbi. Population genetic structure of the armorhead, Pseudopentaceros wheeleri, in the North Pacific Ocean: Application of the polymerase chain reaction to fisheries problems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11): 2386–2391, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-263.

Havens:1992:AAE

[834] Karl E. Havens. Acid and aluminum effects on sodium homeostasis and survival of acid-sensitive and acid-tolerant cladocera. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2392–2398, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-264.

Brauner:1992:ESD

[835] C. J. Brauner, J. M. Shrimpton, and D. J. Randall. Effect of shortduration seawater exposure on plasma ion concentrations and swimming performance in Coho salmon (Oncorhynchus kisutch) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2399–2405, November 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-265.

Johnsen:1992:IFC

[836] Peter B. Johnsen and Steven W. Lloyd. Influence of fat content on uptake and depuration of the off-flavor 2-methylisoborneol by channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11): 2406–2411, November 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-266.

Hudson:1992:BPF

[837] J. J. Hudson, J. C. Roff, and B. K. Burnison. Bacterial productivity in forested and open streams in Southern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2412–2422, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-267.

Nakashima:1992:PCM

[838] Brian S. Nakashima. Patterns in coastal migration and stock structure of capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(11):2423–2429, November 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-268.

Fee:1992:INO

[839] E. J. Fee and R. E. Hecky. Introduction to the Northwest Ontario Lake Size Series (NOLSS). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12): 2434–2444, December 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-269.

Fee:1992:ELS

[840] E. J. Fee, J. A. Shearer, E. R. DeBruyn, and E. U. Schindler. Effects of lake size on phytoplankton photosynthesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2445–2459, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-270.

Comeau:1992:MGM

[841] Michel Comeau and Gérard Y. Conan. Morphometry and gonad maturity of male snow crab, Chionoecetes opilio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2460–2468, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-271.

Cowey:1992:FDP

[842] C. B. Cowey and C. Y. Cho. Failure of dietary putrescine to enhance the growth of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2469–2473, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-272.

Nelson:1992:MSY

[843] Jay A. Nelson and John J. Magnuson. Metabolic stores of yellow perch (Perca flavescens): Comparison of populations from an acidic, dystrophic lake and circumneutral, mesotrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2474–2482, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-273.

Petticrew:1992:WFC

[844] Ellen L. Petticrew and Jacob Kalff. Water flow and clay retention in submerged macrophyte beds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (12):2483–2489, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f92-274.

Rowan:1992:EMD

[845] D. J. Rowan, J. Kalff, and J. B. Rasmussen. Estimating the mud deposition boundary depth in lakes from wave theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2490–2497, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-275.

Schindler:1992:NRS

[846] Daniel E. Schindler. Nutrient regeneration by sockeye salmon (Oncorhynchus nerka) fry and subsequent effects on zooplankton and phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2498–2506, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-276.

Havens:1992:AEA

[847] Karl E. Havens. Acidification effects on the algal–zooplankton interface. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2507–2514, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-277.

Lo:1992:IRA

[848] N. Chyan huei Lo, Larry D. Jacobson, and James L. Squire. Indices of relative abundance from fish spotter data based on delta-lognornial models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2515–2526, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f92-278.

Zwanenburg:1992:MDD

[849] K. C. T. Zwanenburg, P. Bentzen, and J. M. Wright. Mitochondrial DNA differentiation in Western North Atlantic populations of haddock (Melanogrammus aeglefinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49 (12):2527–2537, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f92-279.

Stott:1992:GPS

[850] Wendylee Stott, Moira M. Ferguson, and Ross F. Tallman. Genetic population structure of American plaice (Hippoglossoides platessoides) from the Gulf of St. Lawrence, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2538–2545, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-280.

Hammill:1992:HSC

[851] M. O. Hammill, G. B. Stenson, and R. A. Myers. Hooded seal (Cystophora cristata) pup production in the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2546–2550, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-281.

Petersen:1992:FRC

[852] James H. Petersen and Donald L. DeAngelis. Functional response and capture timing in an individual-based model: Predation by northern squawfish (Ptychocheilus oregonensis) on juvenile salmonids in the Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2551– 2565, December 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-282.

Boisclair:1992:RBF

[853] Daniel Boisclair. Relationship between feeding and activity rates for actively foraging juvenile brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2566–2573, December 1992. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-283.

Dillon:1992:PGH

[854] Robert T. Dillon, Jr. and John J. Manzi. Population genetics of the hard clam, Mercenaria mercenaria, at the northern limit of its range. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2574–2578, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f92-284.

Guckert:1992:PRI

[855] James B. Guckert, Stephen C. Nold, Harry L. Boston, and David C. White. Periphyton response in an industrial receiving stream: Lipidbased physiological stress analysis and pattern recognition of microbial community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12): 2579–2587, December 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-285.

Valerio:1992:SNA

[856] Paul F. Valerio, Sally V. Goddard, Ming H. Kao, and Garth L. Fletcher. Survival of Northern Atlantic cod (Gadus morhua) eggs and larvae when exposed to ice and low temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2588–2595, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-286.

Kraft:1992:EPN

[857] Clifford E. Kraft. Estimates of phosphorus and nitrogen cycling by fish using a bioenergetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2596–2604, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-287.

Watson:1992:SRB

[858] Susan Watson, Edward McCauley, and John A. Downing. Sigmoid relationships between phosphorus, algal biomass, and algal community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2605–2610, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f92-288.

Ramcharan:1992:MPP

[859] Charles W. Ramcharan, Dianna K. Padilla, and Stanley I. Dodson. Models to predict potential occurrence and density of the zebra mussel, Dreissena polymorpha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2611– 2620, December 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-289.

Johnson:1992:LHJ

[860] S. W. Johnson, J. F. Thedinga, and K. V. Koski. Life history of juvenile ocean-type Chinook salmon (Oncorhynchus tshawytscha) in the Situk River, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12): 2621–2629, December 1992. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f92-290.

Feltmate:1992:RBD

[861] Blair W. Feltmate, D. Dudley Williams, and Adriane Montgomerie. Relationship between diurnal activity patterns, cryptic coloration, and subsequent avoidance of predaceous fish by perlid stoneflies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2630–2634, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-291.

Brattey:1992:LAS

[862] John Brattey and Claude A. Bishop. Larval Anisakis simplex (Nematoda: Ascaridoidea) infection in the musculature of Atlantic cod, Gadus morhua, from Newfoundland and Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2635–2647, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f92-292.

Rogers:1992:NDG

[863] Jean Beyer Rogers and Ellen K. Pikitch. Numerical definition of groundfish assemblages caught off the coasts of Oregon and Washington using commercial fishing strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 49(12):2648–2656, December 1992. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f92-293.

Martel:1993:DRZ

[864] André Martel. Dispersal and recruitment of zebra mussel (Dreissena polymorpha) in a nearshore area in West-central Lake Erie: The significance of postmetamorphic drifting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (1):3–12, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-001.

Haag:1993:RSF

[865] Wendell R. Haag, David J. Berg, David W. Garton, and Jerry L. Farris. Reduced survival and fitness in native bivalves in response to fouling by the introduced zebra mussel (Dreissena polymorpha) in Western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):13–19, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-002.

Lewis:1993:SMR

[866] Andrew R. Lewis and J. Howard Choat. Spawning mode and reproductive output of the tropical cephalopod Idiosepius pygmaeus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):20–28, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-003.

Richardson:1993:CIM

[867] William B. Richardson and Stephen T. Threlkeld. Complex interactions of multiple aquatic consumers: an experimental mesocosm manipulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):29–42, January 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-004.

Harding:1993:MAL

[868] Gareth Harding, Ellen Kenchington, and Zhensui Zheng. Morphometrics of American lobster (Homarus americanus) larvae in relation to stock determinations in the maritimes, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):43–52, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-005.

Pepin:1993:AES

[869] Pierre Pepin. Application of empirical size-dependent models of larval fish vital rates to the study of production: Accuracy and association with adult stock dynamics in a comparison among species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):53–59, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-006.

Ramakrishna:1993:CMI

[870] N. R. Ramakrishna, M. D. B. Burt, and B. M. MacKinnon. Cell-mediated immune response of rainbow trout (Oncorhynchus mykiss) to larval Pseudoterranova decipiens (Nematoda; Ascaridoidea) following sensitization to live sealworm, sealworm extract, and nonhomologous extracts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):60–65, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-007.

Johannes:1993:PAC

[871] Mark R. S. Johannes. Prey aggregation is correlated with increased predation pressure in lake fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):66–73, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-008.

Deegan:1993:NET

[872] Linda A. Deegan. Nutrient and energy transport between estuaries and coastal marine ecosystems by fish migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):74–79, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-009.

Wirgin:1993:MDV

[873] Isaac I. Wirgin, Tun-Liang Ong, Lorraine Maceda, John R. Waldman, David Moore, and Simon Courtenay. Mitochondrial DNA variation in striped bass (Morone saxatilis) from Canadian rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):80–87, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-010.

Mather:1993:UGM

[874] Martha E. Mather and Roy A. Stein. Using growth/mortality tradeoffs to explore a crayfish species replacement in stream riffles and pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):88–96, January 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-011.

Madenjian:1993:APL

[875] Charles P. Madenjian, Stephen R. Carpenter, Gary W. Eck, and Michael A. Miller. Accumulation of PCBs by lake trout (Salvelinus namaycush): an individual-based model approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):97–109, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-012.

Downing:1993:PFP

[876] John A. Downing and Céline Plante. Production of fish populations in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):110–120, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-013.

Gensemer:1993:PTM

[877] Robert W. Gensemer, Ralph E. H. Smith, Hamish C. Duthie, and Sherry L. Schiff. pH tolerance and metal toxicity in populations of the planktonic diatom Asterionella: Influences of synthetic and natural dissolved organic carbon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1): 121–132, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-014.

Rice:1993:GRV

[878] James A. Rice, Thomas J. Miller, Kenneth A. Rose, Larry B. Crowder, Elizabeth A. Marschall, Anett S. Trebitz, and Donald L. DeAngelis. Growth rate variation and larval survival: Inferences from an individualbased size-dependent predation model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):133–142, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-015.

Jackson:1993:SPR

[879] Leland J. Jackson, Jacob Kalff, and Joseph B. Rasnnussen. Sediment pH and redox potential affect the bioavailability of Al, Cu, Fe, Mn, and Zn to rooted aquatic macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (1):143–148, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-016.

Havens:1993:PFWa

[880] Karl E. Havens. Pelagic food web structure in Adirondack mountain, USA, lakes of varying acidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (1):149–155, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-017.

Palace:1993:IAA

[881] V. P. Palace, H. S. Majewski, and J. F. Klaverkamp. Interactions among antioxidant defenses in liver of rainbow trout (Oncorhynchus mykiss) exposed to cadmium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1): 156–162, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-018.

McGurk:1993:RBP

[882] M. D. McGurk, A. J. Paul, K. O. Coyle, D. A. Ziemann, and L. J. Haldorson. Relationships between prey concentration and growth, condition, and mortality of Pacific herring, Clupea pallasi, larvae in an Alaskan subarctic embayment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1): 163–180, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-019.

Yoshimoto:1993:CDV

[883] Stacey S. Yoshimoto and Raymond P. Clarke. Comparing dynamic versions of the Schaefer and Fox production models and their application to lobster fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1): 181–189, January 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-020.

Oluyedun:1993:SPB

[884] O. A. Oluyedun, S. O. Ajayi, G. W. VanLoon, and P. Sly. Sedimentary phosphorus in the Bay of Quinte, Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):190–197, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-021.

Audet:1993:BME

[885] Céline Audet and Chris M. Wood. Branchial morphological and endocrine responses of rainbow trout (Oncorhynchus mykiss) to a long-term sublethal acid exposure in which acclimation did not occur. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):198–209, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-022.

Vuorinen:1993:GMD

[886] J. A. Vuorinen, R. A. Bodaly, J. D. Reist, L. Bernatchez, and J. J. Dodson. Genetic and morphological differentiation between dwarf and normal size forms of lake whitefish (Coregonus clupeafotmis) in Como Lake, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):210–216, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-023.

Houston:1993:CEO

[887] Arthur Houston, Shelley Blahut, Ajmal Murad, and Preetha Amirtharaj. Changes in erythron organization during prolonged cadmium exposure: an indicator of heavy metal stress? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(1):217–222, January 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-024.

Bohlin:1993:OTS

[888] Torgny Bohlin, Claes Dellefors, and Ulo Faremo. Optimal time and size for smolt migration in wild sea trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):224–232, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-025.

Gregory:1993:SPB

[889] R. S. Gregory and T. G. Northcote. Surface, planktonic, and benthic foraging by juvenile Chinook salmon (Oncorhynchus tshawytscha) in turbid laboratory conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2): 233–240, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-026.

Gregory:1993:ETP

[890] Robert S. Gregory. Effect of turbidity on the predator avoidance behaviour of juvenile Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):241–246, February 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-027.

Sorum:1993:CPP

[891] Henning Sørum, Egil Myhr, Betty M. Zwicker, and Atle Lillehaug. Comparison by plasmid profiling of Vibrio salmonicida strains isolated from diseased fish from different North European and Canadian coastal areas of the Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2): 247–250, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-028.

Brodziak:1993:ESC

[892] Jon Brodziak. An extension of stock composition analysis to include marking data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):251–257, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-029.

Wright:1993:RPR

[893] Richard F. Wright, Erik Lotse, and Arne Semb. RAIN project: Results after 8 years of experimentally reduced acid deposition to a whole catchment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):258–268, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-030.

Louchouarn:1993:GMT

[894] Patrick Louchouarn, Marc Lucotte, Alfonso Mucci, and Pierre Pichet. Geochemistry of mercury in two hydroelectric reservoirs in Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):269–281, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-031.

delGiorgio:1993:BBP

[895] Paul A. del Giorgio and Robert H. Peters. Balance between phytoplankton production and plankton respiration in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):282–289, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-032.

Ishida:1993:RCA

[896] Yukimasa Ishida, Soto o Ito, Masahide Kaeriyama, Skip McKinnell, and Kazuya Nagasawa. Recent changes in age and size of chum salmon (Oncorhynchus keta) in the North Pacific Ocean and possible causes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):290–295, February 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-033.

Conley:1993:PRS

[897] Daniel J. Conley and Claire L. Schelske. Potential role of sponge spicules in influencing the silicon biogeochemistry of Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):296–302, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-034.

Wildish:1993:SCD

[898] D. J. Wildish, P. D. Keizer, A. J. Wilson, and J. L. Martin. Seasonal changes of dissolved oxygen and plant nutrients in seawater near salmonid net pens in the macrotidal Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):303–311, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-035.

Dolloff:1993:PRO

[899] C. Andrew Dolloff. Predation by river otters (Lutra canadensis) on juvenile Coho salmon (Oncorhynchus kisutch) and Dolly Varden (Salvelinus malma) in Southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (2):312–315, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-036.

Anderson:1993:QDL

[900] J. Stewart Anderson, Santosh P. Lall, Derek M. Anderson, and Mary A. McNiven. Quantitative dietary lysine requirement of Atlantic salmon (Salmo salar) fingerlings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (2):316–322, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-037.

Gillis:1993:MDF

[901] Darren M. Gillis, Randall M. Peterman, and Albert V. Tyler. Movement dynamics in a fishery: Application of the ideal free distribution to spatial allocation of effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (2):323–333, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-038.

Sameoto:1993:CAO

[902] D. Sameoto, N. Cochrane, and A. Herman. Convergence of acoustic, optical, and net-catch estimates of euphausiid abundance: Use of artificial light to reduce net. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (2):334–346, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-039.

Hankin:1993:EIA

[903] David G. Hankin, Jay W. Nicholas, and Timothy W. Downey. Evidence for inheritance of age of maturity in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):347–358, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-040.

Herman:1993:IMS

[904] A. W. Herman, D. F. Knox, J. Conrad, and M. R. Mitchell. Instruments for measuring subice algal profiles and productivity in situ. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):359–369, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-041.

Bird:1993:PGS

[905] David F. Bird and Jacob Kalff. Protozoan grazing and the size– activity structure of limnetic bacterial communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):370–380, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-042.

Dumbauld:1993:UOS

[906] Brett R. Dumbauld, David A. Armstrong, and Trent L. McDonald. Use of oyster shell to enhance intertidal habitat and mitigate loss of Dungeness crab (Cancer magister) caused by dredging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):381–390, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-043.

MacFarlane:1993:LDR

[907] R. Bruce MacFarlane, Elizabeth C. Norton, and Michael J. Bowers. Lipid dynamics in relation to the annual reproductive cycle in yellowtail rockfish (Sebastes flavidus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2): 391–401, February 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-044.

Cohen:1993:AIA

[908] Yosef Cohen, Paul Radomski, and Ron Moen. Assessing the interdependence of assemblages from Rainy Lake fisheries data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):402–409, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-045.

Blais:1993:DPC

[909] Jules M. Blais and Edward J. Maly. Differential predation by Chaoborus americanus on males and females of two species of Diaptomus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):410–415, February 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-046.

Jamieson:1993:MSD

[910] G. S. Jamieson and A. Phillips. Megalopal spatial distribution and stock separation in Dungeness crab (Cancer magister). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):416–429, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-047.

Berthiaume:1993:MLD

[911] L. Berthiaume, J. Heppell, M. Désy, L. Leblanc, R. Lallier, R. Bailey, and J.-D. Dutil. Manifestation of lymphocystis disease in American plaice (Hippoglossoides platessoides) exposed to low salinities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):430–434, February 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-048.

Heath:1993:GEI

[912] Daniel D. Heath, Nicholas J. Bernier, John W. Heath, and George K. Iwama. Genetic, environmental, and interaction effects on growth and stress response of Chinook salmon (Oncorhynchus tshawytscha) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):435–442, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-049.

Tousignant:1993:LBT

[913] M. E. Tousignant, J. L. Boisvert, and A. Chalifour. Loss of Bacillus thuringiensis var. israelensis larvicidal activity and its distribution in benthic substrates and hyporheic zone of streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):443–451, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-050.

U:1993:LCLa

[914] F. H. U. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):452–453, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-051.

U:1993:CRPa

[915] F. H. U. Corrections to recent publications/Corrections sur récentes publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(2):454, February 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-052.

Parsons:1993:DGR

[916] G. Jay Parsons, Shawn M. C. Robinson, John C. Roff, and Michael J. Dadswell. Daily growth rates as indicated by valve ridges in postlarval giant scallop (Placopecten magellanicus) (Bivalvia: Pectinidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):456–464, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-053.

Hoogenboezem:1993:PRS

[917] Wim Hoogenboezem, Eddy H. R. R. Lammens, Peter J. MacGillavry, and Ferdinand A. Sibbing. Prey retention and sieve adjustment in filterfeeding bream (Abramis brama) (cyprinidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):465–471, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-054.

Hoogenboezem:1993:IMF

[918] Wim Hoogenboezem and Jos G. M. van den Boogaart. Importance of mucus in filter-feeding of bream (Abramis brama). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):472–479, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-055.

Johnsson:1993:HDR

[919] Jörgen I. Johnsson, W. Craig Clarke, and Ruth E. Withler. Hybridization with domesticated rainbow trout (Oncorhynchus mykiss) reduces seasonal variation in growth of steelhead trout (O. mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):480–487, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-056.

Chauvet:1993:BIC

[920] E. Chauvet, N. Giani, and M. O. Gessner. Breakdown and invertebrate colonization of leaf litter in two contrasting streams, significance of oligochaetes in a large river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (3):488–495, March 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-057.

Aalto:1993:WDS

[921] S. K. Aalto and G. E. Newsome. Winds and the demic structure of a population of yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):496–501, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-058.

Foster:1993:ENR

[922] A. R. Foster, D. F. Houlihan, and S. I. Hall. Effects of nutritional regime on correlates of growth rate in juvenile Atlantic cod (Gadus morhua): Comparison of morphological and biochemical measurements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):502–512, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-059.

Shardlow:1993:CAD

[923] Thomas F. Shardlow. Components analysis of a density-dependent catchability coefficient in a salmon hook and line fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):513–520, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-060.

Parkman:1993:MMS

[924] Helena Parkman and Markus Meili. Mercury in macroinvertebrates from Swedish forest lakes: Influence of lake type, habitat, life cycle, and food quality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):521–534, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-061.

Dorazio:1993:PST

[925] Robert M. Dorazio. Prerelease stratification in tag-recovery models with time dependence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):535–541, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-062.

Martoja:1993:LCD

[926] Micheline Martoja and Christiane Marcaillou. Localisation cytologique du cuivre et de quelques autres métaux dans la glande digestive de la seiche, Sepia officinalis L. (mollusque céphalopode). (French) [Cytological localization of copper and some other metals in the digestive gland of cuttlefish, Sepia officinalis L. (cephalopod mollusk)]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):542–550, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-063.

Neilson:1993:RBA

[927] John D. Neilson, John F. Kearney, Peter Perley, and Harry Sampson. Reproductive biology of Atlantic halibut (Hippoglossus hippoglossus) in Canadian waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):551–563, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-064.

McGarvey:1993:SPA

[928] Richard McGarvey, Fredric M. Serchuk, and Ian A. McLaren. Spatial and parent-age analysis of stock–recruitment in the Georges Bank sea scallop (Placopecten magellanicus) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):564–574, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-065.

Bulger:1993:BTS

[929] Arthur J. Bulger, Leif Lien, B. J. Cosby, and Arne Henriksen. Brown trout (Salmo trutta) status and chemistry from the Norwegian Thousand Lake Survey: Statistical analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):575–585, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-066.

Martens:1993:SSP

[930] Dennis W. Martens and James A. Servizi. Suspended sediment particles inside gills and spleens of juvenile Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):586–590, March 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-067.

Whyte:1993:NCP

[931] John N. C. Whyte, Richard J. Beamish, Norma G. Ginther, and Chrys-Ellen Neville. Nutritional condition of the Pacific lamprey (Lampetra tridentata) deprived of food for periods of up to two years. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):591–599, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-068.

Koenings:1993:SAS

[932] Jeffery P. Koenings, Harold J. Geiger, and James J. Hasbrouck. Smolt-to-adult survival patterns of sockeye salmon (Oncorhynchus nerka): Effects of smolt length and geographic latitude when entering the sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):600–611, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-069.

Hannah:1993:IEV

[933] Robert W. Hannah. Influence of environmental variation and spawning stock levels on recruitment of ocean shrimp (Pandalus jordani). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):612–622, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-070.

Lamontagne:1993:ECD

[934] S. Lamontagne and J. B. Rasmussen. Estimating crayfish density in lakes using quadrats: Maximizing precision and efficiency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):623–626, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-071.

Smith:1993:IAU

[935] Eric P. Smith, David R. Orvos, and John Cairns, Jr. Impact assessment using the before-after-control-impact (BACI) model: Concerns and comments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):627–637, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-072.

Rodriguez:1993:FSC

[936] Marco A. Rodríguez, Pierre Magnan, and Sylvain Lacasse. Fish species composition and lake abiotic variables in relation to the abundance and size structure of cladoceran zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):638–647, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-073.

Ashton:1993:FAC

[937] H. J. Ashton, D. O. Farkvam, and B. E. March. Fatty acid composition of lipids in the eggs and alevins from wild and cultured Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (3):648–655, March 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-074.

McCormick:1993:MNG

[938] Stephen D. McCormick. Methods for nonlethal gill biopsy and measurement of Na^{+}, K^{+}-ATPase activity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):656–658, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-075.

Guenette:1993:MDV

[939] S. Guénette, R. Fortin, and E. Rassart. Mitochondrial DNA variation in lake sturgeon (Acipenser fulvescens) from the St. Lawrence River and James Bay drainage basins in Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):659–664, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-076.

Triska:1993:DIN

[940] Frank J. Triska, Catherine M. Pringle, Gary W. Zellweger, John H. Duff, and Ronald J. Avanzino. Dissolved inorganic nitrogen composition, transformation, retention, and transport in naturally phosphaterich and phosphate-poor tropical streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):665–675, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-077.

Holtby:1993:MEA

[941] L. Blair Holtby, Douglas P. Swain, and G. Michael Allan. Mirrorelicited agonistic behaviour and body morphology as predictors of dominance status in juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(3):676–684, March 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-078.

Hornstrom:1993:PCP

[942] Einar Hörnström, Christina Ekström, Elisabeth Fröberg, and Jörgen Ek. Plankton and chemical–physical development in six Swedish West Coast lakes under acidic and limed conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):688–702, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-079.

Young:1993:IGS

[943] Paciencia S. Young and Joseph J. Cech, Jr. Improved growth, swimming performance, and muscular development in exercise-conditioned youngof-the-year striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):703–707, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-080.

Cronin:1993:MDV

[944] Matthew A. Cronin, William J. Spearman, Richard L. Wilmot, John C. Patton, and John W. Bickham. Mitochondrial DNA variation in Chinook (Oncorhynchus tshawytscha) and chum salmon (O. keta) detected by restriction enzyme analysis of polymerase chain reaction (PCR) products. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):708–715, April 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-081.

vandenBerg:1993:SZR

[945] Coen van den Berg, Jos G. M. van den Boogaart, Ferdinand A. Sibbing, Eddy H. R. R. Lammens, and Jan W. M. Osse. Shape of zooplankton and retention in filter-feeding: a quantitative comparison between industrial sieves and the branchial sieves of common bream (Abramis brama) and white bream (Blicca bjoerkna). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):716–724, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-082.

Swain:1993:DDG

[946] D. P. Swain and E. J. Wade. Density-dependent geographic distribution of Atlantic cod (Gadus morhua) in the Southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):725–733, April 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-083.

Sinclair:1993:PRC

[947] A. F. Sinclair. Partial recruitment considerations in setting catch quotas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):734–742, April 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-084.

Wahl:1993:OCP

[948] Christina M. Wahl, Edward L. Mills, William N. McFarland, and Joseph S. DeGisi. Ontogenetic changes in prey selection and visual acuity of the yellow perch, Perca flavescens. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):743–749, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-085.

Stone:1993:ABO

[949] Robert P. Stone, Charles E. O’Clair, and Thomas C. Shirley. Aggregating behavior of ovigerous female red king crab, Paralithodes camtschaticus, in Auke Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (4):750–758, April 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-086.

Salonius:1993:EER

[950] Kira Salonius and George K. Iwama. Effects of early rearing environment on stress response, immune function, and disease resistance in juvenile Coho (Oncorhynchus kisutch) and Chinook salmon (O. tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):759–766, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-087.

Dixon:1993:VCP

[951] B. Dixon, W. Kimmins, and B. Pohajdak. Variation in colour of Pseudoterranova decipiens (Nematoda; Anisakidae) larvae correlates with haemoglobin concentration in the pseudocoelomic fluid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):767–771, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-088.

Kenchington:1993:EFF

[952] T. J. Kenchington. Estimation of the functional form of selectivity by baited hooks: a graphical approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):772–780, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-089.

Miner:1993:IIT

[953] Jeffrey G. Miner and Roy A. Stein. Interactive influence of turbidity and light on larval bluegill (Lepomis macrochirus) foraging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):781–788, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-090.

MacKinnon:1993:HRA

[954] Barbara M. MacKinnon. Host response of Atlantic salmon (Salmo salar) to infection by sea lice (Caligus elongatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):789–792, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-091.

MacKinnon:1993:RCL

[955] Barbara M. MacKinnon. Response of the copepodid larvae of Caligus elongatus to light, and the ultrastructure of the eyes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):793–799, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-092.

Rodriguez:1993:CSL

[956] Marco A. Rodríguez and Pierre Magnan. Community structure of lacustrine macrobenthos: Do taxon-based and size-based approaches yield similar insights? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):800–815, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f93-
093.

Eilers:1993:MIC

[957] J. M. Eilers, D. H. Landers, A. D. Newell, M. E. Mitch, M. Morrison, and J. Ford. Major ion chemistry of lakes on the Kenai Peninsula, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):816–826, April 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-094.

Bence:1993:IAS

[958] James R. Bence, Ana Gordoa, and Joseph E. Hightower. Influence of age-selective surveys on the reliability of stock synthesis assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):827–840, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-095.

Davies:1993:IOS

[959] Ronald W. Davies and V. Kalarani. Influence of overwintering stresses on respiration throughout the life history of the freshwater leech, Nephelopsis obscura. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):841–845, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-096.

Servizi:1993:EBB

[960] J. A. Servizi, R. W. Gordon, D. W. Martens, W. L. Lockhart, D. A. Metner, I. H. Rogers, J. R. McBride, and R. J. Norstrom. Effects of biotreated bleached kraft mill effluent on fingerling Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (4):846–857, April 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-097.

Trites:1993:UCR

[961] Andrew W. Trites and Anne E. York. Unexpected changes in reproductive rates and mean age at first birth during the decline of the Pribilof northern fur seal (Callorhinus ursinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):858–864, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-098.

Rada:1993:WLB

[962] Ronald G. Rada, David E. Powell, and James G. Wiener. Whole-lake burdens and spatial distribution of mercury in surficial sediments in Wisconsin seepage lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (4):865–873, April 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-099.

Hilborn:1993:CTI

[963] Ray Hilborn, Ellen K. Pikitch, and Robert C. Francis. Current trends in including risk and uncertainty in stock assessment and harvest decisions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):874–880, April 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-100.

U:1993:LCLb

[964] F. H. U. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(4):881–884, April 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-101.

Reid:1993:IAT

[965] D. G. Reid and E. J. Simmonds. Image analysis techniques for the study of fish school structure from acoustic survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):886–893, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-102.

Starr:1993:ECG

[966] Michel Starr, John H. Himmelman, and Jean-Claude Therriault. Environmental control of green sea urchin, Strongylocentrotus droebachiensis, spawning in the St. Lawrence Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):894–901, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-103.

Simonin:1993:SBB

[967] Howard A. Simonin, Walter A. Kretser, Dale W. Bath, Mark Olson, and James Gallagher. In situ bioassays of brook trout (Salvelinus fontinalis) and blacknose dace (Rhinichthys atratulus) in Adirondack streams affected by episodic acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):902–912, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f93-104.

Otteraa:1993:FGS

[968] Håkon Otterå. Feeding, growth, and survival of Atlantic cod (Gadus morhua) larvae reared in replicate plastic enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):913–924, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-105.

Helminen:1993:YCF

[969] Harri Helminen, Heikki Auvinen, Arto Hirvonen, Jouko Sarvala, and Jorma Toivonen. Year-class fluctuations of vendace (Coregonus albula) in Lake Pyhäjärvi, Southwest Finland, during 1971–90. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):925–931, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-106.

Hsu:1993:ICE

[970] Sheau-Yu Hsu and Frederick William Goetz. Inhibition of chorion expansion and preservation of fertility in goldfish (Carassius auratus) eggs by protease inhibitors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):932–935, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-107.

Wehr:1993:EEM

[971] John D. Wehr. Effects of experimental manipulations of light and phosphorus supply on competition among picoplankton and nanoplankton in an oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):936–945, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-108.

Takamura:1993:NAE

[972] Noriko Takamura, Ja-Le Li, He-Quan Yang, Xue-Bao Zhu, and Taizo Miura. A novel approach to evaluate feeding by mixed cyprinid species in a Chinese integrated fish culture pond using measurements of chlorophyll derivatives and photosynthesis in gut contents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):946–952, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-109.

StLouiss:1993:ECC

[973] Vincent L. St. Louiss. Element concentrations in chironomids and their abundance in the littoral zone of acidified lakes in Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):953–963, May 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-110.

Tanasichuk:1993:CSL

[974] R. W. Tanasichuk, A. H. Kristofferson, and D. V. Gillman. Comparison of some life history characteristics of Pacific herring (Clupea pallasi) from the Canadian Pacific Ocean and Beaufort Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):964–971, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-111.

Ramlal:1993:SMM

[975] P. S. Ramlal, C. A. Kelly, J. W. M. Rudd, and A. Furutani. Sites of methyl mercury production in remote Canadian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):972–979, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-112.

Bodaly:1993:MCF

[976] R. A. Bodaly, J. W. M. Rudd, R. J. P. Fudge, and C. A. Kelly. Mercury concentrations in fish related to size of remote Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):980–987, May 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-113.

Justesen:1993:PTC

[977] Nils Petter Berg Justesen, Torbjørn Dall-Larsen, and Leiv Klungsøyr. Proton transport and chloride cells in the gill of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):988–995, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-114.

Wang:1993:RBG

[978] Shiao Y. Wang, Judith L. Lum, Mark G. Carls, and Stanley D. Rice. Relationship between growth and total nucleic acids in juvenile pink salmon, Oncorhynchus gorbuscha, fed crude oil contaminated food. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):996–1001, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-115.

Beamish:1993:PSP

[979] Richard J. Beamish and Daniel R. Bouillon. Pacific salmon production trends in relation to climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):1002–1016, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-116.

Bunt:1993:PRP

[980] Christopher M. Bunt, Hugh J. Maclsaac, and W. Gary Sprules. Pumping rates and projected filtering impacts of juvenile zebra mussels (Dreissena polymorpha) in Western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1017–1022, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-117.

James:1993:SRR

[981] William F. James and John W. Barko. Sediment resuspension, redeposition, and focusing in a small dimictic reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1023–1028, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-118.

Welch:1993:SDS

[982] David W. Welch and Yukimasa Ishida. On the statistical distribution of salmon in the sea: Application of the negative binomial distribution, and the influence of sampling effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1029–1038, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-119.

Ennevor:1993:ULE

[983] Bridget C. Ennevor and R. M. Beames. Use of lanthanide elements to mass mark juvenile salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1039–1044, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-120.

Lebo:1993:SVN

[984] Martin E. Lebo, John E. Reuter, Charles R. Goldman, Cathryn L. Rhodes, Nancy Vucinich, and Dan Mosely. Spatial variations in nutrient and particulate matter concentrations in Pyramid Lake, Nevada, USA, during a dry period. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):1045–1054, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-121.

Tyler:1993:EWV

[985] Jeffrey A. Tyler. Effects of water velocity, group size, and prey availability on the stream-drift capture efficiency of blacknose dace, Rhinicthys atratulus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1055–1061, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-122.

Campana:1993:SDU

[986] Steven E. Campana and John M. Casselman. Stock discrimination using otolith shape analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):1062–1083, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-123.

Kessler:1993:MST

[987] Richard K. Kessler and James H. Thorp. Microhabitat segregation of the threatened spotted darter (Etheostoma maculatum) and closely related orangefin darter (E. bellum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (5):1084–1091, May 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-124.

Newbound:1993:CRD

[988] G. C. Newbound, D. J. Speare, K. L. Hammell, M. L. Kent, V. E. Ostland, and G. S. Traxler. Chehalis River disease: a unique gill disease of salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1092–1100, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f93-
125.

Brezonik:1993:EAL

[989] P. L. Brezonik, J. G. Eaton, T. M. Frost, P. J. Garrison, T. K. Kratz, C. E. Mach, J. H. McCormick, J. A. Perry, W. A. Rose, C. J. Sampson, B. C. L. Shelley, W. A. Swenson, and K. E. Webster. Experimental acidification of Little Rock Lake, Wisconsin: Chemical and biological changes over the pH range 6.1 to 4.7. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1101–1121, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-126.

U:1993:LCLc

[990] F. H. U. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(5):1122–1126, May 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-127.

Bohlin:1993:TSR

[991] Torgny Bohlin, Claes Dellefors, and Ulo Faremo. Timing of sea-run brown trout (Salmo trutta) smolt migration: Effects of climatic variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1132–1136, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-128.

Sakai:1993:IPF

[992] Masahiro Sakai, M. K. Soliman, T. Yoshida, and Masanori Kobayashi. Identification of pathogenic fish bacteria using the API ZYM system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1137–1141, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-129.

Soma:1993:EAS

[993] Yuko Soma, Takashi Imaizumi, Kei ichi Yagi, and Sei ichi Kasuga. Estimation of algal succession in lake water using HPLC analysis of pigments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1142–1146, June 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-130.

Boulding:1993:GVO

[994] Elizabeth C. Boulding, John D. C. Boom, and Andrew T. Beckenbach. Genetic variation in one bottlenecked and two wild populations of the Japanese scallop (Patinopecten yessoensis): Empirical parameter estimates from coding regions of mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1147–1157, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-131.

Fjeld:1993:UPA

[995] Eirik Fjeld and Sigurd Rognerud. Use of path analysis to investigate mercury accumulation in brown trout (Salmo trutta) in Norway and the influence of environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1158–1167, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-132.

Unwin:1993:HSP

[996] Martin J. Unwin and Thomas P. Quinn. Homing and straying patterns of Chinook salmon (Oncorhynchus tshawytscha) from a New Zealand hatchery: Spatial distribution of strays and effects of release date. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1168–1175, June 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-133.

Ostrofsky:1993:ETL

[997] M. L. Ostrofsky. Effect of tannins on leaf processing and conditioning rates in aquatic ecosystems: an empirical approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1176–1180, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-134.

Schwarz:1993:ESS

[998] Carl J. Schwarz, Richard E. Bailey, James R. Irvine, and Frank C. Dalziel. Estimating salmon spawning escapement using capture– recapture methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6): 1181–1197, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f93-135. See erratum [2311].

Fausch:1993:EAM

[999] Kurt D. Fausch. Experimental analysis of microhabitat selection by juvenile steelhead (Oncorhynchus mykiss) and Coho salmon (O. kisutch) in a British Columbia stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (6):1198–1207, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-136.

Skalski:1993:FLH

[1000] John R. Skalski, Annette Hoffmann, Bruce H. Ransom, and Tracey W. Steig. Fixed-location hydroacoustic monitoring designs for estimating fish passage using stratified random and systematic sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1208–1221, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-137.

Kenchington:1993:UBG

[1001] E. Kenchington, K. S. Naidu, D. L. Roddick, D. I. Cook, and E. Zouros. Use of biochemical genetic markers to discriminate between adductor muscles of the sea scallop (Placopecten magellanicus) and the Iceland scallop (Chlamys islandica). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (6):1222–1228, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-138.

Weisberg:1993:UHP

[1002] Sanford Weisberg. Using hard-part increment data to estimate age and environmental effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6): 1229–1237, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-139.

Locke:1993:EEA

[1003] Andrea Locke and W. Gary Sprules. Effects of experimental acidification on zooplankton population and community dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1238–1247, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-140.

Bates:1993:EAN

[1004] Stephen S. Bates, Jean Worms, and John C. Smith. Effects of ammonium and nitrate on growth and domoic acid production by Nitzschia pungens in batch culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (6):1248–1254, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-141.

Swain:1993:ADD

[1005] Douglas P. Swain. Age- and density-dependent bathymetric pattern of Atlantic cod (Gadus morhua) in the Southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1255–1264, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-142.

Dempson:1993:OSL

[1006] J. B. Dempson and T. R. Porter. Occurrence of sea lamprey, Petromyzon marinus, in a Newfoundland river, with additional records from the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6): 1265–1269, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-143.

Marks:1993:IEN

[1007] Jane Claire Marks and Rex L. Lowe. Interactive effects of nutrient availability and light levels on the periphyton composition of a large oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1270– 1278, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-144.

Mather:1993:DIE

[1008] Martha E. Mather and Roy A. Stein. Direct and indirect effects of fish predation on the replacement of a native crayfish by an invading congener. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1279–1288, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-145.

Knight:1993:CPF

[1009] Roger L. Knight and Bruce Vondracek. Changes in prey fish populations in Western Lake Erie, 1969–88, as related to walleye, Stizostedion vitreum, predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6): 1289–1298, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-146.

Beamish:1993:ESD

[1010] F. William H. Beamish. Environmental sex determination in southern brook lamprey, Ichthyomyzon gagei. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1299–1307, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-147.

Thiebaux:1993:SBS

[1011] M. L. Thiebaux and L. M. Dickie. Structure of the body-size spectrum of the biomass in aquatic ecosystems: a consequence of allometry in predator–prey interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (6):1308–1317, June 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-148.

Coutant:1993:SMS

[1012] Charles C. Coutant and C. H. Chen. Strontium microstructure in scales of freshwater and estuarine striped bass (Morone saxatilis) detected by laser ablation mass spectrometry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1318–1323, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-149.

Plante:1993:RSP

[1013] Céline Plante and John A. Downing. Relationship of salmonine production to lake trophic status and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1324–1328, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-150.

Wood:1993:DEM

[1014] Chris C. Wood, N. Brent Hargreaves, Dennis T. Rutherford, and Brian T. Emmett. Downstream and early marine migratory behaviour of sockeye salmon (Oncorhynchus nerka) smolts entering Barkley Sound, Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1329–1337, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-151.

McGuire:1993:FRV

[1015] Shawn McGuire and David J. Currie. Factors related to the variation in mixing depth among meromictic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1338–1342, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-152.

U:1993:LCLd

[1016] F. H. U. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(6):1343–1345, June 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-153.

Hickie:1993:TTM

[1017] B. E. Hickie, N. J. Hutchinson, D. G. Dixon, and P. V. Hodson. Toxicity of trace metal mixtures to alevin rainbow trout (Oncorhynchus mykiss) and larval fathead minnow (Pimephales promelas) in soft, acidic water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1348–1355, July 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-154.

Welsh:1993:EPD

[1018] P. G. Welsh, J. F. Skidmore, D. J. Spry, D. G. Dixon, P. V. Hodson, N. J. Hutchinson, and B. E. Hickie. Effect of pH and dissolved organic carbon on the toxicity of copper to larval fathead minnow (Pimephales promelas) in natural lake waters of low alkalinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1356–1362, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-155.

Beck:1993:SVD

[1019] Gregor Gilpin Beck, Michael O. Hammill, and Thomas G. Smith. Seasonal variation in the diet of harp seals (Phoca groenlandica) from the Gulf of St. Lawrence and Western Hudson Strait. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1363–1371, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-156.

Beck:1993:EBC

[1020] Gregor Gilpin Beck, Thomas G. Smith, and Michael O. Hammill. Evaluation of body condition in the Northwest Atlantic harp seal (Phoca groenlandica). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1372–1381, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-157.

Hartmann:1993:PCS

[1021] Hans J. Hartmann, Hassan Taleb, Lotfi Aleya, and Nicole Lair. Predation on ciliates by the suspension-feeding calanoid copepod Acanthodiaptomus denticornis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (7):1382–1393, July 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-158.

Garcia-Berthou:1993:MAC

[1022] Emili García-Berthou and Ramon Moreno-Amich. Multivariate analysis of covariance in morphometric studies of the reproductive cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1394–1399, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-159.

Grahl-Nielsen:1993:DBD

[1023] Otto Grahl-Nielsen, Olav Mjaavatten, and Einar Tvedt. Distinguishing between different populations of harp seal (Phoca groenlandica) by chemometry of the fatty acid profiles in the jaw bone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1400–1404, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-160.

Miller:1993:MTO

[1024] Michael A. Miller. Maternal transfer of organochlorine compounds in salmonines to their eggs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (7):1405–1413, July 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-161.

Quinn:1993:VLH

[1025] Thomas P. Quinn and Martin J. Unwin. Variation in life history patterns among New Zealand Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1414–1421, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-162.

Hudon:1993:WIA

[1026] Christiane Hudon and Pierre Fradette. Wind-induced advection of larval decapods into Baie de Plaisance (Îles de la Madeleine, Québec). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1422–1434, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-163.

Treble:1993:DEM

[1027] R. J. Treble, R. W. Day, and T. J. Quinn II. Detection and effects on mortality estimates of changes in tag loss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1435–1441, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-164.

Trippel:1993:MTL

[1028] Edward A. Trippel and F. William H. Beamish. Multiple trophic level structuring in Salvelinus–Coregonus assemblages in boreal forest lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1442–1455, July 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-165.

StJohn:1993:HRP

[1029] M. A. St. John, S. G. Marinone, J. Stronach, P. J. Harrison, J. Fyfe, and R. J. Beamish. A horizontally resolving physical–biological model of nitrate concentration and primary productivity in the Strait of Georgia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1456–1466, July 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-166.

Dawe:1993:CRI

[1030] Earl G. Dawe, John M. Hoenig, and Xucai Xu. Change-in-ratio and index-removal methods for population assessment and their application to snow crab (Chionoecetes opilio). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1467–1476, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-167.

Kortelainen:1993:CTO

[1031] Pirkko Kortelainen. Content of total organic carbon in Finnish lakes and its relationship to catchment characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1477–1483, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-168.

DiDonato:1993:SRA

[1032] Guy T. DiDonato and David M. Lodge. Species replacements among Orconectes crayfishes in Wisconsin lakes: The role of predation by fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1484–1488, July 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-169.

Goldman:1993:DIS

[1033] Charles R. Goldman, Alan D. Jassby, and Scott H. Hackley. Decadal, interannual, and seasonal variability in enrichment bioassays at Lake Tahoe, California–Nevada, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1489–1496, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-170.

Freda:1993:TAB

[1034] Joseph Freda and D. Gordon McDonald. Toxicity of amphibian breeding ponds in the Sudbury region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (7):1497–1503, July 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-171.

Stevens:1993:MMP

[1035] Bradley G. Stevens, William E. Donaldson, Jan. A. Haaga, and J. Eric Munk. Morphometry and maturity of paired Tanner crabs, Chionoecetes bairdi, from shallow- and deepwater environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1504–1516, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-172.

Chmielewski:1993:CEB

[1036] C. M. Chmielewski and R. J. Hall. Changes in the emergence of blackflies (Diptera: Simuliidae) over 50 years from Algonquin Park streams: Is acidification the cause? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (7):1517–1529, July 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-173.

Schweigert:1993:EMR

[1037] Jacob F. Schweigert and Carl J. Schwarz. Estimating migration rates for Pacific herring (Clupea pallasi) using tag-recovery data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1530–1540, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-174.

Fahrig:1993:PSS

[1038] Lenore Fahrig, George R. Lilly, and Daniel S. Miller. Predator stomachs as sampling tools for prey distribution: Atlantic cod (Gadus morhua) and capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (7):1541–1547, July 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-175.

Santucci:1993:FIS

[1039] Victor J. Santucci, Jr. and David H. Wahl. Factors influencing survival and growth of stocked walleye (Stizostedion vitreum) in a centrarchiddominated impoundment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (7):1548–1558, July 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-176.

Stevens:1993:NMP

[1040] T. A. Stevens, R. E. Withler, S. H. Goh, and T. D. Beacham. A new multilocus probe for DNA fingerprinting in Chinook salmon (Oncorhynchus tshawytscha), and comparisons with a single-locus probe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1559–1567, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-177.

U:1993:LCLe

[1041] F. H. U. Letters and comments/letters et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(7):1568–1573, July 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-178.

Myers:1993:DDJ

[1042] Ransom A. Myers and Noel G. Cadigan. Density-dependent juvenile mortality in marine demersal fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (8):1576–1590, August 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-179.

Myers:1993:JNM

[1043] Ransom A. Myers and Noel G. Cadigan. Is juvenile natural mortality in marine demersal fish variable? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1591–1598, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-180.

Myers:1993:SRA

[1044] Ransom A. Myers, Kenneth F. Drinkwater, Nicholas J. Barrowman, and James W. Baird. Salinity and recruitment of Atlantic cod (Gadus morhua) in the Newfoundland region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1599–1609, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-181.

Gottgens:1993:QIL

[1045] J. F. Gottgens and T. L. Crisman. Quantitative impacts of lake-level stabilization on material transfer between water and sediment in Newnans Lake, Florida. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8): 1610–1616, August 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-182.

Rijnsdorp:1993:RBJ

[1046] A. D. Rijnsdorp. Relationship between juvenile growth and the onset of sexual maturity of female North Sea plaice, Pleuronectes platessa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1617–1631, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-183.

Nicieza:1993:RAS

[1047] A. G. Nicieza and F. Braña. Relationships among smolt size, marine growth, and sea age at maturity of Atlantic salmon (Salmo salar) in Northern Spain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1632– 1640, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f93-184.

Ward:1993:HMV

[1048] Robert D. Ward and Nicholas G. Elliott. Heterozygosity and morphological variability in the orange roughy, Hoplostethus atlanticus (Teleostei: Trachichthyidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1641– 1649, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f93-185.

Walters:1993:SMP

[1049] Carl J. Walters, Norm Hall, Rhys Brown, and Chris Chubb. Spatial model for the population dynamics and exploitation of the Western Australian rock lobster, Panulirus cygnus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1650–1662, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-186.

Lancaster:1993:CSF

[1050] Jill Lancaster and Alan G. Hildrew. Characterizing in-stream flow refugia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1663–1675, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f93-
187.

Vijayan:1993:HCH

[1051] M. M. Vijayan, A. G. Maule, C. B. Schreck, and T. W. Moon. Hormonal control of hepatic glycogen metabolism in food-deprived, continuously swimming Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1676–1682, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-188.

Miller:1993:RCD

[1052] P. A. Miller, R. P. Lanno, M. E. McMaster, and D. G. Dixon. Relative contributions of dietary and waterborne copper to tissue copper burdens and waterborne-copper tolerance in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1683–1689, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f93-
189.

Rice:1993:FAH

[1053] J. C. Rice. Forecasting abundance from habitat measures using nonparametric density estimation methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1690–1698, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-190.

Kieffer:1993:FFP

[1054] J. D. Kieffer and P. W. Colgan. Foraging flexibility in pumpkinseed (Lepomis gibbosus); influence of habitat structure and prey type. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1699–1705, August 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-191.

Flamarique:1993:SCS

[1055] Iñigo Novales Flamarique and Craig W. Hawryshyn. Spectral characteristics of salmonid migratory routes from Southern Vancouver Island (British Columbia). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8): 1706–1716, August 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-192.

Hontela:1993:ELA

[1056] A. Hontela, J. B. Rasmussen, K. Lederis, H. V. Tra, and G. Chevalier. Elevated levels of arginine vasotocin in the brain of brook trout (Salvelinus fontinalis) from acid lakes: a field test of a potential biomarker for acid stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1717–1727, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f93-
193.

Nurnberg:1993:IBT

[1057] Gertrud K. Nürnberg and Peter J. Dillon. Iron budgets in temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1728–1737, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f93-
194.

Devlin:1993:SSS

[1058] Robert H. Devlin. Sequence of sockeye salmon type 1 and 2 growth hormone genes and the relationship of rainbow trout with Atlantic and Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1738– 1748, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f93-195.

Overholtz:1993:HSF

[1059] William J. Overholtz. Harvesting strategies and fishing mortality reference point comparisons for the Northwest Atlantic stock of Atlantic mackerel (Scomber scombrus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1749–1756, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-196.

Carpenter:1993:CDM

[1060] Stephen R. Carpenter, Richard C. Lathrop, and Alejandro Muñoz del Rio. Comparison of dynamic models for edible phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1757–1767, August 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-197.

Bowen:1993:SGV

[1061] W. D. Bowen, J. W. Lawson, and B. Beck. Seasonal and geographic variation in the species composition and size of prey consumed by grey seals (Halichoerus grypus) on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1768–1778, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-198.

Malley:1993:ANI

[1062] D. F. Malley, P. C. Williams, M. P. Stainton, and B. W. Hauser. Application of near-infrared reflectance spectroscopy in the measurement of carbon, nitrogen, and phosphorus in seston from oligotrophic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1779–1785, August 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-199.

Tang:1993:ISE

[1063] Ming Tang and Daniel Boisclair. Influence of the size of enclosures on the swimming characteristics of juvenile brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1786–1793, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-200.

Gobeil:1993:MSS

[1064] Charles Gobeil and Daniel Cossa. Mercury in sediments and sediment pore water in the Laurentian Trough. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1794–1800, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-201.

Kieser:1993:BCR

[1065] R. Kieser, T. J. Mulligan, Laura J. Richards, and B. M. Leaman. Bias correction of rockfish school cross section widths from digitized echo sounder data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1801– 1811, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f93-202.

U:1993:CRPb

[1066] F. H. U. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(8):1812, August 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-203.

Staurnes:1993:EAW

[1067] Magne Staurnes, Per Blix, and Ola B. Reite. Effects of acid water and aluminum on parr–smolt transformation and seawater tolerance in Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (9):1816–1827, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f93-204.

Todd:1993:GVG

[1068] Thomas N. Todd and Charles O. Hatcher. Genetic variability and glacial origins of yellow perch (Perca flavescens) in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1828–1834, September 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-205.

Johnston:1993:MFF

[1069] Thomas A. Johnston and J. A. Mathias. Mortality of first-feeding postlarval walleye (Stizostedion vitreum) in culture ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1835–1843, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-206.

Fritz:1993:DAI

[1070] S. C. Fritz, S. Juggins, and R. W. Battarbee. Diatom assemblages and ionic characterization of lakes of the Northern Great Plains, North America: a tool for reconstructing past salinity and climate fluctuations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1844–1856, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-207.

Cote:1993:IDD

[1071] Jean Côté, John H. Himmelman, Michel Claereboudt, and John C. Bonardelli. Influence of density and depth on the growth of juvenile sea scallops (Placopecten magellanicus) in suspended culture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1857–1869, September 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-208.

Hinch:1993:RLF

[1072] Scott G. Hinch and Nicholas C. Collins. Relationships of littoral fish abundance to water chemistry and macrophyte variables in Central Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1870–1878, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-209.

Lazzari:1993:IRC

[1073] Mark A. Lazzari, David K. Stevenson, and Richard F. Shaw. Influence of residual circulation and vertical distribution on the abundance and horizontal transport of larval Atlantic herring (Clupea harengus) in a Maine estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1879– 1890, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-210.

Delong:1993:IFT

[1074] Michael D. Delong, R. Brent Summers, and James H. Thorp. Influence of food type on the growth of a riverine amphipod, Gammarus fasciatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1891–1896, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-211.

Dey:1993:EHN

[1075] William P. Dey, Thomas H. Peck, Charlie E. Smith, and Guat-Lian Kreamer. Epizoology of hepatic neoplasia in Atlantic tomcod (Microgadus tomcod) from the Hudson River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1897–1907, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-212.

Bertram:1993:RAC

[1076] Douglas F. Bertram and Gary W. Kaiser. Rhinoceros auklet (Cerorhinca monocerata) nestling diet may gauge Pacific sand lance (Ammodytes hexapterus) recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9): 1908–1915, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-213.

Schnute:1993:ACD

[1077] Jon T. Schnute and Ray Hilborn. Analysis of contradictory data sources in fish stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9): 1916–1923, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-214.

Pepin:1993:MMG

[1078] Pierre Pepin and Steven M. Carr. Morphological, meristic, and genetic analysis of stock structure in juvenile Atlantic cod (Gadus morhua) from the Newfoundland shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9): 1924–1933, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-215.

Cohen:1993:WLR

[1079] Yosef Cohen and Paul Radomski. Water level regulations and fisheries in Rainy Lake and the Namakan Reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1934–1945, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-216.

Urban:1993:TSS

[1080] N. R. Urban and P. L. Brezonik. Transformations of sulfur in sediment microcosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1946– 1960, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-217.

Wahl:1993:CPC

[1081] David H. Wahl and Roy A. Stein. Comparative population characteristics of muskellunge (Esox masquinongy), northern pike (E. lucius), and their hybrid (E. masquinongy × E. lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1961–1968, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-218.

Boisclair:1993:GEF

[1082] Daniel Boisclair and François Marchand. The guts to estimate fish daily ration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1969–1975, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-219.

Pace:1993:RAE

[1083] Michael L. Pace, Stephen B. Baines, Hélène Cyr, and John A. Downing. Relationships among early life stages of Morone americana and Morone saxatilis from long-term monitoring of the Hudson River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1976–1985, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-220.

Youngson:1993:SEF

[1084] A. F. Youngson, J. H. Webb, C. E. Thompson, and D. Knox. Spawning of escaped farmed Atlantic salmon (Salmo salar): Hybridization of females with brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1986–1990, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-221.

Dugan:1993:IIS

[1085] Jenifer E. Dugan and Gary E. Davis. Introduction to the International Symposium on Marine Harvest Refugia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):1991–1992, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-222.

Armstrong:1993:TRB

[1086] David A. Armstrong, Thomas C. Wainwright, Gregory C. Jensen, Paul A. Dinnel, and Helle B. Andersen. Taking refuge from bycatch issues: Red king crab (Paralithodes camtschaticus) and trawl fisheries in the Eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9): 1993–2000, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-223.

Shepherd:1993:WAS

[1087] S. A. Shepherd and L. D. Brown. What is an abalone stock: Implications for the role of refugia in conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):2001–2009, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-224.

Tegner:1993:SCA

[1088] Mia J. Tegner. Southern California abalones: Can stocks be rebuilt using marine harvest refugia? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9): 2010–2018, September 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-225.

Carr:1993:CIR

[1089] Mark H. Carr and Daniel C. Reed. Conceptual issues relevant to marine harvest refuges: Examples from temperate reef fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):2019–2028, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-226.

Dugan:1993:AMR

[1090] Jenifer E. Dugan and Gary E. Davis. Applications of marine refugia to coastal fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):2029–2042, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-227.

Hilborn:1993:LES

[1091] Ray Hilborn and John Winton. Learning to enhance salmon production: Lessons from the salmonid enhancement program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(9):2043–2056, September 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-228.

Walters:1993:RLC

[1092] Carl J. Walters and Francis Juanes. Recruitment limitation as a consequence of natural selection for use of restricted feeding habitats and predation risk taking by juvenile fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2058–2070, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-229.

Hesslein:1993:RSC

[1093] R. H. Hesslein, K. A. Hallard, and P. Ramlal. Replacement of sulfur, carbon, and nitrogen in tissue of growing broad whitefish (Coregonus nasus) in response to a change in diet traced by δ ^{34}S, δ ^{13}C, and δ ^{15}N. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2071–2076, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-230.

Bayley:1993:QLE

[1094] Peter B. Bayley. Quasi-likelihood estimation of marked fish recapture. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2077–2085, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-231.

Locke:1993:BWE

[1095] A. Locke, D. M. Reid, H. C. van Leeuwen, W. G. Sprules, and J. T. Carlton. Ballast wafer exchange as a means of controlling dispersal of freshwater organisms by ships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2086–2093, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-232.

Young:1993:EEC

[1096] Paciencia S. Young and Joseph J. Cech, Jr. Effects of exercise conditioning on stress responses and recovery in cultured and wild young-of-the-year striped bass, Morone saxatilis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2094–2099, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-233.

Fernandez:1993:FCY

[1097] Miriam Fernandez, David Armstrong, and Oscar Iribarne. First cohort of young-of-the-year Dungeness crab, Cancer magister, reduces abundance of subsequent cohorts in intertidal shell habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2100–2105, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-234.

Cheng:1993:DPS

[1098] Jin-Hua Cheng and Ernest S. Chang. Determinants of postmolt size in the American lobster (Homarus americanus). I. D_{1}^{3} is the critical stage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2106–2111, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-235.

Lockwood:1993:PRA

[1099] S. Fournier Lockwood, Robert E. Dillinger, Jr., Tim P. Birt, John M. Green, and Thomas P. Snyder. Phylogenetic relationships among members of the Coregoninae inferred from direct sequencing of PCR-amplified mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2112–2118, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-236.

Riehle:1993:CHU

[1100] Michael D. Riehle and J. S. Griffith. Changes in habitat use and feeding chronology of juvenile rainbow trout (Oncorhynchus mykiss) in fall and the onset of winter in Silver Creek, Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2119–2128, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-237.

Rose:1993:UOF

[1101] G. A. Rose and W. C. Leggett. Use of oceanographic forecasts and echosounders to guide and enhance an inshore gillnet fishery for Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2129–2136, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-238.

Taylor:1993:RPV

[1102] Derek J. Taylor and Paul D. N. Hebert. A reappraisal of phenotypic variation in Daphnia galeata mendotae: The role of interspecific hybridization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2137–2146, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-239.

Sainte-Marie:1993:RCF

[1103] Bernard Sainte-Marie. Reproductive cycle and fecundity of primiparous and multiparous female snow crab, Chionoecetes opilio, in the Northwest Gulf of Saint Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2147–2156, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-240.

Trudel:1993:SED

[1104] Marc Trudel and Daniel Boisclair. An in situ evaluation of the dayto-day variation in the quantity of food consumed by fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2157–2165, October 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-241.

Pepin:1993:ASD

[1105] Pierre Pepin. An appraisal of the size-dependent mortality hypothesis for larval fish: Comparison of a multispecies study with an empirical review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2166–2174, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-242.

Arthur:1993:UPS

[1106] J. R. Arthur and E. Albert. Use of parasites for separating stocks of Greenland halibut (Reinhardtius hippoglossoides) in the Canadian Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2175–2181, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-243.

Measures:1993:GLA

[1107] Lena N. Measures and Luci Bossé. Gammarus lawrencianus (Amphipoda) as intermediate host of Echinorhynchus salmonis (Acanthocephala) in an estuarine environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2182–2184, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-244.

Meffe:1993:LDD

[1108] Gary K. Meffe and Franklin F. Snelson, Jr. Lipid dynamics during reproduction in two livebearing fishes, Gambusia holbrooki and Poecilia latipinna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2185–2191, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-245.

Rasmussen:1993:PSS

[1109] J. B. Rasmussen. Patterns in the size structure of littoral zone macroinvertebrate communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2192–2207, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-246.

Carrick:1993:APN

[1110] Hunter J. Carrick, Claire L. Schelske, Frederick J. Aldridge, and Michael F. Coveney. Assessment of phytoplankton nutrient limitation in productive waters: Application of dilution bioassays. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2208–2221, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-247.

Devito:1993:IRW

[1111] Kevin J. Devito and Peter J. Dillon. Importance of runoff and winter anoxia to the P and N dynamics of a beaver pond. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2222–2234, October 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-248.

Smith:1993:PCA

[1112] Ralph E. H. Smith and Julie A. Maly. Photosynthetic carbon assimilation and allocation by surface and deep phytoplankton in Colpoys Bay (Western Georgian Bay). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2235–2244, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-249.

Parrott:1993:PTS

[1113] Joanne L. Parrott and John B. Sprague. Patterns in toxicity of sublethal mixtures of metals and organic chemicals determined by Microtox(R) and by DNA, RNA, and protein content of fathead minnows (Pimephales promelas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2245–2253, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-250.

Paerl:1993:ERA

[1114] Hans W. Paerl. Emerging role of atmospheric nitrogen deposition in coastal eutrophication: Biogeochemical and trophic perspectives. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10):2254–2269, October 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-251.

Beamish:1993:CEF

[1115] Richard J. Beamish. Climate and exceptional fish production off the West Coast of North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(10): 2270–2291, October 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-252.

Domm:1993:PTS

[1116] S. Domm, R. W. McCauley, E. Kott, and J. D. Ackerman. Physiological and taxonomic separation of two dreissenid mussels in the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2294– 2297, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-253.

Dermott:1993:ILE

[1117] R. Dermott and M. Munawar. Invasion of Lake Erie offshore sediments by Dreissena, and its ecological implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2298–2304, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-254.

Mills:1993:CEP

[1118] Edward L. Mills, Ron M. Dermott, Edward F. Roseman, Donna Dustin, Eric Mellina, David Bruce Conn, and Adrian P. Spidle. Colonization, ecology, and population structure of the “Quagga” mussel (Bivalvia: Dreissenidae) in the Lower Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2305–2314, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-255.

Gislason:1993:ECR

[1119] Henrik Gislason. Effect of changes in recruitment levels on multispecies long-term predictions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):

2315–2322, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-256.

Heide-Jorgensen:1993:IRA

[1120] M. P. Heide-Jørgensen, H. Lassen, J. Teilmann, and R. A. Davis. An index of the relative abundance of wintering belugas, Delphinapterus leucas, and narwhals, Monodon monoceros, off West Greenland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2323–2335, November 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-257.

Bevelhimer:1993:BAD

[1121] Mark S. Bevelhimer and S. Marshall Adams. A bioenergetics analysis of diel vertical migration by Kokanee salmon, Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2336–2349, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-258.

Kline:1993:RET

[1122] Thomas C. Kline, Jr., John J. Goering, Ole A. Mathisen, Patrick H. Poe, Patrick L. Parker, and Richard S. Scalan. Recycling of elements transported upstream by runs of Pacific salmon: II. δ ^{15}N and δ ^{13}C evidence in the Kvichak River Watershed, Bristol Bay, Southwestern Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2350–2365, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-259.

Jackson:1993:GZW

[1123] George D. Jackson. Growth zones within the statolith microstructure of the deepwater squid Moroteuthis ingens (Cephalopoda: Onychoteuthidae): Evidence for a habitat shift? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2366–2374, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-260.

Maekawa:1993:MPA

[1124] Koji Maekawa, Teruaki Hino, Shigeru Nakano, and William W. Smoker. Mate preference in anadromous and landlocked Dolly Varden (Salvelinus malma) females in two Alaskan streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2375–2379, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-261.

Beecher:1993:PMS

[1125] Hal A. Beecher, Thom H. Johnson, and John P. Carleton. Predicting microdistributions of steelhead (Oncorhynchus mykiss) parr from depth and velocity preference criteria: Test of an assumption of the instream flow incremental methodology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11): 2380–2387, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-262.

Somers:1993:AMC

[1126] Keith M. Somers and Donald A. Jackson. Adjusting mercury concentration for fish-size covariation: a multivariate alternative to bivariate regression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2388–2396, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-263.

Grewe:1993:MDV

[1127] Peter M. Grewe, Charles C. Krueger, Charles F. Aquadro, Eldredge Bermingham, Harold L. Kincaid, and Bernie May. Mitochondrial DNA variation among lake trout (Salvelinus namaycush) strains stocked into Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2397– 2403, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-264.

Arts:1993:ERL

[1128] Michael T. Arts, Richard D. Robarts, and Marlene S. Evans. Energy reserve lipids of zooplanktonic crustaceans from an oligotrophic saline lake in relation to food resources and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2404–2420, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-265.

Paloheimo:1993:EEE

[1129] J. E. Paloheimo and Yong Chen. Estimation of effective effort from catch-at-age data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2421– 2428, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-266.

Stenson:1993:PPH

[1130] G. B. Stenson, R. A. Myers, M. O. Hammill, I-H. Ni, W. G. Warren, and M. C. S. Kingsley. Pup production of harp seals, Phoca groenlandica, in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11): 2429–2439, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-267.

McKinley:1993:SVP

[1131] R. S. McKinley, T. D. Singer, J. S. Ballantyne, and G. Power. Seasonal variation in plasma nonesterified fatty acids of lake sturgeon (Acipenser fulvescens) in the vicinity of hydroelectric facilities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2440–2447, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-268.

Youson:1993:ICF

[1132] J. H. Youson, J. A. Holmes, J. A. Guchardi, J. G. Seelye, R. E. Beaver, J. E. Gersmehl, S. A. Sower, and F. W. H. Beamish. Importance of condition factor and the influence of water temperature and photoperiod on metamorphosis of sea lamprey Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2448–2456, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-269.

Hutchings:1993:GVS

[1133] Jeffrey A. Hutchings, Ransom A. Myers, and George R. Lilly. Geographic variation in the spawning of Atlantic cod, Gadus morhua, in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11): 2457–2467, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-270.

Hutchings:1993:EAS

[1134] Jeffrey A. Hutchings and Ransom A. Myers. Effect of age on the seasonality of maturation and spawning of Atlantic cod, Gadus morhua, in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11): 2468–2474, November 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-271.

Unwin:1993:SCW

[1135] M. J. Unwin and D. H. Lucas. Scale characteristics of wild and hatchery Chinook salmon (Oncorhynchus tshawytscha) in the Rakaia River, New Zealand, and their use in stock identification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2475–2484, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-272.

Xu:1993:ETS

[1136] Xucai Xu and Russell B. Millar. Estimation of trap selectivity for male snow crab (Chionoecetes opilio) using the SELECT modeling approach with unequal sampling effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50 (11):2485–2490, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f93-273.

Salvado:1993:PFT

[1137] Carlos A. M. Salvadó. Population field theory with applications to tag analysis and fishery modeling: the empirical green function. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2491–2512, November 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-274.

Millie:1993:MPA

[1138] David F. Millie, Hans W. Paerl, and James P. Hurley. Microalgal pigment assessments using high-performance liquid chromatography: a synopsis of organismal and ecological applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(11):2513–2527, November 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-275.

Balm:1993:ART

[1139] P. H. M. Balm and T. G. Pottinger. Acclimation of rainbow trout (Oncorhynchus mykiss) to low environmental pH does not involve an activation of the pituitary-interrenal axis, but evokes adjustments in branchial ultrastructure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2532– 2541, December 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-276.

Bjornsson:1993:SSS

[1140] Björn Björnsson. Swimming speed and swimming metabolism of Atlantic cod (Gadus morhua) in relation to available food: a laboratory study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2542–2551, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-277.

Skiftesvik:1993:CBA

[1141] Anne Berit Skiftesvik and Øivind Bergh. Changes in behaviour of Atlantic halibut (Hippoglossus hippoglossus) and turbot (Scophthalmus maximus) yolk-sac larvae induced by bacterial infections. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2552–2557, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-278.

Jenkins:1993:SVF

[1142] Gregory P. Jenkins, Megan Shaw, and Bryce D. Stewart. Spatial variation in food-limited growth of juvenile greenback flounder, Rhombosolea tapirina: Evidence from otolith daily increments and otolith scaling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2558–2567, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-279.

Rodriguez-Ariza:1993:BIO

[1143] A. Rodríguez-Ariza, J. Peinado, C. Pueyo, and J. López-Barea. Biochemical indicators of oxidative stress in fish from polluted littoral areas.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2568–2573, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-280.

Hammar:1993:APP

[1144] Johan Hammar, Per Larsson, and Maris Klavins. Accumulation of persistent pollutants in normal and dwarfed Arctic char (Salvelinus alpinus sp. complex). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2574– 2580, December 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-281.

DeBlois:1993:MMB

[1145] E. M. DeBlois and W. C. Leggett. Match/mismatch between the abundance of marine fish eggs and invertebrate predators: an analysis of Calliopius laeviusculus (Amphipoda: Gammaridae) population growth relative to the seasonal spawning cycle of capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2581–2590, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-282.

Wang:1993:ICC

[1146] Y. Wang and R. D. Evans. Influence of calcium concentrations on cadmium uptake by the freshwater mussel Ellptio complanata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2591–2596, December 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-283.

Polacheck:1993:FSP

[1147] Tom Polacheck, Ray Hilborn, and André E. Punt. Fitting surplus production models: Comparing methods and measuring uncertainty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2597–2607, December 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-284.

Pearcy:1993:BPP

[1148] William G. Pearcy, Joseph P. Fisher, and Mary M. Yoklavich. Biology of the Pacific pomfret (Brama japonica) in the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2608–2625, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-285.

Patalas:1993:SVC

[1149] K. Patalas and A. Salki. Spatial variation of crustacean plankton in lakes of different size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12): 2626–2640, December 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-286.

Jackson:1993:FBI

[1150] Donald A. Jackson and Harold H. Harvey. Fish and benthic invertebrates: Community concordance and community–environment relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2641–2651, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-287.

Wilson:1993:NHB

[1151] Chris C. Wilson and Paul D. N. Hebert. Natural hybridization between Arctic char (Salvelinus alpinus) and lake trout (S. namaycush) in the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12): 2652–2658, December 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-288.

Cattaneo:1993:SSB

[1152] Antonella Cattaneo. Size spectra of benthic communities in Laurentian streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2659–2666, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-289.

Playle:1993:CCBa

[1153] Richard C. Playle, D. George Dixon, and Kent Burnison. Copper and cadmium binding to fish gills: Modification by dissolved organic carbon and synthetic ligands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):

2667–2677, December 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-290.

Playle:1993:CCBb

[1154] Richard C. Playle, D. George Dixon, and Kent Burnison. Copper and cadmium binding to fish gills: Estimates of metal–gill stability constants and modelling of metal accumulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2678–2687, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-291.

Havens:1993:PFWb

[1155] Karl E. Havens. Pelagic food web structure in acidic Adirondack mountain, New York, lakes of varying humic content. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2688–2691, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-292.

Pennuto:1993:BRD

[1156] Christopher M. Pennuto and Frank deNoyelies, Jr. Behavioral responses of Drunella coloradensis (Ephemeroptera) nymphs to short-term pH reductions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2692–2697, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f93-293.

Meador:1993:TMP

[1157] J. P. Meador, U. Varanasi, P. A. Robisch, and S-L. Chan. Toxic metals in pilot whales (Globicephala melaena) from standings in 1986 and 1990 on Cape Cod, Massachusetts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2698–2706, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f93-294.

Tuchman:1993:RIM

[1158] Nancy C. Tuchman. Relative importance of microbes versus macroinvertebrate shredders in the process of leaf decay in lakes of differing pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2707–2712, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f93-295.

Hurley:1993:CSA

[1159] James P. Hurley and Paul J. Garrison. Composition and sedimentation of aquatic pigments associated with deep plankton in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2713–2722, December 1993. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f93-296.

Fogarty:1993:RDR

[1160] Michael J. Fogarty. Recruitment distributions revisited. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2723–2728, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-297.

deYoung:1993:RDA

[1161] B. deYoung and G. A. Rose. On recruitment and distribution of Atlantic cod (Gadus morhua) off Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12):2729–2741, December 1993. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f93-298.

Stephenson:1993:MAH

[1162] R. L. Stephenson, D. E. Lane, D. G. Aldous, and R. Nowak. Management of the 4WX Atlantic herring (Clupea harengus) fishery: an evaluation of recent events. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 50(12): 2742–2757, December 1993. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f93-299.

Tsai:1994:SSA

[1163] H. J. Tsai, H. M. Chen, and C. F. Lo. Secretory synthesis of active recombinant fish growth hormone by insect cells using a baculovirus vector. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):1–7, January 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-001.

Cech:1994:NSP

[1164] Joseph J. Cech, Jr., Daniel T. Castleberry, Todd E. Hopkins, and James H. Petersen. Northern squawfish, Ptychocheilus oregonensis, O_{2} consumption rate and respiration model: Effects of temperature and body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):8–12, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-002.

Cech:1994:TCE

[1165] Joseph J. Cech, Jr., Daniel T. Castleberry, and Todd E. Hopkins. Temperature and CO_{2} effects on blood O_{2} equilibria in northern squawfish, Ptychocheilus oregonensis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (1):13–19, January 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-003.

Weiher:1994:AAP

[1166] Evan R. Weiher, Charles W. Boylen, and Paul A. Bukaveckas. Alterations in aquatic plant community structure following liming of an acidic Adirondack lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):20–24, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-004.

Baines:1994:RBS

[1167] Stephen B. Baines and Michael L. Pace. Relationships between suspended particulate matter and sinking flux along a trophic gradient and implications for the fate of planktonic primary production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):25–36, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-005.

Ralph:1994:SCM

[1168] Stephen C. Ralph, Geoffrey C. Poole, Loveday L. Conquest, and Robert J. Naiman. Stream channel morphology and woody debris in logged and unlogged basins of Western Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):37–51, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-006.

Junger:1994:QUS

[1169] Michel Junger and Dolors Planas. Quantitative use of stable carbon isotope analysis to determine the trophic base of invertebrate communities in a boreal forest lotic system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (1):52–61, January 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-007.

Kornfield:1994:MDC

[1170] Irv Kornfield and Frederick W. Kircheis. Mitochondrial DNA and conservation of an Aboriginal Arctic char (Salvelinus alpinus oquassa) from Floods Pond, Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (1):62–67, January 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-008.

Rieman:1994:UOM

[1171] B. E. Rieman, D. L. Myers, and R. L. Nielsen. Use of otolith microchemistry to discriminate Oncorhynchus nerka of resident and anadromous origin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):68–77, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-009.

Methven:1994:TVS

[1172] David A. Methven and Christopher Bajdik. Temporal variation in size and abundance of juvenile Atlantic cod (Gadus morhua) at an inshore site off Eastern Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):78–90, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-010.

Friedland:1994:UOM

[1173] K. D. Friedland and D. G. Reddin. Use of otolith morphology in stock discriminations of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):91–98, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-011.

Barange:1994:DCV

[1174] M. Barange, I. Hampton, S. C. Pillar, and M. A. Soule. Determination of composition and vertical structure of fish communities using in situ measurements of acoustic target strength. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):99–109, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-012.

Mace:1994:RBC

[1175] Pamela M. Mace. Relationships between common biological reference points used as thresholds and targets of fisheries management strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):110–122, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-013.

Ouellet:1994:VDN

[1176] Patrick Ouellet and Denis Lefaivre. Vertical distribution of northern shrimp (Pandalus borealis) larvae in the Gulf of St Lawrence; implications for trophic interactions and transport. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):123–132, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-014.

Sadovy:1994:EPR

[1177] Yvonne Sadovy and Kenneth P. Severin. Elemental patterns in red hind (Epinephelus guttatus) otoliths from Bermuda and Puerto Rico reflect growth rate, not temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):133–141, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-015.

Wroblewski:1994:OAA

[1178] J. S. Wroblewski, Wade L. Bailey, and Kristine A. Howse. Observations of adult Atlantic cod (Gadus morhua) overwintering in nearshore waters of Trinity Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (1):142–150, January 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-016.

Locke:1994:ZCW

[1179] Andrea Locke, W. Gary Sprules, W. Keller, and J. Roger Pitblado. Zooplankton communities and water chemistry of Sudbury area lakes: Changes related to pH recovery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):151–160, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-017.

Bertrand:1994:SDP

[1180] Normand Bertrand and Warwick F. Vincent. Structure and dynamics of photosynthetic picoplankton across the saltwater transition zone of the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (1):161–171, January 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-018.

Fontaine:1994:SDH

[1181] Pierre-Michel Fontaine, Mike O. Hammill, Cyrille Barrette, and Michael C. Kingsley. Summer diet of the harbour porpoise (Phocoena phocoena) in the estuary and the Northern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):172–178, January 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-019.

Alexander:1994:TTE

[1182] James E. Alexander, Jr., James H. Thorp, and Ronald D. Fell. Turbidity and temperature effects on oxygen consumption in the zebra mussel (Dreissena polymorpha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1): 179–184, January 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-020.

Aguilar:1994:MRO

[1183] Carmen Aguilar and Kenneth H. Nealson. Manganese reduction in Oneida Lake, New York: Estimates of spatial and temporal manganese flux. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):185–196, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-021.

Tsiger:1994:LHF

[1184] V. V. Tsiger, V. I. Skirin, N. I. Krupyanko, K. A. Kashkin, and A. Yu. Semenchenko. Life history forms of male masu salmon (Oncorhynchus masou) in South Primor’e, Russia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):197–208, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-022.

Johnson:1994:LTE

[1185] Lionel Johnson. Long-term experiments on the stability of two fish populations in previously unexploited Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):209–225, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-023.

Johnson:1994:PPE

[1186] Lionel Johnson. Pattern and process in ecological systems: a step in the development of a general ecological theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(1):226–246, January 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-024.

Moran:1994:EAC

[1187] Paloma Morán, Alberto M. Pendás, Eva Garcia-Vázquez, Jorge T. Izquierdo, and Dennis T. Rutherford. Electrophoretic assessment of the contribution of transplanted Scottish Atlantic salmon (Salmo salar) to the Esva River (Northern Spain). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):248–252, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-025.

Holmes:1994:TNL

[1188] John A. Holmes and Ping Lin. Thermal niche of larval sea lamprey, Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):

253–262, February 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-026.

Xiao:1994:GMC

[1189] Yongshun Xiao. Growth models with corrections for the retardative effects of tagging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):263–267, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-027.

McLaughlin:1994:FMR

[1190] Robert L. McLaughlin, James W. A. Grant, and Donald L. Kramer. Foraging movements in relation to morphology, water-column use, and diet for recently emerged brook trout (Salvelinus fontinalis) in still-water pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):268–279, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-028.

Park:1994:AEH

[1191] E.-H. Park, H.-H. Chang, W. N. Joo, H.-S. Chung, and H.-S. Kwak. Assessment of the estuarine hermaphroditic fish Rivulus marmoratus as a useful euryhaline species for acute toxicity tests as shown using cadmium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):280–285, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-029.

Juinio:1994:ERG

[1192] Marie Antonette R. Juinio and J. Stanley Cobb. Estimation of recent growth of field-caught postlarval American lobsters, Homarus americanus, from RNA:DNA ratios. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):286–294, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-030.

Jackson:1994:TBT

[1193] John K. Jackson, Bernard W. Sweeney, Thomas L. Bott, J. Denis Newbold, and Louis A. Kaplan. Transport of Bacillus thuringiensis var. israelensis and its effect on drift and benthic densities of nontarget macroinvertebrates in the Susquehanna River, Northern Pennsylvania. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):295–314, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-031.

Goksoyr:1994:EDI

[1194] Anders Goksøyr, Marit Bjørnevik, and Amund Maage. Effects of dietary iron concentrations on the cytochrome P450 system of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2): 315–320, February 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-032.

Pihl:1994:CDD

[1195] Leif Pihl. Changes in the diet of demersal fish due to eutrophicationinduced hypoxia in the Kattegat, Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):321–336, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://

cdnsciencepub.com/doi/full/10.1139/f94-033.

Gagnon:1994:SEB

[1196] Monique M. Gagnon, Julian J. Dodson, Peter V. Hodson, Glen Van Der Kraak, and John H. Carey. Seasonal effects of bleached kraft mill effluent on reproductive parameters of white sucker (Catostomus commersoni) populations of the St. Maurice River, Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):337–347, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-034.

Kenchington:1994:FAS

[1197] E. L. Kenchington and W. E. Full. Fourier analysis of sea scallop (Placopecten magellanicus) shells in determining population structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):348–356, February 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-035.

Richards:1994:AGL

[1198] Laura J. Richards, Jon T. Schnute, and Jeff Fargo. Application of a generalized logit model to condition data for trawl-caught Pacific halibut, Hippoglossus stenolepis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2): 357–364, February 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-036.

Arthur:1994:ERD

[1199] A. D. Arthur and D. G. Dixon. Effects of rearing density on the growth response of juvenile fathead minnow (Pimephales promelas) under toxicant-induced stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (2):365–371, February 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-037.

Rankin:1994:ACT

[1200] M. G. Rankin and D. G. Dixon. Acute and chronic toxicity of waterborne arsenite to rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):372–380, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-038.

Cabana:1994:PFC

[1201] Gilbert Cabana, Alain Tremblay, Jacob Kalff, and Joseph B. Rasmussen. Pelagic food chain structure in Ontario lakes: a determinant of mercury levels in lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):381–389, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-039.

Mazumder:1994:PCRa

[1202] Asit Mazumder. Phosphorus–chlorophyll relationships under contrasting herbivory and thermal stratification: Predictions and patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):390–400, February 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-040.

Mazumder:1994:PCRb

[1203] Asit Mazumder. Phosphorus–chlorophyll relationships under contrasting zooplankton community structure: Potential mechanisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):401–407, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-041.

Warren:1994:MMS

[1204] William G. Warren. Maturity and molt status of crustacea: Determining classification without prior criteria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):408–416, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-042.

Xu:1994:CLS

[1205] Shizhong Xu, Carol J. Kobak, and Peter E. Smouse. Constrained least squares estimation of mixed population stock composition from mtDNA haplotype frequency data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (2):417–425, February 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-043.

Wagemann:1994:CHM

[1206] R. Wagemann and R. E. A. Stewart. Concentrations of heavy metals and selenium in tissues and some foods of walrus (Odobenus rosmarus rosmarus) from the Eastern Canadian Arctic and sub-arctic, and associations between metals, age, and gender. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):426–436, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-044.

Heath:1994:HCT

[1207] Susan Heath, Wayne A. Bennett, James Kennedy, and Thomas L. Beitinger. Heat and cold tolerance of the fathead minnow, Pimephales promelas, exposed to the synthetic pyrethroid cyfluthrin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):437–440, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-045.

Thomson:1994:CSI

[1208] Keith A. Thomson, W. James Ingraham, Jr., Michael C. Healey, Paul H. LeBlond, Cornelius Groot, and Christopher G. Healey. Computer simulations of the influence of ocean currents on Fraser River sockeye salmon (Oncorhynchus nerka) return times. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):441–449, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-046.

Cowan:1994:DFP

[1209] Cathy A. Cowan and Barbara L. Peckarsky. Diel feeding and positioning periodicity of a grazing mayfly in a trout stream and a fishless stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):450–459, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-047.

Hargreaves:1994:RPS

[1210] N. B. Hargreaves, D. M. Ware, and G. A. McFarlane. Return of Pacific sardine (Sardinops sagax) to the British Columbia Coast in 1992. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):460–463, February 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-048.

Hodson:1994:STV

[1211] P. V. Hodson, M. Castonguay, C. M. Couillard, C. Desjardins, E. Pelletier, and R. McLeod. Spatial and temporal variations in chemical contamination of American eels, Anguilla rostrata, captured in the estuary of the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (2):464–478, February 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-049.

Castonguay:1994:WRA

[1212] M. Castonguay, P. V. Hodson, C. M. Couillard, M. J. Eckersley, J.-D. Dutil, and G. Verreault. Why is recruitment of the American eel, Anguilla rostrata, declining in the St. Lawrence River and Gulf? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(2):479–488, February 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-050.

Danylchuk:1994:SRP

[1213] Andy J. Danylchuk and Michael G. Fox. Seasonal reproductive patterns of pumpkinseed (Lepomis gibbosus) populations with varying body size characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):490–500, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-051.

Ogura:1994:CKO

[1214] Miki Ogura and Soto o Ito. Change in the known ocean distribution of Japanese chum salmon, Oncorhynchus keta, in relation to the progress of stock enhancement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (3):501–505, March 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-052.

Panfili:1994:SVG

[1215] Jacques Panfili, Marie-Claude Ximénès, and Alain J. Crivelli. Sources of variation in growth of the European eel (Anguilla anguilla) estimated from otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):506–515, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-053.

Mooij:1994:GRF

[1216] W. M. Mooij, E. H. R. R. Lammens, and W. L. T. Van Densen. Growth rate of 0+ fish in relation to temperature, body size, and food in shallow eutrophic Lake Tjeukemeer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (3):516–526, March 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-054.

Wilson:1994:MCPa

[1217] Rod W. Wilson, Harold L. Bergman, and Chris M. Wood. Metabolic costs and physiological consequences of acclimation to aluminum in juvenile rainbow trout (Oncorhynchus mykiss). 1: Acclimation specificity, resting physiology, feeding, and growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):527–535, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-055.

Wilson:1994:MCPb

[1218] Rod W. Wilson, Harold L. Bergman, and Chris M. Wood. Metabolic costs and physiological consequences of acclimation to aluminum in juvenile rainbow trout (Oncorhynchus mykiss). 2: Gill morphology, swimming performance, and aerobic scope. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):536–544, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-056.

Gauldie:1994:DRS

[1219] R. W. Gauldie, K. Xi, and S. K. Sharma. Developing a Raman spectral method for measuring the strontium and calcium concentrations of fish otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):545–551, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-057.

Labelle:1994:LME

[1220] M. Labelle. A likelihood method for estimating Pacific salmon escapement based on fence counts and mark–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):552–566, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-058.

Plisetskaya:1994:LGE

[1221] Erika M. Plisetskaya, Thomas W. Moon, Donald A. Larsen, Glen D. Foster, and Walton W. Dickhoff. Liver glycogen, enzyme activities, and pancreatic hormones in juvenile Atlantic salmon (Salmo salar) during their first summer in seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (3):567–576, March 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-059.

Tallman:1994:HSG

[1222] Ross F. Tallman and Michael C. Healey. Homing, straying, and gene flow among seasonally separated populations of chum salmon (Oncorhynchus keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):577–588, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-060.

Perry:1994:IHA

[1223] R. Ian Perry and Stephen J. Smith. Identifying habitat associations of marine fishes using survey data: an application to the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):589–602, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-061.

Forbes:1994:SSS

[1224] L. Scott Forbes and Randall M. Peterman. Simple size-structured models of recruitment and harvest in Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):603–616, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-062.

DeLafontaine:1994:ZBS

[1225] Yves De Lafontaine. Zooplankton biomass in the Southern Gulf of St. Lawrence: Spatial patterns and the influence of freshwater runoff. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):617–635, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-063.

Bosakowski:1994:AFE

[1226] Thomas Bosakowski and Eric J. Wagner. Assessment of fin erosion by comparison of relative fin length in hatchery and wild trout in Utah. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):636–641, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-064.

Shackell:1994:ASS

[1227] Nancy L. Shackell, Peter A. Shelton, John M. Hoenig, and James E. Carscadden. Age- and sex-specific survival of Northern Grand Bank capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (3):642–649, March 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-065.

Brylinsky:1994:IFT

[1228] Michael Brylinsky, Jamie Gibson, and Donald C. Gordon, Jr. Impacts of flounder trawls on the intertidal habitat and community of the Minas Basin, Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (3):650–661, March 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-066.

Korman:1994:DEB

[1229] Josh Korman, David R. Marmorek, Gilles L. Lacroix, Peter G. Amiro, John A. Ritter, Walton D. Watt, Richard E. Cutting, and Donald C. E. Robinson. Development and evaluation of a biological model to assess regional-scale effects of acidification on Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):662–680, March 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-067.

Campeau:1994:RIA

[1230] Suzanne Campeau, Henry R. Murkin, and Rodger D. Titman. Relative importance of algae and emergent plant litter to freshwater marsh invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):681–692, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-068.

Borgmann:1994:PSC

[1231] U. Borgmann and D. M. Whittle. Particle-size-conversion efficiency, invertebrate production, and potential fish production in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):693–700, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-069.

Hinch:1994:SAA

[1232] Scott G. Hinch, Keith M. Somers, and Nicholas C. Coliins. Spatial autocorrelation and assessment of habitat–abundance relationships in littoral zone fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):701–712, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-070.

Walters:1994:CBP

[1233] Carl Walters and Donald Ludwig. Calculation of Bayes posterior probability distributions for key population parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):713–722, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-071.

Gascuel:1994:MSA

[1234] Didier Gascuel. Une méthode simple d’ajustement des clés taille/âge : application aux captures d’albacores (Thunnus albacares) de l’Atlantique Est. (French) [A simple method of adjusting the size/age keys: application to yellowfin tuna (Thunnus albacares) catches from the East Atlantic]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):723–733, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-072.

Anganuzzi:1994:ESS

[1235] Alejandro Anganuzzi, Ray Hilborn, and John R. Skalski. Estimation of size selectivity and movement rates from mark–recovery data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):734–742, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-073.

Simpson:1994:RSF

[1236] James J. Simpson. Remote sensing in fisheries: a tool for better management in the utilization of a renewable resource. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(3):743–771, March 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-074.

Miller:1994:ESD

[1237] Loren M. Miller and Anne R. Kapuscinski. Estimation of selection differentials from fish scales: a step towards evaluating genetic alteration of fish size in exploited populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (4):774–783, April 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f94-075. See [1878] and reply [1879].

Benfieid:1994:APP

[1238] Mark C. Benfieid and David V. Aldrich. Avoidance of pentachlorophenol by postlarval brown shrimp (Penaeus aztecus) (Decapoda, Penaeidae) in a laminar-flow choice chamber. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):784–791, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-076.

VanOffelen:1994:SDI

[1239] Henry K. Van Offelen, Charles C. Krueger, Carl L. Schofield, and Chris Keleher. Survival, distribution, and ion composition in two strains of brook trout (Salvelinus fontinalis) fry after exposure to episodic pH depressions in an Adirondack lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):792–799, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f94-077.

Madenjian:1994:WPC

[1240] Charles P. Madenjian, Stephen R. Carpenter, and Peter S. Rand. Why are the PCB concentrations of salmonine individuals from the same lake so highly variable? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (4):800–807, April 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-078.

Marshall:1994:GRG

[1241] C. Tara Marshall and Kenneth T. Frank. Geographic responses of groundfish to variation in abundance: Methods of detection and their interpretation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):808–816, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f94-
079.

Cordue:1994:ACR

[1242] P. L. Cordue and R. I. C. C. Francis. Accuracy and choice in risk estimation for fisheries assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (4):817–829, April 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-080.

Moltschaniwskyj:1994:MTG

[1243] N. A. Moltschaniwskyj. Muscle tissue growth and muscle fibre dynamics in the tropical loliginid squid Photololigo sp. (Cephalopoda: Loliginidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):830–835, April 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-081.

Beeman:1994:ASJ

[1244] J. W. Beeman, D. W. Rondorf, and M. E. Tilson. Assessing smoltification of juvenile spring Chinook salmon (Oncorhynchus tshawytscha) using changes in body morphology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (4):836–844, April 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-082.

Vandermeulen:1994:AOS

[1245] J. H. Vandermeulen and J. G. Singh. ARROW oil spill, 1970–90: Persistence of 20-yr weathered bunker C fuel oil. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):845–855, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-083.

Martel:1994:DSR

[1246] André Martel, Andrea F. Mathieu, C. Scott Findlay, Stephen J. Nepszy, and Joseph H. Leach. Daily settlement rates of the zebra mussel, Dreissena polymorpha, on an artificial substrate correlate with veliger abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):856–861, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-084.

Lebo:1994:IVN

[1247] Martin E. Lebo, John E. Reuter, Charles R. Goldman, and Cathryn L. Rhodes. Interannual variability of nitrogen limitation in a desert lake: Influence of regional climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (4):862–872, April 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-085.

DeMelo:1994:AVS

[1248] Rita De Melo and Paul D. N. Hebert. Allozymic variation and species diversity in North American bosminidae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):873–880, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-086.

Levasseur:1994:OBD

[1249] M. Levasseur, M. D. Keller, E. Bonneau, D. D’Amours, and W. K. Bellows. Oceanographic basis of a DMS-related Atlantic cod (Gadus morhua) fishery problem: Blackberry feed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):881–889, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-087.

Fechhelm:1994:ECW

[1250] Robert G. Fechhelm, James D. Bryan, William B. Griffiths, William J. Wilson, and Benny J. Gallaway. Effect of coastal winds on the summer dispersal of young least cisco (Coregonus sardinella) from the Colville River to Prudhoe Bay, Alaska: a simulation model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):890–899, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-088.

McGarvey:1994:ASO

[1251] Richard McGarvey. An age-structured open-access fishery model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):900–912, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-089.

Bremigan:1994:GDL

[1252] Mary T. Bremigan and Roy A. Stein. Gape-dependent larval foraging and zooplankton size: Implications for fish recruitment across systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):913–922, April 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-090.

Miskimmin:1994:LTI

[1253] Brenda M. Miskimmin and David W. Schindler. Long-term invertebrate community response to toxaphene treatment in two lakes: 50-yr records reconstructed from lake sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):923–932, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-091.

Monosson:1994:ICP

[1254] Emily Monosson and John J. Stegeman. Induced cytochrome P4501A in winter flounder, Pleuronectes americanus, from offshore and coastal sites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):933–941, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-092.

Jensen:1994:LPM

[1255] A. L. Jensen. Larkin’s predation model of lake trout (Salvelinus namaycush) extinction with harvesting and sea lamprey (Petromyzon marinus) predation: a qualitative analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):942–945, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-093.

Walters:1994:POS

[1256] Carl Walters and André Punt. Placing odds on sustainable catch using virtual population analysis and survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):946–958, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-094.

Siddall:1994:IPH

[1257] Mark E. Siddall, Lena N. Measures, and Sherwin S. Desser. Infection with the piroplasm Haemohormidium terranovae in relation to haematocrit and mortality of American plaice (Hippoglossoides platessoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):959–964, April 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-095.

Bradford:1994:TAC

[1258] Michael J. Bradford. Trends in the abundance of Chinook salmon (Oncorhynchus tshawytscha) of the Nechako River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):965–973, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-096.

Rice:1994:PGD

[1259] Stanley D. Rice, Robert E. Thomas, and Adam Moles. Physiological and growth differences in three stocks of underyearling sockeye salmon (Oncorhynchus nerka) on early entry into seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):974–980, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-097.

Anonymous:1994:LCLa

[1260] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(4):981–983, April 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-098.

Henderson:1994:RTW

[1261] Bryan A. Henderson and Stephen J. Nepszy. Reproductive tactics of walleye (Stizostedion vitreum) in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):986–997, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-099.

Gribble:1994:MSY

[1262] N. Gribble and M. Dredge. Mixed-species yield-per-recruit simulations of the effect of seasonal closure on a Central Queensland coastal prawn trawling ground. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):998–1011, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f94-
100.

Blom:1994:CSG

[1263] Geir Blom, Terje Svåsand, Knut E. Jørstad, Håkon Otterå, Ole I. Paulsen, and Jens C. Holm. Comparative survival and growth of two strains of Atlantic cod (Gadus morhua) through the early life stages in a marine pond. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1012– 1023, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-101.

Mellina:1994:PDA

[1264] Eric Mellina and Joseph B. Rasmussen. Patterns in the distribution and abundance of zebra mussel (Dreissena polymorpha) in rivers and lakes in relation to substrate and other physicochemical factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1024–1036, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-102.

Richardson:1994:EBI

[1265] John S. Richardson and Chris J. Perrin. Effects of the bacterial insecticide Bacillus thuringiensis var. kurstaki (Btk) on a stream benthic community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1037–1045, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-103.

Swain:1994:FDC

[1266] D. P. Swain and A. F. Sinclair. Fish distribution and catchability: What is the appropriate measure of distribution? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1046–1054, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-104.

Wen:1994:PFL

[1267] Yuan Hua Wen, Alain Vézina, and Robert Henry Peters. Phosphorus fluxes in limnetic cladocerans: Coupling of allometry and compartmental analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1055–1064, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-105.

StLouis:1994:IWS

[1268] Vincent L. St. Louis, John W. M. Rudd, Carol A. Kelly, Ken G. Beaty, Nicholas S. Bloom, and Robert J. Flett. Importance of wetlands as sources of methyl mercury to boreal forest ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1065–1076, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-106.

Fore:1994:SPI

[1269] Leska S. Fore, James W. Karr, and Loveday L. Conquest. Statistical properties of an index of biological integrity used to evaluate water resources. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1077–1087, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-107.

Carignan:1994:RPN

[1270] Richard Carignan, Stéphane Lorrain, and Ken Lum. A 50-yr record of pollution by nutrients, trace metals, and organic chemicals in the St Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (5):1088–1100, May 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-108.

Grewe:1994:MDD

[1271] Peter M. Grewe, Adam J. Smolenski, and Robert D. Ward. Mitochondrial DNA diversity in jackass morwong (Nemadactylus macropterus: Teleostei) from Australian and New Zealand waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1101–1109, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-109.

Claxton:1994:CFM

[1272] W. T. Claxton, C. K. Govind, and R. W. Elner. Chela function, morphometric maturity, and the mating embrace in male snow crab, Chionoecetes opilio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1110–1118, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f94-
110.

Krohn:1994:USV

[1273] Martha M. Krohn and Daniel Boisdair. Use of a stereo-video system to estimate the energy expenditure of free-swimming fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1119–1127, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-111.

Pierce:1994:LFC

[1274] Clay L. Pierce, Joseph B. Rasmussen, and William C. Leggett. Littoral fish communities in Southern Quebec lakes: Relationships with limnological and prey resource variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1128–1138, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-112.

Scrivener:1994:JCS

[1275] J. Charles Scrivener, Thomas G. Brown, and Bruce C. Andersen. Juvenile Chinook salmon (Oncorhynchus tshawytscha) utilization of Hawks Creek, a small and nonnatal tributary of the Upper Fraser River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1139–1146, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-113.

Stephenson:1994:EEC

[1276] Malcolm Stephenson, Gregory Mierle, Ronald A. Reid, and Gerald L. Mackie. Effects of experimental and cultural lake acidification on littoral benthic macroinvertebrate assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1147–1161, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-114.

Hutcheson:1994:PRP

[1277] Michael S. Hutcheson and Patrick L. Stewart. A possible relict population of Mesodesma deauratum (Turton): Bivalvia (Mesodesmatidae) from the Southeast Shoal, Grand Banks of Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1162–1168, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-115.

Winters:1994:LSW

[1278] G. H. Winters and J. P. Wheeler. Length-specific weight as a measure of growth success of adult Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1169–1179, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-116.

Hanson:1994:RPT

[1279] Mark A. Hanson and Malcolm G. Butler. Responses of plankton, turbidity, and macrophytes to biomanipulation in a shallow prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1180–1188, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-117.

Koutnik:1994:PSD

[1280] Michael A. Koutnik and Dianna K. Padilla. Predicting the spatial distribution of Dreissena polymorpha (zebra mussel) among inland lakes of Wisconsin: Modeling with a GIS. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1189–1196, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-118.

Petersen:1994:DPN

[1281] James H. Petersen, Dena M. Gadomski, and Thomas P. Poe. Differential predation by northern squawfish (Ptychocheilus oregonensis) on live and dead juvenile salmonids in the Bonneville Dam Tailrace (Columbia River). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1197–1204, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-119.

Kleinow:1994:CPB

[1282] Kevin M. Kleinow, Herman H. Jarboe, Kathy E. Shoemaker, and Kevin J. Greenless. Comparative pharmacokinetics and bioavailability of oxolinic acid in channel catfish (Ictalurus punctatus) and rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (5):1205–1211, May 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-120.

Anonymous:1994:LCLb

[1283] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(5):1212–1213, May 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-121.

Toline:1994:GDA

[1284] C. Anna Toline and Allan J. Baker. Genetic differentiation among populations of the northern redbelly dace (Phoxinus eos) in Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1218–1228, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-122.

Kristoffersen:1994:IPG

[1285] Knut Kristoffersen, Morten Halvorsen, and Lisbeth Jørgensen. Influence of parr growth, lake morphology, and freshwater parasites on the degree of anadromy in different populations of Arctic char (Salvelinus alpinus) in Northern Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6): 1229–1246, June 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-123.

Attwood:1994:VDG

[1286] Colin G. Attwood and Bruce A. Bennett. Variation in dispersal of Galjoen (Coracinus capensis) (Teleostei: Coracinidae) from a marine reserve. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1247–1257, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-124.

Lagardere:1994:IWP

[1287] J. P. Lagardère, M. L. Bégout, J. Y. Lafaye, and J. P. Villotte. Influence of wind-produced noise on orientation in the sole (Solea solea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1258–1264, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-125.

Tamai:1994:LMP

[1288] Yoichi Tamai, Hisako Kojima, Kumiko Takayama-Abe, and Atsushi Suzuki. Lipids and myelin proteins in the brains of coelacanth (Latimeria chalumnae), lungfish (Lepidosiren paradoxa and Protopterus aetiopicus), bichir (Polypterus senegalus), and sturgeon (Acipenser ruthenus) (Osteichthyes): Phylogenetic implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1265–1272, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-126.

Stevens:1994:AMT

[1289] Bradley G. Stevens, Jan A. Haaga, and William E. Donaldson. Aggregative mating of Tanner crabs, Chionoecetes bairdi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1273–1280, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-127.

Axler:1994:PND

[1290] Richard P. Axler, Charles Rose, and Craig A. Tikkanen. Phytoplankton nutrient deficiency as related to atmospheric nitrogen deposition in Northern Minnesota acid-sensitive lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1281–1296, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-128.

deYoung:1994:ADM

[1291] Brad deYoung, John Anderson, Richard J. Greatbatch, and Paul Fardy. Advection–diffusion modelling of larval capelin (Mallotus villosus) dispersion in Conception Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1297–1307, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-129.

Hudon:1994:LSA

[1292] Christiane Hudon. Large-scale analysis of Atlantic Nova Scotia American lobster (Homarus americanus) landings with respect to habitat, temperature, and wind conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (6):1308–1321, June 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-130.

Pollard:1994:ABR

[1293] Susan M. Pollard, Roy G. Danzmann, and Ross R. Claytor. Association between the regulatory locus PGM-1r* and life-history types of juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1322–1329, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-131.

Guy:1994:DTP

[1294] Martha Guy, William D. Taylor, and John C. H. Carter. Decline in total phosphorus in the surface waters of lakes during summer stratification, and its relationship to size distribution of particles and sedimentation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1330–1337, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-132.

Gu:1994:SCN

[1295] Binhe Gu, Donald M. Schell, and Vera Alexander. Stable carbon and nitrogen isotopic analysis of the plankton food web in a subarctic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1338–1344, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-133.

Gundersen:1994:PHH

[1296] Deke T. Gundersen, Sjahrul Bustaman, Wayne K. Seim, and Lawrence R. Curtis. pH, hardness, and humic acid influence aluminum toxicity to rainbow trout (Oncorhynchus mykiss) in weakly alkaline waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1345–1355, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-134.

Cope:1994:CMB

[1297] W. Gregory Cope, James G. Wiener, Mark T. Steingraeber, and Gary J. Atchison. Cadmium, metal-binding proteins, and growth in bluegill (Lepomis macrochirus) exposed to contaminated sediments from the Upper Mississippi River Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6): 1356–1367, June 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-135.

Vignier:1994:IMF

[1298] V. Vignier, J. H. Vandermeulen, J. Singh, and D. Mossman. Interannual mixed function oxidase (MFO) activity in winter flounder (Pleuronectes americanus) from a coal tar contaminated estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1368–1375, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-136.

Lamontagne:1994:HSF

[1299] S. Lamontagne and D. W. Schindler. Historical status of fish populations in Canadian Rocky Mountain lakes inferred from subfossil Chaoborus (Diptera: Chaoboridae) mandibles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1376–1383, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-137.

Stow:1994:PCT

[1300] Craig A. Stow, Stephen R. Carpenter, and James F. Amrhein. PCB concentration trends in Lake Michigan Coho (Oncorhynchus kisutch) and Chinook salmon (O. tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1384–1390, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-138.

Yang:1994:ACS

[1301] Xiuwei Yang and Terry A. Dick. Arctic char (Salvelinus alpinus) and rainbow trout (Oncorhynchus mykiss) differ in their growth and lipid metabolism in response to dietary polyunsaturated fatty acids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1391–1400, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-139.

Perry:1994:EED

[1302] R. Ian Perry, Max Stocker, and Jeff Fargo. Environmental effects on the distributions of groundfish in Hecate Strait, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1401–1409, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-140.

Walker:1994:TTD

[1303] Mary K. Walker, Philip M. Cook, Allan R. Batterman, Brian C. Butterworth, Christine Berini, John J. Libal, Lawrence C. Hufnagle, and Richard E. Peterson. Translocation of 2,3,7,8-tetrachlorodibenzo-p-dioxin from adult female lake trout (Salvelinus namaycush) to oocytes: Effects on early life stage development and sac fry survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1410–1419, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-141.

Chen:1994:RRA

[1304] Y. Chen, D. A. Jackson, and J. E. Paloheimo. Robust regression approach to analyzing fisheries data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1420–1429, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-142.

Taylor:1994:PSI

[1305] Eric B. Taylor, Terry D. Beacham, and Masahide Kaeriyama. Population structure and identification of North Pacific Ocean chum salmon (Oncorhynchus keta) revealed by an analysis of minisateliite DNA variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1430–1442, June 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-143.

Pringle:1994:QEA

[1306] Catherine M. Pringle and Gail A. Blake. Quantitative effects of atyid shrimp (Decapoda: Atyidae) on the depositional environment in a tropical stream: Use of electricity for experimental exclusion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1443–1450, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-144.

Wipfli:1994:LTB

[1307] Mark S. Wipfli, Richard W. Merritt, and William W. Taylor. Low toxicity of the black fly larvicide Bacillus thuringiensis var. israelensis to early stages of brook trout (Salvelinus fontinalis), brown trout (Salmo trutta), and steelhead trout (Oncorhynchus my kiss) following direct and indirect exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(6):1451–1458, June 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-145.

Tuljapurkar:1994:NFD

[1308] Shripad Tuljapurkar, Carl Boe, and Kenneth W. Wachter. Nonlinear feedback dynamics in fisheries: Analysis of the deriso–schnute model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1462–1473, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-146.

Rosenberg:1994:NRD

[1309] Gary Rosenberg and Michael L. Ludyanskiy. A nomenclatural review of Dreissena (Bivalvia: Dreissenidae), with identification of the Quagga mussel as Dreissena bugensis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (7):1474–1484, July 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-147.

Spidle:1994:IGL

[1310] Adrian P. Spidle, J. Ellen Marsden, and Bernie May. Identification of the Great Lakes Quagga mussel as Dreissena bugensis from the Dnieper River, Ukraine, on the basis of allozyme variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1485–1489, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-148.

Vainola:1994:GZM

[1311] Risto Väinölä, Bruce J. Riddoch, Robert D. Ward, and Roger I. Jones. Genetic zoogeography of the Mysis relicta species group (Crustacea:

Mysidacea) in Northern Europe and North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1490–1505, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-149.

Pelletier:1994:SDP

[1312] Dominique Pelletier and Ana M. Parma. Spatial distribution of Pacific halibut (Hippoglossus stenolepis): an application of geostatistics to longline survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (7):1506–1518, July 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-150.

Young:1994:OEC

[1313] Paciencia S. Young and Joseph J. Cech, Jr. Optimum exercise conditioning velocity for growth, muscular development, and swimming performance in young-of-the-year striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1519–1527, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-151.

Young:1994:EDE

[1314] Paciencia S. Young and Joseph J. Cech, Jr. Effects of different exercise conditioning velocities on the energy reserves and swimming stress responses in young-of-the-year striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1528–1534, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-152.

Cave:1994:PSM

[1315] James D. Cave and William J. Gazey. A preseason simulation model for fisheries on Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1535–1549, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-153.

Cumming:1994:WDA

[1316] B. F. Cumming, K. A. Davey, J. P. Smol, and H. J. B. Birks. When did acid-sensitive Adirondack lakes (New York, USA) begin to acidify and are they still acidifying? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (7):1550–1568, July 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-154.

Lambert:1994:EIL

[1317] Yvan Lambert, Jean-Denis Dutil, and Jean Munro. Effects of intermediate and low salinity conditions on growth rate and food conversion of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (7):1569–1576, July 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-155.

Clemons:1994:CTE

[1318] J. H. Clemons, M. R. van den Heuvel, J. J. Stegeman, D. G. Dixon, and N. C. Bols. Comparison of toxic equivalent factors for selected dioxin and furan congeners derived using fish and mammalian liver cell lines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1577–1584, July 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-156.

Xiao:1994:BGM

[1319] Yongshun Xiao. von Bertalanffy growth models with variability in, and correlation between, K and L_{∞}. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1585–1590, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-157.

Marzolf:1994:IDU

[1320] Erich R. Marzolf, Patrick J. Mulholland, and Alan D. Steinman. Improvements to the diurnal upstream–downstream dissolved oxygen change technique for determining whole-stream metabolism in small streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1591–1599, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-158. See

comment [2562] and reply [2563].

Grantham:1994:LAH

[1321] Brian A. Grantham and Brenda J. Hann. Leeches (Annelida: Hirudinea) in the Experimental Lakes Area, Northwestern Ontario, Canada: Patterns of species composition in relation to environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1600–1607, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-159.

Boom:1994:MDV

[1322] J. D. G. Boom, E. G. Boulding, and A. T. Beckenbach. Mitochondrial DNA variation in introduced populations of Pacific oyster, Crassostrea gigas, in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (7):1608–1614, July 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-160.

Charpentier:1994:ECV

[1323] Bernadette Charpentier and Antoine Morin. Effect of current velocity on ingestion rates of black fly larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1615–1619, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-161.

Gagen:1994:DMM

[1324] Charles J. Gagen, William E. Sharpe, and Robert F. Carline. Downstream movement and mortality of brook trout (Salvelinus fontinalis) exposed to acidic episodes in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1620–1628, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-162.

Abrahams:1994:RPI

[1325] Mark V. Abrahams. Risk of predation and its influence on the relative competitive abilities of two species of freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1629–1633, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-163.

Thorp:1994:ZAO

[1326] James H. Thorp, A. Ross Black, Kim H. Haag, and John D. Wehr. Zooplankton assemblages in the Ohio River: Seasonal, tributary, and navigation dam effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (7):1634–1643, July 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-164.

Shirvell:1994:ECS

[1327] C. S. Shirvell. Effect of changes in streamflow on the microhabitat use and movements of sympatric juvenile Coho salmon (Oncorhynchus kisutch) and Chinook salmon (O. tshawytscha) in a natural stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1644–1652, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-165.

Smith:1994:DMB

[1328] Thomas G. Smith and A. R. Martin. Distribution and movements of belugas, Delphinapterus leucas, in the Canadian High Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1653–1663, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-166.

Cury:1994:ONE

[1329] Philippe Cury. Obstinate nature: an ecology of individuals. thoughts on reproductive behavior and biodiversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(7):1664–1673, July 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-167. See comment [1852].

Schnute:1994:GFD

[1330] Jon T. Schnute. A general framework for developing sequential fisheries models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1676–1688, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f94-
168.

Hearn:1994:CPB

[1331] William S. Hearn and George M. Leigh. Comparing polynomial and von Bertalanffy growth functions for fitting tag–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1689–1691, August 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-169.

Jensen:1994:INP

[1332] J. P. Jensen, E. Jeppesen, K. Olrik, and P. Kristensen. Impact of nutrients and physical factors on the shift from cyanobacterial to chlorophyte dominance in shallow Danish lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1692–1699, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-170.

Hyllner:1994:IPC

[1333] Sven Johan Hyllner, Birgitta Norberg, and Carl Haux. Isolation, partial characterization, induction, and the occurrence in plasma of the major vitelline envelope proteins in the Atlantic halibut (Hippoglossus hippoglossus) during sexual maturation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1700–1707, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-171.

Fjeld:1994:IEF

[1334] Eirik Fjeld, Sigurd Rognerud, and Eiliv Steinnes. Influence of environmental factors on heavy metal concentration in lake sediments in Southern Norway indicated by path analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1708–1720, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-172.

Ogle:1994:DFA

[1335] D. H. Ogle, G. R. Spangler, and S. M. Shroyer. Determining fish age from temporal signatures in growth increments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1721–1727, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-173.

Powell:1994:EIC

[1336] Mark D. Powell, David J. Speare, and Neil MacNair. Effects of intermittent chloramine-T exposure on growth, serum biochemistry, and fin condition of juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1728–1736, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-174.

Robinson:1994:MPF

[1337] Clifford L. K. Robinson and Daniel M. Ware. Modelling pelagic fish and plankton trophodynamics off Southwestern Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1737–1751, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-175.

Juanes:1994:RGH

[1338] Francis Juanes and David O. Conover. Rapid growth, high feeding rates, and early piscivory in young-of-the-year bluefish (Pomatomus saltatrix). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1752–1761, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-176.

Graves:1994:GAS

[1339] John E. Graves and Jan R. McDowell. Genetic analysis of striped marlin (Tetrapturus audax) population structure in the Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1762–1768, August 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-177.

Jackson:1994:MSP

[1340] Leland J. Jackson, David J. Rowan, R. Jack Cornett, and Jacob Kalff. Myriophyllum spicatum pumps essential and nonessential trace elements from sediments to epiphytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (8):1769–1773, August 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-178.

Cheng:1994:DPS

[1341] Jin-Hua Cheng and Ernest S. Chang. Determinants of postmolt size in the American lobster (Homarus americanus). II. Folding of premolt cuticle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1774–1779, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f94-
179.

Kruzynski:1994:PBQ

[1342] G. M. Kruzynski and I. K. Birtwell. A predation bioassay to quantify the ecological significance of sublethal responses of juvenile Chinook salmon (Oncorhynchus tshawytscha) to the antisapstain fungicide TCMTB. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1780–1790, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-180.

Mitton:1994:CPD

[1343] C. J. A. Mitton and D. G. McDonald. Consequences of pulsed DC electrofishing and air exposure to rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1791–1798, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-181.

Mitton:1994:EEA

[1344] C. J. A. Mitton and D. G. McDonald. Effects of electroshock, air exposure, and forced exercise on swim performance in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8): 1799–1803, August 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-182.

Minns:1994:IBI

[1345] Charles K. Minns, Victor W. Cairns, Robert G. Randall, and James E. Moore. An index of biotic integrity (IBI) for fish assemblages in the littoral zone of Great Lakes’ areas of concern. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1804–1822, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-183.

Hoenig:1994:BRA

[1346] John M. Hoenig, William G. Warren, and Max Stocker. Bayesian and related approaches to fitting surplus production models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1823–1831, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-184.

Hosn:1994:ICS

[1347] Wafa Aboul Hosn and John A. Downing. Influence of cover on the spatial distribution of littoral-zone fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1832–1838, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-185.

Cass:1994:EDF

[1348] A. J. Cass and C. C. Wood. Evaluation of the depensatory fishing hypothesis as an explanation for population cycles in Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1839–1854, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-186.

Davis:1994:SDN

[1349] Ronald B. Davis, Dennis S. Anderson, Stephen A. Norton, Jesse Ford, P. Roger Sweets, and Jeffrey S. Kahl. Sedimented diatoms in Northern New England lakes and their use as pH and alkalinity indicators. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1855–1876, August 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-187.

Hall:1994:RBC

[1350] Ronald J. Hall. Responses of benthic communities to episodic acid disturbances in a lake outflow stream at the Experimental Lakes Area, Ontario.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1877–1892, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-188.

Kjesbu:1994:ORF

[1351] O. S. Kjesbu and J. C. Holm. Oocyte recruitment in first-time spawning Atlantic cod (Gadus morhua) in relation to feeding regime. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1893–1898, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-189.

Bergh:1994:SIM

[1352] Øivind Bergh, Kjell E. Naas, and Torstein Harboe. Shift in the intestinal microflora of Atlantic halibut (Hippoglossus hippoglossus) larvae during first feeding. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8): 1899–1903, August 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-190.

Hu:1994:CRP

[1353] L. L. Hu. Corrections to recent publications/corrections sur récents publications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(8):1904, August 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-191.

Ventling-Schwank:1994:TBC

[1354] Andrea R. Ventling-Schwank and David M. Livingstone. Transport and burial as a cause of whitefish (Coregonus sp.) egg mortality in a eutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1908– 1919, September 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-192.

Bohlin:1994:PFS

[1355] Torgny Bohlin, Claes Dellefors, and Ulo Faremo. Probability of first sexual maturation of male parr in wild sea-run brown trout (Salmo trutta) depends on condition factor 1 yr in advance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1920–1926, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-193.

Gyselman:1994:FAA

[1356] E. C. Gyselman. Fidelity of anadromous Arctic char (Salvelinus alpinus) to Nauyuk Lake, N.W.T., Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1927–1934, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-194.

Adare:1994:SCT

[1357] Katherine I. Adare and David C. Lasenby. Seasonal changes in the total lipid content of the opossum shrimp, Mysis relicta (Malacostraca: Mysidacea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1935– 1941, September 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-195.

Campana:1994:OEF

[1358] Steven E. Campana, Anthony J. Fowler, and Cynthia M. Jones. Otolith elemental fingerprinting for stock identification of Atlantic cod (Gadus morhua) using laser ablation ICPMS. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1942–1950, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-196.

Wagemann:1994:SWD

[1359] R. Wagemann, M. J. Capel, R. Hesslein, and M. Stephenson. Sediment– water distribution coefficients and speciation of cadmium in a Canadian Shield lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1951– 1958, September 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-197.

Brooker:1994:OMD

[1360] Amanda L. Brooker, Doug Cook, Paul Bentzen, Jonathan M. Wright, and Roger W. Doyle. Organization of microsatellites differs between mammals and cold-water teleost fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1959–1966, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-198.

Johansen:1994:SET

[1361] J. A. Johansen, C. J. Kennedy, R. M. Sweeting, A. P. Farrell, and B. A. McKeown. Sublethal effects of tetrachloroguaiacol on juvenile rainbow trout, Oncorhynchus mykiss, following acute and chronic exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1967–1974, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-199.

Kiessling:1994:ISE

[1362] A. Kiessling, D. A. Higgs, B. S. Dosanjh, and J. G. Eales. Influence of sustained exercise at two ration levels on growth and thyroid function of all-female Chinook salmon (Oncorhynchus tshawytscha) in seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1975–1984, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-200.

Morin:1994:CSF

[1363] Pierre-Philippe Morin, Øivind Andersen, Egil Haug, and Kjell B. Døving. Changes in serum free thyroxine, prolactin, and olfactory activity during induced smoltification in Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1985–1992, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-201.

Stahl:1994:ILG

[1364] Thomas P. Stahl and Roy A. Stein. Influence of larval gizzard shad (Dorosoma cepedianum) density on piscivory and growth of young-of-year saugeye (Stizostedion vitreum × S. canadense). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):1993–2002, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-202.

Amyot:1994:ASF

[1365] Marc Amyot, Bernadette Pinel-Alloul, and Peter G. C. Campbell. Abiotic and seasonal factors influencing trace metal levels (Cd, Cu, Ni, Pb, and Zn) in the freshwater amphipod Gammarus fasciatus in two fluvial lakes of the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2003–2016, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-203.

Walton:1994:SRC

[1366] William E. Walton, Stephen S. Easter, Jr., Celeste Malinoski, and Nelson G. Hairston, Jr. Size-related change in the visual resolution of sunfish (Lepomis spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2017– 2026, September 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-204.

Mullen:1994:EMS

[1367] Ashley J. Mullen. Effects of movement on stock assessment in a restricted-range fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9): 2027–2033, September 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-205.

Vezina:1994:IMA

[1368] Alain F. Vézina and Michael L. Pace. An inverse model analysis of planktonic food webs in experimental lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2034–2044, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-206.

Holmes:1994:LTI

[1369] J. A. Holmes, F. W. H. Beamish, J. G. Seelye, S. A. Sower, and J. H. Youson. Long-term influence of water temperature, photoperiod, and food deprivation on metamorphosis of sea lamprey, Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2045–2051, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-207.

Chow-Fraser:1994:THE

[1370] P. Chow-Fraser, D. O. Trew, D. Findlay, and M. Stainton. A test of hypotheses to explain the sigmoidal relationship between total phosphorus and chlorophyll a concentrations in Canadian lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2052–2065, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-208.

Chen:1994:MWS

[1371] Y. Chen and H. H. Harvey. Maturation of white sucker, Catostomus commersoni, populations in Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2066–2076, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-209.

Johnston:1994:FEW

[1372] Thomas A. Johnston and J. A. Mathias. Feeding ecology of walleye, Stizostedion vitreum, larvae: Effects of body size, zooplankton abundance, and zooplankton community composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2077–2089, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-210.

Parks:1994:RBM

[1373] J. W. Parks, P. C. Craig, and G. W. Ozburn. Relationships between mercury concentrations in walleye (Stizostedion vitreum) and northern pike (Esox lucius): Implications for modelling and biomonitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2090–2104, September 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-211.

Jackson:1994:PMM

[1374] A. E. Jackson, A. S. W. deFreitas, L. Hooper, A. Mallet, and J. A. Walter. Phosphorus metabolism monitored by ^{31}P NMR in juvenile sea scallop (Placopecten magellanicus) overwintering in pearl nets at a Nova Scotian aquaculture site. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2105– 2114, September 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-212.

McDaniels:1994:CFM

[1375] Timothy L. McDaniels, Michael Healey, and Richard Kyle Paisley. Cooperative fisheries management involving First Nations in British Columbia: an adaptive approach to strategy design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2115–2125, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-213.

Hutchings:1994:WCL

[1376] Jeffrey A. Hutchings and Ransom A. Myers. What can be learned from the collapse of a renewable resource? Atlantic cod, Gadus morhua, of Newfoundland and Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2126–2146, September 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-214.

Anonymous:1994:LCLc

[1377] Anonymous. Letters and comments/lettres et commentaires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(9):2147–2151, September 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-215.

Hughes:1994:WDA

[1378] Nicholas F. Hughes and James B. Reynolds. Why do Arctic grayling (Thymallus arcticus) get bigger as you go upstream? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2154–2163, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-216.

Whalen:1994:SGA

[1379] K. G. Whalen and G. W. LaBar. Survival and growth of Atlantic salmon (Salmo salar) fry stocked at varying densities in the White River, Vermont. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2164–2169, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f94-

Shrimpton:1994:DMS

[1380] J. Mark Shrimpton, Nicholas J. Bernier, George K. Iwama, and David J. Randall. Differences in measurements of smolt development between wild and hatchery-reared juvenile Coho salmon (Oncorhynchus kisutch) before and after saltwater exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2170–2178, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-218.

Shrimpton:1994:CCD

[1381] J. Mark Shrimpton, Nicholas J. Bernier, and David J. Randall. Changes in cortisol dynamics in wild and hatchery-reared juvenile Coho salmon (Oncorhynchus kisutch) during smoltification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2179–2187, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-219.

Brauner:1994:ESD

[1382] C. J. Brauner, G. K. Iwama, and D. J. Randall. The effect of shortduration seawater exposure on the swimming performance of wild and hatchery-reared juvenile Coho salmon (Oncorhynchus kisutch) during smoltification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2188– 2194, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f94-220.

Sandstrom:1994:IRC

[1383] Olof Sandström. Incomplete recovery in a coastal fish community exposed to effluent from a modernized Swedish bleached kraft mill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2195–2202, October 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-221.

Barker:1994:BSW

[1384] D. E. Barker, R. A. Khan, and R. Hooper. Bioindicators of stress in winter flounder, Pleuronectes americanus, captured adjacent to a pulp and paper mill in St. George’s Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2203–2209, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-222.

Muotka:1994:DSS

[1385] Timo Muotka and Antti Penttinen. Detecting small-scale spatial patterns in lotic predator–prey relationships: Statistical methods and a case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2210–2218, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-223.

Robarts:1994:CHB

[1386] Richard D. Robarts, Michael T. Arts, Marlene S. Evans, and Marley J. Waiser. The coupling of heterotrophic bacterial and phytoplankton production in a hypertrophic, shallow prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2219–2226, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-224.

Nalepa:1994:DNU

[1387] Thomas F. Nalepa. Decline of native unionid bivalves in Lake St. Clair after infestation by the zebra mussel, Dreissena polymorpha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2227–2233, October 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-225.

Schloesser:1994:DDU

[1388] Don W. Schloesser and Thomas F. Nalepa. Dramatic decline of unionid bivalves in offshore waters of Western Lake Erie after infestation by the zebra mussel, Dreissena polymorpha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2234–2242, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-226.

Hecky:1994:AEL

[1389] R. E. Hecky, D. M. Rosenberg, and P. Campbell. The 25th anniversary of the Experimental Lakes Area and the history of Lake 227. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2243–2246, October 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-227.

Hendzel:1994:RCF

[1390] L. L. Hendzel, R. E. Hecky, and D. L. Findlay. Recent changes of N_{2}- fixation in Lake 227 in response to reduction of the N : P loading ratio. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2247–2253, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-228.

Findlay:1994:RBF

[1391] D. L. Findlay, R. E. Hecky, L. L. Hendzel, M. P. Stainton, and G. W. Regehr. Relationship between N_{2}-fixation and heterocyst abundance and its relevance to the nitrogen budget of Lake 227. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2254–2266, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-229.

Kling:1994:ASS

[1392] Hedy J. Kling, David L. Findlay, and Jiri Komárek. Aphanizomenon schindleri sp.nov.: a new nostocacean cyanoprokaryote from the Experimental Lakes Area, northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2267–2273, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-230.

Wolfe:1994:MDF

[1393] Brent Wolfe, Hedy J. Kling, Gregg J. Brunskill, and Paul Wilkinson. Multiple dating of a freeze core from Lake 227, an experimentally fertilized lake with varved sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2274–2285, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-231.

Leavitt:1994:CFP

[1394] P. R. Leavitt and D. L. Findlay. Comparison of fossil pigments with 20 years of phytoplankton data from eutrophic Lake 227, Experimental Lakes Area, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10): 2286–2299, October 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-232.

Zeeb:1994:RDC

[1395] Barbara A. Zeeb, Catherine E. Christie, John P. Smol, David L. Findlay, Hedy J. Kling, and H. J. B. Birks. Responses of diatom and chrysophyte assemblages in Lake 227 sediments to experimental eutrophication. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2300–2311, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-233.

Hann:1994:CCR

[1396] Brenda J. Hann, Peter R. Leavitt, and Philip S. S. Chang. Cladocera community response to experimental eutrophication in Lake 227 as recorded in laminated sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2312–2321, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-234.

Leavitt:1994:PAW

[1397] P. R. Leavitt, B. J. Hann, J. P. Smol, B. A. Zeeb, C. E. Christie, B. Wolfe, and H. J. Kling. Paleolimnological analysis of whole-lake experiments: an overview of results from Experimental Lakes Area Lake 227. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2322–2332, October 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-235.

West:1994:DFA

[1398] I. F. West and R. W. Gauldie. Determination of fish age using ^{210}Pb : ^{226}Ra disequilibrium methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2333–2340, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-236.

Gauldie:1994:MBF

[1399] R. W. Gauldie. The morphological basis of fish age estimation methods based on the otolith of Nemadactylus macropterus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10):2341–2362, October 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-237.

Pinkerton:1994:LFC

[1400] Evelyn W. Pinkerton. Local fisheries co-management: a review of international experiences and their implications for salmon management in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(10): 2363–2378, October 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-238.

Sutton:1994:SOD

[1401] Trent M. Sutton and Stephen H. Bowen. Significance of organic detritus in the diet of larval lampreys in the Great Lakes Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2380–2387, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-239.

Rowan:1994:BRF

[1402] David J. Rowan and Joseph B. Rasmussen. Bioaccumulation of radiocesium by fish: the influence of physicochemical factors and trophic structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2388–2410, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-240.

Leavitt:1994:FPR

[1403] P. R. Leavitt, D. E. Schindler, A. J. Paul, A. K. Hardie, and D. W. Schindler. Fossil pigment records of phytoplankton in trout-stocked alpine lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2411– 2423, November 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-241.

Keller:1994:CZC

[1404] W. Keller and M. Conlon. Crustacean zooplankton communities and lake morphometry in Precambrian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2424–2434, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-242.

Stemberger:1994:ZAR

[1405] Richard S. Stemberger and James M. Lazorchak. Zooplankton assemblage responses to disturbance gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2435–2447, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-243.

Munro:1994:PRA

[1406] J. Munro, C. Audet, M. Besner, and J.-D. Dutil. Physiological response of American plaice (Hippoglossoides platessoides) exposed to low salinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2448–2456, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-244.

Poister:1994:RNE

[1407] David Poister, David E. Armstrong, and James P. Hurley. A 6-yr record of nutrient element sedimentation and recycling in three North temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2457–2466, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-245.

Hudon:1994:BED

[1408] Christiane Hudon. Biological events during ice breakup in the Great Whale River (Hudson Bay). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (11):2467–2481, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f94-246.

Zia:1994:RGG

[1409] S. Zia and D. G. McDonald. Role of the gills and gill chloride cells in metal uptake in the freshwater-adapted rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2482–2492, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-247.

Yang:1994:HPC

[1410] C. Z. Yang and L. J. Albright. The harmful phytoplankter Chaetoceros concavicornis causes high mortalities and leucopenia in Chinook salmon (Oncorhynchus tshawytscha) and Coho salmon (O. kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2493–2500, November 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-248.

Post:1994:VFY

[1411] John R. Post and Donald J. McQueen. Variability in first-year growth of yellow perch (Perca flavescens): Predictions from a simple model, observations, and an experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2501–2512, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-249.

Ryan:1994:NDL

[1412] Phil A. Ryan and Terry R. Marshall. A niche definition for lake trout (Salvelinus namaycush) and its use to identify populations at risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2513–2519, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-250.

Paul:1994:RRP

[1413] Andrew J. Paul and David W. Schindler. Regulation of rotifers by predatory calanoid copepods (subgenus Hesperodiaptomus) in lakes of the Canadian Rocky Mountains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2520–2528, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-251.

Tranvik:1994:MAA

[1414] Lars J. Tranvik, Wilhelm Granéli, and Gunnar Gahnström. Microbial activity in acidified and limed humic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2529–2536, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-252.

Sampson:1994:ENF

[1415] David B. Sampson. Estimating the number of fish landed from their total weight c and a sample average weight. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2537–2548, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-253.

Forrester:1994:DPD

[1416] Graham E. Forrester. Diel patterns of drift by five species of mayfly at different levels of fish predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2549–2557, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-254.

Trudel:1994:SCD

[1417] Marc Trudel and Daniel Boisclair. Seasonal consumption by dace (Phoxinus eos × P. neogaeus): a comparison between field and bioenergetic model estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2558– 2567, November 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-255.

Sprules:1994:IMT

[1418] W. Gary Sprules, Stephen B. Brandt, and Mohiuddin Munawar. Introduction: Multiple trophic level comparisons of Lakes Michigan and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2568–2569, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-256.

Johengen:1994:TST

[1419] T. H. Johengen, Ora E. Johannsson, G. Laird Pernie, and E. S. Millard. Temporal and seasonal trends in nutrient dynamics and biomass measures in lakes Michigan and Ontario in response to phosphorus control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2570–2578, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-257.

Millard:1994:CPL

[1420] E. S. Millard and P. E. Sager. Comparison of phosphorus, light climate, and photosynthesis between two culturally eutrophied bays: Green Bay, Lake Michigan, and the Bay of Quinte, Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2579–2590, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-258.

Johannsson:1994:MRA

[1421] Ora E. Johannsson, Lars G. Rudstam, and David C. Lasenby. Mysis relicta: Assessment of metalimnetic feeding and implications for competition with fish in lakes Ontario and Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2591–2602, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-259.

Sprules:1994:SBS

[1422] W. Gary Sprules and Andy P. Goyke. Size-based structure and production in the pelagia of Lakes Ontario and Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2603–2611, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-260.

Jackson:1994:AFP

[1423] George D. Jackson. Application and future potential of statolith increment analysis in squids and sepioids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2612–2625, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-261.

Gowan:1994:RMR

[1424] Charles Gowan, Michael K. Young, Kurt D. Fausch, and Stephen C. Riley. Restricted movement in resident stream salmonids: a paradigm lost? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(11):2626–2637, November 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-262.

Richards:1994:RDQ

[1425] Laura J. Richards and Bernard A. Megrey. Recent developments in the quantitative analysis of fisheries data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2640–2641, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-263.

Kleiber:1994:MEF

[1426] P. Kleiber and J. Hampton. Modeling effects of FADs and islands on movement of skipjack tuna (Katsuwonus pelamis): Estimating parameters from tagging data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12): 2642–2653, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-264.

Thompson:1994:CGS

[1427] Grant G. Thompson. Confounding of gear selectivity and the natural mortality rate in cases where the former is a nonmonotone function of age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2654–2664, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-265.

Collie:1994:MPP

[1428] Jeremy S. Collie and Paul D. Spencer. Modeling predator–prey dynamics in a fluctuating environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2665–2672, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-266.

McAllister:1994:BAS

[1429] Murdoch K. McAllister, Ellen K. Pikitch, Andre E. Punt, and Ray Hilborn. A Bayesian approach to stock assessment and harvest decisions using the sampling/importance resampling algorithm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2673–2687, December 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-267.

Sampson:1994:FTT

[1430] David B. Sampson. Fishing tactics in a two-species fisheries model: The bioeconomics of bycatch and discarding. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2688–2694, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-268.

Megrey:1994:SOH

[1431] Bernard A. Megrey, Anne Babcock Hollowed, and Rebecca T. Baldwin. Sensitivity of optimum harvest strategy estimates to alternative definitions of risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2695– 2704, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-269.

Walters:1994:UGP

[1432] Carl Walters. Use of gaming procedures in evaluation of management experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2705– 2714, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-270.

Rosenberg:1994:URE

[1433] Andrew A. Rosenberg and Victor R. Restrepo. Uncertainty and risk evaluation in stock assessment advice for U.S. marine fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2715–2720, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-271.

Hecky:1994:IEL

[1434] R. E. Hecky, P. Campbell, and D. M. Rosenberg. Introduction to experimental lakes and natural processes: 25 years of observing natural ecosystems at the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2721–2722, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-272.

Beaty:1994:STS

[1435] K. G. Beaty. Sediment transport in a small stream following two successive forest fires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2723– 2733, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-273.

Schindler:1994:SES

[1436] E. U. Schindler, E. R. DeBruyn, E. J. Fee, and J. A. Shearer. Sensitivity of estimates of seasonal phytoplankton photosynthesis to sampling frequency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2734–2738, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f94-274.

Campbell:1994:PBS

[1437] P. Campbell. Phosphorus budgets and stoichiometry during the openwater season in two unmanipulated lakes in the Experimental Lakes Area, Northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12): 2739–2755, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f94-275. See comment [1931].

Fee:1994:ELS

[1438] E. J. Fee, R. E. Hecky, G. W. Regehr, L. L. Hendzel, and P. Wilkinson. Effects of lake size on nutrient availability in the mixed layer during summer stratification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12): 2756–2768, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-276.

Guildford:1994:ELS

[1439] S. J. Guildford, L. L. Hendzel, H. J. Kling, E. J. Fee, G. G. C. Robinson, R. E. Hecky, and S. E. M. Kasian. Effects of lake size on phytoplankton nutrient status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12): 2769–2783, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-277.

Turner:1994:RNC

[1440] Michael A. Turner, E. Todd Howell, Gordon G. C. Robinson, Paul Campbell, Robert E. Hecky, and Eva U. Schindler. Roles of nutrients in controlling growth of epilithon in oligotrophic lakes of low alkalinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2784–2793, December 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-278.

Findlay:1994:BLE

[1441] D. L. Findlay, S. E. M. Kasian, L. L. Hendzel, G. W. Regehr, E. U. Schindler, and J. A. Shearer. Biomanipulation of Lake 221 in the Experimental Lakes Area (ELA): Effects on phytoplankton and nutrients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2794–2807, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f94-279.

Fleming:1994:PDS

[1442] Ian A. Fleming, Bror Jonsson, and Mart R. Gross. Phenotypic divergence of sea-ranched, farmed, and wild salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2808–2824, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f94-280.

Hansson:1994:PAR

[1443] Lars-Anders Hansson, Lars G. Rudstam, Timothy B. Johnson, Patricia Soranno, and Yvonne Allen. Patterns in algal recruitment from sediment to water in a dimictic, eutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2825–2833, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-281.

Goddard:1994:OAN

[1444] Sally V. Goddard, J. S. Wroblewski, C. T. Taggart, K. A. Howse, W. L. Bailey, M. H. Kao, and G. L. Fletcher. Overwintering of adult Northern Atlantic cod (Gadus morhua) in cold inshore waters as evidenced by plasma antifreeze glycoprotein levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2834–2842, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-282.

Beamish:1994:RBF

[1445] Richard J. Beamish, Chrys-Ellen M. Neville, Barbara L. Thomson, Paul J. Harrison, and Mike St. John. A relationship between Fraser River discharge and interannual production of Pacific salmon (Oncorhynchus spp.) and Pacific herring (Clupea pallasi) in the Strait of Georgia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2843–2855, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-283.

Magnan:1994:DVF

[1446] Pierre Magnan, Marco A. Rodríguez, Pierre Legendre, and Sylvain Lacasse. Dietary variation in a freshwater fish species: Relative contributions of biotic interactions, abiotic factors, and spatial structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2856–2865, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f94-284.

Miloslavich:1994:DEF

[1447] Patricia Miloslavich and Louise Dufresne. Development and effect of female size on egg and juvenile production in the neogastropod Buccinum cyaneum from the Saguenay Fjord. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12):2866–2872, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f94-285.

Chick:1994:PAC

[1448] John H. Chick and Carole C. Mclvor. Patterns in the abundance and composition of fishes among beds of different macrophytes: Viewing a littoral zone as a landscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(12): 2873–2882, December 1994. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f94-286.

Gharrett:1994:IGS

[1449] A. J. Gharrett and W. W. Smoker. Introduction to genetics of subarctic fish and shellfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):1–3, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Hartl:1994:KAM

[1450] Daniel L. Hartl. Keynote address: Macro and micro in molecular evolution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):4–8, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Smoker:1994:GAS

[1451] William W. Smoker, Anthony J. Gharrett, Michael S. Stekoll, and John E. Joyce. Genetic aanalysis of size in an anadromous population of pink salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):9–15, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Galbreath:1994:VFP

[1452] Peter F. Galbreath and Gary H. Thorgaard. Viability and freshwater performance of Atlantic salmon (Salmo salar) × brown trout (Salmo trutta) triploid hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):16–24, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Joyce:1994:SFS

[1453] John E. Joyce, Ron Heintz, William W. Smoker, and A. J. Gharrett. Survival to fry and seawater tolerance of diploid and triploid hybrids between Chinook (Oncorhynchus tshawytscha), chum (O. keta), and pink salmon (O. gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):25–30, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Habicht:1994:TCS

[1454] Christopher Habicht, James E. Seeb, Richard B. Gates, Irvin R. Brock, and Carmen A. Olito. Triploid Coho salmon outperform diploid and triploid hybrids between Coho salmon and Chinook salmon during their first year. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):31–37, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Teplitz:1994:PPA

[1455] R. L. Teplitz, J. E. Joyce, S. I. Doroshov, and B. H. Min. A preliminary ploidy analysis of diploid and triploid salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):38–41, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Miller:1994:SEF

[1456] Gary D. Miller, James E. Seeb, Brian G. Bue, and Samuel Sharr. Saltwater exposure at fertilization induces ploidy alterations. including mosaicism, in salmonid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51 (S1):42–49, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Kondzela:1994:GRA

[1457] C. M. Kondzela, C. M. Guthrie, S. L. Hawkins, C. d. Russell, J. H. Helle, and A. J. Gharrett. Genetic relationships among chum salmon populations in southeast Alaska and northern British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):50–64, December 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Phelps:1994:GDP

[1458] S. R. Phelps, L. L. LeClair, S. Young, and H. L. Blankenship. Genetic diversity patterns of chum salmon in the Pacific Northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):65–83, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Wilmot:1994:GSS

[1459] Richard L. Wilmot, Rebecca J. Everett, William J. Spearman, Ramone Baccus, et al. Genetic stock structure of western Alaska chum salmon and a comparison with Russian Far East stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):84–94, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Winans:1994:DCO

[1460] Gary A. Winans, Paul B. Aebersold, Shigehiko Urawa, and Nataly V. Varnavskaya. Determining continent of origin of chum salmon (Oncorhynchus keta) using genetic stock identification techniques: Status of allozyme baseline in Asia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):95–113, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Wood:1994:BGS

[1461] Chris C. Wood, Brian E. Riddell, Dennis T. Rutherford, and Ruth E. Withler. Biochemical genetic survey of sockeye salmon (Oncorhynchus nerka) in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):114–131, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Varnavskaya:1994:GVS

[1462] Nataly V. Varnavskaya, Chris C. Wood, and Rebecca J. Everett. Genetic variation in sockeye salmon (Oncorhynchus nerka) populations of Asia and North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):132– 146, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Varnavskaya:1994:GDS

[1463] Nataly V. Varnavskaya, Chris C. Wood, Rebecca J. Everett, Richard L. Wilmot, et al. Genetic differentiation of subpopulations of sockeye salmon (Oncorhynchus nerka) within lakes of Alaska, British Columbia, and Kamchatka, Russia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1): 147–157, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Shaklee:1994:ECO

[1464] James B. Shaklee and Natalya V. Varnavskaya. Electrophoretic characterization of odd-year pink salmon (Oncorhynchus gorbuscha) populations from the Pacific coast of Russia, and comparison with selected North American populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):158–171, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). See erratum [1779].

Adams:1994:VMD

[1465] Noah S. Adams, William J. Spearman, Carl V. Burger, Kenneth P. Currens, Carl B. Schreck, and Hiram W. Li. Variation in mitochondrial DNA and allozymes discriminates early and late forms of Chinook salmon (Oncorhynchus tshawytscha) in the Kenai and Kasilof Rivers, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):172–181, December 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Crane:1994:GRA

[1466] Penelope A. Crane, Lisa W. Seeb, and James E. Seeb. Genetic relationships among Salvelinus species inferred from allozyme data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):182–197, December 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Phillips:1994:SSG

[1467] Ruth B. Phillips, Susan A. Manley, and Thomas J. Daniels. Systematics of the salmonid genus Salvelinus inferred from ribosomal DNA sequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):198–204, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Gold:1994:MDD

[1468] John R. Gold, Linda R. Richardson, Carol Furman, and Feng Sun. Mitochondrial DNA diversity and population structure in marine fish species from the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1): 205–214, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

King:1994:PSE

[1469] Tim L. King, Rocky Ward, and Earl G. Zimmerman. Population structure of eastern oysters (Crassostrea virginica) inhabiting the Laguna Madre, Texas, and adjacent bay systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):215–222, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Zhivotovsky:1994:GDP

[1470] Lev A. Zhivotovsky, A. J. Gharrett, A. J. McGregor, M. K. Glubokovsky, and Marcus W. Feldman. Gene differentiation in Pacific salmon (Oncorhynchus sp.): Facts and models with reference to pink salmon (O. gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):223–232, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Jorstad:1994:GCB

[1471] K. E. Jørstad, G. Dahle, and O. I. Paulsen. Genetic comparison between Pacific herring (Clupea pallasi) and a Norwegian fjord stock of Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):233–239, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Bernatchez:1994:PRA

[1472] Louis Bernatchez and Julian J. Dodson. Phylogenetic relationships among Palearctic and Nearctic whitefish (Coregonus sp.) populations as revealed by mitochondrial DNA variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):240–251, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). See erratum [1632].

Spruell:1994:CTA

[1473] Paul Spruell, Shawn A. Cummings, Yousok Kim, and Gary H. Thorgaard. Comparison of three anadromous rainbow trout (Oncorhynchus mykiss) populations using DNA fingerprinting and mixed DNA samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):252–257, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Leung:1994:DFC

[1474] F. C. Leung, M. Welt, R. D. Quesenberry, and X-Z. Shen. DNA fingerprinting and cloning of hypervariable minisatellite repeats in salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):258–266, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Withler:1994:IFR

[1475] R. E. Withler, T. D. Beacham, R. F. Watkins, and T. A. Stevens. Identification of farm-reared and native Chinook salmon (Oncorhynchus tshawytscha) on the West Coast of Vancouver Island, British Columbia, using the nuclear DNA probe B2-2. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):267–276, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Hartley:1994:DSD

[1476] S. E. Hartley and W. S. Davidson. Distribution of satellite DNA sequences isolated from Arctic char, Salvelinus alpinus, in the genus Salvelinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):277–283, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Danzmann:1994:HDM

[1477] Roy G. Danzmann, Moira M. Ferguson, and David M. Heculuck. Heterogeneity in the distribution of mitochondrial DNA haplotypes in female rainbow trout spawning in different seasons. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):284–289, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Nielsen:1994:DGD

[1478] Jennifer L. Nielsen, Christina Gan, and W. Kelley Thomas. Differences in genetic diversity for mitochondrial DNA between hatchery and wild populations of Oncorhynchus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):290–297, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Gresswell:1994:LHO

[1479] Robert E. Gresswell, William J. Liss, and Gary L. Larson. Life-history organization of Yellowstone cutthroat trout (Oncorhynchus clarki bouvieri) in Yellowstone Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1): 298–309, December 1994. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Waples:1994:GRA

[1480] Robin S. Waples and Chi Do. Genetic risk associated with supplementation of Pacific salmonids: Captive broodstock programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 51(S1):310–329, December 1994. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Levy:1995:STS

[1481] David A. Levy and Allen D. Cadenhead. Selective tidal stream transport of adult sockeye salmon (Oncorhynchus nerka) in the Fraser River Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):1–12, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-001.

Handy:1995:CIC

[1482] R. D. Handy. Comparison of intermittent and continuous exposure to mercuric chloride in rainbow trout (Oncothynchus mykiss), goldfish (Carassius auratus), and the fathead minnow (Pimephales ptomelas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):13–22, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-002.

Sheldon:1995:PAC

[1483] Andrew L. Sheldon and Gary K. Meffe. Path analysis of collective properties and habitat relationships of fish assemblages in coastal plain streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):23–33, January 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-003.

Riley:1995:TPR

[1484] Stephen C. Riley and Kurt D. Fausch. Trout population response to habitat enhancement in six northern Colorado streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):34–53, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-004.

Chang:1995:EAN

[1485] Cecily C. Y. Chang and Richard Petersen. Evidence of autumn nitrogen limitation and contribution of picoplankton to carbon fixation in Lake Tahoe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):54–62, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-005.

Vanderploeg:1995:PMT

[1486] Henry A. Vanderploeg, James R. Liebig, and Thomas F. Nalepa. From picoplankton to microplankton: temperature-driven filtration by the unionid bivalve Lampsilis radiata siliquoidea in Lake St. Clair. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):63–74, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-006.

Hansson:1995:GCE

[1487] Sture Hansson and Lars G. Rudstam. Gillnet catches as an estimate of fish abundance: a comparison between vertical gillnet catches and hydroacoustic abundances of Baltic Sea herring (Clupea harengus) and sprat (Sptattus sptattus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (1):75–83, January 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-007.

Hayes:1995:EBE

[1488] Daniel B. Hayes, Jon K. T. Brodziak, and Joseph B. O’Gorman. Efficiency and bias of estimators and sampling designs for determining length–weight relationships of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):84–92, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub. com/doi/full/10.1139/f95-008. See comment [2153] and reply [2154].

Elser:1995:FAI

[1489] J. J. Elser, F. S. Lubnow, E. R. Marzolf, M. T. Brett, G. Dion, and C. R. Goldman. Factors associated with interannual and intraannual variation in nutrient limitation of phytoplankton growth in Castle Lake, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):93–104, January 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-009.

Kocik:1995:EJS

[1490] John F. Kocik and William W. Taylor. Effect of juvenile steelhead (Oncorhynchus mykiss) on age-o and age-1 brown trout (Salmo trutta) survival and growth in a sympatric nursery stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):105–114, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-444.

Anderson:1995:ADP

[1491] J. T. Anderson, E. L. Dalley, and J. E. Carscadden. Abundance and distribution of pelagic 0-group cod (Gadus morhua) in Newfoundland waters: inshore versus offshore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (1):115–125, January 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-011.

Ackerman:1995:WJM

[1492] Josef Daniel Ackerman, C. Ross Ethier, Jan K. Spelt, D. Grant Allen, and Catherine M. Cottrell. A wall jet to measure the attachment strength of zebra mussels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):126–135, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-012.

Marschall:1995:DDS

[1493] Elizabeth A. Marschall and Larry B. Crowder. Density-dependent survival as a function of size in juvenile salmonids in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):136–140, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/f95-013. See erratum [1734].

Yurista:1995:BAP

[1494] Peder M. Yurista and Kimberly L. Schulz. Bioenergetic analysis of prey consumption by Bythotrephes cederstroemi in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):141–150, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-014.

Havel:1995:IEC

[1495] John E. Havel, William R. Mabee, and John R. Jones. Invasion of the exotic cladoceran Daphnia lumholtzi into North American reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):151–160, January 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-015.

McClelland:1995:LCT

[1496] Grant McClelland, Jean-Michel Weber, Georges Zwingelstein, and Gérard Brichon. Lipid composition off tissue and plasma in two Mediterranean fishes, the gilt-head sea bream (Chrysophyrys auratus) and the European seabass (Dicentratchus labrx). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):161–170, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-016.

OConnell:1995:MDA

[1497] M. O’Connell, D. O. F. Skibinski, and J. A. Beardmore. Mitochondrial DNA and allozyme variation in Atlantic salmon (Salmo salar) populations in Wales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):171–178, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-017.

Bernatchez:1995:IFA

[1498] Louis Bernatchez, Hélène Glémet, Chris C. Wilson, and Roy G. Danzmann. Introgression and fixation of Arctic char (Salvelinus alpinus) mitochondrial genome in an allopatric population of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):179–185, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
018.

Keeley:1995:AEC

[1499] Ernest R. Keeley and James W. A. Grant. Allometric and environmental correlates of territory size in juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):186–196, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-019.

Clair:1995:RPS

[1500] T. A. Clair, P. J. Dillon, J. Ion, M. Papineau, D. S. Jeffries, and R. J. Vet. Regional precipitation and surface water chemistry trends in southeastern Canada (1983–1991). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (1):197–212, January 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-020.

Ensign:1995:ULT

[1501] William E. Ensign, Paul L. Angermeier, and C. Andrew Dolloff. Use of line transect methods to estimate abundance of benthic stream fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):213–222, January 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-021.

Myers:1995:TSB

[1502] Ransom A. Myers and N. J. Barrowman. Time series bias in the estimation of density-dependent mortality in stock-recruitment models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(1):223–232, January 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-022.

Worthington:1995:VRB

[1503] D. O. Worthington, P. J. Doherty, and A. J. Fowler. Variation in the relationship between otolith weight and age: implications for the estimation of age of two tropical damselfish (Pomacentrus moluccensis and P. wardi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):233–242, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-023.

Berge:1995:DST

[1504] Åse I. Berge, Arne Berg, Trygg Barnung, Tom Hansen, Hans Jørgen Fyhn, and Sigurd O. Stefansson. Development of salinity tolerance in underyearling smolts of Atlantic salmon (Salmo salar) reared under different photoperiods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2): 243–251, February 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-024.

Wang:1995:MLA

[1505] You-Gan Wang, Mervyn R. Thomas, and Ian F. Somers. A maximum likelihood approach for estimating growth from tag–recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):252–259, February 1995.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-025.

Somerton:1995:SST

[1506] David A. Somerton and Bert S. Kikkawa. A stock survey technique using the time to capture individual fish on longlines. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):260–267, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-026.

Renaud:1995:HVR

[1507] Claud B. Renaud, Klaus L. E. Kaiser, and Michael E. Comba. Historical versus recent levels of organochlorine contaminants in lamprey larvae of the St. Lawrence River Basin, Québec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):268–275, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-027.

Renaud:1995:CBL

[1508] Claude B. Renaud, Klaus L. E. Kaiser, Michael E. Comba, and Janice L. Metcalfe-Smith. Comparison between lamprey ammocoetes and bivalve molluscs as biomonitors of organochlorine contaminants. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):276–282, February 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-028.

Booth:1995:ELS

[1509] Richard K. Booth, James D. Kieffer, Bruce L. Tufts, Kevin Davidson, and Alex T. Bielak. Effects of late-season catch and release angling on anaerobic metabolism, acid–base status, survival, and gamete viability in wild Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):283–290, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-029.

Frederick:1995:CFH

[1510] Shane W. Frederick and Randall M. Peterman. Choosing fisheries harvest policies: when does uncertainty matter? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):291–306, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-030.

Ferguson:1995:GEC

[1511] Moira M. Ferguson, Arunas P. Liskauskas, and Roy G. Danzmann. Genetic and environmental correlates of variation in body weight of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (2):307–314, February 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-031.

Miller:1995:TIC

[1512] R. J. Miller and J. T. Addison. Trapping interactions of crabs and American lobster in laboratory tanks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):315–324, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-032.

Hsieh:1995:WIU

[1513] William W. Hsieh, Daniel M. Ware, and Richard E. Thomson. Windinduced upwelling along the west coast of North America, 1899–1988. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):325–334, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-033.

Banneheka:1995:ERF

[1514] Sarath G. Banneheka, Richard D. Routledge, Ian C. Guthrie, and James C. Woodey. Estimation of in-river fish passage using a combination of transect and stationary hydroacoustic sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):335–343, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-034.

Zhang:1995:DOM

[1515] Z. Zhang, R. J. Beamish, and B. E. Riddell. Differences in otolith microstructure between hatchery-reared and wild Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (2):344–352, February 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-035.

Joseph:1995:EAF

[1516] Elaine M. Joseph, François M. M. Morel, and Neil M. Price. Effects of aluminum and fluoride on phosphorus acquisition by Chlamydomonas reinhardtii. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):353–357, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-036.

Janssen:1995:ULL

[1517] John Janssen, Warren R. Jones, Audrey Whang, and Philip E. Oshel. Use of the lateral line in particulate feeding in the dark by juvenile alewife (Alosa pseudoharengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (2):358–363, February 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-037.

Hoenig:1995:ADB

[1518] J. M. Hoenig, M. J. Morgan, and C. A. Brown. Analysing differences between two age determination methods by tests of symmetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):364–368, February 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-038.

Swartzman:1995:RTW

[1519] Gordon Swartzman, Emily Silverman, and Neal Williamson. Relating trends in walleye pollock (Theragra chalcogramma) abundance in the Bering Sea to environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):369–380, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-039.

Madenjian:1995:RAZ

[1520] Charles P. Madenjian. Removal of algae by the zebra mussel (Dreissena polymorpha) population in western Lake Erie: a bioenergetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):381–390, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-040.

Smith:1995:ADG

[1521] David C. Smith, Simon G. Robertson, Gwen E. Fenton, and Stephen A. Short. Age determination and growth of orange roughy (Hoplostethus atlanticus): a comparison of annulus counts with radiometric ageing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):391–401, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-041.

Gillis:1995:ITR

[1522] Darren M. Gillis, Randall M. Peterman, and Ellen K. Pikitch. Implications of trip regulations for high-grading; a model of the behavior of fishermen. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):402–415, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-042.

LaBaugh:1995:RAB

[1523] James W. LaBaugh. Relation of algal biovolume to chlorophyll a in selected lakes and wetlands in the north-central United States. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(2):416–424, February 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-043.

Stites:1995:PDG

[1524] David L. Stites, Arthur C. Benke, and David M. Gillespie. Population dynamics, growth, and production of the Asiatic clam, Corbicula fluminea, in a blackwater river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (2):425–437, February 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-044.

Bankey:1995:RCP

[1525] Laura A. Bankey, Peter A. Van Veld, Dennis L. Borton, Larry LaFleur, and John J. Stegeman. Responses of cytochrome P4501A in freshwater fish exposed to bleached kraft mill effluent in experimental stream channels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):439–447, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-045.

McLaren:1995:PFD

[1526] Ian A. McLaren and Patricia Avendaño. Prey field and diet of larval cod on Western Bank, Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):448–463, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-046a.

Koel:1995:SHF

[1527] Todd M. Koel and John J. Peterka. Survival to hatching of fishes in sulfate-saline waters, Devils Lake, North Dakota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):464–469, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-047.

Ricciardi:1995:AET

[1528] Anthony Ricciardi, Robert Serrouya, and Frederick G. Whoriskey. Aerial exposure tolerance off zebra and quagga mussels (Bivalvia: Dreissenidae): implications for overland dispersal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):470–477, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-048.

Hammill:1995:GCN

[1529] M. O. Hammill, M. C. S. Kingsley, G. G. Beck, and T. G. Smith. Growth and condition in the Northwest Atlantic harp seal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):478–488, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-049.

Welch:1995:UTL

[1530] D. W. Welch, A. I. Chigirinsky, and Y. Ishida. Upper thermal limits on the oceanic distribution of Pacific salmon (Oncorhynchus spp.) in the spring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):489–503, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-050.

Appenzeller:1995:ELM

[1531] A. R. Appenzeller and W. C. Leggett. An evaluation of light-mediated vertical migration of fish based on hydroacoustic analysis of the diel vertical movements of rainbow smelt (Osmerus mordax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):504–511, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-051.

Ndayibagira:1995:ETD

[1532] Aristocle Ndayibagira, Marie-Josée Cloutier, Perry D. Anderson, and Philip A. Spear. Effects of 3,3’,4,4’-tetrachlorobiphenyl on the dynamics of vitamin A in brook trout (Salvelinus fontinalis) and intestinal retinoid concentrations in lake sturgeon (Acipenser fulvescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):512–520, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-052.

Silvert:1995:DMP

[1533] W. L. Silvert and A. D. Cembella. Dynamic modelling of phycotoxin kinetics in the blue mussel, Mytilus edulis, with implications for other marine invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (3):521–531, March 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-053.

Ogura:1995:HBV

[1534] Miki Ogura and Yukimasa Ishida. Homing behavior and vertical movements of four species of Pacific salmon (Oncorhynchus spp.) in the central Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):532–540, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-054.

Hop:1995:SEA

[1535] Haakon Hop, Vance L. Trudeau, and Mark Graham. Spawning energetics of Arctic cod (Boreogadus saida) in relation to seasonal development of the ovary and plasma sex steroid levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):541–550, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-055.

Blomqvist:1995:PWC

[1536] Peter Blomqvist, Russell T. Bell, Hans Olofsson, Ulrika Stensdotter, and Katarina Vrede. Plankton and water chemistry in Lake Njupfatet before and after liming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):551–565, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-056.

Jacobson:1995:SRM

[1537] Larry D. Jacobson and Alec D. MacCall. Stock-recruitment models for Pacific sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):566–577, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub. com/doi/full/10.1139/f95-057. See erratum [1711], comment [2915], and reply [2916].

Fitzsimons:1995:IZM

[1538] J. D. Fitzsimons, J. H. Leach, S. J. Nepszy, and V. W. Cairns. Impacts of zebra mussel on walleye (Stizostedion vitreum) reproduction in western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):578–586, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-058.

Vitt:1995:SVW

[1539] Dale H. Vitt, Suzanne E. Bayley, and Tai-Long Jin. Seasonal variation in water chemistry over a bog-rich fen gradient in Continental Western Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):587–606, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-059.

Cocchio:1995:EWC

[1540] L. A. Cocchio, F. W. H. Beamish, and D. W. Rodgers. Effects of water chemistry and temperature on radiocesium dynamics in rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (3):607–613, March 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-060.

Healey:1995:PAN

[1541] M. C. Healey and U. Reinhardt. Predator avoidance in naive and experienced juvenile Chinook and Coho salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):614–622, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-061.

Kreeger:1995:DOS

[1542] Karen Young Kreeger. Differences in the onset of salinity tolerance between juvenile Chinook salmon from two coastal Oregon river systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):623–630, March 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-062.

Randall:1995:FPF

[1543] R. G. Randall, C. K. Minns, and J. R. M. Kelso. Fish production in freshwaters: Are rivers more productive than lakes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):631–643, March 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-063.

Beamish:1995:PSP

[1544] Richard J. Beamish and Chrys-Ellen M. Neville. Pacific salmon and Pacific herring mortalities in the Fraser River plume caused by river lamprey (Lampetra ayresi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (3):644–650, March 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-064.

France:1995:CES

[1545] Robert France. Critical examination of stable isotope analysis as a means for tracing carbon pathways in stream ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(3):651–656, March 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-065. See com-

ment [1978].

Fletcher:1995:AOW

[1546] W. J. Fletcher. Application of the otolith weight–age relationship for the pilchard, Sardinops sagax neopilchardus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):657–664, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-066.

Koljonen:1995:DBR

[1547] Marja-Liisa Koljonen. Distinguishing between resident and migrating Atlantic salmon (Salmo salar) stocks by genetic stock composition analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):665–674, April 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-067.

Nathanailides:1995:IPT

[1548] C. Nathanailides, N. C. Stickland, and O. Lopez-Albors. Influence of prehatch temperature on the development of muscle cellularity in posthatch Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):675–680, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-068.

Heuch:1995:DVM

[1549] Peter Andreas Heuch, Aengus Parsons, and Karin Boxaspen. Diel vertical migration: a possible host-finding mechanism in salmon louse (Lepeophtheirus salmonis) copepodids? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):681–689, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f95-069.

Couillard:1995:FTFa

[1550] Y. Couillard, P. G. C. Campbell, A. Tessier, J. C. Auclair, and J. Pellerin-Massicotte. Field transplantation of a freshwater bivalve, Pyganodon grandis, across a metal contamination gradient. I. Temporal changes in metallothionein and metal (Cd, Cu, and Zn) concentrations in soft tissues. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):690–702, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-070.

Couillard:1995:FTFb

[1551] Y. Couillard, P. G. C. Campbell, J. G. Auclair, and J. Pellerin-Massicotte. Field transplantation of a freshwater bivalve, Pyganodon grandis, across a metal contamination gradient. II. Metallothionein response to Cd and Zn exposure, evidence for cytotoxicity, and links to effects at higher levels of biological organization. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):703–715, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-071.

Crone:1995:SDS

[1552] Paul R. Crone. Sampling design and statistical considerations for the commercial groundfish fishery of Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):716–732, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-072.

Hessen:1995:RHZ

[1553] Dag O. Hessen, Tom Andersen, and Bjørn A. Faafeng. Replacement of herbivore zooplankton species along gradients of ecosystem productivity and fish predation pressure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (4):733–742, April 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-073.

Stokesbury:1995:BPV

[1554] Kevin D. E. Stokesbury and John H. Himmelman. Biological and physical variables associated with aggregations of the giant scallop Placopecten magellanicus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):743–753, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-

LaBaugh:1995:GCW

[1555] James W. LaBaugh, Donald O. Rosenberry, and Thomas C. Winter. Groundwater contribution to the water and chemical budgets of Williams Lake, Minnesota, 1980–1991. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (4):754–767, April 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-075.

Barnhisel:1995:SSF

[1556] D. Rae Barnhisel and Heather A. Harvey. Size-specific fish avoidance of the spined crustacean Bythotrephes: field support for laboratory predictions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):768–775, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-076.

Tsai:1995:ESI

[1557] H. J. Tsai, F. S. Tseng, and I. C. Liao. Electroporation of sperm to introduce foreign DNA into the genome of loach (Misgurnus anguillicaudatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):776–787, April 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-077.

Prairie:1995:NVE

[1558] Yves T. Prairie, David F. Bird, and Robert H. Peters. Natural variability and the estimation of empirical relationships: a reassessment of regression methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (4):788–798, April 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-078.

Prairie:1995:UST

[1559] Yves T. Prairie and C. Tara Marshall. On the use of structured timeseries to detect and test hypotheses about within-lakes relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):799–803, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-079.

Basu:1995:LSD

[1560] B. K. Basu and F. R. Pick. Longitudinal and seasonal development of planktonio chlorophyll a in the Rideau River, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):804–815, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-080.

DeWalle:1995:EFD

[1561] David R. DeWalle, Bryan R. Swistock, and William E. Sharpe. Episodic flow–duration analysis: a method of assessing toxic exposure of brook trout (Salvelinus fontinalis) to episodic increases in aluminum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):816–827, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-081.

Cattaneo:1995:TVC

[1562] Antonella Cattaneo and Yves T. Prairie. Temporal variability in the chemical characteristics along the Rivière de l’Achigan: How many samples are necessary to describe stream chemistry? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):828–835, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-082.

Marsden:1995:GSA

[1563] J. Ellen Marsden, Adrian Spidle, and Bernie May. Genetic similarity among zebra mussel populations within North America and Europe. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):836–847, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-083.

McLaughlin:1995:CNA

[1564] Robert L. McLaughlin, Moira M. Ferguson, and David L. G. Noakes. Concentrations of nucleic acids and protein as indices of nutritional status for recently emerged brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):848–854, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-084.

Baker:1995:EIT

[1565] Peter Fritz Baker, Franklin K. Ligon, and Terence P. Speed. Estimating the influence of temperature on the survival of Chinook salmon smolts (Oncorhynchus tshawytscha) migrating through the Sacramento– San Joaquin River Delta of California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):855–863, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-085.

Borgmann:1995:KEA

[1566] U. Borgmann and W. P. Norwood. Kinetics of excess (above background) copper and zinc in Hyalella azteca and their relationship to chronic toxicity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):864–874, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-086.

Borgmann:1995:ETB

[1567] U. Borgmann and W. P. Norwood. EDTA toxicity and background concentrations of copper and zinc in Hyalella azteca. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(4):875–881, April 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-087.

Millar:1995:FFH

[1568] R. B. Millar. The functional form of hook and gillnet selection curves cannot be determined from comparative catch data alone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):883–891, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-088.

Araki:1995:ADR

[1569] Kazuo Araki, Hisako Shinma, Hiroyuki Nagoya, Ichiro Nakayama, and Hiroshi Onozato. Androgenetic diploids of rainbow trout (Oncorhynchus mykiss) produced by fused sperm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):892–896, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f95-089.

Robarts:1995:PPP

[1570] Richard D. Robarts, Michael T. Arts, and David B. Donald. Phytoplankton primary production of three temporary northern prairie wetlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):897–902, May 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-090.

Sainte-Marie:1995:GMB

[1571] Bernard Sainte-Marie, Sylvain Raymond, and Jean-Claude Brêthes. Growth and maturation of the benthic stages of male snow crab, Chionoecetes opilio (Brachyura: Majidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):903–924, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-091.

Mills:1995:PAA

[1572] Edward L. Mills, Connie Adams, Robert O’Gorman, Randall W. Owens, and Edward F. Roseman. Planktivory by alewife (Alosa pseudoharengus) and rainbow smelt (Osmerus mordax) on microcrustacean zooplankton and dreissenid (Bivalvia: Dreissenidae) veligers in southern Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):925–935, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-092.

Angermeier:1995:ENS

[1573] Paul L. Angermeier and Roy A. Smogor. Estimating number of species and relative abundances in stream-fish communities: effects of sampling effort and discontinuous spatial distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):936–949, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-093.

Walker:1995:CDQ

[1574] Ian R. Walker, Susan E. Wilson, and John P. Smol. Chironomidae (Diptera): quantitative palaeosalinity indicators for lakes of western Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):950–960, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-094.

Wirgin:1995:MSA

[1575] Isaac Wirgin, Mark Pedersen, Steven Maceda, Brian Jessop, Simon Courtenay, and John R. Waldman. Mixed-stock analysis of striped bass in two rivers of the Bay of Fundy as revealed by mitochondrial DMA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):961–970, May 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-095.

Miltner:1995:IFP

[1576] Robert J. Miltner, Steve W. Ross, and Martin H. Posey. Influence of food and predation on the depth distribution of juvenile spot (Leiostomus xanthurus) in tidal nurseries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (5):971–982, May 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-096.

Brown:1995:SEC

[1577] Richard S. Brown and William C. Mackay. Spawning ecology of cutthroat trout (Oncorhynchus clarki) in the Ram River, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):983–992, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-097.

Chen:1995:RRA

[1578] Y. Chen and J. E. Paloheimo. A robust regression analysis of recruitment in fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):993–1006, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
098.

Marshall:1995:DDH

[1579] C. Tara Marshall and Kenneth T. Frank. Density-dependent habitat selection by juvenile haddock (Melanogrammus aeglefinus) on the southwestern Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (5):1007–1017, May 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-099.

Hannah:1995:VGS

[1580] Robert W. Hannah. Variation in geographic stock area, catchability, and natural mortality of ocean shrimp (Pandalus jordani): some new evidence for a trophic interaction with Pacific hake (Merluccius productus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1018–1029, May 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-100.

Kurokawa:1995:DDO

[1581] Tadahide Kurokawa, Hirohiko Kagawa, Hiromi Ohta, Hideki Tanaka, Koichi Okuzawa, and Keiji Hirose. Development of digestive organs and feeding ability in larvae of Japanese eel (Anguilla japonica). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1030–1036, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-101.

Evans:1995:PVA

[1582] Marlene S. Evans, Richard D. Robarts, and Michael T. Arts. Predicted versus actual determinations of algal production, algal biomass, and zooplankton biomass in a hypereutrophic, hyposaline prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1037–1049, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-102.

McKinnell:1995:ASE

[1583] Skip McKinnell. Age-specific effects of sockeye abundance on adult body size of selected British Columbia sockeye stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1050–1063, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-103.

Westgate:1995:DBH

[1584] Andrew J. Westgate, Andrew J. Head, Per Berggren, Heather N. Koopman, and David E. Gaskin. Diving behaviour of harbour porpoises, Phocoena phocoena. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (5):1064–1073, May 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-104.

Rudstam:1995:PCB

[1585] Lars G. Rudstam, Paul E. Peppard, Thomas W. Fratt, Richard E. Bruesewitz, Daniel W. Coble, Fred A. Copes, and James F. Kitchell. Prey consumption by the burbot (Lota lota) population in Green Bay, Lake Michigan, based on a bioenergetics model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1074–1082, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-105.

Miller:1995:IBA

[1586] Thomas J. Miller, Tomasz Herra, and William C. Leggett. An individualbased analysis of the variability of eggs and their newly hatched larvae of Atlantic cod (Gadus morhua) on the Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1083–1093, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-106.

Millie:1995:TTS

[1587] David F. Millie, Bryan T. Vinyard, Michael C. Baker, and Craig S. Tucker. Testing the temporal and spatial validity of site-specific models derived from airborne remote sensing of phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1094–1107, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-107.

Leland:1995:DPY

[1588] Harry V. Leland. Distribution of phytobenthos in the Yakima River Basin, Washington, in relation to geology, land use and other environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1108– 1129, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-108.

Payne:1995:AVP

[1589] Barry S. Payne, Andrew C. Miller, Erica D. Hubertz, and Jin Lei. Adaptive variation in palp and gill size of the zebra mussel (Dreissena polymorpha) and Asian clam (Corbicula fluminea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(5):1130–1134, May 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-109.

Castell:1995:PIG

[1590] John D. Castell, Linda D. Boston, Robert J. Miller, and Trevor Kenchington. The potential identification of the geographic origin of lobster eggs from various wild stocks based on fatty acid composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1135–1140, June 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-110.

Schroder:1995:MSF

[1591] Steven L. Schroder, Curtis M. Knudsen, and Eric C. Volk. Marking salmon fry with strontium chloride solutions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1141–1149, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-111.

Mallatt:1995:QUG

[1592] Jon Mallatt, Steven J. Lampa, J. Franklin Bailey, Marc A. Evans, and Wayne Tate. Quantitative ultrastructure of gill epithelial cells in the larval lamprey Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1150–1164, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-112.

Mallatt:1995:FGS

[1593] Jon Mallatt, Steven J. Lampa, J. Franklin Bailey, Marc A. Evans, and Steve Brumbaugh. A fish gill system for quantifying the ultrastructural effects of environmental stressors: methylmercury, Kepone(R), and heat shock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1165–1182, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-113.

Murray:1995:RQP

[1594] Thomas E. Murray and Peter H. Rich. Respiratory quotients predict increases in reducing products in the hypolimnion of a stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1183–1189, June 1995.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-114.

Murray:1995:CBI

[1595] Thomas E. Murray. The correlation between iron sulfide precipitation and hypolimnetic phosphorus accumulation during one summer in a softwater lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1190–1194, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
115.

Hobson:1995:CTS

[1596] Keith A. Hobson and Harold E. Welch. Cannibalism and trophic structure in a high Arctic lake: insights from stable-isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1195–1201, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-116.

Holland:1995:TNC

[1597] Ruth E. Holland, Thomas H. Johengen, and Alfred M. Beeton. Trends in nutrient concentrations in Hatchery Bay, western Lake Erie, before and after Dreissena polymorpha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (6):1202–1209, June 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-117.

Rosel:1995:VMC

[1598] P. E. Rosel, A. E. Dizon, and M. G. Haygood. Variability of the mitochondrial control region in populations of the harbour porpoise, Phocoena, on interoceanic and regional scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1210–1219, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-118.

Watras:1995:MRR

[1599] C. J. Watras, K. A. Morrison, and N. S. Bloom. Mercury in remote Rocky Mountain lakes of Glacier National Park, Montana, in comparison with other temperate North American regions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1220–1228, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-119.

Zheng:1995:LBP

[1600] J. Zheng, M. C. Murphy, and G. H. Kruse. A length-based population model and stock–recruitment relationships for red king crab, Paralithodes camtschaticus, in Bristol Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1229–1246, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-120.

Brett:1995:IMS

[1601] Michael T. Brett, Charles R. Goldman, Frederick S. Lubnow, Anke Müller-Solger, Astrid Bracher, Darren Brandt, and Odette Brandt. Impact of a major soil fumigant spill on the planktonic ecosystem of Shasta Lake, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1247– 1256, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-121.

Bergstedt:1995:SGD

[1602] Roger A. Bergstedt and William D. Swink. Seasonal growth and duration of the parasitic life stage of the landlocked sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1257–1264, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-122.

Myers:1995:SAC

[1603] Ransom A. Myers and Noel G. Cadigan. Statistical analysis of catchat-age data with correlated errors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1265–1273, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-123.

Myers:1995:WIN

[1604] Ransom A. Myers and Noel G. Cadigan. Was an increase in natural mortality responsible for the collapse off northern cod? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1274–1285, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-124.

Chamberland:1995:APP

[1605] G. Chamberland, D. Bélanger, N. Larivière, L. Vermette, J. F. Klaverkamp, and J. S. Blais. Abnormal porphyrin profile in mussels exposed to low concentrations of cadmium in an experimental Precambrian Shield lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1286–1293, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
125.

Lochmann:1995:LCC

[1606] Steve E. Lochmann, Gary L. Maillet, Kenneth T. Frank, and Christopher T. Taggart. Lipid class composition as a measure of nutritional condition in individual larval Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1294–1306, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-126.

Baumgartner:1995:PGE

[1607] Jeffrey V. Baumgartner. Phenotypic, genetic, and environmental integration of morphology in a stream population of the threespine stickleback, Gasterosteus aculeatus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (6):1307–1317, June 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-127.

Songster-Alpin:1995:CET

[1608] M. S. Songster-Alpin and R. L. Klotz. A comparison of electron transport system activity in stream and beaver pond sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1318–1326, June 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-128.

Bradford:1995:CRP

[1609] Michael J. Bradford. Comparative review of Pacific salmon survival rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(6):1327–1338, June 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-129.

vandenHeuvel:1995:HED

[1610] M. R. van den Heuvel, D. G. Dixon, K. R. Munkittrick, M. R. Servos, and G. J. Van Der Kraak. Hepatic 7-ethoxyresorufin-O-deethylase activity, plasma steroid hormone concentrations, and liver bioassayderived 2,3,7,8-TCDD toxic equivalent concentrations in wild white sucker (Catostomus commersoni) caged in bleached kraft pulp mill effluent. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1339–1350, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-130.

DeLeo:1995:SAS

[1611] G. A. De Leo and M. Gatto. A size and age-structured model of the European eel (Anguilla anguilla L.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1351–1367, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-131.

Wang:1995:AIV

[1612] Yoy-Gan Wang and Mervyn R. Thomas. Accounting for individual variability in the von Bertalanffy growth model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1368–1375, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-132.

Devlin:1995:PGT

[1613] Robert H. Devlin, Timothy Y. Yesaki, Edward M. Donaldson, Shao Jun Du, and Choy-Leong Hew. Production of germline transgenic Pacific salmonids with dramatically increased growth performance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1376–1384, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-133.

Gudjonsson:1995:RGS

[1614] S. Gudjonsson, S. M. Einarsson, Th. Antonsson, and G. Gudbergsson. Relation of grilse to salmon ratio to environmental changes in several wild stocks of Atlantic salmon (Salmo salar) in Iceland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1385–1398, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-134.

Kalish:1995:ABR

[1615] John M. Kalish. Application of the bomb radiocarbon chronometer to the validation of redfish Centroberyx affinis age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1399–1405, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-135.

Reid:1995:STV

[1616] R. A. Reid, K. M. Somers, and S. M. David. Spatial and temporal variation in littoral-zone benthic invertebrates from three south-central Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1406–1420, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
136.

Fowler:1995:EAEa

[1617] Anthony J. Fowler, Steven E. Campana, Simon R. Thorrold, and Cynthia M. Jones. Experimental assessment of the effect of temperature and salinity on elemental composition of otoliths using solution-based ICPMS. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1421–1430, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-137.

Fowler:1995:EAEb

[1618] Anthony J. Fowler, Steven E. Campana, Simon R. Thorrold, and Cynthia M. Jones. Experimental assessment of the effect of temperature and salinity on elemental composition of otoliths using laser ablation ICPMS. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1431–1441, July 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-138.

Emlen:1995:PVS

[1619] John M. Emlen. Population viability of the Snake River Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (7):1442–1448, July 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-139.

Ricciardi:1995:PII

[1620] A. Ricciardi, J. B. Rasmussen, and F. G. Whoriskey. Predicting the intensity and impact of Dreissena infestation on native unionid bivalves from Dreissena field density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (7):1449–1461, July 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-140.

Mandrak:1995:BPF

[1621] Nicholas E. Mandrak. Biogeographic patterns of fish species richness in Ontario lakes in relation to historical and environmental factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1462–1474, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-141.

Pepin:1995:VCT

[1622] P. Pepin, J. A. Helbig, R. Laprise, E. Colbourne, and T. H. Shears. Variations in the contribution of transport to changes in planktonic animal abundance: a study of the flux of fish larvae in Conception Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1475–1486, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
142.

Wood:1995:MMU

[1623] Brian M. Wood and Mark B. Bain. Morphology and microhabitat use in stream fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1487–1498, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
143.

Minns:1995:AHR

[1624] Charles K. Minns. Allometry of home range size in lake and river fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1499–1508, July 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-144.

Johannsson:1995:RMR

[1625] Ora E. Johannsson. Response of Mysis relicta population dynamics and productivity to spatial and seasonal gradients in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1509–1522, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-145.

Walters:1995:EHS

[1626] Carl Walters. Estimation of historical stock sizes and recruitment anomalies from relative abundance time series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1523–1534, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-146.

Bell:1995:SIC

[1627] Kim N. I. Bell, Joseph A. Brown, and Pierre Pepin. Seasonal, inverse cycling of length- and age-at-recruitment in the diadromous gobies Sicydium punctatum and Sicydium antillarum in Dominica, West Indies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1535–1545, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-147.

Rand:1995:ELW

[1628] Peter S. Rand, Donald J. Stewart, Brian F. Lantry, Lars G. Rudstam, Ora E. Johannsson, Andrew P. Goyke, Stephen B. Brandt, Robert O’Gorman, and Gary W. Eck. Effect of lake-wide planktivory by the pelagic prey fish community in Lakes Michigan and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1546–1563, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-148.

Eby:1995:PRP

[1629] Lisa A. Eby, Lars G. Rudstam, and James F. Kitchell. Predator responses to prey population dynamics: an empirical analysis based on lake trout growth rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (7):1564–1571, July 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-149.

Mason:1995:SEB

[1630] Doran M. Mason, Andrew Goyke, and Stephen B. Brandt. A spatially explicit bioenergetics measure of habitat quality for adult salmonines: Comparison between Lakes Michigan and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1572–1583, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-750.

Olver:1995:TDC

[1631] C. H. Olver, B. J. Shuter, and C. K. Minns. Toward a definition of conservation principles for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1584–1594, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-751. See [2360] and re-

ply [2361].

Bernatchez:1995:EPR

[1632] Louis Bernatchez and Julian J. Dodson. Erratum: Phylogenic relationships among Palearctic and Nearctic whitefish (Coregonus sp.) populations as revealed by mtDNA variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(7):1595, July 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-752. See [1472].

Matilainen:1995:MMD

[1633] Tuula Matilainen and Matti Verta. Mercury methylation and demethylation in aerobic surface waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (8):1597–1608, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-753.

Watanabe:1995:PDJ

[1634] Yoshiro Watanabe, Hiromu Zenitani, and Ryo Kimura. Population decline off the Japanese sardine Sardinops melanostictus owing to recruitment failures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1609– 1616, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f95-154.

McVeigh:1995:MDD

[1635] H. P. McVeigh, R. A. Hynes, and A. Ferguson. Mitochondrial DNA differentiation of sympatric populations of brown trout, Salmo trutta L., from Lough Melvin, Ireland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (8):1617–1622, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-755.

Truscott:1995:EAL

[1636] R. Truscott, C. R. McCrohan, S. E. R. Bailey, and K. N. White. Effect of aluminium and lead on activity in the freshwater pond snail Lymnaea stagnalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1623–1629, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-756.

Bryan:1995:ECF

[1637] Miehael D. Bryan, Gary J. Atchison, and Mark B. Sandheinrich. Effects of cadmium on the foraging behavior and growth of juvenile bluegill, Lepomis macrochirus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8): 1630–1638, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-757.

Brown:1995:EPS

[1638] Larry R. Brown and Anne M. Brasher. Effect of predation by Sacramento squawfish (Ptychocheilus grandis) on habitat choice of California roach (Lavinia symmetricus) and rainbow trout (Oncorhynchus mykiss) in artificial streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8): 1639–1646, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-758.

Hartman:1995:CED

[1639] Kyle J. Hartman and Stephen B. Brandt. Comparative energetics and the development of bioenergetics models for sympatric estuarine piscivores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1647–1666, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-759.

Harfmart:1995:PDI

[1640] Kyle J. Harfmart and Stephen B. Brandt. Predatory demand and impact of striped bass, bluefish, and weakfish in the Chesapeake Bay: applications of bioenergetics models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (8):1667–1687, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-760.

Epifanio:1995:MDD

[1641] John M. Epifanio, Bonnie L. Brown, Peter E. Smouse, and Carol J. Kobak. Mitochondrial DNA divergence among popylations of American shad (Alosa sapidissima): how much variation is enough for mixed-stock analysis? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1688–1702, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-761.

Branchaud:1995:VST

[1642] Alain Branchaud, Andrée Gendron, Réjean Fortin, Philip A. Spear, and Perry D. Anderson. Vitamin A stores, teratogenesis, and EROD activity in white sucker, Catostomus commersoni, from Rivière des prairies near Montréal and a reference site. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (8):1703–1713, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-762.

Landry:1995:ESS

[1643] François Landry, Thomas J. Miller, and William C. Leggett. The effects of small-scale turbulence on the ingestion rate of fathead minnow (Pimephales promelas) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1714–1719, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-763.

Bremer:1995:MDC

[1644] Jaime R. Alvarado Bremer, Allan J. Baker, and Jaime Mejuto. Mitochondrial DNA control region sequences indicate extensive mixing of swordfish (Xiphias gladius) populations in the Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1720–1732, August 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-764.

Curry:1995:GSS

[1645] R. Allen Curry and David L. G. Noakes. Groundwater and the selection of spawning sites by brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1733–1740, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-765.

Curry:1995:GIE

[1646] R. Allen Curry, David L. G. Noakes, and George E. Morgan. Groundwater and the incubation and emergence of brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1741–1749, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
766.

Devito:1995:SMB

[1647] Kevin J. Devito. Sulphate mass balances of Precambrian Shield wetlands; the influence of catchment hydrogeology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1750–1760, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-767.

Aldridge:1995:OCN

[1648] David W. Aldridge, Barry S. Payne, and Andrew C. Miller. Oxygen consumption, nitrogenous excretion, and filtration rates of Dreissena polymorpha at acclimation temperatures between 20 and 32 ^{o}C. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1761–1767, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-768.

Parent:1995:MDF

[1649] Serge Parent and Lynn M. Schriml. A model for the determination of fish species at risk based upon life-history traits and ecological data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1768–1781, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-769.

Auclair:1995:IIU

[1650] J. C. Auclair. Implications of increased UV-B induced photoreduction: iron (II) enrichment stimulates picocyanobacterial growth and the microbial food web in clear-water acidic Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1782–1788, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-770.

Richmond:1995:CFL

[1651] Ann D. Richmond and Kurt D. Fauseh. Characteristics and function of large woody debris in subalpine Rocky Mountain streams in northern Colorado. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1789–1802, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-771.

Havens:1995:PLA

[1652] Karl E. Havens. Particulate light attenuation in a large subtropical lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1803–1811, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-172.

France:1995:OSP

[1653] Robert L. France, Robert H. Peters, and Robert H. Rishikof. Optimal sampling period for predicting phosphores and chlorophyll concentrations in north-temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (8):1812–1815, August 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-173.

Cornwell:1995:BOI

[1654] Jeffrey C. Cornwell, J. C. Stevenson, and C. Neill. Biogeochemieal origin of δ ^{34}S isotopic signatures in a prairie marsh. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(8):1816–1820, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-174.

Bower:1995:DAI

[1655] S. M. Bower. Differences in acquired and innate resistance among Pacific salmon (Oncorhynchus spp.) against the haemoflagellate Cryptobia salmositica and numbers delivered by the leech vector Piscicola salmositica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):1–6, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Cone:1995:REM

[1656] David K. Cone and Richard L. Raesly. Redescription of Myxobolus rhinichthidis (Myxosporea) parasitizing Rhinichthys cataractae, with a revised taxonomic list of species of Myxobolus known from North American freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1): 7–12, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Kent:1995:PKD

[1657] M. L. Kent, M. Higgins, D. J. Whitaker, and H. Yokoyama. Proliferative kidney disease and Sphaerospora oncorhynchi in wild-caught salmonids from the Puntledge River system, Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):13–17, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Urawa:1995:ERC

[1658] Shigehiko Urawa. Effects of rearing conditions on growth and mortality of juvenile chum salmon (Oncorhynchus keta) infected with Ichthyobodo necator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):18–23, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Beverley-Burton:1995:OMS

[1659] Mary Beverley-Burton. Origins of the Monogenea of selected major taxa of Nearctic freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (S1):24–34, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Euzet:1995:ODD

[1660] Louis Euzet, Guy Oliver, and Mohamed Hedi Ktari. Organisation et développement du système osmorégulateur chez les Monogenea. (French) [Organization and development of the osmoregulatory system in Monogenea]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):35–51, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Gusev:1995:SPF

[1661] A. V. Gusev. Some pathways and factors of monogenean microevolution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):52–56, August 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Malashenko:1995:PPA

[1662] A. A. Malashenko and V. A. Roytman. A population-phenetic analysis of morphological polymorphism in Polystoma integerrimum (Monogenea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):57–61, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Ogawa:1995:REB

[1663] Kazuo Ogawa, Melba G. Bondad-Reantaso, and Hisatsugu Wakabayashi. Redescription of Benedenia epinepheli (Yamaguti, 1937) Meserve, 1938 (Monogenea: Capsalidae) from cultured and aquarium marine fishes of Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):62–70, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Bray:1995:ESO

[1664] Rodney A. Bray. Steringophorus Odhner, 1905 (Digenea: Fellodistomidae) in deep-sea fishes from the northeastern Atlantic, with the description of Steringophorus margolisi n.sp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):71–77, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Ching:1995:ECD

[1665] Hilda Lei Ching. Evaluation of characters of the digenean family Gymnophallidae Morozov, 1955. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):78–83, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Dailey:1995:RPT

[1666] Murray D. Dailey and Robert Morris. Relationship of parasites (Trematoda: Spirorchidae) and their eggs to the occurrence of fibropapillomas in the green turtle (Chelonia mydas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):84–89, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Gibson:1995:EAM

[1667] David I. Gibson. Allopodocotyle margolisi n.sp. (Digenea: Opecoelidae) from the deep-sea fish Coryphaenoides (Chalinura) mediterraneus in the northeastern Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1): 90–94, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Kamegai:1995:NDE

[1668] Shunya Kamegai and Jun Araki. A new digenean, Didymocystis margolisi n.sp. (Didymozoidae: Didymozoinae), from the skipjack tuna, Katsuwonus pelamis, of Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (S1):95–97, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Overstreet:1995:NSM

[1669] Robin M. Overstreet and Richard W. Heard. A new species of Megalophallus (Digenea: Microphallidae) from the clapper rail, other birds, and the littoral isopod Ligia baudiniana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):98–104, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Boily:1995:GVA

[1670] France Boily and David J. Marcogliese. Geographical variations in abundance of larval anisakine nematodes in Atlantic cod (Gadus morhua) and American plaice (Hippoglossoides platessoides) from the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):105–115, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Brattey:1995:ILE

[1671] John Brattey. Identification of larval Contracaecum osculatum s.l. and Phocascaris sp. (Nematoda: Ascaridoidea) from marine fishes by allozyme electrophoresis and discriminant function analysis of morphometric data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):116–128, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). See erratum [1827].

Despres:1995:DDM

[1672] L. Després, M. L. Adamson, and T. E. McDonald. Development of a diagnostic molecular marker for Philonema spp. (Nematoda; Dracunculoidea) infecting salmonids in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):129–133, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Koie:1995:DTS

[1673] Marianne Køie, Bjørn Berland, and Michäel D. B. Burt. Development to third-stage larvae occurs in the eggs of Anisakis simplex and Pseudotetranova decipiens (Nematoda, Ascaridoidea, Anisakidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):134–139, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

McClelland:1995:EIF

[1674] G. McClelland. Experimental infection of fish with larval sealworm, Pseudotertanova decipiens (Nematoda, Anisakinae), transmitted by amphipods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):140–155, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Measures:1995:NME

[1675] Lena N. Measures and Henry Hong. The number of moults in the egg of sealworm, Pseudoterranova decipiens (Nematoda: Ascaridoidea): an ultrastructural study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):156– 160, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Moravec:1995:EPM

[1676] František Moravec, Victor M. Vidal-Martínez, and Leopoldina Aguirre-Macedo. Philometra margolisi n.sp. (Nematoda: Philometridae) from the gonads of the red grouper, Epinephelus morio (Pisces: Serranidae), in Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):161–165, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Johnson:1995:TNC

[1677] S. C. Johnson and Z. Kabata. Two new copepod species of the genus Pseudotaeniacanthus Yamaguti and Yamasu, 1959 (Poecilostomatoida: Taeniacanthidae), parasitic on moray eels in the Red Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):166–174, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

MacKinnon:1995:PPC

[1678] B. M. MacKinnon. The poor potential of cunner, Tautogolabrus adspersus, to act as cleaner fish in removing sea lice (Caligus elongatus) from farmed salmon in eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):175–177, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Nagasawa:1995:RJH

[1679] Kazuya Nagasawa, Masamichi Yamamoto, Yasunori Sakurai, and Akira Kumagai. Rediscovery in Japan and host association of Salmincola carpionis (Copepoda: Lernaeopodidae), a parasite of wild and reared freshwater salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):178– 185, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Curtis:1995:PES

[1680] Mark A. Curtis, Michel Bérubé, and Andreas Stenzel. Parasitological evidence for specialized foraging behavior in lake-resident Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):186– 194, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Khan:1995:PBI

[1681] R. A. Khan and C. Tuck. Parasites as biological indicators of stocks of Atlantic cod (Gadus morhua) off Newfoundland, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):195–201, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Konovalov:1995:PIB

[1682] S. M. Konovalov. Parasites as indicators of biological processes, with special reference to sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):202–212, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

MacKenzie:1995:PHE

[1683] K. MacKenzie and M. Longshaw. Parasites of the hakes Merluccius australis and M. hubbsi in the waters around the Falkland Islands, southern Chile, and Argentina, with an assessment of their potential value as biological tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):213– 224, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Sewell:1995:SCM

[1684] Kim B. Sewell and Robert J. G. Lester. Stock composition and movement of gemfish, Rexea solandri, as indicated by parasites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):225–232, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Valtonen:1995:ITP

[1685] E. Tellervo Valtonen and Markku Julkunen. Influence of the transmission of parasites from prey fishes on the composition of the parasite community of a predatory fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (S1):233–245, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

Arthur:1995:PCE

[1686] J. R. Arthur, E. Albert, and F. Boily. Parasites of capelin (Mallotus villosus) in the St. Lawrence estuary and gulf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(S1):246–253, August 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

McIntosh:1995:CPR

[1687] Angus R. McIntosh and Colin R. Townsend. Contrasting predation risks presented by introduced brown trout and native common river galaxias in New Zealand streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9): 1821–1833, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-175.

Tupper:1995:EHS

[1688] M. Tupper and R. G. Boutilier. Effects of habitat on settlement, growth, and postsettlement survival of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1834–1841, September 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-176.

Sempeski:1995:SRC

[1689] P. Sempeski and P. Gaudin. Size-related changes in diel distribution of young grayling (Thymallus thymallus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1842–1848, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-177.

Myers:1995:SSV

[1690] R. A. Myers, G. Mertz, and N. J. Barrowman. Spatial scales of variability in cod recruitment in the North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1849–1862, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-778.

McConnell:1995:PML

[1691] Stewart K. McConnell, Patrick O’Reilly, Lorraine Hamilton, Jonathan M. Wright, and Paul Bentzen. Polymorphic microsatellite loci from Atlantic salmon (Salmo salar): genetic differentiation of North American and European populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1863– 1872, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-779.

Mallet:1995:CGS

[1692] André L. Mallet and Claire E. Carver. Comparative growth and survival patterns of Mytilus trossulus and Mytilus edulis in Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1873–1880, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-780.

Morin:1995:TEV

[1693] Antoine Morin, Marco A. Rodríguez, and Daniel Nadon. Temporal and environmental variation in the biomass spectrum of benthic invertebrates in streams: an application of thin-plate splines and relative warp analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1881–1892, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-781.

DAngelo:1995:EUG

[1694] Donna J. D’Angelo, Judy L. Meyer, Leslie M. Howard, Stanley V. Gregory, and Linda R. Ashkenas. Ecological uses for genetic algorithms: predicting fish distributions in complex physical habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1893–1908, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-782.

Hare:1995:EAG

[1695] Jonathan A. Hare and Robert K. Cowen. Effect of age, growth rate, and ontogeny on the otolith size–fish size relationship in bluefish, Pomatomus saltatrix, and the implications for back-calculation of size in fish early life history stages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1909– 1922, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-783.

Moore:1995:MBW

[1696] A. Moore, E. C. E. Potter, N. J. Milner, and S. Bamber. The migratory behaviour of wild Atlantic salmon (Salmo salar) smolts in the estuary of the River Conwy, North Wales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1923–1935, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-784.

Couture:1995:RLL

[1697] Suzanne Couture. Response of the Laflamme Lake watersheds Quebec, to reduced sulphate deposition (1981–1992). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1936–1943, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-785.

Overholtz:1995:ECN

[1698] William J. Overholtz, Steven F. Edwards, and Jon K. T. Brodziak. Effort control in the New England groundfish fishery: a bioeconomic perspective. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1944–1957, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-786.

Sorensen:1995:RIB

[1699] Peter W. Sorensen, Tim Essington, Dana E. Weigel, and James R. Cardwell. Reproductive interactions between sympatric brook and brown trout in a small Minnesota stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1958–1965, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-787.

Clements:1995:IEB

[1700] William H. Clements and Peter M. Kiffney. The influence of elevation on benthic community responses to heavy metals in Rocky Mountain streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1966–1977, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-788.

Anderson:1995:AOD

[1701] John T. Anderson and Brad de Young. Application of a one-dimensional model to vertical distributions of cod eggs on the northeastern Newfoundland shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1978– 1989, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-789.

Phillips:1995:IAR

[1702] Glenn Phillips and Joshua Lipton. Injury to aquatic resources caused by metals in Montana’s Clark Fork River Basin: historic perspective and overview. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1990– 1993, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-190.

Woodward:1995:MCB

[1703] Daniel F. Woodward, Aïda M. Farag, Harold L. Bergman, Aaron J. De-Lonay, Edward E. Little, Charlie E. Smiths, and Frederic T. Barrows. Metals-contaminated benthic invertebrates in the Clark Fork River, Montana: effects on age-0 brown trout and rainbow trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):1994–2004, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-791.

Marr:1995:RSB

[1704] J. C. A. Marr, H. L. Bergman, M. Parker, J. Lipton, D. Cacela, W. Erickson, and G. R. Phillips. Relative sensitivity of brown and rainbow trout to pulsed exposures of an acutely lethal mixture of metals typical of the Clark Fork River, Montana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2005–2015, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-792.

Marr:1995:DRS

[1705] J. C. A. Marr, H. L. Bergman, J. Lipton, and C. Hogstrand. Differences in relative sensitivity of naive and metals-acclimated brown and rainbow trout exposed to metals representative of the Clark Fork River, Montana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2016–2030, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-793.

Woodward:1995:BTA

[1706] Daniel F. Woodward, James A. Hansen, Harold L. Bergman, Aaron J. DeLonay, and Edward E. Little. Brown trout avoidance of metals in water characteristic of the Clark Fork River, Montana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2031–2037, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-794.

Farag:1995:PIF

[1707] Aïda M. Farag, Mark A. Stansbury, Harold L. Bergman, Christer Hogstrand, and Elizabeth MacConnell. The physiological impairment of free-ranging brown trout exposed to metals in the Clark Fork River, Montana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2038–2050, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-795.

Stephenson:1995:FMS

[1708] Robert L. Stephenson and Daniel E. Lane. Fisheries management sciences: a plea for conceptual change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2051–2056, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-796. See comments [1709, 1710].

Caddy:1995:CFM

[1709] J. F. Caddy. Comment — fisheries management science: a plea for conceptual change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2057– 2058, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f95-797. See [1708, 1710].

Gauldie:1995:CFM

[1710] R. W. Gauldie. Comment — fisheries management science: a plea for conceptual change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2059– 2061, September 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f95-798. See [1708, 1709].

Jacobson:1995:ESR

[1711] Larry D. Jacobson and A. D. MacCall. Erratum: Stock-recruitment models for Pacific sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(9):2062, September 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-799. See [1537].

Schnute:1995:IEP

[1712] Jon T. Schnute and Laura J. Richards. The influence of error on population estimates from catch-age models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2063–2077, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-800.

Ankley:1995:IPH

[1713] G. T. Ankley, M. K. Schubauer-Berigan, and P. D. Monson. Influence of pH and hardness on toxicity of ammonia to the amphipod Hyalella azteca. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2078–2083, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
801.

Gibert:1995:HES

[1714] Janine Gibert, Sandrine Plénet, Pierre Marmonier, and Vladimir Vanek. Hydrological exchange and sediment characteristics in a riverbank: relationship between heavy metals and invertebrate community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2084–2097, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-202.

Hannah:1995:FOS

[1715] Robert W. Hannah, Stephen A. Jones, and Michele R. Long. Fecundity of the ocean shrimp (Pandalus jordani). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2098–2107, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-803.

Spidle:1995:LTT

[1716] Adrian P. Spidle, Bernie May, and Edward L. Mills. Limits to tolerance of temperature and salinity in the quagga mussel (Dreissena bugensis) and the zebra mussel (Dreissena polymorpha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2108–2119, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-804.

Wallace:1995:ILA

[1717] J. Bruce Wallace, Jackson R. Webster, and Judith L. Meyer. Influence of log additions on physical and biotic characteristics of a mountain stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2120–2137, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-805.

Tang:1995:RBR

[1718] Ming Tang and Daniel Boisclair. Relationship between respiration rate of juvenile brook trout (Salvelinus fontinalis), water temperature, and swimming characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (10):2138–2145, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f95-806.

Rowan:1995:OAS

[1719] David J. Rowan, Joseph B. Rasmussen, and Jacob Kalff. Optimal allocation of sampling effort in lake sediment studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2146–2158, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-207.

Hesthagen:1995:DHU

[1720] T. Hesthagen, O. Hegge, J. Skurdal, and B. K. Dervo. Differences in habitat utilization among native, native stocked, and non-native stocked brown trout (Salmo trutta) in a hydroelectric reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2159–2167, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-808.

Weis:1995:SPP

[1721] Judith S. Weis and Peddrick Weis. Swimming performance and predator avoidance by mummichog (Fundulus heteroclitus) larvae after embryonic or larval exposure to methylmercury. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2168–2173, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-809.

Korman:1995:ETA

[1722] Josh Korman, Randall M. Peterman, and Carl J. Walters. Empirical and theoretical analyses of correction of time-series bias in stock-recruitment relationships of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2174–2189, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-810.

Naesje:1995:SFP

[1723] Tor Næsje, Bror Jonssons, and Jostein Skurdal. Spring flood: a primary cue for hatching of river spawning Coregoninae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2190–2196, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-811.

Stemberger:1995:PRA

[1724] Richard S. Stemberger. Pleistocene refuge areas and postglacial dispersal of copepods of the northeastern United States. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2197–2210, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-812.

Rosenberg:1995:ICR

[1725] D. M. Rosenberg, R. E. Hecky, and P. Campbell. Introduction to contaminants research at the Experimental Lakes Area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2211–2212, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-813.

Allan:1995:HRU

[1726] Craig J. Allan. Hydrochemical response of upland Precambrian Shield catchments to additions off H_{2}SO_{4} and NH_{4}NO_{3} during snowmelt. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2213–2228, October 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-814.

Rosenberg:1995:EAP

[1727] David M. Rosenberg, Allen P. Wiens, Bohdan Bilyj, and Llwellyn Armstrong. Experimental acidification of a poor fen in northwestern Ontario: effects on emergence of Chironomidae (Diptera). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2229–2237, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-815.

Turner:1995:DLA

[1728] M. A. Turner, D. W. Schindler, D. L. Findlay, M. B. Jackson, and G. G. C. Robinson. Disruption of littoral algal associations by experimental lake acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10): 2238–2250, October 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-815b.

Turner:1995:GCB

[1729] Michael A. Turner, Leif J. Sigurdson, David L. Findlay, E. Todd Howell, Gordon G. C. Robinson, and John F. Brewster. Growth characteristics of bloom-forming filamentous green algae in the littoral zone of an experimentally acidified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10): 2251–2263, October 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-816.

Turner:1995:EEB

[1730] Michael A. Turner, Bruce E. Townsend, Gordon G. C. Robinson, Brenda J. Hann, and John A. Amaral. Ecological effects of blooms of filamentous green algae in the littoral zone of an acid lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2264–2275, October 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-817.

Bird:1995:FAH

[1731] G. A. Bird, M. Stephenson, R. Roshon, W. J. Schwartz, and M. Motycka. Fate of ^{60}Co and ^{134}Cs added to the hypolimnion of a Canadian Shield lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2276–2289, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
818.

Bird:1995:DMC

[1732] G. A. Bird, M. J. Rosentreter, and W. J. Schwartz. Deformities in the menta of chironomid larvae from the Experimental Lakes Area, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2290–2295, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-819.

Kimura:1995:PRB

[1733] Daniel K. Kimura and Craig R. Kastelle. Perspectives on the relationship between otolith growth and the conversion of isotope activity ratios to fish ages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2296–2303, October 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
820.

Marschall:1995:EDD

[1734] Elizabeth A. Marschall and Larry B. Crowder. Erratum: Densitydependent survival as a function of size in juvenile salmonids in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(10):2304, October 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-821. See [1493].

Taggart:1995:MDM

[1735] John B. Taggart, Eric Verspoor, Paul T. Galvin, Paloma Morán, and Andrew Ferguson. A minisatellite DNA marker for discriminating between European and North American Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2305–2311, November 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-822.

Wahl:1995:EHS

[1736] David H. Wahl. Effect of habitat selection and behavior on vulnerability to predation of introduced fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2312–2319, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-823.

Worthington:1995:DME

[1737] D. G. Worthington, A. J. Fowler, and P. J. Doherty. Determining the most efficient method of age determination for estimating the age structure of a fish population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11): 2320–2326, November 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-224.

Mather:1995:RNP

[1738] Martha E. Mather, Michael J. Vanni, Thomas E. Wissing, Scott A. Davis, and Maynard H. Schaus. Regeneration of nitrogen and phosphorus by bluegill and gizzard shad: effect of feeding history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2327–2338, November 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-825.

Sunyer:1995:PRD

[1739] J. O. Sunyer, E. Gómez, L. Tort, V. Navarro, and J. Quesada. Physiological responses and depression of humoral components of the immune system in gilthead sea bream (Sparus aurata) following daily acute stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2339–2346, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-826.

Haarding:1995:EMA

[1740] Karin C. Hårding and Tero Härkönen. Estimating mean age at sexual maturity in the crabeater seal (Lobodon carcinophagus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2347–2352, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-827.

Dew:1995:NSD

[1741] C. Braxton Dew. The nonrandom size distribution and size-selective transport of age-o Atlantic tomcod (Microgadus tomcod) in the lower Hudson River estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11): 2353–2366, November 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-828.

Kock:1995:ATM

[1742] G. Köck, R. Hofer, and S. Wögrath. Accumulation of trace metals (Cd, Pb, Cu, Zn) in Arctic char (Salvelinus alpinus) from oligotrophic alpine lakes: relation to alkalinity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2367–2376, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-829.

DeBlois:1995:EFA

[1743] E. M. DeBlois and G. A. Rose. Effect of foraging activity on the shoal structure of cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2377–2387, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-830.

Reavie:1995:PEH

[1744] Euan D. Reavie, John P. Smol, and N. Bruce Carmichael. Postsettlement eutrophication histories of six British Columbia (Canada) lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2388–2401, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-831.

Whiles:1995:MPH

[1745] Matt R. Whiles and J. Bruce Wallace. Macroinvertebrate production in a headwater stream during recovery from anthropogenic disturbance and hydrologic extremes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11): 2402–2422, November 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-832.

Johnston:1995:LFD

[1746] Thomas A. Johnston, Marc N. Gaboury, Richard A. Janusz, and Laureen R. Janusz. Larval fish drift in the Valley River, Manitoba: influence of abiotic and biotic factors, and relationships with future year-class strengths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2423–2431, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-833.

Hamilton:1995:IFA

[1747] Steven J. Hamilton and Terry A. Haines. Influence of fluoride on aluminum toxicity to Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2432–2444, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-834.

Campana:1995:SIT

[1748] S. E. Campana, R. K. Mohn, S. J. Smith, and G. A. Chouinard. Spatial implications of a temperature-based growth model for Atlantic cod (Gadus morhua) off the eastern coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2445–2456, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-835. See [2076] and re-

ply [2051].

Bonardelli:1995:EAC

[1749] John C. Bonardelli and John H. Himmelman. Examination of assumptions critical to body component indices: application to the giant scallop Placopecten magellanicus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (11):2457–2469, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f95-836.

Back:1995:MIZ

[1750] Richard C. Back, Vanessa Visman, and Carl J. Watras. Microhomogenization of individual zooplankton species improves mercury and methylmercury determinations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2470–2475, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-837.

Berejikian:1995:EHW

[1751] Barry A. Berejikian. The effects of hatchery and wild ancestry and experience on the relative ability of steelhead trout fry (Oncorhynchus mykiss) to avoid a benthic predator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2476–2482, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-838.

Montgomery:1995:TDM

[1752] Shelagh Montgomery, Marc Lucotte, Pierre Pichet, and Alfonso Mucci. Total dissolved mercury in the water column of several natural and artificial aquatic systems of Northern Quebec (Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2483–2492, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-839.

Grondin:1995:MLP

[1753] Agnès Grondin, Marc Lucotte, Benoit Fortin, and Alfonso Mucci. Mercury and lead profiles and burdens in soils of Quebec (Canada) before and after flooding. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2493– 2506, November 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-840.

Mucci:1995:MRF

[1754] Alfonso Mucci, Shelagh Montgomery, Marc Lucotte, Yanick Plourde, Pierre Pichet, and Huu Van Tra. Mercury remobilization from flooded soils in a hydroelectric reservoir of northern Quebec, La Grande-2: results of a soil resuspension experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2507–2517, November 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-841.

Stein:1995:FWR

[1755] Roy A. Stein, Dennis R. DeVries, and John M. Dettmers. Food-web regulation by a planktivore: exploring the generality of the trophic cascade hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(11):2518– 2526, November 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-842.

Suuronen:1995:SVC

[1756] Petri Suuronen, Timo Turunen, Marko Kiviniemi, and Juha Karjalainen. Survival of vendace (Coregonus albula) escaping from a trawl cod end. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2527–2533, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-843.

Sanderson:1995:PRD

[1757] S. Laurie Sanderson and Joseph J. Cech, Jr. Particle retention during respiration and particulate feeding in the suspension-feeding blackfish, Orthodon microlepidotus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2534–2542, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-244.

Bridcut:1995:DVF

[1758] Emily E. Bridcut and Paul S. Giller. Diet variability and foraging strategies in brown trout (Salmo trutta): an analysis from subpopulations to individuals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2543–2552, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-845.

Mellina:1995:IZM

[1759] Eric Mellina, Joseph B. Rasmussen, and Edward L. Mills. Impact of zebra mussel (Dreissena polymorpha) on phosphorus cycling and chlorophyll in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2553– 2573, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-246.

Morrison:1995:ERC

[1760] Heather Morrison, Rodica Lazar, G. Douglas Haffner, and Tamara Yankovich. Elimination rate constants of 36 PCBs in zebra mussels (Dreissena polymorpha) and exposure dynamics in the Lake St. Clair– Lake Erie corridor. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2574–2582, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-847.

Gundersen:1995:AHS

[1761] Deke T. Gundersen and Lawrence R. Curtis. Acclimation to hard or soft water at weakly alkaline pH influences gill permeability and gill surface calcium binding in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2583–2593, December 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-848.

Stirling:1995:RBE

[1762] Ian Stirling and Nils Are Øritsland. Relationships between estimates of ringed seal (Phoca hispida) and polar bear (Ursus maritimus) populations in the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2594–2612, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-849.

Kenchington:1995:CTM

[1763] E. Kenchington, D. Landry, and C. J. Bird. Comparison of taxa of the mussel Mytilus (Bivalvia) by analysis of the nuclear small-subunit rRNA gene sequence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2613–2620, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-850.

Heath:1995:PBN

[1764] Daniel D. Heath, Paul D. Rawson, and Thomas J. Hilbish. PCR-based nuclear markers identify alien blue mussel (Mytilus spp.) genotypes on the west coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52 (12):2621–2627, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f95-851.

Paul:1995:DIE

[1765] A. J. Paul, D. W. Schindler, A. K. Hardie, and P. R. Leavitt. Direct and indirect effects of predation by a calanoid copepod (subgenus: Hesperodiaptomus) and of nutrients in a fishless alpine lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2628–2638, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-852.

Villanueva:1995:ERG

[1766] Roger Villanueva. Experimental rearing and growth of planktonic Octopus vulgaris from hatching to settlement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2639–2650, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-853.

Hinch:1995:PEC

[1767] Scott G. Hinch, Michael C. Healey, Ron E. Diewert, Michael A. Henderson, Keith A. Thomson, Roy Hourston, and Francis Juanes. Potential effects of climate change on marine growth and survival of Fraser River sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2651–2659, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-854.

Kiriluk:1995:URS

[1768] Richard M. Kiriluk, Mark R. Servos, D. Michael Whittle, Gilbert Cabana, and Joseph B. Rasmussen. Using ratios of stable nitrogen and carbon isotopes to characterize the biomagnification of DDE, mirex, and PCB in a Lake Ontario pelagic food web. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2660–2674, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-855.

Scully:1995:PFH

[1769] N. M. Scully, D. J. McQueen, D. R. S. Lean, and W. J. Cooper. Photochemical formation of hydrogen peroxide in lakes: effects of dissolved organic carbon and ultraviolet radiation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2675–2681, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-856.

Berglund:1995:ESS

[1770] Ingemar Berglund. Effects of size and spring growth on sexual maturation in 1+ Atlantic salmon (Salmo salar) male parr: interactions with smoltification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2682–2694, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-857.

Ricciardi:1995:LSE

[1771] Anthony Ricciardi, Fred L. Snyder, David O. Kelch, and Henry M. Reiswig. Lethal and sublethal effects of sponge overgrowth on introduced dreissenid mussels in the Great Lakes–St. Lawrence River system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2695–2703, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-858.

Meffe:1995:GDT

[1772] Gary K. Meffe, Stephen C. Weeks, Margaret Mulvey, and K. L. Kandl. Genetic differences in thermal tolerance of eastern mosqyitofish (Gambusia holbrooki; Poeciliidae) from ambient and thermal ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2704–2711, December 1995. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-259.

Cadrin:1995:DAL

[1773] Steven X. Cadrin. Discrimination of American lobster (Homarus americanus) stocks off southern New England on the basis of secondary sex character allometry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2712–2723, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-860.

Payne:1995:AHS

[1774] J. F. Payne, L. L. Fancey, J. Hellou, M. J. King, and G. L. Fletcher. Aliphatic hydrocarbons in sediments: a chronic toxicity study with winter flounder (Pleuronectes americanus) exposed to oil well drill cuttings. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2724–2735, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-861.

Cain:1995:IGC

[1775] D. J. Cain, S. N. Luoma, and E. V. Axtmann. Influence of gut content in immature aquatic insects on assessments of environmental metal contamination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2736– 2746, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-862.

Bigelow:1995:GPP

[1776] Keith A. Bigelow, Juble T. Jones, and Gerard T. DiNardo. Growth of the Pacific pomfret, Brama japonica: a comparison between otolith and length–frequency (MULTIFAN) analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2747–2756, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-863.

Hammill:1995:GSH

[1777] M. O. Hammill and J. F. Gosselin. Grey seal (Halichoerus grypus) from the Northwest Atlantic: female reproductive rates, age at first birth, and age of maturity in males. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12): 2757–2761, December 1995. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-864.

Adkison:1995:PDP

[1778] Milo D. Adkison. Population differentiation in Pacific salmons: local adaptation genetic drift, or the environment? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2762–2777, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-865.

Shaklee:1995:EEC

[1779] James B. Shaklee and Natalya V. Varnavskaya. Erratum: Electrophoretic characterization of odd-year pink salmon (Oncorhynchus gorbuscha) populations from the Pacific coast of Russia, and comparison with selected North American populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 52(12):2778, December 1995. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-866. See [1464].

Hall:1996:PAL

[1780] Roland I. Hall and John P. Smol. Paleolimnological assessment of longterm water-quality changes in south-central Ontario lakes affected by cottage development and acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):1–17, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-171.

Boersma:1996:SPM

[1781] Maarten Boersma, Onno F. R. van Tongeren, and Wolf M. Mooij. Seasonal patterns in the mortality of Daphnia species in a shallow lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):18–28, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-154a.

Lei:1996:FDZ

[1782] Jin Lei, Barry S. Payne, and Shiao Y. Wang. Filtration dynamics of the zebra mussel, Dreissena polymorpha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):29–37, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-164.

Hatcher:1996:DMP

[1783] Bruce G. Hatcher, Robert E. Scheibling, Myriam A. Barbeau, Alan W. Hennigar, Lawrence H. Taylor, and Anthony J. Windust. Dispersion and mortality of a population of sea scallop (Placopecten magellanicus) seeded in a tidal channel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):38–54, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-170.

Watanabe:1996:OES

[1784] Yoshiro Watanabe, Hiromu Zenitani, and Ryo Kimura. Offshore expansion of spawning of the Japanese sardine, Sardinops melanostictus, and its implication for egg and larval survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):55–61, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-153.

Larsson:1996:PPS

[1785] Per Larsson, Cecilia Backe, Gudrun Bremle, Anders Eklöv, and Lennart Okla. Persistent pollutants in a salmon population (Salmo salar) of the southern Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):62–69, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-157.

Hightower:1996:CRC

[1786] Joseph E. Hightower. Comparison of reliability of catch-at-age analyses using auxiliary survey estimates of relative or absolute recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):70–79, January 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-168.

Graham:1996:SCH

[1787] W. Douglas Graham, John E. Thorpe, and Neil B. Metcalfe. Seasonal current holding performance of juvenile Atlantic salmon in relation to temperature and smolting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):80–86, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-167.

Clay:1996:STS

[1788] Allen Clay and Martin Castonguay. In situ target strengths of Atlantic cod (Gadus morhua) and Atlantic mackerel (Scomber scombrus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):87–98, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-201.

VanNieuwenhuyse:1996:PCR

[1789] Erwin E. Van Nieuwenhuyse and John R. Jones. Phosphorus chlorophyll relationship in temperate streams and its variation with stream catchment area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):99–105, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-166.

Swain:1996:RBG

[1790] Douglas P. Swain and Roderick Morin. Relationships between geographic distribution and abundance of American plaice (Hippoglossoides platessoides) in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):106–119, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-163.

Morris:1996:MRT

[1791] Dianne B. Morris, Kenny R. Richard, and Jonathan M. Wright. Microsatellites from rainbow trout (Oncorhynchus mykiss) and their use for genetic study of salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):120–126, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-161.

Henderson:1996:RWL

[1792] Bryan A. Henderson, Jonathan L. Wong, and Stephen J. Nepszy. Reproduction of walleye in Lake Erie: allocation of energy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):127–133, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-162.

Wainman:1996:BSP

[1793] B. C. Wainman, S. S. Hincks, N. K. Kaushik, and G. L. Mackie. Biofilm and substrate preference in the dreissenid larvae of Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):134–140, January 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-213.

Craig:1996:ETD

[1794] J. Kevin Craig, Chris J. Foote, and Chris C. Wood. Evidence for temperature-dependent sex determination in sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1): 141–147, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-152.

Walters:1996:FER

[1795] Carl Walters and Ana M. Parma. Fixed exploitation rate strategies for coping with effects of climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):148–158, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-151.

Mertz:1996:ECA

[1796] G. Mertz and R. A. Myers. An extended cohort analysis: incorporating the effect of seasonal catches. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):159–163, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-156.

Bilby:1996:INC

[1797] Robert E. Bilby, Brian R. Fransen, and Peter A. Bisson. Incorporation of nitrogen and carbon from spawning Coho salmon into the trophic system of small streams: evidence from stable isotopes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):164–173, January 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-159.

Uutala:1996:PRL

[1798] Allen J. Uutala and John P. Smol. Paleolimnological reconstructions of long-term changes in fisheries status in Sudbury area lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):174–180, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-150.

Miller:1996:SIC

[1799] Kristina M. Miller, Ruth E. Withler, and Terry D. Beacham. Stock identification of Coho salmon (Oncorhynchus kisutch) using minisatellite DNA variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1): 181–195, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-165.

Skalski:1996:RAE

[1800] John R. Skalski. Regression of abundance estimates from mark recapture surveys against environmental covariates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):196–204, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-169.

Metcalfe-Smith:1996:IBF

[1801] Janice L. Metcalfe-Smith, Roger H. Green, and Lee C. Grapentine. Influence of biological factors on concentrations of metals in the tissues of freshwater mussels (Elliptio complanata and Lampsilis radiata radiata) from the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (1):205–219, January 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-158.

Limburg:1996:GMY

[1802] Karen E. Limburg. Growth and migration of 0-year American shad (Alosa sapidissima) in the Hudson River estuary: otolith microstructural analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(1):220–238, January 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-160.

Evans:1996:UEO

[1803] Geoffrey T. Evans. Using the elementary operations of sequential population analysis to display problems in catch or survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):239–243, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-191.

Goldstein:1996:CMC

[1804] Robert M. Goldstein, Mark E. Brigham, and Joseph C. Stauffer. Comparison of mercury concentrations in liver, muscle, whole bodies, and composites of fish from the Red River of the North. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):244–252, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-203.

Hamel:1996:EDS

[1805] Jean-Francois Hamel and Annie Mercier. Early development, settlement, growth, and spatial distribution of the sea cucumber Cucumaria frondosa (Echinodermata: Holothuroidea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):253–271, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-186.

Smith:1996:DBP

[1806] Richard J. Smith, Keith A. Hobson, Heather N. Koopman, and David M. Lavigne. Distinguishing between populations of fresh- and salt-water harbour seals (Phoca vitulina) using stable-isotope ratios and fatty acid profiles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):272–279, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-192.

Machias:1996:APF

[1807] A. Machias and N. Tsimenides. Anatomical and physiological factors affecting the swim-bladder cross-section of the sardine Sardina pilchardus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):280–287, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-188.

Schwinghamer:1996:QIT

[1808] P. Schwinghamer, J. Y. Guigné, and W. C. Siu. Quantifying the impact of trawling on benthic habitat structure using high resolution acoustics and chaos theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2): 288–296, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-277.

Hudson:1996:STV

[1809] Lori A. Hudson and Jan J. H. Ciborowski. Spatial and taxonomic variation in incidence of mouthpart deformities in midge larvae (Diptera: Chironomidae: Chironomini). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (2):297–304, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-199.

Fraser:1996:IBA

[1810] Sandra Fraser, Vytenis Gotceitas, and Joseph A. Brown. Interactions between age-classes of Atlantic cod and their distribution among bottom substrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):305–314, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-183.

Almond:1996:SSS

[1811] M. Jane R. Almond, Ellen Bentzen, and William D. Taylor. Size structure and species composition of plankton communities in deep Ontario lakes with and without Mysis relicta and planktivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):315–325, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-184.

Zhou:1996:ALD

[1812] Shengying Zhou, Robert G. Ackman, and Carol Morrison. Adipocytes and lipid distribution in the muscle tissue of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):326–332, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-197.

Brattey:1996:GVW

[1813] J. Brattey and W. S. Davidson. Genetic variation within Pseudoterranova decipiens (Nematoda: Ascaridoidea) from Canadian Atlantic marine fishes and seals: characterization by RFLP analysis of genomic DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):333–341, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-196.

Colbourne:1996:DFB

[1814] John Kenneth Colbourne, Bryan D. Neff, Jonathan M. Wright, and Mart R. Gross. DNA fingerprinting of bluegill sunfish (Lepomis macrochirus) using (GT)_{n} microsatellites and its potential for assessment of mating success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (2):342–349, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-179.

Jenkins:1996:SVR

[1815] Gregory P. Jenkins, Melissa J. Wheatley, and Alistair G. B. Poore. Spatial variation in recruitment, growth, and feeding of postsettlement King George whiting, Sillaginodes punctata, associated with seagrass beds of Port Phillip Bay, Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (2):350–359, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-195.

Cargnelli:1996:TDF

[1816] Luca M. Cargnelli and Mart R. Gross. The temporal dimension in fish recruitment: birth date, body size, and size-dependent survival in a sunfish (bluegill: Lepomis macrochirus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):360–367, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-193.

Gong:1996:CVS

[1817] Bang Q. Gong, Anthony P. Farrell, Anders Kiessling, and David Higgs. Coronary vascular smooth muscle responses to swimming challenges in juvenile salmonid fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):

368–371, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-182.

Veinott:1996:IAP

[1818] Geoffrey I. Veinott and R. Jack Cornett. Identification of annually produced opaque bands in the shell of the freshwater mussel Elliptio complanata using the seasonal cycle of δ ^{18}O. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):372–379, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-187.

Beacham:1996:SIC

[1819] Terry D. Beacham, Ruth E. Withler, and Tracy A. Stevens. Stock identification of Chinook salmon (Oncorhynchus tshawytscha) using minisatellite DNA variation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2): 380–394, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-198.

Post:1996:URF

[1820] John R. Post, Rena Vandenbos, and Donald J. McQueen. Uptake rates of food-chain and waterborne mercury by fish: field measurements, a mechanistic model, and an assessment of uncertainties. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):395–407, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/f95-178. See erratum [1906].

Matschak:1996:ITM

[1821] T. W. Matschak and N. C. Stickland. The influence of temperature on mRNA levels for muscle contractile protein and a proto-oncogene associated with cell division in Atlantic salmon (Salmo salar L.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):408–413, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-194.

Kinas:1996:BFS

[1822] Paul Gerhard Kinas. Bayesian fishery stock assessment and decision making using adaptive importance sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):414–423, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-189.

Bernatchez:1996:MDD

[1823] Louis Bernatchez and Sylvain Martin. Mitochondrial DNA diversity in anadromous rainbow smelt, Osmerus mordax Mitchill: a genetic assessment of the member-vagrant hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):424–433, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-180.

Dittman:1996:TIN

[1824] Andrew H. Dittman, Thomas P. Quinn, and Gabrielle A. Nevitt. Timing of imprinting to natural and artificial odors by Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2): 434–442, February 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-185.

Snodgrass:1996:FAO

[1825] Joel W. Snodgrass, A. Lawrence Bryan, Jr., Robert F. Lide, and Gordon M. Smith. Factors affecting the occurrence and structure of fish assemblages in isolated wetlands of the upper coastal plain, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):443–454, February 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-200.

Bigler:1996:RST

[1826] Brian S. Bigler, David W. Welch, and John H. Helle. A review of size trends among North Pacific salmon (Oncorhynchus spp.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):455–465, February 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-181.

Brattey:1996:EIL

[1827] John Brattey. Erratum: Identification of larval Contracaecum osculatum s.l. and Phocascaris sp. (Nematoda: Ascaridoidea) from marine fishes by allozyme electrophoresis and discriminant function analysis of morphometric data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(2):466, February 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-010. See

[1671].

Ward:1996:EAM

[1828] F. J. Ward and J. A. T. Boubée. Effect of age at maturity on lifespan and generation time of New Zealand common smelt, Retropinna retropinna (Richardson). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (3):467–473, March 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-218.

Dougherty:1996:MDA

[1829] J. D. Dougherty, W. S. Moore, and J. L. Ram. Mitochondrial DNA analysis of round goby (Neogobius melanostomus) and tubenose goby (Proterorhinus marmoratus) in the Great Lakes basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):474–480, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-211.

Beamish:1996:GSJ

[1830] F. W. H. Beamish, J.-A. Jebbink, A. Rossiter, and D. L. G. Noakes. Growth strategy of juvenile lake sturgeon (Acipenser fulvescens) in a northern river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):481–489, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-222.

Prairie:1996:EPP

[1831] Y. T. Prairie. Evaluating the predictive power of regression models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):490–492, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-204.

Vyverman:1996:DBD

[1832] W. Vyverman, R. Vyverman, V. S. Rajendran, and P. Tyler. Distribution of benthic diatom assemblages in Tasmanian highland lakes and their possible use as indicators of environmental changes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):493–508, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-220.

Araki:1996:EEH

[1833] K. Araki, N. Fujikawa, I. Nakayama, H. Nagoya, and H. Onozato. Early expression of a hatching enzyme gene in masu salmon (Oncorhynchus masou) embryos. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):509–512, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-210.

Beland:1996:RBR

[1834] K. F. Beland. The relation between redd counts and Atlantic salmon (Salmo salar) parr populations in the Dennys River, Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):513–519, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/f95-216. See erratum [2027].

Welch:1996:PDG

[1835] H. E. Welch, T. D. Siferd, and P. Bruecker. Population densities, growth, and respiration of the chaetognath Parasagitta elegans in the Canadian high Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):520–527, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f95-227.

Hobson:1996:SCN

[1836] K. A. Hobson, D. M. Schell, D. Renouf, and E. Noseworthy. Stable carbon and nitrogen isotopic fractionation between diet and tissues of captive seals: implications for dietary reconstructions involving marine mammals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):528–533, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-209.

Josephson:1996:ABE

[1837] D. C. Josephson and W. D. Youngs. Association between emigration and age structure in populations of brook trout (Salvelinus fontinalis) in Adirondack lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (3):534–541, March 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-223.

Mathur:1996:TPS

[1838] D. Mathur, P. G. Heisey, E. T. Euston, J. R. Skalski, and S. Hays. Turbine passage survival estimation for Chinook salmon smolts (Oncorhynchus tshawytscha) at a large dam on the Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):542–549, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-206.

Guderley:1996:PSA

[1839] H. Guderley, J.-D. Dutil, and D. Pelletier. The physiological status of Atlantic cod, Gadus morhua, in the wild and the laboratory: estimates of growth rates under field conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):550–557, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-219.

Chen:1996:NFM

[1840] X. H. Chen and P. A. Shelton. A nonparametric forecast model of inshore Atlantic cod (Gadus morhua) landings based on biomass, cumulative landings, and water temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):558–562, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-208.

Bene:1996:EMC

[1841] C. Béné. Effects of market constraints, the remuneration system, and resource dynamics on the spatial distribution of fishing effort. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):563–571, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-212.

Lancaster:1996:IDL

[1842] J. Lancaster, A. G. Hildrew, and C. Gjerlov. Invertebrate drift and longitudinal transport processes in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):572–582, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-217.

Heist:1996:GPS

[1843] E. J. Heist, J. A. Musick, and J. E. Graves. Genetic population structure of the shortfin mako (Isurus oxyrinchus) inferred from restriction fragment length polymorphism analysis of mitochondrial DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):583–588, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-245.

Lacroix:1996:TEA

[1844] G. L. Lacroix and J. Korman. Timing of episodic acidification in Atlantic salmon rivers influences evaluation of mitigative measures and recovery forecasts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):589–599, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-224a.

Jacobson:1996:BDN

[1845] L. D. Jacobson and R. D. Vetter. Bathymetric demography and niche separation of thornyhead rockfish: Sebastolobus alascanus and Sebastolobus altivelis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):600–609, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-207a.

Kjesbu:1996:VAE

[1846] O. S. Kjesbu, P. Solemdal, P. Bratland, and M. Fonn. Variation in annual egg production in individual captive Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):610–620, March 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-215.

Zhang:1996:RDP

[1847] Y. Zhang and E. E. Prepas. Regulation of the dominance of planktonic diatoms and cyanobacteria in four eutrophic hardwater lakes by nutrients, water column stability, and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):621–633, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-205.

Ruzzante:1996:GDB

[1848] D. E. Ruzzante, C. T. Taggart, D. Cook, and S. Goddard. Genetic differentiation between inshore and offshore Atlantic cod (Gadus morhua) off Newfoundland: microsatellite DNA variation and antifreeze level. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):634–645, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-228.

Arnott:1996:NPR

[1849] D. L. Arnott and M. J. Vanni. Nitrogen and phosphorus recycling by the zebra mussel (Dreissena polymorpha) in the western basin of Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):646–659, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-214.

Martel:1996:GRI

[1850] G. Martel. Growth rate and influence of predation risk on territoriality in juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):660–669, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-225.

Peters:1996:ARE

[1851] R. H. Peters, G. E. Ball, R. Carignan, P. D. N. Hebert, and E. E. Prepas. An assessment of research in evolution and ecology supported by the Natural Sciences and Engineering Research Council of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):670–680, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-226.

Baras:1996:CHE

[1852] E. Baras. Commentaire a l’hypothèse de l’éternel retour de Cury (1994) : proposition d’un mécanisme fonctionnel dynamique. (French) [Commentary on the eternal return hypothesis of Cury (1994): proposal of a dynamic functional mechanism]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(3):681–684, March 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-221. See [1329].

Wainright:1996:SVT

[1853] S. C. Wainright, C. M. Fuller, R. H. Michener, and R. A. Richards. Spatial variation of trophic position and growth rate of juvenile striped bass(Morone saxatilis) in the Delaware River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):685–692, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-229.

Beamish:1996:LHA

[1854] F. W. H. Beamish and S. Lowartz. Larval habitat of American brook lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):693–700, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-232.

Powell:1996:RAB

[1855] M. D. Powell and S. F. Perry. Respiratory and acid-base disturbances in rainbow trout (Oncorhynchus mykiss) blood during exposure to chloramine t, paratoluenesulphonamide, and hypochlorite. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):701–708, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-250.

Cortes:1996:CDT

[1856] E. Cortés and G. R. Parsons. Comparative demography of two populations of the bonnethead shark (Sphyrna tiburo). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):709–718, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-240.

Hume:1996:JSR

[1857] J. M. B. Hume, K. S. Shortreed, and K. F. Morton. Juvenile sockeye rearing capacity of three lakes in the Fraser River system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):719–733, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-237.

Rowan:1996:MBC

[1858] David J. Rowan and Joseph B. Rasmussen. Measuring the bioenergetic cost of fish activity in situ using a globally dispersed radiotracer (^{137}Cs). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):734–745, April 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-046.

Copp:1996:WDF

[1859] G. H. Copp and V. Kovác. When do fish with indirect development become juveniles? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):746–752, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
252.

Kandl:1996:REM

[1860] K. L. Kandl and A. J. Thompson. Responses of eastern mosquitofish (Gambusia holbrooki) to acute salinity stress: do family effects obscure inbreeding and genotype effects? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):753–760, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-239.

Huppert:1996:DPP

[1861] D. D. Huppert, G. M. Ellis, and B. Noble. Do permit prices reflect the discounted value of fishing? Evidence from Alaska’s commercial salmon fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):761–768, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-242.

Simila:1996:ODK

[1862] T. Similä, J. C. Holst, and I. Christensen. Occurrence and diet of killer whales in northern Norway: seasonal patterns relative to the distribution and abundance of Norwegian spring-spawning herring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):769–779, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-253.

Kock:1996:SPM

[1863] G. Köck, M. Triendl, and R. Hofer. Seasonal patterns of metal accumulation in Arctic char (Salvelinus alpinus) from an oligotrophic alpine lake related to temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (4):780–786, April 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-243.

Letcher:1996:VSL

[1864] B. H. Letcher, J. A. Rice, L. B. Crowder, and K. A. Rose. Variability in survival of larval fish: disentangling components with a generalized individual-based model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (4):787–801, April 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-241.

Wilson:1996:GPT

[1865] R. W. Wilson, C. M. Wood, and D. F. Houlihan. Growth and protein turnover during acclimation to acid and aluminum in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):802–811, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f95-235.

Hill:1996:MSB

[1866] W. R. Hill, A. J. Stewart, and G. E. Napolitano. Mercury speciation and bioaccumulation in lotic primary producers and primary consumers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):812–819, April 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-231.

Jensen:1996:BHL

[1867] A. L. Jensen. Beverton and Holt life history invariants result from optimal trade-off of reproduction and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):820–822, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-233.

Hendry:1996:GEP

[1868] A. P. Hendry, T. P. Quinn, and F. M. Utter. Genetic evidence for the persistence and divergence of native and introduced sockeye salmon (Oncorhynchus nerka) within Lake Washington, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):823–832, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-238.

Scribner:1996:ICN

[1869] K. T. Scribner, J. R. Gust, and R. L. Fields. Isolation and characterization of novel salmon microsatellite loci: cross-species amplification and population genetic applications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):833–841, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-254.

Bachmann:1996:RBT

[1870] R. W. Bachmann, B. L. Jones, D. D. Fox, M. Hoyer, L. A. Bull, and D. E. Canfield, Jr. Relations between trophic state indicators and fish in Florida (U.S.A.) lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (4):842–855, April 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-236.

Findlay:1996:EIP

[1871] D. L. Findlay and S. E. M. Kasian. The effect of incremental pH recovery on the Lake 223 phytoplankton community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):856–864, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-251.

Roseman:1996:ECB

[1872] E. F. Roseman, E. L. Mills, J. L. Forney, and L. G. Rudstam. Evaluation of competition between age-0 yellow perch (Perca flavescens) and gizzard shad (Dorosoma cepedianum) in Oneida Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):865–874, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-249.

Gu:1996:RBS

[1873] B. Gu, C. L. Schelske, and M. Brenner. Relationship between sediment and plankton isotope ratios (δ ^{13}C and δ ^{15}N) and primary productivity in Florida lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (4):875–883, April 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-248.

Arsenault:1996:OHS

[1874] D. J. Arsenault and J. H. Himmelman. Ontogenic habitat shifts of the Iceland scallop, Chlamys islandica (Müller, 1776), in the northern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):884–895, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f95-
234.

Grubaugh:1996:LCM

[1875] J. W. Grubaugh, J. B. Wallace, and E. S. Houston. Longitudinal changes of macroinvertebrate communities along an Appalachian stream continuum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):896–909, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f95-247.

Lee:1996:MOT

[1876] J. R. Post J. A. Lee. Metabolic ontogeny of teleost fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):910–923, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-278.

Livingstone:1996:PHO

[1877] D. M. Livingstone and D. M. Imboden. The prediction of hypolimnetic oxygen profiles: a plea for a deductive approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):924–932, April 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f95-230. See com-

ment [2151] and reply [2152].

Sheridan:1996:CES

[1878] A. K. Sheridan. Comment — Estimation of selection differentials from fish scales: a step towards evaluating genetic alteration of fish size in exploited populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4): 933, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f95-257. See [1237] and reply [1879].

Miller:1996:RES

[1879] L. M. Miller and A. R. Kapuscinski. Reply — Estimation of selection differentials from fish scales: a step towards evaluating genetic alteration of fish size in exploited populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):934–935, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f95-256. See [1237, 1878, 1237, 1880].

McAllister:1996:CSM

[1880] M. K. McAllister and D. M. Gillis. Commentary on Sheridan (1996) and Miller and Kapuscinski (1996) concerning “Estimation of selection differentials from fish scales: a step towards evaluating genetic alteration of fish size in exploited populations”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):936–939, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-045. See [1237, 1881, 1879, 1878].

Kathuria:1996:CMG

[1881] A. Kathuria, G. N. G. Gordon, and A. K. Sheridan. Commentary on McAllister and Gillis (1996) concerning “Estimation of selection differentials from fish scales: a step towards evaluating genetic alteration of fish size in exploited populations”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(4):940–941, April 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-072. See [1880].

Hutchings:1996:STV

[1882] J. A. Hutchings. Spatial and temporal variation in the density of northern cod and a review of hypotheses for the stock’s collapse. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):943–962, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-097.

Vignaux:1996:ASS

[1883] M. Vignaux. Analysis of spatial structure in fish distribution using commercial catch and effort data from the New Zealand hoki fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):963–973, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-035.

Tringali:1996:GSS

[1884] M. D. Tringali and T. M. Bert. The genetic stock structure of common snook (Centropomus undecimalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):974–984, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-027.

Hindar:1996:LWA

[1885] A. Hindar, F. Kroglund, E. Lydersen, A. Skiple, and R. Høgberget. Liming of wetlands in the acidified Lake Røynelandsvatn catchment in southern Norway: effects on stream water chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):985–993, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-032.

Cyr:1996:BSS

[1886] H. Cyr and R. H. Peters. Biomass-size spectra and the prediction of fish biomass in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):994–1006, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
033.

Hovgaard:1996:TSA

[1887] H. Hovgård. A two-step approach to estimating selectivity and fishing power of research gill nets used in Greenland waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1007–1013, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-022.

Hovgaard:1996:ETD

[1888] H. Hovgård. Effect of twine diameter on fishing power of experimental gill nets used in Greenland waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1014–1017, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-023.

Premdas:1996:ETE

[1889] P. D. Premdas and C. D. Metcalfe. Experimental transmission of epidermal lip papillomas in white sucker, Catostomus commersoni. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1018–1029, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-034.

Stead:1996:ERS

[1890] S. M. Stead, D. F. Houlihan, H. A. McLay, and R. Johnstone. Effect of ration and seawater transfer on food consumption and growth of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (5):1030–1037, May 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-040.

Vitic:1996:MDA

[1891] R. Vitic and C. Strobeck. Mitochondrial DNA analysis of populations of lake trout (Salvelinus namaycush) in west-central Canada: implications for stock identification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (5):1038–1047, May 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-36.

Evans:1996:APA

[1892] M. S. Evans, M. T. Arts, and R. D. Robarts. Algal productivity, algal biomass, and zooplankton biomass in a phosphorus-rich, saline lake: deviations from regression model predictions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1048–1060, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-037.

Montgomery:1996:SBS

[1893] D. R. Montgomery, J. M. Buffington, N. P. Peterson, D. Schuett-Hames, and T. P. Quinn. Stream-bed scour, egg burial depths, and the influence of salmonid spawning on bed surface mobility and embryo survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1061–1070, May 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-028.

Rose:1996:SWF

[1894] K. A. Rose, J. A. Tyler, R. C. Chambers, G. Klein-MacPhee, and D. J. Danila. Simulating winter flounder population dynamics using coupled individual-based young-of-the-year and age-structured adult models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1071–1091, May 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-031.

Millard:1996:PPT

[1895] E. S. Millard, D. D. Myles, O. E. Johannsson, and K. M. Ralph. Phytoplankton photosynthesis at two index stations in Lake Ontario 1987– 1992: assessment of the long-term response to phosphorus control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1092–1111, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-029.

Millard:1996:SPD

[1896] E. S. Millard, D. D. Myles, O. E. Johannsson, and K. M. Ralph. Seasonal phosphorus deficiency of Lake Ontario phytoplankton at two index stations: light versus. phosphorus limitation of growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1112–1124, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-030.

Perin:1996:EPF

[1897] S. Perin, F. R. Pick, D. R. S. Lean, and A. Mazumder. Effects of planktivorous fish and nutrient additions on primary production of shallow versus deep (stratified) lake enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1125–1132, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-024.

Larocque:1996:SAR

[1898] I. Larocque, A. Mazumder, M. Proulx, D. R. S. Lean, and F. R. Pick. Sedimentation of algae: relationships with biomass and size distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1133–1142, May 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-025.

Strayer:1996:ASE

[1899] D. L. Strayer, J. Powell, P. Ambrose, L. C. Smith, M. L. Pace, and D. T. Fischer. Arrival, spread, and early dynamics of a zebra mussel (Dreissena polymorpha) population in the Hudson River estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1143–1149, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-038.

Fortin:1996:DGB

[1900] R. Fortin, P. Dumont, and S. Guénette. Determinants of growth and body condition of lake sturgeon (Acipenser fulvescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1150–1156, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-043.

Weeks:1996:RRQ

[1901] S. C. Weeks. A reevaluation of the Red Queen model for the maintenance of sex in a clonal-sexual fish complex (Poeciliidae: Poeciliopsis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1157–1164, May 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-041.

McKay:1996:PPS

[1902] S. J. McKay, R. H. Devlin, and M. J. Smith. Phylogeny of Pacific salmon and trout based on growth hormone type-2 and mitochondrial NADH dehydrogenase subunit 3 DNA sequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1165–1176, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-042.

Clemons:1996:CPI

[1903] J. H. Clemons, L. E. J. Lee, C. R. Myers, D. G. Dixon, and N. C. Bols. Cytochrome P4501A1 induction by polychlorinated biphenyls (PCBs) in liver cell lines from rat and trout and the derivation of toxic equivalency factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1177–1185, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-039.

Pelletier:1996:DMP

[1904] D. Pelletier and P. Magal. Dynamics of a migratory population under different fishing effort allocation schemes in time and space. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1186–1199, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f95-255.

Pitcher:1996:ORJ

[1905] T. J. Pitcher. Obituary: Raymond John Heaphy Beverton, CBE, FRS. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1200–1201, May 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-145.

Post:1996:EUR

[1906] J. R. Post, R. Vandenbos, and D. J. McQueen. Erratum: Uptake rates of food-chain and waterborne mercury by fish: field measurements, a mechanistic model, and an assessment of uncertainties. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(5):1202, May 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-122. See [1820].

Jay:1996:DBT

[1907] C. V. Jay. Distribution of bottom-trawl fish assemblages over the continental shelf and upper slope of the U.S. west coast, 1977–1992. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1203–1225, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-055.

Paradis:1996:VFE

[1908] A. R. Paradis, P. Pepin, and J. A. Brown. Vulnerability of fish eggs and larvae to predation: review of the influence of the relative size of prey and predator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1226–1235, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-051.

Persson:1996:SEM

[1909] J. Persson and L. Håkanson. A simple empirical model to predict deepwater turnover time in coastal waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1236–1245, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-048.

Williams:1996:DCE

[1910] P. J. Williams, J. A. Brown, V. Gotceitas, and P. Pepin. Developmental changes in escape response performance of five species of marine larval fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1246–1253, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-052.

Weyhenmeyer:1996:ISA

[1911] G. A. Weyhenmeyer. The influence of stratification on the amount and distribution of different settling particles in Lake Erken. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1254–1262, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-057.

Welsh:1996:EAC

[1912] P. G. Welsh, J. L. Parrott, D. G. Dixon, P. V. Hodson, D. J. Spry, and G. Mierle. Estimating acute copper toxicity to larval fathead minnow (Pimephales promelas) in soft water from measurements of dissolved organic carbon, calcium, and pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1263–1271, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-063.

Xiao:1996:FEE

[1913] Y. Xiao. A framework for evaluating experimental designs for estimating rates of fish movement from tag recoveries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1272–1280, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-050.

Schnute:1996:MOA

[1914] J. T. Schnute and A. R. Kronlund. A management oriented approach to stock recruitment analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (6):1281–1293, June 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-069.

Trippel:1996:IMH

[1915] E. A. Trippel, J. Y. Wang, M. B. Strong, L. S. Carter, and J. D. Conway. Incidental mortality of harbour porpoise (Phocoena phocoena) by the gillnet fishery in the lower Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1294–1300, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-060.

Yan:1996:RCZ

[1916] N. D. Yan, W. Keller, K. M. Somers, T. W. Pawson, and R. E. Girard. Recovery of crustacean zooplankton communities from acid and metal contamination: comparing manipulated and reference lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1301–1327, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-065.

Yan:1996:DGE

[1917] N. D. Yan, P. G. Welsh, H. Lin, D. J. Taylor, and J.-M. Filion. Demographic and genetic evidence of the long-term recovery of (Daphnia galeata mendotae) (Crustacea: Daphniidae) in Sudbury lakes following additions of base: the role of metal toxicity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1328–1344, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-066.

Claytor:1996:SMA

[1918] R. R. Claytor. In-season management of Atlantic salmon (Salmo salar): an example using southern Gulf of St. Lawrence rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1345–1359, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-064.

Shannon:1996:ODR

[1919] J. P. Shannon, D. W. Blinn, P. L. Benenati, and K. P. Wilson. Organic drift in a regulated desert river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1360–1369, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-062.

Gowan:1996:MBT

[1920] C. Gowan and K. D. Fausch. Mobile brook trout in two high-elevation Colorado streams: reevaluating the concept of restricted movement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1370–1381, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-058.

Desrosiers:1996:EDE

[1921] R. R. Desrosiers, J. Désilets, and F. Dubé. Early developmental events following fertilization in the giant scallop Placopecten magellanicus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1382–1392, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-071.

Elliott:1996:ETP

[1922] J. K. Elliott and W. C. Leggett. The effect of temperature on predation rates of a fish (Gasterosteus aculeatus) and a jellyfish (Aurelia aurita) on larval capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1393–1402, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-049.

Young:1996:SMH

[1923] M. K. Young. Summer movements and habitat use by Colorado River cutthroat trout (Oncorhynchus clarki pleuriticus) in small, montane streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1403–1408, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-054.

Sanderson:1996:RDP

[1924] B. L. Sanderson and T. M. Frost. Regulation of dinoflagellate populations: relative importance of grazing, resource limitation, and recruitment from sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6): 1409–1417, June 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-059.

Ballantyne:1996:PNF

[1925] J. S. Ballantyne, F. Mercure, M. F. Gerrits, G. Van Der Kraak, S. McKinley, D. W. Martens, S. G. Hinch, and R. E. Diewert. Plasma nonesterified fatty acid profiles in male and female sockeye salmon, (Oncorhynchus nerka), during the spawning migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1418–1426, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-056.

Plakas:1996:UTD

[1926] S. M. Plakas, K. R. El Said, G. R. Stehly, W. H. Gingerich, and J. L. Allen. Uptake, tissue distribution, and metabolism of malachite green in the channel catfish (Ictalurus punctatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1427–1433, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-061.

Ricciardi:1996:IDI

[1927] A. Ricciardi, F. G. Whoriskey, and J. B. Rasmussen. Impact of the(Dreissena) invasion on native unionid bivalves in the upper St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1434– 1444, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-068.

Mitchell:1996:SIN

[1928] M. J. Mitchell, E. L. Mills, N. Idrisi, and R. Michener. Stable isotopes of nitrogen and carbon in an aquatic food web recently invaded by (Dreissena polymorpha) (Pallas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (6):1445–1450, June 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-053.

Stoermer:1996:PER

[1929] E. F. Stoermer, G. Emmert, M. L. Julius, and C. L. Schelske. Paleolimnologic evidence of rapid recent change in Lake Erie’s trophic status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1451–1458, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-067.

Gauldie:1996:LAF

[1930] R. W. Gauldie, S. K. Sharma, and C. E. Helsley. LIDAR applications to fisheries monitoring problems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1459–1468, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-070.

Nurnberg:1996:CPB

[1931] G. K. Nürnberg. Comment: Phosphorus budgets and stoichiometry during the open-water season in two unmanipulated lakes in the Experimental Lakes Area, northwestern Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(6):1469–1471, June 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-073. See [1437].

Frank:1996:REC

[1932] K. T. Frank, J. E. Carscadden, and J. E. Simon. Recent excursions of capelin (Mallotus villosus) to the Scotian Shelf and Flemish Cap during anomalous hydrographic conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1473–1486, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-078.

Hawkins:1996:CSV

[1933] D. K. Hawkins and T. P. Quinn. Critical swimming velocity and associated morphology of juvenile coastal cutthroat trout (Oncorhynchus clarki clarki), steelhead trout (Oncorhynchus mykiss), and their hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1487–1496, July 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-085.

Maravelias:1996:SAM

[1934] C. D. Maravelias, D. G. Reid, E. J. Simmonds, and J. Haralabous. Spatial analysis and mapping of acoustic survey data in the presence of high local variability: geostatistical application to North Sea herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1497–1505, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
079.

Wanzenbock:1996:CBM

[1935] J. Wanzenböck, M. Zaunreiter, C. M. Wahl, and D. L. G. Noakes. Comparison of behavioural and morphological measures of visual resolution during ontogeny of roach (Rutilus rutilus) and yellow perch (Perca flavescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1506–1512, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
082.

Hudon:1996:CNO

[1936] C. Hudon, R. Morin, J. Bunch, and R. Harland. Carbon and nutrient output from the Great Whale River (Hudson Bay) and a comparison with other rivers around Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (7):1513–1525, July 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-080.

Pomeroy:1996:CSE

[1937] P. P. Pomeroy, N. Green, A. J. Hall, M. Walton, K. Jones, and J. Harwood. Congener-specific exposure of grey seal (Halichoerus grypus) pups to chlorinated biphenyls during lactation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1526–1534, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-087.

Regnell:1996:MMP

[1938] O. Regnell, A. Tunlid, G. Ewald, and O. Sangfors. Methyl mercury production in freshwater microcosms affected by dissolved oxygen levels: role of cobalamin and microbial community composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1535–1545, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-086.

Johnsson:1996:DGH

[1939] J. I. Johnsson, E. Petersson, E. Jönsson, B. T. Björnsson, and T. Järvi. Domestication and growth hormone alter antipredator behaviour and growth patterns in juvenile brown trout, Salmo trutta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1546–1554, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/f96-090. See erratum [2028].

Quinn:1996:IHC

[1940] T. P. Quinn and N. P. Peterson. The influence of habitat complexity and fish size on over-winter survival and growth of individually marked juvenile Coho salmon (Oncorhynchus kisutch) in Big Beef Creek, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1555–1564, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-092.

Winter:1996:ECH

[1941] A. Winter, J. J. H. Ciborowski, and T. B. Reynoldson. Effects of chronic hypoxia and reduced temperature on survival and growth of burrowing mayflies, (Hexagenia limbata) (Ephemeroptera: Ephemeridae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1565–1571, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-093.

Paloheimo:1996:EFM

[1942] J. E. Paloheimo and Y. Chen. Estimating fish mortalities and cohort sizes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1572–1579, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-077.

Wilson:1996:ARS

[1943] S. E. Wilson, B. F. Cumming, and J. P. Smol. Assessing the reliability of salinity inference models from diatom assemblages: an examination of a 219-lake data set from western North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1580–1594, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-094.

Kiffney:1996:MIE

[1944] P. M. Kiffney. Main and interactive effects of invertebrate density, predation, and metals on a Rocky Mountain stream macroinvertebrate community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1595–1601, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-091.

Marmorek:1996:IPA

[1945] D. R. Marmorek, R. M. MacQueen, C. H. R. Wedeles, J. Korman, P. J. Blancher, and D. K. McNicol. Improving pH and alkalinity estimates for regional-scale acidification models: incorporation of dissolved organic carbon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1602–1608, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-081.

Mason:1996:EAP

[1946] D. M. Mason and S. B. Brandt. Effect of alewife predation on survival of larval yellow perch in an embayment of Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1609–1617, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-076.

Mertz:1996:IFR

[1947] G. Mertz and R. A. Myers. Influence of fecundity on recruitment variability of marine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (7):1618–1625, July 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-089.

Summers:1996:RAD

[1948] R. Brent Summers, James H. Thorp, James E. Alexander, Jr., and Ronald D. Fell. Respiratory adjustment of dreissenid mussels (Dreissena polymorpha and Dreissena bugensis) in response to chronic turbidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1626–1631, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-096.

Wang:1996:MDA

[1949] J. Y. Wang, D. E. Gaskin, and B. N. White. Mitochondrial DNA analysis of the harbour porpoise, (Phocoena phocoena), subpopulations in North American waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (7):1632–1645, July 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-095.

Arvola:1996:EPA

[1950] L. Arvola, P. Kankaala, T. Tulonen, and A. Ojala. Effects of phosphorus and allochthonous humic matter enrichment on metabolic processes and community structure of plankton in a boreal lake (Lake Pääjärvi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1646–1662, July 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-083.

Kankaala:1996:CBP

[1951] P. Kankaala, L. Arvola, T. Tulonen, and A. Ojala. Carbon budget for the pelagic food web of the euphotic zone in a boreal lake (Lake Pääjärvi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1663–1674, July 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-084.

Xiao:1996:HDS

[1952] Y. Xiao. How does somatic growth rate affect otolith size in fishes? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1675–1682, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-074.

Gavaris:1996:PSR

[1953] S. Gavaris. Population stewardship rights: decentralized management through explicit accounting of the value of uncaught fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1683–1691, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-088.

Kalff:1996:ORH

[1954] J. Kalff. Obituary: Robert Henry Peters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(7):1692–1694, July 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-208.

Staurnes:1996:STE

[1955] M. Staurnes, L. P. Hansen, K. Fugelli, and Ø. Haraldstad. Short-term exposure to acid water impairs osmoregulation, seawater tolerance, and subsequent marine survival of smolts of Altantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1695–1704, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-099.

Mitchell:1996:ADP

[1956] J. S. Mitchell, R. C. Bailey, and R. W. Knapton. Abundance of (Dreissena polymorpha and Dreissena bugensis) in a warmwater plume: effects of depth and temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (8):1705–1712, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-111.

McGovern:1996:FAS

[1957] J. C. McGovern and J. E. Olney. Factors affecting survival of early life stages and subsequent recruitment of striped bass on the Pamunkey River, Virginia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1713– 1726, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f96-100.

Bernard:1996:PQS

[1958] R. L. Bernard and K. W. Myers. The performance of quantitative scale pattern analysis in the identification of hatchery and wild steelhead (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8): 1727–1735, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-102.

Morgan:1996:PRH

[1959] J. D. Morgan, S. K. Balfry, M. M. Vijayan, and G. K. Iwama. Physiological responses to hyposaline exposure and handling and confinement stress in juvenile dolphin (mahimahi: Coryphaena hippurus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1736–1740, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-113.

Vandermeulen:1996:SVS

[1960] J. H. Vandermeulen and D. Mossman. Sources of variability in seasonal hepatic microsomal oxygenase activity in winter flounder (Pleuronectes americanus) from a coal tar contaminated estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1741–1753, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-026.

Holz:1996:EMS

[1961] J. C. Holz and K. D. Hoagland. Experimental microcosm study of the effects of phosphorus reduction on plankton community structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1754–1764, August 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-104.

Officer:1996:SVC

[1962] R. A. Officer, A. S. Gason, T. I. Walker, and J. G. Clement. Sources of variation in counts of growth increments in vertebrae from gummy shark, (Mustelus antarcticus, and school shark, Galeorhinus galeus): implications for age determination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (8):1765–1777, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-103.

Lonergan:1996:MTI

[1963] S. P. Lonergan and J. B. Rasmussen. A multi-taxonomic indicator of acidification: isolating the effects of pH from other water-chemistry variables. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1778–1787, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
101.

Ogle:1996:BAA

[1964] D. H. Ogle, R. C. Pruitt, G. R. Spangler, and M. J. Cyterski. A Bayesian approach to assigning probabilities to fish ages determined from temporal signatures in growth increments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1788–1794, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-110.

Meeuwig:1996:CPL

[1965] J. J. Meeuwig and R. H. Peters. Circumventing phosphorus in lake management: a comparison of chlorophyll a predictions from land-use and phosphorus-loading models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (8):1795–1806, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-107.

Skulason:1996:GBL

[1966] S. Skúlason, S. S. Snorrason, D. L. G. Noakes, and M. M. Ferguson. Genetic basis of life history variations among sympatric morphs of Arctic char Salvelinus alpinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8): 1807–1813, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-098.

Biro:1996:DRF

[1967] P. A. Biro, M. S. Ridgway, and R. L. McLaughlin. Does the rate of foraging attempts predict ingestion rate for young-of-the-year brook trout (Salvelinus fontinalis) in the field? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1814–1820, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-109.

Moran:1996:PSB

[1968] J. D. W. Moran, J. R. Arthur, and M. D. B. Burt. Parasites of sharpbeaked redfishes (Sebastes fasciatus and Sebates mentella) collected from the Gulf of St. Lawrence, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1821–1826, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-115.

AAtland:1996:ABA

[1969] Å. Åtland and B. T. Barlaup. Avoidance behaviour of Atlantic salmon (Salmo salar L.) fry in waters of low pH and elevated aluminium concentration: laboratory experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1827–1834, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-112.

Barlaup:1996:EMB

[1970] B. T. Barlaup and Å. Åtland. Episodic mortality of brown trout (Salmo trutta L.) caused by sea-salt-induced acidification in western Norway: effects on different life stages within three populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1835–1843, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-108.

Trebitz:1996:EPV

[1971] A. S. Trebitz and N. Nibbelink. Effect of pattern of vegetation removal on growth of bluegill: a simple model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1844–1851, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-105.

Xiao:1996:GME

[1972] Y. Xiao. A general model for estimating tag-specific shedding rates and tag interactions from exact or pooled times at liberty for a double tagging experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8): 1852–1861, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-106.

Malley:1996:CWL

[1973] D. F. Malley. Cadmium whole-lake experiment at the Experimental Lakes Area: an anachoronism? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1862–1870, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-116.

Holoka:1996:AAT

[1974] M. H. Holoka and R. V. Hunt. Automated addition of toxicant to a whole lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1871–1875, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
117.

Lawrence:1996:MYE

[1975] S. G. Lawrence, M. H. Holoka, R. V. Hunt, and R. H. Hesslein. Multi-year experimental additions of cadmium to a lake epilimnion and resulting water column cadmium concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1876–1887, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-118.

Stephenson:1996:SEA

[1976] M. Stephenson, L. Bendell-Young, G. A. Bird, G. J. Brunskill, P. J. Curtis, W. L. Fairchild, M. H. Holoka, R. V. Hunt, S. G. Lawrence, M. F. Motycka, W. J. Schwartz, M. A. Turner, and P. Wilkinson. Sedimentation of experimentally added cadmium and ^{109}Cd in Lake 382, Experimental Lakes Area, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (8):1888–1902, August 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-119.

Findlay:1996:LTE

[1977] D. L. Findlay, S. E. M. Kasian, and E. U. Schindler. Long-term effects of low cadmium concentrations on a natural phytoplankton community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1903–1912, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-120.

Doucett:1996:CCE

[1978] R. R. Doucett, D. R. Barton, K. R. A. Guiguer, G. Power, and R. J. Drimmie. Comment: Critical examination of stable isotope analysis as a means for tracing carbon pathways in stream ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1913–1915, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-114. See [1545].

France:1996:CCL

[1979] R. L. France. Carbon-13 conundrums: limitations and cautions in the use of stable isotope analysis in stream ecotonal research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(8):1916–1919, August 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-121.

Schneider:1996:PSL

[1980] C. P. Schneider, R. W. Owens, R. A. Bergstedt, and R. O’Gorman. Predation by sea lamprey (Petromyzon marinus) on lake trout (Salvelinus namaycush) in southern Lake Ontario, 1982–1992. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1921–1932, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-129.

Galvin:1996:PGA

[1981] P. Galvin, J. Taggart, A. Ferguson, M. O’Farrell, and T. Cross. Population genetics of Atlantic salmon (Salmo salar) in the River Shannon system in Ireland: an appraisal using single locus minisatellite (VNTR) probes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1933–1942, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-130.

Stormer:1996:UNN

[1982] J. Stormer, F. B. Jensen, and J. C. Rankin. Uptake of nitrite, nitrate, and bromide in rainbow trout, (Oncorhynchus mykiss): effects on ionic balance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1943–1950, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-142.

Brodeur:1996:MAP

[1983] R. D. Brodeur and M. T. Wilson. Mesoscale acoustic patterns of juvenile walleye pollock (Theragra chalcogramma) in the western Gulf of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1951–1963, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-127.

Fitzhugh:1996:MCV

[1984] G. R. Fitzhugh, L. B. Crowder, and J. P. Monaghan, Jr. Mechanisms contributing to variable growth in juvenile southern flounder (Paralichthys lethostigma). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1964– 1973, September 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-125.

Kotak:1996:MLC

[1985] B. G. Kotak, R. W. Zurawell, E. E. Prepas, and C. F. B. Holmes. Microcystin-LR concentration in aquatic food web compartments from lakes of varying trophic status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1974–1985, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-135.

Chambers:1996:MSD

[1986] R. C. Chambers and K. G. Waiwood. Maternal and seasonal differences in egg sizes and spawning characteristics of captive Atlantic cod, Gadus morhua. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):1986– 2003, September 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-141.

Berejikian:1996:EHW

[1987] B. A. Berejikian, S. B. Mathews, and T. P. Quinn. Effects of hatchery and wild ancestry and rearing environments on the development of agonistic behavior in steelhead trout (Oncorhynchus mykiss) fry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2004–2014, September 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-133.

Gauldie:1996:ETV

[1988] R. W. Gauldie. Effects of temperature and vaterite replacement on the chemistry of metal ions in the otoliths of Oncorhynchus tshawytscha. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2015–2026, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-123.

Skaala:1996:GIT

[1989] Ø. Skaala, K. E. Jørstad, and R. Borgstrom. Genetic impact on two wild brown trout (Salmo trutta) populations after release of non-indigenous hatchery spawners. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2027– 2035, September 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-126.

Brobbel:1996:PEC

[1990] M. A. Brobbel, M. P. Wilkie, K. Davidson, J. D. Kieffer, A. T. Bielak, and B. L. Tufts. Physiological effects of catch and release angling in Atlantic salmon (Salmo salar) at different stages of freshwater migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2036–2043, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-143.

Cavaletto:1996:SVL

[1991] J. F. Cavaletto, T. F. Nelepa, R. Dermott, W. S. Gardner, M. A. Quigley, and G. A. Lang. Seasonal variation of lipid composition, weight, and length in Diporeia spp. (Amphipoda) from lakes Michigan and Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2044–2051, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-144.

Marks:1996:OEP

[1992] R. E. Marks, F. Juanes, J. A. Hare, and D. O. Conover. Occurrence and effect of the parasitic isopod, Lironeca ovalis (Isopoda: Cyymothoidae), on young-of-the-year bluefish, Pomatomus saltatrix (Pisces: Pomatomidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2052– 2057, September 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f95-155.

Doucett:1996:SIA

[1993] R. R. Doucett, G. Power, D. R. Barton, R. J. Drimmie, and R. A. Cunjak. Stable isotope analysis of nutrient pathways leading to Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2058–2066, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-132.

Lin:1996:ESI

[1994] F. Lin and K. Dabrowski. Effects of sperm irradiation and heat shock on induction of gynogenesis in muskellunge (Esox masquinongy). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2067–2075, September 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-140.

Sulkin:1996:EST

[1995] S. D. Sulkin, E. Mojica, and G. L. McKeen. Elevated summer temperature effects on megalopal and early juvenile development in the Dungeness crab, Cancer magister. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2076–2079, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-137.

Guiney:1996:ATD

[1996] P. D. Guiney, P. M. Cook, J. M. Casselman, J. D. Fitzsimmons, H. A. Simonin, E. W. Zabel, and R. E. Peterson. Assessment of 2,3,7,8-tetrachlorodibenzo-p-dioxin induced sac fry mortality in lake trout (Salvelinus namaycush) from different regions of the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2080–2092, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-139.

Blaikie:1996:EAL

[1997] H. B. Blaikie and S. R. Kerr. Effect of activity level on apparent heat increment in Atlantic cod, Gadus morhua. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2093–2099, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-124.

Dueck:1996:MPS

[1998] L. A. Dueck and R. G. Danzmann. Matriarchal population structure of introduced rainbow trout (Oncorhynchus mykiss) in the Lake Ontario watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2100– 2114, September 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-138.

Horbowy:1996:DBF

[1999] J. Horbowy. The dynamics of Baltic fish stocks on the basis of a multispecies stock-production model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2115–2125, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-128.

Vignaux:1996:AVM

[2000] M. Vignaux. Analysis of vessel movements and strategies using commercial catch and effort data from the New Zealand hoki fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2126–2136, September 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-131.

Lanno:1996:CTW

[2001] R. P. Lanno and D. G. Dixon. Chronic toxicity of waterborne thiocyanate to rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2137–2146, September 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-134.

Nathanailides:1996:MSM

[2002] C. Nathanailides. Metabolic specialization of muscle during development in cold-water and warmwater fish species exposed to different thermal conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(9):2147– 2155, September 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-136.

Reed:1996:ACE

[2003] W. J. Reed and C. M. Simons. Analyzing catch — effort data by means of the Kalman filter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10): 2157–2166, October 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-176.

Rundle:1996:STV

[2004] H. D. Rundle and D. A. Jackson. Spatial and temporal variation in littoral-zone fish communities: a new statistical approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2167–2176, October 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-183.

Adams:1996:DGB

[2005] S. M. Adams, K. D. Ham, M. S. Greeley, R. F. LeHew, D. E. Hinton, and C. F. Saylor. Downstream gradients in bioindicator responses: point source contaminant effects on fish health. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2177–2187, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-191.

Skalski:1996:SET

[2006] J. R. Skalski, G. E. Johnson, C. M. Sullivan, E. Kudera, and M. W. Erho. Statistical evaluation of turbine bypass efficiency at Wells Dam on the Columbia River, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2188–2198, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-179.

Young:1996:IVD

[2007] R. G. Young and A. D. Huryn. Interannual variation in discharge controls ecosystem metabolism along a grassland river continuum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2199–2211, October 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-186.

MacLean:1996:BKT

[2008] R. S. MacLean, U. Borgmann, and D. G. Dixon. Bioaccumulation kinetics and toxicity of lead in (Hyalella azteca) (Crustacea, Amphipoda). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2212–2220, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-193.

Westcott:1996:EFA

[2009] K. Westcott and J. Kalff. Environmental factors affecting methyl mercury accumulation in zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2221–2228, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-178.

Hansen:1996:FEG

[2010] M. M. Hansen and K-L. D. Mensberg. Founder effects and genetic population structure of brown trout(Salmo trutta) in a Danish river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2229–2237, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-188. See

erratum [2183].

Engaas:1996:ESS

[2011] A. Engås, S. Løkkeborg, E. Ona, and A. V. Soldal. Effects of seismic shooting on local abundance and catch rates of cod (Gadus morhua) and haddock (Melanogrammus aeglefinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2238–2249, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-177.

Lathrop:1996:WCL

[2012] R. C. Lathrop, S. R. Carpenter, and L. G. Rudstam. Water clarity in Lake Mendota since 1900: responses to differing levels of nutrients and herbivory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2250–2261, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-187.

Harwood:1996:DAB

[2013] L. A. Harwood, S. Innes, P. Norton, and M. C. S. Kingsley. Distribution and abundance of beluga whales in the Mackenzie estuary, southeast Beaufort Sea, and west Amundsen Gulf during late July 1992. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2262–2273, October 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-180.

Brown:1996:TSM

[2014] B. L. Brown, J. M. Epifanio, P. E. Smouse, and C. J. Kobak. Temporal stability of mtDNA haplotype frequencies in American shad stocks: to pool or not to pool across years? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2274–2283, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-189.

Klerks:1996:EZM

[2015] P. L. Klerks, P. C. Fraleigh, and J. E. Lawniczak. Effects of zebra mussel (Dreissena polymorpha) on seston levels and sediment deposition in western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10): 2284–2291, October 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-190.

OReilly:1996:RAG

[2016] P. T. O’Reilly, L. C. Hamilton, S. K. McConnell, and J. M. Wright. Rapid analysis of genetic variation in Atlantic salmon (Salmo salar) by PCR multiplexing of dinucleotide and tetranucleotide microsatellites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2292–2298, October 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-192.

Schlenk:1996:CHB

[2017] D. Schlenk, E. J. Perkins, G. Hamilton, Y. S. Zhang, and W. Layher. Correlation of hepatic biomarkers with whole animal and population– community metrics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10): 2299–2309, October 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-185.

Taylor:1996:STD

[2018] F. J. R. Taylor and R. Haigh. Spatial and temporal distributions of microplankton during the summers of 1992–1993 in Barkley Sound, British Columbia, with emphasis on harmful species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2310–2322, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-181.

Bernard:1996:ESH

[2019] D. R. Bernard and J. E. Clark. Estimating salmon harvest with codedwire tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2323–2332, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
182.

Nathanailides:1996:CEA

[2020] C. Nathanailides. Are changes in enzyme activities of fish muscle during cold acclimation significant? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2333–2336, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-184.

Brown:1996:ISE

[2021] E. D. Brown, B. L. Norcross, and J. W. Short. Introduction to studies on the effects of the (Exxon Valdez) oil spill on early life history stages of Pacific herring, (Clupea pallasi), in Prince William Sound, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2337–2342, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-211.

McGurk:1996:ELM

[2022] M. D. McGurk and E. D. Brown. Egg–larval mortality of Pacific herring in Prince William Sound, Alaska, after the Exxon Valdez oil spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2343–2354, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-172.

Hose:1996:SEE

[2023] J. E. Hose, M. D. McGurk, G. D. Marty, D. E. Hinton, E. D. Brown, and T. T. Baker. Sublethal effects of the (Exxon Valdez) oil spill on herring embryos and larvae: morphological, cytogenetic, and histopathological assessments, 1989–1991. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10): 2355–2365, October 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-174.

Kocan:1996:PHC

[2024] R. M. Kocan, J. E. Hose, E. D. Brown, and T. T. Baker. Pacific herring (Clupea pallasi) embryo sensitivity to Prudhoe Bay petroleum hydrocarbons: laboratory evaluation and in situ exposure at oiled and unoiled sites in Prince William Sound. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2366–2375, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-173.

Norcross:1996:DAM

[2025] B. L. Norcross, J. E. Hose, M. Frandsen, and E. D. Brown. Distribution, abundance, morphological condition, and cytogenetic abnormalities of larval herring in Prince William Sound, Alaska, following the (Exxon Valdez) oil spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10): 2376–2387, October 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-212.

Kocan:1996:RSH

[2026] R. M. Kocan, G. D. Marty, M. S. Okihiro, E. D. Brown, and T. T. Baker. Reproductive success and histopathology of individual Prince William Sound Pacific herring 3 years after the(Exxon Valdez) oil spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2388–2393, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-175.

Beland:1996:ERB

[2027] K. F. Beland. Erratum: The relation between redd counts and Atlantic salmon (Salmo salar) parr populations in the Dennys River, Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2394, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-294. See [1834].

Johnsson:1996:EDG

[2028] J. I. Johnsson, E. Petersson, E. Jönsson, B. T. Björnsson, and T. Järvi. Erratum: Domestication and growth hormone alter antipredator behavior and growth patterns in juvenile brown trout, (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(10):2395, October 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-295. See [1939].

Bentzen:1996:RFW

[2029] E. Bentzen, D. R. S. Lean, W. D. Taylor, and D. Mackay. Role of food web structure on lipid bioaccumulation of organic contaminants by lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2397–2407, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-196.

Leis:1996:FGH

[2030] A. L. Leis and M. G. Fox. Feeding, growth, and habitat associations of young-of-year walleye (Stizostedion vitreum) in a river affected by a mine tailings spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2408–2417, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-203.

Domanico:1996:SVR

[2031] M. J. Domanico, R. B. Phillips, and J. F. Schweigert. Sequence variation in the ribosomal DNA of Pacific (Clupea pallasi) and Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2418–2423, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-194.

Wood:1996:EDC

[2032] A. W. Wood, B. D. Johnston, A. P. Farrell, and C. J. Kennedy. Effects of didecyldimethylammonium chloride (DDAC) on the swimming performance, gill morphology, disease resistance, and biochemistry of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2424–2432, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-201.

Pinel-Alloul:1996:ALW

[2033] B. Pinel-Alloul, N. Bourbonnais, and P. Dutilleul. Among-lake and within-lake variations of autotrophic pico- and nanoplankton biomass in six Quebec lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2433–2445, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-197.

Adams:1996:WSF

[2034] C. E. Adams and F. A. Huntingford. What is a successful fish? Determinants of competitive success in Arctic char (Salvelinus alpinus) in different social contexts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2446–2450, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-195.

Savenkoff:1996:DOC

[2035] C. Savenkoff, A. F. Vézina, T. T. Packard, N. Silverberg, J.-C. Therriault, W. Chen, C. Bérubé, A. Mucci, B. Klein, F. Mesplé, J.-E. Tremblay, L. Legendre, J. Wesson, and R. G. Ingram. Distributions of oxygen, carbon, and respiratory activity in the deep layer of the Gulf of St.

Lawrence and their implications for the carbon cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2451–2465, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-198.

Roy:1996:NMO

[2036] R. Roy, R. Knowles, and M. N. Charlton. Nitrification and methane oxidation at the sediment surface in Hamilton Harbour (Lake Ontario). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2466–2472, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-199.

Hughes:1996:NTD

[2037] N. F. Hughes and L. H. Kelly. New techniques for 3-D video tracking of fish swimming movements in still or flowing water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2473–2483, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-200.

Hughes:1996:HME

[2038] N. F. Hughes and L. H. Kelly. A hydrodynamic model for estimating the energetic cost of swimming maneuvers from a description of their geometry and dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2484–2493, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-204.

Barbeau:1996:DJS

[2039] M. A. Barbeau, B. G. Hatcher, R. E. Scheibling, A. W. Hennigar, L. H. Taylor, and A. C. Risk. Dynamics of juvenile sea scallop (Placopecten magellanicus) and their predators in bottom seeding trials in Lunenburg Bay, Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2494–2512, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-202.

Nero:1996:LFA

[2040] R. W. Nero and M. E. Huster. Low frequency acoustic imaging of Pacific salmon on the high seas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11): 2513–2523, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-206.

Kvenseth:1996:EDA

[2041] A. M. Kvenseth, K. Pittman, and J. V. Helvik. Eye development in Atlantic halibut (Hippoglossus hippoglossus): differentiation and development of the retina from early yolk sac stages through metamorphosis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2524–2532, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-205.

Wang:1996:EMI

[2042] Y.-G. Wang and N. R. Loneragan. An extravariation model for improving confidence intervals of population size estimates from removal data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2533–2539, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-207.

Kennicutt:1996:GMO

[2043] M. C. Kennicutt II, R. H. Green, P. Montagna, and P. F. Roscigno. Gulf of Mexico Offshore Operations Monitoring Experiment (GOOMEX) Phase I: Sublethal responses to contaminant exposure-introduction and overview. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2540–2553, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-213.

Kennicutt:1996:GPS

[2044] M. C. Kennicutt II, P. N. Boothe, T. L. Wade, S. T. Sweet, R. Rezak, F. J. Kelly, J. M. Brooks, B. J. Presley, and D. A. Wiesenburg. Geochemical patterns in sediments near offshore production platforms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2554–2566, November 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-214.

Montagna:1996:BIL

[2045] P. Montagna and D. E. Harper, Jr. Benthic infaunal long-term response to offshore production platforms in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2567–2588, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-215.

Ellis:1996:EGP

[2046] M. S. Ellis, E. A. Wilson-Ormond, and E. N. Powell. Effects of gasproducing platforms on continental shelf macroepifauna in the northwestern Gulf of Mexico: abundance and size structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2589–2605, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-216.

McDonald:1996:SDR

[2047] S. J. McDonald, K. L. Willett, J. Thomsen, K. B. Beatty, K. Connor, T. R. Narasimhan, C. M. Erickson, and S. H. Safe. Sublethal detoxification responses to contaminant exposure associated with offshore production platforms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2606– 2617, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-217.

Carr:1996:SPT

[2048] R. S. Carr, D. C. Chapman, B. J. Presley, J. M. Biedenbach, L. Robertson, P. N. Boothe, R. Kilada, T. Wade, and P. Montagna. Sediment porewater toxicity assessment studies in the vicinity of offshore oil and gas production platforms in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2618–2628, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-218.

Green:1996:IMS

[2049] R. H. Green and P. Montagna. Implications for monitoring: study designs and interpretation of results. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2629–2636, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-219.

Peterson:1996:ECE

[2050] C. H. Peterson, M. C. Kennicutt II, R. H. Green, P. Montagna, D. E. Harper, Jr., E. N. Powell, and P. F. Roscigno. Ecological consequences of environmental perturbations associated with offshore hydrocarbon production: a perspective on long-term exposures in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2637–2654, November 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-220.

Mohn:1996:CEO

[2051] R. K. Mohn. Comment on elementary operations of sequential population analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(11):2655– 2656, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f96-209. See reply [2052].

Evans:1996:RM

[2052] G. T. Evans. Reply to Mohn. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53 (11):2657, November 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-210. See [2051].

Stottlemyer:1996:PSS

[2053] J. R. Stottlemyer and D. Toczydlowski. Precipitation, snowpack, streamwater ion chemistry, and flux in a northern Michigan watershed, 1982– 1991. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2659–2672, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-236.

Handeland:1996:OSA

[2054] S. O. Handeland, T. Järvi, A. Fernö, and S. O. Stefansson. Osmotic stress, antipredatory behaviour, and mortality of Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2673– 2680, December 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-227.

Fox:1996:CCF

[2055] D. S. Fox and R. M. Starr. Comparison of commercial fishery and research catch data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2681– 2694, December 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-230.

Ruzzante:1996:STV

[2056] D. E. Ruzzante, C. T. Taggart, and D. Cook. Spatial and temporal variation in the genetic composition of a larval cod (Gadus morhua) aggregation: cohort contribution and genetic stability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2695–2705, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-235.

Bentzen:1996:MPP

[2057] P. Bentzen, C. T. Taggart, D. E. Ruzzante, and D. Cook. Microsatellite polymorphism and the population structure of Atlantic cod (Gadus morhua) in the northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2706–2721, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-238.

Mohn:1996:GSP

[2058] R. Mohn and W. D. Bowen. Grey seal predation on the eastern Scotian Shelf: modelling the impact on Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2722–2738, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-239.

Webb:1996:VTS

[2059] J. H. Webb and H. A. McLay. Variation in the time of spawning of Atlantic salmon (Salmo salar) and its relationship to temperature in the Aberdeenshire Dee, Scotland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2739–2744, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-240.

Russ:1996:TTS

[2060] G. R. Russ, D. C. Lou, and B. P. Ferreira. Temporal tracking of a strong cohort in the population of a coral reef fish, the coral trout, (Plectropomus leopardus) Serranidae: Epinephelinae, in the central Great Barrier Reef, Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2745– 2751, December 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-233.

Dockray:1996:EES

[2061] J. J. Dockray, S. D. Reid, and C. M. Wood. Effects of elevated summer temperatures and reduced pH on metabolism and growth of juvenile rainbow trout (Oncorhynchus mykiss) on unlimited ration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2752–2763, December 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-232.

Wilson:1996:POL

[2062] C. C. Wilson and P. D. N. Hebert. Phylogeographic origins of lake trout (Salvelinus namaycush) in eastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2764–2775, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-223.

Jones:1996:CGB

[2063] M. W. Jones, D. Clay, and R. G. Danzmann. Conservation genetics of brook trout (Salvelinus fontinalis): population structuring in Fundy National Park, New Brunswick, and eastern Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2776–2791, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-237.

Johnson:1996:LTC

[2064] T. B. Johnson and J. F. Kitchell. Long-term changes in zooplanktivorous fish community composition: implications for food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2792–2803, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-244.

Measures:1996:ETS

[2065] L. N. Measures. Effect of temperature and salinity on development and survival of eggs and free-living larvae of sealworm (Pseudoterranova decipiens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2804–2807, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-241.

Birkeland:1996:CPR

[2066] K. Birkeland. Consequences of premature return by sea trout (Salmo trutta) infested with the salmon louse (Lepeophtheirus salmonis Krøyer): migration, growth, and mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2808–2813, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-231.

Gu:1996:SIE

[2067] B. Gu, D. M. Schell, X. Huang, and F. Yie. Stable isotope evidence for dietary overlap between two planktivorous fishes in aquaculture ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2814–2818, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-242.

Ramcharan:1996:MEP

[2068] C. W. Ramcharan, R. L. France, and D. J. McQueen. Multiple effects of planktivorous fish on algae through a pelagic trophic cascade. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2819–2828, December 1996. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-224.

Marcogliese:1996:DAS

[2069] D. J. Marcogliese, F. Boily, and M. O. Hammill. Distribution and abundance of stomach nematodes (Anisakidae) among grey seals (Halichoerus grypus) and harp seals (Phoca groenlandica) in the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2829–2836, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-243.

Perez:1996:EGD

[2070] J. A. A. Perez, R. K. O’Dor, P. Beck, and E. G. Dawe. Evaluation of gladius dorsal surface structure for age and growth studies of the shortfinned squid, (Illex illecebrosus) (Teuthoidea: Ommastrephidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2837–2846, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-228.

Nagelkerke:1996:EFZ

[2071] L. A. J. Nagelkerke and F. A. Sibbing. Efficiency of feeding on zebra mussel (Dreissena polymorpha) by common bream (Abramis brama), white bream (Blicca bjoerkna), and roach (Rutilus rutilus): the effects of morphology and behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12): 2847–2861, December 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-229.

Headon:1996:DRM

[2072] C. M. Headon, R. J. Hall, and G. Mierle. Dynamics of radiolabelled methylmercury in crayfish (Orconectes virilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2862–2869, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-222.

Patterson:1996:MID

[2073] K. R. Patterson. Modelling the impact of disease-induced mortality in an exploited population: the outbreak of the fungal parasite (Ichthyophonus hoferi) in the North Sea herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2870–2887, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-234.

Johnson:1996:DIS

[2074] S. C. Johnson, R. B. Blaylock, J. Elphick, and K. D. Hyatt. Disease induced by the sea louse (Lepeophteirus salmonis) (Copepoda: Caligidae) in wild sockeye salmon (Oncorhynchus nerka) stocks of Alberni Inlet, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):

2888–2897, December 1996. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-226.

Beamesderfer:1996:EBB

[2075] R. C. P. Beamesderfer, D. L. Ward, and A. A. Nigro. Evaluation of the biological basis for a predator control program on northern squawfish (Ptychocheilus oregonensis) in the Columbia and Snake Rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2898–2908, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-225.

Sinclair:1996:CSI

[2076] A. F. Sinclair and D. P. Swain. Comment: Spatial implications of a temperature-based growth model for Atlantic cod (Gadus morhua) off the eastern coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2909–2911, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-293. See [1748] and reply [2051].

Campana:1996:RSI

[2077] S. E. Campana, R. K. Mohn, S. J. Smith, and G. A. Chouinard. Reply: Spatial implications of a temperature-based growth model for Atlantic cod (Gadus morhua) off the eastern coast of Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 53(12):2912–2914, December 1996. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-271. See [2076].

Myers:1997:DEG

[2078] R. A. Myers and J. M. Hoenig. Direct estimates of gear selectivity from multiple tagging experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):1–9, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-262.

Klump:1997:SPC

[2079] J. V. Klump, D. N. Edgington, P. E. Sager, and D. M. Robertson. Sedimentary phosphorus cycling and a phosphorus mass balance for the Green Bay (Lake Michigan) ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):10–26, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-247.

Harvey:1997:HDI

[2080] B. C. Harvey and R. J. Nakamoto. Habitat-dependent interactions between two size-classes of juvenile steelhead in a small stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):27–31, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-249.

Helminen:1997:RLP

[2081] H. Helminen and J. Sarvala. Responses of Lake Pyhäjärvi (southwestern Finland) to variable recruitment of the major planktivorous fish, vendace (Coregonus albula). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (1):32–40, January 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-248.

Prepas:1997:AEC

[2082] E. E. Prepas, B. G. Kotak, L. M. Campbell, J. C. Evans, S. E. Hrudey, and C. F. B. Holmes. Accumulation and elimination of cyanobacterial hepatotoxins by the freshwater clam Anodonta grandis simpsoniana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):41–46, January 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-261.

Lu:1997:MDD

[2083] G. Lu, S. Li, and L. Bernatchez. Mitochondrial DNA diversity, population structure, and conservation genetics of four native carps within the Yangtze River, China. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (1):47–58, January 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-266.

Spencer:1997:OHF

[2084] P. D. Spencer. Optimal harvesting of fish populations with nonlinear rates of predation and autocorrelated environmental variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):59–74, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-260.

Hendry:1997:VAL

[2085] A. P. Hendry and T. P. Quinn. Variation in adult life history and morphology among Lake Washington sockeye salmon (Oncorhynchus nerka) populations in relation to habitat features and ancestral affinities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):75–84, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-257.

Persson:1997:CPE

[2086] L. Persson. Competition, predation and environmental factors as structuring forces in freshwater fish communities: Sumari (1971) revisited. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):85–88, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-250.

Clarke:1997:ITR

[2087] K. D. Clarke, R. Knoechel, and P. M. Ryan. Influence of trophic role and life-cycle duration on timing and magnitude of benthic macroinvertebrate response to whole-lake enrichment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):89–95, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-267.

Fukushima:1997:DSL

[2088] M. Fukushima and W. W. Smoker. Determinants of stream life, spawning efficiency, and spawning habitat in pink salmon in the Auke Lake system, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):96–104, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-258.

Tollit:1997:SSD

[2089] D. J. Tollit, M. J. Steward, P. M. Thompson, G. J. Pierce, M. B. Santos, and S. Hughes. Species and size differences in the digestion of otoliths and beaks: implications for estimates of pinniped diet composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):105–119, January 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-264.

Perry:1997:ICP

[2090] W. L. Perry, D. M. Lodge, and G. A. Lamberti. Impact of crayfish predation on exotic zebra mussels and native invertebrates in a lake-outlet stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):120–125, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-255.

Williams:1997:GRF

[2091] H. A. Williams and M. K. Lowe. Growth rates of four Hawaiian deep slope fishes: a comparison of methods for estimating age and growth from otolith microincrement widths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):126–136, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-245.

Dawe:1997:PSG

[2092] E. G. Dawe and P. C. Beck. Population structure, growth, and sexual maturation of short-finned squid (Illex illecebrosus) at Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):137–146, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-263.

Gagnon:1997:BSB

[2093] C. Gagnon and N. S. Fisher. The bioavailability of sediment-bound Cd, Co, and Ag to the mussel Mytilus edulis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):147–156, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-256.

Doblander:1997:ONN

[2094] C. Doblander and R. Lackner. Oxidation of nitrite to nitrate in isolated erythrocytes: a possible mechanism for adaptation to environmental nitrite. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):157–161, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-254.

Anras:1997:DAR

[2095] M-L. Bégout Anras, J.-P. Lagardére, and J.-Y. Lafaye. Diel activity rhythm of seabass tracked in a natural environment: group effects on swimming patterns and amplitudes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):162–168, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-253.

Stefansson:1997:SEM

[2096] G. Stefansson and O. K. Palsson. Statistical evaluation and modelling of the stomach contents of Icelandic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):169–181, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-246.

Zydlewski:1997:OST

[2097] J. Zydlewski and S. D. McCormick. The ontogeny of salinity tolerance in the American shad, Alosa sapidissima. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):182–189, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-251.

Hamel:1997:CDM

[2098] P. Hamel, P. Magnan, P. East, M. Lapointe, and P. Laurendeau. Comparison of different models to predict the in situ embryonic developmental rate of fish, with special reference to white sucker (Catostomus commersoni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):190–197, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-252.

Ouellet:1997:SAC

[2099] P. Ouellet, Y. Lambert, and M. Castonguay. Spawning of Atlantic cod (Gadus morhua) in the northern Gulf of St. Lawrence: a study of adult and egg distributions and characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):198–210, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-259.

Ouellet:1997:CVD

[2100] P. Ouellet. Characteristics and vertical distribution of Atlantic cod (Gadus morhua) eggs in the northern Gulf of St. Lawrence, and the possible effect of cold water temperature on recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):211–223, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-297.

Zhong:1997:FCP

[2101] Y. Zhong and G. Power. Fisheries in China: progress, problems, and prospects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(1):224–238, January 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f96-
265.

Sainte-Marie:1997:LBA

[2102] B. Sainte-Marie, J.-M. Sevigny, and Y. Gauthier. Laboratory behavior of adolescent and adult males of the snow crab (Chionoecetes opilio) (Brachyura: Majidae) mated noncompetitively and competitively with primiparous females. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2): 239–248, February 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-287.

Horgan:1997:CRF

[2103] M. J. Horgan and E. L. Mills. Clearance rates and filtering activity of zebra mussel (Dreissena polymorpha): implications for freshwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):249–255, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-276.

Leino:1997:RCR

[2104] R. L. Leino and J. H. McCormick. Reproductive characteristics of the ruffe, Gymnocephalus cernuus, in the St. Louis River estuary on western Lake Superior: a histological examination of the ovaries over one annual cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):256–263, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-291.

Witt:1997:EIA

[2105] J. D. S. Witt, P. D. N. Hebert, and W. B. Morton. Echinogammarus ischnus: another crustacean invader in the Laurentian Great Lakes basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):264–268, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-292.

Deegan:1997:EFD

[2106] L. A. Deegan, B. J. Peterson, H. Golden, C. C. McIvor, and M. C. Miller. Effects of fish density and river fertilization on algal standing stocks, invertebrate communities, and fish production in an Arctic river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):269–283, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-268.

McAllister:1997:BSA

[2107] M. K. McAllister and J. N. Ianelli. Bayesian stock assessment using catch-age data and the sampling–importance resampling algorithm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):284–300, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-285.

McAllister:1997:BAC

[2108] M. K. McAllister and E. K. Pikitch. A Bayesian approach to choosing a design for surveying fishery resources: application to the eastern Bering Sea trawl survey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2): 301–311, February 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-286.

Bergstedt:1997:HMF

[2109] L. C. Bergstedt and E. P. Bergersen. Health and movements of fish in response to sediment sluicing in the Wind River, Wyoming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):312–319, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-269.

Waiser:1997:IHF

[2110] M. J. Waiser and R. D. Robarts. Impacts of a herbicide and fertilizers on the microbial community of a saline prairie lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):320–329, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-280.

Pienitz:1997:PCLa

[2111] R. Pienitz, J. P. Smol, and D. R. S. Lean. Physical and chemical limnology of 59 lakes located between the southern Yukon and the Tuktoyaktuk Peninsula, Northwest Territories (Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):330–346, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-274.

Pienitz:1997:PCLb

[2112] Reinhard Pienitz, John P. Smol, and David R. S. Lean. Physical and chemical limnology of 24 lakes located between Yellowknife and Contwoyto Lake, Northwest Territories (Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):347–358, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-275.

Shuter:1997:FAP

[2113] Brian J. Shuter and K. K. Ing. Factors affecting the production of zooplankton in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2): 359–377, February 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f96-270. See errata [2388, 2609].

Wehr:1997:END

[2114] J. D. Wehr and J. H. Thorp. Effects of navigation dams, tributaries, and littoral zones on phytoplankton communities in the Ohio River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):378–395, February 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-283.

Mair:1997:GMS

[2115] G. C. Mair, J. S. Abucay, T. A. Abella, J. A. Beardmore, and D. O. F. Skibinski. Genetic manipulation of sex ratio for the large-scale production of all-male tilapia Oreochromis niloticus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):396–404, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-282.

Letcher:1997:SSD

[2116] B. H. Letcher, J. A. Rice, L. B. Crowder, and F. P. Binkowski. Sizeand species-dependent variability in consumption and growth rates of larvae and juveniles of three freshwater fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):405–414, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-273.

Au:1997:DMP

[2117] D. W. Au and S. E. Smith. A demographic method with population density compensation for estimating productivity and yield per recruit of the leopard shark (Triakis semifasciata). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):415–420, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-288.

Baxter:1997:ENH

[2118] J. S. Baxter, E. B. Taylor, R. H. Devlin, J. Hagen, and J. D. McPhail. Evidence for natural hybridization between Dolly Varden (Salvelinus malma) and bull trout (Salvelinus confluentus) in a northcentral British Columbia watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2): 421–429, February 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-289.

Johnsson:1997:DPG

[2119] J. I. Johnsson, J. Blackburn, W. C. Clarke, and R. E. Withler. Does presmolt growth rate in steelhead trout (Oncorhynchus mykiss) and Coho salmon (Oncorhynchus kisutch) predict growth rate in seawater? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):430–433, February 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-277.

Park:1997:PLR

[2120] J. C. Park, M. Aizaki, T. Fukushima, and A. Otsuki. Production of labile and refractory dissolved organic carbon by zooplankton excretion: an experimental study using large outdoor continuous flow-through ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):434–443, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-278.

Silverstein:1997:EMA

[2121] Jeffrey T. Silverstein, H. Shimma, and H. Ogata. Early maturity in amago salmon (Oncorhynchus masu ishikawai): an association with energy storage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):444–451, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-284.

Herrnkind:1997:TRL

[2122] W. F. Herrnkind and M. J. Butler I. V. A test of recruitment limitation and the potential for artificial enhancement of spiny lobster (Panulirus argus) populations in Florida. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):452–463, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-281.

Stow:1997:BOE

[2123] C. A. Stow, S. R. Carpenter, and R. C. Lapthrop. A Bayesian observation error model to predict cyanobacterial biovolume from spring total phosphorus in Lake Mendota, Wisconsin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):464–473, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-279.

Grafton:1997:FIS

[2124] R. Q. Grafton and H. W. Nelson. Fishers’ individual salmon harvesting rights: an option for Canada’s Pacific fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):474–482, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-290.

Gilbertson:1997:CKS

[2125] M. Gilbertson. Are causes knowable? Some consequences of successional versus toxicological interpretations of the Great Lakes Water Quality Agreement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):483–495, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-296.

Sainte-Marie:1997:CIL

[2126] B. Sainte-Marie. Comment — an improved link between industry, management and science: review of case history of the southwestern Gulf of St. Lawrence snow crab fisher. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):496–500, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-001. See reply [2127].

Moriyasu:1997:RIL

[2127] M. Moriyasu, J. S. Loch, and J. B. Jones. Reply — an improved link between industry, management and science: review of case history of the southwestern Gulf of St. Lawrence snow crab fisher. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(2):501–502, February 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-002. See [2126].

Wilkie:1997:ITU

[2128] M. P. Wilkie, M. A. Brobbel, K. G. Davidson, L. Forsyth, and B. L. Tufts. Influences of temperature upon the postexercise physiology of Atlantic salmon(Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (3):503–511, March 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-305.

Baker:1997:APE

[2129] S. M. Baker and D. J. Hornbach. Acute physiological effects of zebra mussel (Dreissena polymorpha) infestation on two unionid mussels, Actiononaias ligamentina and Amblema plicata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):512–519, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-319.

Maki-Petays:1997:SCH

[2130] A. Maki-Petäys, T. Muotka, A. Huusko, P. Tikkanen, and P. Kreivi. Seasonal changes in habitat use and preference by juvenile brown trout, Salmo trutta, in a northern boreal river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):520–530, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-311.

Toole:1997:STD

[2131] C. L. Toole, D. F. Markle, and C. J. Donohoe. Settlement timing, distribution, and abundance of Dover sole (Microstomus pacificus) on an outer continental shelf nursery area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):531–542, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-304.

Beamish:1997:PFR

[2132] R. J. Beamish, C-E. M. Neville, and A. J. Cass. Production of Fraser River sockeye salmon (Oncorhynchus nerka) in relation to decadal-scale changes in the climate and the ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):543–554, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-310.

Lafontaine:1997:IGS

[2133] P. Lafontaine and J. J. Dodson. Intraspecific genetic structure of white sucker (Catostomus commersoni) in northeastern North America as revealed by mitochondrial DNA polymorphism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):555–565, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-318.

Lauff:1997:ETR

[2134] R. F. Lauff and C. M. Wood. Effects of training on respiratory gas exchange, nitrogenous waste excretion, and fuel usage during aerobic swimming in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):566–571, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-315.

Valtonen:1997:EPF

[2135] E. T. Valtonen, J. C. Holmes, and M. Koskivaara. Eutrophication, pollution and fragmentation: effects on parasite communities in roach (Rutilus rutilus) and perch (Perca fluviatilis) in four lakes in central Finland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):572–585, March 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-306.

Beisner:1997:ITF

[2136] B. E. Beisner, E. McCauley, and F. J. Wrona. The influence of temperature and food chain length on plankton predator–prey dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):586–595, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-312.

Morin:1997:TFO

[2137] P.-P. Morin, T. J. Hara, and J. G. Eales. Thyroid function and olfactory responses to L-alanine during induced smoltification in Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (3):596–602, March 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-309.

Baldigo:1997:ESA

[2138] B. P. Baldigo and P. S. Murdoch. Effect of stream acidification and inorganic aluminum on mortality of brook trout (Salvelinus fontinalis) in the Catskill Mountains, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):603–615, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-314.

Smith:1997:BCL

[2139] S. J. Smith. Bootstrap confidence limits for groundfish trawl survey estimates of mean abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (3):616–630, March 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-303.

Fukuwaka:1997:SAE

[2140] M. Fukuwaka and M. Kaeriyama. Scale analyses to estimate somatic growth in sockeye salmon, Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):631–636, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-307.

Fischer:1997:IEL

[2141] J. M. Fischer and T. M. Frost. Indirect effects of lake acidification on Chaoborus population dynamics: the role of food limitation and predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):637–646, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-313.

Bushmann:1997:SSC

[2142] P. J. Bushmann and J. Atema. Shelter sharing and chemical courtship signals in the lobster, Homarus americanus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):647–654, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-302.

Hankin:1997:DIF

[2143] D. G. Hankin, T. H. Butler, P. W. Wild, and Q.-L. Xue. Does intense fishing on males impair mating success of female Dungeness crabs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):655–669, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-308.

Pepin:1997:DDA

[2144] P. Pepin and J. A. Helbig. Distribution and drift of Atlantic cod (Gadus morhua) eggs and larvae on the northeast Newfoundland shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):670–685, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-317.

Robarts:1997:IPG

[2145] R. D. Robarts. Introduction to papers from the GAP VI workshop: effects of physical forcing on primary production processes in inland and marine environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3): 686, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-324.

Beardall:1997:EUR

[2146] J. Beardall, T. Berman, S. Markager, R. Martinez, and V. Montecino. The effects of ultraviolet radiation on respiration and photosynthesis in two species of microalgae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (3):687–696, March 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-320.

Maske:1997:SUR

[2147] H. Maske and M. Latasa. Solar ultraviolet radiation dependent decrease of particle light absorption and pigments in lake phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):697–704, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-321.

Ferreyra:1997:PRN

[2148] G. A. Ferreyra, S. Demers, P. A. del Giorgio, and J.-P. Chanut. Physiological responses of natural plankton communities to ultraviolet-B radiation in Redberry Lake (Saskatchewan, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):705–714, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-322.

Bott:1997:ETM

[2149] T. L. Bott, J. T. Brock, A. Baatrup-Pedersen, P. A. Chambers, W. K. Dodds, K. T. Himbeault, J. R. Lawrence, D. Planas, E. Snyder, and G. M. Wolfaardt. An evaluation of techniques for measuring periphyton metabolism in chambers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (3):715–725, March 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-323.

Cortes:1997:CRM

[2150] E. Cortés. A critical review of methods of studying fish feeding based on analysis of stomach contents: application to elasmobranch fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):726–738, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-316.

Molot:1997:CPH

[2151] L. A. Molot. Comment — the prediction of hypolimnetic oxygen profiles: a plea for a deductive approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):739, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-300. See [1877] and reply [2152].

Livingstone:1997:RPH

[2152] D. M. Livingstone and D. M. Imboden. Reply — the prediction of hypolimnetic oxygen profiles: a plea for a deductive approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):740–741, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-301. See [1877, 2151].

Wang:1997:CEB

[2153] Y.-G. Wang. Comment: Efficiency and bias of estimators and sampling designs for determining length–weight relationships of fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):742–743, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-298. See [1488] and re-

ply [2154].

Hayes:1997:REB

[2154] D. B. Hayes and J. K. T. Brodziack. Reply: Efficiency and bias of estimators and sampling designs for determining length–weight relationships of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(3):744–745, March 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-299. See

[1488, 2153].

Blanchfield:1997:RTU

[2155] P. J. Blanchfield and M. S. Ridgway. Reproductive timing and use of redd sites by lake-spawning brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):747–756, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-344.

Mesa:1997:PAA

[2156] M. G. Mesa and J. J. Warren. Predator avoidance ability of juvenile Chinook salmon (Oncorhynchus tshawytscha) subjected to sublethal exposures of gas-supersaturated water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):757–764, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-326.

Pepin:1997:VCE

[2157] P. Pepin and T. H. Shears. Variability and capture efficiency of bongo and Tucker trawl samplers in the collection of ichthyoplankton and other macrozooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):765– 773, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f96-347.

Nilo:1997:CHD

[2158] P. Nilo, P. Dumont, and R. Fortin. Climatic and hydrological determinants of year-class strength of St. Lawrence River lake sturgeon (Acipenser fulvescens). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (4):774–780, April 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-330.

Carscadden:1997:EFL

[2159] J. Carscadden, B. S. Nakashima, and K. T. Frank. Effects of fish length and temperature on the timing of peak spawning in capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):781–787, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-331.

Hwang:1997:BPP

[2160] S.-J. Hwang and R. T. Heath. Bacterial productivity and protistan bacterivory in coastal and offshore communities of Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):788–799, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-345.

Rajwani:1997:AMT

[2161] K. N. Rajwani and C. J. Schwarz. Adjusting for missed tags in salmon escapement surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (4):800–808, April 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-337.

Zaranko:1997:AEM

[2162] D. T. Zaranko, D. G. Farara, and F. G. Thompson. Another exotic mollusc in the Laurentian Great Lakes: the New Zealand native Potamopyrgus antipodarum (Gray 1843) (Gastropoda, Hydrobiidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):809–814, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-343.

Jackson:1997:RHC

[2163] C. J. Jackson, D. J. Marcogliese, and M. D. B. Burt. Role of hyperbenthic crustaceans in the transmission of marine helminth parasites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):815–820, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-329.

Plourde:1997:CSP

[2164] Y. Plourde, M. Lucotte, and P. Pichet. Contribution of suspended particulate matter and zooplankton to MeHg contamination of the food chain in midnorthern Quebec (Canada) reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):821–831, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-340.

Tremblay:1997:ATM

[2165] A. Tremblay and M. Lucotte. Accumulation of total mercury and methyl mercury in insect larvae of hydroelectric reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):832–841, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-339.

Measures:1997:HAS

[2166] L. N. Measures, J.-F. Gosselin, and E. Bergeron. Heartworm, Acanthocheilonema spirocauda (Leidy, 1858), infections in Canadian phocid seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):842–846, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-342.

Walter:1997:ROM

[2167] E. E. Walter, J. P. Scandol, and M. C. Healey. A reappraisal of the ocean migration patterns of Fraser River sockeye salmon (Oncorhynchus nerka) by individual-based modelling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):847–858, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-336.

Briggs:1997:SAM

[2168] C. T. Briggs and J. R. Post. In situ activity metabolism of rainbow trout (Oncorhynchus mykiss): estimates obtained from telemetry of axial muscle electromyograms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (4):859–866, April 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-338.

Post:1997:EDD

[2169] J. R. Post, M. R. S. Johannes, and D. J. McQueen. Evidence of densitydependent cohort splitting in age-0 yellow perch, (Perca flavescens): potential behavioural mechanisms and population-level consequences.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):867–875, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-335.

Magnusson:1997:MEM

[2170] K. G. Magnússon and T. Aspelund. A model for estimating meal frequency and meal size from stomach data with an application to Atlantic cod (Gadus morhua) feeding on capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):876–889, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-341.

Gould:1997:CEM

[2171] W. R. Gould and K. H. Pollock. Catch-effort maximum likelihood estimation of important population parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):890–897, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-327.

Gould:1997:EME

[2172] W. R. Gould, L. A. Stefanski, and K. H. Pollock. Effects of measurement error on catch—effort estimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):898–906, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-328.

Douglas:1997:ADA

[2173] D. J. Douglas, E. R. Kenchington, C. J. Bird, R. Pocklington, B. Bradford, and W. Silvert. Accumulation of domoic acid by the sea scallop (Placopecten magellanicus) fed cultured cells of toxic Pseudo-nitzschia multiseries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):907–913, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f96-333.

Lunn:1997:DAR

[2174] N. J. Lunn, I. Stirling, and S. N. Nowicki. Distribution and abundance of ringed (Phoca hispida) and bearded seals (Erignathus barbatus) in western Hudson Bay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (4):914–921, April 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-346.

Dermott:1997:CDB

[2175] Ronald Dermott and Dave Kerec. Changes to the deepwater benthos of eastern Lake Erie since the invasion of Dreissena: 1979–1993. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):922–930, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f96-332.

Hilderbrand:1997:ELW

[2176] R. H. Hilderbrand, A. D. Lemly, C. A. Dolloff, and K. L. Harpster. Effects of large woody debris placement on stream channels and benthic macroinvertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (4):931–939, April 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-334.

Carney:1997:MDR

[2177] B. L. Carney, A. K. Gray, and A. J. Gharrett. Mitochondrial DNA restriction site variation within and among five populations of Alaskan Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):940–949, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f96-325.

Ricker:1997:CAA

[2178] W. E. Ricker. Cycles of abundance among Fraser River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (4):950–968, April 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-047.

Gilbert:1997:TNR

[2179] D. J. Gilbert. Towards a new recruitment paradigm for fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):969–977, April 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f96-272. See com-

ment [2181].

Myers:1997:CRP

[2180] R. A. Myers. Comment and reanalysis: paradigms for recruitment studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):978–981, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-048. See [2182].

Francis:1997:CHS

[2181] R. I. C. C. Francis. Comment: How should fisheries scientists and managers react to uncertainty about stock-recruit relationships? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):982–983, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-049. See [2179].

Hilborn:1997:CRP

[2182] R. Hilborn. Comment: Recruitment paradigms for fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):984–985, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-050. See [2180].

Hansen:1997:EFE

[2183] M. M. Hansen and K-L. D. Mensberg. Erratum: Founder effects and genetic population structure of brown trout (Salmo trutta) in a Danish river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(4):986, April 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-074. See

erratum [2010].

Cook:1997:EE

[2184] D. G. Cook. Editorial/Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):??, May 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/davecook.

Jensen:1997:ORB

[2185] A. L. Jensen. Origin of relation between K and L_{inf} and synthesis of relations among life history parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):987–989, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-007.

Gunderson:1997:TBR

[2186] D. R. Gunderson. Trade-off between reproductive effort and adult survival in oviparous and viviparous fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):990–998, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-019.

Welch:1997:MZB

[2187] H. E. Welch, T. D. Siferd, and P. Bruecker. Marine zooplanktonic and benthic community respiration rates at resolute, Canadian high Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):999–1005, May 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-006.

Johnston:1997:WPV

[2188] T. A. Johnston. Within-population variability in egg characteristics of walleye (Stizostedion vitreum) and white sucker (Catostomus commersoni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1006–1014, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-008.

Yin:1997:EFF

[2189] K. Yin, P. J. Harrison, and R. J. Beamish. Effects of a fluctuation in Fraser River discharge of primary production in the central Strait of Georgia, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1015–1024, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-010.

Dunham:1997:RBB

[2190] J. B. Dunham and G. L. Vinyard. Relationships between body mass, population density, and the self-thinning rule in stream-living salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1025–1030, May 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-012.

Stow:1997:EPL

[2191] C. A. Stow, L. J. Jackson, and J. F. Amrhein. An examination of the PCB: lipid relationship among individual fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1031–1038, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-014.

Berntssen:1997:RSM

[2192] M. H. G. Berntssen, F. Kroglund, B. O. Rosseland, and S. E. Wendelaar Bonga. Responses of skin mucous cells to aluminium exposure at low pH in Atlantic salmon (Salmo salar) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1039–1045, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-015.

Borgmann:1997:TAZ

[2193] U. Borgmann and W. P. Norwood. Toxicity and accumulation of zinc and copper in Hyalella azteca exposed to metal-spiked sediments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1046–1054, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-020.

Borgmann:1997:ITA

[2194] U. Borgmann and W. P. Norwood. Identification of the toxic agent in metal-contaminated sediments from Manitouwadge Lake, Ontario, using toxicity–accumulation relationships in Hyalella azteca. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1055–1063, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-021.

Youngson:1997:CPP

[2195] A. F. Youngson, A. I. Mitchell, P. T. Noack, and L. M. Laird. Carotenoid pigment profiles distinguish anadromous and nonanadromous brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (5):1064–1066, May 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-023.

Watanabe:1997:LPD

[2196] Y. Watanabe, Y. Oozeki, and D. Kitagawa. Larval parameters determining preschooling juvenile production of Pacific saury (Cololabis saira) in the northwestern Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (5):1067–1076, May 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-013.

Swain:1997:SST

[2197] D. P. Swain. Sex-specific temperature distribution of American plaice (Hippoglossoides platessoides) and its relation to age and abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1077–1087, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-017.

Paterson:1997:CPF

[2198] M. J. Paterson, D. Findlay, K. Beaty, W. Findlay, E. U. Schindler, M. Stainton, and G. McCullough. Changes in the planktonic food web of a new experimental reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (5):1088–1102, May 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-018.

Fechhelm:1997:SGP

[2199] R. G. Fechhelm, J. D. Bryan, W. B. Griffiths, and L. R. Martin. Summer growth patterns of northern Dolly Varden (Salvelinus malma) smolts from the Prudhoe Bay region of Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1103–1110, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-022.

Pepin:1997:SDV

[2200] P. Pepin and J. T. Anderson. Scale-dependent variations in the precision of larval fish abundance estimates: a study of Sebastes sp. on Flemish Cap. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1111–1120, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-004.

Zheng:1997:AHS

[2201] J. Zheng, M. C. Murphy, and G. H. Kruse. Analysis of harvest strategies for red king crab, Paralithodes camtschaticus, in Bristol Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1121–1134, May 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-003.

Fraser:1997:APS

[2202] B. G. Fraser and D. D. Williams. Accuracy and precision in sampling hyporheic fauna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1135– 1141, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-024.

VanderZanden:1997:CTP

[2203] M. J. Vander Zanden, G. Cabana, and J. B. Rasmussen. Comparing trophic position of freshwater fish calculated using stable nitrogen isotope ratios (δ ^{15}N) and literature dietary data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1142–1158, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-016.

Cox:1997:CSM

[2204] S. P. Cox and S. G. Hinch. Changes in size at maturity of Fraser River sockeye salmon (Oncorhynchus nerka) (1952–1993) and associations with temperature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (5):1159–1165, May 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-009.

Stanley:1997:SSV

[2205] D. R. Stanley and C. A. Wilson. Seasonal and spatial variation in the abundance and size distribution of fishes associated with a petroleum platform in the northern Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1166–1176, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-005.

Gross:1997:FOL

[2206] H. P. Gross, W. A. Wurtsbaugh, C. Luecke, and P. Budy. Fertilization of an oligotrophic lake with a deep chlorophyll maximum: predicting the effect on primary productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (5):1177–1189, May 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-011.

Kerr:1997:LGL

[2207] S. R. Kerr and R. A. Ryder. The Laurentian Great Lakes experience: a prognosis for the fisheries of Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1190–1197, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-076.

Hutchings:1997:SII

[2208] J. A. Hutchings, C. Walters, and R. L. Haedrich. Is scientific inquiry incompatible with government information control? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(5):1198–1210, May 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-051.

Anonymous:1997:EE

[2209] Anonymous. Editorial/Éditorial. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):??, June 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/roffedit.

Rader:1997:FCD

[2210] Russell B. Rader. A functional classification of the drift: traits that influence invertebrate availability to salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1211–1234, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-025.

Unwin:1997:CLH

[2211] M. J. Unwin and G. J. Glova. Changes in life history parameters in a naturally spawning population of Chinook salmon (Oncorhynchus tshawytscha) associated with releases of hatchery-reared fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1235–1245, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-031.

Unwin:1997:FAS

[2212] M. J. Unwin. Fry-to-adult survival of natural and hatchery-produced Chinook salmon (Oncorhynchus tshawytscha) from a common origin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1246–1254, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-032.

France:1997:EDT

[2213] R. L. France and R. H. Peters. Ecosystem differences in the trophic enrichment of 13C in aquatic food webs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1255–1258, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-044.

Wipfli:1997:TIS

[2214] Mark S. Wipfli. Terrestrial invertebrates as salmonid prey and nitrogen sources in streams: contrasting old-growth and young-growth riparian forests in southeastern Alaska, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1259–1269, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-034.

Cabal:1997:EPR

[2215] J. Cabal, L. R. Harris, and E. J. H. Head. Egg production rates of Calanus finmarchicus in the Northwest Atlantic (Labrador Sea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1270–1279, June 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-028.

Claxton:1997:DFC

[2216] W. Trevor Claxton, André Martel, Ronald M. Dermott, and Elizabeth G. Boulding. Discrimination of field-collected juveniles of two introduced dreissenids (Dreissena polymorpha and Dreissena bugensis) using mitochondrial DNA and shell morphology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1280–1288, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-029.

McCann:1997:BLH

[2217] Kevin McCann and Brian Shuter. Bioenergetics of life history strategies and the comparative allometry of reproduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1289–1298, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-026.

France:1997:LWL

[2218] Robert France. Land water linkages: influences of riparian deforestation on lake thermocline depth and possible consequences for cold stenotherms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1299– 1305, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-030.

Gotceitas:1997:UEB

[2219] Vytenis Gotceitas, Sandra Fraser, and Joseph A. Brown. Use of eelgrass beds (Zostera marina) by juvenile Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1306–1319, June 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-033.

Wong:1997:TMB

[2220] Allan H. K. Wong, Donald J. McQueen, D. Dudley Williams, and Eric Demers. Transfer of mercury from benthic invertebrates to fishes in lakes with contrasting fish community structures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1320–1330, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-035.

Inoue:1997:JMS

[2221] Mikio Inoue, Shigeru Nakano, and Futoshi Nakamura. Juvenile masu salmon (Oncorhynchus masou) abundance and stream habitat relationships in northern Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (6):1331–1341, June 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-036.

Merrick:1997:DDS

[2222] Richard L. Merrick, M. Kathryn Chumbley, and G. Vernon Byrd. Diet diversity of Steller sea lions (Eumetopias jubatus) and their population decline in Alaska: a potential relationship. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1342–1348, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-037.

Quinn:1997:TFM

[2223] Thomas P. Quinn, Sayre Hodgson, and Charles Peven. Temperature, flow, and the migration of adult sockeye salmon (Oncorhynchus nerka) in the Columbia River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6): 1349–1360, June 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-038.

Peake:1997:RSP

[2224] S. Peake, F. W. H. Beamish, R. S. McKinley, D. A. Scruton, and C. Katopodis. Relating swimming performance of lake sturgeon, Acipenser fulvescens, to fishway design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1361–1366, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-039.

Bukaveckas:1997:EBA

[2225] Paul A. Bukaveckas and William Shaw. Effects of base addition and brook trout (Salvelinus fontinalis) introduction on the plankton community of an acidic Adirondack lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1367–1376, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-040.

Smith:1997:FAS

[2226] Stephen J. Smith, Sara J. Iverson, and W. D. Bowen. Fatty acid signatures and classification trees: new tools for investigating the foraging ecology of seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6): 1377–1386, June 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f97-041. See [2879] and reply [2880].

Greene:1997:TMM

[2227] Michael W. Greene and Richard M. Kocan. Toxicological mechanisms of a multicomponent agricultural seed protectant in the rainbow trout (Oncorhynchus mykiss) and fathead minnow (Pimephales promelas). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1387–1390, June 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-042.

OConnell:1997:DRT

[2228] Michael O’Connell, Roy G. Danzmann, Jean-Marie Cornuet, Jonathan M. Wright, and Moira M. Ferguson. Differentiation of rainbow trout (Oncorhynchus mykiss) populations in Lake Ontario and the evaluation of the stepwise mutation and infinite allele mutation models using microsatellite variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (6):1391–1399, June 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-043.

Myers:1997:SSI

[2229] R. A. Myers, G. Mertz, and J. Bridson. Spatial scales of interannual recruitment variations of marine, anadromous, and freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1400–1407, June 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-045.

Gries:1997:NAJ

[2230] Gabe Gries, Kevin G. Whalen, Francis Juanes, and Donna L. Parrish. Nocturnal activity of juvenile Atlantic salmon (Salmo salar) in late summer: evidence of diel activity partitioning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1408–1413, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-046.

Medley:1997:RIP

[2231] Paul A. H. Medley and Christopher H. Ninnes. A recruitment index and population model for spiny lobster (Panulirus argus) using catch and effort data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1414–1421, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-027.

Doubleday:1997:CSI

[2232] William G. Doubleday, D. B. Atkinson, and J. Baird. Comment: Scientific inquiry and fish stock assessment in the Canadian Department of Fisheries and Oceans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (6):1422–1426, June 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f97-244. See reply [2234].

Healey:1997:CIP

[2233] M. C. Healey. Comment: The interplay of policy, politics, and science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1427–1429, June 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-245. See [2234].

Hutchings:1997:RSI

[2234] Jeffrey A. Hutchings, Richard L. Haedrich, and Carl Walters. Reply: Scientific inquiry and fish stock assessment in the Canadian Department of Fisheries and Oceans and Reply: The interplay of policy, politics, and science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(6):1430–1431, June 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-246. See

[2232, 2233].

Labelle:1997:OSE

[2235] M. Labelle, C. J. Walters, and B. Riddell. Ocean survival and exploitation of Coho salmon (Oncorhynchus kisutch) stocks from the east coast of Vancouver Island, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1433–1449, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-052.

Murdoch:1997:MDE

[2236] M. H. Murdoch and P. D. N. Hebert. Mitochondrial DNA evidence of distinct glacial refugia for brown bullhead (Ameiurus nebulosus) in the Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1450– 1460, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-053.

Lochmann:1997:ACL

[2237] S. E. Lochmann, C. T. Taggart, D. A. Griffin, K. R. Thompson, and G. L. Maillet. Abundance and condition of larval cod (Gadus morhua) at a convergent front on Western Bank, Scotian Shelf. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1461–1479, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-054.

Derby:1997:PFC

[2238] C. E. Derby and J. R. Lovvorn. Predation on fish by cormorants and pelicans in a cold-water river: a field and modeling study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1480–1493, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-055.

Sajdak:1997:PRA

[2239] S. L. Sajdak and R. B. Phillips. Phylogenetic relationships among Coregonus species inferred from the DNA sequence of the first internal transcribed spacer (ITS1) of ribosomal DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1494–1503, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-057.

Mayer:1997:RBP

[2240] C. M. Mayer and D. H. Wahl. The relationship between prey selectivity and growth and survival in a larval fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1504–1512, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-056.

Seip:1997:DMS

[2241] K. L. Seip. Defining and measuring species interactions in aquatic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1513–1519, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-058.

Eklov:1997:EHC

[2242] P. Eklöv. Effects of habitat complexity and prey abundance on the spatial and temporal distributions of perch (Perca fluviatilis) and pike (Esox lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1520–1531, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-059.

Hosn:1997:USA

[2243] W. Aboul Hosn, P. Dutilleul, and D. Boisclair. Use of spectral analysis to estimate short-term periodicities in growth rates of brook trout Salvelinus frontinalis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1532–1541, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-060.

May:1997:GVM

[2244] B. May, C. C. Krueger, and H. L. Kincaid. Genetic variation at microsatellite loci in sturgeon: primer sequence homology in Acipenser and Scaphirhynchus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1542–1547, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-061.

Patton:1997:GVM

[2245] J. C. Patton, B. J. Gallaway, R. G. Fechhelm, and M. A. Cronin. Genetic variation of microsatellite and mitochondrial DNA markers in broad whitefish (Coregonus nasus) in the Colville and Sagavanirktok Rivers in northern Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1548–1556, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-062.

Vinebrooke:1997:PAI

[2246] R. D. Vinebrooke and M. D. Graham. Periphyton assemblages as indicators of recovery in acidified Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1557–1568, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-063.

Neuman:1997:STV

[2247] J. Neuman, D. L. Pearl, P. J. Ewins, R. Black, D. V. Weseloh, M. Pike, and K. Karwowski. Spatial and temporal variation in the diet of doublecrested cormorants (Phalacrocorax auritus) breeding on the lower Great Lakes in the early 1990s. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1569–1584, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-064.

Bradford:1997:IVD

[2248] M. J. Bradford and G. C. Taylor. Individual variation in dispersal behaviour of newly emerged Chinook salmon (Oncorhynchus tshawytscha) from the Upper Fraser River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1585–1592, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-065.

Whittier:1997:CDN

[2249] T. R. Whittier, D. B. Halliwell, and S. G. Paulsen. Cyprinid distributions in Northeast U.S.A. lakes: evidence of regional-scale minnow biodiversity losses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1593–1607, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-066.

Mertz:1997:IEN

[2250] G. Mertz and R. A. Myers. Influence of errors in natural mortality estimates in cohort analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1608–1612, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-067.

Miller:1997:SDP

[2251] R. J. Miller. Spatial differences in the productivity of American lobster in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1613–1618, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-068.

Trevorrow:1997:DMS

[2252] M. V. Trevorrow. Detection of migrating salmon in the Fraser River using 100-kHz sidescan sonars. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1619–1629, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-069.

Klerks:1997:EEZ

[2253] P. L. Klerks, P. C. Fraleigh, and J. E. Lawniczak. Effects of the exotic zebra mussel (Dreissena polymorpha) on metal cycling in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1630–1638, July 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-071.

Bedard:1997:SPM

[2254] C. Bédard and R. Knowles. Some properties of methane oxidation in a thermally stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (7):1639–1645, July 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f97-072. See erratum [2389].

Richards:1997:VCA

[2255] L. J. Richards, J. T. Schnute, and N. Olsen. Visualizing catch–age analysis: a case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1646– 1658, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-073.

Furgal:1997:URP

[2256] J. A. Furgal and R. E. H. Smith. Ultraviolet radiation and photosynthesis by Georgian Bay phytoplankton of varying nutrient and photoadaptive status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1659–1667, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-075.

Bell:1997:CRD

[2257] K. N. I. Bell. Complex recruitment dynamics with Doppler-like effects caused by shifts and cycles in age-at-recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1668–1681, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-070.

Welch:1997:OLM

[2258] D. W. Welch. Obituary: Leo Margolis, 1927–1997. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(7):1682–1684, July 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-126.

DeVries:1997:RSE

[2259] P. DeVries. Riverine salmonid egg burial depths: review of published data and implications for scour studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1685–1698, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-090.

Francis:1997:RFM

[2260] R. I. C. C. Francis and R. Shotton. “Risk” in fisheries management: a review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1699–1715, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
100.

Ramsay:1997:RME

[2261] P. M. Ramsay, S. D. Rundle, M. J. Attrill, M. G. Uttley, P. R. Williams, P. S. Elsmere, and A. Abada. A rapid method for estimating biomass size spectra of benthic metazoan communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1716–1724, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-081.

Duplisea:1997:DFB

[2262] D. E. Duplisea, S. R. Kerr, and L. M. Dickie. Demersal fish biomass size spectra on the Scotian Shelf, Canada: species replacement at the shelfwide scale. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1725– 1735, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f97-077.

Lam:1997:SEO

[2263] A. K-Y. Lam and E. E. Prepas. In situ evaluation of options for chemical treatment of hepatotoxic cyanobacterial blooms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1736–1742, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-078.

Brazner:1997:PFA

[2264] J. C. Brazner and E. W. Beals. Patterns in fish assemblages from coastal wetland and beach habitats in Green Bay, Lake Michigan: a multivariate analysis of abiotic and biotic forcing factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1743–1761, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-079.

Harding:1997:GRA

[2265] G. C. Harding, E. L. Kenchington, C. J. Bird, D. S. Pezzack, and D. C. Landry. Genetic relationships among subpopulations of the American lobster (Homarus americanus) as revealed by random amplified polymorphic DNA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1762–1771, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-082.

Hop:1997:BAC

[2266] H. Hop, W. M. Tonn, and H. E. Welch. Bioenergetics of Arctic cod (Boreogadus saida) at low temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1772–1784, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-086.

Foote:1997:SSM

[2267] C. J. Foote, G. S. Brown, and C. C. Wood. Spawning success of males using alternative mating tactics in sockeye salmon, Oncorhynchus nerka. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1785–1795, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-080.

Billerbeck:1997:LVV

[2268] J. M. Billerbeck, G. Ortí, and D. O. Conover. Latitudinal variation in vertebrate number has a genetic basis in the Atlantic silverside, Menidia menidia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1796–1801, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-084.

Bowman:1997:DTR

[2269] M. F. Bowman and R. C. Bailey. Does taxonomic resolution affect the multivariate description of the structure of freshwater benthic macroinvertebrate communities? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (8):1802–1807, August 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-085.

Warrillow:1997:DSM

[2270] J. A. Warrillow, D. C. Josephson, W. D. Youngs, and C. C. Krueger. Differences in sexual maturity and fall emigration between diploid and triploid brook trout (Salvelinus fontinalis) in an Adirondack lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1808–1812, August 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-087.

Vis:1997:GHR

[2271] M. L. Vis, S. M. Carr, W. R. Bowering, and W. S. Davidson. Greenland halibut (Reinhardtius hippoglossoides) in the North Atlantic are genetically homogeneous. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8): 1813–1821, August 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-088.

Heithaus:1997:GVC

[2272] M. R. Heithaus and R. H. Laushman. Genetic variation and conservation of stream fishes: influence of ecology, life history, and water quality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1822–1836, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-083.

Soleng:1997:STG

[2273] A. Soleng and T. A. Bakke. Salinity tolerance of Gyrodactylus salaris (Platyhelminthes, Monogenea): laboratory studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1837–1845, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-089.

Marty:1997:HCE

[2274] G. D. Marty, J. E. Hose, M. D. McGurk, E. D. Brown, and D. E. Hinton. Histopathology and cytogenetic evaluation of Pacific herring larvae exposed to petroleum hydrocarbons in the laboratory or in Prince William Sound, Alaska, after the Exxon Valdez oil spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1846–1857, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-091.

Zhou:1997:PTR

[2275] S. Zhou and T. C. Shirley. Performance of two red king crab pot designs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1858–1864, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-094.

Domanico:1997:PAP

[2276] M. J. Domanico, R. B. Phillips, and T. H. Oakley. Phylogenetic analysis of Pacific salmon (genus Oncorhynchus) using nuclear and mitochondrial DNA sequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8): 1865–1872, August 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-092.

King:1997:RTS

[2277] J. R. King, B. J. Shuter, and A. P. Zimmerman. The response of the thermal stratification of South Bay (Lake Huron) to climatic variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1873–1882, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-093. See

erratum [2587].

Soranno:1997:IPL

[2278] P. A. Soranno, S. R. Carpenter, and R. C. Lathrop. Internal phosphorus loading in Lake Mendota: response to external loads and weather. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1883–1893, August 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-095.

Keeley:1997:ADS

[2279] E. R. Keeley and J. W. A. Grant. Allometry of diet selectivity in juvenile Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1894–1902, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-096.

Thayer:1997:ZMD

[2280] S. A. Thayer, R. C. Haas, R. D. Hunter, and R. H. Kushler. Zebra mussel (Dreissena polymorpha) effects on sediment, other zoobenthos, and the diet and growth of adult yellow perch (Perca flavescens) in pond enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1903–1915, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
101.

Couillard:1997:CBP

[2281] C. M. Couillard, P. V. Hodson, and M. Castonguay. Correlations between pathological changes and chemical contamination in American eels, Anguilla rostrata, from the St. Lawrence River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1916–1927, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-097.

Hebert:1997:TDG

[2282] P. D. N. Hebert and T. L. Finston. Taxon diversity in the genus Holopedium (Crustacea: Cladocera) from the lakes of eastern North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1928–1936, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-098.

Harvey:1997:LSD

[2283] M. Harvey, J.-C. Therriault, and N. Simard. Late-summer distribution of phytoplankton in relation to water mass characteristics in Hudson Bay and Hudson Strait (Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (8):1937–1952, August 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-099.

He:1997:CMB

[2284] J. He and D. J. Stewart. Comment — measuring the bioenergetic cost of fish activity in situ using a globally dispersed radiotracer (^{137}Cs). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1953–1954, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-127. See [1858]

and reply [2285].

Rowan:1997:RMB

[2285] D. J. Rowan and J. B. Rasmussen. Reply — measuring the bioenergetic cost of fish activity in situ using a globally dispersed radiotracer (^{137}Cs). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(8):1955–1956, August 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-128. See [1858,

2284].

Larsson:1997:GRH

[2286] D. G. J. Larsson, C. C. Mylonas, Y. Zohar, and L. W. Crim. Gonadotropin-releasing hormone analogue (GnRH-A) induces multiple ovulations of high-quality eggs in a cold-water, batch-spawning teleost, the yellowtail flounder (Pleuronectes ferrugineus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):1957–1964, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-102.

Soranno:1997:FAT

[2287] P. A. Soranno. Factors affecting the timing of surface scums and epilimnetic blooms of blue-green algae in a eutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):1965–1975, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-104.

Liermann:1997:DFS

[2288] M. Liermann and R. Hilborn. Depensation in fish stocks: a hierarchic Bayesian meta-analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9): 1976–1984, September 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-105.

Yin:1997:FCT

[2289] K. Yin, P. J. Harrison, R. H. Goldblatt, M. A. St. John, and R. J. Beamish. Factors controlling the timing of the spring bloom in the Strait of Georgia estuary, British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):1985–1995, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-106.

Gauthier:1997:EID

[2290] S. Gauthier and D. Boisclair. The energetic implications of the diel onshore-offshore migration by dace(Phoxinus eos × P. neogaeus) in a small oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9): 1996–2006, September 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-107.

Rejwan:1997:STD

[2291] C. Rejwan, B. J. Shuter, M. S. Ridgway, and N. C. Collins. Spatial and temporal distributions of smallmouth bass (Micropterus dolomieu) nests in Lake Opeongo, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2007–2013, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-108.

Reid:1997:ECE

[2292] S. D. Reid, J. J. Dockray, T. K. Linton, D. G. McDonald, and C. M. Wood. Effects of chronic environmental acidification and a summer global warming scenario: protein synthesis in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9): 2014–2024, September 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-109.

Smith:1997:EIV

[2293] E. B. Smith, F. M. Williams, and C. R. Fisher. Effects of intrapopulation variability on von Bertalanffy growth parameter estimates from equal mark–recapture intervals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2025–2032, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-110.

Jansson:1997:HBA

[2294] H. Jansson and T. Öst. Hybridization between Atlantic salmon (Salmo salar) and brown trout (S. trutta) in a restored section of the River Dälalven, Sweden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2033– 2039, September 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-111.

Hamel:1997:TSA

[2295] P. Hamel, P. Magnan, M. Lapointe, and P. East. Timing of spawning and assessment of a degree-day model to predict the in situ embryonic developmental rate of white sucker, Catostomus commersoni. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2040–2048, September 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-103.

Hincks:1997:EPC

[2296] S. S. Hincks and G. L. Mackie. Effects of pH, calcium, alkalinity, hardness, and chlorophyll on the survival, growth, and reproductive success of zebra mussel (Dreissena polymorpha) in Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2049–2057, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-114.

Korman:1997:UET

[2297] J. Korman and P. S. Higgins. Utility of escapement time series data for monitoring the response of salmon populations to habitat alteration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2058–2067, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-115.

Campana:1997:CAP

[2298] S. E. Campana, S. R. Thorrold, C. M. Jones, D. Günther, M. Tubrett, H. Longerich, S. Jackson, N. M. Halden, J. M. Kalish, P. Piccoli, H. de Pontual, H. Troadec, J. Panfili, D. H. Secor, K. P. Severin, S. H. Sie, R. Thresher, W. J. Teesdale, and J. L. Campbell. Comparison of accuracy, precision, and sensitivity in elemental assays of fish otoliths using the electron microprobe, proton-induced X-ray emission, and laser ablation inductively coupled plasma mass spectrometry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2068–2079, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-117.

Mackas:1997:AEP

[2299] D. L. Mackas, R. Kieser, M. Saunders, D. R. Yelland, R. M. Brown, and D. F. Moore. Aggregation of euphausiids and Pacific hake (Merluccius productus) along the outer continental shelf off Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2080–2096, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-113.

Myers:1997:UMR

[2300] R. A. Myers, M. O. Hammill, and G. B. Stenson. Using mark-recapture to estimate the numbers of a migrating stage-structured population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2097–2104, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-116.

Prepas:1997:IAL

[2301] E. E. Prepas, K. M. Field, T. P. Murphy, W. L. Johnson, J. M. Burke, and W. M. Tonn. Introduction to the Amisk Lake Project: oxygenation of a deep, eutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2105– 2110, September 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-119.

Prepas:1997:EHO

[2302] E. E. Prepas and J. M. Burke. Effects of hypolimnetic oxygenation on water quality in Amisk Lake, Alberta, a deep, eutrophic lake with high internal phosphorus loading rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2111–2120, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-125.

Lawrence:1997:EWO

[2303] Gregory A. Lawrence, J. M. Burke, T. P. Murphy, and E. E. Prepas. Exchange of water and oxygen between the two basins of Amisk Lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2121–2132, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-235.

Webb:1997:IEW

[2304] D. J. Webb, R. D. Robarts, and E. E. Prepas. Influence of extended water column mixing during the first 2 years of hypolimnetic oxygenation on the phytoplankton community of Amisk Lake, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2133–2145, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-120.

Field:1997:IAD

[2305] K. M. Field and E. E. Prepas. Increased abundance and depth distribution of pelagic crustacean zooplankton during hypolimnetic oxygenation in a deep, eutrophic Alberta lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2146–2156, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-121.

Dinsmore:1997:IHOa

[2306] W. P. Dinsmore and E. E. Prepas. Impact of hypolimnetic oxygenation on profundal macroinvertebrates in a eutrophic lake in central Alberta. I. Changes in macroinvertebrate abundance and diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2157–2169, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-122.

Dinsmore:1997:IHOb

[2307] W. P. Dinsmore and E. E. Prepas. Impact of hypolimnetic oxygenation on profundal macroinvertebrates in a eutrophic lake in central Alberta. II. Changes in Chironomus spp. abundance and biomass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2170–2181, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-123.

Aku:1997:IHO

[2308] P. M. K. Aku, L. G. Rudstam, and W. M. Tonn. Impact of hypolimnetic oxygenation on the vertical distribution of cisco (Coregonus artedi) in Amisk Lake, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9): 2182–2195, September 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-124.

Aku:1997:CPS

[2309] P. M. K. Aku and W. M. Tonn. Changes in population structure, growth, and biomass of cisco (Coregonus artedi) during hypolimnetic oxygenation of a deep, eutrophic lake, Amisk Lake, Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2196–2206, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-118.

Mahon:1997:DFS

[2310] R. Mahon. Does fisheries science serve the needs of managers of small stocks in developing countries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2207–2213, September 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-112.

Schwarz:1997:EES

[2311] C. J. Schwarz, R. E. Bailey, J. R. Irvine, and F. C. Dalziel. Erratum: Estimating salmon spawning escapement using capture-recapture methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(9):2214, September 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-262. See [998].

DArcy:1997:ICT

[2312] Pierre D’Arcy and Richard Carignan. Influence of catchment topography on water chemistry in southeastern Québec Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2215–2227, October 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-129.

Hershberger:1997:VDC

[2313] P. K. Hershberger, J. E. Rensel, A. L. Matter, and F. B. Taub. Vertical distribution of the chloromonad flagellate heterosigma carterae in columns: implications for bloom development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2228–2234, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-131.

LePage:1997:PVS

[2314] C. LePage and P. Cury. Population viability and spatial fish reproductive strategies in constant and changing environments: an individual-based modelling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10): 2235–2246, October 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-132.

French:1997:RFN

[2315] T. D. French and P. A. Chambers. Reducing flows in the Nechako River (British Columbia, Canada): potential response of the macrophyte community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2247–2254, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
133.

Marchal:1997:MGO

[2316] P. Marchal. Managing growth overfishing with multiannual compromise strategies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2255–2276, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-135.

Weaver:1997:DLF

[2317] M. J. Weaver, J. J. Magnuson, and M. K. Clayton. Distribution of littoral fishes in structurally complex macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2277–2289, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-130.

Dawson:1997:QHL

[2318] K. A. Dawson, R. L. Eshenroder, M. E. Holey, and C. Ward. Quantification of historic lake trout (Salvelinus namaycush) spawning aggregations in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):

2290–2302, October 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-136.

Yamauchi:1997:PDF

[2319] A. Yamauchi and Y. Matsumiya. Population dynamics and fishery policy for migratory resources with two migration paths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2303–2313, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-137.

Cardinale:1997:CDE

[2320] B. J. Cardinale, T. M. Burton, and V. J. Brady. The community dynamics of epiphytic midge larvae across the pelagic–littoral interface: do animals respond to changes in the abiotic environment? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2314–2322, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-138.

Hammond:1997:BIT

[2321] T. Hammond. A Bayesian interpretation of target strength data from the Grand Banks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10): 2323–2333, October 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-139.

Leppard:1997:CSM

[2322] G. G. Leppard, M. M. West, D. T. Flannigan, J. Carson, and J. N. A. Lott. A classification scheme for marine organic colloids in the Adriatic Sea: colloid speciation by transmission electron microscopy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2334–2349, October 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-140.

Stockwell:1997:TDA

[2323] J. D. Stockwell and O. E. Johannsson. Temperature-dependent allometric models to estimate zooplankton production in temperate freshwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2350–2360, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-

141. See erratum [2608].

Peterson:1997:TIN

[2324] B. J. Peterson, M. Bahr, and G. W. Kling. A tracer investigation of nitrogen cycling in a pristine tundra river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2361–2367, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-142.

McKinnell:1997:PSA

[2325] S. McKinnell, J. J. Pella, and M. L. Dahlberg. Population-specific aggregations of steelhead trout (Oncorhynchus mykiss) in the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2368–2376, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
143.

Zydlewski:1997:LHA

[2326] J. Zydlewski and S. D. McCormick. The loss of hyperosmoregulatory ability in migrating juvenile American shad, Alosa sapidissima. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2377–2387, October 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-144.

Lambert:1997:CER

[2327] Y. Lambert and J.-D. Dutil. Condition and energy reserves of Atlantic cod (Gadus morhua) during the collapse of the northern Gulf of St. Lawrence stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10): 2388–2400, October 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-145.

Conover:1997:CVG

[2328] D. O. Conover, J. J. Brown, and A. Ehtisham. Countergradient variation in growth of young striped bass (Morone saxatilis) from different latitudes 1. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2401–2409, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
147.

Pickova:1997:EEC

[2329] J. Pickova, P. C. Dutta, P.-O. Larsson, and A. Kiessling. Early embryonic cleavage pattern, hatching success, and egg-lipid fatty acid composition: comparison between two cod (Gadus morhua) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2410–2416, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-148.

Lovett-Doust:1997:OCS

[2330] J. Lovett-Doust, L. Lovett-Doust, M. Biernacki, T. K. Mal, and R. Lazar. Organic contaminants in submersed macrophytes drifting in the Detroit River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2417–2427, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
150.

Roe:1997:DPS

[2331] S. L. Roe and H. J. MacIsaac. Deepwater population structure and reproductive state of quagga mussels (Dreissena bugensis) in Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2428–2433, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-151.

Fontaine:1997:GTM

[2332] P.-M. Fontaine, J. J. Dodson, L. Bernatchez, and A. Slettan. A genetic test of metapopulation structure in Atlantic salmon (Salmo salar) using microsatellites. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10): 2434–2442, October 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-152.

Bergeron:1997:LOC

[2333] E. Bergeron, L. N. Measures, and J. Huot. Lungworm (Otostrongylus circumlitus) infections in ringed seals (Phoca hispida) from eastern Arctic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2443–2448, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f97-
153.

Myers:1997:EDD

[2334] R. A. Myers, M. J. Bradford, J. M. Bridson, and G. Mertz. Estimating delayed density-dependent mortality in sockeye salmon (Oncorhynchus nerka): a meta-analytic approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2449–2462, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-146.

Minns:1997:QNN

[2335] C. K. Minns. Quantifying “no net loss” of productivity of fish habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(10):2463–2473, October 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-149.

Sandal:1997:FMO

[2336] Leif K. Sandal and Stein I. Steinshamn. A feedback model for the optimal management of renewable natural capital stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2475–2482, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-175.

Hansen:1997:GSL

[2337] Michael J. Hansen, Charles P. Madenjian, James H. Selgeby, and Thomas E. Helser. Gillnet selectivity for lake trout (Salvelinus namaycush) in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11): 2483–2490, November 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-156.

Pinckney:1997:ECP

[2338] James L. Pinckney, David F. Millie, Brian T. Vinyard, and Hans W. Paerl. Environmental controls of phytoplankton bloom dynamics in the Neuse River Estuary, North Carolina, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2491–2501, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-165.

Walton:1997:EPS

[2339] William E. Walton, Jamie A. Emiley, and Nelson G. Hairston, Jr. Effect of prey size on the estimation of behavioral visual resolution of bluegill (Lepomis macrochirus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11): 2502–2508, November 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-164.

Minshall:1997:PRL

[2340] G. Wayne Minshall, Christopher T. Robinson, and Deron E. Lawrence. Postfire responses of lotic ecosystems in Yellowstone National Park, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2509–2525, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-160.

VanWinkle:1997:ECT

[2341] Webb Van Winkle, Kenneth A. Rose, Brian J. Shuter, Henriette I. Jager, and Brady D. Holcomb. Effects of climatic temperature change on growth, survival, and reproduction of rainbow trout: predictions from a simulation model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11): 2526–2542, November 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-167.

Campbell:1997:ANO

[2342] Peter G. C. Campbell, Michael R. Twiss, and Kevin J. Wilkinson. Accumulation of natural organic matter on the surfaces of living cells: implications for the interaction of toxic solutes with aquatic biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2543–2554, November 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-161.

Whitledge:1997:ESE

[2343] Gregory W. Whitledge and Charles F. Rabeni. Energy sources and ecological role of crayfishes in an Ozark stream: insights from stable isotopes and gut analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11): 2555–2563, November 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-173.

Leblanc:1997:MRP

[2344] Judith Leblanc, Catherine M. Couillard, and Jean-Claude F. Brêthes. Modifications of the reproductive period in mummichog (Fundulus heteroclitus) living downstream from a bleached kraft pulp mill in the Miramichi Estuary, New Brunswick, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2564–2573, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-159.

Perrin:1997:PLB

[2345] Chris J. Perrin and John S. Richardson. N and P limitation of benthos abundance in the Nechako River, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2574–2583, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-163.

Dube:1997:NLH

[2346] Monique G. Dubé, Joseph M. Culp, and Garry J. Scrimgeour. Nutrient limitation and herbivory: processes influenced by bleached kraft pulp mill effluent. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2584– 2595, November 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-168.

Ricciardi:1997:RZM

[2347] Anthony Ricciardi, Fred G. Whoriskey, and Joseph B. Rasmussen. The role of the zebra mussel (Dreissena polymorpha) in structuring macroinvertebrate communities on hard substrata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2596–2608, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-174.

Sabo:1997:CBS

[2348] J. L. Sabo and G. B. Pauley. Competition between stream-dwelling cutthroat trout (Oncorhynchus clarki) and Coho salmon (Oncorhynchus kisutch): effects of relative size and population origin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2609–2617, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-155.

Einfalt:1997:PSJ

[2349] Lisa M. Einfalt and David H. Wahl. Prey selection by juvenile walleye as influenced by prey morphology and behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2618–2626, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-172.

Payne:1997:CML

[2350] Brigid A. Payne and Michel F. Lapointe. Channel morphology and lateral stability: effects on distribution of spawning and rearing habitat for Atlantic salmon in a wandering cobble-bed river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2627–2636, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-171.

Rippey:1997:ADI

[2351] Brian Rippey, N. John Anderson, and Robert H. Foy. Accuracy of diatom-inferred total phosphorus concentrations and the accelerated eutrophication of a lake due to reduced flushing and increased internal loading. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2637–2646, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-158.

Clifton:1997:CAQ

[2352] David R. Clifton and Russell J. Rodriguez. Characterization and application of a quantitative DNA marker that discriminates sex in Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2647–2652, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-177.

Speare:1997:EIH

[2353] David J. Speare and Garth J. Arsenault. Effects of intermittent hydrogen peroxide exposure on growth and columnaris disease prevention of juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2653–2658, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-169.

Stockwell:1997:RCB

[2354] Jason D. Stockwell and Brett M. Johnson. Refinement and calibration of a bioenergetics-based foraging model for Kokanee (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2659–2676, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-176.

Officer:1997:CGS

[2355] Rickard A. Officer, Rob W. Day, John G. Clement, and Lauren P. Brown. Captive gummy sharks, Mustelus antarcticus, form hypermineralised bands in their vertebrae during winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2677–2683, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-157.

Geiger:1997:VFS

[2356] Harold J. Geiger, William W. Smoker, Lev A. Zhivotovsky, and A. J. Gharrett. Variability of family size and marine survival in pink salmon (Oncorhynchus gorbuscha) has implications for conservation biology and human use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2684–2690, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-154.

Taylor:1997:ZSR

[2357] William D. Taylor and John C. H. Carter. Zooplankton size and its relationship to trophic status in deep Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2691–2699, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-166.

Ruzzante:1997:GDB

[2358] Daniel E. Ruzzante, Christopher T. Taggart, Doug Cook, and Sally V. Goddard. Genetic differentiation between inshore and offshore Atlantic cod (Gadus morhua) off Newfoundland: a test and evidence of temporal stability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2700– 2708, November 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-170.

Keenleyside:1997:DRF

[2359] Miles H. A. Keenleyside. Development of research on fish behaviour as part of fisheries science in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2709–2719, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-162.

Crawford:1997:CTD

[2360] Stephen S. Crawford and Bruce Morito. Comment: Toward a definition of conservation principles for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2720–2723, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-236. See [1631] and re-

ply [2361].

Shuter:1997:RTD

[2361] B. J. Shuter, C. K. Minns, and C. H. Olver. Reply: Toward a definition of conservation principles for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(11):2724–2725, November 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-237. See [1631, 2360].

Borchers:1997:IPD

[2362] D. L. Borchers, S. T. Buckland, I. G. Priede, and S. Ahmadi. Improving the precision of the daily egg production method using generalized additive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2727– 2742, December 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-134.

Stoeckmann:1997:SEB

[2363] Ann M. Stoeckmann and David W. Garton. A seasonal energy budget for zebra mussels (Dreissena polymorpha) in western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2743–2751, December 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-184.

Brodeur:1997:ICS

[2364] Julie C. Brodeur, Graham Sherwood, Joseph B. Rasmussen, and Alice Hontela. Impaired cortisol secretion in yellow perch (Perca flavescens) from lakes contaminated by heavy metals: in vivo and in vitro assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2752–2758, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-200.

Elliott:1997:ITS

[2365] J. K. Elliott and W. C. Leggett. Influence of temperature on sizedependent predation by a fish (Gasterosteus aculeatus) and a jellyfish (Aurelia aurita) on larval capelin (Mallotus villosus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2759–2766, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-190.

Berglund:1997:FIO

[2366] Olof Berglund, Per Larsson, Christer Brönmark, Larry Greenberg, Anders Eklöv, and Lennart Okla. Factors influencing organochlorine uptake in age-0 brown trout (Salmo trutta) in lotic environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2767–2774, December 1997. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-199.

Holmgren:1997:SRG

[2367] K. Holmgren, H. Wickström, and P. Clevestam. Sex-related growth of European eel, Anguilla anguilla, with focus on median silver eel age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2775–2781, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-193.

Tanasichuk:1997:IBO

[2368] R. W. Tanasichuk. Influence of biomass and ocean climate on the growth of Pacific herring (Clupea pallasi) from the southwest coast of Vancouver Island. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2782–2788, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-181.

Molot:1997:CMB

[2369] Lewis A. Molot and Peter J. Dillon. Colour–mass balances and colour– dissolved organic carbon relationships in lakes and streams in central Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2789–2795, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-196.

Chambers:1997:WOC

[2370] P. A. Chambers, G. J. Scrimgeour, and A. Pietroniro. Winter oxygen conditions in ice-covered rivers: the impact of pulp mill and municipal effluents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2796–2806, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-188.

Jackson:1997:QQS

[2371] Donald A. Jackson and Harold H. Harvey. Qualitative and quantitative sampling of lake fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2807–2813, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-182.

Wirgin:1997:MSA

[2372] Isaac I. Wirgin, John R. Waldman, Lorraine Maceda, Joseph Stabile, and Victor J. Vecchio. Mixed-stock analysis of Atlantic coast striped bass (Morone saxatilis) using nuclear DNA and mitochondrial DNA markers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2814–2826, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-195.

Bremset:1997:DSA

[2373] Gunnbjørn Bremset and Ole Kristian Berg. Density, size-at-age, and distribution of young Atlantic salmon (Salmo salar) and brown trout (Salmo trutta) in deep river pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2827–2836, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-185.

Murphy:1997:DMJ

[2374] Michael L. Murphy, K. V. Koski, J. Mitchel Lorenz, and John F. Thedinga. Downstream migrations of juvenile Pacific salmon (Oncorhynchus spp.) in a glacial transboundary river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2837–2846, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-178.

Saborowski:1997:AAG

[2375] Reinhard Saborowski, Jean-François Koprivnjak, Mitchell M. Sisak, Gerrit Sahling, Friedrich Buchholz, Ken R. Lum, and Rolf Schneider. Ascorbic acid in the gonads of North Sea dab (Limanda limanda) during the reproductive cycle. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12): 2847–2852, December 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-183.

Hudon:1997:IWL

[2376] Christiane Hudon. Impact of water level fluctuations on St. Lawrence River aquatic vegetation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12): 2853–2865, December 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-201.

Chiasson:1997:HBP

[2377] Wayne B. Chiasson, David L. G. Noakes, and F. William H. Beamish. Habitat, benthic prey, and distribution of juvenile lake sturgeon (Acipenser fulvescens) in northern Ontario rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2866–2871, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-194.

Sivertsen:1997:GEF

[2378] Knut Sivertsen. Geographic and environmental factors affecting the distribution of kelp beds and barren grounds and changes in biota associated with kelp reduction at sites along the Norwegian coast. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2872–2887, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-186.

Wilhelm:1997:NRS

[2379] Frank M. Wilhelm, David C. Lasenby, Ralph M. Wilhelm, and Ron Plante. A new recorder for simultaneously recording the activity and oxygen consumption of small benthic invertebrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2888–2893, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-179.

Alanara:1997:DNF

[2380] Anders Alanärä and Eva Brännäs. Diurnal and nocturnal feeding activity in Arctic char (Salvelinus alpinus) and rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2894–2900, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-187.

Koponen:1997:IIV

[2381] Kari Koponen, Ossi Ritola, Sirpa Huuskonen, Dan Linder, Katalin Monostory, and Pirjo Lindström-Seppä. Intrastrain and interstrain variability in biotransformation enzyme activities of rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12): 2901–2906, December 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-191.

Jackson:1997:AGM

[2382] G. D. Jackson, J. W. Forsythe, R. F. Hixon, and R. T. Hanlon. Age, growth, and maturation of Lolliguncula brevis (Cephalopoda:

Loliginidae) in the northwestern Gulf of Mexico with a comparison of length-frequency versus statolith age analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2907–2919, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-192.

Spencer:1997:ENP

[2383] Paul D. Spencer and Jeremy S. Collie. Effect of nonlinear predation rates on rebuilding the Georges Bank haddock (Melanogrammus aeglefinus) stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2920–2929, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-197.

Ion:1997:CLS

[2384] J. Ion, Y. de Lafontaine, P. Dumont, and L. Lapierre. Contaminant levels in St. Lawrence River yellow perch (Perca flavescens): spatial variation and implications for monitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2930–2946, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-198.

Hogaasen:1997:PLT

[2385] Helga Rachel Høgåsen and Patrick Prunet. Plasma levels of thyroxine, prolactin, and cortisol in migrating and resident wild Arctic char, Salvelinus alpinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2947– 2954, December 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-233.

Allard:1997:SDF

[2386] Jacques Allard and Ghislain A. Chouinard. A strategy to detect fish discarding by combining onboard and onshore sampling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2955–2963, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-180.

Mobrand:1997:ADE

[2387] Lars E. Mobrand, James A. Lichatowich, Lawrence C. Lestelle, and Thomas S. Vogel. An approach to describing ecosystem performance “through the eyes of salmon”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (12):2964–2973, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f97-189.

Shuter:1997:EFA

[2388] B. J. Shuter and K. K. Ing. Erratum: Factors affecting the production of zooplankton in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54 (12):2974, December 1997. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-289. See [2113].

Bedard:1997:ESP

[2389] Charles Bédard and Roger Knowles. Erratum: Some properties of methane oxidation in a thermally stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 54(12):2975, December 1997. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-300. See [2254].

Ruzzante:1998:CSM

[2390] Daniel E. Ruzzante. A comparison of several measures of genetic distance and population structure with microsatellite data: bias and sampling variance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):1–14, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-203.

Duffy:1998:PDP

[2391] W. G. Duffy. Population dynamics, production, and prey consumption of fathead minnows (Pimephales promelas) in prairie wetlands: a bioenergetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):15–27, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-204.

Cobb:1998:TIB

[2392] Susan E. Cobb and Mary C. Watzin. Trophic interactions between yellow perch (Perca flavescens) and their benthic prey in a littoral zone community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):28–36, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-205.

Gillis:1998:IIA

[2393] D. M. Gillis and R. M. Peterman. Implications of interference among fishing vessels and the ideal free distribution to the interpretation of CPUE. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):37–46, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-206.

Berg:1998:ESV

[2394] Ole Kristian Berg, Eivind Thronæs, and Gunnbjørn Bremset. Energetics and survival of virgin and repeat spawning brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):47–53, January 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-208.

Caraco:1998:ECC

[2395] N. F. Caraco and R. Miller. Effects of CO_{2} on competition between a cyanobacterium and eukaryotic phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):54–62, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-202.

Mallory:1998:CTS

[2396] M. L. Mallory, D. K. McNicol, D. A. Cluis, and C. Laberge. Chemical trends and status of small lakes near Sudbury, Ontario, 1983–1995: evidence of continued chemical recovery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):63–75, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-213.

Fahrig:1998:RET

[2397] L. Fahrig, S. E. Pope, K. M. Henein, and G. A. Rose. Relative effects of trap versus trawl fisheries on population dynamics of the northern cod (Gadus morhua) stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (1):76–85, January 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-215.

Robb:1998:ABD

[2398] Christina A. Robb and Randall M. Peterman. Application of Bayesian decision analysis to management of a sockeye salmon (Oncorhynchus nerka) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):86–98, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-220.

Sullivan:1998:IPH

[2399] Patrick J. Sullivan and Suzanne Dyner Rebert. Interpreting Pacific halibut catch statistics in the British Columbia individual quota program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):99–115, January 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-221.

Nislow:1998:FFB

[2400] Keith H. Nislow, Carol Folt, and Marco Seandel. Food and foraging behavior in relation to microhabitat use and survival of age-0 Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):116–127, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-222.

Dower:1998:EGF

[2401] John F. Dower, Pierre Pepin, and William C. Leggett. Enhanced gut fullness and an apparent shift in size selectivity by radiated shanny (Ulvaria subbifurcata) larvae in response to increased turbulence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):128–142, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-225.

Power:1998:DMR

[2402] M. Power. Decision-making for the restoration of Atlantic salmon populations damaged by acid precipitation: the interplay of biology and economics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):143–149, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-226.

Lougheed:1998:FRZ

[2403] Vanessa L. Lougheed and Patricia Chow-Fraser. Factors that regulate the zooplankton community structure of a turbid, hypereutrophic Great Lakes wetland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):150–161, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-227.

Casas:1998:IFO

[2404] Maria C. Casas. Increment formation in otoliths of slow-growing winter flounder (Pleuronectes americanus) larvae in cold water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):162–169, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-231.

Sellers:1998:PDL

[2405] Todd J. Sellers, Brian R. Parker, David W. Schindler, and William M. Tonn. Pelagic distribution of lake trout (Salvelinus namaycush) in small Canadian Shield lakes with respect to temperature, dissolved oxygen, and light. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):170–179, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-232.

Dorn:1998:FSF

[2406] Martin W. Dorn. Fine-scale fishing strategies of factory trawlers in a midwater trawl fishery for Pacific hake (Merluccius productus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):180–198, January 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-234.

Hearn:1998:PPS

[2407] William S. Hearn, Kenneth H. Pollock, and Elizabeth N. Brooks. Preand post-season tagging models: estimation of reporting rate and fishing and natural mortality rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (1):199–205, January 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-243.

Steinman:1998:CPA

[2408] Alan D. Steinman, Karl E. Havens, J. William Louda, Nancy M. Winfree, and Earl W. Baker. Characterization of the photoautotrophic algal and bacterial communities in a large, shallow, subtropical lake using HPLC-PDA based pigment analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (1):206–219, January 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-239.

Thorp:1998:ROR

[2409] James H. Thorp, James E. Alexander, Jr., Bonny L. Bukaveckas, Gary A. Cobbs, and Kurt L. Bresko. Responses of Ohio River and Lake Erie dreissenid molluscs to changes in temperature and turbidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):220–229, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-242.

Campbell:1998:EFR

[2410] Christine E. Campbell, Roy Knoechel, and Douglas Copeman. Evaluation of factors related to increased zooplankton biomass and altered species composition following impoundment of a Newfoundland reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):230–238, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-248.

Mercier:1998:RTS

[2411] Annie Mercier, Émilien Pelletier, and Jean-François Hamel. Response of temperate sea anemones to butyltin contamination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):239–245, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-249.

Chaput:1998:RST

[2412] G. Chaput, J. Allard, F. Caron, J. B. Dempson, C. C. Mullins, and M. F. O’Connell. River-specific target spawning requirements for Atlantic salmon (Salmo salar) based on a generalized smolt production model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):246–261, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-252.

Comeau:1998:GSD

[2413] Michel Comeau, Gérard Y. Conan, Francesc Maynou, Guy Robichaud, Jean-Claude Therriault, and Michel Starr. Growth, spatial distribution, and abundance of benthic stages of the snow crab (Chionoecetes opilio) in Bonne Bay, Newfoundland, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(1):262–279, January 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-255.

Schwarz:1998:USP

[2414] Carl James Schwarz and Carolyn Gail Taylor. Use of the stratified-Petersen estimator in fisheries management: estimating the number of pink salmon (Oncorhynchus gorbuscha) spawners in the Fraser River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):281–296, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-238.

Haag:1998:REF

[2415] Wendell R. Haag and Melvin L. Warren, Jr. Role of ecological factors and reproductive strategies in structuring freshwater mussel communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):297–306, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-210.

Rand:1998:DSD

[2416] Peter S. Rand and Donald J. Stewart. Dynamics of salmonine diets and foraging in Lake Ontario, 1983–1993: a test of a bioenergetic model prediction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):307–317, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-253.

Rand:1998:PFE

[2417] Peter S. Rand and Donald J. Stewart. Prey fish exploitation, salmonine production, and pelagic food web efficiency in Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):318–327, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-254.

Lange:1998:TKF

[2418] U. Lange, R. Saborowski, D. Siebers, F. Buchholz, and L. Karbe. Temperature as a key factor determining the regional variability of the xenobiotic-inducible ethoxyresorufin-O-deethylase activity in the liver of dab (Limanda limanda). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2): 328–338, February 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f97-273. See erratum [2586].

Stemberger:1998:FTM

[2419] Richard S. Stemberger and Celia Y. Chen. Fish tissue metals and zooplankton assemblages of northeastern U.S. lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):339–352, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-223.

Jensen:1998:SRA

[2420] A. L. Jensen. Simulation of relations among fish life history parameters with a bioenergetics-based population model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):353–357, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-259.

Clifford:1998:GCA

[2421] S. L. Clifford, P. McGinnity, and A. Ferguson. Genetic changes in Atlantic salmon (Salmo salar) populations of Northwest Irish rivers resulting from escapes of adult farm salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):358–363, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-229.

Jones:1998:DGP

[2422] Megan E. B. Jones and Christopher T. Taggart. Distribution of gill parasite (Lernaeocera branchialis) infection in Northwest Atlantic cod (Gadus morhua) and parasite-induced host mortality: inferences from tagging data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):364–375, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-209.

Cordue:1998:DOE

[2423] Patrick L. Cordue. Designing optimal estimators for fish stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):376–386, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-228.

Nibbelink:1998:IVG

[2424] Nathan P. Nibbelink and Stephen R. Carpenter. Interlake variation in growth and size structure of bluegill (Lepomis macrochirus): inverse analysis of an individual-based model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):387–396, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-207.

Wang:1998:IFM

[2425] You-Gan Wang. An improved fabens method for estimation of growth parameters in the von Bertalanffy model with individual asymptotes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):397–400, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-211.

Madon:1998:EIS

[2426] Sharook P. Madon, Daniel W. Schneider, James A. Stoeckel, and Richard E. Sparks. Effects of inorganic sediment and food concentrations on energetic processes of the zebra mussel, Dreissena polymorpha: implications for growth in turbid rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):401–413, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-214.

Cyr:1998:CCD

[2427] Hélène Cyr. Cladoceran- and copepod-dominated zooplankton communities graze at similar rates in low-productivity lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):414–422, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-217.

Hammill:1998:PPP

[2428] M. O. Hammill, G. B. Stenson, R. A. Myers, and W. T. Stobo. Pup production and population trends of the grey seal (Halichoerus grypus) in the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (2):423–430, February 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-218.

Limburg:1998:AMA

[2429] Karin E. Limburg. Anomalous migrations of anadromous herrings revealed with natural chemical tracers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):431–437, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-219.

Girard:1998:RIT

[2430] Caroline Girard, Julie C. Brodeur, and Alice Hontela. Responsiveness of the interrenal tissue of yellow perch (Perca flavescens) from contaminated sites to an ACTH challenge test in vivo. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):438–450, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-224.

Mooij:1998:SAS

[2431] W. M. Mooij and E. H. van Nes. Statistical analysis of the somatic growth rate of 0+ fish in relation to temperature under natural conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):451–458, February 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-240.

Alunno-Bruscia:1998:AAO

[2432] Marianne Alunno-Bruscia and Bernard Sainte-Marie. Abdomen allometry, ovary development, and growth of female snow crab, Chionoecetes opilio (Brachyura, Majidae), in the northwestern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):459–477, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-241.

Mertz:1998:SFF

[2433] G. Mertz and R. A. Myers. A simplified formulation for fish production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):478–484, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-216.

Giberson:1998:SHH

[2434] Donna J. Giberson and Daniel Caissie. Stream habitat hydraulics: interannual variability in three reaches of Catamaran Brook, New Brunswick. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):485–494, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-247.

Hill:1998:HZC

[2435] Alan R. Hill and Donna J. Lymburner. Hyporheic zone chemistry and stream-subsurface exchange in two groundwater-fed streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):495–506, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-250.

Soleng:1998:PDG

[2436] Arnulf Soleng, Tor A. Bakke, and Lars P. Hansen. Potential for dispersal of Gyrodactylus salaris (Platyhelminthes, Monogenea) by sea-running stages of the Atlantic salmon (Salmo salar): field and laboratory studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):507–514, February 1998.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-251.

Schnute:1998:AMF

[2437] Jon T. Schnute and Laura J. Richards. Analytical models for fishery reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(2):515–528, February 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-212.

Starr:1998:CSA

[2438] Paul Starr, John H. Annala, and Ray Hilborn. Contested stock assessment: two case studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (2):529–537, February 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-230.

Anonymous:1998:FBS

[2439] Anonymous. The Frank Bernard Scholarship in fisheries science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):??, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-024.

Reinert:1998:LTR

[2440] Thomas R. Reinert, Julie Wallin, Mary C. Griffin, Michael J. Conroy, and Michael J. Van Den Avyle. Long-term retention and detection of oxytetracycline marks applied to hatchery-reared larval striped bass, Morone saxatilis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):539–543, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-280.

Paterson:1998:DLS

[2441] Michael J. Paterson, D. C. G. Muir, B. Rosenberg, E. J. Fee, C. Anema, and W. Franzin. Does lake size affect concentrations of atmospherically derived polychlorinated biphenyls in water, sediment, zooplankton, and fish? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):544–553, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-276.

Kinnison:1998:PSV

[2442] Michael Kinnison, Martin Unwin, Nelson Boustead, and Thomas Quinn. Population-specific variation in body dimensions of adult Chinook salmon (Oncorhynchus tshawytscha) from New Zealand and their source population, 90 years after introduction. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):554–563, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-303.

Barbin:1998:ROH

[2443] Gayle P. Barbin. The role of olfaction in homing and estuarine migratory behavior of yellow-phase American eels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):564–575, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-274.

Linton:1998:CER

[2444] Tyler K. Linton, I. J. Morgan, P. J. Walsh, and Chris M. Wood. Chronic exposure of rainbow trout (Oncorhynchus mykiss) to simulated climate warming and sublethal ammonia: a year-long study of their appetite, growth, and metabolism. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (3):576–586, March 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-302.

Quinlan:1998:QIP

[2445] Roberto Quinlan, John P. Smol, and Roland I. Hall. Quantitative inferences of past hypolimnetic anoxia in south-central Ontario lakes using fossil midges (Diptera: Chironomidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):587–596, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-279.

Puel:1998:PSR

[2446] O. Puel, F. Galgani, C. Dalet, and P. Lassus. Partial sequence of the 24S rRNA and polymerase chain reaction based assay for the toxic dinoflagellate Dinophysis acuminata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (3):597–604, March 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-288.

Moles:1998:EOL

[2447] Adam Moles and Brenda L. Norcross. Effects of oil-laden sediments on growth and health of juvenile flatfishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):605–610, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-278.

Brown:1998:APR

[2448] Grant E. Brown and R. Jan F. Smith. Acquired predator recognition in juvenile rainbow trout (Oncorhynchus mykiss): conditioning hatchery-reared fish to recognize chemical cues of a predator. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):611–617, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-261.

Fukushima:1998:EEL

[2449] M. Fukushima, T. J. Quinn, and W. W. Smoker. Estimation of eggs lost from superimposed pink salmon (Oncorhynchus gorbuscha) redds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):618–625, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-260.

Elliott:1998:LCV

[2450] Joel K. Elliott and William C. Leggett. Larval condition and vulnerability to predation: an analysis based on mixed-prey experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):626–630, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-282.

Huse:1998:SSL

[2451] Geir Huse. Sex-specific life history strategies in capelin (Mallotus villosus)? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):631–638, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-275.

Kreutzweiser:1998:ZCR

[2452] David P. Kreutzweiser, John M. Gunn, Dean G. Thompson, Heather G. Pollard, and Marvin J. Faber. Zooplankton community responses to a novel forest insecticide, tebufenozide (RH-5992), in littoral lake enclosures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):639–648, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-277.

Peragon:1998:RBG

[2453] Juan Peragón, Juan B. Barroso, Manuel de la Higuera, and José A. Lupiáñez. Relationship between growth and protein turnover rates and nucleic acids in the liver of rainbow trout (Oncorhynchus mykiss) during development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):649–657, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f97-283.

Zabel:1998:MRM

[2454] Richard W. Zabel, James J. Anderson, and Pamela A. Shaw. A multiplereach model describing the migratory behavior of Snake River yearling Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):658–667, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-281.

Vinebrooke:1998:FPI

[2455] Rolf D. Vinebrooke, Roland I. Hall, Peter R. Leavitt, and Brian F. Cumming. Fossil pigments as indicators of phototrophic response to salinity and climatic change in lakes of western Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):668–681, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-263.

Peake:1998:RES

[2456] S. Peake and R. S. McKinley. A re-evaluation of swimming performance in juvenile salmonids relative to downstream migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):682–687, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-264.

deBisthoven:1998:RBH

[2457] Luc Janssens de Bisthoven, Jaap F. Postma, Peter Parren, Klaas R. Timmermans, and Frans Ollevier. Relations between heavy metals in aquatic sediments and in Chironomus larvae of Belgian lowland rivers and their morphological deformities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (3):688–703, March 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-265.

Watters:1998:NME

[2458] George Watters and Alistair J. Hobday. A new method for estimating the morphometric size at maturity of crabs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):704–714, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-266.

Estoup:1998:PAU

[2459] Arnaud Estoup, Karim Gharbi, Magali SanCristobal, Claude Chevalet, Pierrick Haffray, and René Guyomard. Parentage assignment using microsatellites in turbot (Scophthalmus maximus) and rainbow trout (Oncorhynchus mykiss) hatchery populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):715–723, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-268.

Furgal:1998:ECP

[2460] J. A. Furgal, W. D. Taylor, and R. E. H. Smith. Environmental control of photosynthate allocation in the phytoplankton of Georgian Bay (Lake Huron). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):726–736, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-267.

Fox:1998:BSR

[2461] Michael G. Fox and Alain J. Crivelli. Body size and reproductive allocation in a multiple spawning centrarchid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):737–748, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-269.

Chen:1998:PEM

[2462] Y. Chen and N. Andrew. Parameter estimation in modelling the dynamics of fish stock biomass: are currently used observation-error estimators reliable? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):749–760, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-270.

Skalski:1998:ESW

[2463] John R. Skalski. Estimating season-wide survival rates of outmigrating salmon smolt in the Snake River, Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):761–769, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-272.

Jonsdottir:1998:PGS

[2464] Ólöf D. B. Jónsdóttir, Albert K. Imsland, and Gunnar Nævdal. Population genetic studies of northern shrimp, Pandalus borealis, in Icelandic waters and the Denmark Strait. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(3):770–780, March 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-271.

Adams:1998:ESG

[2465] Noah S. Adams, Dennis W. Rondorf, Scott D. Evans, Joseph E. Kelly, and Russell W. Perry. Effects of surgically and gastrically implanted radio transmitters on swimming performance and predator avoidance of juvenile Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):781–787, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-285.

Dutil:1998:NAE

[2466] Jean-Denis Dutil, Yvan Lambert, Helga Guderley, Pierre U. Blier, Dany Pelletier, and Martine Desroches. Nucleic acids and enzymes in Atlantic cod (Gadus morhua) differing in condition and growth rate trajectories.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):788–795, April 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-294.

Marchand:1998:IFD

[2467] François Marchand and Daniel Boisclair. Influence of fish density on the energy allocation pattern of juvenile brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):796–805, April 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-301.

Graham:1998:TOB

[2468] Mark D. Graham and Rolf D. Vinebrooke. Trade-offs between herbivore resistance and competitiveness in periphyton of acidified lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):806–814, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-309.

Ericson:1998:DAH

[2469] G. Ericson, E. Lindesjöö, and L. Balk. DNA adducts and histopathological lesions in perch (Perca fluviatilis) and northern pike (Esox lucius) along a polycyclic aromatic hydrocarbon gradient on the Swedish coastline of the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (4):815–824, April 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-296.

Gosselin:1998:LNM

[2470] Jean-François Gosselin, Lena N. Measures, and Jean Huot. Lungworm (Nematoda: Metastrongyloidea) infections in Canadian phocids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):825–834, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-306.

Krabbenhoft:1998:MMD

[2471] David P. Krabbenhoft, Cynthia C. Gilmour, Janina M. Benoit, Christopher L. Babiarz, Anders W. Andren, and James P. Hurley. Methyl mercury dynamics in littoral sediments of a temperate seepage lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):835–844, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-304.

Boyd:1998:DMF

[2472] I. L. Boyd, D. J. McCafferty, K. Reid, R. Taylor, and T. R. Walker. Dispersal of male and female Antarctic fur seals (Arctocephalus gazella). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):845–852, April 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-314.

McLaughlin:1998:GAF

[2473] Robert L. McLaughlin and David L. G. Noakes. Going against the flow: an examination of the propulsive movements made by young brook trout in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):853–860, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-308.

Sanderson:1998:PRN

[2474] S. Laurie Sanderson, Mark E. Mort, and Joseph J. Cech, Jr. Particle retention by non-suspension-feeding cyprinid fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):861–868, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-292.

Kidd:1998:ETP

[2475] Karen A. Kidd, David W. Schindler, Raymond H. Hesslein, and Derek C. G. Muir. Effects of trophic position and lipid on organochlorine concentrations in fishes from subarctic lakes in Yukon Territory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):869–881, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-307.

Helser:1998:IDD

[2476] Thomas E. Helser and Jon K. T. Brodziak. Impacts of density-dependent growth and maturation on assessment advice to rebuild depleted U.S. silver hake (Merluccius bilinearis) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):882–892, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-290.

Rempel:1998:PFD

[2477] L. L. Rempel and D. G. Smith. Postglacial fish dispersal from the Mississippi refuge to the Mackenzie River basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):893–899, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-257.

Ma:1998:ERM

[2478] Y. Ma, O. S. Kjesbu, and T. Jørgensen. Effects of ration on the maturation and fecundity in captive Atlantic herring (Clupea harengus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):900–908, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-305.

Smokorowski:1998:QUR

[2479] K. E. Smokorowski, D. C. Lasenby, and R. D. Evans. Quantifying the uptake and release of cadmium and copper by the opossum shrimp Mysis relicta preying upon the cladoceran Daphnia magna using stable isotope tracers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):909–916, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-284.

Zeller:1998:MRP

[2480] Dirk C. Zeller and Garry R. Russ. Marine reserves: patterns of adult movement of the coral trout (Plectropomus leopardus (Serranidae)). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):917–924, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-311.

Sheehy:1998:ULR

[2481] Matt Sheehy, Nicolavito Caputi, Christopher Chubb, and Mark Belchier. Use of lipofuscin for resolving cohorts of western rock lobster (Panulirus cygnus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):925–936, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-016.

Welch:1998:TLO

[2482] D. W. Welch, Y. Ishida, and K. Nagasawa. Thermal limits and ocean migrations of sockeye salmon (Oncorhynchus nerka): long-term consequences of global warming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (4):937–948, April 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f98-023. See erratum [2585].

Richards:1998:MCC

[2483] Laura J. Richards and Jon T. Schnute. Model complexity and catch-age analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):949–957, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-005.

Bukaveckas:1998:PRN

[2484] Paul Bukaveckas and William Shaw. Phytoplankton responses to nutrient and grazer manipulations among northeastern lakes of varying pH. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):958–966, April 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-293.

Cyr:1998:EWD

[2485] Hélène Cyr. Effects of wave disturbance and substrate slope on sediment characteristics in the littoral zone of small lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):967–976, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-298.

Bennett:1998:LTC

[2486] Shannon N. Bennett, Martin L. Adamson, and Leo Margolis. Long-term changes in parasites of sockeye salmon (Oncorhynchus nerka) smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):977–986, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-295.

Bird:1998:FAH

[2487] G. A. Bird, W. J. Schwartz, and M. Motycka. Fate of ^{60}Co and ^{134}Cs added to the hypolimnion of a Canadian Shield lake: accumulation in biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):987–998, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-287.

Garnier-Laplace:1998:TMA

[2488] J. Garnier-Laplace, F. Vray, J. P. Baudin, and L. Jourd’heuil. The TRANSAQUA model applied to zebra mussels (Dreissena polymorpha) used for the biomonitoring of ^{137}Cs and ^{106}Ru contamination. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):999–1009, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-299.

Wilson:1998:PPD

[2489] Christopher C. Wilson and Paul D. N. Hebert. Phylogeography and postglacial dispersal of lake trout (Salvelinus namaycush) in North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):1010–1024, April 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-286.

Li:1998:GHD

[2490] Gang Li and Dennis Hedgecock. Genetic heterogeneity, detected by PCR-SSCP, among samples of larval Pacific oysters (Crassostrea gigas) supports the hypothesis of large variance in reproductive success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):1025–1033, April 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-312.

Naslund:1998:BTS

[2491] Ingemar Näslund, Erik Degerman, and Fredrik Nordwall. Brown trout (Salmo trutta) habitat use and life history in Swedish streams: possible effects of biotic interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (4):1034–1042, April 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-313.

Stewart:1998:EAC

[2492] Timothy W. Stewart, Jeffrey G. Miner, and Rex L. Lowe. An experimental analysis of crayfish (Orconectes rusticus) effects on a Dreissenadominated benthic macroinvertebrate community in western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(4):1043–1050, April 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-022.

Leland:1998:BID

[2493] Harry V. Leland and Steven V. Fend. Benthic invertebrate distributions in the San Joaquin River, California, in relation to physical and chemical factors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1051–1067, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-316.

Bussieres:1998:DAV

[2494] Dany Bussières, Marthe Monique Gagnon, Julian Dodson, and Peter V. Hodson. Does annual variation in growth and sexual maturation of white sucker (Catostomus commersoni) confound comparisons between pulp mill contaminated and reference rivers? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1068–1077, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-014.

Carignan:1998:MPP

[2495] Richard Carignan, Anne-Marie Blais, and Chantal Vis. Measurement of primary production and community respiration in oligotrophic lakes using the Winkler method. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (5):1078–1084, May 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-319.

Madenjian:1998:SDP

[2496] Charles P. Madenjian, George E. Noguchi, Robert C. Haas, and Kathrin S. Schrouder. Sexual difference in polychlorinated biphenyl accumulation rates of walleye (Stizostedion vitreum). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1085–1092, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-009.

Brodersen:1998:RTS

[2497] Klaus P. Brodersen, Melbourne C. Whiteside, and Claus Lindegaard. Reconstruction of trophic state in Danish lakes using subfossil chydorid (Cladocera) assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (5):1093–1103, May 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-008.

Urbani:1998:SCP

[2498] Nicola Urbani, Bernard Sainte-Marie, Jean-Marie Sévigny, David Zadworny, and Urs Kuhnlein. Sperm competition and paternity assurance during the first breeding period of female snow crab (Chionoecetes opilio) (Brachyura: Majidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (5):1104–1113, May 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-006.

Atwell:1998:BBM

[2499] Lisa Atwell, Keith A. Hobson, and Harold E. Welch. Biomagnification and bioaccumulation of mercury in an Arctic marine food web: insights from stable nitrogen isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1114–1121, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-001.

Hurst:1998:WMY

[2500] Thomas P. Hurst and David O. Conover. Winter mortality of young-of-the-year Hudson River striped bass (Morone saxatilis): size-dependent patterns and effects on recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1122–1130, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-017. See erratum [2656].

Dobson:1998:IDO

[2501] Evan P. Dobson and Gerald L. Mackie. Increased deposition of organic matter, polychlorinated biphenyls, and cadmium by zebra mussels (Dreissena polymorpha) in western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1131–1139, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-321.

Davis:1998:CBL

[2502] Bruce M. Davis and Thomas N. Todd. Competition between larval lake herring (Coregonus artedi) and lake whitefish (Coregonus clupeaformis) for zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1140– 1148, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-004.

Schultz:1998:DWS

[2503] Eric T. Schultz, David O. Conover, and Amir Ehtisham. The dead of winter: size-dependent variation and genetic differences in seasonal mortality among Atlantic silverside (Atherinidae: Menidia menidia) from different latitudes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1149–1157, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-320.

Friedland:1998:SCR

[2504] Kevin D. Friedland, David G. Reddin, Nobumichi Shimizu, Ruth E. Haas, and Alan F. Youngson. Strontium:calcium ratios in Atlantic salmon (Salmo salar) otoliths and observations on growth and maturation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1158–1168, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-011.

Lathrop:1998:PLR

[2505] Richard C. Lathrop, Stephen R. Carpenter, Craig A. Stow, Patricia A. Soranno, and John C. Panuska. Phosphorus loading reductions needed to control blue-green algal blooms in Lake Mendota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1169–1178, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-317.

Nonacs:1998:MFN

[2506] Peter Nonacs, Paul E. Smith, and Marc Mangel. Modeling foraging in the northern anchovy (Engraulis mordax): individual behavior can predict school dynamics and population biology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1179–1188, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-010.

Lougheed:1998:PEC

[2507] Vanessa L. Lougheed, Barb Crosbie, and Patricia Chow-Fraser. Predictions on the effect of common carp (Cyprinus carpio) exclusion on water quality, zooplankton, and submergent macrophytes in a Great Lakes wetland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1189–1197, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-315.

McDonald:1998:ACS

[2508] D. G. McDonald, W. J. McFarlane, and C. L. Milligan. Anaerobic capacity and swim performance of juvenile salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1198–1207, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-002.

McDonald:1998:CPJ

[2509] D. G. McDonald, C. L. Milligan, W. J. McFarlane, S. Croke, S. Currie, B. Hooke, R. B. Angus, B. L. Tufts, and K. Davidson. Condition and performance of juvenile Atlantic salmon (Salmo salar): effects of rearing practices on hatchery fish and comparison with wild fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1208–1219, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-003.

Vignaux:1998:FSM

[2510] M. Vignaux, G. A. Vignaux, S. Lizamore, and D. Gresham. Fine-scale mapping of fish distribution from commercial catch and effort data using maximum entropy tomography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1220–1227, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-297.

Klotz:1998:IBP

[2511] R. L. Klotz. Influence of beaver ponds on the phosphorus concentration of stream water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (5):1228–1235, May 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-318.

Smith:1998:EGM

[2512] Barry D. Smith, Louis W. Botsford, and Stephen R. Wing. Estimation of growth and mortality parameters from size frequency distributions lacking age patterns: the red sea urchin (Strongylocentrotus franciscanus) as an example. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (5):1236–1247, May 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-015.

Sinclair:1998:ETF

[2513] Alan F. Sinclair. Estimating trends in fishing mortality at age and length directly from research survey and commercial catch data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1248–1263, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-324.

Bernatchez:1998:MGD

[2514] Louis Bernatchez, J. Brian Dempson, and Sylvain Martin. Microsatellite gene diversity analysis in anadromous Arctic char, Salvelinus alpinus, from Labrador, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (5):1264–1272, May 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-325.

Mason:1998:CPF

[2515] Doran M. Mason, Timothy B. Johnson, and James F. Kitchell. Consequences of prey fish community dynamics on lake trout (Salvelinus namaycush) foraging efficiency in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1273–1284, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-007.

Cartwright:1998:QCT

[2516] Margaret A. Cartwright, David A. Beauchamp, and Mason D. Bryant. Quantifying cutthroat trout (Oncorhynchus clarki) predation on sockeye salmon (Oncorhynchus nerka) fry using a bioenergetics approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1285–1295, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-018.

Horne:1998:SSA

[2517] John K. Horne and Clarence S. Clay. Sonar systems and aquatic organisms: matching equipment and model parameters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(5):1296–1306, May 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f97-322.

Grant:1998:DFC

[2518] Scott M. Grant and Joseph A. Brown. Diel foraging cycles and interactions among juvenile Atlantic cod (Gadus morhua) at a nearshore site in Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6): 1307–1316, June 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-291.

Grant:1998:NSL

[2519] Scott M. Grant and Joseph A. Brown. Nearshore settlement and localized populations of age 0 Atlantic cod (Gadus morhua) in shallow coastal waters of Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (6):1317–1327, June 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-310.

Helser:1998:EGS

[2520] Thomas E. Helser, James P. Geaghan, and Richard E. Condrey. Estimating gillnet selectivity using nonlinear response surface regression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1328–1337, June 1998.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-036.

Ghan:1998:BCV

[2521] David Ghan, K. D. Hyatt, and J. D. McPhail. Benefits and costs of vertical migration by the freshwater copepod Skistodiaptomus oregonensis: testing hypotheses through population comparison. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1338–1349, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-028.

Ghan:1998:TSP

[2522] David Ghan, J. D. McPhail, and K. D. Hyatt. The temporal-spatial pattern of vertical migration by the freshwater copepod Skistodiaptomus oregonensis relative to predation risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1350–1363, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-027.

Hrabik:1998:PER

[2523] Thomas R. Hrabik, John J. Magnuson, and Ann S. McLain. Predicting the effects of rainbow smelt on native fishes in small lakes: evidence from long-term research on two lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1364–1371, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-032.

Marteinsdottir:1998:ISR

[2524] Gudrun Marteinsdottir and Kristjan Thorarinsson. Improving the stockrecruitment relationship in Icelandic cod (Gadus morhua) by including age diversity of spawners. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (6):1372–1377, June 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-035.

Kirsch:1998:DEF

[2525] Penny E. Kirsch, Sara J. Iverson, W. Don Bowen, Stephen R. Kerr, and Robert G. Ackman. Dietary effects on the fatty acid signature of whole Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1378–1386, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-019.

Hendry:1998:ITD

[2526] Andrew P. Hendry, Jay E. Hensleigh, and Reg R. Reisenbichler. Incubation temperature, developmental biology, and the divergence of sockeye salmon (Oncorhynchus nerka) within Lake Washington. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1387–1394, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-020.

McHenry:1998:CQC

[2527] Michael L. McHenry, Eric Shott, Robert H. Conrad, and Glenn B. Grette. Changes in the quantity and characteristics of large woody debris in streams of the Olympic Peninsula, Washington, U.S.A. (1982–1993). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1395–1407, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-013.

Link:1998:EVS

[2528] Michael R. Link and Randall M. Peterman. Estimating the value of inseason estimates of abundance of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1408–1418, June 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-039.

Trevorrow:1998:DAH

[2529] Mark V. Trevorrow and Ross R. Claytor. Detection of Atlantic herring (Clupea harengus) schools in shallow waters using high-frequency sidescan sonars. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1419–1429, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f98-
042.

Methven:1998:GIP

[2530] David A. Methven and David C. Schneider. Gear-independent patterns of variation in catch of juvenile Atlantic cod (Gadus morhua) in coastal habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1430–1442, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-025.

Rott:1998:MOP

[2531] Eugen Rott, Hamish C. Duthie, and Eveline Pipp. Monitoring organic pollution and eutrophication in the Grand River, Ontario, by means of diatoms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1443–1453, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-038.

Mittelbach:1998:OPE

[2532] Gary G. Mittelbach and Lennart Persson. The ontogeny of piscivory and its ecological consequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (6):1454–1465, June 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-041.

Hoenig:1998:MTS

[2533] John M. Hoenig, Nicholas J. Barrowman, William S. Hearn, and Kenneth H. Pollock. Multiyear tagging studies incorporating fishing effort data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1466–1476, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-256.

Hoenig:1998:MTD

[2534] John M. Hoenig, Nicholas J. Barrowman, Kenneth H. Pollock, Elizabeth N. Brooks, William S. Hearn, and Thomas Polacheck. Models for tagging data that allow for incomplete mixing of newly tagged animals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1477–1483, June 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f97-258.

Skalski:1998:UPI

[2535] John R. Skalski, Steven G. Smith, Robert N. Iwamoto, John G. Williams, and Annette Hoffmann. Use of passive integrated transponder tags to estimate survival of migrant juvenile salmonids in the Snake and Columbia rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1484–1493, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f97-323.

Kline:1998:NII

[2536] Thomas C. Kline, Jr., William J. Wilson, and John J. Goering. Natural isotope indicators of fish migration at Prudhoe Bay, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1494–1502, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-012.

Wipfli:1998:ISC

[2537] Mark S. Wipfli, John Hudson, and John Caouette. Influence of salmon carcasses on stream productivity: response of biofilm and benthic macroinvertebrates in southeastern Alaska, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1503–1511, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-031.

Rognerud:1998:CTE

[2538] Sigurd Rognerud, Trond Skotvold, Eirik Fjeld, Stephen A. Norton, and Anders Hobæk. Concentrations of trace elements in recent and preindustrial sediments from Norwegian and Russian Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1512–1523, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-026.

Foote:1998:ERB

[2539] Chris J. Foote and Gayle S. Brown. Ecological relationship between freshwater sculpins (genus Cottus) and beach-spawning sockeye salmon (Oncorhynchus nerka) in Iliamna Lake, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1524–1533, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-034.

Reid:1998:ECS

[2540] Scott D. Reid, T. K. Linton, J. J. Dockray, D. G. McDonald, and C. M. Wood. Effects of chronic sublethal ammonia and a simulated summer global warming scenario: protein synthesis in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6): 1534–1544, June 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-021.

Richards:1998:RIA

[2541] Laura J. Richards and Jean-Jacques Maguire. Recent international agreements and the precautionary approach: new directions for fisheries management science. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(6):1545– 1552, June 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-043.

Ferguson:1998:RGM

[2542] Moira M. Ferguson and Roy G. Danzmann. Role of genetic markers in fisheries and aquaculture: useful tools or stamp collecting? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1553–1563, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-096.

Johnson:1998:ETC

[2543] Paul D. Johnson and Robert F. McMahon. Effects of temperature and chronic hypoxia on survivorship of the zebra mussel (Dreissena polymorpha) and Asian clam (Corbicula fluminea). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1564–1572, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-030.

Arnott:1998:CZS

[2544] Shelley E. Arnott, John J. Magnuson, and Norman D. Yan. Crustacean zooplankton species richness: single- and multiple-year estimates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1573–1582, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-040.

Gregory:1998:IVI

[2545] T. Ryan Gregory and Chris M. Wood. Individual variation and interrelationships between swimming performance, growth rate, and feeding in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1583–1590, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-044.

Couture:1998:BCG

[2546] Patrice Couture, Jean-Denis Dutil, and Helga Guderley. Biochemical correlates of growth and condition in juvenile Atlantic cod (Gadus morhua) from Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (7):1591–1598, July 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-045.

Mesa:1998:VPP

[2547] Matthew G. Mesa, Thomas P. Poe, Alec G. Maule, and Carl B. Schreck. Vulnerability to predation and physiological stress responses in juvenile Chinook salmon (Oncorhynchus tshawytscha) experimentally infected with Renibacterium salmoninarum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1599–1606, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-049.

Cotter:1998:MEV

[2548] A. J. R. Cotter. Method for estimating variability due to sampling of catches on a trawl survey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (7):1607–1617, July 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-054.

Hughes:1998:PDE

[2549] Robert M. Hughes, Philip R. Kaufmann, Alan T. Herlihy, Thomas M. Kincaid, Lou Reynolds, and David P. Larsen. A process for developing and evaluating indices of fish assemblage integrity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1618–1631, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-060.

Stratoudakis:1998:DPC

[2550] Yorgos Stratoudakis, Robert J. Fryer, and Robin M. Cook. Discarding practices for commercial gadoids in the North Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1632–1644, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-029.

Robins:1998:IGP

[2551] Carolyn M. Robins, You-Gan Wang, and David Die. The impact of global positioning systems and plotters on fishing power in the northern prawn fishery, Australia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (7):1645–1651, July 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-037.

Mazumder:1998:NCS

[2552] Asit Mazumder and Karl E. Havens. Nutrient-chlorophyll-secchi relationships under contrasting grazer communities of temperate versus subtropical lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1652–1662, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-050.

Meekan:1998:OFS

[2553] M. G. Meekan, J. J. Dodson, S. P. Good, and D. A. J. Ryan. Otolith and fish size relationships, measurement error, and size-selective mortality during the early life of Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1663–1673, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-047.

Meekan:1998:DMS

[2554] M. G. Meekan, D. A. J. Ryan, J. J. Dodson, and S. P. Good. Does mortality select size or growth potential in young sockeye salmon (Oncorhynchus nerka)? A simulation based on otolith–fish size relationships. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1674–1681, July 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-046.

Zeldis:1998:DHM

[2555] John R. Zeldis, Rob C. Murdoch, Patrick L. Cordue, and Mike J. Page. Distribution of hoki (Macruronus novaezelandiae) eggs, larvae, and adults off Westland, New Zealand, and the design of an egg production survey to estimate hoki biomass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1682–1694, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-048.

Scharf:1998:PJP

[2556] Frederick S. Scharf, Jeffrey A. Buckel, Francis Juanes, and David O. Conover. Predation by juvenile piscivorous bluefish (Pomatomus saltatrix): the influence of prey to predator size ratio and prey type on predator capture success and prey profitability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1695–1703, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-056.

Mahon:1998:ABD

[2557] Robin Mahon, Stephen K. Brown, Kees C. T. Zwanenburg, D. Bruce Atkinson, Kenneth R. Buja, Larry Claflin, Geoffrey D. Howell, Mark E. Monaco, Robert N. O’Boyle, and Michael Sinclair. Assemblages and biogeography of demersal fishes of the east coast of North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1704–1738, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-065.

Pigeon:1998:MAS

[2558] Dany Pigeon, Julian J. Dodson, and Louis Bernatchez. A mtDNA analysis of spatiotemporal distribution of two sympatric larval populations of rainbow smelt (Osmerus mordax) in the St. Lawrence River estuary, Quebec, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1739– 1747, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-057.

Scribner:1998:DAM

[2559] Kim T. Scribner, Penelope A. Crane, William J. Spearman, and Lisa W. Seeb. DNA and allozyme markers provide concordant estimates of population differentiation: analyses of U.S. and Canadian populations of Yukon River fall-run chum salmon (Oncorhynchus keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1748–1758, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-063.

Ricciardi:1998:PII

[2560] Anthony Ricciardi and Joseph B. Rasmussen. Predicting the identity and impact of future biological invaders: a priority for aquatic resource management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1759–1765, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-066.

Marshall:1998:SBS

[2561] C. Tara Marshall, Olav Sigurd Kjesbu, Nathalia A. Yaragina, Per Solemdal, and Øyvind Ulltang. Is spawner biomass a sensitive measure of the reproductive and recruitment potential of Northeast Arctic cod? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1766–1783, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-062.

Young:1998:CID

[2562] Roger G. Young and Alexander D. Huryn. Comment: Improvements to the diurnal upstream–downstream dissolved oxygen change technique for determining whole-stream metabolism in small streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1784–1785, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-052. See [1320].

Marzolf:1998:RID

[2563] Erich R. Marzolf, Patrick J. Mulholland, and Alan D. Steinman. Reply: Improvements to the diurnal upstream–downstream dissolved oxygen change technique for determining whole-stream metabolism in small streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(7):1786–1787, July 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-051. See

[1320, 2562].

Attayde:1998:AIP

[2564] José Luiz Attayde and Reinaldo Luiz Bozelli. Assessing the indicator properties of zooplankton assemblages to disturbance gradients by canonical correspondence analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (8):1789–1797, August 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-033.

Schwarcz:1998:SCI

[2565] H. P. Schwarcz, Y. Gao, S. Campana, D. Browne, M. Knyf, and U. Brand. Stable carbon isotope variations in otoliths of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1798–1806, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-053.

Berg:1998:FDW

[2566] Neil Berg, Ann Carlson, and David Azuma. Function and dynamics of woody debris in stream reaches in the central Sierra Nevada, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1807–1820, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-064.

Hinch:1998:SSE

[2567] Scott G. Hinch and Peter S. Rand. Swim speeds and energy use of upriver-migrating sockeye salmon (Oncorhynchus nerka): role of local environment and fish characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1821–1831, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-067.

Rand:1998:SSE

[2568] Peter S. Rand and Scott G. Hinch. Swim speeds and energy use of upriver-migrating sockeye salmon (Oncorhynchus nerka): simulating metabolic power and assessing risk of energy depletion. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1832–1841, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-068.

Piet:1998:IEP

[2569] G. J. Piet. Impact of environmental perturbation on a tropical fish community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1842–1853, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f98-
073.

Anderson:1998:PTS

[2570] Charles S. Anderson. Partitioning total size selectivity of gill nets for walleye (Stizostedion vitreum) into encounter, contact, and retention components. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1854– 1863, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f98-070.

Murata:1998:MES

[2571] Shigenori Murata, Nobuyoshi Takasaki, Toshio Okazaki, Takanori Kobayashi, Ken ichi Numachi, Kun-Hsing Chang, and Norihiro Okada. Molecular evidence from short interspersed elements (SINEs) that Oncorhynchus masou (cherry salmon) is monophyletic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1864–1870, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-081.

Litvak:1998:SMP

[2572] Matthew K. Litvak and Edward A. Trippel. Sperm motility patterns of Atlantic cod (Gadus morhua) in relation to salinity: effects of ovarian fluid and egg presence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8): 1871–1877, August 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-093.

Mackas:1998:IVD

[2573] David L. Mackas, Robert Goldblatt, and Alan G. Lewis. Interdecadal variation in developmental timing of Neocalanus plumchrus populations at Ocean Station P in the subarctic North Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1878–1893, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-080.

Taniguchi:1998:TMC

[2574] Yoshinori Taniguchi, Frank J. Rahel, Douglas C. Novinger, and Kenneth G. Gerow. Temperature mediation of competitive interactions among three fish species that replace each other along longitudinal stream gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1894–1901, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-072.

Harvey:1998:ILW

[2575] Bret C. Harvey. Influence of large woody debris on retention, immigration, and growth of coastal cutthroat trout (Oncorhynchus clarki clarki) in stream pools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8): 1902–1908, August 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-083.

Bilby:1998:RJC

[2576] Robert E. Bilby, Brian R. Fransen, Peter A. Bisson, and Jason K. Walter. Response of juvenile Coho salmon (Oncorhynchus kisutch) and steelhead (Oncorhynchus mykiss) to the addition of salmon carcasses to two streams in southwestern Washington, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1909–1918, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-094.

MacLeod:1998:ELI

[2577] Neil A. MacLeod and David R. Barton. Effects of light intensity, water velocity, and species composition on carbon and nitrogen stable isotope ratios in periphyton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8): 1919–1925, August 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-075.

Larsson:1998:TPB

[2578] Per Larsson, Lennart Okla, and Gertrud Cronberg. Turnover of polychlorinated biphenyls in an oligotrophic and a eutrophic lake in relation to internal lake processes and atmospheric fallout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1926–1937, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-061.

Butturini:1998:APR

[2579] Andrea Butturini and Francesc Sabater. Ammonium and phosphate retention in a Mediterranean stream: hydrological versus temperature control. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1938–1945, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f98-
071.

Kinnison:1998:ESF

[2580] Michael T. Kinnison, Martin J. Unwin, William K. Hershberger, and Thomas P. Quinn. Egg size, fecundity, and development rate of two introduced New Zealand Chinook salmon (Oncorhynchus tshawytscha) populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1946–1953, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f98-
074.

McIntosh:1998:DDP

[2581] Angus R. McIntosh and Colin R. Townsend. Do different predators affect distance, direction, and destination of movements by a stream mayfly? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1954–1960, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-082.

Walls:1998:LHV

[2582] Mari Walls and Anne-Mari Ventelä. Life history variability in response to temperature and Chaoborus exposure in three Daphnia pulex clones. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1961–1970, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-088.

Heino:1998:MEF

[2583] Mikko Heino. Management of evolving fish stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1971–1982, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-089.

Bernard:1998:PPE

[2584] David R. Bernard, Robert P. Marshall, and John E. Clark. Planning programs to estimate salmon harvest with coded-wire tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1983–1995, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-069.

Welch:1998:ETL

[2585] D. W. Welch, Y. Ishida, and K. Nagasawa. Erratum: Thermal limits and ocean migrations of sockeye salmon (Oncorhynchus nerka): longterm consequences of global warming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1996, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-023e. See [2482].

Lange:1998:ETK

[2586] U. Lange, R. Saborowski, D. Siebers, F. Buchholz, and L. Karbe. Erratum: Temperature as a key factor determining the regional variability of the xenobiotic-inducible ethoxyresorufin-O-deethylase activity in the liver of dab (Limanda limanda). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1997–1998, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f97-273e. See [2418].

King:1998:ERT

[2587] Jacquelynne R. King, Brian J. Shuter, and Ann P. Zimmerman. Erratum: The response of the thermal stratification of South Bay (Lake Huron) to climatic variability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(8):1999, August 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f98-097. See [2277].

Anonymous:1998:PSI

[2588] Anonymous. Publishing of special issues and supplements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):??, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-920.

Anonymous:1998:SIS

[2589] Anonymous. The second international sysposium on krill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):??, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-091.

Brooks:1998:EFN

[2590] Elizabeth N. Brooks, Kenneth H. Pollock, John M. Hoenig, and William S. Hearn. Estimation of fishing and natural mortality from tagging studies on fisheries with two user groups. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2001–2010, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-058.

Thompson:1998:CUM

[2591] C. E. Thompson, W. R. Poole, M. A. Matthews, and A. Ferguson. Comparison, using minisatellite DNA profiling, of secondary male contribution in the fertilisation of wild and ranched Atlantic salmon (Salmo salar) ova. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2011–2018, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-076.

Neji:1998:EAV

[2592] H. Néji and J. de la Noüe. Effect of animal and vegetal protein diets on feed intake and apparent digestibility of nutrients in rainbow trout (Oncorhynchus mykiss) infected by Aeromonas salmonicida, with and without chronic hypoxia stress. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2019–2027, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-077.

Stevens:1998:RMS

[2593] E. Don Stevens, Arnold Sutterlin, and Todd Cook. Respiratory metabolism and swimming performance in growth hormone transgenic Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2028– 2035, September 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-078.

Hall:1998:MMA

[2594] B. D. Hall, D. M. Rosenberg, and A. P. Wiens. Methyl mercury in aquatic insects from an experimental reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2036–2047, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-079.

Hebert:1998:QGV

[2595] K. P. Hebert, P. L. Goddard, W. W. Smoker, and A. J. Gharrett. Quantitative genetic variation and genotype by environment interaction of embryo development rate in pink salmon (Oncorhynchus gorbuscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2048–2057, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-084.

Madon:1998:LER

[2596] Sharook P. Madon. Low-effort regression estimation of daily ration in young walleye, Stizostedion vitreum. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2058–2066, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-085.

Coronado:1998:STF

[2597] Claribel Coronado and Ray Hilborn. Spatial and temporal factors affecting survival in Coho salmon (Oncorhynchus kisutch) in the Pacific Northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2067–2077, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-090.

Hewitt:1998:HMF

[2598] L. M. Hewitt, K. R. Munkittrick, G. J. Van Der Kraak, I. M. Scott, L. P. Schleen, and M. R. Servos. Hepatic mixed function oxygenase activity and vitellogenin induction in fish following a treatment of the lampricide 3-trifluoromethyl-4-nitrophenol (TFM). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2078–2086, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-092.

Drenner:1998:SEN

[2599] Ray W. Drenner, Kirsten L. Gallo, Robert M. Baca, and J. Durward Smith. Synergistic effects of nutrient loading and omnivorous fish on phytoplankton biomass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9): 2087–2096, September 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-095.

Hawkins:1998:ESD

[2600] Denise K. Hawkins and Chris J. Foote. Early survival and development of coastal cutthroat trout (Oncorhynchus clarki clarki), steelhead (Oncorhynchus mykiss), and reciprocal hybrids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2097–2104, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-099.

Fournier:1998:MCL

[2601] Daid A. Fournier, John Hampton, and John R. Sibert. MULTIFAN-CL: a length-based, age-structured model for fisheries stock assessment, with application to South Pacific albacore, Thunnus alalunga. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2105–2116, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-100.

Marmorek:1998:AIA

[2602] D. R. Marmorek, G. L. Lacroix, J. Korman, I. Parnell, and W. D. Watt. Assessing the impacts of acidification on Atlantic salmon (Salmo salar): a simple model of stream chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2117–2126, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-102.

Pyper:1998:CMA

[2603] Brian J. Pyper and Randall M. Peterman. Comparison of methods to account for autocorrelation in correlation analyses of fish data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2127–2140, September 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-104. See erra-

tum [2657].

Lonzarich:1998:EHI

[2604] David George Lonzarich, Melvin L. Warren, Jr., and Mary Ruth Elger Lonzarich. Effects of habitat isolation on the recovery of fish assemblages in experimentally defaunated stream pools in Arkansas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2141–2149, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-113.

Wehr:1998:EDO

[2605] John D. Wehr, Dale A. Holen, Molli M. MacDonald, and Sean P. Lonergan. Effects of different organic carbon sources on a freshwater plankton community. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2150– 2160, September 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-098.

Shuter:1998:GLH

[2606] B. J. Shuter, M. L. Jones, R. M. Korver, and N. P. Lester. A general, life history based model for regional management of fish stocks: the inland lake trout (Salvelinus namaycush) fisheries of Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2161–2177, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-055.

Bisbal:1998:COC

[2607] Gustavo A. Bisbal and Willis E. McConnaha. Consideration of ocean conditions in the management of salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9):2178–2186, September 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-108.

Stockwell:1998:ETD

[2608] Jason D. Stockwell and Ora E. Johannsson. Erratum: Temperaturedependent allometric models to estimate zooplankton production in temperate freshwater lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(9): 2187, September 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f97-141e. See [2323].

Shuter:1998:EFA

[2609] Brian J. Shuter and K. K. Ing. Erratum: Factors affecting the production of zooplankton in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (9):2188, September 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f96-289e. See [2113].

Helbig:1998:PIPa

[2610] J. A. Helbig and P. Pepin. Partitioning the influence of physical processes on the estimation of ichthyoplankton mortality rates. I. Theory. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2189–2205, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-087.

Helbig:1998:PIPb

[2611] J. A. Helbig and P. Pepin. Partitioning the influence of physical processes on the estimation of ichthyoplankton mortality rates. II. Application to simulated and field data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10): 2206–2220, October 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-086.

Chapman:1998:EMA

[2612] Peter M. Chapman, Feiyue Wang, Colin Janssen, Guido Persoone, and Herbert E. Allen. Ecotoxicology of metals in aquatic sediments: binding and release, bioavailability, risk assessment, and remediation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2221–2243, October 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-145.

Lowerre-Barbieri:1998:MSD

[2613] Susan K. Lowerre-Barbieri, James M. Lowerre, and Luiz R. Barbieri. Multiple spawning and the dynamics of fish populations: inferences from an individual-based simulation model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2244–2254, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-105.

Downton:1998:RBA

[2614] Mary W. Downton and Kathleen A. Miller. Relationships between Alaskan salmon catch and North Pacific climate on interannual and interdecadal time scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10): 2255–2265, October 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-106.

Erkinaro:1998:THS

[2615] Jaakko Erkinaro, Eero Niemelä, Aila Saari, Yuri Shustov, and Lisbeth Jørgensen. Timing of habitat shift by Atlantic salmon parr from fluvial to lacustrine habitat: analysis of age distribution, growth, and scale characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2266–2273, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f98-
107.

Hurlbut:1998:MMD

[2616] Thomas Hurlbut and Douglas Clay. Morphometric and meristic differences between shallow- and deep-water populations of white hake (Urophycis tenuis) in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2274–2282, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-110.

Smedbol:1998:OIS

[2617] R. K. Smedbol, D. C. Schneider, J. S. Wroblewski, and D. A. Methven. Outcome of an inshore spawning event by northern Atlantic cod (Gadus morhua) at a low stock level. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55 (10):2283–2291, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f98-115.

Cooke:1998:SPN

[2618] Sandra E. Cooke and Ellie E. Prepas. Stream phosphorus and nitrogen export from agricultural and forested watersheds on the boreal plain. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2292–2299, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-118.

Carls:1998:EVH

[2619] M. G. Carls, G. D. Marty, T. R. Meyers, R. E. Thomas, and S. D. Rice. Expression of viral hemorrhagic septicemia virus in prespawning Pacific herring (Clupea pallasi) exposed to weathered crude oil. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2300–2309, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-116.

Essington:1998:HRR

[2620] Timothy E. Essington, Peter W. Sorensen, and Dean G. Paron. High rate of redd superimposition by brook trout (Salvelinus fontinalis) and brown trout (Salmo trutta) in a Minnesota stream cannot be explained by habitat availability alone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10): 2310–2316, October 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-109.

Hudon:1998:MAA

[2621] Christiane Hudon. Metal accumulation in American wild celery (Vallisneria americana Michx.) in the St. Lawrence River: effects of water depth and exposure to current. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2317–2328, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-114.

Casey:1998:DVT

[2622] Jill M. Casey and Ransom A. Myers. Diel variation in trawl catchability: is it as clear as day and night? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2329–2340, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-120.

Gilles:1998:GDI

[2623] André Gilles, Rémi Chappaz, Laurent Cavalli, Mathias Lörtscher, and Eric Faure. Genetic differentiation and introgression between putative subspecies of Leuciscus soufia (Teleostei: Cyprinidae) of the region of the Mediterranean Alps. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10): 2341–2354, October 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-103.

Myers:1998:SDU

[2624] Ransom A. Myers, Gordon Mertz, Jessica M. Bridson, and Michael J. Bradford. Simple dynamics underlie sockeye salmon (Oncorhynchus nerka) cycles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(10):2355– 2364, October 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f98-059.

Owen:1998:BNB

[2625] S. F. Owen, D. F. Houlihan, M. J. Rennie, and J. H. van Weerd. Bioenergetics and nitrogen balance of the European eel (Anguilla anguilla) fed at high and low ration levels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2365–2375, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-119.

Silverstein:1998:EGF

[2626] Jeffrey T. Silverstein, Karl D. Shearer, Walton W. Dickhoff, and Erika M. Plisetskaya. Effects of growth and fatness on sexual development of Chinook salmon (Oncorhynchus tshawytscha) parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2376–2382, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-111.

Nickelson:1998:PVC

[2627] Thomas E. Nickelson and Peter W. Lawson. Population viability of Coho salmon, Oncorhynchus kisutch, in Oregon coastal basins: application of a habitat-based life cycle model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2383–2392, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-123.

Wang:1998:EIV

[2628] You-Gan Wang and Nick Ellis. Effect of individual variability on estimation of population parameters from length-frequency data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2393–2401, November 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-134.

Nalepa:1998:DBM

[2629] Thomas F. Nalepa, David J. Hartson, David L. Fanslow, Gregory A. Lang, and Stephen J. Lozano. Declines in benthic macroinvertebrate populations in southern Lake Michigan, 1980–1993. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2402–2413, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-112.

Garvey:1998:OGS

[2630] James E. Garvey, Russell A. Wright, and Roy A. Stein. Overwinter growth and survival of age-0 largemouth bass (Micropterus salmoides): revisiting the role of body size. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2414–2424, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-124.

Benoit:1998:CLM

[2631] Hugues P. Benoît, John R. Post, Eric A. Parkinson, and N. Tom Johnston. Colonization by lentic macroinvertebrates: evaluating colonization processes using artificial substrates and appraising applicability of the technique. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2425–2435, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-125.

James-Pirri:1998:IST

[2632] Mary-Jane James-Pirri, J. Stanley Cobb, and Richard A. Wahle. Influence of settlement time and size on postsettlement growth in the American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2436–2446, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-131.

McGeer:1998:PEW

[2633] James C. McGeer and Chris M. Wood. Protective effects of water Cl^{−} on physiological responses to waterborne silver in rainbow trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2447–2454, November 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-133.

Wada:1998:RSR

[2634] Tokio Wada and Larry D. Jacobson. Regimes and stock-recruitment relationships in Japanese sardine (Sardinops melanostictus), 1951–1995. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2455–2463, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-135.

Clair:1998:CRC

[2635] Thomas A. Clair, James Ehrman, and Kaz Higuchi. Changes to the runoff of Canadian ecozones under a doubled CO_{2} atmosphere. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2464–2477, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-129.

Bradford:1998:INA

[2636] David F. Bradford, Scott D. Cooper, Thomas M. Jenkins, Jr., Kim Kratz, Orlando Sarnelle, and Aaron D. Brown. Influences of natural acidity and introduced fish on faunal assemblages in California alpine lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2478–2491, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-128.

Burwen:1998:ESA

[2637] Deborah L. Burwen and Steven J. Fleischman. Evaluation of side-aspect target strength and pulse width as potential hydroacoustic discriminators of fish species in rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11): 2492–2502, November 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-136.

Peterman:1998:PCS

[2638] Randall M. Peterman, Brian J. Pyper, Michael F. Lapointe, Milo D. Adkison, and Carl J. Walters. Patterns of covariation in survival rates of British Columbian and Alaskan sockeye salmon (Oncorhynchus nerka) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2503–2517, November 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-179.

He:1998:OER

[2639] Jixiang He and Donald J. Stewart. Ontogeny of energetic relationships and potential effects of tissue turnover: a comparative modeling study on lake trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(11):2518– 2532, November 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-117.

Walters:1998:SRR

[2640] Carl Walters and Bruce Ward. Is solar radiation responsible for declines in marine survival rates of anadromous salmonids that rear in small streams? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2533– 2538, December 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-203.

Grant:1998:WRF

[2641] W. Stewart Grant, A.-M. Clark, and Brian W. Bowen. Why restriction fragment length polymorphism analysis of mitochondrial DNA failed to resolve sardine (Sardinops) biogeography: insights from mitochondrial DNA cytochrome b sequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2539–2547, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-127.

Swain:1998:SVH

[2642] D. P. Swain, G. A. Chouinard, R. Morin, and K. F. Drinkwater. Seasonal variation in the habitat associations of Atlantic cod (Gadus morhua) and American plaice (Hippoglossoides platessoides) from the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12): 2548–2561, December 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-122.

Jarvinen:1998:ICF

[2643] Marko Järvinen and Kalevi Salonen. Influence of changing food web structure on nutrient limitation of phytoplankton in a highly humic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2562–2571, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-137.

Gopalan:1998:ERE

[2644] Gouthaman Gopalan, David A. Culver, Lin Wu, and Bruce K. Trauben. Effects of recent ecosystem changes on the recruitment of young-of-the-year fish in western Lake Erie. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2572–2579, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-130.

Lady:1998:ESR

[2645] James M. Lady and John R. Skalski. Estimators of stream residence time of Pacific salmon (Oncorhynchus spp.) based on release-recapture data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2580–2587, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-132.

Post:1998:IAA

[2646] David M. Post, James F. Kitchell, and James R. Hodgson. Interactions among adult demography, spawning date, growth rate, predation, overwinter mortality, and the recruitment of largemouth bass in a northern lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2588–2600, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-139.

Hobson:1998:SCN

[2647] Keith A. Hobson and Don M. Schell. Stable carbon and nitrogen isotope patterns in baleen from eastern Arctic bowhead whales (Balaena mysticetus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2601–2607, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-142.

Augustin:1998:SMA

[2648] N. H. Augustin, D. L. Borchers, E. D. Clarke, S. T. Buckland, and M. Walsh. Spatiotemporal modelling for the annual egg production method of stock assessment using generalized additive models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2608–2621, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-143.

McAllister:1998:MEF

[2649] Murdoch K. McAllister. Modeling the effects of fish migration on bias and variance in area-swept estimates of biomass: a vector-based approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2622–2641, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-126.

McAllister:1998:UBD

[2650] M. K. McAllister and G. P. Kirkwood. Using Bayesian decision analysis to help achieve a precautionary approach for managing developing fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2642–2661, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-121.

Sirois:1998:OBB

[2651] Pascal Sirois, Frédéric Lecomte, and Julian J. Dodson. An otolith-based back-calculation method to account for time-varying growth rate in rainbow smelt (Osmerus mordax) larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2662–2671, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-170.

Pelletier:1998:IRS

[2652] D. Pelletier. Intercalibration of research survey vessels in fisheries: a review and an application. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12): 2672–2690, December 1998. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-151.

Walters:1998:EQM

[2653] Carl Walters. Evaluation of quota management policies for developing fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2691–2705, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f98-172.

Hansson:1998:MSF

[2654] Sture Hansson. Methods of studying fish feeding: a comment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2706–2707, December 1998. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-158.

Cortes:1998:MSF

[2655] Enric Cortés. Methods of studying fish feeding: reply. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2708, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-159.

Hurst:1998:EWM

[2656] Thomas P. Hurst and David O. Conover. Erratum: Winter mortality of young-of-the-year Hudson River striped bass (Morone saxatilis): size-dependent patterns and effects on recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2709, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-202. See [2500].

Pyper:1998:ECM

[2657] Brian J. Pyper and Randall M. Peterman. Erratum: Comparison of methods to account for autocorrelation in correlation analyses of fish data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 55(12):2710, December 1998. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-201. See

[2603].

Anonymous:1999:OND

[2658] Anonymous. Obituary/nécrologie: Dr. Robert Haynes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):0–1, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-900.

MacIsaac:1999:ILO

[2659] Hugh J. MacIsaac, Igor A. Grigorovich, James A. Hoyle, Norman D. Yan, and Vadim E. Panov. Invasion of Lake Ontario by the Ponto–Caspian predatory cladoceran Cercopagis pengoi. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):1–5, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-210.

Gerlotto:1999:TDT

[2660] François Gerlotto, Marc Soria, and Pierre Fréon. From two dimensions to three: the use of multibeam sonar for a new approach in fisheries acoustics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):6–12, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-138.

Hara:1999:VGS

[2661] Toshiaki J. Hara, Joachim Carolsfeld, and Shoji Kitamura. The variability of the gustatory sensibility in salmonids, with special reference to strain differences in rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):13–24, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-140.

Leising:1999:LAC

[2662] Andrew W. Leising and Peter J. S. Franks. Larval Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) growth on Georges Bank: a model with temperature, prey size, and turbulence forcing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):25–36, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-144.

Meyer:1999:BSA

[2663] Renate Meyer and Russell B. Millar. Bayesian stock assessment using a state–space implementation of the delay difference model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):37–52, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-146.

Lamon:1999:RDP

[2664] E. Conrad Lamon III, S. R. Carpenter, and C. A. Stow. Rates of decrease of polychlorinated biphenyl concentrations in five species of Lake Michigan salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (1):53–59, January 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-147.

Poister:1999:IGT

[2665] David Poister, David E. Armstrong, and James P. Hurley. Influences of grazing on temporal patterns of algal pigments in suspended and sedimenting algae in a north temperate lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):60–69, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-153.

Persson:1999:DSF

[2666] Anders Persson and Lars-Anders Hansson. Diet shift in fish following competitive release. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (1):70–78, January 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-141.

Whalen:1999:EMP

[2667] Kevin G. Whalen and Donna L. Parrish. Effect of maturation on parr growth and smolt recruitment of Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):79–86, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-154.

Whalen:1999:EIF

[2668] Kevin G. Whalen, Donna L. Parrish, and Martha E. Mather. Effect of ice formation on selection of habitats and winter distribution of postyoung-of-the-year Atlantic salmon parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):87–96, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-156.

Hutchings:1999:SBA

[2669] Jeffrey A. Hutchings, Todd D. Bishop, and Carolyn R. McGregor-Shaw. Spawning behaviour of Atlantic cod, Gadus morhua: evidence of mate competition and mate choice in a broadcast spawner. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):97–104, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-216.

Outridge:1999:SDA

[2670] P. M. Outridge and R. E. A. Stewart. Stock discrimination of Atlantic walrus (Odobenus rosmarus rosmarus) in the eastern Canadian Arctic using lead isotope and element signatures in teeth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):105–112, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-155.

Trippel:1999:MHP

[2671] Edward A. Trippel, Michael B. Strong, John M. Terhune, and Jeremy D. Conway. Mitigation of harbour porpoise (Phocoena phocoena) by-catch in the gillnet fishery in the lower Bay of Fundy. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):113–123, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-162.

Lohman:1999:NSC

[2672] Kirk Lohman and John R. Jones. Nutrient–sestonic chlorophyll relationships in northern Ozark streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):124–130, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-160.

Dixit:1999:AWQ

[2673] Sushil S. Dixit, John P. Smol, Donald F. Charles, Robert M. Hughes, Steven G. Paulsen, and Gary B. Collins. Assessing water quality changes in the lakes of the northeastern United States using sediment diatoms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):131–152, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-148.

Nicholls:1999:RCP

[2674] K. H. Nicholls, G. J. Hopkins, and S. J. Standke. Reduced chlorophyll to phosphorus ratios in nearshore Great Lakes waters coincide with the establishment of dreissenid mussels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(1):153–161, January 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-149.

Arnott:1999:IVS

[2675] Shelley E. Arnott, Norman D. Yan, John J. Magnuson, and Thomas M. Frost. Interannual variability and species turnover of crustacean zooplankton in Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (1):162–172, January 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-152.

Lutcavage:1999:RPS

[2676] Molly E. Lutcavage, Richard W. Brill, Gregory B. Skomal, Bradford C. Chase, and Paul W. Howey. Results of pop-up satellite tagging of spawning size class fish in the Gulf of Maine: do North Atlantic bluefin tuna spawn in the mid-Atlantic? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (2):173–177, February 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-016.

Ciereszko:1999:PRA

[2677] Andrzej Ciereszko, Konrad Dabrowski, Feng Lin, and Li Liu. Protective role of ascorbic acid against damage to male germ cells in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2): 178–183, February 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-161.

Brodeur:1999:IHR

[2678] Julie C. Brodeur, Trine Ytrestøyl, Bengt Finstad, and R. Scott McKinley. Increase of heart rate without elevation of cardiac output in adult Atlantic salmon (Salmo salar) exposed to acidic water and aluminium. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):184–190, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-168.

Power:1999:LMA

[2679] James H. Power and E. Barry Moser. Linear model analysis of net catch data using the negative binomial distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):191–200, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-150.

Duston:1999:EWF

[2680] J. Duston and R. L. Saunders. Effect of winter food deprivation on growth and sexual maturity of Atlantic salmon (Salmo salar) in seawater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):201–207, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-165.

Patterson:1999:EUH

[2681] K. R. Patterson. Evaluating uncertainty in harvest control law catches using Bayesian Markov chain Monte Carlo virtual population analysis with adaptive rejection sampling and including structural uncertainty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):208–221, February 1999.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-157.

Witting:1999:LFM

[2682] David A. Witting, Kenneth W. Able, and Michael P. Fahay. Larval fishes of a Middle Atlantic Bight estuary: assemblage structure and temporal stability. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):222–230, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-175.

McCarthy:1999:EWT

[2683] Ian D. McCarthy, Erlend Moksness, Dimitri A. Pavlov, and Dominic F. Houlihan. Effects of water temperature on protein synthesis and protein growth in juvenile Atlantic wolffish (Anarhichas lupus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):231–241, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-171.

Clark:1999:DCG

[2684] William G. Clark, Steven R. Hare, Ana M. Parma, Patrick J. Sullivan, and Robert J. Trumble. Decadal changes in growth and recruitment of Pacific halibut (Hippoglossus stenolepis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):242–252, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-163.

Holland:1999:EMF

[2685] Daniel S. Holland and Jon G. Sutinen. An empirical model of fleet dynamics in New England trawl fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):253–264, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-169.

Riera:1999:SDC

[2686] Joan L. Riera, John E. Schindler, and Tim K. Kratz. Seasonal dynamics of carbon dioxide and methane in two clear-water lakes and two bog lakes in northern Wisconsin, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):265–274, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-182.

Buckel:1999:IAB

[2687] Jeffrey A. Buckel, David O. Conover, Nancy D. Steinberg, and Kim A. McKown. Impact of age-0 bluefish (Pomatomus saltatrix) predation on age-0 fishes in the Hudson River estuary: evidence for densitydependent loss of juvenile striped bass (Morone saxatilis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):275–287, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-173.

Gillis:1999:BIR

[2688] D. M. Gillis. Behavioral inferences from regulatory observer data: catch rate variation in the Scotian Shelf silver hake (Merluccius bilinearis) fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):288–296, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-166.

Beacham:1999:AMD

[2689] Terry D. Beacham and Chris C. Wood. Application of microsatellite DNA variation to estimation of stock composition and escapement of Nass River sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):297–310, February 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-167.

Gal:1999:LDV

[2690] Gideon Gal, Ellis R. Loew, Lars G. Rudstam, and Ali M. Mohammadian. Light and diel vertical migration: spectral sensitivity and light avoidance by Mysis relicta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (2):311–322, February 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-174.

Hebert:1999:STD

[2691] Craig E. Hebert, J. Laird Shutt, Keith A. Hobson, and D. V. Chip Weseloh. Spatial and temporal differences in the diet of Great Lakes herring gulls (Larus argentatus): evidence from stable isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):323–338, February 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-189.

Garcia:1999:IWC

[2692] Edenise Garcia and Richard Carignan. Impact of wildfire and clearcutting in the boreal forest on methyl mercury in zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(2):339–345, February 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-164.

Marshall:1999:EIV

[2693] C. Tara Marshall and Kenneth T. Frank. The effect of interannual variation in growth and condition on haddock recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):347–355, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-019.

Marshall:1999:IDD

[2694] C. Tara Marshall and Kenneth T. Frank. Implications of densitydependent juvenile growth for compensatory recruitment regulation of haddock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):356–363, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-018.

Willette:1999:PFM

[2695] T. M. Willette, R. T. Cooney, and K. Hyer. Predator foraging mode shifts affecting mortality of juvenile fishes during the subarctic spring bloom. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):364–376, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-185.

Montgomery:1999:CTS

[2696] David R. Montgomery, Eric M. Beamer, George R. Pess, and Thomas P. Quinn. Channel type and salmonid spawning distribution and abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):377–387, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-181.

Sand-Jensen:1999:SGD

[2697] Kaj Sand-Jensen, Tenna Riis, Stiig Markager, and Warwick F. Vincent. Slow growth and decomposition of mosses in Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):388–393, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-184.

Gaertner:1999:IFB

[2698] Daniel Gaertner, Mauricio Pagavino, and Jesus Marcano. Influence of fishers’ behaviour on the catchability of surface tuna schools in the Venezuelan purse-seiner fishery in the Caribbean Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):394–406, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-191.

Richards:1999:NCD

[2699] Jeff G. Richards, B. Kent Burnison, and Richard C. Playle. Natural and commercial dissolved organic matter protects against the physiological effects of a combined cadmium and copper exposure on rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (3):407–418, March 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-187.

Marty:1999:HAP

[2700] Gary D. Marty, Mark S. Okihiro, Evelyn D. Brown, David Hanes, and David E. Hinton. Histopathology of adult Pacific herring in Prince William Sound, Alaska, after the Exxon Valdez oil spill. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):419–426, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-178.

Currie:1999:AST

[2701] David J. Currie, Pam Dilworth-Christie, and François Chapleau. Assessing the strength of top-down influences on plankton abundance in unmanipulated lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (3):427–436, March 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-190.

Maravelias:1999:HSC

[2702] Christos D. Maravelias. Habitat selection and clustering of a pelagic fish: effects of topography and bathymetry on species dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):437–450, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-176.

Levine:1999:INP

[2703] S. N. Levine and D. W. Schindler. Influence of nitrogen to phosphorus supply ratios and physicochemical conditions on cyanobacteria and phytoplankton species composition in the Experimental Lakes Area, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):451–466, March 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-183.

Stewart:1999:ACF

[2704] A. R. Stewart. Accumulation of Cd by a freshwater mussel (Pyganodon grandis) is reduced in the presence of Cu, Zn, Pb, and Ni. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):467–478, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-180.

Gregory:1999:IBI

[2705] T. Ryan Gregory and Chris M. Wood. Interactions between individual feeding behaviour, growth, and swimming performance in juvenile rainbow trout (Oncorhynchus mykiss) fed different rations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):479–486, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-186.

Chafiq:1999:IAS

[2706] Mohamed Chafiq, Janine Gibert, and Cécile Claret. Interactions among sediments, organic matter, and microbial activity in the hyporheic zone of an intermittent stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (3):487–495, March 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-208.

Guyette:1999:ACC

[2707] Richard P. Guyette and William G. Cole. Age characteristics of coarse woody debris (Pinus strobus) in a lake littoral zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):496–505, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-177.

Beamish:1999:RDR

[2708] R. J. Beamish, G. A. McFarlane, and R. E. Thomson. Recent declines in the recreational catch of Coho salmon (Oncorhynchus kisutch) in the Strait of Georgia are related to climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(3):506–515, March 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-195.

Beamish:1999:RCN

[2709] R. J. Beamish, D. J. Noakes, G. A. McFarlane, L. Klyashtorin, V. V. Ivanov, and V. Kurashov. The regime concept and natural trends in the production of Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (3):516–526, March 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-200.

Kershner:1999:MSV

[2710] Mark W. Kershner, Denise M. Schael, Roger L. Knight, Roy A. Stein, and Elizabeth A. Marschall. Modeling sources of variation for growth and predatory demand of Lake Erie walleye (Stizostedion vitreum), 1986– 1995. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):527–538, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-193.

Currie:1999:IES

[2711] David R. Currie and Gregory D. Parry. Impacts and efficiency of scallop dredging on different soft substrates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):539–550, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-194.

Schwarz:1999:EET

[2712] Carl James Schwarz and Wayne T. Stobo. Estimation and effects of tagmisread rates in capture-recapture studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):551–559, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-196.

Chetelat:1999:PBC

[2713] J. Chételat, F. R. Pick, A. Morin, and P. B. Hamilton. Periphyton biomass and community composition in rivers of different nutrient status. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):560–569, April 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-197.

Gallagher:1999:RBC

[2714] Sean P. Gallagher and Mark F. Gard. Relationship between Chinook salmon (Oncorhynchus tshawytscha) redd densities and PHABSIMpredicted habitat in the Merced and Lower American rivers, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):570–577, April 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-198.

Hard:1999:ASV

[2715] Jeffrey J. Hard and William R. Heard. Analysis of straying variation in Alaskan hatchery Chinook salmon (Oncorhynchus tshawytscha) following transplantation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (4):578–589, April 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f98-199.

Fry:1999:TPI

[2716] Brian Fry, Patricia L. Mumford, Franklin Tam, Don D. Fox, Gary L. Warren, Karl E. Havens, and Alan D. Steinman. Trophic position and individual feeding histories of fish from Lake Okeechobee, Florida. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):590–600, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-204.

Walters:1999:MSA

[2717] Carl J. Walters and Ramón Bonfil. Multispecies spatial assessment models for the British Columbia groundfish trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):601–628, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-205.

Kuikka:1999:MED

[2718] S. Kuikka, M. Hildén, H. Gislason, S. Hansson, H. Sparholt, and O. Varis. Modeling environmentally driven uncertainties in Baltic cod (Gadus morhua) management by Bayesian influence diagrams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):629–641, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-206.

Foster:1999:HPA

[2719] Eugene P. Foster and Lawrence R. Curtis. 2,2’,4,4’,5,5’- and 3,3’,4,4’,5,5’- hexachlorobiphenyl pretreatments alter the biliary excretion of a challenge dose of 7,12-[^{3}H]dimethylbenz[a]anthracene in rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (4):642–649, April 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-065.

Jeremiason:1999:ARP

[2720] Jeff D. Jeremiason, Steven J. Eisenreich, and Michael J. Paterson. Accumulation and recycling of PCBs and PAHs in artificially eutrophied Lake 227. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):650–660, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f98-192.

Emmons:1999:ACM

[2721] Edward E. Emmons, Martin J. Jennings, and Clayton Edwards. An alternative classification method for northern Wisconsin lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):661–669, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-008.

Allin:1999:BME

[2722] C. J. Allin and R. W. Wilson. Behavioural and metabolic effects of chronic exposure to sublethal aluminum in acidic soft water in juvenile rainbow trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):670–678, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-002.

Gonzalez:1999:MUA

[2723] María J. González and Amy Downing. Mechanisms underlying amphipod responses to zebra mussel (Dreissena polymorpha) invasion and implications for fish-amphipod interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):679–685, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-211.

Mackinson:1999:NAA

[2724] Steven Mackinson, Marcelo Vasconcellos, and Nathaniel Newlands. A new approach to the analysis of stock-recruitment relationships: “modelfree estimation” using fuzzy logic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):686–699, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f98-215.

Kolok:1999:IVP

[2725] A. S. Kolok. Interindividual variation in the prolonged locomotor performance of ectothermic vertebrates: a comparison of fish and herpetofaunal methodologies and a brief review of the recent fish literature. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):700–710, April 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-026.

Pearson:1999:WDP

[2726] W. H. Pearson, R. A. Elston, R. W. Bienert, A. S. Drum, and L. D. Antrim. Why did the Prince William Sound, Alaska, Pacific herring (Clupea pallasi) fisheries collapse in 1993 and 1994? Review of hypotheses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(4):711–737, April 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f98-207.

Anonymous:1999:CTR

[2727] Anonymous. Concordance of taxonomic richness patterns across multiple assemblages in lakes of the northeastern United States. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):739–747, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-214.

Allen:1999:GMR

[2728] Yvonne C. Allen, Bruce A. Thompson, and Charles W. Ramcharan. Growth and mortality rates of the zebra mussel, Dreissena polymorpha, in the Lower Mississippi River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):748–759, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f98-212.

Morrison:1999:MMM

[2729] Kenneth A. Morrison and Carl J. Watras. Mercury and methyl mercury in freshwater seston: direct determination at picogram per litre levels by dual filtration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (5):760–766, May 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-029.

Wold:1999:ESC

[2730] Amy K. Fisher Wold and Anne E. Hershey. Effects of salmon carcass decomposition on biofilm growth and wood decomposition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):767–773, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-030.

Wang:1999:CSM

[2731] Dacheng Wang, Yves Couillard, G. C. Campbell, and Pierre Jolicoeur. Changes in subcellular metal partitioning in the gills of freshwater bivalves (Pyganodon grandis) living along an environmental cadmium gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):774–784, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-003.

Benoit:1999:IRT

[2732] Hugues P. Benoît and Pierre Pepin. Interaction of rearing temperature and maternal influence on egg development rates and larval size at hatch in yellowtail flounder (Pleuronectes ferrugineus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):785–794, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-213.

Oliveira:1999:LHC

[2733] Kenneth Oliveira. Life history characteristics and strategies of the American eel, Anguilla rostrata. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (5):795–802, May 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-001.

Rocha-Olivares:1999:EOC

[2734] Axayácatl Rocha-Olivares and Russell D. Vetter. Effects of oceanographic circulation on the gene flow, genetic structure, and phylogeography of the rosethorn rockfish (Sebastes helvomaculatus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):803–813, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-004.

Beyers:1999:EPC

[2735] Daniel W. Beyers, James A. Rice, William H. Clements, and Catherine J. Henry. Estimating physiological cost of chemical exposure: integrating energetics and stress to quantify toxic effects in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):814–822, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-006.

Beyers:1999:EBS

[2736] Daniel W. Beyers, James A. Rice, and William H. Clements. Evaluating biological significance of chemical exposure to fish using a bioenergeticsbased stressor-response model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):823–829, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f99-005.

Berejikian:1999:CAS

[2737] Barry A. Berejikian, R. Jan F. Smith, E. Paul Tezak, Steven L. Schroder, and Curtis M. Knudsen. Chemical alarm signals and complex hatchery rearing habitats affect antipredator behavior and survival of Chinook salmon (Oncorhynchus tshawytscha) juveniles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):830–838, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-010.

Bourque:1999:ETC

[2738] Marie-Claude Bourque, Paul H. LeBlond, and Patrick F. Cummins. Effects of tidal currents on Pacific salmon migration: results from a fineresolution coastal model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (5):839–846, May 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-015.

King:1999:SCT

[2739] Jacquelynne R. King, Brian J. Shuter, and Ann P. Zimmerman. Signals of climate trends and extreme events in the thermal stratification pattern of multibasin Lake Opeongo, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):847–852, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-020.

Kallio-Nyberg:1999:ESS

[2740] Irma Kallio-Nyberg, Heikki Peltonen, and Hannu Rita. Effects of stockspecific and environmental factors on the feeding migration of Atlantic salmon (Salmo salar) in the Baltic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):853–861, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-022.

Lake:1999:AES

[2741] Randal G. Lake and Scott G. Hinch. Acute effects of suspended sediment angularity on juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):862–867, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-024.

Xiao:1999:EFN

[2742] Yongshun Xiao, John D. Stevens, and Grant J. West. Estimation of fishing and natural mortalities from tag experiments with exact or grouped times at liberty. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):868–874, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f99-
007.

Tucker:1999:UMC

[2743] S. Tucker and J. B. Rasmussen. Using ^{137}Cs to measure and compare bioenergetic budgets of juvenile Atlantic salmon (Salmo salar) and brook trout (Salvelinus fontinalis) in the field. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):875–887, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-025.

Hilborn:1999:ESE

[2744] Ray Hilborn, Brian G. Bue, and Samuel Sharr. Estimating spawning escapements from periodic counts: a comparison of methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):888–896, May 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-013.

Miller:1999:CMD

[2745] Robert J. Miller. Courage and the management of developing fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(5):897–905, May 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-011.

Fletcher:1999:SDS

[2746] W. J. Fletcher and Neil R. Sumner. Spatial distribution of sardine (Sardinops sagax) eggs and larvae: an application of geostatistics and resampling to survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (6):907–914, June 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-012.

Hauer:1999:LWD

[2747] F. Richard Hauer, Geoffrey C. Poole, John T. Gangemi, and Colden V. Baxter. Large woody debris in bull trout (Salvelinus confluentus) spawning streams of logged and wilderness watersheds in northwest Montana. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):915–924, June 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-014.

Sibert:1999:ADR

[2748] John R. Sibert, John Hampton, David A. Fournier, and Peter J. Bills. An advection–diffusion–reaction model for the estimation of fish movement parameters from tagging data, with application to skipjack tuna (Katsuwonus pelamis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):925–938, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f99-
017.

Roper:1999:EAC

[2749] Brett B. Roper and Dennis L. Scarnecchia. Emigration of age-0 Chinook salmon (Oncorhynchus tshawytscha) smolts from the upper South Umpqua River basin, Oregon, U.S.A. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):939–946, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-023.

Nolan:1999:ILN

[2750] D. T. Nolan, P. Reilly, and S. E. Wendelaar Bonga. Infection with low numbers of the sea louse Lepeophtheirus salmonis induces stressrelated effects in postsmolt Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):947–959, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-021.

Smith:1999:PAD

[2751] Barry D. Smith. A probabilistic analysis of decision-making about trip duration by Strait of Georgia sport anglers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):960–972, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-027.

Noto:1999:PDJ

[2752] Masayuki Noto and Ichiro Yasuda. Population decline of the Japanese sardine, Sardinops melanostictus, in relation to sea surface temperature in the Kuroshio Extension. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (6):973–983, June 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-028.

Dagorn:1999:TTA

[2753] Laurent Dagorn and Pierre Fréon. Tropical tuna associated with floating objects: a simulation study of the meeting point hypothesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):984–993, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f98-209.

Mallet:1999:GMA

[2754] J. P. Mallet, S. Charles, H. Persat, and P. Auger. Growth modelling in accordance with daily water temperature in European grayling (Thymallus thymallus L.). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):994–1000, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f99-
031.

Ostrovsky:1999:OMP

[2755] I. Ostrovsky and Y. Z. Yacobi. Organic matter and pigments in surface sediments: possible mechanisms of their horizontal distributions in a stratified lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6): 1001–1010, June 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-032.

Johnston:1999:RRT

[2756] N. T. Johnston, M. D. Stamford, K. I. Ashley, and K. Tsumura. Responses of rainbow trout (Oncorhynchus mykiss) and their prey to inorganic fertilization of an oligotrophic montane lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1011–1025, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-034.

Prajneshu:1999:BGM

[2757] Prajneshu and R. Venugopalan. von Bertalanffy growth model in a random environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1026– 1030, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-035.

Schaller:1999:CPP

[2758] Howard A. Schaller, Charles E. Petrosky, and Olaf P. Langness. Contrasting patterns of productivity and survival rates for stream-type Chinook salmon (Oncorhynchus tshawytscha) populations of the Snake and Columbia rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1031– 1045, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/

f99-037. See comments [2911].

Pyper:1999:PCL

[2759] Brian J. Pyper, Randall M. Peterman, Michael F. Lapointe, and Carl J. Walters. Patterns of covariation in length and age at maturity of British Columbia and Alaska sockeye salmon (Oncorhynchus nerka) stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1046–1057, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-038.

Beland:1999:ESU

[2760] France Béland, Howard I. Browman, Carolina Alonso Rodriguez, and Jean-François St-Pierre. Effect of solar ultraviolet radiation (280–400 nm) on the eggs and larvae of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1058–1067, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-039.

Brannas:1999:FSN

[2761] Eva Brännäs and Torleif Eriksson. Floating, switching, or nonswitching as different behaviours when Arctic char (Salvelinus alpinus) are visiting two feeding tanks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (6):1068–1077, June 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-042.

Meyer:1999:BBS

[2762] Renate Meyer and Russell B. Millar. BUGS in Bayesian stock assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1078–1087, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-043.

Comeau:1999:FDE

[2763] Michel Comeau, Michel Starr, Gérard Y. Conan, Guy Robichaud, and Jean-Claude Therriault. Fecundity and duration of egg incubation for multiparous female snow crabs (Chionoecetes opilio) in the fjord of Bonne Bay, Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6): 1088–1095, June 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-044.

Fulton:1999:SSI

[2764] Elizabeth Fulton, David Kault, Bruce Mapstone, and Marcus Sheaves. Spawning season influences on commercial catch rates: computer simulations and Plectropomus leopardus, a case in point. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1096–1108, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-046.

Hall:1999:LSR

[2765] Roland I. Hall, Peter R. Leavitt, Aruna S. Dixit, Roberto Quinlan, and John P. Smol. Limnological succession in reservoirs: a paleolimnological comparison of two methods of reservoir formation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1109–1121, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-047.

Aubin-Horth:1999:CAR

[2766] N. Aubin-Horth, J. Gingras, and D. Boisclair. Comparison of activity rates of 1+ yellow perch (Perca flavescens) from populations of contrasting growth rates using underwater video observations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1122–1132, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-048.

Rooper:1999:HFC

[2767] Christopher N. Rooper, Lewis J. Haldorson, and Terrance J. Quinn II. Habitat factors controlling Pacific herring (Clupea pallasi) egg loss in Prince William Sound, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(6):1133–1142, June 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-050.

Roy:1999:RIW

[2768] Isabelle Roy and Landis Hare. Relative importance of water and food as cadmium sources to the predatory insect Sialis velata (Megaloptera). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1143–1149, July 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-036.

Reckhow:1999:WQP

[2769] Kenneth H. Reckhow. Water quality prediction and probability network models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1150–1158, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-040.

Leonard:1999:EMD

[2770] Jill B. K. Leonard and Stephen D. McCormick. Effects of migration distance on whole-body and tissue-specific energy use in American shad (Alosa sapidissima). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (7):1159–1171, July 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-041.

Unwin:1999:ESP

[2771] M. J. Unwin, M. T. Kinnison, and T. P. Quinn. Exceptions to semelparity: postmaturation survival, morphology, and energetics of male Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1172–1181, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-045.

Simard:1999:RKA

[2772] Yvan Simard and Diane Lavoie. The rich krill aggregation of the Saguenay–St. Lawrence Marine Park: hydroacoustic and geostatistical biomass estimates, structure, variability, and significance for whales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1182–1197, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-063.

Neff:1999:MES

[2773] Bryan D. Neff, Peng Fu, and Mart R. Gross. Microsatellite evolution in sunfish (Centrarchidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (7):1198–1205, July 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-068.

Reinhardt:1999:PRB

[2774] Ulrich G. Reinhardt. Predation risk breaks size-dependent dominance in juvenile Coho salmon (Oncorhynchus kisutch) and provides growth opportunities for risk-prone individuals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1206–1212, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-064.

vandenHeuvel:1999:EOSa

[2775] M. R. van den Heuvel, M. Power, M. D. MacKinnon, T. Van Meer, E. P. Dobson, and D. G. Dixon. Effects of oil sands related aquatic reclamation on yellow perch (Perca flavescens). I. Water quality characteristics and yellow perch physiological and population responses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1213–1225, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-062.

vandenHeuvel:1999:EOSb

[2776] M. R. van den Heuvel, M. Power, M. D. MacKinnon, and D. G. Dixon. Effects of oil sands related aquatic reclamation on yellow perch (Perca flavescens). II. Chemical and biochemical indicators of exposure to oil sands related waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (7):1226–1233, July 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-061.

Gould:1999:USE

[2777] W. R. Gould, L. A. Stefanski, and K. H. Pollock. Use of simulation– extrapolation estimation in catch–effort analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1234–1240, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-052.

Piet:1999:RPA

[2778] G. J. Piet, J. S. Pet, W. A. H. P. Guruge, J. Vijverberg, and W. L. T. Van Densen. Resource partitioning along three niche dimensions in a size-structured tropical fish assemblage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1241–1254, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-033.

Hearn:1999:ETR

[2779] William S. Hearn, Thomas Polacheck, Kenneth H. Pollock, and Wade Whitelaw. Estimation of tag reporting rates in agestructured multicomponent fisheries where one component has observers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1255–1265, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-059.

Campbell:1999:MSD

[2780] H. F. Campbell and A. J. Hand. Modeling the spatial dynamics of the U.S. purse-seine fleet operating in the western Pacific tuna fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1266–1277, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-009.

Millar:1999:MSS

[2781] Russell B. Millar, Brian H. McArdle, and Shelton J. Harley. Modeling the size of snapper (Pagrus auratus) using temperature-modified growth curves. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1278–1284, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-057.

Lundvall:1999:SDP

[2782] David Lundvall, Richard Svanbäck, Lennart Persson, and Pär Byström. Size-dependent predation in piscivores: interactions between predator foraging and prey avoidance abilities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1285–1292, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-058.

Vogel:1999:ELP

[2783] Jason L. Vogel and David A. Beauchamp. Effects of light, prey size, and turbidity on reaction distances of lake trout (Salvelinus namaycush) to salmonid prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1293– 1297, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-071.

Bult:1999:DDH

[2784] Tammo P. Bult, Stephen C. Riley, Richard L. Haedrich, R. John Gibson, and Jan Heggenes. Density-dependent habitat selection by juvenile Atlantic salmon (Salmo salar) in experimental riverine habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1298–1306, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-074.

Hovgaard:1999:GSN

[2785] Holger Hovgård, Hans Lassen, Niels Madsen, Thomas Moth Poulsen, and David Wileman. Gillnet selectivity for North Sea Atlantic cod (Gadus morhua): model ambiguity and data quality are related. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1307–1316, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-070.

Welch:1999:ALB

[2786] David W. Welch and J. Paige Eveson. An assessment of light-based geoposition estimates from archival tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(7):1317–1327, July 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-049.

Andrews:1999:AIE

[2787] Allen H. Andrews, Kenneth H. Coale, Jocelyn L. Nowicki, Craig Lundstrom, Zenon Palacz, Erica J. Burton, and Gregor M. Cailliet. Application of an ion-exchange separation technique and thermal ionization mass spectrometry to ^{226}Ra determination in otoliths for radiometric age determination of long-lived fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1329–1338, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-053.

Andrews:1999:AGP

[2788] Allen H. Andrews, Gregor M. Cailliet, and Kenneth H. Coale. Age and growth of the Pacific grenadier (Coryphaenoides acrolepis) with age estimate validation using an improved radiometric ageing technique. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1339–1350, August 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-054.

Smith:1999:AUR

[2789] R. E. H. Smith, J. A. Furgal, M. N. Charlton, B. M. Greenberg, V. Hiriart, and C. Marwood. Attenuation of ultraviolet radiation in a large lake with low dissolved organic matter concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1351–1361, August 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-051.

Reid:1999:ITP

[2790] Scott M. Reid, Michael G. Fox, and Thomas H. Whillans. Influence of turbidity on piscivory in largemouth bass (Micropterus salmoides). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1362–1369, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-056.

Jorgensen:1999:ETP

[2791] Even H. Jørgensen and Johannes Wolkers. Effect of temperature on the P4501A response in winter- and summer-acclimated Arctic char (Salvelinus alpinus) after oral benzo[a]pyrene exposure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1370–1375, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-075.

Gaggiotti:1999:ELH

[2792] Oscar E. Gaggiotti and Russel D. Vetter. Effect of life history strategy, environmental variability, and overexploitation on the genetic diversity of pelagic fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (8):1376–1388, August 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-060.

Morrison:1999:AFW

[2793] Heather A. Morrison, D. Michael Whittle, Chris D. Metcalfe, and Arthur J. Niimi. Application of a food web bioaccumulation model for the prediction of polychlorinated biphenyl, dioxin, and furan congener concentrations in Lake Ontario aquatic biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1389–1400, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-066.

Wilhelm:1999:EGL

[2794] Frank M. Wilhelm and David W. Schindler. Effects of Gammarus lacustris (Crustacea: Amphipoda) on plankton community structure in an alpine lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1401–1408, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-067.

Fabrizio:1999:MDD

[2795] Mary C. Fabrizio, James D. Nichols, James E. Hines, Bruce L. Swanson, and Stephen T. Schram. Modeling data from double-tagging experiments to estimate heterogeneous rates of tag shedding in lake trout (Salvelinus namaycush). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8): 1409–1419, August 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-069.

Maki-Petays:1999:DJB

[2796] Aki Mäki-Petäys, Timo Muotka, and Ari Huusko. Densities of juvenile brown trout (Salmo trutta) in two subarctic rivers: assessing the predictive capability of habitat preference indices. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1420–1427, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-072.

Doyon:1999:RAH

[2797] Christian Doyon, Réjean Fortin, and Philip A. Spear. Retinoic acid hydroxylation and teratogenesis in lake sturgeon (Acipenser fulvescens) from the St. Lawrence River and abitibi region, Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1428–1436, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-086.

Pinel-Alloul:1999:LSS

[2798] Bernadette Pinel-Alloul, Catherine Guay, Nadine Angeli, Pierre Legendre, Pierre Dutilleul, Gérard Balvay, Daniel Gerdeaux, and Jean Guillard. Large-scale spatial heterogeneity of macrozooplankton in Lake of Geneva. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1437–1451, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f99-
055.

Eisenmann:1999:STH

[2799] Heinrich Eisenmann, Peter Burgherr, and Elisabeth I. Meyer. Spatial and temporal heterogeneity of an epilithic streambed community in relation to the habitat templet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (8):1452–1460, August 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-092.

Kohler:1999:IOA

[2800] Stephan Köhler, Jakub Hruška, and Kevin Bishop. Influence of organic acid site density on pH modeling of Swedish lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1461–1470, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-097.

Marr:1999:BAT

[2801] J. C. A. Marr, J. Lipton, D. Cacela, J. A. Hansen, J. S. Meyer, and H. L. Bergman. Bioavailability and acute toxicity of copper to rainbow trout (Oncorhynchus mykiss) in the presence of organic acids simulating natural dissolved organic carbon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (8):1471–1483, August 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-089.

MacRae:1999:SOA

[2802] Russell K. MacRae, Ann S. Maest, and Joseph S. Meyer. Selection of an organic acid analogue of dissolved organic matter for use in toxicity testing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1484–1493, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/f99-
090.

Borgmann:1999:ATL

[2803] Uwe Borgmann and Warren P. Norwood. Assessing the toxicity of lead in sediments to Hyalella azteca: the significance of bioaccumulation and dissolved metal. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8): 1494–1503, August 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-073.

Phillips:1999:CPA

[2804] Ruth B. Phillips, Linda I. Gudex, Kathleen M. Westrich, and Alfred DeCicco. Combined phylogenetic analysis of ribosomal ITS1 sequences and new chromosome data supports three subgroups of Dolly Varden char (Salvelinus malma). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (8):1504–1511, August 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-103.

Routledge:1999:CFS

[2805] Richard D. Routledge and James R. Irvine. Chance fluctuations and the survival of small salmon stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(8):1512–1519, August 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-093.

Shelton:1999:SDD

[2806] Peter A. Shelton and Brian P. Healey. Should depensation be dismissed as a possible explanation for the lack of recovery of the northern cod (Gadus morhua) stock? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1521– 1524, September 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-124.

Chen:1999:IAD

[2807] Y. Chen and D. Fournier. Impacts of atypical data on Bayesian inference and robust Bayesian approach in fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1525–1533, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-076.

Sanderson:1999:CDY

[2808] Beth L. Sanderson, Thomas R. Hrabik, John J. Magnuson, and David M. Post. Cyclic dynamics of a yellow perch (Perca flavescens) population in an oligotrophic lake: evidence for the role of intraspecific interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1534–1542, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-077.

Whalen:1999:NHU

[2809] Kevin G. Whalen and Donna L. Parrish. Nocturnal habitat use of Atlantic salmon parr in winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (9):1543–1550, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f99-078.

Ackerman:1999:EVF

[2810] Josef Daniel Ackerman. Effect of velocity on the filter feeding of dreissenid mussels (Dreissena polymorpha and Dreissena bugensis): implications for trophic dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (9):1551–1561, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f99-079.

Paradis:1999:DES

[2811] A. R. Paradis, M. Pépin, and P. Pepin. Disentangling the effects of sizedependent encounter and susceptibility to predation with an individualbased model for fish larvae. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (9):1562–1575, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f99-080.

Knapp:1999:IPP

[2812] Roland A. Knapp and Haiganoush K. Preisler. Is it possible to predict habitat use by spawning salmonids? A test using California golden trout (Oncorhynchus mykiss aguabonita). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1576–1584, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-081.

Morita:1999:EMG

[2813] Kentaro Morita, Shoichiro Yamamoto, Yoshinobu Takashima, Takashi Matsuishi, Yasuji Kanno, and Kinya Nishimura. Effect of maternal growth history on egg number and size in wild white-spotted char (Salvelinus leucomaenis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (9):1585–1589, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f99-082.

Gallagher:1999:NME

[2814] M. Gallagher and C. P. Nolan. A novel method for the estimation of age and growth in rajids using caudal thorns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1590–1599, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-084.

Wipfli:1999:ISS

[2815] Mark S. Wipfli, John P. Hudson, Dominic T. Chaloner, and John P. Caouette. Influence of salmon spawner densities on stream productivity in Southeast Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1600– 1611, September 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-087.

Hutchings:1999:IGS

[2816] Jeffrey A. Hutchings. Influence of growth and survival costs of reproduction on Atlantic cod, Gadus morhua, population growth rate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1612–1623, September 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-088.

Lowell:1999:CEM

[2817] Richard B. Lowell and Joseph M. Culp. Cumulative effects of multiple effluent and low dissolved oxygen stressors on mayflies at cold temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1624–1630, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-091.

Lee:1999:DDR

[2818] P. A. Lee, R. Haase, S. J. de Mora, J.-P. Chanut, and M. Gosselin. Dimethylsulfoxide (DMSO) and related sulfur compounds in the Saguenay Fjord, Quebec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1631– 1638, September 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-094.

Parent:1999:RCDa

[2819] Serge Parent and Antoine Morin. Role of copepod-dominated meiofauna in the nitrification process of a cold marine mesocosm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1639–1648, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-096.

McCormick:1999:TRL

[2820] Stephen D. McCormick, Richard A. Cunjak, Brian Dempson, Michael F. O’Dea, and Judith B. Carey. Temperature-related loss of smolt characteristics in Atlantic salmon (Salmo salar) in the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1649–1658, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-099.

Laroche:1999:GSD

[2821] J. Laroche, J. D. Durand, Y. Bouvet, B. Guinand, and B. Brohon. Genetic structure and differentiation among populations of two cyprinids, Leuciscus cephalus and Rutilus rutilus, in a large European river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1659–1667, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-100.

Babcock:1999:FBA

[2822] Russ Babcock and John Keesing. Fertilization biology of the abalone Haliotis laevigata: laboratory and field studies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1668–1678, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-106.

Wilhelm:1999:CGL

[2823] Frank M. Wilhelm, Jeff J. Hudson, and David W. Schindler. Contribution of Gammarus lacustris to phosphorus recycling in a fishless alpine lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1679–1686, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-107.

Lewis:1999:IGS

[2824] David Bruce Lewis and John J. Magnuson. Intraspecific gastropod shell strength variation among north temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1687–1695, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-110.

Nash:1999:SAF

[2825] Caedmon H. Nash, John S. Richardson, and Scott G. Hinch. Spatial autocorrelation and fish production in freshwaters: a comment on Randall et al. (1995). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1696– 1699, September 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-105.

Young:1999:MDP

[2826] Kyle A. Young. Managing the decline of Pacific salmon: metapopulation theory and artificial recolonization as ecological mitigation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1700–1706, September 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-113.

Gordon:1999:OND

[2827] Donald C. Gordon, Kent D. Gilkinson, and Stephen J. Kerr. Obituary/nécrologie: Dr. Peter John Schwinghamer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(9):1707–1710, September 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-138.

Omair:1999:FOT

[2828] M. Omair, Henry A. Vanderploeg, David J. Jude, and Gary L. Fahnenstiel. First observations of tumor-like abnormalities (exophytic lesions) on Lake Michigan zooplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1711–1715, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-146.

Pyper:1999:RAA

[2829] Brian J. Pyper and Randall M. Peterman. Relationship among adult body length, abundance, and ocean temperature for British Columbia and Alaska sockeye salmon (Oncorhynchus nerka), 1967–1997. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1716–1720, October 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-167.

Clark:1999:EEN

[2830] William G. Clark. Effects of an erroneous natural mortality rate on a simple age-structured stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1721–1731, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-085.

Fowler:1999:ERP

[2831] G. M. Fowler and W. T. Stobo. Effects of release parameters on recovery rates of tagged groundfish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1732–1751, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-098.

LeBreton:1999:LSA

[2832] Gregory T. LeBreton, F. William H. Beamish, and Robert G. Wallace. Lake sturgeon (Acipenser fulvescens) growth chronologies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1752–1756, October 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-101.

Perrein-Ettajani:1999:EMA

[2833] H. Perrein-Ettajani, J. C. Amiard, J. Haure, and C. Renaud. Effets des métaux (Ag, Cd, Cu) sur la composition biochimique et compartimentation de ces métaux chez deux microalgues Skeletonema costatum et Tetraselmis suecica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10): 1757–1765, October 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-102.

Koljonen:1999:PLD

[2834] Marja-Liisa Koljonen, Håkan Jansson, Tiit Paaver, Oleg Vasin, and Jarmo Koskiniemi. Phylogeographic lineages and differentiation pattern of Atlantic salmon (Salmo salar) in the Baltic Sea with management implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1766–1780, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-104.

Crosbie:1999:PLU

[2835] Barb Crosbie and Patricia Chow-Fraser. Percentage land use in the watershed determines the water and sediment quality of 22 marshes in the Great Lakes basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10): 1781–1791, October 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-109.

Vanier:1999:ERB

[2836] Claire Vanier, Dolores Planas, and Michel Sylvestre. Empirical relationships between polychlorinated biphenyls in sediments and submerged rooted macrophytes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10): 1792–1800, October 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-111.

Holmes:1999:SPM

[2837] Robert M. Holmes, Alain Aminot, Roger Kérouel, Bethanie A. Hooker, and Bruce J. Peterson. A simple and precise method for measuring ammonium in marine and freshwater ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1801–1808, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-128.

Morales-Baquero:1999:NPR

[2838] Rafael Morales-Baquero, Presentación Carrillo, Isabel Reche, and Pedro Sánchez-Castillo. Nitrogen–phosphorus relationship in high mountain lakes: effects of the size of catchment basins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1809–1817, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/f99-130. See erratum [2881].

Dutil:1999:GCE

[2839] J.-D. Dutil, M. Castonguay, D. Gilbert, and D. Gascon. Growth, condition, and environmental relationships in Atlantic cod (Gadus morhua) in the northern Gulf of St. Lawrence and implications for management strategies in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1818–1831, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-140.

Marchand:1999:CCM

[2840] Caroline Marchand, Yvan Simard, and Yves Gratton. Concentration of capelin (Mallotus villosus) in tidal upwelling fronts at the head of the Laurentian Channel in the St. Lawrence estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1832–1848, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-117.

Narimatsu:1999:SDD

[2841] Yoji Narimatsu and Hiroyuki Munehara. Spawn date dependent survival and growth in the early life stages of Hypoptychus dybowskii (Gasterosteiformes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1849– 1855, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-

7533 (electronic). URL https://cdnsciencepub.com/doi/full/10.
1139/f99-118.

Chang:1999:MNC

[2842] Cecily C. Y. Chang, James Langston, Mike Riggs, Donald H. Campbell, Steven R. Silva, and Carol Kendall. A method for nitrate collection for δ ^{15}N and δ ^{18}O analysis from waters with low nitrate concentrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1856–1864, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-126.

Wahlstrom:1999:PPC

[2843] Eva Wahlström and Erika Westman. Planktivory by the predacious cladoceran Bythotrephes longimanus: effects on zooplankton size structure and abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10): 1865–1872, October 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-123.

Campana:1999:MMO

[2844] S. E. Campana, G. A. Chouinard, J. M. Hanson, and A. Fréchet. Mixing and migration of overwintering Atlantic cod (Gadus morhua) stocks near the mouth of the Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1873–1881, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-083.

Scott:1999:PEM

[2845] Beth Scott, Gudrun Marteinsdottir, and Peter Wright. Potential effects of maternal factors on spawning stock–recruitment relationships under varying fishing pressure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10): 1882–1890, October 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-125.

Seppa:1999:NSR

[2846] Teija Seppä, Nina Peuhkuri, Heikki Hirvonen, Anssi Laurila, Jorma Piironen, and Esa Ranta. Narrow size regime among individuals favors rapid growth in Arctic char (Salvelinus alpinus) juveniles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1891–1897, October 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-121.

Brown:1999:EDW

[2847] Randy Brown and Kenneth P. Severin. Elemental distribution within polymorphic inconnu (Stenodus leucichthys) otoliths is affected by crystal structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1898–1903, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-127.

Sheehy:1999:NPG

[2848] M. R. J. Sheehy, R. C. A. Bannister, J. F. Wickins, and P. M. J. Shelton. New perspectives on the growth and longevity of the European lobster (Homarus gammarus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10): 1904–1915, October 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-116.

Landry:1999:SOL

[2849] François Landry, John R. Post, and Eric A. Parkinson. Spatial ontogeny of lentic age-0 rainbow trout, Oncorhynchus mykiss: whole-lake manipulations of population size structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1916–1928, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-129.

Imai:1999:EIL

[2850] Akio Imai, Takehiko Fukushima, and Kazuo Matsushige. Effects of iron limitation and aquatic humic substances on the growth of Microcystis aeruginosa. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1929–1937, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-131.

Parent:1999:RCDb

[2851] Serge Parent and Antoine Morin. The role of copepod-dominated meiofauna in the mineralization of organic matter in a cold marine mesocosm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1938–1948, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-160.

Neis:1999:FAW

[2852] Barbara Neis, David C. Schneider, Lawrence Felt, Richard L. Haedrich, Johanne Fischer, and Jeffrey A. Hutchings. Fisheries assessment: what can be learned from interviewing resource users? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1949–1963, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-115.

Cuddington:1999:IBM

[2853] K. Cuddington and P. R. Leavitt. An individual-based model of pigment flux in lakes: implications for organic biogeochemistry and paleoecology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(10):1964–1977, October 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-108.

Veinott:1999:CSP

[2854] Geoff Veinott, Tom Northcote, Marvin Rosenau, and R. Douglas Evans. Concentrations of strontium in the pectoral fin rays of the white sturgeon (Acipenser transmontanus) by laser ablation sampling–inductively coupled plasma–mass spectrometry as an indicator of marine migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):1981–1990, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-120.

Goddard:1999:PDA

[2855] Sally V. Goddard, Ming H. Kao, and Garth L. Fletcher. Population differences in antifreeze production cycles of juvenile Atlantic cod (Gadus morhua) reflect adaptations to overwintering environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):1991–1999, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-122.

Jaffe:1999:SED

[2856] Jules S. Jaffe, Mark D. Ohman, and Alex De Robertis. Sonar estimates of daytime activity levels of Euphausia pacifica in Saanich Inlet. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2000–2010, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-132.

Squires:1999:SSP

[2857] Dale Squires and James Kirkley. Skipper skill and panel data in fishing industries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2011– 2018, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-135.

Stead:1999:FCG

[2858] Selina M. Stead, Dominic F. Houlihan, H. Anne McLay, and Ray Johnstone. Food consumption and growth in maturing Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2019– 2028, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-136.

Allen:1999:CTC

[2859] Andrew P. Allen, Thomas R. Whittier, David P. Larsen, Philip R. Kaufmann, Raymond J. O’Connor, Robert M. Hughes, Richard S. Stemberger, Sushil S. Dixit, Ralph O. Brinkhurst, Alan T. Herlihy, and Steven G. Paulsen. Concordance of taxonomic composition patterns across multiple lake assemblages: effects of scale, body size, and land use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2029–2040, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-139.

Holomuzki:1999:IBD

[2860] Joseph R. Holomuzki, Robert W. Pillsbury, and Sapna B. Khandwala. Interplay between dispersal determinants of larval hydropsychid caddisflies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2041–2050, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-141.

Corten:1999:RSA

[2861] Ad Corten. The reappearance of spawning Atlantic herring (Clupea harengus) on Aberdeen Bank (North Sea) in 1983 and its relationship to environmental conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56 (11):2051–2061, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/f99-142.

Gao:1999:ICO

[2862] Y. W. Gao and R. J. Beamish. Isotopic composition of otoliths as a chemical tracer in population identification of sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2062–2068, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-145.

Planque:1999:TRA

[2863] B. Planque and T. Frédou. Temperature and the recruitment of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11): 2069–2077, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-114.

Kroglund:1999:WQR

[2864] Frode Kroglund and Magne Staurnes. Water quality requirements of smolting Atlantic salmon (Salmo salar) in limed acid rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2078–2086, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-119.

Birtwell:1999:SPS

[2865] I. K. Birtwell, R. Fink, D. Brand, R. Alexander, and C. D. McAllister. Survival of pink salmon (Oncorhynchus gorbuscha) fry to adulthood following a 10-day exposure to the aromatic hydrocarbon water-soluble fraction of crude oil and release to the Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2087–2098, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-134.

Otterlei:1999:TSD

[2866] Erling Otterlei, Gunnar Nyhammer, Arild Folkvord, and Sigurd O. Stefansson. Temperature- and size-dependent growth of larval and early juvenile Atlantic cod (Gadus morhua): a comparative study of Norwegian coastal cod and northeast Arctic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2099–2111, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-168.

Alsop:1999:IWC

[2867] Derek H. Alsop and Chris M. Wood. Influence of waterborne cations on zinc uptake and toxicity in rainbow trout, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2112–2119, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-156.

Pelletier:1999:CEJa

[2868] Émilien Pelletier, Behzad Mostajir, Suzanne Roy, Michel Gosselin, Yves Gratton, Jean-Pierre Chanut, Claude Belzile, Serge Demers, and Delphine Thibault. Crue éclair de juillet 1996 dans la région du Saguenay (Québec). 1. Impacts sur la colonne d’eau de la baie des Ha! Ha! et du fjord du Saguenay. (French) [Light flood in July 1996 in the Saguenay (Quebec) region. 1. Impacts on the water column of Ha! Ha! Bay and the Saguenay fjord]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11): 2120–2135, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-144.

Pelletier:1999:CEJb

[2869] Émilien Pelletier, Bruno Deflandre, Christian Nozais, Guglielmo Tita, Gaston Desrosiers, Jean-Pierre Gagné, and Alfonso Mucci. Crue éclair de juillet 1996 dans la région du Saguenay (Québec). 2. Impacts sur les sédiments et le biote de la baie des Ha! Ha! et du fjord du Saguenay. (French) [Light flood in July 1996 in the Saguenay (Quebec) region. 2. Impacts on sediment and biota in Ha! Ha! Bay and the Saguenay Fjord]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2136–2147, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-143.

Rutherford:1999:IBM

[2870] Edward S. Rutherford, Kenneth A. Rose, Edward L. Mills, John L. Forney, Christine M. Mayer, and Lars G. Rudstam. Individual-based model simulations of a zebra mussel (Dreissena polymorpha) induced energy shunt on walleye (Stizostedion vitreum) and yellow perch (Perca flavescens) populations in Oneida Lake, New York. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2148–2160, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-133.

Harvey:1999:ILW

[2871] Bret C. Harvey, Rodney J. Nakamoto, and Jason L. White. Influence of large woody debris and a bankfull flood on movement of adult resident coastal cutthroat trout (Oncorhynchus clarki) during fall and winter. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2161–2166, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-154.

Fry:1999:USI

[2872] Brian Fry. Using stable isotopes to monitor watershed influences on aquatic trophodynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11): 2167–2171, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-152.

Doucett:1999:ESM

[2873] R. R. Doucett, R. K. Booth, G. Power, and R. S. McKinley. Effects of the spawning migration on the nutritional status of anadromous Atlantic salmon (Salmo salar): insights from stable-isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2172–2180, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-147.

Bosley:1999:EPA

[2874] Keith L. Bosley and Sam C. Wainright. Effects of preservatives and acidification on the stable isotope ratios (^{15}N :^{14}N,^{13}C :^{12}C) of two species of marine animals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2181– 2185, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-153.

Owen:1999:SNC

[2875] J. S. Owen, M. J. Mitchell, and R. H. Michener. Stable nitrogen and carbon isotopic composition of seston and sediment in two Adirondack lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2186–2192, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-150.

Kidd:1999:ENP

[2876] Karen A. Kidd, Michael J. Paterson, Raymond H. Hesslein, Derek C. G. Muir, and Robert E. Hecky. Effects of northern pike (Esox lucius) additions on pollutant accumulation and food web structure, as determined by δ ^{13}C and δ ^{15}N, in a eutrophic and an oligotrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2193–2202, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-148.

Gorokhova:1999:ESV

[2877] Elena Gorokhova and Sture Hansson. An experimental study on variations in stable carbon and nitrogen isotope fractionation during growth of Mysis mixta and Neomysis integer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2203–2210, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-149.

Leggett:1999:BIC

[2878] M. F. Leggett, M. R. Servos, R. Hesslein, O. Johannsson, E. S. Millard, and D. G. Dixon. Biogeochemical influences on the carbon isotope signatures of Lake Ontario biota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2211–2218, November 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-151.

Grahl-Nielsen:1999:CFA

[2879] Otto Grahl-Nielsen. Comment: Fatty acid signatures and classification trees: new tools for investigating the foraging ecology of seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2219–2223, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-191. See [2226]

and reply [2880].

Smith:1999:RFA

[2880] Stephen J. Smith, Sara J. Iverson, and W. D. Bowen. Reply: Fatty acid signatures and classification trees: new tools for investigating the foraging ecology of seals. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11): 2224–2226, November 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/

10.1139/f99-192. See [2226, 2879].

Morales-Baquero:1999:ENP

[2881] Rafael Morales-Baquero, Presentación Carrillo, Isabel Reche, and Pedro Sánchez-Castillo. Erratum: Nitrogen–phosphorus relationship in high mountain lakes: effects of the size of catchment basins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(11):2227, November 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-244. See [2838].

Gislason:1999:RCP

[2882] Davíd Gíslason, Moira M. Ferguson, Skúli Skúlason, and Sigurdur S. Snorrason. Rapid and coupled phenotypic and genetic divergence in Icelandic Arctic char (Salvelinus alpinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2229–2234, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-245.

Tucker:1999:DPA

[2883] Strahan Tucker, Ivano Pazzia, David Rowan, and Joseph B. Rasmussen. Detecting pan-Atlantic migration in salmon (Salmo salar) using ^{137}Cs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2235–2239, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-267.

Lalonde:1999:MMC

[2884] Janick D. Lalonde, François Chapleau, Anik Brind’Amour, and Lara Louise Ridgway. Les métaux (Mn, Cu, Cd, Zn et Hg) des meuniers noirs (Catostomus commersoni) et des grands brochets (Esox lucius) de l’amont et de l’aval du barrage Moses–Saunders sur le fleuve Saint-Laurent près de Cornwall (Ontario) et Massena (New York). (French) [The metals (Mn, Cu, Cd, Zn and Hg) of black suckers (Catostomus commersoni) and northern pike (Esox lucius) from the upstream and downstream of the Moses–Saunders Dam on the St. Lawrence River near Cornwall (Ontario) and Massena (New York)]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2240–2247, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-137.

Wilson:1999:MDS

[2885] A. B. Wilson, K.-A. Naish, and E. G. Boulding. Multiple dispersal strategies of the invasive quagga mussel (Dreissena bugensis) as revealed by microsatellite analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12): 2248–2261, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-162.

Kann:1999:EPE

[2886] Jacob Kann and Val H. Smith. Estimating the probability of exceeding elevated pH values critical to fish populations in a hypereutrophic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2262–2270, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-158.

Black:1999:EDS

[2887] Kerry P. Black and Gregory D. Parry. Entrainment, dispersal, and settlement of scallop dredge sediment plumes: field measurements and numerical modelling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2271– 2281, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-159.

Gutreuter:1999:EFP

[2888] Steve Gutreuter, Andrew D. Bartels, Kevin Irons, and Mark B. Sandheinrich. Evaluation of the flood-pulse concept based on statistical models of growth of selected fishes of the Upper Mississippi River system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2282–2291, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-161.

Hughes:1999:PPR

[2889] Nicholas F. Hughes. Population processes responsible for larger-fish-upstream distribution patterns of Arctic grayling (Thymallus arcticus) in interior Alaskan runoff rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2292–2299, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-157.

Scheibling:1999:DGE

[2890] Robert E. Scheibling, Allan W. Hennigar, and Toby Balch. Destructive grazing, epiphytism, and disease: the dynamics of sea urchin–kelp interactions in Nova Scotia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12): 2300–2314, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-163.

Rakitin:1999:SCF

[2891] Ana Rakitin, Moira M. Ferguson, and Edward A. Trippel. Sperm competition and fertilization success in Atlantic cod (Gadus morhua): effect of sire size and condition factor on gamete quality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2315–2323, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-164.

Waddy:1999:TMM

[2892] S. L. Waddy and D. E. Aiken. Timing of the metamorphic molt of the American lobster (Homarus americanus) is governed by a populationbased, photoperiodically entrained daily rhythm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2324–2330, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-165.

Hatch:1999:RIS

[2893] Lorin K. Hatch, John E. Reuter, and Charles R. Goldman. Relative importance of stream-borne particulate and dissolved phosphorus fractions to Lake Tahoe phytoplankton. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2331–2339, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-166.

Docker:1999:PLG

[2894] Margaret F. Docker, John H. Youson, Richard J. Beamish, and Robert H. Devlin. Phylogeny of the lamprey genus Lampetra inferred from mitochondrial cytochrome b and ND3 gene sequences. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2340–2349, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-171.

Shackell:1999:DEL

[2895] Nancy L. Shackell, Kenneth T. Frank, Brian Petrie, David Brickman, and Jennifer Shore. Dispersal of early life stage haddock (Melanogrammus aeglefinus) as inferred from the spatial distribution and variability in length-at-age of juveniles. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2350–2361, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/f99-172.

Giannico:1999:IFD

[2896] Guillermo R. Giannico and Michael C. Healey. Ideal free distribution theory as a tool to examine juvenile Coho salmon (Oncorhynchus kisutch) habitat choice under different conditions of food abundance and cover. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2362–2373, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic).

URL https://cdnsciencepub.com/doi/full/10.1139/f99-173.

Ryding:1999:MRR

[2897] Kristen E. Ryding and John R. Skalski. Multivariate regression relationships between ocean conditions and early marine survival of Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12): 2374–2384, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-177.

Chen:1999:NNM

[2898] D. G. Chen and D. M. Ware. A neural network model for forecasting fish stock recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12): 2385–2396, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-178.

Valdimarsson:1999:ETD

[2899] Sveinn K. Valdimarsson and Neil B. Metcalfe. Effect of time of day, time of year, and life history strategy on time budgeting in juvenile Atlantic salmon, Salmo salar. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12): 2397–2403, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-179.

Myers:1999:MRR

[2900] Ransom A. Myers, Keith G. Bowen, and Nicholas J. Barrowman. Maximum reproductive rate of fish at low population sizes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2404–2419, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-201.

Zakardjian:1999:BMI

[2901] Bruno A. Zakardjian, Jeffrey A. Runge, Stephane Plourde, and Yves Gratton. A biophysical model of the interaction between vertical migration of crustacean zooplankton and circulation in the Lower St. Lawrence Estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2420–2432, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-095.

Robinson:1999:SOR

[2902] Clifford K. L. Robinson and Daniel M. Ware. Simulated and observed response of the southwest Vancouver Island pelagic ecosystem to oceanic conditions in the 1990s. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12): 2433–2443, December 1999. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/f99-170.

Reiss:1999:USL

[2903] C. S. Reiss, I. A. McLaren, and P. A. Avendaño. Utility of storage lipid volumes in inferring recent trophic history of copepods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2444–2449, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-196.

Wu:1999:MSC

[2904] Julia Qiuying Wu and William W. Hsieh. A modelling study of the 1976 climate regime shift in the North Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2450–2462, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-195.

Head:1999:DCS

[2905] Erica J. H. Head, Leslie R. Harris, and Brian Petrie. Distribution of Calanus spp. on and around the Nova Scotia shelf in April: evidence for an offshore source of Calanus finmarchicus to the central and western regions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2463–2476, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
f99-193.

Perry:1999:SCI

[2906] R. Ian Perry, Peter A. Thompson, David L. Mackas, Paul J. Harrison, and Douglas R. Yelland. Stable carbon isotopes as pelagic food web tracers in adjacent shelf and slope regions off British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2477–2486, December 1999. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/f99-194.

Riessen:1999:PIL

[2907] Howard P. Riessen. Predator-induced life history shifts in Daphnia: a synthesis of studies using meta-analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2487–2494, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-155.

Trippel:1999:FMB

[2908] Edward A. Trippel. The first marine biological station in Canada: 100 years of scientific research at St. Andrews. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):2495–2507, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f99-190.

Anonymous:1999:AIV

[2909] Anonymous. Author index: Volume 56, 1999. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):??, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfasauthor99.

Anonymous:1999:CV

[2910] Anonymous. Contents: Volume 56, 1999. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 56(12):??, December 1999. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfascontents99.

Zabel:2000:CCP

[2911] Richard W. Zabel and John G. Williams. Comments on “Contrasting patterns of productivity and survival rates for stream-type Chinook salmon (Oncorhynchus tshawytscha) populations of the Snake and Columbia rivers” by Schaller et al. (1999). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(8):1739–1741, August 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-099. See [2758].

Buckley:2000:CLA

[2912] Lawrence J. Buckley, R. Gregory Lough, Myron A. Peck, and Francisco E. Werner. Comment: Larval Atlantic cod and haddock growth models, metabolism, ingestion, and temperature effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1957–1960, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-146. See [2662].

Leising:2000:RBA

[2913] Andrew W. Leising and Peter J. S. Franks. Reply to Buckley et al.’s “Comment: Larval Atlantic cod and haddock growth models, metabolism, ingestion, and temperature effects”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 57(9):1961–1963, September 2000. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/f00-145. See [2662, 2912].

McClatchie:2010:RAS

[2914] Sam McClatchie, Ralf Goericke, Guillermo Auad, and Kevin Hill. Reassessment of the stock–recruit and temperature–recruit relationships for Pacific sardine (Sardinops sagax). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 67(11):1782–1790, November 2010. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub. com/doi/full/10.1139/F10-101. See [1537] and comment [2916, 2915].

Lindegren:2013:TDP

[2915] Martin Lindegren and David M. Checkley. Temperature dependence of Pacific sardine (Sardinops sagax) recruitment in the California Current ecosystem revisited and revised. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(2):245–252, February 2013. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.

com/doi/full/10.1139/cjfas-2012-0211. See [1537, 2916].

Jacobson:2013:CTD

[2916] Larry D. Jacobson and Sam McClatchie. Comment on temperaturedependent stock–recruit modeling for Pacific sardine (Sardinops sagax) in Jacobson and MacCall (1995), McClatchie et al. (2010), and lindegren and checkley (2013). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 70(10): 1566–1569, October 2013. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2013-0128. See [2914].